### v1

# On an entropy inequality for quadratic forms and applications

Abstract
Let $\phi(x,y)$ be a non-degenerate rational quadratic form. Let $X$ and $Y$ be independent $(s,C)$-Frostman random variables whose ranges are contained in $[-c_{1},c_{1}]$, with $10<s<$ 1, $C,c_{1}\geq 1$. We prove that there exist a positive constant $\epsilon=\epsilon(s,\phi)$ and an integer $N=N(s,C,c_{1},\phi)$ such that
$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon)$
for all $n>N$. The proof introduces a novel multi‑step entropy framework, combining the submodularity formula, the discretized entropy Balog-Szemerédi-Gowers theorem, and state-of-the-art results on the Falconer distance problem, to reduce general forms to a diagonal core case. As an application, we derive a result on a discretized sum-product type problem. In particular, for a $\delta$-separated set $A\subset[0,1]$ of cardinality $\delta^{-s}$, satisfying some non-concentration conditions, there exists $\epsilon=\epsilon(s,\phi)>0$ such that
$\mathcal{E}_{\delta}(A+A)+\mathcal{E}_{\delta}(\phi(A,A))\gg\delta^{-\epsilon}(\#A)$
for all $\delta$ small enough. Here by $\mathcal{E}_{\delta}(A)$ we mean the $\delta$-covering number of $A$.
1 Introduction
We say a random variable $X$ is $(s,C)$-Frostman if its distribution measure is $(s,C)$-Frostman, which means that for any $x$ and $r>0$ we have
$P(X\in B(x,r))\leq Cr^{s}.$
For each $n\in\mathbb{N}$, let $D_{n}\colon\mathbb{R}\to\mathbb{R}$ be the function defined by $D_{n}(x)=\frac{k}{2^{n}}$ if $k$ is the unique integer such that $\frac{k}{2^{n}}\leq x<\frac{k+1}{2^{n}}$. For any random variable $X$, $D_{n}(X)$ defines a discrete random variable that takes values in the set of dyadic rationals. We write $H_{n}(X):=H(D_{n}(X))$, where the entropy of a discrete random variable is defined by
$H(D_{n}(X))=-\sum_{x}P(D_{n}(X)=x)\log P(D_{n}(X)=x).$
In this paper, we establish a general entropy inequality that both sharpens our theoretical understanding and points toward a range of applications. Our study is motivated by recent applications of entropy methods to discretized sum-product estimates [11] due to Máthé and O’Regan, and to the recent solution of the polynomial Freiman-Ruzsa conjecture [1, 2] due to Gowers, Green, Manners, and Tao.
More precisely, we are interested in bounding the following quantity $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ from below, where $\phi(x,y)=a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ is a non-degenerate rational quadratic form, i.e. $a_{2}^{2}-4a_{1}a_{3}\neq 0$, $X$ and $Y$ are independent $(s,C)$-Frostman random variables whose ranges are bounded, and $s\in(0,1)$. We state our main result as follows.
Theorem 1.1.
Let $\phi(x,y)$ be a non-degenerate rational quadratic form. Let $X$ and $Y$ be independent $(s,C)$-Frostman random variables whose ranges are contained in $[-c_{1},c_{1}]$, with $10<s<$ 1, and $C,c_{1}\geq 1$. Then there exist a positive constant $\epsilon=\epsilon(s,\phi)$ and an integer $N=N(s,C,c_{1},\phi)$ such that for all $n>N,$ we have
$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon).$ (1)
To show the sharpness of this theorem, we summarize the main statements of several examples in Section A:
  1. 1.
    Any $(s,C)$-Frostman random variable $Z$ has the property that $H_{n}(Z)\geq ns-\log C+O(1)$ (see Example A.1 for an example matching this bound). The lower bound in (1) says that the discretized entropies of $X+Y$ and $\phi(X,Y)$ cannot be small simultaneously. We show in Examples A.1 and A.2 that at least one of the terms $H_{n}(X+Y)$ or $H_{n}(\phi(X,Y))$ can be as small as $(1+o(1))ns$.
  2. 2.
    The positive $\epsilon$ might not exist if one of the conditions of the theorem is not satisfied, including when $s=0$ or $1$ (see Example A.1), the Frostman constant $C$ grows with $n$ (see Example A.4), or the polynomial is degenerate (see Example A.3).
  3. 3.
    In the statement of Theorem 1.1, the fact that $N$ depends on $C$ , and the condition that $X$ and $Y$ are independent are both essential (see Examples A.4 and A.5, respectively).
  4. 4.
    For any $10<s<$ 1 and any sufficiently large $n\in\mathbb{N}$, there exists an $(s,O(1))$-Frostman random variable $X$ with range contained in $[0,1]$, such that
    $\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}=(1+o(1))H_{n}(X)$
    as $n\rightarrow\infty$, where $Y$ is i.i.d. with $X$. In other words, the term $ns$ can not be replaced by $H_{n}(X)$ in the right-hand side of (1).
  5. 5.
    The value of $\epsilon$ is not explicit, and an inspection of the proofs suggests it would be small.
We pause here to add several further remarks. Our proof of Theorem 1.1 presents a connection between $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ and the entropy of the “distance” function, i.e. $H_{n}\big{(}(X_{1}-Y_{1})^{2}+(X_{2}-Y_{2})^{2}\big{)}$, where $(X_{1},Y_{1})$, $(X_{2},Y_{2})$ are i.i.d. copies of $(X,Y)$. When $a_{1}=a_{3}=0$ and $a_{2}=1$, i.e., $\phi(x,y)=xy$, Máthé and O’Regan [11] established a different type of connection by replacing the entropy of the “distance” function with that of the “slope” function. By combining this with a generalization of Marstrand’s theorem due to Orponen, Shmerkin, and Wang in [19], they were able to prove Theorem 1.1 for the full range of $s$ for this specific bilinear form function. We do not see how their method could be adapted to the form of $\phi$ with $a_{1},a_{3}\neq 0$. Moreover, in terms of applications, i.e. sum-product type problems over $\mathbb{R}$ or $\mathbb{F}_{p}$, the two functions $xy$ and $a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ appear to exhibit distinct behaviors. It would be interesting to explore more connections between $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ and the entropy of other functions.
The next theorem presents an application of Theorem 1.1. Here, by $\underline{\dim}_{B}$ and $H^{s}$, we mean the lower box dimension and the $s$-Hausdorff measure, respectively.
Theorem 1.2.
Let $\phi(x,y)$ be a non-degenerate rational quadratic form. For $10<s<$ 1, there exists a positive constant $\epsilon=\epsilon(s{\color[rgb]{0,0,0},\phi})$ such that for every Borel set $A\subset{\color[rgb]{0,0,0}[0,1]}$ with $H^{s}(A)>0$, we have
$\max\left\{\underline{\dim}_{B}(A+A),\underline{\dim}_{B}(\phi(A,A))\right\}\geq s+\epsilon.$
This theorem is of independent interest, particularly when $\phi(x,y)$ is a non-bilinear form, and might be considered as an application of the Falconer distance problem. If $\phi(x,y)=xy$, then stronger results have been obtained by using a variety of techniques, see [4, 10, 11, 16] and references therein for current progress. Given this line of research, it is reasonable to expect different proofs of Theorem 1.2 with other methods. In the prime field setting, such a result has been proved by using tools from incidence geometry, see [8, Theorem 1.5] and [14, Theorem 1.3], for example.
For $A\subset\mathbb{R}$ and $\delta>0$, by $\mathcal{E}_{\delta}(A)$ we mean the $\delta$-covering number of $A$. The above theorem can be rewritten in the language of covering numbers.
Theorem 1.3.
Let $\phi(x,y)$ be a non-degenerate rational quadratic form. Given $10<s<$ 1 and a constant $C\geq 1$. Then there exist $\epsilon=\epsilon(s{\color[rgb]{0,0,0},\phi})>0$ and $\delta_{0}=\delta_{0}(s,C,\phi)>0$ such that the following holds for all $\delta\leq\delta_{0}$. Let $A\subset{\color[rgb]{0,0,0}[0,1]}$ be a $\delta$-separated set of cardinality $\delta^{-s}$. Suppose that for all intervals $I$of length at least $\delta$, we have the non-concentration estimate
#$\#(A\cap I)\leq C|I|^{s}(\#A).$
Then we have
$\mathcal{E}_{\delta}(A+A)+\mathcal{E}_{\delta}(\phi(A,A))\gg C^{-O(1)}\delta^{-\epsilon}(\#A).$
We now sketch the main ideas in the proof of Theorem 1.1.
Our base case is corresponding to the diagonal forms $\phi(X,Y)=\alpha X^{2}+\beta Y^{2}$ (Lemma 4.1), with independent $(s,C)$-Frostman random variables $X,Y$. Roughly speaking, by applying the submodularity formula (Lemma 2.18) on certain algebraic relations between the sum function $X+Y$ and the distance function $(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}$, we are able to reduce the problem to bounding the entropy of the distance function from below. In particular, let $W$ be an $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $10<s<$ 1, the constant $C\geq 1$ is bounded and does not depend on $n$, and $R>0$. Let $W^{\prime}$ be a random variable which is i.i.d with $W$. Then there exists a constant $\epsilon_{0}=\epsilon_{0}(s)>0$ such that
$H_{n}(|W-W^{\prime}|^{2})\geq n\left(s+\epsilon_{0}\right)$ (2)
for all $n\geq N_{0}$, where $N_{0}=N_{0}(s,C,R)$. Note that if the Frostman constant $C$ depends on $n$, then the inequality (2) might not be true (see Example A.5). To prove (2), we make use of results on the Falconer distance problem. While classical estimates taken from [12, Section 4.2] are sufficient in the case $s>1/2$, we invoke the recent work of Shmerkin and Wang [20] for $s\leq 1/2$.
For the general form $\phi(X,Y)$, the proof becomes much more complicated. The first step is to transform it to a diagonal form $\tilde{\phi}(\tilde{X},\tilde{Y})$, but the transformed variables $\tilde{X}$ and $\tilde{Y}$ are no longer independent. To proceed further, we introduce a novel approach, which can be briefly described as follows. By applying the discretized entropy version of Balog-Szemerédi-Gowers theorem (see Theorem 3.1), we obtain new random variables $X^{\prime\prime}$ and $Y^{\prime\prime}$ that are conditionally independent, and thus reduce the problem accordingly. Generally speaking, there exists a random variable $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$ such that
(i) $(X^{\prime},Y^{\prime})\equiv(\tilde{X},\tilde{Y})$;
(ii) $X^{\prime}$ and $X^{\prime\prime}$ are conditionally i.i.d. given $D_{n}(Y^{\prime})$;
(iii) $Y^{\prime}$ and $Y^{\prime\prime}$ are conditionally i.i.d. given $D_{n}(X^{\prime})$;
(iv) $(X^{\prime\prime},Y^{\prime})$ and $Y^{\prime\prime}$ are conditionally independent given $D_{n}(X^{\prime})$.
Then, given that $D_{n}(X^{\prime},Y^{\prime})=(x,y)$ with $x,y$ being dyadic rationals satisfying $P\big{(}D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}>0$, in the proof we show that
$\begin{split}\big{(}X^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}\text{ is $\left(s,O\left(C^{2}\cdot 2^{-ns}/P_{E}(D_{n}(Y^{\prime})=y)\right)\right)$-Frostman},\\ \big{(}Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}\text{ is $\left(s,O\left(C^{2}\cdot 2^{-ns}/P_{E}(D_{n}(X^{\prime})=x)\right)\right)$-Frostman}.\end{split}$ (3)
The problem now is reduced to the case of random variables $\big{(}X^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}$ and $\big{(}Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}$, and the corresponding entropy
$\max\left\{H_{n}\left(X^{\prime\prime}+Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\right),\,H_{n}\left(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\right)\right\}.$
It is clear that we have many pairs $(x,y)$, and for each pair, the corresponding Frostman constants in (3) are not uniformly bounded. Besides technical issues including Lipschitz conditions and partitions of measure supports, the main difficulty arises because the Frostman constants may grow with $n$. Note that the idea of scaling the Frostman constants to $1$ does not help, since it would change the supports of variables from bounded to unbounded.
We address this challenge by explicitly computing the entropy of the distance function. Instead of (2), we have the following for both small and large dimensions:
Large dimension: Let $W$ be an $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $11/2<s<$ 1, $C\geq 1$ and $R>0$. Let $W^{\prime}$ be a random variable which is i.i.d with $W$. Then for any number $t$ satisfying $t\leq 1$ and $s<t<2s-1/2$, we have that for all $n\in\mathbb{N},$
$H_{n}(|W-W^{\prime}|)\geq nt-\log C+O(|\log R|)+O(|\log(2s-t-1/2)|)+O_{t}(1).$ (4)
Small dimension: Let $X^{\prime\prime}$, $Y^{\prime\prime}$ be independent random variables with ranges contained in $[-c,c]$, with $c>0$. Suppose that $X^{\prime\prime}$ and $Y^{\prime\prime}$ are $(s,C_{1})$-Frostman and $(s,C_{2})$-Frostman, respectively, where $20<s\leq 1/$ 2 and $C_{1},C_{2}\geq 1$. Let $(X^{\prime\prime\prime},Y^{\prime\prime\prime})$ be i.i.d. with $(X^{\prime\prime},Y^{\prime\prime})$. Then for any $t$ with $s<t<\frac{s+\varphi(2s)}{2}$, for all $n\in\mathbb{N}$, we have
$H_{n}\big{(}|(X^{\prime\prime},Y^{\prime\prime})-(X^{\prime\prime\prime},Y^{\prime\prime\prime})|\big{)}\geq nt+O_{s}(\log n)+O_{s,t}(\log C_{1}C_{2})+O_{s,c}(1),$ (5)
where
$\varphi(u)=\frac{u}{2}+\frac{\sqrt{4+u^{2}}-2}{2},\quad(0<u\leq 1).$ (6)
By (4) and (5), the error terms involving the Frostman constants in (3) are controlled up to logarithmic level. Then they can be further converted to certain entropy-terms that are feasible to handle. The proof of (4) is essentially the same as that of (2). To prove (5), we adapt the Shmerkin–Wang approach with appropriate modifications, while clarifying, simplifying, and quantifying key steps. Note that our inspection is not verbatim: for instance, in the thin‐tube step we must work with sets having a Cartesian‐product structure. This is why the estimate (5) holds for $W=X^{\prime\prime}\times Y^{\prime\prime}$ rather than in the general form of (4).
Plugging these inequalities in our method, the theorem follows.
Our final comment is that entropy methods have also been used in a range of problems in Additive Combinatorics and Geometric Measure Theory, for example, see [3, 6, 9, 15, 17, 20, 21, 22] and references therein.
2 Preliminaries
2.1 Notations
Let $(\Omega,\mathfrak{B},P)$ be a probability space. A random variable $X$ is a measurable function from $(\Omega,\mathfrak{B},P)$ to another measuable space $(\Omega^{\prime},\mathfrak{B}^{\prime})$, on which a pushforward probability measure $\mu$ is induced, i.e., $\mu(B^{\prime})=P(X\in B^{\prime})$ for any $B^{\prime}\in\mathfrak{B}^{\prime}$. We will say that $\mu$ is the associated distribution measure of $X$. We denote the range of $X$ by $range(X)$, which collects all the possible values that $X$ can take. In this paper, all the measurable spaces will be taken to be subspaces of ${\mathbb{R}}^{m}$, equipped with the Borel $\sigma$-algebra or its sub-$\sigma$-algebra, and the Euclidean norm ||$\cdot$|| ⋅ |. The dimension $m$ is always regarded as an absolute constant. The range of any random variable is assumed to be a bounded set. When two random variables $X$ and $Y$ have the same distribution, we write $X\equiv Y$.
More generally, let $\mu$ be a Borel measure. Its support, the smallest closed set whose complement has zero measure, is denoted by $spt(\mu)$. When $\mu$ is the distribution measure associated to a random variable $X$, we also write $range_{+}(X)=spt(\mu)$. That is to say, $X$ takes value in $range_{+}(X)$ with probability $1$.
For a measure $\mu$ and a Borel set $E$ with $\mu(E)>0$, we define the conditional measure of $\mu$ with respect to $E$ by $\mu_{E}(B)=\frac{\mu(B\cap E)}{\mu(E)}$ for any Borel set $B$. In particular, when $\mu=P$, the conditional probability measure given an event $E$ is denoted by $P_{E}$. For a random varaible $X$ and such an $E$, we denote by $X_{E}$ the random variable satisfying $X_{E}\equiv(X|E)$ (see Lemma B.2). That is to say,
$P_{E}(X_{E}\in B)=\frac{P(\{X\in B\}\cap E)}{P(E)}$
for any $B\in\mathfrak{B}$.
For a measurable function $f:\,(\Omega^{\prime},\mathcal{B}^{\prime})\rightarrow(\Omega^{\prime\prime},\mathfrak{B}^{\prime\prime})$ between measurable spaces, the random variable $f(X)$ is defined to be the composition of $f$ and $X$. The domain and range of $f$ are denoted by $domain(f)$ and $range(f)$, respectively. Throughout this paper, all the functions are assumed to be measurable. And whenever we write $f(X)$ with $X$ a random variable and $f$ a function, we always mean that $domain(f)\supseteq range_{+}(X)$, where the values taken on a zero-measure set is neglectable.
In this paper, by $f=O_{\kappa}(g)$, $f\ll_{\kappa}g$ or $g\gg_{\kappa}f$, we mean that there exists an absolute positive constant $C=C(\kappa)$ depending on the parameter(s) $\kappa$ such that $|f|\leq Cg$. All logarithms we use here have base $2$.
2.2 Entropy and basic properties
Our primary tools will involve entropy. We begin by recalling some basic definitions and elementary consequences thereof. More discussions can be found in [6, 17, 7, 11, 22].
Definition 2.1 (Entropy of partitions).
Let $(\Omega,\mathcal{B},\mu)$ be a measure space. For a countable partition $\mathcal{A}$ of $\Omega$, the entropy of $\mathcal{A}$, denoted $H(\mu,\mathcal{A}),$ is defined to be
$H(\mu,\mathcal{A})=-\sum_{A\in\mathcal{A}}\mu(A)\log\mu(A).$
For two countable partitions $\mathcal{A}$ and $\mathcal{B}$, the conditional entropy of $\mathcal{A}$ given $\mathcal{B},$ denoted $H(\mu,\mathcal{A}|\mathcal{B}),$ is defined as
$H(\mu,\mathcal{A}|\mathcal{B})=\sum_{B\in\mathcal{B}}\mu(B)H(\mu_{B},\mathcal{A}).$
In above definition, the convention $00\log 0=$ 0 is used for convenience. We also say $\mu(B)H(\mu_{B},\mathcal{A})=0$ when $\mu(B)=0$, although $\mu_{B}$ is not defined in this situation. Intuitively, the entropy of a random variable will tell us how hard it is to determine a value. Before stating the first lemma, we introduce the following notation:
$Z(\mathcal{A})=\{A\in\mathcal{A}\mid\mu(A)>0\},\quad\mathcal{A}\land\mathcal{B}=\{A\cap B\mid A\ \in\mathcal{A},B\in\mathcal{B}\}.$
Moreover, the function $\iota(x)=-x\log x$ is non-negative valued, bounded and concave on $[0,1]$. It is increasing on $[0,e^{-1}]$, and decreasing on $[e^{-1},1]$. With this in tow, we record some properties of entropy for partitions, whose proofs are simple exercises.
Lemma 2.2.
For a probability measure $\mu$ and countable partitions $\mathcal{A}$, $\mathcal{B}$ and $\mathcal{C}$, we have the following statements:
(i) $H(\mu,\mathcal{A})\geq 0$ with equality holding if and only if $|Z(\mathcal{A})|=1.$
(ii) $H(\mu,\mathcal{A})\leq\log|Z(\mathcal{A})|$ when $Z(\mathcal{A})$ is a finite set, with equality holding if and only if $\mu(A)=1/|Z(\mathcal{A})|$ for all $A\in Z(\mathcal{A}).$
(iii) $H(\mu,\mathcal{A})+H(\mu,\mathcal{B}|\mathcal{A})=H(\mu,A\land B).$
(iv) $H(\mu,\mathcal{A}|\mathcal{B})\leq H(\mu,\mathcal{A})$, or equivalently $H(\mu,\mathcal{A}\land\mathcal{B})\leq H(\mu,\mathcal{A})+H(\mu,\mathcal{B})$, with equality holding if and only if $\mu(A\cap B)=\mu(A)\mu(B)$ for all $A\in\mathcal{A}$ and $B\in\mathcal{B}$.
(v) $H(\mu,\mathcal{A}\land\mathcal{B}|\mathcal{C})\leq H(\mu,\mathcal{A}|\mathcal{C})+H(\mu,\mathcal{B}|\mathcal{C})$.
(vi) For any Borel set $Y$, we have $\mu(Y)H(\mu_{Y},\mathcal{A})\leq H(\mu,\mathcal{A}).$
For discrete random variables $X$ and $Y$ defined on $(\Omega,\mathfrak{B},P)$, the families
$\mathcal{A}=\big{\{}\{X=x\}:\,x\in range(X)\big{\}},\quad\text{and}\quad\mathcal{B}=\big{\{}\{Y=y\}:\,y\in range(Y)\big{\}}$
give two partitions of $\Omega$. We have the following definitions.
Definition 2.3 (Entropy of random variables).
Let $X$ and $Y$ be discrete random variables, and $E$ be an event. The entropy of $X$ is defined as
$H(X)=-\sum_{x}P(X=x)\log P(X=x).$
The conditional entropy of $X$ given $E$ is defined as
$H(X|E)=H(X_{E})=-\sum\limits_{x}P_{E}(X=x)\log P_{E}(X=x).$
And the conditional entropy of $X$ given $Y$ is defined as
$H(X|Y)=\sum_{y}P(Y=y)H(X|Y=y).$
For the random variable $Z=(X,Y)$, we denote $H(Z)=H(X,Y)$, instead of $H\big{(}(X,Y)\big{)}$. We now state some basic properties of entropy for discrete random variables. These properties are well established in the literature, so we omit their proofs.
Lemma 2.4.
Let $X,Y$ be discrete random variables, and $f$ be a function. The following statements hold.
(i) $H(X)\geq 0$ with equality holding if and only if $|range_{+}(X)|=1$.
(ii) $H(X)\leq\log|range_{+}(X)|$ when $range_{+}(X)$ is a finite set, with equality holding if and only if $X$ is uniformly distributed on $range_{+}(X)$.
(iii) $H(X)+H(Y|X)=H(X,Y)$, and in particular, $H(X)\leq H(X,Y)$.
(iv) $H(X|Y)\leq H(X)$, or equivalently $H(X,Y)\leq H(X)+H(Y)$, with equality holding when $X$ and $Y$ are independent.
(v) $H(f(X)|X)=0$, and furthermore $H(f(X))\leq H(X)$.
As a direct consequence, for discrete random variables $X$ and $Y$,
$H(X)\leq H(X,Y)=H(Y)+H(X|Y).$ (7)
Moreover, Lemma 2.4(iv) can be generalized to conditional entropy: $H(X_{1},X_{2}|Y)\leq H(X_{1}|Y)+H(X_{2}|Y)$, with equality holding if $X_{1},X_{2}$ are conditionally independent given $Y$.
The following lemma shows a concave property of the entropy of a random variable.
Lemma 2.5.
Let $X$ be a random variable and $\mathcal{B}$ be a finite partition of $\Omega$. Then
$H(X)\geq\sum\limits_{B\in\mathcal{B}}P(B)H(X_{B}).$
Proof.
Take $\mathcal{A}=\big{\{}\{X=x\}:\,x\in range(X)\big{\}}$. Then Lemma 2.2(iv) shows that
$H(X)=H(P,\mathcal{A})\geq H(P,\mathcal{A}|\mathcal{B})=\sum\limits_{B\in\mathcal{B}}P(B)H(P_{B},\mathcal{A})=\sum\limits_{B\in\mathcal{B}}P(B)H(X_{B}).$
In the following, we introduce a statement of the submodular inequality, taken from [11].
Lemma 2.6 (Submodularity).
Let $X,Y,Z,W$ be discrete random variables. If $X$ is a function of $Z$, $X$ is also a function of $W$, and furthermore $Y$ is a function of $(Z,W)$, then
$H(X)+H(Y)\leq H(Z)+H(W).$
We often use the following corollary of submodularity.
Corollary 2.7.
Let $X,Y,Z,W$ be discrete random variables. We have
$H(X)+H(Y)\leq H(Z)+H(W)+H(X|Z)+H(X|W)+H(Y|Z,W).$
Proof.
Using Lemma 2.6, we have
$H(X,Z)+H(X,W)\geq H(X,Z,W)+H(X).$
So, by applying Lemma 2.4 (iii) and (7), we obtain
$H(Z)+H(W)+H(X|Z)+H(X|W)+H(Y|Z,W)$
$\quad=H(X,Z)+H(X,W)+H(Y|Z,W)\geq H(Y|Z,W)+H(X,Z,W)+H(X)$
$\quad\geq H(Y|Z,W)+H(Z,W)+H(X)\geq H(Y)+H(X).$
Hence, the corollary follows. ∎
The next lemma is taken from [22], namely, the estimate (22).
Lemma 2.8.
Let $(X,Y)$ be a discrete random variable, and $X^{\prime},X^{\prime\prime}$ be conditionally i.i.d. copies of $X$ given $Y$. Then
$H(X^{\prime},X^{\prime\prime},Y)=2H(X,Y)-H(Y).$
The following is known as the Plünnecke-Ruzsa inequality. A proof can be found several places, for example, [7, Theorem 3.11].
Lemma 2.9 (Plünnecke-Ruzsa inequality).
Let $X,Y_{1},Y_{2},\ldots,Y_{m}$ be independent discrete random variables. Then
$H(X+Y_{1}+\ldots+Y_{m})-H(X)\leq\sum\limits_{i=1}^{m}\big{(}H(X+Y_{i})-H(X)\big{)},$
$H(Y_{1}+\ldots+Y_{m})-H(X)\leq\sum\limits_{i=1}^{m}\big{(}H(X+Y_{i})-H(X)\big{)}.$
The next lemma is helpful in getting an upper bound for certain entropy-terms.
Lemma 2.10.
Let $k\geq 1$ and $20<\lambda\leq 1/$ 2. Suppose that $p_{1},p_{2},\ldots,p_{k}$ are non-negative numbers satisfying $\sum\nolimits_{i=1}^{k}p_{i}\leq\lambda$. Then
$-\sum\limits_{i=1}^{k}p_{i}\log p_{i}\leq\lambda\log k+O(1).$
Proof.
Define $p_{k+1}=1-\sum\nolimits_{i=1}^{k}p_{i}$, which satisfies $p_{k+1}\geq 1-\lambda\geq 1/2$. Then we assume without loss of generality that
$p_{1}\leq p_{2}\leq\ldots\leq p_{k}\leq p_{k+1}.$
Let us consider the the maximum value of the function
$f(p_{1},\ldots,p_{k+1})=-\sum\limits_{i=1}^{k+1}p_{i}\log p_{i},$
which is attainable in the feasible region of $(p_{1},\ldots,p_{k+1})$. For any $11\leq j<i\leq k+$ 1, denote
$\tilde{f}(p_{1},\ldots,p_{j-1},p_{j+1},\ldots,p_{k+1})=f(p_{1},\ldots,p_{j-1},1-\sum\nolimits_{l\neq j}p_{l},p_{j+1},\ldots,p_{k+1}).$
In view of the fact that $p_{i}\geq p_{j}$, one deduces that
$\frac{\partial\tilde{f}}{\partial p_{i}}(p_{1},\ldots,p_{j-1},p_{j+1},\ldots,p_{k+1})=\log\frac{1-\sum\nolimits_{l\neq j}p_{l}}{p_{i}}=\log\frac{p_{j}}{p_{i}}\leq 0.$
That is to say, when the other $k-1$ variables are given, the small $p_{i}$ is, the larger value $\tilde{f}$ and $f$ take. Now one can verify that the maximum value of $f$ is taken at $(p_{1},\ldots,p_{k+1})$ with
$p_{i}=\frac{\lambda}{k}\,\,(1\leq i\leq k),\qquad p_{k+1}=1-\lambda.$
That is to say,
$-\sum\limits_{i=1}^{k+1}p_{i}\log p_{i}\leq k\cdot\left(-\frac{\lambda}{k}\log\frac{\lambda}{k}\right)-(1-\lambda)\log(1-\lambda)\leq\lambda\log k+1.$
It follows that
$-\sum\limits_{i=1}^{k}p_{i}\log p_{i}\leq\lambda\log k+1+p_{k+1}\log p_{k+1}=\lambda\log k+O(1).$
2.3 Discretized entropy and properties
For each $n\in\mathbb{N}$ and $m\in\mathbb{N}$, let us define a function $D_{n,m}:\,{\mathbb{R}}^{m}\rightarrow{\mathbb{R}}^{m}$ as the following. For any $(x_{1},\ldots,x_{m})\in{\mathbb{R}}^{m}$, if $k_{i}$ is the unique integer such that $\frac{k_{i}}{2^{n}}\leq x_{i}<\frac{k_{i}+1}{2^{n}}$ $(i=1,\ldots,m)$, then we define $D_{n,m}(x_{1},\ldots,x_{m})=(\frac{k_{1}}{2^{n}},\ldots,\frac{k_{m}}{2^{n}})$. That is to say, $D_{n,m}(x)$ locates the left endpoints of the dyadic intervals that the $x_{i}$’s fall in. For convenience, we will simplify $D_{n,m}$ as $D_{n}$, where the dimension $m$ is determined from the context. For any random variable $X$, the expression $D_{n}(X)$ is a discrete random variable that takes values in the set of points whose components are all dyadic rationals. Moreover, if $X_{1},\ldots,X_{k}$ are independent (or identically distributed) random variables, then $D_{n}(X_{1}),\ldots,D_{n}(X_{k})$ are also independent (or identically distributed, respectively).
For random variables that are not discrete, we can discretize them with the function $D_{n}$. For convenience, for any random variables $X$ and $Y$, we denote the discretized entropy by
$H_{n}(X):=H(D_{n}(X)),\qquad H_{n}(X|Y):=H(D_{n}(X)\,|\,D_{n}(Y)).$
For a probability measure $\mu$, we introduce similar notation. Let $\mathcal{D}_{n}$ be the collection of all $m$-dimensional cubes $\{x\in{\mathbb{R}}^{m}:\,D_{n}(x)=k2^{-n}\}$ with $k\in\mathbb{Z}^{m}$. And define $H_{n}(\mu)=H_{n}(\mu,\mathcal{D}_{n})$. When $\mu$ is the distribution measure associated to $X$, the above two defintions coincide.
Some properties still hold if we replace $H$ by $H_{n}$. For example,
$H_{n}(X|Y)=H(D_{n}(X)|D_{n}(Y))=H(D_{n}(X),D_{n}(Y))-H(D_{n}(Y))=H_{n}(X,Y)-H_{n}(Y).$
When $X$ and $Y$ are independent, then $D_{n}(X)$ and $D_{n}(Y)$ are also independent, which gives
$H_{n}(X,Y)=H(D_{n}(X),D_{n}(Y))=H(D_{n}(X))+H(D_{n}(Y))=H_{n}(X)+H_{n}(Y).$ (8)
Generally, one has
$H_{n}(X,Y)\leq H_{n}(X)+H_{n}(Y).$ (9)
Combining Lemma B.2, one can verify that the statement in Lemma 2.5 also holds for $H_{n}$: If $X$ is a random variable and $\mathcal{B}$ is a finite partition. Then
$H_{n}(X)\geq\sum\limits_{B\in\mathcal{B}}P(B)H_{n}(X_{B}).$ (10)
Moreover, the discretized entropy only depends on the discretised distribution. That is to say, if $D_{n}(X)$ and $D_{n}(Y)$ are identically distributed, then $H_{n}(X)=H_{n}(Y)$.
Furthermore, if $range_{+}(X)\subseteq B(0,R)\subseteq{\mathbb{R}}^{m}$ for some $R\geq 1$, then $range_{+}(D_{n}(X))$ has at most $O_{m}\big{(}R^{m}2^{mn}\big{)}$ elements. By Lemma 2.4(ii), it satisfies that
$H_{n}(X)\leq mn+m\log R+O_{m}(1).$ (11)
The submodular inequality will still hold if we replace the normal entropy with this special kind of entropy, as long as some additional conditions hold, as seen in the following definition.
Definition 2.11.
Let $\kappa$ be a positive real number, and $f$ be a function.
(i) We say that $f$ is $\kappa$-Lipschitz if $|f(x)-f(y)|\leq\kappa|x-y|$ for any $x,y\in domain(f)$.
(ii) We say that $f$ is $\kappa$-bi-Lipschitz if $f$ is invertible, and both $f$ and $f^{-1}$ are $\kappa$-Lipschitz.
(iii) If $f$ is $\kappa$-Lipschitz, and there exists a decomposition of $domain(f)=\mathcal{R}_{1}\cup\ldots\cup\mathcal{R}_{l}$, where $l$ is a positive integer viewed as an absolute constant, such that each $f|_{\mathcal{R}_{i}}$ has an $\kappa$-Lipschitz inverse $(i=1,\ldots,l)$, then we say that $f$ is weakly $\kappa$-bi-Lipschitz.
Example 2.12.
Let $D$ be an open convex region in ${\mathbb{R}}^{m}$. Let $f:\,D\rightarrow{\mathbb{R}}$ be a function such that $|\frac{\partial f}{\partial x_{i}}(x_{1},\ldots,x_{m})|\leq M$ for all $(x_{1},\ldots,x_{m})\in D$. Then, for any $(x_{1},\ldots,x_{m})$ and $(y_{1},\ldots,y_{m})$ in $D$, there is some $(\xi_{1},\ldots,\xi_{m})\in D$ such that
$|f(x_{1},\ldots,x_{m})-f(y_{1},\ldots,y_{m})|\leq\sum\limits_{i=1}^{m}\left|\frac{\partial f}{\partial x_{i}}(\xi_{1},\ldots,\xi_{m})\right||x_{i}-y_{i}|\leq Mm^{1/2}\left(\sum\limits_{i=1}^{m}|x_{i}-y_{i}|^{2}\right)^{1/2}.$
Therefore $f$ is $O(M)$-Lipschitz, where $m$ is viewed as an absolute constant.
Example 2.13.
Let $\alpha,c_{1},c_{2}$ be real numbers with $\alpha\neq 0$ and $0<c_{2}<c_{1}$. Let $f(x)=\alpha x^{2}$ with domain $[-c_{1},-c_{2}]\cup[c_{2},c_{1}]$. Note that, by the mean value theorem,
$2|\alpha|c_{2}\cdot|x-y|\leq|f(x)-f(y)|\leq 2|\alpha|c_{1}\cdot|x-y|$
[…]
Definition 2.14.
Let $X$, $Y$ be random variables and $\kappa$ be a postive real number. We say that $Y$ is a $\kappa$-Lipschitz function (or $\kappa$-bi-Lipschitz function, or weakly $\kappa$-bi-Lipschitz function) of $X$ if $Y=f(X)$ for some $\kappa$-Lipschitz function (or $\kappa$-bi-Lipschitz function, or weakly $\kappa$-bi-Lipschitz function, respectively) $f$ with $domain(f)\supseteq range_{+}(X)$.
If a function is $\kappa$-Lipschitz with $0<\kappa\leq 2$, then it is also $2$-Lipschitz. Throughout this paper, we always assume without loss of generality that $\kappa\gg 1$, and then $\log\kappa\gg 1$.
Lemma 2.15.
Let $X$ and $Y$ be random variables. We have the following two statements.
(i) If $f$ is $\kappa$-Lipschitz, then for all $n\in\mathbb{N}$,
$H_{n}(f(X)|X)\ll\log\kappa.$
(ii) If $g$ be weakly $\kappa$-bi-Lipschitz, then for all $n\in\mathbb{N}$,
$H_{n}(Y|g(Y)))\ll\log\kappa.$
The above implied constants are independent of $n$, and the concrete expressions of the functions and random variables.
Proof.
We show the proof of (ii) here; (i) follows by similar arguments.
Assume that $domain(g)\subseteq{\mathbb{R}}^{m}$ and $range(g)\subseteq{\mathbb{R}}^{m^{\prime}}$, where $m,m^{\prime}$ are viewed as absolute constants. When $D_{n}(g(Y))$ is determined, one knows that $g(Y)$ takes value in a ball with radius $O_{m^{\prime}}(2^{-n})$. By the definition of weakly $\kappa$-bi-Lipschitz functions, we can locate $Y$ to be lying in the pre-image of this ball, which is a union of at most $O(1)$ balls with radius $O_{m^{\prime}}(\kappa\cdot 2^{-n})$. So $D_{n}(Y)$ has at most $O_{m,m^{\prime}}(\kappa^{m})$ choices. It follows from Lemma 2.4(ii) that
$H_{n}(Y|g(Y))\leq\log(O_{m,m^{\prime}}(\kappa^{m}))\ll\log\kappa.$
The proof is completed. ∎
Lemma 2.16.
We have the following two statements.
(i) Let $X$ and $Y$ be random variables such that $X$ is an $\kappa$-Lipschitz function of $Y$. Then, for all $n\in\mathbb{N}$,
$H_{n}(X)\leq H_{n}(Y)+O(\log\kappa).$
(ii) Let $X$ and $Y$ be random variables such that $X$ is a weakly $\kappa$-bi-Lipschitz function of $Y$. Then, for all $n\in\mathbb{N}$,
$H_{n}(X)=H_{n}(Y)+O(\log\kappa).$
Proof.
Regarding (i), by (7), it is sufficient to prove that $H_{n}(X|Y)\ll\log\kappa$, which is guaranteed by Lemma 2.15(i). Regarding (ii), by (7) and Lemma 2.15(ii), we conclude that
$H_{n}(Y)\leq H_{n}(X)+H_{n}(Y|X)=H_{n}(X)+O(\log\kappa).$
The conclusion then follows. ∎
Example 2.17.
Let $X$ be a random variable with $range_{+}(X)\subseteq[-c_{1},-c_{2}]\cup[c_{2},c_{1}]$, where $0<c_{2}<c_{1}$. Let $\alpha\in{\mathbb{R}}^{\ast}$. One can deduce from Example 2.13 and Lemma 2.16 that
$H_{n}(\alpha X^{2})=H_{n}(X)+O\big{(}\max\{1,\log c_{1},\log(1/c_{2})\}\big{)}+O(\big{|}\log|\alpha|\big{|}).$
The following is the discretized version of submodularity.
Lemma 2.18.
If $X$ is a $\kappa$-Lipschitz function of $Z$, and also a $\kappa$-Lipschitz function of $W$, and furthermore, $Y$ is a $\kappa$-Lipschitz function of $(Z,W)$, then
$H_{n}(X)+H_{n}(Y)\leq H_{n}(Z)+H_{n}(W)+O(\log\kappa)$
for all $n\in\mathbb{N}$.
Proof.
According to Lemma 2.7, we only need to bound $H_{n}(X|Z)+H_{n}(X|W)+H_{n}(Y|Z,W)$ by $O(\log\kappa)$. This is a direct consequence of Lemma 2.15 . ∎
Lemma 2.19.
Let $X$, $Y$ be random variables, and $f$ be a $\kappa$-Lipschitz function with $range(f)\subseteq{\mathbb{R}}^{m}$, where $m$ is viewed as an absolute constant. Suppose that $|X-Y|\leq K$ with probability $1$. Then, for all $n\in\mathbb{N}$ with $2^{-n}\leq K$,
$H_{n}(f(X))\leq H_{n}(f(Y))+m(n+\log K)+O(\log\kappa).$
And, for all $n\in\mathbb{N}$ with $K=O(2^{-n})$,
$H_{n}(f(X))=H_{n}(f(Y))+O(\log\kappa).$
In particular, for all $n\in\mathbb{N}$,
$H_{n}(f(X))=H_{n}(f(D_{n}(X)))+O(\log\kappa).$
Proof.
Since $f$ is $\kappa$-Lipschitz, we have
$|f(X)-f(Y)|\leq\kappa|X-Y|\leq\kappa K$
with probability $1$. By (7), it follows that
$H_{n}(f(X))\leq H_{n}(f(Y))+H_{n}(f(X)|f(Y)).$
When $D_{n}(f(Y)))=t$ is determined, one has $|f(Y)-t|=O_{m}(2^{-n})$ and then $|f(X)-t|\leq\kappa K+O_{m}(2^{-n})$. So $D_{n}(f(X))$ has at most $O\big{(}(\kappa K\cdot 2^{n}+O_{m}(1))^{m}\big{)}$ choices. It follows that
[…]
Then the first conclusion follows. When $K=O(2^{-n})$, it satisfies that $H_{n}(f(X))\leq H_{n}(f(Y))+O(\log\kappa)$. And the second conclusion follows by symmetry. The third conclusion follows from the fact that $|X-D_{n}(X)|=O(2^{-n})$, i.e. $K=O(2^{-n})$. ∎
Lemma 2.20.
Let $l\geq 2$ and $X_{1},X_{2},\ldots,X_{l}$ be random variables. Then for any $n\in\mathbb{N}$,
$H_{n}(X_{1}+X_{2}+\ldots+X_{l})=H(D_{n}(X_{l})+D_{n}(X_{2})+\ldots+D_{n}(X_{l}))+O(\log l).$
Proof.
Note that $\sum\nolimits_{i=1}^{l}D_{n}(X_{i})$ takes values in the dyadic rationals. One has
$H\left(\sum\limits_{i=1}^{l}D_{n}(X_{i})\right)=H\left(D_{n}\Big{(}\sum\limits_{i=1}^{l}D_{n}(X_{i})\Big{)}\right)=H_{n}\left(\sum\limits_{i=1}^{l}D_{n}(X_{i})\right).$
The conclusion then follows from Lemma 2.19 with the $O(l)$-Lipschitz function (see Example 2.12) given by
$(x_{1},\ldots,x_{l})\mapsto x_{1}+\ldots+x_{l}.$
Now we show a discretized version of the Plünnecke-Ruzsa inequality.
Lemma 2.21.
Let $l\geq 1$ and $X,Y_{1},Y_{2},\ldots,Y_{l}$ be independent random variables. Then
$H_{n}(X+Y_{1}+Y_{2}+\ldots+Y_{l})-H_{n}(X)\leq\sum\limits_{i=1}^{l}\big{(}H_{n}(X+Y_{i})-H_{n}(X)\big{)}+O(\log l),$
and
$H_{n}(Y_{1}+Y_{2}+\ldots+Y_{l})-H_{n}(X)\leq\sum\limits_{i=1}^{l}\big{(}H_{n}(X+Y_{i})-H_{n}(X)\big{)}+O(\log l).$
Proof.
Note that $D_{n}(X),D_{n}(Y_{1}),\ldots D_{n}(Y_{l})$ are also independent. The conclusion follows from Lemmas 2.9 and 2.20. ∎
Corollary 2.22.
Let $X,Y,$ and $Z$ be independent random variables. Then, for all $n\in\mathbb{N}$,
$H_{n}(X)$ $\leq H_{n}(X+Y)+O(1),$
$H_{n}(X+Y+Z)$ $\leq H_{n}(X+Z)+H_{n}(Y+Z)-H_{n}(Z)+O(1),$
$H_{n}(X\pm Y)$ $\leq H_{n}(X+Z)+H_{n}(Y+Z)-H_{n}(Z)+O(1).$
Proof.
We only show the proof of the third inequality involving $H_{n}(X-Y)$ here. The others are direct consequences of Lemma 2.21.
By Lemma 2.18,
$H_{n}(X-Y)+H_{n}(X,Y,Z)\leq H_{n}(X+Z,Y+Z)+H_{n}(X,Y)+O(1).$
By the independence of $X,Y,$ and $Z,$ (specifically, by invoking (8) and (9)) it follows that
$H_{n}(X-Y)+H_{n}(X,Y)+H_{n}(Z)\leq H_{n}(X+Z)+H_{n}(Y+Z)+H_{n}(X,Y)+O(1).$
The conclusion follows. ∎
Lemma 2.23.
Let $X$ and $Y$ be independent real random variables and $k,l\in\mathbb{Q}$. There exist non-negative integers $d_{1}=d_{1}(k,l)$ and $d_{2}=d_{2}(k,l)$ such that
$H_{n}(kX+lY)\leq(d_{1}+d_{2}+1)H_{n}(X+Y)-d_{1}H_{n}(X)-d_{2}H_{n}(Y)+O_{k,l}(1).$
Here by $kX$ we mean the product of $k$ and the random variable $X$.
Proof.
The proof is trivial when $k=l=0$. When $l=0$ and $k\neq 0$, we have
$H_{n}(kX)=H_{n}(X)+O_{k}(1)\leq H_{n}(X+Y)+O_{k}(1)$
by Corollary 2.22. Similar conclusion holds when $k=0$. In the following, we assume that $k,l\neq 0$.
First, let us prove the case that $k,l\in\mathbb{Z}^{\ast}$. We apply double induction on $|k|$ and $|l|$. When $|k|=|l|=1$, we claim that
$H_{n}(kX+lY)\leq 3H_{n}(X+Y)-H_{n}(X)-H_{n}(Y)+O(1).$ (12)
Indeed, when $k=l=1$, one can deduce (12) from the fact that
$H_{n}(X+Y)\geq\max\{H_{n}(X),H_{n}(Y)\}+O(1).$
When $k=1$ and $l=-1$, we take $X,Y,Z$ to be independent random variables such that $X$ and $Z$ are identically distributed. Then Corollary 2.22 gives
$H_{n}(X-Y)\leq H_{n}(X+Z)+H_{n}(Y+Z)-H_{n}(Z)+O(1).$
By Corollary 2.22 again, we have
$H_{n}(X+Z)\leq H_{n}(X+Y)+H_{n}(Z+Y)-H_{n}(Y)+O(1).$ (13)
Then (12) follows. Since $H_{n}(W)=H_{n}(-W)$ for any random variable $W$, the cases that $k=-1$ and $l=\pm 1$ are also satisfied.
Now let us assume that the conclusion has been proved for some $k$ and $l$. Without loss of generality, we assume that $k>0$. We aim to prove the conclusion for $k+1$ and $l$. Lemma 2.18 gives us that
$H_{n}((k+1)X+lY+Z)+H_{n}(X,Y,Z)\leq H_{n}(Z+lY,X)+H_{n}(X+Z,kX+lY)+O_{k,l}(1),$
where the random variable $Z$ is taken such that $X,Y,Z$ are independent and $X$, $Z$ are identically distributed. Combining Corollary 2.22, we obtain that
$H_{n}((k+1)X+lY)\leq H_{n}((k+1)X+lY+Z)+O_{k,l}(1)$
$\leq H_{n}(Z+lY,X)+H_{n}(X+Z,kX+lY)-H_{n}(X,Y,Z)+O_{k,l}(1)$
$\leq H_{n}(Z+lY)+H_{n}(X+Z)+H_{n}(kX+lY)-H_{n}(Y)-H_{n}(Z)+O_{k,l}(1).$
Combining (13) and the inductive hypothesis on $k$, one sees that the conclusion holds with some $d_{1}(k,l)$ and $d_{2}(k,l)$ satisfying
$d_{1}(k+1,l)=d_{1}(1,l)+d_{1}(k,l)+1,\quad d_{2}(k,l)=d_{2}(1,l)+d_{2}(k,l)+2.$
By symmetry, we also have the induction on $l$, and the proof is completed for the situation that $k,l\in\mathbb{Z}^{\ast}$.
Second, when $k,l\in\mathbb{Q}^{\ast}$, we can write $k=k^{\prime}/t$ and $l=l^{\prime}/t$ for some $k^{\prime},l^{\prime},t\in\mathbb{Z}^{\ast}$. It follows that
$H_{n}(kX+lY)$ $=H_{n}(k^{\prime}X+l^{\prime}Y)+O_{t}(1)$
$\leq(d_{1}+d_{2}+1)H_{n}(X+Y)-d_{1}H_{n}(X)-d_{2}H_{n}(Y)+O_{k^{\prime},l^{\prime},t}(1)$
for some $d_{1}=d_{1}(k^{\prime},l^{\prime})$ and $d_{2}=d_{2}(k^{\prime},l^{\prime})$. The lemma then follows. ∎
Lemma 2.23 implies the following statement: For given $k,l\in\mathbb{Q}$, one has
$H_{n}(kX+lY)-H_{n}(X)\ll_{k,l}(H_{n}(X+Y)-H_{n}(X))+O_{k,l}(1)$
for any $n\in\mathbb{N}$ and any i.i.d. random variables $X$, $Y$ with bounded range. However, when $k,l$ are $\mathbb{Q}$-linearly independent, such statement may fail. In Example A.6, we would show the following: For any $K>0$, there exists some $N\in\mathbb{N}$ such that: for any $n>N$, there is a random variable $X$ with range contained in $[0,1]$ satisfying
$H_{n}(X+\sqrt{2}Y)-H_{n}(X)\geq K(H_{n}(X+Y)-H_{n}(X)),$
where $Y$ is i.i.d. with $X$.
2.4 Frostman measures
Now we introduce Frostman measures and their basic properties.
Definition 2.24.
Let $s,C>0$. We say a Borel measure $\mu$ on $\mathbb{R}^{m}$ is $(s,C)$-Frostman if for any $x\in\mathbb{R}^{m}$ and $r>0$
$\mu(B(x,r))\leq Cr^{s}.$
Similarly we also say a random variable $X$ is $(s,C)$-Frostman if its distribution measure is $(s,C)$-Frostman, meaning for any $x$ and $r>0$ we have
$P(X\in B(x,r))\leq Cr^{s}.$
For a random variable or a probability measure, when $0\leq s_{2}\leq s_{1}\leq 1$ and $1\leq C_{1}\leq C_{2}$, the $(s_{1},C_{1})$-Frostman condition can imply the $(s_{2},C_{2})$-Frostman condition.
The next lemma presents the Frostman conditions related to the arithmetic of random variables.
Lemma 2.25.
Let $X$ and $Y$ be independent real random variables that are both $(s,C)$-Frostman, where $0<s\leq 1$ and $C>0$. Then the following statements hold for $a,b,c,d\in\mathbb{R}$.
(i) The random variable $aX+bY$ is $(s,O_{a,b}(C))$-Frostman.
(ii) For any invertible linear change of variables $(x_{1},y_{1})=(ax+by,cx+dy)$, the random variable $(X_{1},Y_{1})=(aX+bY,cX+dY)$ is $(2s,O_{a,b,c,d}(C^{2}))$-Frostman.
Proof.
Let $\mu_{X}$, and $\mu_{Y}$ be the distribution measures associated to $X$ and $Y$, respectively.
(i) When $b\neq 0$, for any ball $B(y^{\prime},r)$ one has
[…]
A similar conclusion follows when $a\neq 0$. When $a=b=0$, the conclusion follows automatically.
(ii) For any ball $B((x,y),r)$,
$P((aX+bY,cX+dY)\in B((x,y),r)$ $\leq P(aX+bY\in B(x,r),cX+dY\in B(y,r)).$
If both $aX+bY\in B(x,r)$ and $cX+dY\in B(y,r)$ occur, we must have $X\in B\left(\frac{dx-by}{ad-bc},\frac{(|d|+|b|)r}{|ad-bc|}\right)$ and $Y\in B\left(\frac{-cx+ay}{ad-bc},\frac{(|c|+|a|)r}{|ad-bc|}\right)$. Therefore,
$P(aX+bY\in B(x,r),cX+dY\in B(y,r))\leq C\left(\frac{(|d|+|b|)r}{|ad-bc|}\right)^{s}\cdot C\left(\frac{(|c|+|a|)r}{|ad-bc|}\right)^{s}=O_{a,b,c,d}(C^{2}r^{2s}).$
The lemma now follows. ∎
The next property is taken from [11, Lemma 4.8].
Lemma 2.26.
If $X$ is $(s,C)$-Frostman for some $s,C>0$, then $H_{n}(X)\geq ns-\log C+O(1)$.
The function $x\mapsto x^{2}$ is weakly-bi-Lipschitz in a region that is bounded away from zero. The following lemma shows the relation between a random variable before and after cutting off the values near zero.
Lemma 2.27.
Let $Z$ be a real random variable with bounded range. Assume that there is some $s>0$ and $C\geq 1$ such that $P(|Z|<r)\leq Cr^{s}$ for any $r>0$. For any $0<\eta\leq 2^{-1}(2C)^{-1/s}$, denote by $E$ the event that $|Z|\geq\eta$. Then the random variable $Z_{E}$ satisfies
$H_{n}(Z_{E})=\frac{H_{n}(Z)}{P(E)}+O(C(2\eta)^{s}n)+O(1).$
for any all $n\in\mathbb{N}$.
Proof.
Recall that $Z_{E}$ is the random variable such that $Z_{E}\equiv(Z|E)$, and $P_{E}$ is the conditional probability measure given $E$.
Since $P(|Z|<\eta)\leq C\eta^{s}\leq 1/2$, we have $P(E)\geq 1-C\eta^{s}\geq 1/2$. Assume that $k_{0}$ is a natural number with $(k_{0}-1)2^{-n}\leq\eta<k_{0}2^{-n}$. Combining the conditions in this lemma, one deduces that
$H_{n}(Z_{E})$ $=-\sum\limits_{|k|\geq k_{0}+1}P_{E}(D_{n}(Z_{E})=k2^{-n})\log P_{E}(D_{n}(Z_{E})=k2^{-n})$
$\qquad+O\left(-\sum\limits_{k_{0}-1\leq|k|\leq k_{0}}P_{E}(D_{n}(Z_{E})=k2^{-n})\log P_{E}(D_{n}(Z_{E})=k2^{-n})\right)$
$=-\sum\limits_{|k|\geq k_{0}+1}\frac{P(D_{n}(Z)=k2^{-n})}{P(E)}\big{(}\log P(D_{n}(Z)=k2^{-n})-\log P(E)\big{)}+O(1)$
$=-\frac{1}{P(E)}\sum\limits_{|k|\geq k_{0}+1}P(D_{n}(Z)=k2^{-n})\log P(D_{n}(Z)=k2^{-n})+O(1).$
First, let us consider the case that $\eta>2^{-n+1}$, which gives $(k_{0}+1)2^{-n}\leq\eta+2^{-n+1}<2\eta$. Applying Lemma 2.10 with the fact
$\sum\nolimits_{|k|\leq k_{0}}P(D_{n}(Z)=k2^{-n})\leq P(|Z|<(k_{0}+1)2^{-n})\leq P(|Z|<2\eta)\leq C(2\eta)^{s}\leq 1/2,$
we see that
$-\sum\limits_{|k|\leq k_{0}}P(D_{n}(Z)=k2^{-n})\log P(D_{n}(Z)=k2^{-n})$
$\quad\quad\quad\leq C(2\eta)^{s}\cdot\log\big{(}2\eta 2^{n}+3\big{)}+O(1)\leq C(2\eta)^{s}n+O(1).$
Second, we consider the case that $\eta\leq 2^{-n+1}$. Then $k_{0}\leq 3$ and
$-\sum\nolimits_{|k|\leq k_{0}}P(D_{n}(Z)=k2^{-n})\log P(D_{n}(Z)=k2^{-n})=O(1).$
Finally, we conclude that
$H(Z_{E})$ $=\frac{1}{P(E)}\left(H_{n}(Z)+\sum\limits_{|k|\leq k_{0}}P(D_{n}(Z)=k2^{-n})\log P(D_{n}(Z)=k2^{-n})\right)+O(1)$
$=\frac{H_{n}(Z)}{P(E)}+O(C(2\eta)^{s}n)+O(1).$
In the following lemma, we show that the probability that the distance function takes small values is related to the Frostman condition.
Lemma 2.28.
Let $Z$, $Z^{\prime}$ be independent random variables. Suppose that $Z$ is $(s,C)$-Frostman, where $s,C>0$. Then
$P\big{(}|Z-Z^{\prime}|<r)\leq Cr^{s}.$
for any $r>0$.
Proof.
Denote by $\mu$ the distribution measure associated to $Z^{\prime}$. Since $Z$ is $(s,C)$-Frostman, one sees that
[…]
The following definition of Riesz energy is related to the distance measure.
Definition 2.29.
The $s$-Riesz energy of a Borel measure $\mu$ on $\mathbb{R}^{d}$ is defined by
$I_{s}(\mu)=\int\int\frac{d\mu(x)d\mu(y)}{|x-y|^{s}}.$
For any given measure $\mu$, define the measure $\Delta(\mu)$ by
[…]
for any Borel set $B$. That is to say, $\Delta(\mu)$ is the image of $\mu\times\mu$ under the distance map $(x,y)\mapsto|x-y|$. We now recall a result from Mattila in [12, Section 4.2, page 61].
Lemma 2.30 ([12]).
Let $\mu$ be a probability measure with $\operatorname{spt}(\mu)\subset B(0,R)\subset\mathbb{R}^{2}$, for some $R>0$. Then, for any $0<t\leq 1$ and for real numbers $0<\eta<r<2R$, we have $\Delta(\mu)([r,r+\eta])\ll_{t}r^{1/2}\eta^{t}I_{t+1/2}(\mu)$.
Lemma 2.31.
Let $\mu$ be a $(2s,C)$-Frostman probability measure with $\operatorname{spt}(\mu)\subset B(0,R)\subset\mathbb{R}^{2}$, where $\frac{1}{4}<s\leq 1$, $C\geq 1$ and $R>0$. Then for any $r,\eta,t$ with $0<\eta<r<2R$, $0<t\leq 1$ and $t<2s-\frac{1}{2}$, it satisfies that
$\Delta(\mu)([r,r+\eta])=O_{t}\left(\frac{CR^{2s-t}\eta^{t}}{2s-t-1/2}\right)+O_{t}(R^{-t}\eta^{t}).$
Proof.
Since $\mu$ is $(2s,C)$-Frostman, for $20<t<2s-1/$ 2 and any $x\in\mathbb{R}^{2}$,
[…]
$\qquad=O_{t}\left(\frac{CR^{2s-t-1/2}}{2s-t-1/2}\right)+O_{t}(R^{-t-1/2}):=M.$
Thus
[…]
When $0<\eta<r<2R$ and $0<t\leq 1$, one obtains by Lemma 2.30 that
$\Delta(\mu)([r,r+\eta])\ll_{t}r^{1/2}\eta^{t}I_{t+1/2}(\mu)=O_{t}\left(\frac{CR^{2s-t}\eta^{t}}{2s-t-1/2}\right)+O_{t}(R^{-t}\eta^{t}).$
2.5 Entropy of the distance function
In this subsection, we show some results of the discretized entropy of the distance function of certain measures, together with the square of the distance. These theorems are the key ingredients of the proof of Theorem 1.1.
Theorem 2.32.
Let $W$ be a $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $11/2<s<$ 1, $C\geq 1$ and $R>0$. Let $W^{\prime}$ be a random variable which is i.i.d with $W$. Then for any number $t$ satisfying $t\leq 1$ and $s<t<2s-1/2$, we have
$H_{n}(|W-W^{\prime}|)\geq nt-\log C+O(|\log R|)+O(|\log(2s-t-1/2)|)+O_{t}(1).$ (14)
for all $n\in\mathbb{N}$.
Proof.
Let $\mu$ be the distribution measure associated to $W$. Then $\Delta(\mu)$ is the measure associated to $Z=|W-W^{\prime}|$. Write $N:=\lceil 4R^{2}2^{n}\rceil-1$ and $p_{k}:=P\big{(}D_{n}(Z)=k2^{-n}\big{)}$ for $0\leq k\leq N$. For $0\leq k\leq 3$, Lemma 2.28 shows that
$p_{k}\leq P(D_{n}(Z\leq 2^{-n+2}))\leq C(2^{-n+2})^{2s}\ll C2^{-nt}.$
When $4\leq k\leq N$, it follows from Lemma 2.31 that
$p_{k}$ $=\Delta(\mu)\big{(}[k2^{-n},(k+1)2^{-n}\big{)}=O_{t}\left(\frac{CR^{2s-t}2^{-nt}}{2s-t-1/2}\right)+O_{t}(R^{-t}2^{-nt}).$
It follows that
$H_{n}(Z)$ $=-\sum\limits_{k=0}^{3}p_{k}\log p_{k}-\sum\limits_{k=4}^{N}p_{k}\log p_{k}$
$\geq\sum\limits_{k=0}^{3}p_{k}(nt-\log C+O(1))$
$\qquad+\sum\limits_{k=4}^{N}p_{k}\big{(}nt-\log C+O(|\log R|)+O(|\log(2s-t-1/2)|)+O_{t}(1)\big{)}$
$=nt-\log C+O(|\log R|)+O(|\log(2s-t-1/2)|)+O_{t}(1).$
Remark 2.33.
The parameters $C$ and $R$ in Theorem 2.32 are allowed to be unbounded as $n$ tends to infinity. In Example A.5, we show an example in which $C$ is growing with $n$. This example will tell us that the term $-\log C$ on the right-hand side of (14) is essential.
Remark 2.34.
When $13/4<s<$ 1, we can take $t=1$ in Theorem 2.32. Combining this with (11), we then obtain the identity
$H_{n}(|W-W^{\prime}|)=n+O(\log C)+O(|\log R|)+O_{s}(1).$
Theorem 2.35.
Let $W$ be a $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $11/2<s<$ 1, $C\geq 1$ and $R>0$. Let $W^{\prime}$ be a random variable which is i.i.d with $W$. Then there exists a constant $\epsilon_{0}=\epsilon_{0}(s)>0$ such that
$H_{n}(|W-W^{\prime}|^{2})\geq n(s+\epsilon_{0})+O(\log C)+O(|\log R|)+O_{s}(1)$
for all $n\in\mathbb{N}$.
Proof.
Let $Z$, $t$ be as in Theorem 2.32, with $Z=|W-W^{\prime}|$ and
$H_{n}(Z)\geq nt-\log C+O(|\log R|)+O(|\log(2s-t-1/2)|)+O_{t}(1).$
By Lemma 2.28, one has $P(|Z|<r)\leq Cr^{2s}$ for any $r>0$. Let $\eta$ be a parameter with $0<\eta\leq 2^{-1}(2C)^{-1/2s}$ to be determined later. Denote by $E$ the event that $|Z|\geq\eta$. By Lemma 2.27, we deduce that
$H_{n}(Z_{E})$ $=\frac{H_{n}(Z)}{P(E)}+O(C(2\eta)^{2s}n)+O(1)$
$\geq\frac{nt}{P(E)}+O(\log C)+O(|\log R|)+O(|\log(2s-t-1/2)|)+O_{t}(1)+O(C\eta^{2s}n).$
Since $Z_{E}^{2}$ takes value at least $\eta^{2}$ with probability $1$, the random variable $Z_{E}$ is a $\frac{1}{2\eta}$-Lipschitz function of ${Z_{E}}^{2}$. Combining Lemma 2.16, we have
$H_{n}(Z_{E})\leq H_{n}(Z_{E}^{2})+O\left(\log\frac{1}{2\eta}\right).$
Applying Lemma B.2, one deduces that
$H_{n}(Z^{2})$ $\geq P(E)\,H_{n}(Z_{E}^{2})\geq P(E)\left(H_{n}(Z_{E})+O\left(\log\frac{1}{\eta}\right)+O(1)\right)$
$\geq P(E)\left(\frac{nt}{P(E)}+O(\log C)+O(|\log R|)\right.$
$\qquad\qquad\left.+O(|\log(2s-t-1/2)|)+O_{t}(1)+O(C\eta^{2s}n)+O\left(\log\frac{1}{\eta}\right)\right).$ (15)
Now we take $t=s+\frac{2s-1}{3}$ when $41/2<s\leq 3/$ 4, and take $t=1$ when $13/4<s<$ 1. Moreover, assume that the implied constant in the term $O(C\eta^{2s}n)$ in (15) is $c_{0}$, with $c_{0}\geq 2$. We take $\eta$ to be the number satisfying
$c_{0}C\eta^{2s}=\begin{cases}(2s-1)/6,\quad&\text{if }1/2<s\leq 3/4,\\ (2-2s)/3,\quad&\text{if }3/4<s<1.\\ \end{cases}$
One can verify that the requirements involving $t$ and $\eta$ are satisfied, and $12s-t-1/2\gg_{s}$ 1. Also, $\log(1/\eta)<\log C+O_{s}(1)$. As a result, we conclude that
$H_{n}(Z^{2})\geq n(s+\epsilon_{0})+O(\log C)+O(|\log R|)+O_{s}(1),$
with $\epsilon_{0}=(2s-1)/6$ when $41/2<s\leq 3/$ 4, and $\epsilon_{0}=(1-s)/3$ when $13/4<s<$ 1. ∎
The $\epsilon_{0}$ in the above theorem can be calculated and optimized explicitly. However, we do not pursue that direction in this paper.
The final lemma of this section shows a trivial bound of the discretized entropy of distance function, which is applicable for small $s$. A non-trivial bound for small $s$ is delayed to Section 6 (see Theorem 6.1).
Lemma 2.36.
Let $X$, $X^{\prime}$, $Y$, $Y^{\prime}$ be independent real random variables. Suppose that $X$ and $X^{\prime}$ are both $(s,C)$-Frostman with ranges contained in $[-c,c]$, where $0<s\leq 1$ and $C,c\geq 1$. Then, for any $n\in\mathbb{N}$,
$H_{n}\big{(}(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}\big{)}\geq ns+O_{s}(\log C)+O_{s}(\log n),$
and
$H_{n}\big{(}\sqrt{(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}}\big{)}\geq ns+O_{s}(\log C)+O(\log c)+O_{s}(\log n).$
Proof.
By Corollary 2.22, one has
$H_{n}\big{(}(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}\big{)}\geq H_{n}\big{(}(X-X^{\prime})^{2}\big{)}+O(1).$
By Lemma 2.25, the random variable $Z:=X-X^{\prime}$ is $(s,O(C))$-Frostman. By Lemma 2.26, one has $H_{n}(Z)\geq ns-\log C+O(1)$. Also, $range(Z)\subseteq[-2c,2c]$. Let $E$ be the event that $|Z|\geq\eta$ as in Lemma 2.27. Follow the same arguments as in the proof of Theorem 2.35, we obtain that
$H(Z^{2})\geq ns+O(\log C)+O(C\eta^{s}n)+O\big{(}\log\frac{1}{\eta}\big{)}+O(1).$
The first inequality follows by taking $\eta=2^{-1}\big{(}\frac{\log n}{2Cn}\big{)}^{1/s}$. The second inequality follows from the fact that
$H_{n}\big{(}(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}\big{)}\leq H_{n}\big{(}\sqrt{(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}}\big{)}+O(\log c).$
3 Discretized entropy Balog-Szemerédi-Gowers theorem
In this section, we follow the method introduced by Tao in [22, Theorem 3.1] to obtain a generalized discretized entropy version of the Balog-Szemerédi-Gowers theorem. We do not know whether such a generalization has been studied in the literature, so we provide a complete proof in this section. The Balog-Szemerédi-Gowers theorem (or one of its many variants) says that if one has a finite subset of the integers whose sums line up a lot, then it must not have many sums. More precisely, given $A\subseteq\mathbb{Z},$ whose elements $a_{j}\in A$ have $\gamma|A|^{3}$ relations of the form $a_{1}+a_{2}=a_{3}+a_{4},$ there must be a large subset $A^{\prime}\subseteq A,$ where $|A^{\prime}|\geq C_{1}(\gamma)|A|$ that satisfies $|A^{\prime}+A^{\prime}|\leq C_{2}(\gamma)|A^{\prime}|,$ where the constants $C_{j}$ above are positive and only depend on $\gamma.$
Theorem 3.1.
Let $n$ be a natural number, and $(X,Y)$ be a given random variable. Suppose that
$H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\leq\log K_{1}.$ (16)
Then, there exists a random variable $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$ such that
(i) $(X^{\prime},Y^{\prime})\equiv(X,Y)$;
(ii) $X^{\prime}$ and $X^{\prime\prime}$ are conditionally i.i.d. given $D_{n}(Y^{\prime})$;
(iii) $Y^{\prime}$ and $Y^{\prime\prime}$ are conditionally i.i.d. given $D_{n}(X^{\prime})$;
(iv) $(X^{\prime\prime},Y^{\prime})$ and $Y^{\prime\prime}$ are conditionally independent given $D_{n}(X^{\prime})$.
In particular, the random variables $X^{\prime\prime}$ and $Y^{\prime\prime}$ are conditionally independent given $D_{n}(X^{\prime},Y^{\prime})$. Moreover,
$H_{n}(X^{\prime\prime}|X^{\prime},Y^{\prime})\geq H_{n}(X)-\log K_{1},$ (17)
$H_{n}(Y^{\prime\prime}|X^{\prime},Y^{\prime})\geq H_{n}(Y)-\log K_{1}.$ (18)
In addition, for any weakly $\kappa$-bi-Lipschitz functions $f$ and $g$ with
$H_{n}(f(X)+g(Y))-\frac{H_{n}(X)+H_{n}(Y)}{2}\leq\log K_{2},$ (19)
we have that
$H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|X^{\prime},Y^{\prime})-\frac{H_{n}(X)+H_{n}(Y)}{2}\leq 4\log K_{1}+3\log K_{2}+O(\log\kappa).$ (20)
Remark 3.2.
The original version of entropy Balog-Szemerédi-Gowers theorem in [22] states for $H$ instead of $H_{n}$, does not involve the functions $f$, $g$, and conditions (ii)–(iv) are “given $Y^{\prime}$, $X^{\prime}$, $X^{\prime}$” instead of “given $D_{n}(Y^{\prime})$, $D_{n}(X^{\prime})$, $D_{n}(X^{\prime})$”, respectively. In such a situation, (i)–(iv) is equivalent to the statement that $X^{\prime},Y,X^{\prime\prime},Y^{\prime\prime}$ forms a Markov chain.
The following lemmas are needed in the proof of this theorem.
Lemma 3.3.
Let $n$ be a natural number, and $(X,Y)$ be a given random variable. Then there exists a random variable $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$ such that (i)–(iv) in Theorem 3.1 hold, with $X^{\prime\prime}$ and $Y^{\prime\prime}$ being conditionally independent given $D_{n}(X^{\prime},Y^{\prime})$.
Proof.
Applying Lemma B.1 with $A=0$, $B=X$, $C=Y$ and $h=D_{n}$, we obtain a random variable $(X_{1},Y_{1},X_{2})$ such that: $(X_{1},Y_{1})\equiv(X,Y)$; $X_{2}$ and $X_{1}$ are conditionally i.i.d. given $D_{n}(Y_{1})$.
Next, let us apply Lemma B.1 again, with $A=X_{2}$, $B=Y_{1}$, $C=X_{1}$ and $h=D_{n}$. There exists a random variable $(X_{2}^{\prime},Y_{1}^{\prime},X_{1}^{\prime},Y_{1}^{\prime\prime})$ such that: $(X_{2}^{\prime},Y_{1}^{\prime},X_{1}^{\prime})\equiv(X_{2},Y_{1},X_{1})$, $(Y_{1}^{\prime\prime},D_{n}(X_{1}^{\prime}))\equiv(Y_{1},D_{n}(X_{1}))$; $Y_{1}^{\prime\prime}$ are $Y_{1}^{\prime}$ are conditionally i.i.d. given $D_{n}(X_{1}^{\prime})$; also $(X_{2}^{\prime},Y_{1}^{\prime})$ and $Y_{1}^{\prime\prime}$ are conditionally independent given $D_{n}(X_{1}^{\prime})$.
One can also use $A=(X_{2},X_{1},D_{n}(Y_{1}))$ in the second application of Lemma B.1 to see $(X_{2}^{\prime},X_{1}^{\prime},D_{n}(Y_{1}^{\prime}))\equiv(X_{2},X_{1},D_{n}(Y_{1}))$. Now (i)–(iv) hold by taking $X^{\prime}=X_{1}^{\prime}$, $X^{\prime\prime}=X_{2}^{\prime}$, $Y^{\prime}=Y_{1}^{\prime}$, $Y^{\prime\prime}=Y_{1}^{\prime\prime}$.
Moreover, (iii) implies that $Y^{\prime\prime}$ and $D_{n}(Y^{\prime})$ are conditionally independent given $D_{n}(X^{\prime})$. While (iv) gives that $(X^{\prime\prime},D_{n}(Y^{\prime}))$ and $Y^{\prime\prime}$ are conditionally independent given $D_{n}(X^{\prime},Y^{\prime})$. It follows that $X^{\prime\prime}$ and $Y^{\prime\prime}$ are conditionally independent given $(D_{n}(X^{\prime}),D_{n}(Y^{\prime}))$.
Finally, we mention that the random variables $D_{n}(X,Y)$, $D_{n}(X^{\prime},Y^{\prime})$, $D_{n}(X^{\prime},Y^{\prime\prime})$, $D_{n}(X^{\prime\prime},Y^{\prime})$ all have the same distribution. ∎
Corollary 3.4.
Let $n$ be a natural number and $(X,Y)$ be a given random variable. Let $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$ be the random variable shown in Lemma 3.3. For any $\kappa$-Lipschitz function $h$, the following properties hold.
(i) $H_{n}(h(X^{\prime},Y^{\prime}))=H_{n}(h(X,Y))$.
(ii) $H_{n}(h(X^{\prime},Y^{\prime\prime}))=H_{n}(h(X,Y))+O(\log\kappa)$.
(iii) $H_{n}(h(X^{\prime\prime},Y^{\prime}))=H_{n}(h(X,Y))+O(\log\kappa)$.
Proof.
Since $(X^{\prime},Y^{\prime})\equiv(X,Y)$, then (i) follows. Applying Lemma 2.19, yields
$H_{n}(h(X^{\prime},Y^{\prime}))=H_{n}(h(D_{n}(X^{\prime}),D_{n}(Y^{\prime})))+O(\log\kappa),$
$H_{n}(h(X^{\prime},Y^{\prime\prime}))=H_{n}(h(D_{n}(X^{\prime}),D_{n}(Y^{\prime\prime})))+O(\log\kappa).$
Noting that $(Y^{\prime},D_{n}(X^{\prime}))\equiv(Y^{\prime\prime},D_{n}(X^{\prime}))$, one asserts that
$H_{n}(h(D_{n}(X^{\prime}),D_{n}(Y^{\prime})))=H_{n}(h(D_{n}(X^{\prime}),D_{n}(Y^{\prime\prime}))).$
Combining this with (i), one gets (ii). Claim (iii) follows by a similar argument. ∎
Lemma 3.5.
With $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$ and $f,g$ as stated in Theorem 3.1, we always have
$H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa).$
Proof.
Let $Z$ be a random variable such that $Z,Y^{\prime}$ are conditionally i.i.d. given $D_{n}(X^{\prime},X^{\prime\prime})$.
Recall that $f,g$ are both weakly $\kappa$-bi-Lipschitz. Applying Lemma 2.16 and Lemma 2.18, one can verify that
$H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})+H_{n}(X^{\prime},X^{\prime\prime},Z,Y^{\prime})$
$\qquad=H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})+H_{n}(X^{\prime},X^{\prime\prime},g(Z),Y^{\prime})+O(\log\kappa)$
$\qquad\leq H_{n}(f(X^{\prime})+g(Z),f(X^{\prime\prime})+g(Z),Y^{\prime})+H_{n}(X^{\prime},X^{\prime\prime},Y^{\prime})+O(\log\kappa).$ (21)
Indeed, we have $(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})=f_{1}(f(X^{\prime})+g(Z),f(X^{\prime\prime})+g(Z),Y^{\prime})=f_{2}(X^{\prime},X^{\prime\prime},Y^{\prime})$ and and
$(X^{\prime},X^{\prime\prime},g(Z),Y^{\prime})=f_{3}(f(X^{\prime})+g(Z),f(X^{\prime\prime})+g(Z),Y^{\prime},X^{\prime},X^{\prime\prime}))$
with the $O(\kappa)$-Lipschitz functions
$f_{1}(z_{1},z_{2},z_{3})=(z_{1}-z_{2},z_{3}),$
$f_{2}(z_{1},z_{2},z_{3})=(f(z_{1})-f(z_{2}),z_{3}),$
$f_{3}(z_{1},z_{2},\ldots,z_{5})=(z_{4},z_{5},z_{1}-f(z_{4}),z_{3}).$
Note that $D_{n}(Z)$, $D_{n}(Y^{\prime})$ are conditionally i.i.d. given $D_{n}(X,X^{\prime})$. Also, $D_{n}(X^{\prime})$, $D_{n}(X^{\prime\prime})$ are conditionally i.i.d. given $D_{n}(Y^{\prime})$. By Lemmas 2.8 and 2.19,
$H_{n}(X^{\prime},X^{\prime\prime},Z,Y^{\prime})=2H_{n}(X^{\prime},X^{\prime\prime},Y^{\prime})-H_{n}(X^{\prime},X^{\prime\prime})$
$=\big{(}2H_{n}(X^{\prime},Y^{\prime})-H_{n}(Y^{\prime})\big{)}+H_{n}(X^{\prime},X^{\prime\prime},Y^{\prime})-H_{n}(X^{\prime},X^{\prime\prime})$
$=H_{n}(X^{\prime},X^{\prime\prime},Y^{\prime})-H_{n}(X^{\prime},X^{\prime\prime})+2H_{n}(X,Y)-H_{n}(Y)$
$\geq H_{n}(X^{\prime},X^{\prime\prime},Y^{\prime})+2H_{n}(X,Y)-2H_{n}(X)-H_{n}(Y),$ (22)
where the last step uses the fact $H_{n}(X^{\prime},X^{\prime\prime})\leq H_{n}(X^{\prime})+H_{n}(X^{\prime\prime})=2H_{n}(X)$.
Furthermore, since $(D_{n}(X^{\prime}),D_{n}(X^{\prime\prime}),Z)\equiv(D_{n}(X^{\prime}),D_{n}(X^{\prime\prime}),Y^{\prime})$, we use Lemma 2.19 and Corollary 3.4 to deduce that
$H_{n}(f(X^{\prime\prime})+g(Z))=H_{n}(f(D_{n}(X^{\prime\prime}))+g(D_{n}(Z)))+O(\log\kappa)$
$\qquad=H_{n}(f(D_{n}(X^{\prime\prime}))+g(D_{n}(Y^{\prime})))+O(\log\kappa)=H_{n}(f(X^{\prime\prime})+g(Y^{\prime}))+O(\log\kappa)$
$\qquad=H_{n}(f(X)+g(Y))+O(\log\kappa).$ (23)
Similarly,
$H_{n}(f(X^{\prime})+g(Z))=H_{n}(f(X)+g(Y))+O(\log\kappa).$ (24)
Combining Lemma 2.4(iv) and the fact $H_{n}(Y^{\prime})=H_{n}(Y)$, one concludes that
$H_{n}(f(X^{\prime\prime})+g(Z),f(X^{\prime})+g(Z),Y^{\prime})\leq 2H_{n}(f(X)+g(Y))+H_{n}(Y)+O(\log\kappa).$ (25)
Putting formulae (16), (19), and (21)–(25) together, we get that
$H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})\leq 2H_{n}(f(X)+g(Y))+2H_{n}(X)+2H_{n}(Y)-2H_{n}(X,Y)+O(\log\kappa).$
$\qquad\leq\big{(}2H_{n}(f(X)+g(Y))-H_{n}(X)-H_{n}(Y)\big{)}$
$\qquad\qquad+2\big{(}H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\big{)}+H_{n}(X)+H_{n}(Y)+O(\log\kappa).$
$\qquad\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa).$
We are ready to present a proof of Theorem 3.1
Proof of Theorem 3.1.
First, since $D_{n}(Y^{\prime})$, $D_{n}(Y^{\prime\prime})$ are conditionally i.i.d given $D_{n}(X^{\prime})$, one gets from (16) that
$H_{n}(Y^{\prime\prime}|Y^{\prime},X^{\prime})$ $=H_{n}(Y^{\prime\prime}|X^{\prime})=H_{n}(X^{\prime},Y^{\prime\prime})-H_{n}(X^{\prime})$
$=H_{n}(X,Y)-H_{n}(X)\geq H_{n}(Y)-\log K_{1}.$
So (18) holds, and (17) follows by similar arguments.
For (20), Using Lemma 2.4(iv), one has
$\begin{split}&H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|X^{\prime},Y^{\prime})\leq H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|Y^{\prime})=H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime}),Y^{\prime})-H_{n}(Y^{\prime}).\\ \end{split}$ (26)
Using Lemmas 2.16 and 2.18, we have that
$H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime}),Y^{\prime})+H_{n}(X^{\prime},Y^{\prime},X^{\prime\prime},Y^{\prime\prime})$
$\qquad=H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime}),Y^{\prime})+H_{n}(f(X^{\prime}),Y^{\prime},X^{\prime\prime},Y^{\prime\prime})+O(\log\kappa)$
$\qquad\leq H_{n}(f(X^{\prime\prime})-f(X^{\prime}),f(X^{\prime})+g(Y^{\prime\prime}),Y^{\prime})+H_{n}(X^{\prime\prime},Y^{\prime\prime},Y^{\prime})+O(\log\kappa).$ (27)
Recall that $D_{n}(X^{\prime\prime})$ and $D_{n}(Y^{\prime\prime})$ are conditionally independent given $D_{n}(X^{\prime},Y^{\prime})$. Using Lemma 2.8 and the assumption (16), one has
$H_{n}(X^{\prime},Y^{\prime},X^{\prime\prime},Y^{\prime\prime})$ $=H_{n}(X^{\prime},Y^{\prime})+H_{n}(X^{\prime\prime},Y^{\prime\prime}|X^{\prime},Y^{\prime})$
$=H_{n}(X^{\prime},Y^{\prime})+H_{n}(X^{\prime\prime}|X^{\prime},Y^{\prime})+H_{n}(Y^{\prime\prime}|X^{\prime},Y^{\prime})$
$=H_{n}(X^{\prime\prime},X^{\prime},Y^{\prime})+H_{n}(Y^{\prime\prime},X^{\prime},Y^{\prime})-H_{n}(X^{\prime},Y^{\prime})$
$=\big{(}2H_{n}(X^{\prime},Y^{\prime})-H_{n}(Y^{\prime})\big{)}+\big{(}2H_{n}(X^{\prime},Y^{\prime})-H_{n}(X^{\prime})\big{)}-H_{n}(X^{\prime},Y^{\prime})$
$=H_{n}(X,Y)+H_{n}(X)+H_{n}(Y)-2\big{(}H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\big{)}$
$\geq H_{n}(X,Y)+H_{n}(X)+H_{n}(Y)-2\log K_{1}.$ (28)
Moreover,
$H_{n}(X^{\prime\prime},Y^{\prime\prime},Y^{\prime})\leq H_{n}(X^{\prime\prime},Y^{\prime})+H_{n}(Y^{\prime\prime})=H_{n}(X,Y)+H_{n}(Y).$ (29)
Furthermore, by Corollary 3.4, Lemma 3.5, and (19),
$H_{n}(f(X^{\prime\prime})-f(X^{\prime}),f(X^{\prime})+g(Y^{\prime\prime}),Y^{\prime})$
$\qquad\leq H_{n}(f(X^{\prime\prime})-f(X^{\prime}),Y^{\prime})+H_{n}(f(X^{\prime})+g(Y^{\prime\prime}))$
$\qquad\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa)+\frac{H_{n}(X)+H_{n}(Y)}{2}+\log K_{2}$
$\qquad=\frac{3}{2}(H_{n}(X)+H_{n}(Y))+2\log K_{1}+3\log K_{2}+O(\log\kappa).$ (30)
Combining (26)–(30), one concludes that
$H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|X^{\prime},Y^{\prime})\leq\frac{H_{n}(X)+H_{n}(Y)}{2}+4\log K_{1}+3\log K_{2}+O(\log\kappa).$
Thus, the estimate (20) follows. ∎
4 Proof of Theorem 1.1 with $s>1/2$: diagonal forms
In this section, we consider the diagonal real quadratic form $\phi(x,y)=\alpha x^{2}+\beta y^{2}$. When we write $O_{\phi}(\cdot)$, it means that the implied constant may depend on the coefficients $\alpha$ and $\beta$ that occurred in $\phi$.
Lemma 4.1.
Let $\phi(x,y)=\alpha x^{2}+\beta y^{2}$, where $\alpha,\beta\in\mathbb{R}^{\ast}$. Let $X$, $Y$ be independent random variables with $range_{+}(X)\cup range_{+}(Y)\subseteq[-c_{1},-c_{2}]\cup[c_{2},c_{1}]$, where $0<c_{2}<c_{1}$. Suppose that $X$ and $Y$ are $(s,C_{1})$-Frostman and $(s,C_{2})$-Frostman, respectively, with $11/2<s<$ 1 and $C_{1},C_{2}\geq 1$. Then there exists a constant $\epsilon_{1}=\epsilon_{1}(s)>0$ such that
$\max\{H_{n}(X+Y),~{}H_{n}(\phi(X,Y))\}\geq n(s+\epsilon_{1})+O(\log C_{1}C_{2})+O_{\phi,s,c_{1},c_{2}}(1)$
for all $n\in\mathbb{N}$.
Remark 4.2.
Compared to Theorem 1.1, this provides a stronger statement, namely, for all non-zero real-coefficients $\alpha$ and $\beta$.
Proof.
Write $\phi(x,y)=p(x)+q(y)$, where $p(x)=\alpha x^{2}$ and $q(y)=\beta y^{2}$ are functions whose domains are both taken to be $[-c_{2},-c_{1}]\cup[c_{2},c_{1}]$. One asserts by Example 2.17 that
$H_{n}(p(X))=H_{n}(X)+O_{\phi,c_{1},c_{2}}(1),\quad H_{n}(q(Y))=H_{n}(Y)+O_{\phi,c_{1},c_{2}}(1).$ (31)
Define $W=(\sqrt{|\alpha|}X,\sqrt{|\beta|}Y),$ with $|\alpha|X$ and $|\beta|Y$ being $(s,O_{\phi}(C_{1}))$-Frostman and $(s,O_{\phi}(C_{2}))$-Frostman, respectively. Let $W_{1}=(\sqrt{|\alpha|}X_{1},\sqrt{|\beta|}Y_{1})$, $W_{2}=(\sqrt{|\alpha|}X_{2},\sqrt{|\beta|}Y_{2})$ be i.i.d. copies of $W$. Then $(X_{1},Y_{1})$, $(X_{2},Y_{2})$ are i.i.d. copies of $(X,Y)$, and $X_{1},X_{2},Y_{1},Y_{2}$ are independent.
Let us first consider the case that $\alpha$ and $\beta$ have the same sign. Since $H_{n}(\phi(A,B))=H_{n}(-\phi(A,B))$, we assume without loss of generality that $\alpha,\beta>0$. Then
$|W_{1}-W_{2}|^{2}=\alpha(X_{1}-X_{2})^{2}+\beta(Y_{1}-Y_{2})^{2}.$
Calculation reveals that
$\alpha(X_{1}-X_{2})^{2}+\beta(Y_{1}-Y_{2})^{2}=2(\alpha(X_{1}^{2}+X_{2}^{2})+\beta(Y_{1}^{2}+Y_{2}^{2}))-\alpha(X_{1}+X_{2})^{2}-\beta(Y_{1}+Y_{2})^{2}.$
So $|W_{1}-W_{2}|^{2}=f_{1}(X_{1}+X_{2},Y_{1}+Y_{2},p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}))$, where
$f_{1}(z_{1},z_{2},z_{3})=2z_{3}-\alpha z_{1}^{2}-\beta z_{2}^{2}.$
From Example 2.12 and the fact that $range(X_{1}+X_{2}),range(Y_{1}+Y_{2})\subseteq[-2c_{1},2c_{1}]$, one sees that $f_{1}$ is $O_{\phi,c_{1}}(1)$-Lipschitz. We also have that $|W_{1}-W_{2}|^{2}=f_{2}(X_{1}-X_{2},Y_{1}-Y_{2})$, with the $O_{\phi,c_{1}}(1)$-Lipschitz function
$f_{2}(z_{1},z_{2})=\alpha z_{1}^{2}+\beta z_{2}^{2}.$
Moreover,
$(X_{1},X_{2},Y_{1},Y_{2})=f_{3}(X_{1}+X_{2},Y_{1}+Y_{2},p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}),X_{1}-X_{2},Y_{1}-Y_{2}),$
with the $O(1)$-Lipschitz function
$f_{3}(z_{1},z_{2},z_{3},z_{4},z_{5})=\left(\frac{z_{1}+z_{4}}{2},\frac{z_{1}-z_{4}}{2},\frac{z_{2}+z_{5}}{2},\frac{z_{2}-z_{5}}{2}\right).$
Applying Lemma 2.18 gives
$H_{n}(|W_{1}-W_{2}|^{2})+H_{n}(X_{1},X_{2},Y_{1},Y_{2})$
$\leq H_{n}(X_{1}+X_{2},Y_{1}+Y_{2},p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}))+H_{n}(X_{1}-X_{2},Y_{1}-Y_{2})+O_{\phi,c_{1}}(1)$
$\leq H_{n}(X_{1}+X_{2})+H_{n}(Y_{1}+Y_{2})+H_{n}(p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}))$
$\qquad+H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})+O_{\phi,c_{1}}(1).$ (32)
We now estimate the terms in the last expression. For the first term, apply Corollary 2.22 to get that
$H_{n}(X_{1}+X_{2})$ $\leq H_{n}(X_{1}+Y_{1})+H_{n}(X_{2}+Y_{1})-H_{n}(Y_{1})+O(1),$
$=2H_{n}(X+Y)-H_{n}(Y)+O(1).$
One can argue similarly for $H_{n}(Y_{1}+Y_{2})$, hence,
$H_{n}(X_{1}+X_{2})+H_{n}(Y_{1}+Y_{2})\leq 4H_{n}(X+Y)-H_{n}(X)-H_{n}(Y)+O(1).$ (33)
For the third term, we use Corollary 2.22 repeatedly, and combine Lemma 2.16 and formula (31), to get
$H_{n}(p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}))$
$\leq H_{n}(p(X_{1})+p(X_{2})+q(Y_{1}))+H_{n}(p(X_{1})+q(Y_{1})+q(Y_{2}))-H_{n}(p(X_{1})+q(Y_{1}))+O(1)$
$\leq H_{n}(p(X_{1})+q(Y_{1}))+H_{n}(p(X_{2})+q(Y_{1}))-H_{n}(q(Y_{1}))$
$\qquad+H_{n}(p(X_{1})+q(Y_{2}))+H_{n}(p(X_{1})+q(Y_{1}))-H_{n}(p(X_{1}))$
$\qquad-H_{n}(p(X_{1})+q(Y_{1}))+O(1)$
$=3H_{n}(p(X)+q(Y))-H_{n}(X)-H_{n}(Y)+O_{\phi,c_{1},c_{2}}(1).$ (34)
For the fourth term, applying Corollary 2.22 shows that
$H_{n}(X_{1}-X_{2})\leq H_{n}(X_{1}+Y_{1})+H_{n}(X_{2}+Y_{1})-H_{n}(Y_{1})+O(1).$
So similarly
$H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})\leq 4H_{n}(X+Y)-H_{n}(X)-H_{n}(Y)+O(1).$ (35)
Combining (32)–(35) and Lemma 2.4(iv), we have that
$H_{n}\big{(}|W_{1}-W_{2}|^{2}\big{)}+5H_{n}(X)+5H_{n}(Y)\leq 8H_{n}(X+Y)+3H_{n}(\phi(X,Y))+O_{\phi,c_{1},c_{2}}(1).$
Note that $X$ and $Y$ are independent, and are $(s,C_{1})$-Frostman and $(s,C_{2})$-Frostman, respectively. One asserts that $W$ is $(2s,C_{1}C_{2}(\alpha\beta)^{-s/2})$-Frostman. Noting that $range(W)$ is contained in the ball in ${\mathbb{R}}^{2}$ centered at origin with radius $O_{\phi,c_{1}}(1)$. So we can deduce by Lemma 2.26 and Theorem 2.35 that
$H_{n}\big{(}X\big{)}\geq ns-\log C_{1}+O(1),\quad H_{n}\big{(}Y\big{)}\geq ns-\log C_{2}+O(1),$
$H_{n}\big{(}|W_{1}-W_{2}|^{2}\big{)}\geq n(s+\epsilon_{0})+O(\log C_{1}C_{2})+O_{\phi,s,c_{1},c_{2}}(1)$
for some $\epsilon_{0}=\epsilon_{0}(s)>0$ and all $n\in\mathbb{N}$. It follows that
$n\left(11s+\epsilon_{0}\right)+O(\log C_{1}C_{2})+O_{\phi,s,c_{1},c_{2}}(1)\leq 8H_{n}(X+Y)+3H_{n}(\phi(X,Y)).$
As a consequence,
$\max\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\}$ $\geq n\left(s+\frac{\epsilon_{0}}{11}\right)+O(\log C_{1}C_{2})+O_{\phi,s,c_{1},c_{2}}(1).$
Hence we complete the proof of the lemma with constant $\epsilon_{1}=\epsilon_{0}/11$.
Next, if $\alpha$ and $\beta$ have different signs, the result will follow with a slightly worse constant. Without loss of generality, we assume that $\alpha>0>\beta$. Then
$|W_{1}-W_{2}|^{2}=\alpha(X_{1}-X_{2})^{2}-\beta(Y_{1}-Y_{2})^{2}.$
Note that
$\alpha(X_{1}-X_{2})^{2}-\beta(Y_{1}-Y_{2})^{2}=2(\alpha(X_{1}^{2}+X_{2}^{2})-\beta(Y_{1}^{2}+Y_{2}^{2}))-\alpha(X_{1}+X_{2})^{2}+\beta(Y_{1}+Y_{2})^{2}.$
In this case we use Lemma 2.18 and obtain
$H_{n}(|W-W^{\prime}|^{2})+H_{n}(X_{1},X_{2},Y_{1},Y_{2})$
$\leq H_{n}(X_{1}+X_{2},Y_{1}+Y_{2},p(X_{1})+p(X_{2})-q(Y_{1})-q(Y_{2}))+H_{n}(X_{1}-X_{2},Y_{1}-Y_{2})+O_{\phi,c_{1}}(1)$
$\leq H_{n}(X_{1}+X_{2})+H_{n}(Y_{1}+Y_{2})+H_{n}(p(X_{1})+p(X_{2})-q(Y_{1})-q(Y_{2}))$
$\qquad+H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})+O_{\phi,c_{1}}(1).$
Compared with (34), the second entropy can be bounded by
$H_{n}(p(X_{1})+p(X_{2})-q(Y_{1})-q(Y_{2}))\leq 3H_{n}(p(X_{1})-q(Y_{1}))-H_{n}(X_{1})-H_{n}(Y_{1})+O(1).$
The term $H_{n}(p(X_{1})-q(Y_{1}))$ can be bounded by using (12), which gives
$H_{n}(p(X_{1})-q(Y_{1}))$ $\leq 3H_{n}(p(X_{1})+q(Y_{1}))-H_{n}(p(X_{1}))-H_{n}(q(Y_{1}))+O(1)$
$=3H_{n}(p(X)+q(Y))-H_{n}(X)-H_{n}(q(Y))+O_{\phi,c_{1},c_{2}}(1).$
It leads to
$H_{n}\big{(}|W_{1}-W_{2}|^{2}\big{)}+8H_{n}(X)+8H_{n}(Y)\leq 8H_{n}(X+Y)+9H_{n}(\phi(X,Y))+O_{\phi,c_{1},c_{2}}(1),$
and
$n\left(17s+\epsilon_{0}\right)+O(\log C_{1}C_{2})+O_{\phi,s,c_{1},c_{2}}(1)\leq 8H_{n}(X+Y)+9H_{n}(\phi(X,Y)).$
Finally, the lemma will follow with constant $\epsilon_{1}=\epsilon_{0}/17$. ∎
Remark 4.3.
We emphasize that the parameters $C_{1},C_{2}$ in Lemma 4.1 can be unbounded as $n$ tends to infinity. And, when we apply the above lemma in the next section, the term $O(\log C_{1}C_{2})$ will lead to some entropy-terms. So it cannot be replaced by $O_{C_{1},C_{2}}(1)$.
5 Proof of Theorem 1.1 with $s>1/2$: general forms
In this section, we first consider some special quadratic forms that are more complicated than the diagonal ones. In the following lemma, when we say a constant depends on the quadratic form $\phi$, this means that it depends on the coefficients $\alpha$ and $\beta$ occurred in $\phi$.
Lemma 5.1.
Let $\phi(x,y)=\alpha(x-py)^{2}+\beta(x+y)^{2}$, where $\alpha,\beta\in{\mathbb{R}}^{\ast}$ and $p\in\{0,1\}$. Let $X$ and $Y$ be independent $(s,C)$-Frostman random variables whose ranges are contained in $[-c_{1},c_{1}]$, where $11/2<s<$ 1, $C\geq 1$ and $c_{1}>0$. Then there exists a positive constant $\epsilon_{2}=\epsilon_{2}(s)$ and an $N=N(\phi,s,C,c_{1})$ such that
$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{2})$
for all $n>N$.
Proof.
For simplicity, we write
$\Psi:=\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}-ns,$ (36)
and our aim is to prove that $\Psi\geq\epsilon_{2}n$ for some positive $\epsilon_{2}$.
Since $X$, $Y$ are independent and both $(s,C)$-Frostman, we deduce by Lemma 2.26 Corollary 2.22 and (36) that
$\begin{split}&\min\{H_{n}(X),\,H_{n}(Y)\}\geq ns+O_{C}(1),\\ &\max\{H_{n}(X),\,H_{n}(Y)\}\leq H_{n}(X+Y)+O(1)\leq ns+\Psi+O(1).\end{split}$ (37)
Let us define the random variable $(X_{0},Y_{0})$ as
$(X_{0},Y_{0})=(X-pY,X+Y)$
with $p\in\{0,1\}$. The ranges of $X_{0}$ and $Y_{0}$ are both contained in $[-2c_{1},2c_{1}]$. And $(X_{0},Y_{0})$ is an $O(1)$-bi-Lipschitz function of $(X,Y)$, which, because $X$ and $Y$ are independent, leads to the estimate
$H_{n}(X_{0},Y_{0})=H_{n}(X,Y)+O(1)=H_{n}(X)+H_{n}(Y)+O(1)\leq 2ns+2\Psi+O(1).$ (38)
By Lemma 2.25, one deduces that $X_{0}$ and $Y_{0}$ are $(s,C_{1}^{\prime})$-Frostman and $(s,C_{2}^{\prime})$-Frostman, respectively, with $C_{1}^{\prime},C_{2}^{\prime}=O_{C}(1)$. Similarly, one also has that $X_{0}+Y_{0}=(2X+(1-p)Y)$ is $(s,O_{C}(1))$-Frostman, and $(X_{0},Y_{0})$ is $(2s,O_{C}(1))$-Frostman. We then show below that
$H_{n}(X_{0})\leq ns+O(\Psi)+O_{C}(1),\quad H_{n}(X_{0}+Y_{0})\leq ns+O(\Psi)+O_{C}(1).$ (39)
Indeed, when $p=0$, one sees by (36), (37) and Lemma 2.23 that
$H_{n}(X_{0})=H_{n}(X)\leq ns+\Psi+O(1),$
and
$H_{n}(X_{0}+Y_{0})=H_{n}(2X+Y)$
$\qquad\leq H_{n}(X+Y)+O(H_{n}(X+Y)-H_{n}(X))+O(H(X+Y)-H(Y))+O(1)$
$\qquad\leq ns+O(\Psi)+O_{C}(1).$
When $p=1$, it follows from (12), (36) and (37) that
$H_{n}(X_{0})=H_{n}(X-Y)\leq 3H_{n}(X+Y)-H(X)-H(Y)+O(1)\leq ns+3\Psi+O_{C}(1)$
and
$H_{n}(X_{0}+Y_{0})=H_{n}(2X)=H_{n}(X)+O(1)\leq ns+\Psi+O(1).$
So (39) follows.
[…]
$P(E)=P(|X_{0}|>\eta,|Y_{0}|>\eta)\geq 1-P(|X_{0}|<\eta)-P(|Y_{0}|<\eta)\geq 1-C_{1}^{\prime}\eta^{s}-C_{2}^{\prime}\eta^{s}>\frac{1}{2},$
Lemma B.2 shows that $(X_{1},Y_{1})$ is $(2s,O_{C}(1))$-Frostman. Similarly, all of $X_{1}$, $Y_{1}$ and $X_{1}+Y_{1}$ are $(s,O_{C}(1))$-Frostman. We can apply Lemmas B.2 and 2.26, then combine (36), (38), (39) to get that
$\begin{split}&2ns+O_{C}(1)\leq H_{n}(X_{1},Y_{1})\leq\frac{H_{n}(X_{0},Y_{0})}{P(E)}\leq(1+O_{C}(\eta^{s}))(2ns+2\Psi+O(1)),\\ &ns+O_{C}(1)\leq H_{n}(X_{1})\leq\frac{H_{n}(X_{0})}{P(E)}\leq(1+O_{C}(\eta^{s}))(ns+O(\Psi)+O_{C}(1)),\\ &ns+O_{C}(1)\leq H_{n}(Y_{1})\leq\frac{H_{n}(X+Y)}{P(E)}\leq(1+O_{C}(\eta^{s}))(ns+\Psi),\\ &ns+O_{C}(1)\leq H_{n}(X_{1}+Y_{1})\leq\frac{H_{n}(X_{0}+Y_{0})}{P(E)}\leq(1+O_{C}(\eta^{s}))(ns+O(\Psi)+O_{C}(1)),\end{split}$ (40)
and
$H_{n}(\alpha X_{1}^{2}+\beta Y_{1}^{2})\leq\frac{H_{n}(\alpha X_{0}^{2}+\beta Y_{0}^{2})}{P(E)}=(1+O_{C}(\eta^{s}))H_{n}(\phi(X,Y))\leq(1+O_{C}(\eta^{s}))(ns+\Psi).$ (41)
It follows from (40) and (41) that
$\begin{split}&H_{n}(X_{1})+H_{n}(Y_{1})-H_{n}(X_{1},Y_{1})=O(\Psi)+O_{C}(\sqrt{\eta}n)+O_{C}(1),\\ &H_{n}(X_{1}+Y_{1})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}=O(\Psi)+O_{C}(\sqrt{\eta}n)+O_{C}(1),\\ &H_{n}(\alpha X_{1}^{2}+\beta Y_{1}^{2})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}=O(\Psi)+O_{C}(\sqrt{\eta}n)+O_{C}(1).\end{split}$
Now we write $\log K:=O(\Psi)+O_{C}(\sqrt{\eta}n)+O_{C}(1)$. Applying Theorem 3.1 with the random variable $(X_{1},Y_{1})$, we obtain the random variable $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$, where $X^{\prime\prime}$ and $Y^{\prime\prime}$ are conditionally independent given $D_{n}(X^{\prime},Y^{\prime})$, and
$H_{n}(X^{\prime\prime}|X^{\prime},Y^{\prime})\geq H_{n}(X_{1})-\log K,\quad H_{n}(Y^{\prime\prime}|X^{\prime},Y^{\prime})\geq H_{n}(Y_{1})-\log K.\$ (42)
First, let us take the functions $f$ and $g$ in Theorem 3.1 to be the identity functions. Then
$H_{n}(X^{\prime\prime}+Y^{\prime\prime}|X^{\prime},Y^{\prime})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}\leq 7\log K+O(1).$ (43)
Second, we take the functions $f,g$ in Theorem 3.1 to be $f(x)=\alpha x^{2}$ and $g(y)=\beta y^{2}$. By the construction $(X_{1},Y_{1})\equiv(X_{0},Y_{0}||X_{0}|>\eta,|Y_{0}|>\eta)$, the ranges of $X^{\prime\prime}$ and $Y^{\prime\prime}$ are contained in $[-2c_{1},-\eta]\cup[\eta,2c_{1}]$ with probability $1$. Now, by Example 2.13, the random variables $f(X^{\prime\prime})$ and $g(Y^{\prime\prime})$ are weakly $O_{\phi}(\max\{\eta^{-1},c_{1}\})$-bi-Lipschitz functions of $X^{\prime\prime}$ and $Y^{\prime\prime}$, respectively. It follows that
$H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|X^{\prime},Y^{\prime})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}\leq 7\log K+O_{\phi,c_{1},\eta}(1).$ (44)
Next, conditionally given that $D_{n}(X^{\prime},Y^{\prime})=(x,y)$ with $x,y$ being dyadic rationals satisfying $P\big{(}D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}>0$, we will prove the following results on induced conditional probability measures:
$\begin{split}\big{(}X^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}\text{ is $\left(s,O\left(C^{2}\cdot 2^{-ns}/P_{E}(D_{n}(Y^{\prime})=y)\right)\right)$-Frostman},\\ \big{(}Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}\text{ is $\left(s,O\left(C^{2}\cdot 2^{-ns}/P_{E}(D_{n}(X^{\prime})=x)\right)\right)$-Frostman},\end{split}$ (45)
where the implied constants do not depend on $(x,y)$. We present the proof for $X^{\prime\prime}$, as the one for $Y^{\prime\prime}$ can be done similarly.
For any ball $B(z,r)$, using conditions (i)–(iv) in Theorem 3.1, we have
$P_{E}(X^{\prime\prime}\in B(z,r)|D_{n}(X^{\prime},Y^{\prime})=(x,y))=P_{E}(X^{\prime\prime}\in B(z,r)|D_{n}(Y^{\prime})=y)$
$\qquad=P_{E}(X^{\prime}\in B(z,r)|D_{n}(Y^{\prime})=y)=\frac{P_{E}(X^{\prime}\in B(z,r),D_{n}(Y^{\prime})=y)}{P_{E}(D_{n}(Y^{\prime})=y)}$
$\qquad=\frac{P_{E}(X_{1}\in B(z,r),D_{n}(Y_{1})=y)}{P_{E}(D_{n}(Y^{\prime})=y)}\leq\frac{2P(X-pY\in B(z,r),D_{n}(X+Y)=y)}{P_{E}(D_{n}(Y^{\prime})=y)}.$
Denote by $\mu_{X}$, $\mu_{Y}$ the measures associated to the independent random variables $X$ and $Y$, respectively. Then $\mu_{X}$ and $\mu_{Y}$ are $(s,C)$-Frostman. Furthermore, define $I_{n}(y):=[y,y+2^{-n})$. We have that
[…]
The value inside the integral can be positive only when $I_{n}(y-\tilde{y})\cap B(z+p\tilde{y},r)\not=\emptyset$. Or equivalently, $I_{n}(y)\cap B(z+(p+1)\tilde{y},r)\not=\emptyset$. That is to say,
$z+(p+1)\tilde{y}-r\leq y+2^{-n},\quad z+(p+1)\tilde{y}+r\geq y.$
We can locate such $\tilde{y}$ in an interval $J$ of length at most $2r+2^{n}$. We now estimate the above expression in different ways depending on how $r$ relates to $2^{-n}$. First, for $r\geq 2^{-n}$, and recalling that $\mu_{X}$ and $\mu_{Y}$ are $(s,C)$-Frostman, we have
[…]
Next, for $r<2^{-n}$, we similarly find
[…]
This completes the proof for the claims given in (45)
As the expressions to follow get rather lengthy, we now introduce some shorthand. In what follows, $H_{n}(~{}\cdot~{}|x,y)$ is an abbreviation of $H(D_{n}(\cdot)~{}|D_{n}(X^{\prime},Y^{\prime})=(x,y)).$ By Lemma B.2, there exist random variables, depending on $(x,y)$, that have the same distributions with $(X^{\prime\prime},Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y))$, respectively. These random variables take the same value as $(X^{\prime\prime},Y^{\prime\prime})$ does. By construction, their ranges are contained in $\big{(}[-2c_{1},-\eta]\cup[\eta,2c_{1}]\big{)}^{2}$. Also note that $(X^{\prime\prime}|D_{n}(X_{1},Y_{1})=(x,y))$ and $(Y^{\prime\prime}|D_{n}(X_{1},Y_{1})=(x,y))$ are independent. Recall that the constant $\epsilon_{1}=\epsilon_{1}(s)$ in Lemma 4.1 is independent of the concrete random variables. Finally, recall (45). Now we can apply Lemma 4.1 with these random variables, with $C_{1}=O_{C}\big{(}2^{-ns}/P_{E}(D_{n}(Y^{\prime})=y)\big{)}$, $C_{2}=O_{C}\big{(}2^{-ns}/P_{E}(D_{n}(X^{\prime})=x)\big{)},$ and $c_{2}=\eta$, to deduce that
$\max\left\{H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y),H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|x,y)\right\}\geq n(s+\epsilon_{1})$
$\qquad+O\big{(}-2ns-\log P_{E}(D_{n}(X^{\prime})=x)-\log P_{E}(D_{n}(Y^{\prime})=y)\big{)}+O_{\phi,s,C,c_{1},\eta}(1).$ (46)
By conditional independence, and combining Corollary 2.22, we also have that
$H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y)\geq\max\{H_{n}(X^{\prime\prime}|x,y),H_{n}(Y^{\prime\prime}|x,y)\}+O(1).$
Using a similar argument, Lemma B.2, and Example 2.17, one obtains
$H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|x,y)$ $\geq\max\{H_{n}(\alpha X^{\prime\prime 2}|x,y),H_{n}(\beta Y^{\prime\prime 2}|x,y)\}+O(1)$
$=\max\{H_{n}(X^{\prime\prime}|x,y),H_{n}(Y^{\prime\prime}|x,y)\}+O_{\phi,c_{1},\eta}(1).$
Then
$\begin{split}&\min\{H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y),H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|x,y)\}\geq\frac{H_{n}(X^{\prime\prime}|x,y)+H_{n}(Y^{\prime\prime}|x,y)}{2}+O_{\phi,c_{1},\eta}(1).\end{split}$ (47)
Therefore given the condition that $D_{n}(X_{1},Y_{1})=(x,y)$, it follows from (46) and (47) that
$H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y)+H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|x,y)-\frac{H_{n}(X^{\prime\prime}|x,y)+H_{n}(Y^{\prime\prime}|x,y)}{2}$
$\quad\geq n(s+\epsilon_{1})+O\big{(}-2ns-\log P_{E}(D_{n}(X^{\prime})=x)-\log P_{E}(D_{n}(Y^{\prime})=y)\big{)}+O_{\phi,s,C,c_{1},\eta}(1).$
Now we sum all of these expressions respective to each $(x,y)$ with the weight $P_{E}(D_{n}(X^{\prime},Y^{\prime})=(x,y))$. Note that
$\sum\limits_{x,y}P_{E}(D_{n}(X^{\prime},Y^{\prime})=(x,y))H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y)$
$\qquad=\sum\limits_{x,y}P_{E}(D_{n}(X^{\prime\prime},Y^{\prime\prime})=(x,y))H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y)=H_{n}(X^{\prime\prime}+Y^{\prime\prime}|X^{\prime},Y^{\prime}),\text{ and}$
$-\sum\limits_{x,y}P_{E}(D_{n}(X^{\prime},Y^{\prime})=(x,y))\log P_{E}(D_{n}(X^{\prime})=x)$
$\qquad=-\sum\limits_{x}P_{E}(D_{n}(X^{\prime})=x)\log P_{E}(D_{n}(X^{\prime})=x)=H_{n}(X^{\prime}).$
The other terms behave similarly. Therefore,
$H_{n}(X^{\prime\prime}+Y^{\prime\prime}|X^{\prime},Y^{\prime})+H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|X^{\prime},Y^{\prime})-\frac{H_{n}(X^{\prime\prime}|X^{\prime},Y^{\prime})+H_{n}(Y^{\prime\prime}|X^{\prime},Y^{\prime})}{2}$
$\qquad\qquad\geq n(s+\epsilon_{1})+O(-2ns+H_{n}(X^{\prime})+H_{n}(Y^{\prime}))+O_{\phi,s,C,c_{1},\eta}(1).$ (48)
Substituting from (42)–(44), and further combining with $(X^{\prime},Y^{\prime})\equiv(X_{1},Y_{1})$ we obtain that
$\begin{split}&\left(\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}+7\log K\right)+\left(\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}+7\log K\right)\\ &\qquad-\frac{(H_{n}(X_{1})-\log K)+(H_{n}(Y_{1})-\log K)}{2}-ns+O(H_{n}(X_{1})+H_{n}(Y_{1})-2ns)\\ &\qquad\geq n\epsilon_{1}+O_{\phi,s,C,c_{1},\eta}(1).\end{split}$ (49)
Recalling (40) and the notation $\log K=O(\Psi)+O_{C}(\sqrt{\eta}n)+O_{C}(1)$, the left-hand side of (49) is
$\leq O\left(\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}-ns\right)+15\log K=O(\Psi)+O_{C}(\sqrt{\eta}n)+O_{C}(1).$
So there is some absolute constant $\tilde{C}>0$ such that
$\tilde{C}\Psi+O_{C}(\sqrt{\eta}n)\geq n\epsilon_{1}+O_{\phi,s,C,c_{1},\eta}(1).$
Equivalently,
$\Psi\gg n\epsilon_{1}+O_{C}(\sqrt{\eta}n)+O_{\phi,s,C,c_{1},\eta}(1).$
[…]
Now we are ready to finish the proof of Theorem 1.1.
Proof of Theorem 1.1.
For $\phi(x,y)=a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$, the non-degeneracy gives that $a_{2}^{2}\neq 4a_{1}a_{3}$.
When $a_{3}\neq 0$, we have
$\phi(x,y)=\left(a_{1}-\frac{a_{2}^{2}}{4a_{3}}\right)x^{2}+a_{3}\left(\frac{a_{2}}{2a_{3}}x+y\right)^{2}.$
Take $\tilde{X}=\frac{a_{2}}{2a_{3}}X$ and $\tilde{Y}=Y$. Then $\phi(X,Y)=\tilde{\phi}(\tilde{X},\tilde{Y})$, where
$\tilde{\phi}(\tilde{x},\tilde{y})=\left(\frac{4a_{1}a_{3}^{2}}{a_{2}^{2}}-a_{3}\right)\tilde{x}^{2}+a_{3}(\tilde{x}+\tilde{y})^{2}.$
When $a_{1}\neq 0$, similar arguments lead to $\phi(X,Y)=\tilde{\phi}(\tilde{X},\tilde{Y})$, with $\tilde{X}=\frac{a_{2}}{2a_{1}}Y$, $\tilde{Y}=X$, and
$\tilde{\phi}(\tilde{x},\tilde{y})=\left(\frac{4a_{1}^{2}a_{3}}{a_{2}^{2}}-a_{1}\right)\tilde{x}^{2}+a_{1}(\tilde{x}+\tilde{y})^{2}.$
When $a_{1}=a_{3}=0$, one has $a_{2}\neq 0$ and $\phi(X,Y)=\tilde{\phi}(\tilde{X},\tilde{Y})$, where $\tilde{X}=X$, $\tilde{Y}=Y$, and
$\tilde{\phi}(\tilde{x},\tilde{y})=-\frac{a_{2}}{4}(\tilde{x}-\tilde{y})^{2}+\frac{a_{2}}{4}(\tilde{x}+\tilde{y})^{2}.$
In all cases, the random variables $\tilde{X}$ and $\tilde{Y}$ are independent, and both are $(s,O_{\phi,C}(1))$-Frostman. The ranges of $\tilde{X}$ and $\tilde{Y}$ are both contained in $[-\widetilde{c_{1}},\widetilde{c_{1}}]$, with $\widetilde{c_{1}}=O_{\phi,c_{1}}(1)$.
Applying Lemma 5.1 with the quadratic form $\tilde{\phi}(\tilde{x},\tilde{y})$ and the random variables $\tilde{X}$ and $\tilde{Y}$, we obtain that
$\max\{H_{n}(\xi X+\zeta Y),\,H_{n}(\phi(X,Y))\}\geq n(s+\varepsilon_{2})$
for some non-zero rational numbers $\xi=\xi(\phi)$ and $\zeta=\zeta(\phi)$. Let $\xi=\frac{k}{m}$ and $\zeta=\frac{l}{m}$, with $m$ being the greatest common divisor of $\xi$ and $\zeta$. By Lemmas 2.23 and 2.26, there are some $d_{1}=d_{1}(\xi,\zeta)$ and $d_{2}=d_{2}(\xi,\zeta)$ such that
$H_{n}(\xi X+\zeta Y)$ $\leq(d_{1}+d_{2}+1)H_{n}(X+Y)-d_{1}H_{n}(X)-d_{2}H_{n}(Y)+O_{\xi,\zeta}(1)$
$\leq(d_{1}+d_{2}+1)H_{n}(X+Y)-(d_{1}+d_{2})(ns-\log C)+O_{\phi}(1).$
That is to say,
$H_{n}(X+Y)\geq\frac{(d_{1}+d_{2})ns+H_{n}(\xi X+\zeta Y)}{d_{1}+d_{2}+1}+O_{\phi,C}(1).$
Now the theorem follows by taking $\epsilon=\frac{\epsilon_{2}}{2(d_{1}+d_{2}+1)}$ when $n$ is sufficiently large.
Remark 5.2.
Indeed, our proof of Theorem 1.1 works for any real quadratic form $\phi$ that is congruent over $\mathbb{Q}$ to a real diagonal form, i.e., $\phi(x,y)=\alpha(\xi_{11}x+\xi_{12}y)^{2}+\beta(\xi_{21}x+\xi_{22}y)^{2}$, where $\alpha,\beta\in{\mathbb{R}}^{\ast}$ and $\Xi=\begin{bmatrix}\xi_{11}&\xi_{12}\\ \xi_{21}&\xi_{22}\end{bmatrix}\in GL_{2}(\mathbb{Q})$, together with the constant $\epsilon$ depending both on $s$ and $\Xi$.
6 Proof of Theorem 1.1 with $s\leq 1/2$
The situation is more subtle for this case. It is hard to expect Theorem 2.32 holds true when the dimension gets smaller. We will use the upcoming theorem as a substitution. The core of the proof will follow arguments from [20] by Shmerkin and Wang. However, there are enough differences that we cannot merely apply a single theorem of theirs as a black box, so to address this case, we will introduce various notions and results from their paper, and explicitly mention where our arguments diverge. Recall the definition of $\varphi$ from (6).
Theorem 6.1.
Let $X$, $Y$ be independent random variables with ranges contained in $[-c,c]$, with $c>0$. Suppose that $X$ and $Y$ are $(s,C_{1})$-Frostman and $(s,C_{2})$-Frostman, respectively, where $20<s\leq 1/$ 2 and $C_{1},C_{2}\geq 1$. Let $(X^{\prime},Y^{\prime})$ be i.i.d. with $(X,Y)$. Then for any $t$ with $s<t<\frac{s+\varphi(2s)}{2}$, we have
$H_{n}\big{(}|(X,Y)-(X^{\prime},Y^{\prime})|\big{)}\geq nt+O_{s}(\log n)+O_{s,t}(\log C_{1}C_{2})+O_{s,c}(1)$
for all $n\in\mathbb{N}$.
To compare Theorems 2.32 and 6.1, one may take $W=(X,Y)$ in Theorem 2.32. The random variables $X$ and $Y$ are required to be independent in Theorem 6.1, but may be dependent in Theorem 2.32.
The rest of the proof of Theorem 1.1 for $s\leq 1/2$ will follow the same line as in the case $s>1/2$. So we omit the details. To prove Theorem 6.1, we need the following variant for measures. For two measures $\mu^{\prime},\mu^{\prime\prime}$ whose supports are bounded regions on ${\mathbb{R}}^{2}$, we denote by $\Delta(\mu^{\prime},\mu^{\prime\prime})$ the distance measure between $\mu^{\prime}$ and $\mu^{\prime\prime}$, i.e.,
[…]
for any Borel subset $B$ of ${\mathbb{R}}$. By closely following arguments from Shmerkin and Wang, in [20], we obtain the following entropy estimate.
Theorem 6.2.
Let $\mu_{1}$, $\mu_{2}$, and $\nu$ be $(s,C)$-Frostman probability measures on $[0,1]$, where $20<s\leq 1/$ 2 and $C\geq 1$. Assume that $\text{dist}(\operatorname{spt}(\mu_{1}),\operatorname{spt}(\mu_{2}))\geq C_{d}C^{-1/s}$ with $C_{d}>0$. Then for any $\zeta>0$, we have that
$H_{n}(\Delta(\mu_{1}\times\nu,\mu_{2}\times\nu))\geq n(\varphi(2s)-\zeta)+O_{s,\zeta,C_{d}}(\log C)$
for all sufficiently large $n$.
We emphasize that the term $O_{s,\zeta,C_{d}}(\log C)$ can not be replaced by $O_{s,\zeta,C,C_{d}}(1)$, because in the proof of the main theorem, we are working with infinitely many random variables satisfying (45), where the Frostman constants can be large compared to $n$, and are not uniform in $(x,y)$. So only an error of the form $\log C$ can be turned into entropy-terms as in (48). The main difficulty in this section, which prevents us from applying results in [20] directly, is due to this reason.
Rather than repeating their arguments wholesale with minor modifications, we provide a detailed description of how to prove variants of their results tailored to our purposes, and combine these with a few calculations to prove Theorem 6.2.
6.1 Proof of Theorem 6.1 from Theorem 6.2
We need the following technical lemma.
Lemma 6.3.
Let $10<s<$ 1 and $C\geq 1$. Let $\mu$ be an $(s,C)$-Frostman probability measure with $\operatorname{spt}(\mu)\subset[0,1]$. For any $30<\gamma<1/$ 3, we can find a partition $\{U_{1},U_{2},U_{3}\}$ of $[0,1]$ such that the following hold:
(i) $\text{dist}(U_{1},U_{2})\gg(\gamma/C)^{1/s}$;
(ii) $\mu_{U_{1}}$, $\mu_{U_{2}}$ are $(s,O(C))$-Frostman probability measures;
(iii) $21/2-\gamma\leq\mu(U_{1}),\mu(U_{2})\leq 1/$ 2 and $\mu(U_{3})\leq\gamma$.
Proof.
Let $k=\lfloor\frac{1}{s}\log\frac{C}{\gamma}\rfloor$, which gives $C2^{-ks}\leq\gamma$. Write $q_{i}=\mu\big{(}[0,i2^{-k})\big{)}$ $(i=1,2,\ldots,2^{k})$. In view of
$q_{i+1}-q_{i}=\mu\big{(}[i2^{-k},\,(i+1)2^{-k})\big{)}\leq C2^{-ks}\leq\gamma,$
there exists a unique $i_{0}$ such that
$1/2-\gamma\leq q_{i_{0}}<1/2\leq q_{i_{0}+1}<1/2+\gamma.$
[…]
$1/2-\gamma\leq\mu(U_{1}),\mu(U_{2})\leq 1/2,\quad\mu(U_{3})\leq\gamma.$
By construction, $\text{dist}(U_{1},U_{2})\geq 2^{-k}\gg(\gamma/C)^{1/s}$. And by Lemma B.2, both $\mu_{U_{1}}$ and $\mu_{U_{2}}$ are $(s,O(C))$-Frostman. This completes the proof. ∎
With above results, we are going to prove Theorem 6.1.
Proof of Theorem 6.1.
For simplicity, we denote $Z:=|(X,Y)-(X^{\prime},Y^{\prime})|$, and take $C=C_{1}C_{2}$. The random variables $\frac{X}{2c}+\frac{1}{2}$ and $\frac{Y}{2c}+\frac{1}{2}$ are both $\big{(}s,O((2c)^{s}C)\big{)}$-Frostman, with ranges contained in $[0,1]$. Moreover,
$H_{n}(|(X,Y)-(X^{\prime},Y^{\prime})|)=H_{n}\left(\left|\left(\frac{X}{2c}+\frac{1}{2},\frac{Y}{2c}+\frac{1}{2}\right)-\left(\frac{X^{\prime}}{2c}+\frac{1}{2},\frac{Y^{\prime}}{2c}+\frac{1}{2}\right)\right|\right)+O_{c}(1).$
Without loss of generality, we assume at the beginning of the proof that both $X$ and $Y$ are $(s,C)$-Frostman, with range contained in $[0,1]$. Our aim is to prove that
$H_{n}\big{(}Z\big{)}\geq nt+O_{s}(\log n)+O_{s,t}(\log C).$
Let $\gamma=\gamma(s,t)\in(0,1/3)$ be a parameter to be determined later. We apply Lemma 6.3 on the distribution measure of $X$, denoted by $\mu$. Then there is a partition $\{U_{1},U_{2},U_{3}\}$ of $[0,1]$ such that $\text{dist}(U_{1},U_{2})\gg_{s,t}C^{-1/s}$, both $\mu_{U_{1}}$ and $\mu_{U_{2}}$ are $(s,O(C))$-Frostman, and $21/2-\gamma\leq\mu(U_{1}),\mu(U_{2})\leq 1/$ 2.
Note that $range\big{(}(X,Y,X^{\prime},Y^{\prime})\big{)}\subseteq[0,1]^{4}$. For $i,j=1,2,3$, denote by $E_{ij}$ the event that $(X,Y,X^{\prime},Y^{\prime})\in U_{i}\times[0,1]\times U_{j}\times[0,1]$. Then $\{E_{ij}:\,i,j=1,2,3\}$ forms a partition. By (10), we have
$H_{n}(Z)\geq\sum\limits_{i,j=1}^{3}P(E_{ij})H_{n}\big{(}Z_{E_{ij}}\big{)}=\sum\limits_{i,j=1}^{3}\mu(U_{i})\mu(U_{j})H_{n}\big{(}Z_{E_{ij}}\big{)}.$
Since $X,Y,X^{\prime},Y^{\prime}$ are independent, one can verify that $X_{U_{i}},Y,X_{U_{j}}^{\prime},Y^{\prime}$ are also independent, and $Z_{E_{ij}}\equiv|(X_{U_{i}},Y)-(X_{U_{j}}^{\prime},Y^{\prime})|$.
Now apply Theorem 6.2 on $Z_{E_{12}}$ with $\zeta=\zeta(s,t)>0$ to be determined later. We deduce that
$H_{n}(Z_{E_{21}})=H_{n}(Z_{E_{12}})\geq n(\varphi(2s)-\zeta)+O_{s,t}(\log C)$
for all $n\in\mathbb{N}$. Moreover, by Lemma 2.36, one obtains that
$H_{n}(Z_{E_{11}}),\,H_{n}(Z_{E_{22}})\,\geq ns+O_{s}(\log n)+O_{s}(\log C)$
for all $n\in\mathbb{N}$.
For simplicity, write $a=\mu(U_{1})$ and $b=\mu(U_{2})$, with $21/2-\gamma\leq a,b\leq 1/$ 2. It follows that
$H_{n}(Z)$ $\geq\sum\limits_{i,j=1}^{2}\mu(U_{i})\mu(U_{j})H_{n}\big{(}Z_{E_{ij}}\big{)}$
$\geq 2abn(\varphi(2s)-\zeta)+(a^{2}+b^{2})ns+O_{s}(\log n)+O_{s,t}(\log C)$
$\geq 2(1/2-\gamma)^{2}(s+\varphi(2s)-\zeta)n+O_{s}(\log n)+O_{s,t}(\log C).$
To complete the proof, we consider the inequality
$2(1/2-\gamma)^{2}\left(s+\varphi(2s)-\zeta\right)\geq t.$
Since $s<t<\frac{s+\varphi(2s)}{2}$, then $\frac{t}{s+\varphi(2s)+2t}<\frac{1}{4}$. It is sufficient to take $\zeta=\frac{s+\varphi(2s)}{2}-t$ and select a $\gamma=\gamma(s,t)$ satisfying
$0<\gamma<\min\left\{\frac{1}{2}-\sqrt{\frac{t}{s+\varphi(2s)+2t}}\,,\,\,\,\,\frac{1}{3}\right\}.$
The proof is completed. ∎
6.2 Proof of Theorem 6.2
We begin by recording the results, or slight variants thereof, that we use from Shmerkin and Wang, [20], as well as the requisite material to state them. We then use these to explain how to prove Theorem 6.2. Our proof basically runs parallel to the proof of their Theorem 5.2, but with slightly different hypotheses and a different conclusion.
6.2.1 Material from Shmerkin and Wang
Here we present definitions and results relevant only to this section. Given a set $\Omega,$ we denote by $\mathcal{P}(\Omega)$ the set of probability measures on $\Omega.$ The first lemma we lift from their paper deals with robust measures, which are measures that need not be Frostman, but are close enough for our purposes. Specifically, a measure $\mu\in\mathcal{P}(\mathbb{R}^{d})$ is $(\delta,s,r)$-robust if $\mu(X)>r$ implies that $X$ intersects at least $\delta^{-s}$ cubes from the standard dyadic decomposition of the space into cubes of side length $\delta$. Below, we often use the notation $B_{r}$ to denote an arbitrary ball of radius $r.$
Lemma 6.4.
[Part of Lemma 2.3 [20]] Let $(a_{m})_{m=1}^{\infty}$ be a sequence of positive numbers such that $\sum_{m=1}^{\infty}a_{m}<\infty$. Suppose $\mu\in\mathcal{P}({\mathbb{R}}^{d})$ is $(2^{-m},s,a_{m})$-robust for $m\in[m_{0},m_{1}]$, where $m_{0}<m_{1}\in\mathbb{N}\cup\{+\infty\}$. If $\epsilon:=\sum_{m=m_{0}}^{m_{1}}a_{m}<1$, then there is a set $X$ with $\mu(X)>1-\epsilon$ such that
$\mu_{X}(B_{r})\ll r^{s}\quad\text{for all }r\in[2^{-m_{1}},2^{-m_{0}}].$
[…]
Further, if we have $\mu\in\mathcal{P}(\mathbb{R}^{d})$, we say that the pair of measures $(\mu,\nu)$ has $(t,K,c,\delta)$-thin tubes if
$\mu\left\{x:(x,\nu)\text{ have $(t,K,c,\delta)$-thin tubes }\right\}\geq c.$
We can relax the assumption of probability measures as follows. If $\mu\in\mathcal{M}(\mathbb{R}^{d})$ and $\nu\in\mathcal{M}(\mathbb{R}^{d})$, we say that $(x,\nu)$, $(\mu,\nu)$ have $(t,K,c,\delta)$-thin tubes if the above holds for the associated normalized probability measures, that is, with $\mu/|\mu|$ and $\nu/|\nu|$ in place of $\mu$ and $\nu,$ respectively. Following Shmerkin and Wang, we say that if either of the measures $\mu$ or $\nu$ is trivial, then $(\mu,\nu)$ have $(t,K,c,\delta)$-thin tubes. Moreover, a consequence of Lemma 5.7 in [20] is that this notion of thin tubes is well-defined.
We now refine thin tubes in the following concrete way. Let $\mu,\nu\in\mathcal{P}(\mathbb{R}^{d})$. We say that $(\mu,\nu)$ have $(t,K,c,\delta)$-strong thin tubes if there exists a set $X$ with $\mu(X)\geq c$ and for each $x\in X$ there is a set $Y_{x}$ such that $\nu(Y_{x})\geq c$, the set $\{(x,y):x\in X,y\in Y_{x}\}$ is compact, and
$\nu(Y_{x}\cap T)\leq K\cdot r^{t}\text{ for all }r>\delta\text{ and all open $r$-tubes $T$ through $x$}.$
Again, we extend this definition to $\mu,\nu\in\mathcal{M}(\mathbb{R}^{d})$ by applying the definition to $(\mu/|\mu|,\nu/|\nu|)$. Moreover if either $\mu$ or $\nu$ are trivial, then $(\mu,\nu)$ trivially have $(t,K,c,\delta)$-strong thin tubes.
While it is clear that if $(\mu,\nu)$ have $(t,K,c,\delta)$-strong thing tubes, they have $(t,K,c,\delta)$-thin tubes, the following result from [20] shows that a partial converse also holds.
Lemma 6.5.
[Lemma 5.9 in [20]] Suppose $(\mu,\nu)$ have $(t,K,1-c,\delta)$-thin tubes. Then they have $(t,K^{\prime},c^{\prime},\delta)$-strong thin tubes, where $c^{\prime}>1-3c$, $K^{\prime}=2K$ if $c<1/4$, and $c^{\prime}>(1-\sqrt{c})/2$, $K^{\prime}=(1-\sqrt{c})^{-1}K$ if $c\geq 1/4$.
In [20], they use a result due to Orponen [18], which they call Proposition 5.11. We need something slightly different, so we state and prove explicitly what we will use here. For $v\in\mathbb{R}^{2},$ we will define $P_{v}$ as $P_{v}(x)=x\cdot v,$ so $P_{v}\mu$ is the pullback of $\mu$ through the map $P_{v},$ which means
$P_{v}(\mu)(B(x,r))=\mu\{y:y\cdot v\in B(x,r)\}.$
Proposition 6.6.
[Variant of Proposition 5.11 in [20]] Let $\mu_{1},\mu_{2},\nu$ be $(s,C)$-Frostman probability measures in $[0,1],$then $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has $(s,2C,1,0)$-strong thin tubes.
Proof.
To see this, notice that for any tube $T$ of width $r$, assume that $T$ is defined as $T=\{(x,y)\in\mathbb{R}^{2}:ax+by\in B(z,r)\}$ for some $(a,b)\in S^{1}.$ If $|b|\geq|a|$ then $|b|\geq 1/\sqrt{2}$ so we can bound
[…]
similarly, if $|a|>|b|$ then $|a|>1/\sqrt{2}$ we compute $P_{v}(\mu\times\nu)(B(z,r))$ as
[…]
Therefore, $(\mu_{2}\times\nu)(T)\leq 2Cr^{s}$ for any tube $T$ of width $r$ which immediately gives that $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has $(s,2C,1,0)$-strong thin tubes. ∎
Next we introduce a particularly useful a relaxation of the notion of an $(s,C)$-Frostman measure. In a metric space $X,$ we will call a measure $\mu\in\mathcal{P}(X)$ a $(\delta,s,C)$-measure if $\mu(B_{r})\leq C\cdot r^{s}$ for all $r\in[\delta,1].$ Given $t\in[0,d]$, $s\in[0,k],$ a function $D:[0,d]\to[0,k],$ and $C\geq 1$, we say that a measure $\rho\in\mathcal{P}(\mathbb{G}({\mathbb{R}}^{d},k))$ is $(D;C)$-adapted at scale $\delta$ if whenever $\epsilon>0$ and $\mu$ is a $(\delta,s,\delta^{-\epsilon})$-measure on $[0,1]^{d}$, then for all $s\in[0,d],$ we have
$\rho\{V\in\mathbb{G}(\mathbb{R}^{d},k):P_{V}\mu\text{ is not }(\delta,D(s)-C\epsilon,\delta^{\epsilon})\text{-robust }\}\leq\delta^{\epsilon/C}.$
We now define a family of functions that show up in the next result. Given $s\in\left(0,\varphi(1)\right]$, let
$D_{s}(t)=\left\{\begin{array}[]{lll}\max\{t/2,s+\eta(s,t)\}&\text{ if }&t\in(s,2]\\ t&\text{ if }&t\in[0,s]\end{array}\right..$
We use these new definitions to state the following, which is Lemma 5.16 in Shmerkin and Wang.
Lemma 6.7.
[Lemma 5.16 in [20]] Let $\rho\in\mathcal{P}(S^{1})$ be a $(\delta,s,K)$-measure and fix $\epsilon>0$. If $\delta<K^{-C^{\prime}\epsilon^{-1}}$ for some absolute constant $C^{\prime}$, then $\eta(s,t)$ can be chosen so that $\rho$ is $(D_{s}-\epsilon;1)$-adapted at scale $\delta$.
For a nonnegative number $\sigma,$ a function $f:[0,1]\rightarrow\mathbb{R}$ will be called $\sigma$-superlinear on $[a,b]\subseteq[0,1]$, if for all $x\in[a,b],$ we have that $f(x)\geq f(a)+\sigma\cdot(x-a).$ An interval $[a,b]\subseteq[0,1]$ is $\tau$-allowable if $\tau\leq b-a\leq a.$ Moreover, given $[a,b]\subseteq[0,1]$, we say that a collection of pairwise disjoint intervals $[a_{j},b_{j}]\subseteq[a,b]$ is $\tau$-allowable if each $[a_{j},b_{j}]$ is $\tau$-allowable. Now, given $[a,b]\subseteq[0,1]$, define
$\Sigma_{\tau}(D;f;[a,b],d):=\sup\left\{\sum_{j=1}^{J}(b_{j}-a_{j})D(\sigma_{j})\right\},$
where the supremum is taken over all $\tau$-allowable collections $\{[a_{j},b_{j}]\}$ of sub-intervals of $[a,b]$ and numbers $\sigma_{j}\in[0,d]$ such that $f$ is $\sigma_{j}$-superlinear on $[a_{j},b_{j}]$. Given $u\in[0,d]$, define $\mathcal{L}_{d,u}$ to be the set of all non-decreasing Lipschitz functions $f:[0,1]\to[0,d]$ with Lipschitz constant $\leq d$ that are piecewise linear, with $f(0)=0$ and $f(x)\geq ux$ for all $x\in[0,1]$. Putting all of this together, we define
$\Sigma_{\tau}(D;t,d):=\inf\{\Sigma_{\tau}(D;f;[0,1],d):f\in\mathcal{L}_{d,t}\}.$
Notice that for $t>s,$ we have $\mathcal{L}_{d,t}\subseteq\mathcal{L}_{d,s}$, so $\Sigma_{\tau}(D;t,d)$ is an increasing function of $t$. The next result is Lemma 5.17 from Shmerkin and Wang, which tells us about the continuity of $\Sigma_{\tau}(D;t,d).$
Lemma 6.8.
[Lemma 4.7 from [20]] Fix $\tau$ and $D$ as above. If $D$ is (Lipschitz) continuous, then $s\mapsto\Sigma_{\tau}(D;s,d)$ is (locally Lipschitz) continuous on $(0,d)$.
This result is almost identical to Theorem 4.8 in Shmerkin and Wang, but we modify the hypothesis in (50) below to have a constant, $C_{\mu}$ instead of 1. We then follow the effects of this constant through their proof and arrive at a different conclusion, which is what we will use below. In order to state this, we borrow yet another definition from Shmerkin and Wang. For any $C^{1}$ function $F:U\subseteq[0,1]^{d}\to{\mathbb{R}}^{k}$, and $D$ is the derivative, we define
$V_{F}(x):=\left(\text{ker}(DF(x))\right)^{\perp}.$
Theorem 6.9.
[Variant of Theorem 4.8 in [20]] Let $1\leq k\leq d,$ be natural numbers, $D:[0,d]\to[0,k]$ be a continuous function. Fix $c\in(0,1)$, $C\geq 1$, $0<\delta_{1}\leq 1$. Let $(Y,\nu)$ be a compact measure space, $U\subset\mathbb{R}^{d}$ be a bounded domain, and let $F_{y}:U\rightarrow\mathbb{R}^{k}$ be either the radial projection to $y$ or the distance to $y$.
Let $\mu\in\mathcal{P}(U)$ satisfy
$\mu(B_{r})\leq C_{\mu}r^{s},\quad 0\leq r\leq\delta_{1}.$ (50)
Suppose that there is a set $X^{\prime}$ with $\mu(X^{\prime})>1-c$ such that the following holds for each $x\in X^{\prime}$, there is a set $Y_{x}$ with $\nu(Y_{x})>1-c$, such that the measure
$\rho_{x}=V_{x}\nu_{Y_{x}}\in\mathcal{P}(\mathbb{G}({\mathbb{R}}^{d},k))$
is $(D;C)$-adapted at all scales $\tilde{\delta}\in(0,\delta_{1})$. We also assume that $\{(x,y):x\in X^{\prime},y\in Y_{x}\}$ is compact. Fix $\tau,\epsilon>0$. There are a set $Y_{0}\subset Y$ with $\nu(Y_{0})>\max(1-2c^{1/2},0)$, and for each $y\in Y_{0}$ there is a set $X_{y}\subset X^{\prime}$ with $\mu(X_{y})>\max\{1-2c^{1/2},0\}$, such that
[…]
$\eta\leq\frac{\epsilon\tau}{2},~{}\frac{1}{\tau}\leq\epsilon m,~{}2^{3\epsilon m}\geq\delta_{1}^{-s},\quad 2^{\epsilon m}\geq 1,~{}2^{3\epsilon m}>C_{\mu},\text{ and}$
$\sum_{m\geq m_{0},T|m}J2^{-(\tau\epsilon/C)m}\leq c,~{}J\leq\frac{1}{\tau},~{}T>\frac{2}{\epsilon\tau},~{}\frac{\log(2T+2)}{T}\leq\epsilon.$
Moreover, $\{(x,y):y\in Y_{0},x\in X_{y}\}$ is compact.
This is a technical result taken directly from Shmerkin and Wang.
Lemma 6.10.
Let $\rho\in\mathcal{P}(S^{1})$ be a $(\delta,s,K)$-measure and fix $\epsilon>0$. If $\delta<K^{-C^{\prime}\epsilon^{-1}}$ for some absolute constant $C^{\prime}$, then $\eta(s,t)$ can be chosen so that $\rho$ is $(D_{s}-\epsilon,1)$-adapted at scale $\delta.$
The next result is the combinatorial engine for this part of their argument, which we do not modify.
Proposition 6.11.
[Proposition 5.17 in [20]] Given $\zeta>0$ and $u\in(0,1],$ there are $\xi>0$ and $\tau>0,$ both depending on $u$ and $\zeta,$ such that if $s\in(0,\varphi(u)-\zeta]$ then $\Sigma_{\tau}(D_{s};u,2)\geq s+\xi.$
This lemma will be applied repeatedly below. It is a variant of Lemma 5.18 from Shmerkin and Wang, but there are enough differences that we include a detailed proof.
Lemma 6.12.
[Variant of Lemma 5.18 in [20]] Fix $u\in(0,1]$ and $u_{0}\in(0,u/3)$, $\zeta>0,C>1$ then there is an $\eta$ depending on $\zeta$ and $u$ such that if $s\in[u_{0},\varphi(u)-\zeta]$ then we have the following: If $\mu_{1},\mu_{2}$ are $(u,C)$-Frostman measures in $[0,1]^{2}$ with $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\geq C_{\mu}^{-1}$ and $(\mu_{1},\mu_{2})$ has $(s,K,1-c)$-strong thin tube for some number $K>0,$ then $(\mu_{1},\mu_{2})$ has $(s+\eta,C^{\prime}(KC_{\mu})^{C^{\prime}},1-9c^{1/2},0)$-strong thin tubes for some constant $C^{\prime}$ only depending on $u$ and $\zeta$.
Proof.
First, we will prepare the ingredients needed to apply Theorem 6.9 for the family of functions $\pi:\mathbb{R}^{2}\times\mathbb{R}^{2}\to S^{1}$ defined as $\pi_{x}(y)=\frac{x-y}{|x-y|}$.
By definition, for a fixed $x$, we have that $(\ker D(\pi_{x})(y))^{\perp}=\pi_{x}(y)^{\perp}$ which implies $V_{x}(y)=\pi_{x}(y)^{\perp}$. Using the assumption that $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\gtrsim C_{\mu}^{-1}$ we obtain that $\pi_{x}(B(z,R))\cap\operatorname{spt}\mu_{2}$ will lie inside a tube of length at most $O(C_{\mu}R)$ (see the following figure).
Figure 1: We can see that the tube containing $x$ and $B(z,R)$ has width $\leq C_{\mu}{\rm diam}(\operatorname{spt}\mu_{2})$.
So by the strong thin tubes assumption, we obtain that there is a set $X$ with $\mu_{1}(X)>1-c$ such that for each $x\in X$, there is a set $Y_{x}$ that has $\mu_{2}(Y_{x})>1-c,$ with
$\mu_{2}((V_{x})_{Y_{x}}(B(z,r)))\leq K(C_{\mu}r)^{s}\leq KC_{\mu}r^{s},$
and furthermore, $Y=\{(x,y):x\in X,y\in Y_{x}\}$ is compact.
For some constant $C_{0},$ only depending on the size of the support of $\mu_{1}$ and $\mu_{2}$, we have that by Proposition 6.10 and the method of constructing $D_{s},$ the measure $\mu\circ V_{x}$ is $(t\to D_{s}(t)-\eta)$-adapted at any scales in the range $[0,(K/C_{0})^{-C^{\prime}\eta^{-1}}]$ for any $\eta$ that is small enough. The fact that $\Sigma_{\tau}(D_{s};t,d)$ is continuous as a function in $t$ follows by Lemma 6.8, so by applying Proposition 6.11, we obtain that if $\eta$ is small enough in terms of $u$ and $\zeta,$ then
$\Sigma_{\tau}(D_{s}-\eta;u-\eta,d)=\Sigma_{\tau}(D_{s};u-\eta,d)-\eta\geq s+\xi/2-\eta>s+2\eta.$
This is enough to apply Theorem 6.9 to $Y,\nu,X,$ and the associated projections, $\pi,$ with $\delta_{1}=(KC\mu)^{-C^{\prime}\eta^{-1}}$, provided that $\eta$ is small enough in terms of $u$ and $\zeta.$ This will provide us a $c^{\prime}<2c^{1/2}$, a set $Y^{\prime},$ and $X=\{(y,X_{y}):y\in Y,x\in X_{y}\subset\mathbb{R}^{2}\}$ such that $\mu_{2}(Y^{\prime})\geq 1-c^{\prime}$, and $\mu_{1}(X_{y})>1-c^{\prime}$ for each $y\in Y^{\prime}$, the set $X$ is compact and most importantly:
$\rho_{y}=(\mu_{1}\circ\pi_{y}^{-1})_{X_{y}}\text{ is }(\delta^{\prime},s+\eta,\delta^{\prime\epsilon})\text{-robust for all }\delta^{\prime}<\delta_{0},$
where $\epsilon>0$ depends on $\eta$ and $\tau$ (which, in turn, depend on $u$ and $\zeta$), and $\delta_{0}=C_{1}^{-1}K^{-C_{1}},$ where $C_{1}$ is a constant also depending on $u$ and $\zeta.$
Next, we apply Lemma 6.4 to $\rho_{y}$ to obtain that for each $y$ there is a set $X^{\prime}_{y}\subseteq S_{1}$ for which $(\rho_{y})_{X^{\prime}_{y}}(B(x,r))<r^{s+\eta}$ for every $x<\delta_{0}$ and
$\rho(X^{\prime}_{y})>1-\sum_{m=-\log\delta_{0}}^{\infty}2^{-m\epsilon}\geq 1-\frac{2\delta_{0}}{1-2^{-\epsilon}}\geq 1-c^{1/2}.$
Note that although this does depend on $c$, we choose $C_{1}$ large enough that it will not matter. This implies that $(\rho_{y})_{X_{y}^{\prime}}$ is $(s+\eta,\delta_{1}^{-1})$-Frostman, so the measure $(\mu_{1}\circ P_{v})_{X^{\prime}_{y}}$ is $(s+\eta,O(\delta_{1}^{-1}))$-Frostman for any vector $v\in S^{1}$ because $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\gtrsim C_{\mu}^{-1}$, so the intersection of any tube of width $r$ going through $x\in\operatorname{spt}\mu_{1}$ can be covered by $C_{\mu}$ disjoint arcs of length $r$ of the unit circle centered at $x$ (see the following figure).
By using $\phi_{y}=1_{X^{\prime}_{y}}$ for each $y$, we obtain that $(\mu_{1},\mu_{2})$ has $(s+\eta,10C_{\mu}^{2}\delta^{-1},1-3c^{1/2},0)$-thin tubes. Finally, applying Lemma 6.5 gives us the desired result. ∎
6.2.2 Modifications to their argument
Theorem 6.2 is proved by following the same argument that Shmerkin and Wang prove their Theorem 5.2 in [20], except that their conclusion is about Hausdorff dimension, whereas ours is about entropy. Their proof essentially follows by applying their Proposition 5.15, then their Lemma 4.12. Our proof will begin with a direct argument in place of their Proposition 5.15, then we will use the same ideas from their proof of their Lemma 4.12 up to a statement about the robustness of some measure. From there, Shmerkin and Wang prove statements about Hausdorff dimension, but here we will diverge and write out the details on how to obtain our desired conclusion about entropy.
Proof.
We verify that the hypotheses in the statement allow us to argue as Shmerkin and Wang did in the proof of their Proposition 5.15. They did this by starting with their Proposition 5.11, then iterating their Lemma 5.18. We will start by following the same general approach. In particular, we begin by applying Proposition 6.6 to the measures $\mu_{1}\times\nu$ and $\mu_{2}\times\nu,$ to see that they have $(s,2C,1,0)$-strong thin tubes.
Using Lemma 6.12 with $C_{\mu}=C_{d}C^{\frac{1}{s}},$ we obtain $\eta$ and $C^{\prime}$ only depending on $s$ and $\zeta$ such that if $t<(\varphi(2s)-\zeta/M)$, for some constant $M$ only depending on $s$ and $\zeta,$ which we choose later, and for some $K$ also chosen later, if $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has $(t,K,1-c,0)$-strong thin tubes then it also has $(t+\eta,C^{\prime}\min\{c,K\}^{C^{\prime}},1-9c^{1/2},0)$-strong thin tubes.
Since $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has $(s,2C,1,0)$-strong thin tubes, then it also has $(s,3C,1-c_{0},0)$-strong thin tubes, where
$c_{0}=\left(\frac{\zeta^{2}}{3MC_{\mu}^{2}}\right)^{2^{(\varphi(2s)-s-\zeta/M)/\eta}}.$
Applying Lemma 6.12 $\left\lceil\eta^{-1}(\varphi(2s)-\zeta/M-s)\right\rceil$ times we obtain that $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has $(\varphi(2s)-\zeta/2,(C^{\prime}\min(c_{0}^{-1},C))^{C^{\prime O(1/\eta)}},1-(\zeta/M)^{2},0)$-strong thin tubes, where $K=(C^{\prime}\min(c_{0}^{-1},CC_{\mu}))^{C^{\prime O(1/\eta)}}.$ This concludes what we need from a statement analogous to their Proposition 5.15.
This gives us that there exist $\eta,C^{\prime}>0,$ depending only on $s$ and $\zeta$ so that $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has $(\varphi(2s)-\zeta/2,K,1-(\zeta/M)^{2},0)$-strong thin tubes, where $K=(C^{\prime}\min\{c_{0}^{-1},CC_{\mu}\})^{C^{\prime O(1/\eta)}}.$ Next, we apply Theorem 6.9, with each $F_{y}$ being the function $d_{x}(y)=|x-y|,$ the Euclidean distance between $x$ and $y$. First, we have that $V_{x}(y)$ is the line going through $x,y$ so the measure $V_{x}\mu_{Y_{x}}$ is the same measure as $\mu_{Y_{x}}\circ\pi_{x}^{-1}$.
At this point in their proof of their Theorem 5.2, Shmerkin and Wang apply their Theorem 4.12, and use this to obtain a conclusion about Hausdorff dimension. Instead of applying an analogous result directly, we will follow their proof of Theorem 4.12 up until the statement about robustness that they obtain in the middle of their proof, then pivot to conclude our proof.
Specifically, similar to the proof of Lemma 6.12, using the assumption that $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\gtrsim 1$ we obtain that $\pi_{x}(B(z,R))\cap\operatorname{spt}\mu_{2}$ will lie inside a tube of width at most $O(R).$ So by the definition of strong thin tubes, we obtain that there is a set $X$ with $(\mu_{1}\times\nu)(X)>1-(\zeta/M)^{2}$ such that for each $x\in X$, there is a set $Y_{x}$ that has $(\mu_{2}\times\nu)(Y_{x})>1-(\zeta/M)^{2}$ and
$((\mu_{2}\times\nu)_{Y_{x}}\circ V_{x})(B(z,r))\leq KC_{0}r^{\varphi(2s)-\zeta/2},$
for some constant $C_{0}$ that only depends on the supports of $\mu_{1}$ and $\mu_{2},$ and furthermore, $Y=\{(x,y):x\in X,y\in Y_{x}\}$ is compact.
We now directly mimic a portion of the proof of Lemma 6.12, but with slightly different parameters. By Proposition 6.10 and the method of constructing $D_{\varphi(2s)-\zeta},$ the measure $\mu\circ V_{x}$ is $(t\to D_{\varphi(2s)-\zeta}(t)-\eta_{1})$-adapted at any scales in the range $\left[0,(K/C_{0})^{-C^{\prime}\eta_{1}^{-1}}\right]$ for any $\eta_{1}$ that is small enough. The fact that $\Sigma_{\tau}(D;t,d)$ is continuous as a function in $t$ follows by Lemma 6.8, so by applying Proposition 6.11, we obtain that if $\eta_{1}$ is small enough in terms of $u$ and $\zeta,$ then
$\Sigma_{\tau}(D_{\varphi(2s)-\zeta/2}(t)-\eta_{1});2s-\eta_{1},d)\geq\varphi(2s)-\zeta/2+\xi-2\eta_{1}>\varphi(2s)+\zeta/2.$
This is enough to apply Theorem 6.9 to $Y,\nu,X,$ and the associated projections, $\pi,$ with $\delta_{1}=(KC\mu)^{-C^{\prime}\eta_{1}^{-1}}$, provided that $\eta_{1}$ is small enough in terms of $u$ and $\zeta.$ This will provide us a $c^{\prime}<2c^{1/2}$, a set $Y^{\prime},$ and $X=\{(y,X_{y}):y\in Y,x\in X_{y}\subset\mathbb{R}^{2}\}$ such that $\mu_{2}(Y^{\prime})\geq 1-c^{\prime}$, and $\mu_{1}(X_{y})>1-c^{\prime}$ for each $y\in Y^{\prime}$, the set $X$ is compact and most importantly:
$\rho_{y}=(\mu_{1}\circ\pi_{y}^{-1})_{X_{y}}\text{ is }(\delta^{\prime},s+\eta,\delta^{\prime\epsilon})\text{-robust for all }\delta^{\prime}<\delta_{0},$
where $\epsilon>0$ depends on $\eta$ and $\tau$ (which, in turn, depend on $u$ and $\zeta$), and $\delta_{0}=C_{1}^{-1}(K)^{-C_{1}},$ where $C_{1}$ is a constant also depending on $u$ and $\zeta.$
The analogous statement we get is that there is a $c^{\prime}<2(\zeta/M)$, a set $Y$ and $X=\{(y,x):y\in Y,x\in X_{y}\subset\mathbb{R}^{2}\}$ such that $(\mu_{2}\times\nu)(Y)\geq 1-c^{\prime}$, $(\mu_{1}\times\nu)(X_{y})>1-c^{\prime}$ for each $y\in Y$, the set $X$ is compact and most importantly:
$\rho_{y}=((\mu_{1}\times\nu)_{X_{y}}\circ d_{y}^{-1})\text{ is }(\delta^{\prime},\varphi(2s)-\zeta/2,\delta^{\prime\epsilon})\text{-robust for all }\delta^{\prime}<\delta_{0},$
where $\epsilon>0$ depends on $s$ and $\zeta,$ and $\delta_{0}=C_{1}^{-1}\min\left\{(K)^{-1},\zeta/M\right\}^{C_{1}}$ where $C_{1}>0$ is also a constant depending on $s$ and $\zeta.$
From here, we apply Lemma 6.4 to $\rho_{y}$ to get that for every $y,$ there is a set $X^{\prime}_{y}\subseteq\mathbb{R}$ so that
$\rho_{y}(X^{\prime}_{y})>1-\sum_{m=-\log\delta_{0}}^{\infty}2^{-m\epsilon}\geq 1-\frac{2\delta_{0}}{1-2^{-\epsilon}}>1-(K)^{-1}>1-c^{\prime}$
and $(\rho_{y})_{X^{\prime}_{y}}(B(x,r))<r^{\varphi(2s)-\zeta/2}$ for every $x<\delta_{0}$, for simplicity. This implies that
[…]
Define the map $f(x,y,z,t)=\sqrt{(x-z)^{2}+(y-t)^{2}}$ and the measure $\kappa$ supported in $\mathbb{R}^{4}$ as $\mu_{1}\times\nu\times\mu_{2}\times\nu$. Let the set $X^{\prime}\subset\mathbb{R}^{4}$ be $X^{\prime}:=\{(y,x):y\in Y,x\in X_{y},d(x,y)\in X^{\prime}_{y}\}$ we have that
$\kappa(X^{\prime})=\int_{Y}(\mu_{1}\times\nu)(d_{y}^{-1}(X^{\prime}_{y})\cap X_{y})d(\mu_{2}\times\nu)(y)\geq\int_{Y}(\mu_{1}\times\nu)_{X_{y}}\rho_{y}(X^{\prime}_{y})d(\mu_{2}\times\nu)(y)\geq(1-c^{\prime})^{3}.$
Finally, using (51), and Lemma 2.26, we can calculate the entropy as
$H_{n}(\Delta(\mu_{1}\times\nu,\mu_{2}\times\nu))$ $=H(\kappa,f^{-1}(\mathcal{D}_{n}))$
$\geq\kappa(X^{\prime})H(\kappa_{X^{\prime}},f^{-1}(\mathcal{D}_{n}))$
$\geq(1-c^{\prime})^{3}(n(\varphi(2s)-\zeta/2)-\log(\delta_{0})^{-1}).$
Since $\delta_{0}^{-1}=C_{1}\min\{K,\zeta^{-1}\}^{C_{1}}$ and $K=(C^{\prime}\min\{c_{0}^{-1},C\})^{C^{\prime O(1/\eta)}},$ we have that $\log\delta_{0}^{-1}=C_{2}\log C+C_{2}$ for some large constant $C_{2}$. Furthermore we can choose $M$ large enough in terms of $\zeta$ and $s$ so that $(1-c^{\prime})^{3}(\varphi(2s)-\zeta/2)>(1-2\zeta/M)^{3}(\varphi(2s)-\zeta/2)>(\varphi(2s)-\zeta),$ which concludes the proof. ∎
7 Proof of Theorem 1.2
Since $H^{s}(A)>0$, by the Frostman lemma [12, Theorem 2.7], there exists a probability $(s,1)$-Frostman measure $\mu$ supported on $A$. Let $X$ and $Y$ be i.i.d. random variables that both have distribution measure $\mu$.
Lemma 5.1 gives us
$\max\big{\{}H_{n}(X+Y),\,H_{n}(\phi(X,Y))\big{\}}\geq(1+\epsilon)ns$
for sufficiently large $n$.
By Lemma 2.4(ii), we deduce that
$\log\#\{D_{n}(x)\mid x\in A+A\}\geq H_{n}(X+Y),$
and
$\log\#\{D_{n}(x)\mid x\in\phi(A,A)\}\geq H_{n}(\phi(X,Y)).$
Let $\mathcal{N}_{2^{-n}}(A+A)$ and $\mathcal{N}_{2^{-n}}(\phi(A,A))$ be the least number of intervals of length $2^{-n}$ needed to cover $A+A$ and $\phi(A,A)$, respectively. It is clear that
$\log\mathcal{N}_{2^{-n}}(A+A)\geq H_{n}(X+Y),~{}~{}\log\mathcal{N}_{2^{-n}}(\phi(A,A))\geq H_{n}(\phi(X,Y)).$
Hence,
$\max\left(\frac{\log\mathcal{N}_{2^{-n}}(A+A)}{n},\frac{\log\mathcal{N}_{2^{-n}}(\phi(A,A))}{n}\right)$
$\quad\geq\max\left(\frac{H_{n}(X+Y)}{n},\frac{H_{n}(\phi(X,Y))}{n}\right)\geq(1+\epsilon)s.$
When $n\to\infty$, one obtains
$\max(\underline{\dim}_{B}(A+A),\underline{\dim}_{B}(\phi(A,A)))\geq(1+\epsilon)s.$
This completes the proof.
Appendix A Appendix - Examples
In this appendix, we show details of some examples. Sometimes, we will write $\iota(x)=-x\log x$ for convenience.
Example A.1.
Given $0\leq s\leq 1$, and given $n$ sufficiently large, we will construct an $(s,O(1))$-Frostman random variable $X$ with $range(X)\subseteq[0,1)$, and compare $H_{n}(X)$, $H_{n}(X+Y)$, $H_{n}(X^{2}+Y^{2})$ and $H_{n}(XY)$, where $Y$ is an i.i.d. copy of $X$.
Let $m=m(n)\in\mathbb{N}$ and $l=l(n)\in\mathbb{N}\cup\{0\}$ be parameters to be chosen later. Take
$A=\bigcup\limits_{j=0}^{2^{m}-1}\left[j2^{-m},\,j2^{-m}+2^{-m-l}\right).$
Denote by $\mathcal{L}^{d}$ the $d$-dimensional Lesbegue measure. Then $\mathcal{L}^{1}(A)=2^{-l}$. Let $X$ be a random variable having uniform distribution on $A$, i.e.,
$P(X\in E)=\frac{\mathcal{L}^{1}(E\cap A)}{\mathcal{L}^{1}(A)}$
for any Borel set $E$.
In the following, we will proof the following results.
Claim 1. The random variable $X$ is $\left(s,O(1)\right)$-Frostman whenever $0\leq s\leq\frac{m}{l+m}$.
Claim 2. We have
$H_{n}(X)=\begin{cases}n,&\text{if }n<m,\\ m,\quad&\text{if }m\leq n<m+l,\\ n-l,\quad&\text{if }n\geq m+l.\end{cases}$ (52)
Claim 3. We have $H_{n}(X+Y)=H_{n}(X)+O(1)$ in all the cases in (52).
Claim 4. When $2m\leq n\leq m+l$, one has
$H_{n}(X^{2}+Y^{2})=(2+o(1))m=(2+o(1))H_{n}(X).$
When $n\leq m$, one has
$H_{n}(X^{2}+Y^{2})=(1+o(1))n=(1+o(1))H_{n}(X).$
Claim 5. When $2m\leq n\leq m+l$, one has
$H_{n}(XY)=(2+o(1))m=(2+o(1))H_{n}(X).$
In particular, if we take $n=m+l$ and $s=\frac{m}{l+m}$, then by Claims 2 and 3,
$H_{n}(X)=ns,\quad H_{n}(X+Y)=ns+O(1).$ (53)
The lower bound in Lemma 2.26 is met. The term involving maximum on the left-hand side of (1) can not be replaced by a single term $H_{n}(X+Y)$.
Moreover, when $n\leq m$, it follows from Claims 3 and 4 that
$\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}=(1+o(1))H_{n}(X).$
In such situation, the number $l$ can be any non-negative integer, and $X$ is $(s,O(1))$-Frostman for any $s$ with $0\leq s\leq\frac{m}{m+l}$. So all the $s$ from $(0,1)$ are covered. Therefore, in the right-hand side of (1), the term $ns$ can not be replaced by $H_{n}(X)$.
Furthermore, if we take $n=m+l$ and $s=0$, then one deduces by Claims 2, 3, and 4 that
$\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}\leq(2+o(1))m=o(n)$
when $l/m$ tends to infinity. For $l=0$ and $s=1$, we assert by Claim 2 that $H_{n}(X)=n=ns$. Combining Corollary 2.22, Lemma 2.36 and (11), we conclude that
$\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}=(1+o(1))ns.$
To sum up, the positive constant $\epsilon$ in Theorem 1.1 may not exist when $s=0$ or $1$.
Proof of Claim 1.
Let us consider the probability that $X$ takes a value in any ball $B(x,r)$. Calculation reveals that
$P\big{(}X\in B(x,r)\big{)}\ll\begin{cases}r2^{l},\quad&\text{if }0<r\leq 2^{-l-m},\\ 2^{-m},\quad&\text{if }2^{-l-m}<r\leq 2^{-m},\\ r\quad&\text{if }2^{-m}\leq r<1,\\ 1&\text{if }r\geq 1.\end{cases}$
One can verify that all the terms on the right-hand side of above formula can be bounded by $r^{\frac{m}{l+m}}$. Indeed, the random variable $X$ is $(s,O(1))$-Frostman for any $0\leq s\leq\frac{m}{l+m}$. ∎
Proof of Claim 2.
Note that
$H_{n}(X)=\sum\limits_{i=0}^{2^{n}-1}\iota\big{(}P(D_{n}(x)=i2^{-n})\big{)}.$
When $n<m$, there are $2^{n}$ values of $i$ such that $P(D_{n}(X)=i2^{-n})=2^{-n}$. When $m\leq n<m+l$, there are $2^{m}$ values of $i$ such that $P(D_{n}(X)=i2^{-n})=2^{-m}$. When $n\geq m+l$, there are $2^{n-l}$ values of $i$ such that $P(D_{n}(X)=i2^{-n})=2^{-n+l}$. The conclusion then follows. ∎
Proof of Claim 3.
When $n<m$, one can deduce by (11), together with the fact $range(X+Y)\subseteq[0,2)$, that $H_{n}(X+Y)\leq n+O(1)$. On the other hand, it follows from the independence of $X$ and $Y$ that
$H_{n}(X+Y)\geq H_{n}(X)+O(1)=n+O(1).$
When $n\geq m$, we take $X^{\prime}=D_{m+l}(X)$ and $Y^{\prime}=D_{m+l}(Y)$. Then $|X-X^{\prime}|\leq 2^{-m-l}:=K$, and $X^{\prime}$ have discrete uniform distribution on the set $\{j2^{-m}:\,j=0,1,\ldots,2^{m}-1\}$. Similar conclusions hold for $Y^{\prime}$. Moreover, since $X^{\prime}+Y^{\prime}$ take values on rationals from $2^{-m}\mathbb{Z}$, which is contained in $2^{-n}\mathbb{Z}$, one sees that
$H_{n}(X^{\prime}+Y^{\prime})=H(D_{n}(X^{\prime}+Y^{\prime}))=H(X^{\prime}+Y^{\prime}).$
By Lemma 2.4 (ii) and independence, one sees that
$H(X^{\prime}+Y^{\prime})\leq\log|range_{+}(X^{\prime}+Y^{\prime})|=m+O(1),$
and
$H(X^{\prime}+Y^{\prime})\geq H(X^{\prime})=m.$
Therefore, $H_{n}(X^{\prime}+Y^{\prime})=m+O(1)$.
For $m\leq n<m+l$, one has $K<2^{-n}$. Lemma 2.19 gives that
$H_{n}(X+Y)=H_{n}(X^{\prime}+Y^{\prime})+O(1)=m+O(1).$
For $n\geq m+l$, one has $K\geq 2^{-n}$. It follows from Lemma 2.19 that
$H_{n}(X+Y)\leq H_{n}(X^{\prime}+Y^{\prime})+n-m-l+O(1)=n-l+O(1).$
On the other hand, by independence one has $H_{n}(X+Y)\geq H_{n}(X)+O(1)=n-l+O(1)$. The proof is completed. ∎
Proof of Claim 4.
Recall that $X^{\prime}=D_{m+l}(X)$ and $Y^{\prime}=D_{m+l}(Y)$. Since the function $(x,y)\mapsto x^{2}+y^{2}$ on $[0,1]^{2}$ is $O(1)$-Lipschitz, we obtain by Lemma 2.19 that
$H_{n}(X^{2}+Y^{2})=H_{n}(X^{\prime 2}+Y^{\prime 2})+O(1)$
when $n\leq m+l$.
First, let us consider the case that $2m\leq n\leq m+l$, with $s\leq\frac{m}{m+l}\leq\frac{1}{2}$. The random variables $X^{\prime}$ and $Y^{\prime}$ both have discrete uniform distribution on the set $\{j2^{-m}:\,j=0,1,\ldots,2^{m}-1\}$. After excluding the zero terms, one can obtain that
$H_{n}(X^{\prime 2}+Y^{\prime 2})=\sum\limits_{i=0}^{2^{n+1}}\iota\Big{(}P(D_{n}(X^{\prime 2}+Y^{\prime 2})\quad=i2^{-n})\Big{)}$
$=\sum\limits_{k=0}^{2^{2m+1}}\iota\Big{(}P(X^{\prime 2}+Y^{\prime 2}=k2^{-2m})\Big{)}=\sum\limits_{k\in\mathcal{K}}\iota\big{(}2^{-2m}T(k)\big{)},$
where $\mathcal{K}$ denotes the set consisting of integers $k$ such that $k=j_{1}^{2}+j_{2}^{2}$ for some $10\leq j_{1},j_{2}\leq 2^{m}-$ 1, and
$T(k)=\sum\limits_{0\leq j_{1},j_{2}\leq 2^{m}-1\atop j_{1}^{2}+j_{2}^{2}=k}1.$
Note that $T(k)$ does not exceed the number of representations of $k$ as a sum of two squares, which satisfies that (see Section 18.7 of [5])
$\log T(k)\ll\frac{\log k}{\log\log k}\ll\frac{m}{\log m}.$
On the other hand, one has $T(k)\geq 1$ for all $k\in\mathcal{K}$. So
$\log\left(2^{-2m}T(k)\right)=-2m+O\Big{(}\frac{m}{\log m}\Big{)}.$
Then, we deduce that
$H_{n}(X^{\prime 2}+Y^{\prime 2})=-\sum\limits_{k\in\mathcal{K}}2^{-2m}T(k)\cdot\log\left(2^{-2m}T(k)\right)$
$\qquad=2m\left(1+O\Big{(}\frac{1}{\log m}\Big{)}\right)\cdot 2^{-2m}\sum\limits_{k\in\mathcal{K}}T(k)=(2+o(1))m,$
since $\sum\nolimits_{k\in\mathcal{K}}T(k)=2^{2m}$.
Next, we consider the case that $n\leq m$. The random variables $X^{\prime}$ and $Y^{\prime}$ both have discrete uniform distribution on the set $\{j2^{-n}:\,j=0,1,\ldots,2^{n}-1\}$. By (11), we have $H_{n}(X^{2}+Y^{2})\leq n+O(1)$. On the other hand,
$H_{n}(X^{2}+Y^{2})=H_{n}(X^{\prime 2}+Y^{\prime 2})+O(1)\geq H_{n}(X^{\prime 2})+O(1).$
It is sufficent to show that $H_{n}(X^{\prime 2})=(1+o(1))n$.
For $n\leq i\leq 2^{n}-1$, one has
$P(D_{n}(X^{\prime 2})=i2^{-n})=P\big{(}\sqrt{i2^{-n}}\leq X^{\prime}<\sqrt{(i+1)2^{-n}}\big{)}$
$\quad=\frac{2^{-n/2}}{2\sqrt{i}}\left(1+O\Big{(}\frac{1}{n}\Big{)}\right)+O(2^{-n})=\frac{2^{-n/2-1}}{\sqrt{i}}+O\Big{(}\frac{2^{-n/2}}{n^{3/2}}\Big{)}.$
For $0\leq i<n$,
$P(D_{n}(X^{\prime 2})=i2^{-n})\leq P\big{(}X^{\prime}<\sqrt{n2^{-n}}\big{)}\ll\sqrt{n}2^{-n/2}.$
It follows that
$H_{n}(X^{\prime 2})$ $=\left(\sum\limits_{i=0}^{n-1}+\sum\limits_{i=n}^{2^{n}-1}\right)\iota\big{(}P(D_{n}(X^{\prime 2})\big{)}$
$=O\Big{(}-\sum\limits_{i=0}^{n-1}\sqrt{n}2^{-n/2}\log O\big{(}\sqrt{n}2^{-n/2}\big{)}\Big{)}-(1+o(1))\sum\limits_{i=n}^{2^{n}-1}\frac{2^{-n/2-1}}{\sqrt{i}}\log\frac{2^{-n/2-1}}{\sqrt{i}}$
$=O(1)+(1+o(1))\left((n/2+1)2^{-n/2-1}\sum\limits_{i=1}^{2^{n}-1}\frac{1}{\sqrt{i}}+2^{-n/2-2}\sum\limits_{i=1}^{2^{n}-1}\frac{\log i}{\sqrt{i}}\right)$
$=(1+o(1))\left(\frac{n}{2}\cdot 2^{-\frac{n}{2}-1}\cdot 2\cdot 2^{\frac{n}{2}}+2^{-\frac{n}{2}-2}\cdot 2\cdot 2^{\frac{n}{2}}\cdot n\right)=(1+o(1))n.$
Here we have used the estimate
$\sum\limits_{i=1}^{N}\frac{(\log i)^{h}}{\sqrt{i}}=(1+o(1))2\sqrt{N}(\log N)^{h}$
as $N\rightarrow\infty$, where $h\in\{0,1\}$. The proof is completed. ∎
Proof of Claim 5.
Since the function $(x,y)\mapsto xy$ on $[0,1]^{2}$ is $O(1)$-Lipschitz, we obtain by Lemma 2.19 that
$H_{n}(XY)=H_{n}(X^{\prime}Y^{\prime})+O(1)$
when $n\leq m+l$. We are considering the case that $2m\leq n\leq m+l$, with $s\leq\frac{m}{m+l}\leq\frac{1}{2}$.
Similarly as in the proof of Claim 4, after excluding the zero terms, one can obtain that
$H_{n}(X^{\prime}Y^{\prime})=\sum\limits_{k\in\mathcal{K}^{\prime}}\iota\big{(}2^{-2m}T^{\prime}(k)\big{)},$
where $\mathcal{K}^{\prime}$ denotes the set consisting of integers $k$ such that $k=j_{1}j_{2}$ for some $10\leq j_{1},j_{2}\leq 2^{m}-$ 1, and
$T^{\prime}(k)=\sum\limits_{0\leq j_{1},j_{2}\leq 2^{m}-1\atop j_{1}j_{2}=k}1.$
Denote by $\tau$ the divisor function, i.e., $\tau(k)$ counts the number of divisors of $k$. Then $T^{\prime}(k)\leq\tau(k)$. In [13], it is proved that
$\log\tau(k)\leq\frac{1.5379\log k}{\log\log k}$
for all $n\geq 2$. It follows that
$\log T(k)\ll\frac{\log k}{\log\log k}\ll\frac{m}{\log m}.$
Now, similar arguments as in the proof of Claim 4 leads to the conclusion that $H_{n}(XY)=(2+o(1))m$. ∎
Example A.2.
Let $X$ and $Y$ be the independent $(s,O(1))$-Frostman random variables defined in A.1, with $n=m+l$ and $s=\frac{m}{l+m}$. Consider $\tilde{X}=\sqrt{X}$ and $\tilde{Y}=\sqrt{Y}$. By Example A.1, one sees that
$H_{n}(\tilde{X}^{2}+\tilde{Y}^{2})=H_{n}(X+Y)=(1+o(1))ns.$
Therefore, the term involving maximum on the left-hand side of (1) can not be replaced by a single term $H_{n}(\phi(X,Y))$.
Example A.3.
Let $X$, $Y$ be the independent $(s,O(1))$-Frostman random variables defined in A.1, with $n=m+l$ and $s=\frac{m}{l+m}$. Consider $\tilde{X}=1+X$, which is $(s,O(1))$-Frostman with range contained in $[1,2)$. Let $\tilde{Y}$ be an i.i.d. copy of $X$. Take $\phi(x,y)=(x+y)^{2}$, which is a degenerate quadratic form. By Example A.1, one can deduce that
$H_{n}(\tilde{X})=H_{n}(Y)=ns,\qquad H_{n}(\tilde{X}+Y)=(1+o(1))ns.$
Since $range(\tilde{X}+Y)\subseteq[1,3)$, one deduces from Lemma 2.19 that
$H_{n}\big{(}(\tilde{X}+Y)^{2}\big{)}=H_{n}(\tilde{X}+Y)+O(1)=(1+o(1))ns.$
This gives
$\max\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\}=(1+o(1))ns.$
So, the requirement that $\phi$ is non-degenerate in Theorem 1.1 is essential.
Example A.4.
Assume that the random variables $X$ and $Y$ are independent, $X$ has uniform distribution on the set
$A=\bigcup\limits_{j=0}^{2^{m}-1}\left[1+j2^{-m},\,1+j2^{-m}+2^{-2m}\right),$
and $Y$ has uniform distribution on $[0,2^{-2m}]$. One can verify that $X$ is $(1/2,O(1))$-Frostman, and $Y$ is $(1/2,O(2^{m}))$-Frostman. We take $s=1/2$ and consider $n=2m$.
Claim. We have $\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}=ns+O(1)$.
In this example, we do not obtain a positive constant $\epsilon$ in Theorem 1.1. The reason is that the Frostman constant for $Y$ grows with $m$. More concretely, given $C$ of magnitude $2^{m}$, when $n=2m$ is of magnitude $\log C$, the two entropies may be small simultaneously. Therefore, in the conclusion of Theorem 1.1, it is essential to require $n>N$, where $N$ depends on the Frostman constant $C$.
Proof of Claim.
Let $(X^{\prime},Y^{\prime})=D_{n}(X,Y)$, which has discrete uniform distribution on $\{(1+j2^{-m},0):\,j=0,1,\ldots,2^{m}-1\}$. By Lemma 2.19, one has
$H_{n}(X+Y)=H_{n}(X^{\prime}+Y^{\prime})+O(1),\text{ and }H_{n}(X^{2}+Y^{2})=H_{n}(X^{\prime 2}+Y^{\prime 2})+O(1).$ (54)
Note that $H_{n}(X^{\prime}+Y^{\prime})=H_{n}(X^{\prime})$ and
$H_{n}(X^{\prime 2}+Y^{\prime 2})=H_{n}(X^{\prime 2})=H_{n}(X^{\prime})+O(1).$
The conclusion follows by noting that $H_{n}(X^{\prime})=m=ns$. ∎
Example A.5.
This example is an variant of Example A.4. Let $W=(X,Y)$, and suppose that the joint distribution of $(X,Y)$ is the uniform distribution on the set $A=\bigcup\nolimits_{j=0}^{2^{m}-1}\big{[}j2^{-m},\,j2^{-m}+2^{-2m}\big{)}^{2}$. Note that $X$ and $Y$ are dependent. The marginal distributions of $X$ and $Y$ are both the uniform distribution on $\bigcup\nolimits_{j=0}^{2^{m}-1}\big{[}j2^{-m},\,j2^{-m}+2^{-2m}\big{)}$. So both $X$ and $Y$ are $(1/2,O(1))$-Frostman. Let $W_{1}=(X_{1},Y_{1}),\,(X_{2},Y_{2})$ be i.i.d. copies of $(X,Y)$. We consider $n\leq 2m$.
Claim 1. We have $\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}\leq m+O(1)$.
Claim 2. The random variable $W$ is $\left(s,O(2^{m(2s-1)})\right)$-Frostman for any $1/2\leq s\leq 2$.
Claim 3. For $m\leq n\leq 2m$, we have $H_{n}(|W_{1}-W_{2}|)=(1+o(1))m$.
If we take $n=2m$ in Claim 1, then $\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}\leq n/2+O(1)$. In such a situation, the two entropies can be small simultaneously. Therefore, the requirement of independence of $X$ and $Y$ is essential in Theorem 1.1.
Moreover, for the random variable $W$, its Frostman constant $C$ can grow with $m$. If we consider $n=2m$ in Theorem 2.32, if $21<s<3/$ 2, then one has $H_{n}(|W_{1}-W_{2}|)=(1+o(1))n/2<nt$ for some suitably chosen $t$ with $s/2<t<s-1/2$. That is to say, the term $-\log C$ in Theorem 2.32 is essential.
Proof of Claim 1.
Let $(X^{\prime},Y^{\prime})=D_{2m}(X,Y)$, which has discrete uniform distribution on $\{(j2^{-m},j2^{-m}):\,j=0,1,\ldots,2^{m}-1\}$. Similar to the previous argument, the equalities in (54) hold. Note that
$range(X^{\prime}+Y^{\prime})\subseteq\{j2^{-m+1}:\,j=0,1,\ldots,2^{m}-1\},\text{ and}$
$range(X^{\prime 2}+Y^{\prime 2})\subseteq\{j^{2}2^{-2m+1}:\,j=0,1,\ldots,2^{m}-1\}.$
This gives
$\max\{H_{n}(X^{\prime}+Y^{\prime}),\,H_{n}(X^{\prime 2}+Y^{\prime 2})\}\leq m.$
Proof of Claim 2.
For any ball $B\big{(}(x,y),r\big{)}$, calculation reveals that
$P\big{(}X\in B(x,r)\big{)}\ll\begin{cases}r^{2}2^{3m},\quad&\text{if }0<r\leq 2^{-2m+1/2},\\ 2^{-m},\quad&\text{if }2^{-2m+1/2}<r\leq 2^{-m+1/2},\\ r\quad&\text{if }2^{-m+1/2}\leq r<1,\\ 1&\text{if }r\geq 1.\end{cases}$
One can verify that all the terms on the right-hand side of above formula can be bounded by $2^{m(2s-1)}r^{s}$ for any $1/2\leq s\leq 2$. The random variable $(X,Y)$ is $\left(s,O(2^{m(2s-1))}\right)$-Frostman. ∎
Proof of Claim 3.
Note that $m\leq n\leq 2m$. Put $(X_{i}^{\prime},Y_{i}^{\prime})=D_{n}(X_{i},Y_{i})$ $(i=1,2)$. Since the function
$(x_{1},y_{1},x_{2},y_{2})\mapsto\sqrt{(x_{1}-x_{2})^{2}+(y_{1}-y_{2})^{2}}$
is $O(1)$-Lipschitz, one asserts by Lemma 2.19 that
$H_{n}\big{(}|(X_{1},Y_{1})-(X_{2},Y_{2})|\big{)}=H_{n}\big{(}|(X_{1}^{\prime},Y_{1}^{\prime})-(X_{2}^{\prime},Y_{2}^{\prime})|\big{)}+O(1).$
We can verify that both $(X_{1}^{\prime},Y_{1}^{\prime})$ and $(X_{2}^{\prime},Y_{2}^{\prime})$ have discrete uniform distribution on the set $\{(j2^{-m},j2^{-m}):\,j=0,1,\ldots,2^{m}-1\}$.
By discarding zero-terms, one has
$H_{n}\big{(}|(X_{1}^{\prime},Y_{1}^{\prime})-(X_{2}^{\prime},Y_{2}^{\prime})|\big{)}=H_{n}(\sqrt{2}|X_{1}^{\prime}-X_{2}^{\prime}|)=H_{m}(|X_{1}^{\prime}-X_{2}^{\prime}|)+O(1)=(1+o(1))m.$
Example A.6.
By Lemma 2.23, we have the following statement: Let $k,l\in\mathbb{Q}$ be given. Then
$H_{n}(kX+lY)-H_{n}(X)\ll_{k,l}(H_{n}(X+Y)-H_{n}(X))+O_{k,l}(1)$ (55)
for any $n\in\mathbb{N}$ and any random variable $X$ with bounded range, where $Y$ is i.i.d. with $X$. In this example, we show that the above statement may fail when $k,l$ are $\mathbb{Q}$-independent. In particular, we consider the situation that $(k,l)=(1,\sqrt{2})$.
Let $X,Y$ be the random variables defined in Example A.1, and $X^{\prime}=D_{m+l}(X)$, $Y^{\prime}=D_{m+l}(Y)$, where the parameters are chosen as the following: for any given number $\eta>0$, let $m$ be any sufficiently large natural number, $l$ be a natural number with $(1+\eta/2)m\leq l\leq(1+\eta)m$, and $n=m+l$. By Example A.1, the random variables $X$ and $Y$ are $(s,O(1))$-Frostman, where $s=\frac{m}{l+m}\in[\frac{1}{2+\eta},1/2)$. Also, the equalities in (53) hold.
Claim 1. Denote $A^{\prime}=\{i2^{-m}:\,i=0,1,\ldots,2^{m}-1\}$. For any integer $j$ with $0\leq j<(1+\sqrt{2})2^{n}$, the set $\big{\{}(a,a^{\prime})\in A^{\prime 2}:\,a+\sqrt{2}a^{\prime}\in[j2^{-n},(j+1)2^{-n})\big{\}}$ contains at most one element, provided that $m$ is sufficiently large.
Claim 2. We have $H_{n}(X+\sqrt{2}Y)=2ns+O(1)$.
As a result, we see that $H_{n}(X+Y)-H_{n}(X)=O(1)$, while $H_{n}(X+\sqrt{2}Y)-H_{n}(X)=ns+O(1)$. Therefore (55) fails.
Proof of Claim 1.
Assume on the contrary that the set contains two different elements $(a_{1},a_{1}^{\prime})$ and $(a_{2},a_{2}^{\prime})$ of $A^{\prime 2}$. Write $a_{t}=i_{t}2^{-m}$ and $a_{t}^{\prime}=i_{t}^{\prime}2^{-m}$ $(t=1,2)$. Since
$|(a_{1}+\sqrt{2}a_{1}^{\prime})-(a_{2}+\sqrt{2}a_{2}^{\prime})|<2^{-n},$
one sees that
$|(i_{1}-i_{2})+\sqrt{2}(i_{1}^{\prime}-i_{2}^{\prime})|<2^{-n+m}=2^{-l}\leq 2^{-(1+\eta/2)m}.$
Write $i=i_{1}-i_{2}$ and $i^{\prime}=i_{1}^{\prime}-i_{2}^{\prime}$. Then $|i|,|i^{\prime}|<2^{m}$ and $|i+\sqrt{2}i^{\prime}|<2^{-(1+\eta/2)m}$. Combining the fact that $(i,i^{\prime})\neq(0,0)$, one can deduce that neither $i$ nor $i^{\prime}$ can be 00.
However, the irrational $\sqrt{2}$ is an algebraic number of degree $2$. By Liouville’s approximation theorem, there exists an absolute positive constant $C$ such that
$\left|\sqrt{2}-\frac{p}{q}\right|>\frac{C}{q^{2}}$
for any $p\in\mathbb{Z}$ and $q\in\mathbb{N}$. It follows that $|i+\sqrt{2}i^{\prime}|>C/|i^{\prime}|>C2^{-m}$. This gives $2^{-(1+\eta/2)m}>C2^{-m}$, which is a contradiction when $m$ is sufficiently large. ∎
Proof of Claim 2.
One can deduce by Lemma 2.19 that $H_{n}(X+\sqrt{2}Y)=H_{n}(X^{\prime}+\sqrt{2}Y^{\prime})+O(1).$ So, it is sufficient to prove that $H_{n}(X^{\prime}+\sqrt{2}Y^{\prime})=2m$.
Note that both $X^{\prime}$ and $Y^{\prime}$ have discrete uniform distribution on $A^{\prime}$. Denote by $p_{j}=P\big{(}D_{n}(X^{\prime}+\sqrt{2}Y^{\prime})=j\big{)}$ for $0\leq j<(1+\sqrt{2})2^{n}$. By Claim 1, we see that either $p_{j}=2^{-2m}$ or $p_{j}=0$. Therefore,
$H_{n}(X^{\prime}+\sqrt{2}Y^{\prime})=-\sum\limits_{j}p_{j}\log p_{j}=\sum\limits_{j}p_{j}\cdot 2m=2m.$
The proof is completed. ∎
Appendix B Appendix - Existence of Certain Random Variable
In this appendix, we prove two lemmas concerning the existence of certain random variables.
Lemma B.1.
Let $(A,B,C)$ be a given random variable. Let $h$ be a function such that $h(C)$ is a well-defined discrete random variable. Then there exists a random variable $(A^{\prime},B^{\prime},C^{\prime},B^{\prime\prime})$ such that $(A^{\prime},B^{\prime},C^{\prime})\equiv(A,B,C)$, $(B^{\prime\prime},h(C^{\prime}))\equiv(B,h(C))$, and $(B^{\prime\prime}|A^{\prime},B^{\prime},h(C^{\prime}))\equiv(B^{\prime\prime}|h(C^{\prime}))$. Moreover, $B^{\prime\prime}$ and $B^{\prime}$ are conditionally i.i.d. given $h(C^{\prime})$. Furthermore, $(A^{\prime},B^{\prime})$ and $B^{\prime\prime}$ are conditionally independent given $h(C^{\prime})$.
Proof.
Suppose that $A,B,C$ are defined as measurable functions from a probability space $(\Omega,\mathcal{B},P)$ to some other measurable space. Now we construct a new probability space $(\Omega^{\prime},\mathcal{B}^{\prime},P^{\prime})$. Let
$\Omega=\bigcup_{c\in range(h(C))}\Omega_{c},\quad\Omega_{c}:=\{h(C)=c\}=\{\omega\in\Omega:\,h(C(\omega))=c\}.$
Construct
$\Omega^{\prime}=\bigcup_{c}(\Omega_{c}\times\Omega_{c}).$
Then $(\omega,\nu)\in\Omega^{\prime}$ if and only if $\omega,\nu\in\Omega_{c}$ for some $c$, if and only if $h(C(\omega))=h(C(\nu))$. The $\sigma$-algebra $\mathcal{B}^{\prime}$ can be generated by the related Cartisian products of Borel sets in $\mathcal{B}$, and the probability measure $P^{\prime}$ can be induced by the following: if $E\times F$ is a Borel subset of $\Omega_{c}\times\Omega_{c}$ for some $c$, then
$P^{\prime}(E\times F)=\begin{cases}\frac{P(E)P(F)}{P(\Omega_{c})},\quad&\text{if }P(\Omega_{c})>0,\\ 0,&\text{if }P(\Omega_{c})=0.\end{cases}$
In the following, we assume without loss of generality that $P(\Omega_{c})>0$ for all $c\in range(h(C))$. Now for any random variable $D$ defined on $(\Omega,\mathcal{B},P)$, let $D^{\prime}$ be the following random variable on $(\Omega^{\prime},\mathcal{B}^{\prime},P^{\prime})$:
$D^{\prime}(\omega,\nu)=D(\omega),\quad(\omega,\nu)\in\Omega^{\prime}.$
By this definition, one sees that $\big{(}f(D))^{\prime}=f(D^{\prime})$ for any admissible function $f$. In the following, we verify that $D^{\prime}\equiv D$. For any Borel set $L$,
$P^{\prime}(D^{\prime}\in L)$ $=P^{\prime}(\{(\omega,\nu)\in\Omega^{\prime}:\,D^{\prime}(\omega,\nu)\in L)$
$=\sum\limits_{c}P^{\prime}(\{(\omega,\nu)\in\Omega_{c}\times\Omega_{c}:\,D(\omega)\in L\})$
$=\sum\limits_{c}\frac{P(D\in L,h(C)=c)P(h(C)=c)}{P(h(C)=c)}=P(D\in L).$
In particular, we can take $D=(A,B,C)$ to get $(A^{\prime},B^{\prime},C^{\prime})\equiv(A,B,C)$. All their corresponding marginal distributions are also the same. One has $h(C)\equiv h(C^{\prime})$ as well.
Next, we define $B^{\prime\prime}$ as a function on $(\Omega^{\prime},\mathcal{B}^{\prime},P^{\prime})$, given by
$B^{\prime\prime}(\omega,\nu)=B(\nu),\quad(\omega,\nu)\in\Omega^{\prime}.$
Then, for any Borel sets $L_{1}$, $L_{2}$,
$P^{\prime}(B^{\prime\prime}\in L_{1},D^{\prime}\in L_{2},h(C^{\prime})=c)=P^{\prime}(\{(\omega,\nu)\in\Omega^{\prime}:\,B^{\prime\prime}(\omega,\nu)\in L_{1},D^{\prime}(\omega,\nu)\in L_{2},h(C^{\prime}(\omega,\nu))=c\})$
$\qquad=P^{\prime}(\{(\omega,\nu)\in\Omega_{c}\times\Omega_{c}:\,B(\nu)\in L_{1},D(\omega)\in L_{2}\})=\frac{P(D\in L_{2},h(C)=c)P(B\in L_{1},h(C)=c)}{P(h(C)=c)}.$
When we take $D$ to be trivial (a constant function), then $P^{\prime}(B^{\prime\prime}\in L_{1},h(C^{\prime})=c)=P(B\in L_{1},h(C)=c)$ and then $(B^{\prime\prime},h(C^{\prime}))\equiv(B,h(C))$. Moreover, the previous argument implies that $(D^{\prime},h(C^{\prime}))\equiv(D,h(C))$. We can calculate further to obtain
$P^{\prime}(B^{\prime\prime}\in L_{1},D^{\prime}\in L_{2},h(C^{\prime})=c)$ $=\frac{P(D\in L_{2},h(C)=c)P(B\in L_{1},h(C)=c)}{P(h(C)=c)}$
$=\frac{P^{\prime}(D^{\prime}\in L_{2},h(C^{\prime})=c)P^{\prime}(B^{\prime\prime}\in L_{1},h(C^{\prime})=c)}{P^{\prime}(h(C^{\prime})=c)}$
$=P^{\prime}(D^{\prime}\in L_{2},h(C^{\prime})=c)P^{\prime}(B^{\prime\prime}\in L_{1}|h(C^{\prime})=c).$
So $P^{\prime}(B^{\prime\prime}\in L_{1}|D^{\prime}\in L_{2},h(C^{\prime})=c)=P^{\prime}(B^{\prime\prime}\in L_{1}|h(C^{\prime})=c)$, i.e. $(B^{\prime\prime}|D^{\prime},h(C^{\prime}))\equiv(B^{\prime\prime}|h(C^{\prime}))$. Taking $D=(A,B)$, one gets $(B^{\prime\prime}|A^{\prime},B^{\prime},h(C^{\prime}))\equiv(B^{\prime\prime}|h(C^{\prime}))$.
Moreover, if we take $D=B$, then $(B^{\prime\prime}|B^{\prime},h(C^{\prime}))\equiv(B^{\prime\prime}|h(C^{\prime}))$. That is to say, $B^{\prime\prime}$ and $B^{\prime}$ are conditionally independent given $h(C^{\prime})$. Since $(B^{\prime\prime},h(C^{\prime}))\equiv(B,h(C))\equiv(B^{\prime},h(C^{\prime}))$, one also deduces that $B^{\prime\prime}$, $B$ are conditionally i.i.d. given $h(C^{\prime})$.
Furthermore, it follows from $(B^{\prime\prime}|A^{\prime},B^{\prime},h(C^{\prime}))\equiv(B^{\prime\prime}|h(C^{\prime}))$ that $(A^{\prime},B^{\prime})$ and $B^{\prime\prime}$ are conditionally independent given $h(C^{\prime})$.
Without confusion, we will simplify this $P^{\prime}$ as $P$ in applications for convenience. ∎
Lemma B.2.
Let $X$ be a random variable, and $E$ be an event with $P(E)>0$. Then there is another random variable $X_{E}$ such that $X_{E}\equiv(X|E)$. Moreover, for any function $f$ with $domain(f)\supseteq range(X)$, we have
$H_{n}(f(X_{E}))=H(D_{n}(f(X))|E),$
and
$H_{n}(f(X_{E}))\leq\frac{H_{n}(f(X))}{P(E)}.$
Furthermore, if $X$ is $(s,C)$-Frostman, then $X^{\prime}$ is $(s,C/P(E))$-Frostman.
Proof.
Assume that $X$ is defined on some probability space $(\Omega,\mathcal{B},P)$. The conditional measure given $E$, denoted by $P_{E}$, is given by
$P_{E}(L)=\frac{P(L\cap E)}{P(E)},\quad(L\in\mathcal{B}).$
Define $X_{E}$ on $(\Omega,\mathcal{B},P_{E})$ by
$X_{E}(\omega)=X(\omega),\quad(\omega\in\Omega).$
It is not hard to verify that $X_{E}\equiv(X|E)$, and $f(X^{\prime})=f(X)$. Now
$H_{n}(f(X_{E}))=H(D_{n}(f(X_{E})))=-\sum\limits_{x}P_{E}(D_{n}(f(X^{\prime}))=x)\log_{2}P_{E}(D_{n}(f(X_{E}))=x)=H(D_{n}(f(X))|E).$
Let $\mathcal{A}$ and $\mathcal{B}$ be two partitions of $\Omega$:
$\mathcal{A}=\big{\{}\{D_{n}(f(X))=x\}:\,x\in range(D_{n}(f(X)))\big{\}},\quad\mathcal{B}=\left\{E,E^{c}\right\}.$
With language of entropy of measure, we have
$H_{n}(f(X))=H(P,\mathcal{A}),\quad H_{n}(f(X_{E}))=H(P_{E},\mathcal{A}).$
By Lemma 2.2(iv), we have
$H(P,\mathcal{A})\geq H(P,\mathcal{A}|\mathcal{B})=P(E)H(P_{E},\mathcal{A})+P(E^{c})H(P_{E^{c}},\mathcal{A})\geq P(E)H(P_{E},\mathcal{A}).$
When $X$ is $(s,C)$-Frostman, for any ball $B(x,r)$, we have
$P_{E}(X_{E}\in B(x,r))\leq\frac{P(X\in B(x,r))}{P(E)}\leq\frac{Cr^{s}}{P(E)}.$
Hence $X_{E}$ is $(s,C/P(E))$-Frostman. ∎
Acknowledgements
Thang Pham and Steven Senger would like to thank the Vietnam Institute for Advanced Study in Mathematics (VIASM) and the Erdős Center for the hospitality and for the excellent working condition.

### v2

@@
 Abstract
-Let $\phi(x,y)$ be a non-degenerate rational quadratic form.
-Let $X$ and $Y$ be independent $(s,C)$-Frostman random variables whose ranges are contained in $[-c_{1},c_{1}]$, with $10<s<$ 1, $C,c_{1}\geq 1$.
-We prove that there exist a positive constant $\epsilon=\epsilon(s,\phi)$ and an integer $N=N(s,C,c_{1},\phi)$ such that
+We introduce Frostman conditions for bivariate random variables and study discretized entropy sum–product phenomena in both independent and dependent regimes.
+In the independent setting, we show that for any non-degenerate rational quadratic form $\phi(x,y)$ and $0<s<1$, there exists some positive constant $\epsilon=\epsilon(\phi,s)$ such that any independent marginally $(s,C;s,C)$-Frostman random variable $(X,Y)$ satisfies
 $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon)$
-for all $n>N$.
+for sufficiently large $n$.
 The proof introduces a novel multi‑step entropy framework, combining the submodularity formula, the discretized entropy Balog-Szemerédi-Gowers theorem, and state-of-the-art results on the Falconer distance problem, to reduce general forms to a diagonal core case.
-As an application, we derive a result on a discretized sum-product type problem.
-In particular, for a $\delta$-separated set $A\subset[0,1]$ of cardinality $\delta^{-s}$, satisfying some non-concentration conditions, there exists $\epsilon=\epsilon(s,\phi)>0$ such that
-$\mathcal{E}_{\delta}(A+A)+\mathcal{E}_{\delta}(\phi(A,A))\gg\delta^{-\epsilon}(\#A)$
-for all $\delta$ small enough. Here by $\mathcal{E}_{\delta}(A)$ we mean the $\delta$-covering number of $A$.
+In the dependent setting, we obtain analogous bounds for special polynomials.
+As an application, we derive that there exists $\epsilon=\epsilon(s)>0$ such that, for any $\delta$-separated set $A\subset[0,1]$ of cardinality $\delta^{-s}$ satisfying some non-concentration condition and any subset $G\subseteq A\times A$ with $|G|\geq\delta^{\epsilon}|A|^{2}$, the $\delta$-covering numbers satisfy
+$\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2})\gg_{\alpha,\beta,s}C^{-O_{s}(1)}\delta^{-\epsilon}|A|,$
+whenever $\delta$ is small enough.
 1 Introduction
@@
 $a_{2}^{2}-4a_{1}a_{3}\neq 0$, $X$ and $Y$ are independent $(s,C)$-Frostman random variables whose ranges are bounded, and $s\in(0,1)$.
-We state our main result as follows.
+Throughout this paper, the quadratic polynomial $\phi(x,y)$ is called special if there exist univariate real-analytic functions $h,a,b$ such that $\phi(x,y)=h\bigl{(}a(x)+b(y)\bigr{)}$.
+A direct computation shows that a special quadratic polynomial is one of the following two forms: $\phi(x,y)=\alpha x^{2}+\beta y^{2}$ and $\phi(x,y)=\gamma xy$, for $\alpha,\beta,\gamma\in\mathbb{R}^{*}$.
+For $\phi(x,y)=x\cdot y$, we recall the following theorem due to Máthé and O’Regan in [12].
+Let $X$ and $Y$ be independent $(s,C)$-Frostman random variables whose ranges are contained in $c_{1},c_{1}]$, with $0<s<1$, and $C,c_{1}\geq 1$.
+Then there exists a positive constant $\epsilon=\epsilon(s)$ and an integer $N=N(s,C,c_{1})$ such that for all $n>N,$ we have
+$\max\left\{H_{n}(X+Y),\,H_{n}(X\cdot Y)\right\}\geq n(s+\epsilon).$ (1)
+To state our main theorems, we first introduce the following definitions for bivariate random variables $(X,Y)$.
+Definition 1.1.
+Let $s,s_{1},s_{2}\geq 0$ and $C,C_{1},C_{2}>0$.
+(i) We say that the random variable $(X,Y)$ is jointly $(s_{1},s_{2},C)$-Frostman if
+$P(X\in B(x,r_{1}),\,Y\in B(y,r_{2}))\leq Cr_{1}^{s_{1}}r_{2}^{s_{2}}$
+for any ball $B(x,r_{1})$ and $B(y,r_{2})$.
+(ii) We say that the random variable $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman if $X$ and $Y$ are $(s_{1},C_{1})$-Frostman and $(s_{2},C_{2})$-Frostman, respectively.
+(iii) We say a random variable $X$ is $(s,C)$-Frostman given $Y$ if
+$P(X\in B(x,r),\,Y\in E)\leq Cr^{s}P(Y\in E)$
+for any ball $B(x,r)$ and Borel set $E$.
+Moreover, We say that the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, if $X$ is $(s_{1},C_{1})$-Frostman given $Y$, and $Y$ is $(s_{2},C_{2})$-Frostman given $X$.
+These definitions are not equivalent. At a common exponent level $(s_{1},s_{2})$, we have the implications
+$\boxed{\text{\parbox{116.65646pt}{\centering independent\\ and marginally Frostman\@add@centering}}}\;\Rightarrow\;\boxed{\text{conditionally Frostman}}\;\Rightarrow\;\boxed{\text{jointly Frostman}}\;\Rightarrow\;\boxed{\text{marginally Frostman}}$
+In Section 7, we provide examples illustrating each level:
+  • an independent pair that is marginally (s,O(1);s,O(1))(s,O(1);\,s,O(1))-Frostman for $0\leq s\leq 1$ (Example 7.1);
+  • a dependent pair that is jointly $(1/4,1/4,O(1))$-Frostman, but $(1/2,O(1);1/2,O(1))$-marginally Frostman (Example 7.6);
+  • a dependent pair that is conditionally $(1/4,O(1);1/4,O(1))$-Frostman, but $(1/2,1/2;O(1))$-jointly Frostman (Example 7.7).
+Our first theorem addresses the case that $X$ and $Y$ are independent real random variables, $(X,Y)$ forms marginally $(s_{1},C_{1};S_{2},C_{2})$-Frostman, and $\phi(x,y)$ is a non-degenerate rational quadratic form.
 Theorem 1.2.
-Let $\phi(x,y)$ be a non-degenerate rational quadratic form.
-Let $X$ and $Y$ be independent $(s,C)$-Frostman random variables whose ranges are contained in $[-c_{1},c_{1}]$, with $10<s<$ 1, and $C,c_{1}\geq 1$.
-Then there exist a positive constant $\epsilon=\epsilon(s,\phi)$ and an integer $N=N(s,C,c_{1},\phi)$ such that for all $n>N,$ we have
-$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon).$ (1)
-To show the sharpness of this theorem, we summarize the main statements of several examples in Section A:
+Let $\phi(x,y)$ be a non-degenerate rational quadratic form, $0<s<1$ and $C_{1},C_{2},c>0$.
+There exist a positive constant $\epsilon=\epsilon(\phi,s)$ and an integer $N=N(\phi,s,C_{1},C_{2},c)$ such that the following statement holds: Let $X$ and $Y$ be independent real random variables whose ranges are contained in [c,c][-c,c].
+Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};S_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
+Then
+$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon)$ (2)
+for all $n>N$.
+Here, a constant depends on $\phi$ means that it depends on the coefficients $a_{1}$, $a_{2}$ and $a_{3}$.
+To show the sharpness of this theorem, we summarize the main statements of several examples in Section 7:
   1. 1.
-    Any $(s,C)$-Frostman random variable $Z$ has the property that $H_{n}(Z)\geq ns-\log C+O(1)$ (see Example A.1 for an example matching this bound).
+    Any $(s,C)$-Frostman random variable $Z$ has the property that $H_{n}(Z)\geq ns-\log C+O(1)$ (see Example 7.1 for an example matching this bound).
 The lower bound in (2) says that the discretized entropies of $X+Y$ and $\phi(X,Y)$ cannot be small simultaneously.
-We show in Examples A.1 and A.2 that at least one of the terms $H_{n}(X+Y)$ or $H_{n}(\phi(X,Y))$ can be as small as $(1+o(1))ns$.
+We show in Examples 7.1 and 7.2 that at least one of the terms $H_{n}(X+Y)$ or $H_{n}(\phi(X,Y))$ can be as small as $(1+o(1))ns$.
   2. 2.
-    The positive $\epsilon$ might not exist if one of the conditions of the theorem is not satisfied, including when $s=0$ or $1$ (see Example A.1), the Frostman constant $C$ grows with $n$ (see Example A.4), or the polynomial is degenerate (see Example A.3).
+    The positive $\epsilon$ might not exist if one of the conditions of the theorem is not satisfied, including when $s=0$ or 11 (see Example 7.1), the Frostman constant $C$ grows with $n$ (see Example 7.4), or the polynomial is degenerate (see Example 7.3).
   3. 3.
-    In the statement of Theorem 1.1, the fact that $N$ depends on $C$ , and the condition that $X$ and $Y$ are independent are both essential (see Examples A.4 and A.5, respectively).
+    In the statement of Theorem 1.2, the fact that NN depends on $C$ , and the condition that $X$ and $Y$ are independent are both essential (see Examples 7.4 and 7.5, respectively).
   4. 4.
-    For any $10<s<$ 1 and any sufficiently large $n\in\mathbb{N}$, there exists an $(s,O(1))$-Frostman random variable $X$ with range contained in $[0,1]$, such that
+    For any $0<s<1$ and any sufficiently large $n\in\mathbb{N}$, there exists an $(s,O(1))$-Frostman random variable $X$ with range contained in $[0,1]$, such that
     $\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}=(1+o(1))H_{n}(X)$
@@
 It would be interesting to explore more connections between $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ and the entropy of other functions.
+It is worth noting that the proof of Theorem 1.2 can be extended to more general settings.
+In this version, we present two results for dependent random variables and special polynomials.
+Further results will be included in a forthcoming update.
+Theorem 1.3.
+Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$ and $C_{1},C_{2},c>0$.
+There exist a positive constant $\epsilon=\epsilon(s)$ and an integer $N=N(\alpha,\beta,s,C_{1},C_{2},c)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$.
+Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
+Then
+$\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\geq n(s+\epsilon),$
+and
+$\max\left\{\frac{H_{n}(X+Y)+H_{n}(X-Y)}{2},\,H_{n}(XY)\right\}\geq n(s+\epsilon)$
+for all $n>N$.
+Both Theorems 1.2 and 1.3 are consequences of the following theorem, which only requires a jointly Frostman condition.
+Theorem 1.4.
+Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$, $C\geq 1$ and $c>0$.
+There exist a positive constant $\epsilon=\epsilon(s)$ and an integer $N=N(\alpha,\beta,s,C,c)$ such that the following statement holds: For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
+$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\geq n(s+\epsilon)$
+for all $n>N$.
+Example 7.6 will illustrate that the first term $\frac{H_{n}(X)+H_{n}(Y)}{2}$ in above formula cannot be removed, and cannot be replaced by $\frac{1}{2}H(X,Y)$ as in Theorem 1.3.
+In the following, we discuss some applications of the entropy theorems.
+Covering number: non-degenerate quadratic polynomials.
 The next theorem presents an application of Theorem 1.2. Here, by $\underline{\dim}_{B}$ and $H^{s}$, we mean the lower box dimension and the $s$-Hausdorff measure, respectively.
@@
 Let $\phi(x,y)$ be a non-degenerate rational quadratic form.
-For $10<s<$ 1, there exists a positive constant $\epsilon=\epsilon(s{\color[rgb]{0,0,0},\phi})$ such that for every Borel set $A\subset{\color[rgb]{0,0,0}[0,1]}$ with $H^{s}(A)>0$, we have
-$\max\left\{\underline{\dim}_{B}(A+A),\underline{\dim}_{B}(\phi(A,A))\right\}\geq s+\epsilon.$
-This theorem is of independent interest, particularly when $\phi(x,y)$ is a non-bilinear form, and might be considered as an application of the Falconer distance problem.
-If $\phi(x,y)=xy$, then stronger results have been obtained by using a variety of techniques, see [4, 10, 11, 16] and references therein for current progress.
-Given this line of research, it is reasonable to expect different proofs of Theorem 1.2 with other methods.
-In the prime field setting, such a result has been proved by using tools from incidence geometry, see [8, Theorem 1.5] and [14, Theorem 1.3], for example.
+For $0<s<1$, there exists a positive constant $\epsilon=\epsilon(s{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},\phi})$ such that for every Borel set $A\subset{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[0,1]}$ with $H^{s}(A)>0$, we have
+$\max\left\{\underline{\dim}_{B}(A+A),\overline{\dim}_{B}(\phi(A,A))\right\}\geq s+\epsilon,$
+and
+$\max\left\{\overline{\dim}_{B}(A+A),\underline{\dim}_{B}(\phi(A,A))\right\}\geq s+\epsilon.$
 For $A\subset\mathbb{R}$ and $\delta>0$, by $\mathcal{E}_{\delta}(A)$ we mean the $\delta$-covering number of $A$. The above theorem can be rewritten in the language of covering numbers.
@@
 Let $\phi(x,y)$ be a non-degenerate rational quadratic form.
-Given $10<s<$ 1 and a constant $C\geq 1$.
-Then there exist $\epsilon=\epsilon(s{\color[rgb]{0,0,0},\phi})>0$ and $\delta_{0}=\delta_{0}(s,C,\phi)>0$ such that the following holds for all $\delta\leq\delta_{0}$.
-Let $A\subset{\color[rgb]{0,0,0}[0,1]}$ be a $\delta$-separated set of cardinality $\delta^{-s}$.
-Suppose that for all intervals $I$of length at least $\delta$, we have the non-concentration estimate
-#$\#(A\cap I)\leq C|I|^{s}(\#A).$
+Given $0<s<1$ and $C\geq 1$.
+Then there exist positive constants $\epsilon=\epsilon(\phi,s)$ such that the following holds for all $\delta>0$.
+Let $A\subset[0,1]$ be any $\delta$-separated set of cardinality $\lceil\delta^{-s}\rceil$.
+Suppose that for all intervals $I$ of length $l$ at least $\delta$, we have the non-concentration estimate
+$(A\cap I)\leq Cl^{s}(\#A).$
 Then we have
-$\mathcal{E}_{\delta}(A+A)+\mathcal{E}_{\delta}(\phi(A,A))\gg C^{-O(1)}\delta^{-\epsilon}(\#A).$
-We now sketch the main ideas in the proof of Theorem 1.1.
+$\mathcal{E}_{\delta}(A+A)+\mathcal{E}_{\delta}(\phi(A,A))\gg_{\phi,s}C^{-O_{\phi,s}(1)}\delta^{-\epsilon}(\#A).$
+These two applications are of independent interest, particularly when $\phi(x,y)$ is a non-bilinear form, and might be considered as an application of the Falconer distance problem.
+If $\phi(x,y)=xy$, then stronger results have been obtained by using a variety of techniques, see [5, 11, 12, 17] and references therein for current progress.
+Given this line of research, it is reasonable to expect different proofs of Theorem 1.5 with other methods.
+In the prime field setting, such a result has been proved by using tools from incidence geometry, see [9, Theorem 1.5] and [15, Theorem 1.3], for example.
+When $\phi(x,y)=x^{2}+y^{2}$, after posting this paper to Arxiv, Tuomas Orponen pointed out that one can use his proof of [22, Corollary 1.3] and recent results in [1] or [23] to obtain a quantitative bound for this special case.
+Fractional covering number: special quadratic polynomials.
+For a set $G\subseteq A\times A$, we denote
+$A+_{G}A=\{x+y:\,(x,y)\in G\},\qquad\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2}=\{\alpha x^{2}+\beta y^{2}:\,(x,y)\in G\}.$
+The next theorem presents an application of Theorem 1.4 on fractional covering numbers.
+Theorem 1.7.
+Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$ and $C\geq 1$.
+Then there exist positive constants $\epsilon=\epsilon(s)$ and $\delta_{0}=\delta_{0}(\alpha,\beta,s,C)$ such that the following holds for all $0<\delta\leq\delta_{0}$.
+Let $A\subset{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[0,1]}$ be any $\delta$-separated set of cardinality $\lceil\delta^{-s}\rceil$.
+Suppose that for all intervals $I$ of length $l$ at least $\delta$, we have the non-concentration estimate
+$(A\cap I)\leq Cl^{s}(\#A).$
+Then, for all $G\subset A\times A$ with $|G|\geq\delta^{\epsilon}|A|^{2}$, then
+$\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2})\gg_{\alpha,\beta,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A).$
+Sketch the main ideas in the proof of Theorem 1.2.
 Our base case is corresponding to the diagonal forms $\phi(X,Y)=\alpha X^{2}+\beta Y^{2}$ (Lemma 4.1), with independent $(s,C)$-Frostman random variables $X,Y$.
 Roughly speaking, by applying the submodularity formula (Lemma 2.17) on certain algebraic relations between the sum function $X+Y$ and the distance function $(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}$, we are able to reduce the problem to bounding the entropy of the distance function from below.
-In particular, let $W$ be an $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $10<s<$ 1, the constant $C\geq 1$ is bounded and does not depend on $n$, and $R>0$.
-Let $W^{\prime}$ be a random variable which is i.i.d with $W$.
+In particular, let $W$ be an $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $0<s<1$, the constant $C\geq 1$ is bounded and does not depend on $n$, and $R>0$.
+Let $W^{\prime}$ be a random variable which is i.i.d.
+with $W$.
 Then there exists a constant $\epsilon_{0}=\epsilon_{0}(s)>0$ such that
@@
 for all $n\geq N_{0}$, where $N_{0}=N_{0}(s,C,R)$.
-Note that if the Frostman constant $C$ depends on $n$, then the inequality (2) might not be true (see Example A.5).
+Note that if the Frostman constant $C$ depends on $n$, then the inequality (3) might not be true (see Example 7.5).
 To prove (3), we make use of results on the Falconer distance problem.
-While classical estimates taken from [12, Section 4.2] are sufficient in the case $s>1/2$, we invoke the recent work of Shmerkin and Wang [20] for $s\leq 1/2$.
-For the general form $\phi(X,Y)$, the proof becomes much more complicated.
-The first step is to transform it to a diagonal form $\tilde{\phi}(\tilde{X},\tilde{Y})$, but the transformed variables $\tilde{X}$ and $\tilde{Y}$ are no longer independent.
-To proceed further, we introduce a novel approach, which can be briefly described as follows.
-By applying the discretized entropy version of Balog-Szemerédi-Gowers theorem (see Theorem 3.1), we obtain new random variables $X^{\prime\prime}$ and $Y^{\prime\prime}$ that are conditionally independent, and thus reduce the problem accordingly.
-Generally speaking, there exists a random variable $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$ such that
-(i) $(X^{\prime},Y^{\prime})\equiv(\tilde{X},\tilde{Y})$;
+While classical estimates taken from [13, Section 4.2] are sufficient in the case $s>1/2$, we invoke the recent work of Shmerkin and Wang [24] for $s\leq 1/2$.
+[…]
+[…]
 (ii) $X^{\prime}$ and $X^{\prime\prime}$ are conditionally i.i.d. given $D_{n}(Y^{\prime})$;
@@
 We address this challenge by explicitly computing the entropy of the distance function. Instead of (3), we have the following for both small and large dimensions:
-Large dimension: Let $W$ be an $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $11/2<s<$ 1, $C\geq 1$ and $R>0$.
+Large dimension: Let $W$ be an $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $1/2<s<1$, $C\geq 1$ and $R>0$.
 Let $W^{\prime}$ be a random variable which is i.i.d with $W$.
-Then for any number $t$ satisfying $t\leq 1$ and $s<t<2s-1/2$, we have that for all $n\in\mathbb{N},$
+Then for any number $t$ satisfying $t\leq 1$ and s<t<2s1/2s<t<2s-1/2, we have that for all $n\in\mathbb{N},$
 $H_{n}(|W-W^{\prime}|)\geq nt-\log C+O(|\log R|)+O(|\log(2s-t-1/2)|)+O_{t}(1).$ (5)
-Small dimension: Let $X^{\prime\prime}$, $Y^{\prime\prime}$ be independent random variables with ranges contained in $[-c,c]$, with $c>0$.
-Suppose that $X^{\prime\prime}$ and $Y^{\prime\prime}$ are $(s,C_{1})$-Frostman and $(s,C_{2})$-Frostman, respectively, where $20<s\leq 1/$ 2 and $C_{1},C_{2}\geq 1$.
+Small dimension: Let $X^{\prime\prime}$, $Y^{\prime\prime}$ be independent random variables with ranges contained in [c,c][-c,c], with $c>0$.
+Suppose that $X^{\prime\prime}$ and $Y^{\prime\prime}$ are $(s,C_{1})$-Frostman and $(s,C_{2})$-Frostman, respectively, where $0<s\leq 1/2$ and $C_{1},C_{2}\geq 1$.
 Let $(X^{\prime\prime\prime},Y^{\prime\prime\prime})$ be i.i.d.
@@
 This is why the estimate (6) holds for $W=X^{\prime\prime}\times Y^{\prime\prime}$ rather than in the general form of (5).
-Plugging these inequalities in our method, the theorem follows.
+Plugging these inequalities into our method, the theorem follows.
 Our final comment is that entropy methods have also been used in a range of problems in Additive Combinatorics and Geometric Measure Theory, for example, see [4, 7, 10, 16, 19, 24, 25, 26] and references therein.
@@
 We denote the range of $X$ by $range(X)$, which collects all the possible values that $X$ can take.
-In this paper, all the measurable spaces will be taken to be subspaces of ${\mathbb{R}}^{m}$, equipped with the Borel $\sigma$-algebra or its sub-$\sigma$-algebra, and the Euclidean norm ||$\cdot$|| ⋅ |.
+In this paper, all the measurable spaces will be taken to be subspaces of ${\mathbb{R}}^{m}$, equipped with the Borel $\sigma$-algebra or its sub-$\sigma$-algebra, and the Euclidean norm $|\cdot|$.
 The dimension $m$ is always regarded as an absolute constant.
@@
 In particular, when $\mu=P$, the conditional probability measure given an event $E$ is denoted by $P_{E}$.
-For a random varaible $X$ and such an $E$, we denote by $X_{E}$ the random variable satisfying $X_{E}\equiv(X|E)$ (see Lemma B.2).
+For a random varaible $X$ and such an $E$, we denote by $X_{E}$ the random variable satisfying $X_{E}\equiv(X|E)$ (see Lemma A.2).
 That is to say,
@@
 In this paper, by $f=O_{\kappa}(g)$, $f\ll_{\kappa}g$ or $g\gg_{\kappa}f$, we mean that there exists an absolute positive constant $C=C(\kappa)$ depending on the parameter(s) $\kappa$ such that $|f|\leq Cg$.
-All logarithms we use here have base $2$.
+All logarithms we use here have base $2$. Moreover, we denote $\log_{+}x:=\max\{\log x,\,0\}$ for any $x>0$.
 2.2 Entropy and basic properties
@@
 $H(\mu,\mathcal{A}|\mathcal{B})=\sum_{B\in\mathcal{B}}\mu(B)H(\mu_{B},\mathcal{A}).$
-In above definition, the convention $00\log 0=$ 0 is used for convenience.
+In above definition, the convention $0\log 0=0$ is used for convenience.
 We also say $\mu(B)H(\mu_{B},\mathcal{A})=0$ when $\mu(B)=0$, although $\mu_{B}$ is not defined in this situation.
@@
 Moreover, the function $\iota(x)=-x\log x$ is non-negative valued, bounded and concave on $[0,1]$.
-It is increasing on $[0,e^{-1}]$, and decreasing on $[e^{-1},1]$.
+It is increasing on $0,e^{-1}]$, and decreasing on $e^{-1},1]$.
 With this in tow, we record some properties of entropy for partitions, whose proofs are simple exercises.
@@
 For a probability measure $\mu$ and countable partitions $\mathcal{A}$, $\mathcal{B}$ and $\mathcal{C}$, we have the following statements:
-(i) $H(\mu,\mathcal{A})\geq 0$ with equality holding if and only if $|Z(\mathcal{A})|=1.$
-(ii) $H(\mu,\mathcal{A})\leq\log|Z(\mathcal{A})|$ when $Z(\mathcal{A})$ is a finite set, with equality holding if and only if $\mu(A)=1/|Z(\mathcal{A})|$ for all $A\in Z(\mathcal{A}).$
+(i) $H(\mu,\mathcal{A})\geq 0$ with equality holding if and only if $Z(\mathcal{A})=1.$
+(ii) $H(\mu,\mathcal{A})\leq\log\#Z(\mathcal{A})$ when $Z(\mathcal{A})$ is a finite set, with equality holding if and only if $\mu(A)=1/\#Z(\mathcal{A})$ for all $A\in Z(\mathcal{A}).$
 (iii) $H(\mu,\mathcal{A})+H(\mu,\mathcal{B}|\mathcal{A})=H(\mu,A\land B).$
@@
 Let $X,Y$ be discrete random variables, and $f$ be a function. The following statements hold.
-(i) $H(X)\geq 0$ with equality holding if and only if $|range_{+}(X)|=1$.
-(ii) $H(X)\leq\log|range_{+}(X)|$ when $range_{+}(X)$ is a finite set, with equality holding if and only if $X$ is uniformly distributed on $range_{+}(X)$.
+(i) $H(X)\geq 0$ with equality holding if and only if $range_{+}(X)=1$.
+(ii) $H(X)\leq\log\#\,range_{+}(X)$ when $range_{+}(X)$ is a finite set, with equality holding if and only if $X$ is uniformly distributed on $range_{+}(X)$.
 (iii) $H(X)+H(Y|X)=H(X,Y)$, and in particular, $H(X)\leq H(X,Y)$.
@@
 $H(X^{\prime},X^{\prime\prime},Y)=2H(X,Y)-H(Y).$
-The following is known as the Plünnecke-Ruzsa inequality. A proof can be found several places, for example, [7, Theorem 3.11].
-Lemma 2.9 (Plünnecke-Ruzsa inequality).
-Let $X,Y_{1},Y_{2},\ldots,Y_{m}$ be independent discrete random variables. Then
-$H(X+Y_{1}+\ldots+Y_{m})-H(X)\leq\sum\limits_{i=1}^{m}\big{(}H(X+Y_{i})-H(X)\big{)},$
-$H(Y_{1}+\ldots+Y_{m})-H(X)\leq\sum\limits_{i=1}^{m}\big{(}H(X+Y_{i})-H(X)\big{)}.$
 The next lemma is helpful in getting an upper bound for certain entropy-terms.
 Lemma 2.9.
-Let $k\geq 1$ and $20<\lambda\leq 1/$ 2. Suppose that $p_{1},p_{2},\ldots,p_{k}$ are non-negative numbers satisfying $\sum\nolimits_{i=1}^{k}p_{i}\leq\lambda$. Then
-$-\sum\limits_{i=1}^{k}p_{i}\log p_{i}\leq\lambda\log k+O(1).$
+Let $k\geq 1$ and $0<\lambda\leq 1/2$. Suppose that $p_{1},p_{2},\ldots,p_{k}$ are non-negative numbers satisfying $\sum\nolimits_{i=1}^{k}p_{i}\leq\lambda$. Then
+$\sum\limits_{i=1}^{k}p_{i}\log p_{i}\leq\lambda\log k+O(1).$
 Proof.
@@
 $f(p_{1},\ldots,p_{k+1})=-\sum\limits_{i=1}^{k+1}p_{i}\log p_{i},$
-which is attainable in the feasible region of $(p_{1},\ldots,p_{k+1})$. For any $11\leq j<i\leq k+$ 1, denote
+which is attainable in the feasible region of $(p_{1},\ldots,p_{k+1})$. For any $1\leq j<i\leq k+1$, denote
 $\tilde{f}(p_{1},\ldots,p_{j-1},p_{j+1},\ldots,p_{k+1})=f(p_{1},\ldots,p_{j-1},1-\sum\nolimits_{l\neq j}p_{l},p_{j+1},\ldots,p_{k+1}).$
@@
 $\frac{\partial\tilde{f}}{\partial p_{i}}(p_{1},\ldots,p_{j-1},p_{j+1},\ldots,p_{k+1})=\log\frac{1-\sum\nolimits_{l\neq j}p_{l}}{p_{i}}=\log\frac{p_{j}}{p_{i}}\leq 0.$
-That is to say, when the other $k-1$ variables are given, the small $p_{i}$ is, the larger value $\tilde{f}$ and $f$ take.
+That is to say, when the other k1k-1 variables are given, the small $p_{i}$ is, the larger value $\tilde{f}$ and $f$ take.
 Now one can verify that the maximum value of $f$ is taken at $(p_{1},\ldots,p_{k+1})$ with
@@
 That is to say,
-$-\sum\limits_{i=1}^{k+1}p_{i}\log p_{i}\leq k\cdot\left(-\frac{\lambda}{k}\log\frac{\lambda}{k}\right)-(1-\lambda)\log(1-\lambda)\leq\lambda\log k+1.$
+$\sum\limits_{i=1}^{k+1}p_{i}\log p_{i}\leq k\cdot\left(-\frac{\lambda}{k}\log\frac{\lambda}{k}\right)-(1-\lambda)\log(1-\lambda)\leq\lambda\log k+1.$
 It follows that
-$-\sum\limits_{i=1}^{k}p_{i}\log p_{i}\leq\lambda\log k+1+p_{k+1}\log p_{k+1}=\lambda\log k+O(1).$
+$\sum\limits_{i=1}^{k}p_{i}\log p_{i}\leq\lambda\log k+1+p_{k+1}\log p_{k+1}=\lambda\log k+O(1).$
 2.3 Discretized entropy and properties
@@
 $H_{n}(X,Y)\leq H_{n}(X)+H_{n}(Y).$ (10)
-Combining Lemma B.2, one can verify that the statement in Lemma 2.5 also holds for $H_{n}$: If $X$ is a random variable and $\mathcal{B}$ is a finite partition. Then
+Combining Lemma A.2, one can verify that the statement in Lemma 2.5 also holds for $H_{n}$: If $X$ is a random variable and $\mathcal{B}$ is a finite partition. Then
 $H_{n}(X)\geq\sum\limits_{B\in\mathcal{B}}P(B)H_{n}(X_{B}).$ (11)
@@
 Example 2.12.
-Let $\alpha,c_{1},c_{2}$ be real numbers with $\alpha\neq 0$ and $0<c_{2}<c_{1}$. Let $f(x)=\alpha x^{2}$ with domain $[-c_{1},-c_{2}]\cup[c_{2},c_{1}]$. Note that, by the mean value theorem,
+Let $\alpha,c_{1},c_{2}$ be real numbers with $\alpha\neq 0$ and $0<c_{2}<c_{1}$. Let $f(x)=\alpha x^{2}$ with domain $c_{1},-c_{2}]\cup[c_{2},c_{1}]$. Note that, by the mean value theorem,
 $2|\alpha|c_{2}\cdot|x-y|\leq|f(x)-f(y)|\leq 2|\alpha|c_{1}\cdot|x-y|$
-[…]
+whenever $x,y\in[-c_{1},-c_{2}]$, or $x,y\in[c_{2},c_{1}]$. One asserts that $f$ is weakly $\max\left\{(2|\alpha|c_{2})^{-1},\,2|\alpha|c_{1}\right\}$-bi-Lipschitz.
 Definition 2.13.
@@
 It follows that
-[…]
+$H_{n}(f(X))\leq H_{n}(f(Y))+\log_{+}\big{(}(\kappa K2^{n})^{m}\big{)}+O_{m}(1).$
 Then the first conclusion follows.
@@
 ∎
+The following lemma says that the $n$-discretized operation makes no harm when we study entropy of sums of random variables.
 Lemma 2.19.
-Let $l\geq 2$ and $X_{1},X_{2},\ldots,X_{l}$ be random variables. Then for any $n\in\mathbb{N}$,
-$H_{n}(X_{1}+X_{2}+\ldots+X_{l})=H(D_{n}(X_{l})+D_{n}(X_{2})+\ldots+D_{n}(X_{l}))+O(\log l).$
+Let $k,l\geq 0$ and $X_{1},\ldots,X_{k}$, $Y_{1},\ldots,Y_{l}$ be real random variables. Then for any $n\in\mathbb{N}$,
+$H_{n}(X_{1}+\ldots+X_{k},\,Y_{1}+\ldots,Y_{l})$
+$\quad=H(D_{n}(X_{l})+\ldots+D_{n}(X_{k}),\,D_{n}(Y_{1})+\ldots+D_{n}(Y_{l}))+O_{k,l}(1).$
 Proof.
-Note that $\sum\nolimits_{i=1}^{l}D_{n}(X_{i})$ takes values in the dyadic rationals. One has
-$H\left(\sum\limits_{i=1}^{l}D_{n}(X_{i})\right)=H\left(D_{n}\Big{(}\sum\limits_{i=1}^{l}D_{n}(X_{i})\Big{)}\right)=H_{n}\left(\sum\limits_{i=1}^{l}D_{n}(X_{i})\right).$
-The conclusion then follows from Lemma 2.19 with the $O(l)$-Lipschitz function (see Example 2.12) given by
-$(x_{1},\ldots,x_{l})\mapsto x_{1}+\ldots+x_{l}.$
-Now we show a discretized version of the Plünnecke-Ruzsa inequality.
+Note that $\sum\nolimits_{i=1}^{l}D_{n}(X_{i})$ and $\sum\nolimits_{j=1}^{l}D_{n}(Y_{j})$ takes values in the dyadic rationals. One has
+$H\left(\sum\limits_{i=1}^{k}D_{n}(X_{i}),\,\sum\limits_{j=1}^{l}D_{n}(Y_{j})\right)$ $=H\left(D_{n}\Big{(}\sum\limits_{i=1}^{l}D_{n}(X_{i})\Big{)},\,D_{n}\Big{(}\sum\limits_{j=1}^{l}D_{n}(Y_{j})\Big{)}\right)$
+$=H_{n}\left(\sum\limits_{i=1}^{l}D_{n}(X_{i}),\,\sum\limits_{j=1}^{l}D_{n}(Y_{j})\right).$
+The conclusion then follows from Lemma 2.18 with the $O_{k,l}(1)$-Lipschitz function given by
+$(x_{1},\ldots,x_{k},y_{1},\ldots,y_{l})\mapsto(x_{1}+\ldots+x_{k},\,y_{1}+\ldots+y_{l}).$
+For independent discrete random variables $X$ and $Y$, one has
+$H(X+Y)\geq H(X+Y|Y)=H(X|Y)=H(X).$
+Applying Lemma 2.19, for any independent random variables $X,Y$ and any $n\in\mathbb{N}$,
+$H_{n}(X+Y)=H(D_{n}(X)+D_{n}(Y))+O(1)\geq H_{n}(X)+O(1).$
+As a consequence,
+$\max\{H_{n}(X),H_{n}(Y)\}\leq H_{n}(X+Y)+O(1)$ (13)
+for all $n\in\mathbb{N}$, provided that $X$ and $Y$ are independent. In the rest of this paper, we may apply Lemma 2.19 without mention it.
+In [16], Ruzsa studied some entropy inequalities involving sums of random variables. In the following, more Plünnecke-Ruzsa type inequalities for discretized entropy are presented.
 Lemma 2.20.
-Let $l\geq 1$ and $X,Y_{1},Y_{2},\ldots,Y_{l}$ be independent random variables. Then
-$H_{n}(X+Y_{1}+Y_{2}+\ldots+Y_{l})-H_{n}(X)\leq\sum\limits_{i=1}^{l}\big{(}H_{n}(X+Y_{i})-H_{n}(X)\big{)}+O(\log l),$
-and
-$H_{n}(Y_{1}+Y_{2}+\ldots+Y_{l})-H_{n}(X)\leq\sum\limits_{i=1}^{l}\big{(}H_{n}(X+Y_{i})-H_{n}(X)\big{)}+O(\log l).$
-Proof.
-Note that $D_{n}(X),D_{n}(Y_{1}),\ldots D_{n}(Y_{l})$ are also independent. The conclusion follows from Lemmas 2.9 and 2.20. ∎
-Corollary 2.22.
-Let $X,Y,$ and $Z$ be independent random variables. Then, for all $n\in\mathbb{N}$,
-$H_{n}(X)$ $\leq H_{n}(X+Y)+O(1),$
+Let $X,Y$ and $Z$ be independent real random variables. Then, for all $n\in\mathbb{N}$,
 $H_{n}(X+Y+Z)$ $\leq H_{n}(X+Z)+H_{n}(Y+Z)-H_{n}(Z)+O(1),$
@@
 Proof.
-We only show the proof of the third inequality involving $H_{n}(X-Y)$ here. The others are direct consequences of Lemma 2.21.
 By Lemma 2.17,
+$H_{n}(X+Y+Z)+H_{n}(X,Y,Z)\leq H_{n}(X+Z,Y)+H_{n}(X,Y+Z)+O(1).$
+Combining Lemma 2.4(iv) and the independence, the first inequality follows. Next, it follows from (13) that
+$H_{n}(X+Y)\leq H_{n}(X+Y+Z)+O(1)\leq H_{n}(X+Z)+H_{n}(Y+Z)-H_{n}(Z)+O(1).$
+Another application of Lemma 2.17 gives
 $H_{n}(X-Y)+H_{n}(X,Y,Z)\leq H_{n}(X+Z,Y+Z)+H_{n}(X,Y)+O(1).$
-By the independence of $X,Y,$ and $Z,$ (specifically, by invoking (8) and (9)) it follows that
-$H_{n}(X-Y)+H_{n}(X,Y)+H_{n}(Z)\leq H_{n}(X+Z)+H_{n}(Y+Z)+H_{n}(X,Y)+O(1).$
-The conclusion follows. ∎
+Combining independence, the second inequality involving XYX-Y then follows. ∎
 Lemma 2.21.
-Let $X$ and $Y$ be independent real random variables and $k,l\in\mathbb{Q}$. There exist non-negative integers $d_{1}=d_{1}(k,l)$ and $d_{2}=d_{2}(k,l)$ such that
-$H_{n}(kX+lY)\leq(d_{1}+d_{2}+1)H_{n}(X+Y)-d_{1}H_{n}(X)-d_{2}H_{n}(Y)+O_{k,l}(1).$
-Here by $kX$ we mean the product of $k$ and the random variable $X$.
+Let $X$ and $Y$ be independent real random variables and $k,l\in\mathbb{Q}$. There exist a non-negative integers $d=d(k,l)$ such that
+$H_{n}(kX+lY)\leq dH_{n}(X+Y)-(d-1)(H_{n}(X)-H_{n}(Y))+O_{k,l}(1)$
+for all $n\in\mathbb{N}$. Here by $kX$ we mean the product of $k$ and the random variable $X$.
 Proof.
+We first prove that
+$H_{n}(kX+lY)\leq(d_{1}+d_{2}+1)H_{n}(X+Y)$
+$\qquad-d_{1}H_{n}(X)-d_{2}H_{n}(Y)+(d_{1}+d_{2}+1)I_{n}(X,Y)+O_{k,l}(1)$ (14)
+for some non-negative integers $d_{1}=d_{1}(k,l)$ and $d_{2}=d_{2}(k,l)$.
 The proof is trivial when $k=l=0$. When $l=0$ and $k\neq 0$, we have
 $H_{n}(kX)=H_{n}(X)+O_{k}(1)\leq H_{n}(X+Y)+O_{k}(1)$
-by Corollary 2.22. Similar conclusion holds when $k=0$. In the following, we assume that $k,l\neq 0$.
+by (13). Similar conclusion holds when $k=0$. In the following, we assume that $k,l\neq 0$.
 First, let us prove the case that $k,l\in\mathbb{Z}^{\ast}$. We apply double induction on $|k|$ and $|l|$. When $|k|=|l|=1$, we claim that
 $H_{n}(kX+lY)\leq 3H_{n}(X+Y)-H_{n}(X)-H_{n}(Y)+O(1).$ (15)
-Indeed, when $k=l=1$, one can deduce (12) from the fact that
-$H_{n}(X+Y)\geq\max\{H_{n}(X),H_{n}(Y)\}+O(1).$
-When $k=1$ and $l=-1$, we take $X,Y,Z$ to be independent random variables such that $X$ and $Z$ are identically distributed. Then Corollary 2.22 gives
+Indeed, when $k=l=1$, one can deduce (15) from (13).
+When $k=1$ and l=1l=-1, we take $X,Y,Z$ to be independent random variables such that $X$ and $Z$ are identically distributed.
+Then Corollary 2.20 gives
 $H_{n}(X-Y)\leq H_{n}(X+Z)+H_{n}(Y+Z)-H_{n}(Z)+O(1).$
@@
 $H_{n}(X+Z)\leq H_{n}(X+Y)+H_{n}(Z+Y)-H_{n}(Y)+O(1).$ (16)
-Then (12) follows. Since $H_{n}(W)=H_{n}(-W)$ for any random variable $W$, the cases that $k=-1$ and $l=\pm 1$ are also satisfied.
+Then (15) follows.
+Since $H_{n}(W)=H_{n}(-W)$ for any random variable $W$, the cases that k=1k=-1 and $l=\pm 1$ are also satisfied.
+Since $H_{n}(W)=H_{n}(-W)$ for any random variable $W$, the cases that k=1k=-1 and $l=\pm 1$ are also satisfied.
 Now let us assume that the conclusion has been proved for some $k$ and $l$. Without loss of generality, we assume that $k>0$. We aim to prove the conclusion for $k+1$ and $l$. Lemma 2.17 gives us that
 $H_{n}((k+1)X+lY+Z)+H_{n}(X,Y,Z)\leq H_{n}(Z+lY,X)+H_{n}(X+Z,kX+lY)+O_{k,l}(1),$
-where the random variable $Z$ is taken such that $X,Y,Z$ are independent and $X$, $Z$ are identically distributed. Combining Corollary 2.22, we obtain that
+where the random variable $Z$ is taken such that $X,Y,Z$ are independent and $X$, $Z$ are identically distributed. Combining (13), we obtain that
 $H_{n}((k+1)X+lY)\leq H_{n}((k+1)X+lY+Z)+O_{k,l}(1)$
@@
 $\leq H_{n}(Z+lY)+H_{n}(X+Z)+H_{n}(kX+lY)-H_{n}(Y)-H_{n}(Z)+O_{k,l}(1).$
-Combining (13) and the inductive hypothesis on $k$, one sees that the conclusion holds with some $d_{1}(k,l)$ and $d_{2}(k,l)$ satisfying
+Combining the fact
+$H_{n}(X+Z)\leq H_{n}(X+Y)+H_{n}(Z+Y)-H_{n}(Y)+O(1)$
+and the inductive hypothesis on $k$, one sees that the conclusion holds with some $d_{1}(k,l)$ and $d_{2}(k,l)$ satisfying
 $d_{1}(k+1,l)=d_{1}(1,l)+d_{1}(k,l)+1,\quad d_{2}(k,l)=d_{2}(1,l)+d_{2}(k,l)+2.$
-By symmetry, we also have the induction on $l$, and the proof is completed for the situation that $k,l\in\mathbb{Z}^{\ast}$.
+By symmetry, we also have the induction on $l$, and the proof of (14) is completed for the situation that $k,l\in\mathbb{Z}^{\ast}$.
 Second, when $k,l\in\mathbb{Q}^{\ast}$, we can write $k=k^{\prime}/t$ and $l=l^{\prime}/t$ for some $k^{\prime},l^{\prime},t\in\mathbb{Z}^{\ast}$. It follows that
@@
 $\leq(d_{1}+d_{2}+1)H_{n}(X+Y)-d_{1}H_{n}(X)-d_{2}H_{n}(Y)+O_{k^{\prime},l^{\prime},t}(1)$
-for some $d_{1}=d_{1}(k^{\prime},l^{\prime})$ and $d_{2}=d_{2}(k^{\prime},l^{\prime})$. The lemma then follows. ∎
+for some $d_{1}=d_{1}(k^{\prime},l^{\prime})$ and $d_{2}=d_{2}(k^{\prime},l^{\prime})$. Then (14) follows.
+Next, in view of (15), we can add multiple scalars of $H_{n}(X+Y)-H_{n}(X)$ and $H_{n}(X+Y)-H_{n}(Y)$ on the right-hand side of (14) to get
+$H_{n}(kX+lY)\leq(2d_{1}+2d_{2}+1)H_{n}(X+Y)-(2d_{1}+2d_{2})\frac{H_{n}(X)+H_{n}(Y)}{2}+O_{k,l}(1).$
+The lemma then follows by taking $d=2d_{1}+2d_{2}+1$. ∎
 Lemma 2.21 implies the following statement: For given $k,l\in\mathbb{Q}$, one has
@@
 However, when $k,l$ are $\mathbb{Q}$-linearly independent, such statement may fail.
-In Example A.6, we would show the following: For any $K>0$, there exists some $N\in\mathbb{N}$ such that: for any $n>N$, there is a random variable $X$ with range contained in $[0,1]$ satisfying
+In Example 7.8, we would show the following: For any $K>0$, there exists some $N\in\mathbb{N}$ such that: for any $n>N$, there is a random variable $X$ with range contained in $[0,1]$ satisfying
 $H_{n}(X+\sqrt{2}Y)-H_{n}(X)\geq K(H_{n}(X+Y)-H_{n}(X)),$
@@
 2.4 Frostman measures
-Now we introduce Frostman measures and their basic properties.
-Definition 2.24.
-Let $s,C>0$. We say a Borel measure $\mu$ on $\mathbb{R}^{m}$ is $(s,C)$-Frostman if for any $x\in\mathbb{R}^{m}$ and $r>0$
+Now we recall the definitions of Frostman measures and show some of their properties.
+Definition. Let $s\geq 0$ and $C>0$. We say that a Borel measure $\mu$ on $\mathbb{R}^{m}$ is $(s,C)$-Frostman if for any $x\in\mathbb{R}^{m}$ and $r>0$
 $\mu(B(x,r))\leq Cr^{s}.$
@@
 For a random variable or a probability measure, when $0\leq s_{2}\leq s_{1}\leq 1$ and $1\leq C_{1}\leq C_{2}$, the $(s_{1},C_{1})$-Frostman condition can imply the $(s_{2},C_{2})$-Frostman condition.
+Definition.Let $s,s_{1},s_{2}\geq 0$ and $C,C_{1},C_{2}>0$.
+(i) We say that the random variable $(X,Y)$ is jointly $(s_{1},s_{2},C)$-Frostman, if
+$P(X\in B(x,r_{1}),\,Y\in B(y,r_{2}))\leq Cr_{1}^{s_{1}}r_{2}^{s_{2}}$
+for any ball $B(x,r_{1})$ and $B(y,r_{2})$.
+(ii) We say that the random variable $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, if $X$ and $Y$ are $(s_{1},C_{1})$-Frostman and $(s_{2},C_{2})$-Frostman, respectively.
+(iii) We say a random variable $X$ is $(s,C)$-Frostman given $Y$, if
+$P(X\in B(x,r),\,Y\in E)\leq Cr^{s}P(Y\in E)$
+for any ball $B(x,r)$ and Borel set $E$.
+Moreover, We say that the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, if $X$ is $(s_{1},C_{1})$-Frostman given $Y$, and $Y$ is $(s_{2},C_{2})$-Frostman given $X$.
+The following properties follows directly from the definitions, whose proofs are left as an exercise.
+Lemma 2.22.
+Let $s_{1},s_{2}\geq 0$ and $C,C_{1},C_{2},R>0$.
+(i) Suppose that $X$ and $Y$ are independent random variables such that $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman. Then $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman.
+(ii) Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman. Then it is both marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, and jointly $(s_{1},s_{2},C_{1}C_{2})$-Frostman.
+(iii) Suppose that $(X,Y)$ is jointly $(s_{1},s_{2},C)$-Frostman. Then it is also $(s_{1}+s_{2},C)$-Frostman.
+(iv) Suppose that $(X,Y)$ is a jointly $(s_{1},s_{2},C)$-Frostman random variable whose range is contained in $B(0,R)$, where $R>0$.
+Then $(X,Y)$ is marginally $(s_{1},CR^{s_{2}};s_{2},CR^{s_{1}})$-Frostman.
 The next lemma presents the Frostman conditions related to the arithmetic of random variables.
 Lemma 2.23.
-Let $X$ and $Y$ be independent real random variables that are both $(s,C)$-Frostman, where $0<s\leq 1$ and $C>0$. Then the following statements hold for $a,b,c,d\in\mathbb{R}$.
-(i) The random variable $aX+bY$ is $(s,O_{a,b}(C))$-Frostman.
-(ii) For any invertible linear change of variables $(x_{1},y_{1})=(ax+by,cx+dy)$, the random variable $(X_{1},Y_{1})=(aX+bY,cX+dY)$ is $(2s,O_{a,b,c,d}(C^{2}))$-Frostman.
+Let $(X,Y)$ be a bivariate real random variable that is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, where $0\leq s_{1},s_{2}\leq 1$ and $C_{1},C_{2}>0$.
+Then the following statements hold for all $a,b,c,d\in\mathbb{R}$.
+(i) The random variable $aX+bY$ is $(s_{1},C_{1}^{\prime})$-Frostman when $a\neq 0$, and is $(s_{2},C_{2}^{\prime})$-Frostman when $b\neq 0$.
+Here $C_{1}^{\prime},C_{2}^{\prime}$ are positive numbers satisfying $\log C_{i}^{\prime}=\log C_{i}+O_{a,b}(1)$ $(i=1,2)$.
+(ii) Suppose that $ad-bc\neq 0$.
+Then (aX+bY,cX+dY)(aX+bY,\,cX+dY) is jointly $\big{(}s_{1},s_{2},C^{\prime\prime}\big{)}$-Frostman when $ad\neq 0$, and is jointly $\big{(}s_{2},s_{1},C^{\prime\prime}\big{)}$-Frostman when $bc\neq 0$.
+Here $C^{\prime\prime}$ is a positive number satisfying $\log C^{\prime\prime}=\log C_{1}C_{2}+O_{a,b,c,d}(1)$.
 Proof.
-Let $\mu_{X}$, and $\mu_{Y}$ be the distribution measures associated to $X$ and $Y$, respectively.
-(i) When $b\neq 0$, for any ball $B(y^{\prime},r)$ one has
-[…]
-A similar conclusion follows when $a\neq 0$. When $a=b=0$, the conclusion follows automatically.
-(ii) For any ball $B((x,y),r)$,
-$P((aX+bY,cX+dY)\in B((x,y),r)$ $\leq P(aX+bY\in B(x,r),cX+dY\in B(y,r)).$
-If both $aX+bY\in B(x,r)$ and $cX+dY\in B(y,r)$ occur, we must have $X\in B\left(\frac{dx-by}{ad-bc},\frac{(|d|+|b|)r}{|ad-bc|}\right)$ and $Y\in B\left(\frac{-cx+ay}{ad-bc},\frac{(|c|+|a|)r}{|ad-bc|}\right)$.
-Therefore,
-$P(aX+bY\in B(x,r),cX+dY\in B(y,r))\leq C\left(\frac{(|d|+|b|)r}{|ad-bc|}\right)^{s}\cdot C\left(\frac{(|c|+|a|)r}{|ad-bc|}\right)^{s}=O_{a,b,c,d}(C^{2}r^{2s}).$
-The lemma now follows. ∎
-The next property is taken from [11, Lemma 4.8].
+By Lemma 2.22, the random variable $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman.
+(i) Consider any ball $B(z,r)$. Let $\{I_{i}\}$ be a partition of $\mathbb{R}$} composed of intervals of length $r$. Then
+$P(aX+bY\in B(z,r))=\sum\limits_{i}P(aX+bY\in B(z,r),\,X\in I_{i}):=\sum\limits_{i}q_{i}.$
+Assume that $b\neq 0$. If $q_{i}>0$, then $Y$ is located in an interval $J_{i}$ of length at most $\frac{2(|a|+1)r}{|b|}$. So
+$q_{i}\leq P(X\in I_{i},\,Y\in J_{i})\leq C_{2}\Big{(}\frac{(|a|+2)r}{|b|}\Big{)}^{s_{2}}P(X\in I_{i}).$
+It follows that
+$P(aX+bY\in B(z,r))\leq\sum\limits_{i}C_{2}\Big{(}\frac{(|a|+2)r}{|b|}\Big{)}^{s_{2}}P(X\in I_{i})=C_{2}\Big{(}\frac{(|a|+2)r}{|b|}\Big{)}^{s_{2}}.$
+Hence $aX+bY$ is $(s_{2},C^{\prime}_{2})$-Frostman with $\log C^{\prime}_{2}=\log C_{2}+O_{a,b}(1)$.
+By symmetry, it is also $(s_{1},C^{\prime}_{1})$-Frostman with $\log C_{1}^{\prime}=\log C_{1}+O_{a,b}(1)$, provided that $a\neq 0$.
+(ii) Write $p:=P\big{(}aX+bY\in B(z_{1},r_{1}),\,cX+dY\in B(z_{2},r_{2})\big{)}$ for an arbitrary pair of balls $B(z_{1},r_{1})$ and $B(z_{2},r_{2})$.
+Calculation shows that, if $p>0$, then we must have
+$X\in B\Big{(}\frac{dz_{1}-bz_{2}}{ad-bc},\,\frac{|d|r_{1}+|b|r_{2}}{|ad-bc|}\Big{)},\qquad Y\in B\Big{(}\frac{-cz_{1}+az_{2}}{ad-bc},\,\frac{|c|r_{1}+|a|r_{2}}{|ad-bc|}\Big{)}.$ (17)
+First, let us consider the case $ad\neq 0$. When $r_{1}\leq r_{2}$, let $\{J_{i}\}$ be a partition of $\mathbb{R}$} composed of intervals of length $r_{1}$. We estimate $p$ by
+$p$ $=\sum\limits_{i}P\big{(}aX+bY\in B(z_{1},r_{1}),\,cX+dY\in B(z_{2},r_{2}),\,Y\in J_{i}\big{)}$
+$\leq\sum\limits_{i}P\big{(}aX+bY\in B(z_{1},r_{1}),\,Y\in J_{i}\big{)}:=\sum\limits_{i}p_{i}.$
+One can check that, if $p_{i}>0$, then $X$ is located in an interval $I_{i}$ of length at most $\frac{2(|b|+1)r_{1}}{|a|}$. Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, we have
+$p_{i}\leq P(X\in I_{i},Y\in J_{i})\leq C_{1}\Big{(}\frac{(|b|+2)r_{1}}{|a|}\Big{)}^{s_{1}}P(Y\in J_{i}).$
+Combining (17), it follows that
+$p$ $\leq\sum\limits_{i}C_{1}\Big{(}\frac{(|b|+2)r_{1}}{|a|}\Big{)}^{s_{1}}P(Y\in J_{i})$
+$\leq C_{1}\Big{(}\frac{(|b|+2)r_{1}}{|a|}\Big{)}^{s_{1}}P\left(Y\in B\Big{(}\frac{-cz_{1}+az_{2}}{ad-bc},\,\frac{(|c|+|a|)r_{2}}{|ad-bc|}\Big{)}\right)$
+$\leq C_{1}\Big{(}\frac{(|b|+2)r_{1}}{|a|}\Big{)}^{s_{1}}\cdot C_{2}\Big{(}\frac{(|c|+|a|)r_{2}}{|ad-bc|}\Big{)}^{s_{2}}.$
+When $r_{1}>r_{2}$, let $\{I_{i}^{\prime}\}$ be a partition of $\mathbb{R}$} composed of intervals of length $r_{2}$. We estimate $p$ by
+$p\leq\sum\limits_{i}P\big{(}cX+dY\in B(z_{2},r_{2}),\,X\in I_{i}^{\prime}\big{)}:=\sum\limits_{i}p_{i}^{\prime}.$
+We apply similar arguments as above. For non-zero terms, the value of $Y$ locates in an interval $J_{i}^{\prime}$ of length at most $\frac{2(|c|+1)r_{2}}{|d|}$. So
+$p_{i}^{\prime}\leq P(X_{i}\in I_{i}^{\prime},Y\in J_{i}^{\prime})\leq C_{1}\Big{(}\frac{(|c|+2)r_{2}}{|d|}\Big{)}^{s_{2}}P(X\in I_{i}^{\prime}),$
+and then
+$p$ $\leq C_{1}\Big{(}\frac{(|c|+2)r_{2}}{|d|}\Big{)}^{s_{2}}P\left(X\in B\Big{(}\frac{dz_{1}-bz_{2}}{ad-bc},\,\frac{(|d|+|b|)r_{1}}{|ad-bc|}\Big{)}\right)$
+$\leq C_{1}\Big{(}\frac{(|c|+2)r_{2}}{|d|}\Big{)}^{s_{2}}\cdot C_{1}\Big{(}\frac{dz_{1}-bz_{2}}{ad-bc},\,\frac{(|d|+|b|)r_{1}}{|ad-bc|}\Big{)}^{s_{1}}.$
+We conclude that $(X,Y)$ is jointly $(s_{1},s_{2},C^{\prime\prime})$-Frostman with $\log C^{\prime\prime}=\log C_{1}C_{2}+O_{a,b,c,d}(1)$.
+Second, when $ad=0$, then $bc\neq 0$ since $ad-bc\neq 0$. Similar arguments show that $(X,Y)$ is jointly $(s_{2},s_{1},C^{\prime\prime})$-Frostman. The proof is completed. ∎
+The following lemma shows relations of the discretized entropies and the Frostman conditions.
 Lemma 2.24.
-If $X$ is $(s,C)$-Frostman for some $s,C>0$, then $H_{n}(X)\geq ns-\log C+O(1)$.
+Let $s,s_{1},s_{2}\geq 0$, $C,C_{1},C_{2}>0$ and $X,Y,Z$ be random variables. The following statements hold for all $n\in\mathbb{N}$.
+(i) If $Z$ is $(s,C)$-Frostman, then
+$H_{n}(Z)\geq ns-\log C+O(1).$
+(ii) If $X$ is conditionally $(s,C)$-Frostman given $Y$, then
+$H_{n}(X|Y)\geq ns-\log C+O(1).$
+Moreover, if $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, then
+$H_{n}(X,Y)\geq\frac{1}{2}\big{(}H_{n}(X)+H_{n}(Y)+n(s_{1}+s_{2})-\log C_{1}C_{2}\big{)}+O(1).$
+Proof.
+(i) is taken from [12, Lemma 4.8].
+(ii) Since $X$ is conditionally $(s,C)$-Frostman given $Y$, then
+$H_{n}(X,Y)$ $=-\sum\limits_{x,y}P(D_{n}(X)=x,\,D_{n}(Y)=y)\log P(D_{n}(X)=x,\,D_{n}(Y)=y)$
+$\geq-\sum\limits_{x,y}P(D_{n}(X)=x,\,D_{n}(Y)=y)\log\Big{(}O(C2^{-ns})\cdot P(D_{n}(Y)=y)\Big{)}$
+$=\sum\limits_{x,y}P(D_{n}(X)=x,\,D_{n}(Y)=y)\cdot(ns-\log C+O(1))$
+$\qquad\qquad-\sum\limits_{y}\big{(}\sum\limits_{x}P(D_{n}(X)=x,\,D_{n}(Y)=y)\big{)}\log P(D_{n}(Y)=y)$
+$=ns-\log C+O(1)+H_{n}(Y),$
+i.e., $H_{n}(X|Y)\geq ns-\log C+O(1)$.
+For the conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman random variable $(X,Y)$, we have both
+$H_{n}(X,Y)-H_{n}(Y)\geq ns_{1}-\log C_{1}+O(1),\quad H_{n}(X,Y)-H_{n}(X)\geq ns_{2}-\log C_{2}+O(1).$
+The conclusion then follows.
 The function $x\mapsto x^{2}$ is weakly-bi-Lipschitz in a region that is bounded away from zero.
@@
 Lemma 2.25.
-Let $Z$ be a real random variable with bounded range.
-Assume that there is some $s>0$ and $C\geq 1$ such that $P(|Z|<r)\leq Cr^{s}$ for any $r>0$.
-For any $0<\eta\leq 2^{-1}(2C)^{-1/s}$, denote by $E$ the event that $|Z|\geq\eta$.
-Then the random variable $Z_{E}$ satisfies
-$H_{n}(Z_{E})=\frac{H_{n}(Z)}{P(E)}+O(C(2\eta)^{s}n)+O(1).$
+Let $m\geq 1$.
+Suppose that $Z$ is a random variable with bounded range in ${\mathbb{R}}^{m}$, and $f$ is a function such that $domain(f)\supseteq range_{+}(Z)$.
+Assume that there are numbers $s,C>0$ such that $P(|f(Z)|<r)\leq Cr^{s}$ for any $r>0$.
+Let $0<\eta\leq\min\{2^{-1}(2C)^{-1/s},\,1/2\}$, and denote by $E$ the event that $|Z|\geq\eta$.
+Assume that $P(E)\gg 1$.
+Then the random variable $f(Z)_{E}$ satisfies
+$H_{n}(f(Z)_{E})=\frac{H_{n}(f(Z))}{P(E)}+O(C(2\eta)^{s}n)+O_{m}(1).$
 for any all $n\in\mathbb{N}$.
 Proof.
-Recall that $Z_{E}$ is the random variable such that $Z_{E}\equiv(Z|E)$, and $P_{E}$ is the conditional probability measure given $E$.
-Since $P(|Z|<\eta)\leq C\eta^{s}\leq 1/2$, we have $P(E)\geq 1-C\eta^{s}\geq 1/2$.
-Assume that $k_{0}$ is a natural number with $(k_{0}-1)2^{-n}\leq\eta<k_{0}2^{-n}$.
-Combining the conditions in this lemma, one deduces that
-$H_{n}(Z_{E})$ $=-\sum\limits_{|k|\geq k_{0}+1}P_{E}(D_{n}(Z_{E})=k2^{-n})\log P_{E}(D_{n}(Z_{E})=k2^{-n})$
-$\qquad+O\left(-\sum\limits_{k_{0}-1\leq|k|\leq k_{0}}P_{E}(D_{n}(Z_{E})=k2^{-n})\log P_{E}(D_{n}(Z_{E})=k2^{-n})\right)$
-$=-\sum\limits_{|k|\geq k_{0}+1}\frac{P(D_{n}(Z)=k2^{-n})}{P(E)}\big{(}\log P(D_{n}(Z)=k2^{-n})-\log P(E)\big{)}+O(1)$
-$=-\frac{1}{P(E)}\sum\limits_{|k|\geq k_{0}+1}P(D_{n}(Z)=k2^{-n})\log P(D_{n}(Z)=k2^{-n})+O(1).$
-First, let us consider the case that $\eta>2^{-n+1}$, which gives $(k_{0}+1)2^{-n}\leq\eta+2^{-n+1}<2\eta$. Applying Lemma 2.10 with the fact
-$\sum\nolimits_{|k|\leq k_{0}}P(D_{n}(Z)=k2^{-n})\leq P(|Z|<(k_{0}+1)2^{-n})\leq P(|Z|<2\eta)\leq C(2\eta)^{s}\leq 1/2,$
+Recall that $f(Z)_{E}$ is the random variable such that $f(Z)_{E}\equiv(f(Z)|E)$, and $P_{E}$ is the conditional probability measure given $E$ (see Lemma A.2).
+Assume that $k_{0}$ is a natural number with $(k_{0}-1)2^{-n}\leq\eta<k_{0}2^{-n}$. Combining the conditions in this lemma, one deduces that
+$H_{n}(f(Z)_{E})$ $=-\sum\limits_{k\in\mathbb{Z}^{m}\atop|k|\geq k_{0}+m}P_{E}(D_{n}(f(Z)_{E})=k2^{-n})\log P_{E}(D_{n}(f(Z)_{E})=k2^{-n})$
+$\qquad+O\left(-\sum\limits_{k_{0}-m\leq|k|<k_{0}+m}P_{E}(D_{n}(f(Z)_{E})=k2^{-n})\log P_{E}(D_{n}(f(Z)_{E})=k2^{-n})\right)$
+$=-\sum\limits_{|k|\geq k_{0}+m}\frac{P(D_{n}(f(Z))=k2^{-n})}{P(E)}\big{(}\log P(D_{n}(f(Z))=k2^{-n})-\log P(E)\big{)}+O_{m}(1)$
+$=-\frac{1}{P(E)}\sum\limits_{|k|\geq k_{0}+m}P(D_{n}(f(Z))=k2^{-n})\log P(D_{n}(f(Z))=k2^{-n})+O_{m}(1).$
+First, let us consider the case that $\eta>m2^{-n+2}$, which gives $(k_{0}+2m)2^{-n}\leq\eta+(2m+1)2^{-n}<2\eta$. Applying Lemma 2.9 with the fact
+$\sum\nolimits_{|k|<k_{0}+m}P(D_{n}(f(Z))=k2^{-n})\leq P(|f(Z)|<(k_{0}+2m)2^{-n})\leq P(|f(Z)|<2\eta)\leq C(2\eta)^{s}\leq 1/2,$
 we see that
-$-\sum\limits_{|k|\leq k_{0}}P(D_{n}(Z)=k2^{-n})\log P(D_{n}(Z)=k2^{-n})$
-$\quad\quad\quad\leq C(2\eta)^{s}\cdot\log\big{(}2\eta 2^{n}+3\big{)}+O(1)\leq C(2\eta)^{s}n+O(1).$
-Second, we consider the case that $\eta\leq 2^{-n+1}$. Then $k_{0}\leq 3$ and
-$-\sum\nolimits_{|k|\leq k_{0}}P(D_{n}(Z)=k2^{-n})\log P(D_{n}(Z)=k2^{-n})=O(1).$
+$-\sum\limits_{|k|<k_{0}+m}P(D_{n}(f(Z))=k2^{-n})\log P(D_{n}(f(Z))=k2^{-n})$
+$\quad\quad\quad\leq C(2\eta)^{s}\cdot\log\big{(}2\eta 2^{n}+2m+3\big{)}+O(1)\leq C(2\eta)^{s}n+O_{m}(1).$
+Second, we consider the case that $\eta\leq m2^{-n+2}$. Then $k_{0}\leq 4m+1$ and
+$\sum\nolimits_{|k|<k_{0}+m}P(D_{n}(f(Z))=k2^{-n})\log P(D_{n}(f(Z))=k2^{-n})=O_{m}(1).$
 Finally, we conclude that
-$H(Z_{E})$ $=\frac{1}{P(E)}\left(H_{n}(Z)+\sum\limits_{|k|\leq k_{0}}P(D_{n}(Z)=k2^{-n})\log P(D_{n}(Z)=k2^{-n})\right)+O(1)$
-$=\frac{H_{n}(Z)}{P(E)}+O(C(2\eta)^{s}n)+O(1).$
+$H(f(Z)_{E})$ $=\frac{1}{P(E)}\left(H_{n}(f(Z))+\sum\limits_{|k|<k_{0}+m}P(D_{n}(f(Z))=k2^{-n})\log P(D_{n}(f(Z))=k2^{-n})\right)+O_{m}(1)$
+$=\frac{H_{n}(f(Z))}{P(E)}+O(C(2\eta)^{s}n)+O_{m}(1).$
 In the following lemma, we show that the probability that the distance function takes small values is related to the Frostman condition.
 Lemma 2.26.
-Let $Z$, $Z^{\prime}$ be independent random variables. Suppose that $Z$ is $(s,C)$-Frostman, where $s,C>0$. Then
+Let $Z$, $Z^{\prime}$ be independent random variables. Suppose that $Z$ is $(s,C)$-Frostman, where $s\geq 0$ and $C>0$. Then
 $P\big{(}|Z-Z^{\prime}|<r)\leq Cr^{s}.$
@@
 Denote by $\mu$ the distribution measure associated to $Z^{\prime}$. Since $Z$ is $(s,C)$-Frostman, one sees that
-[…]
+$P(|Z-Z^{\prime}|<r)=\int P(|Z-z^{\prime}|<r)d\mu(z^{\prime})\leq\int Cr^{s}d\mu(z^{\prime})=Cr^{s}.$
+2.5 Entropy of the distance function
 The following definition of Riesz energy is related to the distance measure.
@@
 For any given measure $\mu$, define the measure $\Delta(\mu)$ by
-[…]
-for any Borel set $B$. That is to say, $\Delta(\mu)$ is the image of $\mu\times\mu$ under the distance map $(x,y)\mapsto|x-y|$. We now recall a result from Mattila in [12, Section 4.2, page 61].
+$\Delta(\mu)(B)=\int\mu(\{y\in\Omega:|x-y|\in B\})d\mu(x)$
+for any Borel set $B$. That is to say, $\Delta(\mu)$ is the image of $\mu\times\mu$ under the distance map $(x,y)\mapsto|x-y|$. We now recall a result from Mattila in [13, Section 4.2, page 61].
 Lemma 2.28 ([13]).
@@
 Lemma 2.29.
-Let $\mu$ be a $(2s,C)$-Frostman probability measure with $\operatorname{spt}(\mu)\subset B(0,R)\subset\mathbb{R}^{2}$, where $\frac{1}{4}<s\leq 1$, $C\geq 1$ and $R>0$.
+Let $\mu$ be a $(2s,C)$-Frostman probability measure with $\operatorname{spt}(\mu)\subset B(0,R)\subset\mathbb{R}^{2}$, where $\frac{1}{4}<s\leq 1$ and $C,R>0$.
 Then for any $r,\eta,t$ with $0<\eta<r<2R$, $0<t\leq 1$ and $t<2s-\frac{1}{2}$, it satisfies that
-$\Delta(\mu)([r,r+\eta])=O_{t}\left(\frac{CR^{2s-t}\eta^{t}}{2s-t-1/2}\right)+O_{t}(R^{-t}\eta^{t}).$
+$\Delta(\mu)([r,r+\eta])=O_{t}\left(\Big{(}\frac{C\cdot\max\{R^{2s-t},1\}}{2s-t-1/2}+1\Big{)}\eta^{t}\right).$
 Proof.
-Since $\mu$ is $(2s,C)$-Frostman, for $20<t<2s-1/$ 2 and any $x\in\mathbb{R}^{2}$,
-[…]
-[…]
-$\qquad=O_{t}\left(\frac{CR^{2s-t-1/2}}{2s-t-1/2}\right)+O_{t}(R^{-t-1/2}):=M.$
+Put $\tilde{R}=\max\{R,1\}$. Since $\mu$ is $(2s,C)$-Frostman, for 0<t<2s1/20<t<2s-1/2 and any $x\in\mathbb{R}^{2}$,
+$\int|x-y|^{-t-1/2}d\mu(y)=(t+1/2)\int_{0}^{\infty}\frac{\mu(B(x,\ell))}{\ell^{t+3/2}}d\ell$
+$\qquad\leq(t+1/2)\left(\int_{0}^{2\tilde{R}}C\ell^{2s-t-3/2}d\ell+\int_{2\tilde{R}}^{\infty}\ell^{-t-3/2}d\ell\right)$
+$\qquad=O_{t}\left(\frac{C\tilde{R}^{2s-t-1/2}}{2s-t-1/2}\right)+O_{t}(1):=M.$
 Thus
-[…]
+$I_{t+1/2}(\mu)=\int\int|x-y|^{-t-1/2}d\mu(x)d\mu(y)\leq\int Md\mu(x)=M.$
 When $0<\eta<r<2R$ and $0<t\leq 1$, one obtains by Lemma 2.28 that
-$\Delta(\mu)([r,r+\eta])\ll_{t}r^{1/2}\eta^{t}I_{t+1/2}(\mu)=O_{t}\left(\frac{CR^{2s-t}\eta^{t}}{2s-t-1/2}\right)+O_{t}(R^{-t}\eta^{t}).$
-2.5 Entropy of the distance function
+$\Delta(\mu)([r,r+\eta])\ll_{t}r^{1/2}\eta^{t}I_{t+1/2}(\mu)=O_{t}\left(\frac{C\tilde{R}^{2s-t}\eta^{t}}{2s-t-1/2}\right)+O_{t}(\eta^{t}).$
+The proof is completed. ∎
 In this subsection, we show some results of the discretized entropy of the distance function of certain measures, together with the square of the distance.
@@
 Theorem 2.30.
-Let $W$ be a $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $11/2<s<$ 1, $C\geq 1$ and $R>0$.
+Let $W$ be a $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $1/2<s<1$ and $C,R>0$.
 Let $W^{\prime}$ be a random variable which is i.i.d with $W$.
-Then for any number $t$ satisfying $t\leq 1$ and $s<t<2s-1/2$, we have
-$H_{n}(|W-W^{\prime}|)\geq nt-\log C+O(|\log R|)+O(|\log(2s-t-1/2)|)+O_{t}(1).$ (14)
+Then for any number $t$ satisfying $t\leq 1$ and s<t<2s1/2s<t<2s-1/2, we have
+$H_{n}(|W-W^{\prime}|)\geq nt-\log_{+}C+O(\log_{+}R)+O(|\log(2s-t-1/2)|)+O_{t}(1).$ (18)
 for all $n\in\mathbb{N}$.
@@
 When $4\leq k\leq N$, it follows from Lemma 2.29 that
-$p_{k}$ $=\Delta(\mu)\big{(}[k2^{-n},(k+1)2^{-n}\big{)}=O_{t}\left(\frac{CR^{2s-t}2^{-nt}}{2s-t-1/2}\right)+O_{t}(R^{-t}2^{-nt}).$
+$p_{k}$ $=\Delta(\mu)\big{(}[k2^{-n},(k+1)2^{-n}\big{)}=O_{t}\left(\Big{(}\frac{C\cdot\max\{R^{2s-t},1\}}{2s-t-1/2}+1\Big{)}2^{-nt}\right).$
 It follows that
@@
 $\geq\sum\limits_{k=0}^{3}p_{k}(nt-\log C+O(1))$
-$\qquad+\sum\limits_{k=4}^{N}p_{k}\big{(}nt-\log C+O(|\log R|)+O(|\log(2s-t-1/2)|)+O_{t}(1)\big{)}$
-$=nt-\log C+O(|\log R|)+O(|\log(2s-t-1/2)|)+O_{t}(1).$
+$\qquad+\sum\limits_{k=4}^{N}p_{k}\big{(}nt-\log_{+}C+O(\log_{+}R)+O(|\log(2s-t-1/2)|)+O_{t}(1)\big{)}$
+$\geq nt-\log_{+}C+O(\log_{+}R)+O(|\log(2s-t-1/2)|)+O_{t}(1).$
 Remark 2.31.
 The parameters $C$ and $R$ in Theorem 2.30 are allowed to be unbounded as $n$ tends to infinity.
-In Example A.5, we show an example in which $C$ is growing with $n$.
-This example will tell us that the term $-\log C$ on the right-hand side of (14) is essential.
+In Example 7.5, we show an example in which $C$ is growing with $n$.
+This example will tell us that the term $\log C$ on the right-hand side of (18) is essential.
 Remark 2.32.
-When $13/4<s<$ 1, we can take $t=1$ in Theorem 2.32. Combining this with (11), we then obtain the identity
-$H_{n}(|W-W^{\prime}|)=n+O(\log C)+O(|\log R|)+O_{s}(1).$
+When $3/4<s<1$, we can take $t=1$ in Theorem 2.30. Combining this with (12), we then obtain the identity
+$H_{n}(|W-W^{\prime}|)=n+O(\log_{+}C)+O(\log_{+}R)+O_{s}(1).$
 Theorem 2.33.
-Let $W$ be a $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $11/2<s<$ 1, $C\geq 1$ and $R>0$.
+Let $W$ be a $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $1/2<s<1$ and $C,R>0$.
 Let $W^{\prime}$ be a random variable which is i.i.d with $W$.
 Then there exists a constant $\epsilon_{0}=\epsilon_{0}(s)>0$ such that
-$H_{n}(|W-W^{\prime}|^{2})\geq n(s+\epsilon_{0})+O(\log C)+O(|\log R|)+O_{s}(1)$
+$H_{n}(|W-W^{\prime}|^{2})\geq n(s+\epsilon_{0})+O(\log_{+}C)+O(\log_{+}R)+O_{s}(1)$
 for all $n\in\mathbb{N}$.
@@
 Let $Z$, $t$ be as in Theorem 2.30, with $Z=|W-W^{\prime}|$ and
-$H_{n}(Z)\geq nt-\log C+O(|\log R|)+O(|\log(2s-t-1/2)|)+O_{t}(1).$
+$H_{n}(Z)\geq nt+O(\log_{+}C)+O(\log_{+}R)+O(|\log(2s-t-1/2)|)+O_{t}(1).$
 By Lemma 2.26, one has $P(|Z|<r)\leq Cr^{2s}$ for any $r>0$.
-Let $\eta$ be a parameter with $0<\eta\leq 2^{-1}(2C)^{-1/2s}$ to be determined later.
+Let $\eta$ be a parameter with $0<\eta<\min\{2^{-1}(2C)^{-1/2s},\,1/2\}$ to be determined later.
 Denote by $E$ the event that $|Z|\geq\eta$.
@@
 $H_{n}(Z_{E})$ $=\frac{H_{n}(Z)}{P(E)}+O(C(2\eta)^{2s}n)+O(1)$
-$\geq\frac{nt}{P(E)}+O(\log C)+O(|\log R|)+O(|\log(2s-t-1/2)|)+O_{t}(1)+O(C\eta^{2s}n).$
+$\geq\frac{nt}{P(E)}+O(\log_{+}C)+O(\log_{+}R)+O(|\log(2s-t-1/2)|)+O_{t}(1)+O(C\eta^{2s}n).$
 Since $Z_{E}^{2}$ takes value at least $\eta^{2}$ with probability $1$, the random variable $Z_{E}$ is a $\frac{1}{2\eta}$-Lipschitz function of ${Z_{E}}^{2}$. Combining Lemma 2.15, we have
 $H_{n}(Z_{E})\leq H_{n}(Z_{E}^{2})+O\left(\log\frac{1}{2\eta}\right).$
-Applying Lemma B.2, one deduces that
+Applying Lemma A.2, one deduces that
 $H_{n}(Z^{2})$ $\geq P(E)\,H_{n}(Z_{E}^{2})\geq P(E)\left(H_{n}(Z_{E})+O\left(\log\frac{1}{\eta}\right)+O(1)\right)$
-$\geq P(E)\left(\frac{nt}{P(E)}+O(\log C)+O(|\log R|)\right.$
+$\geq P(E)\left(\frac{nt}{P(E)}+O(\log_{+}C)+O(\log_{+}R)\right$.
 $\qquad\qquad\left.+O(|\log(2s-t-1/2)|)+O_{t}(1)+O(C\eta^{2s}n)+O\left(\log\frac{1}{\eta}\right)\right).$ (19)
-Now we take $t=s+\frac{2s-1}{3}$ when $41/2<s\leq 3/$ 4, and take $t=1$ when $13/4<s<$ 1.
+Now we take $t=s+\frac{2s-1}{3}$ when $1/2<s\leq 3/4$, and take $t=1$ when $3/4<s<1$.
 Moreover, assume that the implied constant in the term $O(C\eta^{2s}n)$ in (19) is $c_{0}$, with $c_{0}\geq 2$.
 We take $\eta$ to be the number satisfying
-$c_{0}C\eta^{2s}=\begin{cases}(2s-1)/6,\quad&\text{if }1/2<s\leq 3/4,\\ (2-2s)/3,\quad&\text{if }3/4<s<1.\\ \end{cases}$
-One can verify that the requirements involving $t$ and $\eta$ are satisfied, and $12s-t-1/2\gg_{s}$ 1. Also, $\log(1/\eta)<\log C+O_{s}(1)$. As a result, we conclude that
-$H_{n}(Z^{2})\geq n(s+\epsilon_{0})+O(\log C)+O(|\log R|)+O_{s}(1),$
-with $\epsilon_{0}=(2s-1)/6$ when $41/2<s\leq 3/$ 4, and $\epsilon_{0}=(1-s)/3$ when $13/4<s<$ 1. ∎
+[…]
+One can verify that the requirements involving $t$ and $\eta$ are satisfied, and $2s-t-1/2\gg_{s}1$. Also, $\log(1/\eta)<\log C+O_{s}(1)$. As a result, we conclude that
+$H_{n}(Z^{2})\geq n(s+\epsilon_{0})+O(\log_{+}C)+O(\log_{+}R)+O_{s}(1),$
+with $\epsilon_{0}=(2s-1)/6$ when $1/2<s\leq 3/4$, and $\epsilon_{0}=(1-s)/3$ when $3/4<s<1$. ∎
 The $\epsilon_{0}$ in the above theorem can be calculated and optimized explicitly. However, we do not pursue that direction in this paper.
@@
 Let $X$, $X^{\prime}$, $Y$, $Y^{\prime}$ be independent real random variables.
-Suppose that $X$ and $X^{\prime}$ are both $(s,C)$-Frostman with ranges contained in $[-c,c]$, where $0<s\leq 1$ and $C,c\geq 1$.
+Suppose that $X$ and $X^{\prime}$ are both $(s,C)$-Frostman with ranges contained in [c,c][-c,c], where $0<s\leq 1$ and $C,c>0$.
 Then, for any $n\in\mathbb{N}$,
-$H_{n}\big{(}(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}\big{)}\geq ns+O_{s}(\log C)+O_{s}(\log n),$
+$H_{n}\big{(}(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}\big{)}\geq ns+O_{s}(\log_{+}C)+O_{s}(\log n),$
 and
-$H_{n}\big{(}\sqrt{(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}}\big{)}\geq ns+O_{s}(\log C)+O(\log c)+O_{s}(\log n).$
+$H_{n}\big{(}\sqrt{(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}}\big{)}\geq ns+O_{s}(\log_{+}C)+O(\log_{+}c)+O_{s}(\log n).$
 Proof.
-By Corollary 2.22, one has
+By (13), one has
 $H_{n}\big{(}(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}\big{)}\geq H_{n}\big{(}(X-X^{\prime})^{2}\big{)}+O(1).$
@@
 Follow the same arguments as in the proof of Theorem 2.33, we obtain that
-$H(Z^{2})\geq ns+O(\log C)+O(C\eta^{s}n)+O\big{(}\log\frac{1}{\eta}\big{)}+O(1).$
-The first inequality follows by taking $\eta=2^{-1}\big{(}\frac{\log n}{2Cn}\big{)}^{1/s}$. The second inequality follows from the fact that
-$H_{n}\big{(}(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}\big{)}\leq H_{n}\big{(}\sqrt{(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}}\big{)}+O(\log c).$
+$H(Z^{2})\geq ns+O(\log_{+}C)+O(C\eta^{s}n)+O\big{(}\log\frac{1}{\eta}\big{)}+O(1).$
+The first inequality follows by taking $\eta=\min\left\{\frac{1}{2}\big{(}\frac{\log n}{2Cn}\big{)}^{1/s},\,\frac{1}{10}\right\}$. The second inequality follows from the fact that
+$H_{n}\big{(}(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}\big{)}\leq H_{n}\big{(}\sqrt{(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}}\big{)}+O(\log_{+}c)+O(1).$
 3 Discretized entropy Balog-Szemerédi-Gowers theorem
-In this section, we follow the method introduced by Tao in [22, Theorem 3.1] to obtain a generalized discretized entropy version of the Balog-Szemerédi-Gowers theorem.
+In this section, we follow the method introduced by Tao in [26, Theorem 3.1] to obtain a generalized discretized entropy version of the Balog-Szemerédi-Gowers theorem.
 We do not know whether such a generalization has been studied in the literature, so we provide a complete proof in this section.
@@
 Remark 3.2.
-The original version of entropy Balog-Szemerédi-Gowers theorem in [22] states for $H$ instead of $H_{n}$, does not involve the functions $f$, $g$, and conditions (ii)–(iv) are “given $Y^{\prime}$, $X^{\prime}$, $X^{\prime}$” instead of “given $D_{n}(Y^{\prime})$, $D_{n}(X^{\prime})$, $D_{n}(X^{\prime})$”, respectively.
+The original version of entropy Balog-Szemerédi-Gowers theorem in [26] states for $H$ instead of $H_{n}$, does not involve the functions $f$, $g$, and conditions (ii)–(iv) are “given $Y^{\prime}$, $^{\prime}, XX^{\prime}$” instead of “given $D_{n}(Y^{\prime})$, $D_{n}(X^{\prime})$, $D_{n}(X^{\prime})$”, respectively.
 In such a situation, (i)–(iv) is equivalent to the statement that $X^{\prime},Y,X^{\prime\prime},Y^{\prime\prime}$ forms a Markov chain.
@@
 Proof.
-Applying Lemma B.1 with $A=0$, $B=X$, $C=Y$ and $h=D_{n}$, we obtain a random variable $(X_{1},Y_{1},X_{2})$ such that: $(X_{1},Y_{1})\equiv(X,Y)$; $X_{2}$ and $X_{1}$ are conditionally i.i.d.
+Applying Lemma A.1 with $A=0$, $B=X$, $C=Y$ and $h=D_{n}$, we obtain a random variable $(X_{1},Y_{1},X_{2})$ such that: $(X_{1},Y_{1})\equiv(X,Y)$; $X_{2}$ and $X_{1}$ are conditionally i.i.d.
 given $D_{n}(Y_{1})$.
-Next, let us apply Lemma B.1 again, with $A=X_{2}$, $B=Y_{1}$, $C=X_{1}$ and $h=D_{n}$.
+Next, let us apply Lemma A.1 again, with $A=X_{2}$, $B=Y_{1}$, $C=X_{1}$ and $h=D_{n}$.
 There exists a random variable $(X_{2}^{\prime},Y_{1}^{\prime},X_{1}^{\prime},Y_{1}^{\prime\prime})$ such that: $(X_{2}^{\prime},Y_{1}^{\prime},X_{1}^{\prime})\equiv(X_{2},Y_{1},X_{1})$, $(Y_{1}^{\prime\prime},D_{n}(X_{1}^{\prime}))\equiv(Y_{1},D_{n}(X_{1}))$; $Y_{1}^{\prime\prime}$ are $Y_{1}^{\prime}$ are conditionally i.i.d.
 given $D_{n}(X_{1}^{\prime})$; also $(X_{2}^{\prime},Y_{1}^{\prime})$ and $Y_{1}^{\prime\prime}$ are conditionally independent given $D_{n}(X_{1}^{\prime})$.
-One can also use $A=(X_{2},X_{1},D_{n}(Y_{1}))$ in the second application of Lemma B.1 to see $(X_{2}^{\prime},X_{1}^{\prime},D_{n}(Y_{1}^{\prime}))\equiv(X_{2},X_{1},D_{n}(Y_{1}))$.
-Now (i)–(iv) hold by taking $X^{\prime}=X_{1}^{\prime}$, $X^{\prime\prime}=X_{2}^{\prime}$, $Y^{\prime}=Y_{1}^{\prime}$, $Y^{\prime\prime}=Y_{1}^{\prime\prime}$.
+One can also use $A=(X_{2},X_{1},D_{n}(Y_{1}))$ in the second application of Lemma A.1 to see $(X_{2}^{\prime},X_{1}^{\prime},D_{n}(Y_{1}^{\prime}))\equiv(X_{2},X_{1},D_{n}(Y_{1}))$.
+Now (i)–(iv) hold by taking $^{\prime}=X_{1}^{\prime}, X′′=X$$^{\prime\prime}=X_{2}^{\prime}$, $^{\prime}=Y_{1}^{\prime}, Y′′=Y$$^{\prime\prime}=Y_{1}^{\prime\prime}$.
 Moreover, (iii) implies that $Y^{\prime\prime}$ and $D_{n}(Y^{\prime})$ are conditionally independent given $D_{n}(X^{\prime})$.
@@
 given $D_{n}(Y^{\prime})$.
-By Lemmas 2.8 and 2.19,
+By Lemmas 2.8 and 2.18,
 $H_{n}(X^{\prime},X^{\prime\prime},Z,Y^{\prime})=2H_{n}(X^{\prime},X^{\prime\prime},Y^{\prime})-H_{n}(X^{\prime},X^{\prime\prime})$
@@
 For (24), Using Lemma 2.4(iv), one has
-$\begin{split}&H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|X^{\prime},Y^{\prime})\leq H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|Y^{\prime})=H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime}),Y^{\prime})-H_{n}(Y^{\prime}).\\ \end{split}$ (26)
-Using Lemmas 2.16 and 2.18, we have that
+[…]
+Using Lemmas 2.15 and 2.17, we have that
 $H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime}),Y^{\prime})+H_{n}(X^{\prime},Y^{\prime},X^{\prime\prime},Y^{\prime\prime})$
@@
 Thus, the estimate (24) follows. ∎
-4 Proof of Theorem 1.1 with $s>1/2$: diagonal forms
+4 Proof of Theorems 1.2-1.4 with $s>1/2$
+4.1 Diagonal forms with independent random variables
 In this section, we consider the diagonal real quadratic form $\phi(x,y)=\alpha x^{2}+\beta y^{2}$.
@@
 Lemma 4.1.
-Let $\phi(x,y)=\alpha x^{2}+\beta y^{2}$, where $\alpha,\beta\in\mathbb{R}^{\ast}$.
-Let $X$, $Y$ be independent random variables with $range_{+}(X)\cup range_{+}(Y)\subseteq[-c_{1},-c_{2}]\cup[c_{2},c_{1}]$, where $0<c_{2}<c_{1}$.
-Suppose that $X$ and $Y$ are $(s,C_{1})$-Frostman and $(s,C_{2})$-Frostman, respectively, with $11/2<s<$ 1 and $C_{1},C_{2}\geq 1$.
-Then there exists a constant $\epsilon_{1}=\epsilon_{1}(s)>0$ such that
-$\max\{H_{n}(X+Y),~{}H_{n}(\phi(X,Y))\}\geq n(s+\epsilon_{1})+O(\log C_{1}C_{2})+O_{\phi,s,c_{1},c_{2}}(1)$
+Let $\alpha,\beta\in\mathbb{R}^{\ast}$, $1/2<s<1$, $C_{1},C_{2}>0$ and $0<c_{2}<c_{1}$.
+There exists a constant $\epsilon_{1}=\epsilon_{1}(s)>0$ such that the following statement holds: Let $X$, $Y$ be independent random variables with $range_{+}(X)\cup range_{+}(Y)\subseteq[-c_{1},-c_{2}]\cup[c_{2},c_{1}]$.
+Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, where $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
+Then
+$\begin{split}&\max\{H_{n}(X+Y),~H_{n}(\alpha X^{2}+\beta Y^{2})\}\geq n(s+\epsilon_{1})\\ &\qquad\qquad\qquad+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi,s}(1).\end{split}$ (35)
 for all $n\in\mathbb{N}$.
-Remark 4.2.
-Compared to Theorem 1.1, this provides a stronger statement, namely, for all non-zero real-coefficients $\alpha$ and $\beta$.
 Proof.
-Write $\phi(x,y)=p(x)+q(y)$, where $p(x)=\alpha x^{2}$ and $q(y)=\beta y^{2}$ are functions whose domains are both taken to be $[-c_{2},-c_{1}]\cup[c_{2},c_{1}]$. One asserts by Example 2.17 that
-$H_{n}(p(X))=H_{n}(X)+O_{\phi,c_{1},c_{2}}(1),\quad H_{n}(q(Y))=H_{n}(Y)+O_{\phi,c_{1},c_{2}}(1).$ (31)
+Write $\phi(x,y)=p(x)+q(y)$, where $p(x)=\alpha x^{2}$ and $q(y)=\beta y^{2}$ are functions whose domains are both taken to be $c_{2},-c_{1}]\cup[c_{2},c_{1}]$. One asserts by Example 2.16 that
+$\begin{split}H_{n}(p(X))&=H_{n}(X)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1),\\ H_{n}(q(Y))&=H_{n}(Y)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1).\end{split}$ (36)
 Define $W=(\sqrt{|\alpha|}X,\sqrt{|\beta|}Y),$ with $|\alpha|X$ and $|\beta|Y$ being $(s,O_{\phi}(C_{1}))$-Frostman and $(s,O_{\phi}(C_{2}))$-Frostman, respectively.
@@
 $f_{1}(z_{1},z_{2},z_{3})=2z_{3}-\alpha z_{1}^{2}-\beta z_{2}^{2}.$
-From Example 2.12 and the fact that $range(X_{1}+X_{2}),range(Y_{1}+Y_{2})\subseteq[-2c_{1},2c_{1}]$, one sees that $f_{1}$ is $O_{\phi,c_{1}}(1)$-Lipschitz.
-We also have that $|W_{1}-W_{2}|^{2}=f_{2}(X_{1}-X_{2},Y_{1}-Y_{2})$, with the $O_{\phi,c_{1}}(1)$-Lipschitz function
+From Example 2.11 and the fact that $range(X_{1}+X_{2}),range(Y_{1}+Y_{2})\subseteq[-2c_{1},2c_{1}]$, one sees that $f_{1}$ is $O_{\phi}(c_{1})$-Lipschitz.
+We also have that $|W_{1}-W_{2}|^{2}=f_{2}(X_{1}-X_{2},Y_{1}-Y_{2})$, with the $O_{\phi}(C_{1})$-Lipschitz function
 $f_{2}(z_{1},z_{2})=\alpha z_{1}^{2}+\beta z_{2}^{2}.$
@@
 $H_{n}(|W_{1}-W_{2}|^{2})+H_{n}(X_{1},X_{2},Y_{1},Y_{2})$
-$\leq H_{n}(X_{1}+X_{2},Y_{1}+Y_{2},p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}))+H_{n}(X_{1}-X_{2},Y_{1}-Y_{2})+O_{\phi,c_{1}}(1)$
+$\leq H_{n}(X_{1}+X_{2},Y_{1}+Y_{2},p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}))$
+$\qquad+H_{n}(X_{1}-X_{2},Y_{1}-Y_{2})+O(\log_{+}c_{1})+O_{\phi}(1)$
 $\leq H_{n}(X_{1}+X_{2})+H_{n}(Y_{1}+Y_{2})+H_{n}(p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}))$
-$\qquad+H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})+O_{\phi,c_{1}}(1).$ (32)
-We now estimate the terms in the last expression. For the first term, apply Corollary 2.22 to get that
+$\qquad+H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})+O(\log_{+}c_{1})+O_{\phi}(1).$ (37)
+We now estimate the terms in the last expression. For the first term, apply Lemma 2.20 to get that
 $H_{n}(X_{1}+X_{2})$ $\leq H_{n}(X_{1}+Y_{1})+H_{n}(X_{2}+Y_{1})-H_{n}(Y_{1})+O(1),$
@@
 $H_{n}(X_{1}+X_{2})+H_{n}(Y_{1}+Y_{2})\leq 4H_{n}(X+Y)-H_{n}(X)-H_{n}(Y)+O(1).$ (38)
-For the third term, we use Corollary 2.22 repeatedly, and combine Lemma 2.16 and formula (31), to get
+For the third term, we use Lemma 2.20 repeatedly, and combine Lemma 2.15 and formula (36), to get
 $H_{n}(p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}))$
@@
 $\qquad-H_{n}(p(X_{1})+q(Y_{1}))+O(1)$
-$=3H_{n}(p(X)+q(Y))-H_{n}(X)-H_{n}(Y)+O_{\phi,c_{1},c_{2}}(1).$ (34)
-For the fourth term, applying Corollary 2.22 shows that
+$=3H_{n}(p(X)+q(Y))-H_{n}(X)-H_{n}(Y)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1).$ (39)
+For the fourth term, applying Lemma 2.20 shows that
 $H_{n}(X_{1}-X_{2})\leq H_{n}(X_{1}+Y_{1})+H_{n}(X_{2}+Y_{1})-H_{n}(Y_{1})+O(1).$
@@
 Combining (37)–(40) and Lemma 2.4(iv), we have that
-$H_{n}\big{(}|W_{1}-W_{2}|^{2}\big{)}+5H_{n}(X)+5H_{n}(Y)\leq 8H_{n}(X+Y)+3H_{n}(\phi(X,Y))+O_{\phi,c_{1},c_{2}}(1).$
-Note that $X$ and $Y$ are independent, and are $(s,C_{1})$-Frostman and $(s,C_{2})$-Frostman, respectively.
-One asserts that $W$ is $(2s,C_{1}C_{2}(\alpha\beta)^{-s/2})$-Frostman.
+$H_{n}\big{(}|W_{1}-W_{2}|^{2}\big{)}+$ $5H_{n}(X)+5H_{n}(Y)\leq 8H_{n}(X+Y)+3H_{n}(\phi(X,Y))$
+$\quad+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1).$
+Since $X$ and $Y$ are independent, and are $(s_{1},C_{1})$-Frostman and $(s_{2},C_{2})$-Frostman, respectively, one asserts that $W$ is $(s_{1}+s_{2},C_{1}C_{2}\alpha^{-s_{1}}\beta^{-s_{2}})$-Frostman.
+Here $\log(\alpha^{-s_{1}}\beta^{-s_{2}})=O_{\phi}(1)$.
 Noting that $range(W)$ is contained in the ball in ${\mathbb{R}}^{2}$ centered at origin with radius $O_{\phi,c_{1}}(1)$.
 So we can deduce by Lemma 2.24 and Theorem 2.33 that
-$H_{n}\big{(}X\big{)}\geq ns-\log C_{1}+O(1),\quad H_{n}\big{(}Y\big{)}\geq ns-\log C_{2}+O(1),$
-$H_{n}\big{(}|W_{1}-W_{2}|^{2}\big{)}\geq n(s+\epsilon_{0})+O(\log C_{1}C_{2})+O_{\phi,s,c_{1},c_{2}}(1)$
+$H_{n}\big{(}X\big{)}\geq ns_{1}-\log C_{1}+O(1),\quad H_{n}\big{(}Y\big{)}\geq ns_{2}-\log C_{2}+O(1),$
+$H_{n}\big{(}|W_{1}-W_{2}|^{2}\big{)}\geq n(s+\epsilon_{0})+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O_{\phi,s}(1)$
 for some $\epsilon_{0}=\epsilon_{0}(s)>0$ and all $n\in\mathbb{N}$. It follows that
-$n\left(11s+\epsilon_{0}\right)+O(\log C_{1}C_{2})+O_{\phi,s,c_{1},c_{2}}(1)\leq 8H_{n}(X+Y)+3H_{n}(\phi(X,Y)).$
-As a consequence,
-$\max\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\}$ $\geq n\left(s+\frac{\epsilon_{0}}{11}\right)+O(\log C_{1}C_{2})+O_{\phi,s,c_{1},c_{2}}(1).$
+$n\left(s+\epsilon_{0}+5s_{1}+5s_{2}\right)\leq 8H_{n}(X+Y)+3H_{n}(\phi(X,Y))$
+$\qquad\qquad+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi,s}(1).$
+Recalling that $s_{1}+s_{2}\geq 2s$, we conclude that
+$\max\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\}\geq n\left(s+\frac{\epsilon_{0}}{11}\right)$
+$\qquad\qquad+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi,s}(1).$
 Hence we complete the proof of the lemma with constant $\epsilon_{1}=\epsilon_{0}/11$.
@@
 $\leq H_{n}(X_{1}+X_{2})+H_{n}(Y_{1}+Y_{2})+H_{n}(p(X_{1})+p(X_{2})-q(Y_{1})-q(Y_{2}))$
-$\qquad+H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})+O_{\phi,c_{1}}(1).$
+$\qquad+H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})+O(\log_{+}c_{1})+O_{\phi}(1).$
 Compared with (39), the second entropy can be bounded by
@@
 $H_{n}(p(X_{1})-q(Y_{1}))\leq 3H_{n}(p(X_{1})+q(Y_{1}))-H_{n}(p(X_{1}))-H_{n}(q(Y_{1}))+O(1)$
-$=3H_{n}(p(X)+q(Y))-H_{n}(X)-H_{n}(q(Y))+O_{\phi,c_{1},c_{2}}(1).$
+$\quad=3H_{n}(p(X)+q(Y))-H_{n}(X)-H_{n}(Y)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1).$
 It leads to
-$H_{n}\big{(}|W_{1}-W_{2}|^{2}\big{)}+8H_{n}(X)+8H_{n}(Y)\leq 8H_{n}(X+Y)+9H_{n}(\phi(X,Y))+O_{\phi,c_{1},c_{2}}(1),$
-and
-$n\left(17s+\epsilon_{0}\right)+O(\log C_{1}C_{2})+O_{\phi,s,c_{1},c_{2}}(1)\leq 8H_{n}(X+Y)+9H_{n}(\phi(X,Y)).$
+$H_{n}\big{(}|W_{1}-W_{2}|^{2}\big{)}+$ $8H_{n}(X)+8H_{n}(Y)\leq 8H_{n}(X+Y)+9H_{n}(\phi(X,Y))$
+$\quad+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1),$
+and then
+$n\left(s+\epsilon_{0}+8s_{1}+8s_{2}\right)\leq 8H_{n}(X+Y)+9H_{n}(\phi(X,Y))$
+$\qquad\qquad+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi,s}(1).$
 Finally, the lemma will follow with constant $\epsilon_{1}=\epsilon_{0}/17$. ∎
@@
 We emphasize that the parameters $C_{1},C_{2}$ in Lemma 4.1 can be unbounded as $n$ tends to infinity.
-And, when we apply the above lemma in the next section, the term $O(\log C_{1}C_{2})$ will lead to some entropy-terms.
+And, when we apply the above lemma in the next subsection, the term $O(\log C_{1}C_{2})$ will lead to some entropy-terms.
 So it cannot be replaced by $O_{C_{1},C_{2}}(1)$.
-5 Proof of Theorem 1.1 with $s>1/2$: general forms
-In this section, we first consider some special quadratic forms that are more complicated than the diagonal ones.
-In the following lemma, when we say a constant depends on the quadratic form $\phi$, this means that it depends on the coefficients $\alpha$ and $\beta$ occurred in $\phi$.
-Lemma 5.1.
-Let $\phi(x,y)=\alpha(x-py)^{2}+\beta(x+y)^{2}$, where $\alpha,\beta\in{\mathbb{R}}^{\ast}$ and $p\in\{0,1\}$.
-Let $X$ and $Y$ be independent $(s,C)$-Frostman random variables whose ranges are contained in $[-c_{1},c_{1}]$, where $11/2<s<$ 1, $C\geq 1$ and $c_{1}>0$.
-Then there exists a positive constant $\epsilon_{2}=\epsilon_{2}(s)$ and an $N=N(\phi,s,C,c_{1})$ such that
-$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{2})$
-for all $n>N$.
+4.2 Diagonal forms with dependent random variables
+Now we prove Theorem 1.4 in the following form, where $X$ and $Y$ are allowed to be dependent.
+We still take $\phi(x,y)=\alpha x^{2}+\beta y^{2}$ in this subsection.
+And a constant depending on $\phi$ means it depends on $\alpha$ and $\beta$.
+Theorem 1.4’.
+Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$, $C\geq 1$ and $c>0$.
+There is a positive constant $\epsilon_{2}=\epsilon_{2}(s)$ such that the following statement holds: For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
+$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\geq n(s+\epsilon_{2})+O_{s}(\log_{+}C)+O(\log_{+}c)+O_{\phi,s}(1).$
 Proof.
 For simplicity, we write
-$\Psi:=\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}-ns,$ (36)
-and our aim is to prove that $\Psi\geq\epsilon_{2}n$ for some positive $\epsilon_{2}$.
-Since $X$, $Y$ are independent and both $(s,C)$-Frostman, we deduce by Lemma 2.26 Corollary 2.22 and (36) that
-$\begin{split}&\min\{H_{n}(X),\,H_{n}(Y)\}\geq ns+O_{C}(1),\\ &\max\{H_{n}(X),\,H_{n}(Y)\}\leq H_{n}(X+Y)+O(1)\leq ns+\Psi+O(1).\end{split}$ (37)
-Let us define the random variable $(X_{0},Y_{0})$ as
-$(X_{0},Y_{0})=(X-pY,X+Y)$
-with $p\in\{0,1\}$.
-The ranges of $X_{0}$ and $Y_{0}$ are both contained in $[-2c_{1},2c_{1}]$.
-And $(X_{0},Y_{0})$ is an $O(1)$-bi-Lipschitz function of $(X,Y)$, which, because $X$ and $Y$ are independent, leads to the estimate
-$H_{n}(X_{0},Y_{0})=H_{n}(X,Y)+O(1)=H_{n}(X)+H_{n}(Y)+O(1)\leq 2ns+2\Psi+O(1).$ (38)
-By Lemma 2.25, one deduces that $X_{0}$ and $Y_{0}$ are $(s,C_{1}^{\prime})$-Frostman and $(s,C_{2}^{\prime})$-Frostman, respectively, with $C_{1}^{\prime},C_{2}^{\prime}=O_{C}(1)$.
-Similarly, one also has that $X_{0}+Y_{0}=(2X+(1-p)Y)$ is $(s,O_{C}(1))$-Frostman, and $(X_{0},Y_{0})$ is $(2s,O_{C}(1))$-Frostman.
-We then show below that
-$H_{n}(X_{0})\leq ns+O(\Psi)+O_{C}(1),\quad H_{n}(X_{0}+Y_{0})\leq ns+O(\Psi)+O_{C}(1).$ (39)
-Indeed, when $p=0$, one sees by (36), (37) and Lemma 2.23 that
-$H_{n}(X_{0})=H_{n}(X)\leq ns+\Psi+O(1),$
-and
-$H_{n}(X_{0}+Y_{0})=H_{n}(2X+Y)$
-$\qquad\leq H_{n}(X+Y)+O(H_{n}(X+Y)-H_{n}(X))+O(H(X+Y)-H(Y))+O(1)$
-$\qquad\leq ns+O(\Psi)+O_{C}(1).$
-When $p=1$, it follows from (12), (36) and (37) that
-$H_{n}(X_{0})=H_{n}(X-Y)\leq 3H_{n}(X+Y)-H(X)-H(Y)+O(1)\leq ns+3\Psi+O_{C}(1)$
-and
-$H_{n}(X_{0}+Y_{0})=H_{n}(2X)=H_{n}(X)+O(1)\leq ns+\Psi+O(1).$
-So (39) follows.
-[…]
-$P(E)=P(|X_{0}|>\eta,|Y_{0}|>\eta)\geq 1-P(|X_{0}|<\eta)-P(|Y_{0}|<\eta)\geq 1-C_{1}^{\prime}\eta^{s}-C_{2}^{\prime}\eta^{s}>\frac{1}{2},$
-Lemma B.2 shows that $(X_{1},Y_{1})$ is $(2s,O_{C}(1))$-Frostman.
-Similarly, all of $X_{1}$, $Y_{1}$ and $X_{1}+Y_{1}$ are $(s,O_{C}(1))$-Frostman.
-We can apply Lemmas B.2 and 2.26, then combine (36), (38), (39) to get that
-$\begin{split}&2ns+O_{C}(1)\leq H_{n}(X_{1},Y_{1})\leq\frac{H_{n}(X_{0},Y_{0})}{P(E)}\leq(1+O_{C}(\eta^{s}))(2ns+2\Psi+O(1)),\\ &ns+O_{C}(1)\leq H_{n}(X_{1})\leq\frac{H_{n}(X_{0})}{P(E)}\leq(1+O_{C}(\eta^{s}))(ns+O(\Psi)+O_{C}(1)),\\ &ns+O_{C}(1)\leq H_{n}(Y_{1})\leq\frac{H_{n}(X+Y)}{P(E)}\leq(1+O_{C}(\eta^{s}))(ns+\Psi),\\ &ns+O_{C}(1)\leq H_{n}(X_{1}+Y_{1})\leq\frac{H_{n}(X_{0}+Y_{0})}{P(E)}\leq(1+O_{C}(\eta^{s}))(ns+O(\Psi)+O_{C}(1)),\end{split}$ (40)
-and
-$H_{n}(\alpha X_{1}^{2}+\beta Y_{1}^{2})\leq\frac{H_{n}(\alpha X_{0}^{2}+\beta Y_{0}^{2})}{P(E)}=(1+O_{C}(\eta^{s}))H_{n}(\phi(X,Y))\leq(1+O_{C}(\eta^{s}))(ns+\Psi).$ (41)
+$\Psi:=\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}-ns,$
+and our aim is to prove that $\Psi\geq\epsilon n$ for some positive $\epsilon$. In particular,
+$\frac{H_{n}(X)+H_{n}(Y)}{2},\,\,H_{n}(X+Y),\,\,H_{n}(\alpha X^{2}+\beta Y^{2})\,\leq\,ns+\Psi.$ (41)
+By Lemma 2.22, the random variables $X$, $Y$, $(X,Y)$ are $(s_{1},C\cdot\max\{1,c\})$-Frostman, $(s_{2},C\cdot\max\{1,c\})$-Frostman and $(s_{1}+s_{2},C)$-Frostman, respectively.
+Let $\eta$ be a parameter to be chosen later, with $0<\eta<\min\{2^{-1}(1/2C)^{1/2s},\,1/10\}$.
+We denote by $E$ the event that $|(X,Y)|\geq\eta$, and by $P_{E}$ the corresponding conditional probability measure.
+Denote by $(X_{1},Y_{1})$ the random variable such that $(X_{1},Y_{1})\equiv(X,Y|E)$.
+Since
+$P(E)$ $=P(|(X,Y)|\geq\eta)=1-P\big{(}(X,Y)\in B((0,0),\eta)\big{)}$
+$\geq 1-C\eta^{s_{1}+s_{2}}\geq 1-C\eta^{2s}>\frac{1}{2},$
+Lemma A.2 shows that the random variables $X_{1}$, $Y_{1}$, $(X_{1},Y_{1})$ are $(s_{1},O(C\cdot\max\{1,c\}))$-Frostman, $(s_{2},O(C\cdot\max\{1,c\}))$-Frostman and $(s_{1}+s_{2},O(C)))$-Frostman, respectively.
+Combining Lemma 2.24(i), one sees that
+$\begin{split}H_{n}(X_{1})+H_{n}(Y_{1})&\geq n(s_{1}+s_{2})-2\log C+O(\log_{+}c)+O(1),\\ H_{n}(X_{1},Y_{1})&\geq n(s_{1}+s_{2})-\log C+O(1).\end{split}$ (42)
+Since $s_{1}+s_{2}\geq 2s$, the right-hand sides of both inequalities in (42) are at least $2ns+O(\log_{+}C)+O(\log_{+}c)+O(1)$.
+On the other hand, We can apply Lemma A.2, then combine (41) to get that
+$\begin{split}&\frac{H_{n}(X_{1})+H(Y_{1})}{2}\leq\frac{H_{n}(X)+H_{n}(Y)}{2P(E)}\leq(1+O(C\eta^{2s}))(ns+\Psi),\\ &H_{n}(X_{1}+Y_{1})\leq\frac{H_{n}(X+Y)}{P(E)}\leq(1+O(C\eta^{2s}))(ns+\Psi),\\ &H_{n}(\alpha X_{1}^{2}+\beta Y_{1}^{2})\leq\frac{H_{n}(\alpha X^{2}+\beta Y^{2})}{P(E)}\leq(1+O(C\eta^{2s}))(ns+\Psi).\end{split}$ (43)
 It follows from (42) and (43) that
-$\begin{split}&H_{n}(X_{1})+H_{n}(Y_{1})-H_{n}(X_{1},Y_{1})=O(\Psi)+O_{C}(\sqrt{\eta}n)+O_{C}(1),\\ &H_{n}(X_{1}+Y_{1})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}=O(\Psi)+O_{C}(\sqrt{\eta}n)+O_{C}(1),\\ &H_{n}(\alpha X_{1}^{2}+\beta Y_{1}^{2})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}=O(\Psi)+O_{C}(\sqrt{\eta}n)+O_{C}(1).\end{split}$
-Now we write $\log K:=O(\Psi)+O_{C}(\sqrt{\eta}n)+O_{C}(1)$.
+$\begin{split}&H_{n}(X_{1})+H_{n}(Y_{1})-H_{n}(X_{1},Y_{1})=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(1),\\ &H_{n}(X_{1}+Y_{1})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(1),\\ &H_{n}(\alpha X_{1}^{2}+\beta Y_{1}^{2})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(1).\end{split}$
+Now we write $\log K:=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(1)$.
 Applying Theorem 3.1 with the random variable $(X_{1},Y_{1})$, we obtain the random variable $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$, where $X^{\prime\prime}$ and $Y^{\prime\prime}$ are conditionally independent given $D_{n}(X^{\prime},Y^{\prime})$, and
-$H_{n}(X^{\prime\prime}|X^{\prime},Y^{\prime})\geq H_{n}(X_{1})-\log K,\quad H_{n}(Y^{\prime\prime}|X^{\prime},Y^{\prime})\geq H_{n}(Y_{1})-\log K.\$ (42)
+$H_{n}(X^{\prime\prime}|X^{\prime},Y^{\prime})\geq H_{n}(X_{1})-\log K,\quad H_{n}(Y^{\prime\prime}|X^{\prime},Y^{\prime})\geq H_{n}(Y_{1})-\log K.\\$ (44)
 First, let us take the functions $f$ and $g$ in Theorem 3.1 to be the identity functions. Then
@@
 Second, we take the functions $f,g$ in Theorem 3.1 to be $f(x)=\alpha x^{2}$ and $g(y)=\beta y^{2}$.
-By the construction $(X_{1},Y_{1})\equiv(X_{0},Y_{0}||X_{0}|>\eta,|Y_{0}|>\eta)$, the ranges of $X^{\prime\prime}$ and $Y^{\prime\prime}$ are contained in $[-2c_{1},-\eta]\cup[\eta,2c_{1}]$ with probability $1$.
-Now, by Example 2.13, the random variables $f(X^{\prime\prime})$ and $g(Y^{\prime\prime})$ are weakly $O_{\phi}(\max\{\eta^{-1},c_{1}\})$-bi-Lipschitz functions of $X^{\prime\prime}$ and $Y^{\prime\prime}$, respectively.
+By the construction $(X_{1},Y_{1})\equiv(X,Y||(X,Y)|>\eta)$, the ranges of $X^{\prime\prime}$ and $Y^{\prime\prime}$ are contained in $c,-\eta/\sqrt{2}]\cup[\eta/\sqrt{2},c]$ with probability $1$.
+Now, by Example 2.12, the random variables $f(X^{\prime\prime})$ and $g(Y^{\prime\prime})$ are weakly $O_{\phi}(\max\{\eta^{-1},c\})$-bi-Lipschitz functions of $X^{\prime\prime}$ and $Y^{\prime\prime}$, respectively.
 It follows that
-$H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|X^{\prime},Y^{\prime})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}\leq 7\log K+O_{\phi,c_{1},\eta}(1).$ (44)
+$H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|X^{\prime},Y^{\prime})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}\leq 7\log K+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi}(1).$ (46)
 Next, conditionally given that $D_{n}(X^{\prime},Y^{\prime})=(x,y)$ with $x,y$ being dyadic rationals satisfying $P\big{(}D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}>0$, we will prove the following results on induced conditional probability measures:
-$\begin{split}\big{(}X^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}\text{ is $\left(s,O\left(C^{2}\cdot 2^{-ns}/P_{E}(D_{n}(Y^{\prime})=y)\right)\right)$-Frostman},\\ \big{(}Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}\text{ is $\left(s,O\left(C^{2}\cdot 2^{-ns}/P_{E}(D_{n}(X^{\prime})=x)\right)\right)$-Frostman},\end{split}$ (45)
+$\begin{split}\big{(}X^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}\text{ is $\left(s_{1},O\left(C\cdot 2^{-ns_{2}}/P_{E}(D_{n}(Y^{\prime})=y)\right)\right)$-Frostman},\\ \big{(}Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}\text{ is $\left(s_{2},O\left(C\cdot 2^{-ns_{1}}/P_{E}(D_{n}(X^{\prime})=x)\right)\right)$-Frostman},\end{split}$ (47)
 where the implied constants do not depend on $(x,y)$. We present the proof for $X^{\prime\prime}$, as the one for $Y^{\prime\prime}$ can be done similarly.
@@
 $\qquad=P_{E}(X^{\prime}\in B(z,r)|D_{n}(Y^{\prime})=y)=\frac{P_{E}(X^{\prime}\in B(z,r),D_{n}(Y^{\prime})=y)}{P_{E}(D_{n}(Y^{\prime})=y)}$
-$\qquad=\frac{P_{E}(X_{1}\in B(z,r),D_{n}(Y_{1})=y)}{P_{E}(D_{n}(Y^{\prime})=y)}\leq\frac{2P(X-pY\in B(z,r),D_{n}(X+Y)=y)}{P_{E}(D_{n}(Y^{\prime})=y)}.$
-Denote by $\mu_{X}$, $\mu_{Y}$ the measures associated to the independent random variables $X$ and $Y$, respectively.
-Then $\mu_{X}$ and $\mu_{Y}$ are $(s,C)$-Frostman.
-Furthermore, define $I_{n}(y):=[y,y+2^{-n})$.
-We have that
-[…]
-The value inside the integral can be positive only when $I_{n}(y-\tilde{y})\cap B(z+p\tilde{y},r)\not=\emptyset$. Or equivalently, $I_{n}(y)\cap B(z+(p+1)\tilde{y},r)\not=\emptyset$. That is to say,
-$z+(p+1)\tilde{y}-r\leq y+2^{-n},\quad z+(p+1)\tilde{y}+r\geq y.$
-We can locate such $\tilde{y}$ in an interval $J$ of length at most $2r+2^{n}$.
-We now estimate the above expression in different ways depending on how $r$ relates to $2^{-n}$.
-First, for $r\geq 2^{-n}$, and recalling that $\mu_{X}$ and $\mu_{Y}$ are $(s,C)$-Frostman, we have
-[…]
-[…]
-Next, for $r<2^{-n}$, we similarly find
-[…]
-[…]
-This completes the proof for the claims given in (45)
+$\qquad=\frac{P_{E}(X_{1}\in B(z,r),D_{n}(Y_{1})=y)}{P_{E}(D_{n}(Y^{\prime})=y)}\leq\frac{2P(X\in B(z,r),D_{n}(Y)=y)}{P_{E}(D_{n}(Y^{\prime})=y)}.$
+Since $(X,Y)$ is jointly $(s_{1},s_{2},C)$-Frostman, we have
+$P(X\in B(z,r),D_{n}(Y)=y)\ll Cr^{s_{1}}2^{-ns_{2}}.$
+This completes the proof for the claims given in (47). We mention that these two Frostman constants in (47) are $\gg 1$.
 As the expressions to follow get rather lengthy, we now introduce some shorthand.
-In what follows, $H_{n}(~{}\cdot~{}|x,y)$ is an abbreviation of $H(D_{n}(\cdot)~{}|D_{n}(X^{\prime},Y^{\prime})=(x,y)).$ By Lemma B.2, there exist random variables, depending on $(x,y)$, that have the same distributions with $(X^{\prime\prime},Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y))$, respectively.
+In what follows, $H_{n}(~\cdot~|x,y)$ is an abbreviation of $H(D_{n}(\cdot)~|D_{n}(X^{\prime},Y^{\prime})=(x,y)).$ By Lemma A.2, there exist random variables, depending on $(x,y)$, that have the same distributions with $(X^{\prime\prime},Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y))$, respectively.
 These random variables take the same value as $(X^{\prime\prime},Y^{\prime\prime})$ does.
-By construction, their ranges are contained in $\big{(}[-2c_{1},-\eta]\cup[\eta,2c_{1}]\big{)}^{2}$.
+By construction, their ranges are contained in $\big{(}[-c,-\eta/\sqrt{2}]\cup[\eta/\sqrt{2},c]\big{)}^{2}$ with probability $1$.
 Also note that $(X^{\prime\prime}|D_{n}(X_{1},Y_{1})=(x,y))$ and $(Y^{\prime\prime}|D_{n}(X_{1},Y_{1})=(x,y))$ are independent.
@@
 Finally, recall (47).
-Now we can apply Lemma 4.1 with these random variables, with $C_{1}=O_{C}\big{(}2^{-ns}/P_{E}(D_{n}(Y^{\prime})=y)\big{)}$, $C_{2}=O_{C}\big{(}2^{-ns}/P_{E}(D_{n}(X^{\prime})=x)\big{)},$ and $c_{2}=\eta$, to deduce that
+Now we can apply Lemma 4.1 with these random variables, with $C_{1}=O\big{(}C\cdot 2^{-ns_{2}}/P_{E}(D_{n}(Y^{\prime})=y)\big{)}$, $C_{2}=O\big{(}C\cdot 2^{-ns_{1}}/P_{E}(D_{n}(X^{\prime})=x)\big{)}$, $c_{1}=c$ and $c_{2}=\eta$, to deduce that
 $\max\left\{H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y),H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|x,y)\right\}\geq n(s+\epsilon_{1})$
-$\qquad+O\big{(}-2ns-\log P_{E}(D_{n}(X^{\prime})=x)-\log P_{E}(D_{n}(Y^{\prime})=y)\big{)}+O_{\phi,s,C,c_{1},\eta}(1).$ (46)
-By conditional independence, and combining Corollary 2.22, we also have that
+$\qquad+O\big{(}\log_{+}C-n(s_{1}+s_{2})-\log P_{E}(D_{n}(X^{\prime})=x)-\log P_{E}(D_{n}(Y^{\prime})=y)\big{)}$
+$\qquad+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1)$ (48)
+for some $\epsilon_{1}=\epsilon_{1}(s)>0$. By conditional independence, and combining (13), we also have that
 $H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y)\geq\max\{H_{n}(X^{\prime\prime}|x,y),H_{n}(Y^{\prime\prime}|x,y)\}+O(1).$
-Using a similar argument, Lemma B.2, and Example 2.17, one obtains
+Using a similar argument, Lemma A.2, and Example 2.16, one obtains
 $H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|x,y)\geq\max\{H_{n}(\alpha X^{\prime\prime 2}|x,y),H_{n}(\beta Y^{\prime\prime 2}|x,y)\}+O(1)$
-$=\max\{H_{n}(X^{\prime\prime}|x,y),H_{n}(Y^{\prime\prime}|x,y)\}+O_{\phi,c_{1},\eta}(1).$
+$\qquad=\max\{H_{n}(X^{\prime\prime}|x,y),H_{n}(Y^{\prime\prime}|x,y)\}+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi}(1).$
 Then
-$\begin{split}&\min\{H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y),H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|x,y)\}\geq\frac{H_{n}(X^{\prime\prime}|x,y)+H_{n}(Y^{\prime\prime}|x,y)}{2}+O_{\phi,c_{1},\eta}(1).\end{split}$ (47)
+$\min\{H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y),H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|x,y)\}$
+$\qquad\geq\frac{H_{n}(X^{\prime\prime}|x,y)+H_{n}(Y^{\prime\prime}|x,y)}{2}+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi}(1).$ (49)
 Therefore given the condition that $D_{n}(X_{1},Y_{1})=(x,y)$, it follows from (48) and (49) that
 $H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y)+H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|x,y)-\frac{H_{n}(X^{\prime\prime}|x,y)+H_{n}(Y^{\prime\prime}|x,y)}{2}$
-$\quad\geq n(s+\epsilon_{1})+O\big{(}-2ns-\log P_{E}(D_{n}(X^{\prime})=x)-\log P_{E}(D_{n}(Y^{\prime})=y)\big{)}+O_{\phi,s,C,c_{1},\eta}(1).$
+$\quad\geq n(s+\epsilon_{1})+O\big{(}-n(s_{1}+s_{2})-\log P_{E}(D_{n}(X^{\prime})=x)-\log P_{E}(D_{n}(Y^{\prime})=y)\big{)}$
+$\qquad\qquad\qquad\quad+O(\log_{+}C)+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1).$
 Now we sum all of these expressions respective to each $(x,y)$ with the weight $P_{E}(D_{n}(X^{\prime},Y^{\prime})=(x,y))$. Note that
@@
 $H_{n}(X^{\prime\prime}+Y^{\prime\prime}|X^{\prime},Y^{\prime})+H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|X^{\prime},Y^{\prime})-\frac{H_{n}(X^{\prime\prime}|X^{\prime},Y^{\prime})+H_{n}(Y^{\prime\prime}|X^{\prime},Y^{\prime})}{2}$
-$\qquad\qquad\geq n(s+\epsilon_{1})+O(-2ns+H_{n}(X^{\prime})+H_{n}(Y^{\prime}))+O_{\phi,s,C,c_{1},\eta}(1).$ (48)
+$\qquad\qquad\geq n(s+\epsilon_{1})+O(-n(s_{1}+s_{2})+H_{n}(X^{\prime})+H_{n}(Y^{\prime}))$
+$\qquad\qquad\qquad\quad+O(\log_{+}C)+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1).$ (50)
 Substituting from (44)–(46), and further combining with $(X^{\prime},Y^{\prime})\equiv(X_{1},Y_{1})$ we obtain that
-$\begin{split}&\left(\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}+7\log K\right)+\left(\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}+7\log K\right)\\ &\qquad-\frac{(H_{n}(X_{1})-\log K)+(H_{n}(Y_{1})-\log K)}{2}-ns+O(H_{n}(X_{1})+H_{n}(Y_{1})-2ns)\\ &\qquad\geq n\epsilon_{1}+O_{\phi,s,C,c_{1},\eta}(1).\end{split}$ (49)
-Recalling (40) and the notation $\log K=O(\Psi)+O_{C}(\sqrt{\eta}n)+O_{C}(1)$, the left-hand side of (49) is
-$\leq O\left(\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}-ns\right)+15\log K=O(\Psi)+O_{C}(\sqrt{\eta}n)+O_{C}(1).$
+$\begin{split}&\left(\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}+7\log K\right)+\left(\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}+7\log K\right)\\ &\qquad-\frac{(H_{n}(X_{1})-\log K)+(H_{n}(Y_{1})-\log K)}{2}-ns+O(H_{n}(X_{1})+H_{n}(Y_{1})-n(s_{1}+s_{2}))\\ &\qquad\geq n\epsilon_{1}+O(\log_{+}C)+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1).\end{split}$ (51)
+Recalling (43), the fact $s_{1}+s_{2}\geq 2s$, and the notation $\log K=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(1)$, the left-hand side of (51) is
+$\leq O\left(\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}-ns\right)+15\log K$
+$=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1).$
 So there is some absolute constant $\tilde{C}>0$ such that
-$\tilde{C}\Psi+O_{C}(\sqrt{\eta}n)\geq n\epsilon_{1}+O_{\phi,s,C,c_{1},\eta}(1).$
-Equivalently,
-$\Psi\gg n\epsilon_{1}+O_{C}(\sqrt{\eta}n)+O_{\phi,s,C,c_{1},\eta}(1).$
-[…]
-Now we are ready to finish the proof of Theorem 1.1.
-Proof of Theorem 1.1.
-For $\phi(x,y)=a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$, the non-degeneracy gives that $a_{2}^{2}\neq 4a_{1}a_{3}$.
-When $a_{3}\neq 0$, we have
-$\phi(x,y)=\left(a_{1}-\frac{a_{2}^{2}}{4a_{3}}\right)x^{2}+a_{3}\left(\frac{a_{2}}{2a_{3}}x+y\right)^{2}.$
-Take $\tilde{X}=\frac{a_{2}}{2a_{3}}X$ and $\tilde{Y}=Y$. Then $\phi(X,Y)=\tilde{\phi}(\tilde{X},\tilde{Y})$, where
-$\tilde{\phi}(\tilde{x},\tilde{y})=\left(\frac{4a_{1}a_{3}^{2}}{a_{2}^{2}}-a_{3}\right)\tilde{x}^{2}+a_{3}(\tilde{x}+\tilde{y})^{2}.$
-When $a_{1}\neq 0$, similar arguments lead to $\phi(X,Y)=\tilde{\phi}(\tilde{X},\tilde{Y})$, with $\tilde{X}=\frac{a_{2}}{2a_{1}}Y$, $\tilde{Y}=X$, and
-$\tilde{\phi}(\tilde{x},\tilde{y})=\left(\frac{4a_{1}^{2}a_{3}}{a_{2}^{2}}-a_{1}\right)\tilde{x}^{2}+a_{1}(\tilde{x}+\tilde{y})^{2}.$
-When $a_{1}=a_{3}=0$, one has $a_{2}\neq 0$ and $\phi(X,Y)=\tilde{\phi}(\tilde{X},\tilde{Y})$, where $\tilde{X}=X$, $\tilde{Y}=Y$, and
-$\tilde{\phi}(\tilde{x},\tilde{y})=-\frac{a_{2}}{4}(\tilde{x}-\tilde{y})^{2}+\frac{a_{2}}{4}(\tilde{x}+\tilde{y})^{2}.$
-In all cases, the random variables $\tilde{X}$ and $\tilde{Y}$ are independent, and both are $(s,O_{\phi,C}(1))$-Frostman.
-The ranges of $\tilde{X}$ and $\tilde{Y}$ are both contained in $[-\widetilde{c_{1}},\widetilde{c_{1}}]$, with $\widetilde{c_{1}}=O_{\phi,c_{1}}(1)$.
-Applying Lemma 5.1 with the quadratic form $\tilde{\phi}(\tilde{x},\tilde{y})$ and the random variables $\tilde{X}$ and $\tilde{Y}$, we obtain that
-$\max\{H_{n}(\xi X+\zeta Y),\,H_{n}(\phi(X,Y))\}\geq n(s+\varepsilon_{2})$
-for some non-zero rational numbers $\xi=\xi(\phi)$ and $\zeta=\zeta(\phi)$.
-Let $\xi=\frac{k}{m}$ and $\zeta=\frac{l}{m}$, with $m$ being the greatest common divisor of $\xi$ and $\zeta$.
-By Lemmas 2.23 and 2.26, there are some $d_{1}=d_{1}(\xi,\zeta)$ and $d_{2}=d_{2}(\xi,\zeta)$ such that
-$H_{n}(\xi X+\zeta Y)$ $\leq(d_{1}+d_{2}+1)H_{n}(X+Y)-d_{1}H_{n}(X)-d_{2}H_{n}(Y)+O_{\xi,\zeta}(1)$
-$\leq(d_{1}+d_{2}+1)H_{n}(X+Y)-(d_{1}+d_{2})(ns-\log C)+O_{\phi}(1).$
-That is to say,
-$H_{n}(X+Y)\geq\frac{(d_{1}+d_{2})ns+H_{n}(\xi X+\zeta Y)}{d_{1}+d_{2}+1}+O_{\phi,C}(1).$
-Now the theorem follows by taking $\epsilon=\frac{\epsilon_{2}}{2(d_{1}+d_{2}+1)}$ when $n$ is sufficiently large.
+$\tilde{C}\Psi\geq n\epsilon_{1}+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1).$
+Recall that we require $0<\eta<\min\{2^{-1}(1/2C)^{1/2s},1/10\}$, $0<\epsilon_{1}<1$, and $\epsilon_{1}$ only depends on $s$.
+Now assume that the implied constant in the term $O(C\eta^{2s}n)$ is $C^{\prime}$, with $C^{\prime}\geq 1$.
+We then take $\eta=\min\Big{\{}\frac{1}{2}\big{(}\frac{\epsilon_{1}}{2C^{\prime}C}\big{)}^{1/2s},\,\frac{1}{10}\Big{\}}$.
+Then $C^{\prime}C\eta^{2s}n\leq n\epsilon_{1}/2$.
+Therefore,
+$\Psi\gg n\epsilon_{1}+O_{s}(\log_{+}C)+O(\log_{+}c)+O_{\phi,s}(1).$
+The proof is completed. ∎
+Next, we prove Theorem 1.3 in the following form.
+Theorem 1.3’.
+Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$ and $C_{1},C_{2},c>0$.
+There is a positive constants $\epsilon_{3}=\epsilon_{3}(s)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$.
+Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
+Then
+(i) For all $n\in\mathbb{N}$,
+$\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\geq n(s+\epsilon_{3})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{s}(1).$
+(ii) For all $n\in\mathbb{N}$,
+$\max\left\{\frac{H_{n}(X+Y)+H_{n}(X-Y)}{2},\,H_{n}(XY)\right\}\geq n(s+\epsilon_{3})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{s}(1).$
+Proof.
+(i) Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, by Lemma 2.22 it is also jointly $(s_{1},s_{2},C_{1}C_{2})$-Frostman.
+Applying Theorem 1.4’, there is some $\epsilon_{2}=\epsilon_{2}(s)$ such that
+$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}$
+$\qquad\qquad\geq n(s+\epsilon_{2})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{s}(1)$
+for all $n\in\mathbb{N}$. In view of $s_{1}+s_{2}\geq 2s$, Lemma 2.24(ii) shows that
+$H_{n}(X,Y)\geq\frac{H_{n}(X)+H_{n}(Y)}{2}+ns+O(\log_{+}C_{1}C_{2})+O(1).$
+The conclusion then follows with $\epsilon_{3}=\epsilon_{2}/2$.
+(ii) Notice that
+$xy=\frac{1}{4}(x+y)^{2}-\frac{1}{4}(x-y)^{2}.$
+Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, by Lemma 2.23 it is also jointly $(s_{1},s_{2},C^{\prime\prime})$-Frostman, with $\log C^{\prime\prime}=\log C_{1}C_{2}+O(1)$.
+Applying Theorem 1.4’ with $\alpha=1/4$, $\beta=-1/4$ and the random variable (X+Y,XY)(X+Y,X-Y), we obtain
+$\max\left\{\frac{H_{n}(X+Y)+H_{n}(X-Y)}{2},\,H_{n}(2X),\,H_{n}(XY)\right\}$ (52)
+$\qquad\qquad\geq n(s+\epsilon_{2})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1)$
+for some $\epsilon_{2}=\epsilon_{2}(s)$.
+If we switch the role of $X$ and $Y$, then similar conclusion holds, with $H_{n}(2X)$ replaced by $H_{n}(2Y)$.
+Therefore, the left-hand side of (52) can be replaced by
+$\max\left\{\frac{H_{n}(X+Y)+H_{n}(X-Y)}{2},\,\frac{H_{n}(2X)+H_{n}(2Y)}{2},\,H_{n}(XY)\right\}.$
+By Lemma 2.24, combining $s_{1}+s_{2}\geq 2s$, the second item satisfies
+$H_{n}(X,Y)\geq\frac{H_{n}(2X)+H_{n}(2Y)+O(1)}{2}+ns+O(\log_{+}C_{1}C_{2})+O_{s}(1).$
+Combining the fact
+$H_{n}(X,Y)=H_{n}(X+Y,X-Y)+O(1)\leq H_{n}(X+Y)+H_{n}(X-Y)+O(1),$
+the first conclusion follows by taking $\epsilon_{3}=\epsilon_{2}/2$. ∎
+4.3 The general forms
+Now we are ready to prove Theorem 1.2.
+It actually holds for any non-degenerate real quadratic form $\phi(x,y)=a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ that is congruent over $\mathbb{Q}$ to a non-degenerate real diagonal form, i.e.,
+$\phi(x,y)=\alpha(\xi_{11}x+\xi_{12}y)^{2}+\beta(\xi_{21}x+\xi_{22}y)^{2}$ (53)
+for some $\alpha,\beta\in{\mathbb{R}}^{\ast}$ and $\xi_{11},\xi_{12},\xi_{21},\xi_{22}\in\mathbb{Q}$ with $\xi_{11}\xi_{22}-\xi_{12}\xi_{21}\neq 0$.
+Here, all the constants $\alpha,\beta,\xi_{11},\xi_{12},\xi_{21},\xi_{22}$ depends on the coefficients $a_{1},a_{2},a_{3}$ in $\phi(x,y)$.
+Theorem 1.2’ Let $\phi(x,y)$ be in the form (53).
+Let $0<s<1$, $C\geq 1$ and $c>0$.
+There is a positive constant $\epsilon_{4}=\epsilon_{4}(\phi,s)$ such that the following statement holds: Let $X$ and $Y$ be independent real random variables with ranges contained in [c,c][-c,c].
+Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, where $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
+Then
+$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(x,y))\right\}\geq n(s+\epsilon)+O_{s}(\log_{+}C)+O(\log_{+}c)+O_{\phi,s}(1)$
+for all $n\in\mathbb{N}$.
+Proof.
+By Lemma 2.23, the random variable $Z:=(\xi_{11}X+\xi_{12}Y,\xi_{21}X+\xi_{22}Y)$ is jointly $(s_{1},s_{2},C^{\prime\prime})$-Frostman, where $\log C^{\prime\prime}=\log C_{1}C_{2}+O_{\phi}(1)$.
+Moreover, the range of $Z$ is contained in $c^{\prime\prime},c^{\prime\prime}]^{2}$ with $c^{\prime\prime}=O_{\phi}(c)$.
+Applying Theorem 1.4’ with $Z$, we obtain some positive number $\epsilon_{2}=\epsilon_{2}(s)$ such that
+$\max\left\{\frac{H_{n}(\xi_{11}X+\xi_{12}Y)+H_{n}(\xi_{21}X+\xi_{22}Y)}{2},\,H_{n}\big{(}(\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big{)},\,H_{n}\big{(}\phi(X,Y)\big{)}\right\}$
+$\qquad\geq n(s+\epsilon_{2})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1)$ (54)
+for all $n\in\mathbb{N}$.
+Next, using Lemma 2.21, let us assume that $d=d(\xi_{11},\xi_{12},\xi_{21},\xi_{22})=d(\phi)$ is the positive number satisfying
+$H_{n}\big{(}(\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big{)}\leq dH_{n}(X+Y)-(d-1)\frac{H_{n}(X)+H_{n}(Y)}{2}+O(1).$ (55)
+Combining Lemma 2.24(i) and the fact that $s_{1}+s_{2}\geq 2s$, we see that the right-hand side of (55) is
+$\leq dH_{n}(X+Y)-(d-1)\frac{(ns_{1}-\log C_{1}+O(1))+(ns_{2}-\log C_{2}+O(1))}{2}+O(1)$
+$\leq dH_{n}(X+Y)-(d-1)ns+O_{\phi}(\log_{+}C_{1}C_{2})+O_{\phi}(1).$
+This leads to
+$H_{n}(X+Y)\geq\frac{H_{n}\big{(}(\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big{)}+(d-1)ns}{d}+O_{\phi}(\log_{+}C_{1}C_{2})+O_{\phi}(1).$
+Similarly, we can deduce that
+$H_{n}(X+Y)\geq\frac{H_{n}(\xi_{11}X+\xi_{12}Y)+(d^{\prime}-1)ns}{d^{\prime}}+O_{\phi}(\log_{+}C_{1}C_{2})+O_{\phi}(1),$
+and also
+$H_{n}(X+Y)\geq\frac{H_{n}(\xi_{21}X+\xi_{22}Y)+(d^{\prime\prime}-1)ns}{d^{\prime\prime}}+O_{\phi}(\log_{+}C_{1}C_{2})+O_{\phi}(1)$
+for some positive numbers $d^{\prime}=d^{\prime}(\phi)$ and $d^{\prime\prime}=d^{\prime\prime}(\phi)$. Now we deduce from (54) that
+$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{5})+O_{\phi,s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1)$
+for $\epsilon_{4}=\epsilon_{2}/(d+d^{\prime}+d^{\prime\prime})$ and all $n\in\mathbb{N}$. Theorem 1.2 then follows. ∎
 Remark 4.3.
-Indeed, our proof of Theorem 1.1 works for any real quadratic form $\phi$ that is congruent over $\mathbb{Q}$ to a real diagonal form, i.e., $\phi(x,y)=\alpha(\xi_{11}x+\xi_{12}y)^{2}+\beta(\xi_{21}x+\xi_{22}y)^{2}$, where $\alpha,\beta\in{\mathbb{R}}^{\ast}$ and $\Xi=\begin{bmatrix}\xi_{11}&\xi_{12}\\ \xi_{21}&\xi_{22}\end{bmatrix}\in GL_{2}(\mathbb{Q})$, together with the constant $\epsilon$ depending both on $s$ and $\Xi$.
+In the conclusions of Lemma 4.1 and Theorems 1.2’-1.4’, the implied constants in the terms $O_{s}(\log_{+}C_{1}C_{2})$, $O_{\phi,s}(\log_{+}C_{1}C_{2})$ or $O_{s}(\log_{+}C)$ can be independent of $s$ if either of the following two situations happens: (1) we are only considering the case $1/2<s<1$; (2) the range of random variables are bounded away from 0 by some absolute constant.
+We preserve the dependence on $s$ in these terms for the uniformity with the case $0<s\leq 1/2$.
 5 Proof of Theorem 1.2 with $s\leq 1/2$
@@
 Theorem 5.1.
-Let $X$, $Y$ be independent random variables with ranges contained in $[-c,c]$, with $c>0$.
-Suppose that $X$ and $Y$ are $(s,C_{1})$-Frostman and $(s,C_{2})$-Frostman, respectively, where $20<s\leq 1/$ 2 and $C_{1},C_{2}\geq 1$.
+Let $X$, $Y$ be independent random variables with ranges contained in [c,c][-c,c], with $c>0$.
+Suppose that $X$ and $Y$ are $(s,C_{1})$-Frostman and $(s,C_{2})$-Frostman, respectively, where $0<s\leq 1/2$ and $C_{1},C_{2}\geq 1$.
 Let $(X^{\prime},Y^{\prime})$ be i.i.d.
@@
 for all $n\in\mathbb{N}$.
-To compare Theorems 2.32 and 6.1, one may take $W=(X,Y)$ in Theorem 2.32. The random variables $X$ and $Y$ are required to be independent in Theorem 6.1, but may be dependent in Theorem 2.32.
+To compare Theorems 2.30 and 5.1, one may take $W=(X,Y)$ in Theorem 2.30. The random variables $X$ and $Y$ are required to be independent in Theorem 5.1, but may be dependent in Theorem 2.30.
 The rest of the proof of Theorem 1.2 for $s\leq 1/2$ will follow the same line as in the case $s>1/2$.
@@
 For two measures $\mu^{\prime},\mu^{\prime\prime}$ whose supports are bounded regions on ${\mathbb{R}}^{2}$, we denote by $\Delta(\mu^{\prime},\mu^{\prime\prime})$ the distance measure between $\mu^{\prime}$ and $\mu^{\prime\prime}$, i.e.,
-[…]
+$\Delta(\mu^{\prime},\mu^{\prime\prime})(B)=\int\mu^{\prime}(\{x\colon|x-y|\in B\})d\mu^{\prime\prime}(y)$
 for any Borel subset $B$ of $\mathbb{R}$}. By closely following arguments from Shmerkin and Wang, in [24], we obtain the following entropy estimate.
 Theorem 5.2.
-Let $\mu_{1}$, $\mu_{2}$, and $\nu$ be $(s,C)$-Frostman probability measures on $[0,1]$, where $20<s\leq 1/$ 2 and $C\geq 1$.
+Let $\mu_{1}$, $\mu_{2}$, and $\nu$ be $(s,C)$-Frostman probability measures on $[0,1]$, where $0<s\leq 1/2$ and $C\geq 1$.
 Assume that $\text{dist}(\operatorname{spt}(\mu_{1}),\operatorname{spt}(\mu_{2}))\geq C_{d}C^{-1/s}$ with $C_{d}>0$.
@@
 Lemma 5.3.
-Let $10<s<$ 1 and $C\geq 1$.
+Let $0<s<1$ and $C\geq 1$.
 Let $\mu$ be an $(s,C)$-Frostman probability measure with $\operatorname{spt}(\mu)\subset[0,1]$.
-For any $30<\gamma<1/$ 3, we can find a partition $\{U_{1},U_{2},U_{3}\}$ of $[0,1]$ such that the following hold:
+For any $0<\gamma<1/3$, we can find a partition $\{U_{1},U_{2},U_{3}\}$ of $[0,1]$ such that the following hold:
 (i) $\text{dist}(U_{1},U_{2})\gg(\gamma/C)^{1/s}$;
 (ii) $\mu_{U_{1}}$, $\mu_{U_{2}}$ are $(s,O(C))$-Frostman probability measures;
-(iii) $21/2-\gamma\leq\mu(U_{1}),\mu(U_{2})\leq 1/$ 2 and $\mu(U_{3})\leq\gamma$.
+(iii) $1/2-\gamma\leq\mu(U_{1}),\mu(U_{2})\leq 1/2$ and $\mu(U_{3})\leq\gamma$.
 Proof.
@@
 $1/2-\gamma\leq q_{i_{0}}<1/2\leq q_{i_{0}+1}<1/2+\gamma.$
-[…]
+Take $U_{1}=[0,\,i_{0}2^{-k})$, $U_{2}=[(i_{0}+1)2^{-k},\,1]$ and $U_{3}=[i_{0}2^{-k},\,(i_{0}+1)2^{-k})$. Then
 $1/2-\gamma\leq\mu(U_{1}),\mu(U_{2})\leq 1/2,\quad\mu(U_{3})\leq\gamma.$
-By construction, $\text{dist}(U_{1},U_{2})\geq 2^{-k}\gg(\gamma/C)^{1/s}$. And by Lemma B.2, both $\mu_{U_{1}}$ and $\mu_{U_{2}}$ are $(s,O(C))$-Frostman. This completes the proof. ∎
+By construction, $\text{dist}(U_{1},U_{2})\geq 2^{-k}\gg(\gamma/C)^{1/s}$. And by Lemma A.2, both $\mu_{U_{1}}$ and $\mu_{U_{2}}$ are $(s,O(C))$-Frostman. This completes the proof. ∎
 With above results, we are going to prove Theorem 5.1.
@@
 We apply Lemma 5.3 on the distribution measure of $X$, denoted by $\mu$.
-Then there is a partition $\{U_{1},U_{2},U_{3}\}$ of $[0,1]$ such that $\text{dist}(U_{1},U_{2})\gg_{s,t}C^{-1/s}$, both $\mu_{U_{1}}$ and $\mu_{U_{2}}$ are $(s,O(C))$-Frostman, and $21/2-\gamma\leq\mu(U_{1}),\mu(U_{2})\leq 1/$ 2.
+Then there is a partition $\{U_{1},U_{2},U_{3}\}$ of $[0,1]$ such that $\text{dist}(U_{1},U_{2})\gg_{s,t}C^{-1/s}$, both $\mu_{U_{1}}$ and $\mu_{U_{2}}$ are $(s,O(C))$-Frostman, and $1/2-\gamma\leq\mu(U_{1}),\mu(U_{2})\leq 1/2$.
 Note that $range\big{(}(X,Y,X^{\prime},Y^{\prime})\big{)}\subseteq[0,1]^{4}$.
@@
 for all $n\in\mathbb{N}$.
-For simplicity, write $a=\mu(U_{1})$ and $b=\mu(U_{2})$, with $21/2-\gamma\leq a,b\leq 1/$ 2. It follows that
+For simplicity, write $a=\mu(U_{1})$ and $b=\mu(U_{2})$, with $1/2-\gamma\leq a,b\leq 1/2$. It follows that
 $H_{n}(Z)$ $\geq\sum\limits_{i,j=1}^{2}\mu(U_{i})\mu(U_{j})H_{n}\big{(}Z_{E_{ij}}\big{)}$
@@
 $\mu_{X}(B_{r})\ll r^{s}\quad\text{for all }r\in[2^{-m_{1}},2^{-m_{0}}].$
-[…]
+We now record some related concepts that we will need in the results below.
+Basically, given two measures, $\mu$ and $\nu,$ these definitions will quantify how much mass from $\nu$ can assign to a projection associated to a point in the support of $\mu.$ Specifically, for $x\in\mathbb{R}^{d},$ and a measure $\nu\in\mathcal{P}(\mathbb{R}^{d}),$ the pair $(x,\nu)$ is said to have $(t,K,c,\delta)$-thin tubes if there exists a Borel function $\phi_{x}:\mathbb{R}^{d}\to\mathbb{R}$ such that $0\leq\phi_{x}\leq 1$, $\int\phi_{x}d\nu\geq c$, and
 […]
@@
 We say that $(\mu,\nu)$ have $(t,K,c,\delta)$-strong thin tubes if there exists a set $X$ with $\mu(X)\geq c$ and for each $x\in X$ there is a set $Y_{x}$ such that $\nu(Y_{x})\geq c$, the set $\{(x,y):x\in X,y\in Y_{x}\}$ is compact, and
-$\nu(Y_{x}\cap T)\leq K\cdot r^{t}\text{ for all }r>\delta\text{ and all open $r$-tubes $T$ through $x$}.$
+[…]
 Again, we extend this definition to $\mu,\nu\in\mathcal{M}(\mathbb{R}^{d})$ by applying the definition to $(\mu/|\mu|,\nu/|\nu|)$.
@@
 Proposition 5.6.
-[Variant of Proposition 5.11 in [20]] Let $\mu_{1},\mu_{2},\nu$ be $(s,C)$-Frostman probability measures in $[0,1],$then $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has $(s,2C,1,0)$-strong thin tubes.
+[Variant of Proposition 5.11 in [24]] Let $\mu_{1},\mu_{2},\nu$ be $(s,C)$-Frostman probability measures in [0,1],[0,1],then $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has $(s,2C,1,0)$-strong thin tubes.
 Proof.
 To see this, notice that for any tube $T$ of width $r$, assume that $T$ is defined as $T=\{(x,y)\in\mathbb{R}^{2}:ax+by\in B(z,r)\}$ for some $(a,b)\in S^{1}.$ If $|b|\geq|a|$ then $|b|\geq 1/\sqrt{2}$ so we can bound
-[…]
+$P_{v}(\mu_{2}\times\nu)(B(z,r))=\int\nu\left(B\left(\frac{z-ax}{b},\frac{r}{b}\right)\right)d\mu_{2}(x)\leq\int C\left(\frac{r}{b}\right)^{s}d\mu_{2}(x)\leq 2Cr^{s}$
 similarly, if $|a|>|b|$ then $|a|>1/\sqrt{2}$ we compute $P_{v}(\mu\times\nu)(B(z,r))$ as
-[…]
+$P_{v}(\mu_{2}\times\nu)(B(z,r))=\int\mu_{2}\left(B\left(\frac{z-by}{a},\frac{r}{a}\right)\right)d\nu(y)\leq\int C\left(\frac{r}{a}\right)^{s}d\nu(y)\leq 2Cr^{s}$
 Therefore, $(\mu_{2}\times\nu)(T)\leq 2Cr^{s}$ for any tube $T$ of width $r$ which immediately gives that $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has $(s,2C,1,0)$-strong thin tubes. ∎
 Next we introduce a particularly useful a relaxation of the notion of an $(s,C)$-Frostman measure.
-In a metric space $X,$ we will call a measure $\mu\in\mathcal{P}(X)$ a $(\delta,s,C)$-measure if $\mu(B_{r})\leq C\cdot r^{s}$ for all $r\in[\delta,1].$ Given $t\in[0,d]$, $s\in[0,k],$ a function $D:[0,d]\to[0,k],$ and $C\geq 1$, we say that a measure $\rho\in\mathcal{P}(\mathbb{G}({\mathbb{R}}^{d},k))$ is $(D;C)$-adapted at scale $\delta$ if whenever $\epsilon>0$ and $\mu$ is a $(\delta,s,\delta^{-\epsilon})$-measure on $[0,1]^{d}$, then for all $s\in[0,d],$ we have
+In a metric space $X,$ we will call a measure $\mu\in\mathcal{P}(X)$ a $(\delta,s,C)$-measure if $\mu(B_{r})\leq C\cdot r^{s}$ for all $r\in[\delta,1].$ Given $t\in[0,d]$, $s\in[0,k],$ a function $D:[0,d]\to[0,k],$ and $C\geq 1$, we say that a measure $\rho\in\mathcal{P}(\mathbb{G}({\mathbb{R}}^{d},k))$ is $(D;C)$-adapted at scale $\delta$ if whenever $\epsilon>0$ and $\mu$ is a $(\delta,s,\delta^{-\epsilon})$-measure on $0,1]^{d}$, then for all $s\in[0,d],$ we have
 $\rho\{V\in\mathbb{G}(\mathbb{R}^{d},k):P_{V}\mu\text{ is not }(\delta,D(s)-C\epsilon,\delta^{\epsilon})\text{-robust }\}\leq\delta^{\epsilon/C}.$
@@
 If $\delta<K^{-C^{\prime}\epsilon^{-1}}$ for some absolute constant $C^{\prime}$, then $\eta(s,t)$ can be chosen so that $\rho$ is $(D_{s}-\epsilon;1)$-adapted at scale $\delta$.
-For a nonnegative number $\sigma,$ a function $f:[0,1]\rightarrow\mathbb{R}$ will be called $\sigma$-superlinear on $[a,b]\subseteq[0,1]$, if for all $x\in[a,b],$ we have that $f(x)\geq f(a)+\sigma\cdot(x-a).$ An interval $[a,b]\subseteq[0,1]$ is $\tau$-allowable if $\tau\leq b-a\leq a.$ Moreover, given $[a,b]\subseteq[0,1]$, we say that a collection of pairwise disjoint intervals $[a_{j},b_{j}]\subseteq[a,b]$ is $\tau$-allowable if each $[a_{j},b_{j}]$ is $\tau$-allowable.
-Now, given $[a,b]\subseteq[0,1]$, define
+For a nonnegative number $\sigma,$ a function $f:[0,1]\rightarrow\mathbb{R}$ will be called $\sigma$-superlinear on $a,b]\subseteq[0,1]$, if for all $x\in[a,b],$ we have that $f(x)\geq f(a)+\sigma\cdot(x-a).$ An interval $a,b]\subseteq[0,1]$ is $\tau$-allowable if $\tau\leq b-a\leq a.$ Moreover, given $a,b]\subseteq[0,1]$, we say that a collection of pairwise disjoint intervals $a_{j},b_{j}]\subseteq[a,b]$ is $\tau$-allowable if each $a_{j},b_{j}]$ is $\tau$-allowable.
+Now, given $a,b]\subseteq[0,1]$, define
 $\Sigma_{\tau}(D;f;[a,b],d):=\sup\left\{\sum_{j=1}^{J}(b_{j}-a_{j})D(\sigma_{j})\right\},$
-where the supremum is taken over all $\tau$-allowable collections $\{[a_{j},b_{j}]\}$ of sub-intervals of $[a,b]$ and numbers $\sigma_{j}\in[0,d]$ such that $f$ is $\sigma_{j}$-superlinear on $[a_{j},b_{j}]$.
+where the supremum is taken over all $\tau$-allowable collections $\{[a_{j},b_{j}]\}$ of sub-intervals of $[a,b]$ and numbers $\sigma_{j}\in[0,d]$ such that $f$ is $\sigma_{j}$-superlinear on $a_{j},b_{j}]$.
 Given $u\in[0,d]$, define $\mathcal{L}_{d,u}$ to be the set of all non-decreasing Lipschitz functions $f:[0,1]\to[0,d]$ with Lipschitz constant $\leq d$ that are piecewise linear, with $f(0)=0$ and $f(x)\geq ux$ for all $x\in[0,1]$.
@@
 There are a set $Y_{0}\subset Y$ with $\nu(Y_{0})>\max(1-2c^{1/2},0)$, and for each $y\in Y_{0}$ there is a set $X_{y}\subset X^{\prime}$ with $\mu(X_{y})>\max\{1-2c^{1/2},0\}$, such that
-[…]
+$F_{y}(\mu_{X_{y}})\text{ is }(2^{-m},\Sigma_{\tau}(D;s,d)-O(\epsilon^{1/2}),2^{-\eta m/2})\text{-robust, provided that},$
 $\eta\leq\frac{\epsilon\tau}{2},~\frac{1}{\tau}\leq\epsilon m,~2^{3\epsilon m}\geq\delta_{1}^{-s},\quad 2^{\epsilon m}\geq 1,~2^{3\epsilon m}>C_{\mu},\text{ and}$
@@
 Lemma 5.12.
-[Variant of Lemma 5.18 in [20]] Fix $u\in(0,1]$ and $u_{0}\in(0,u/3)$, $\zeta>0,C>1$ then there is an $\eta$ depending on $\zeta$ and $u$ such that if $s\in[u_{0},\varphi(u)-\zeta]$ then we have the following: If $\mu_{1},\mu_{2}$ are $(u,C)$-Frostman measures in $[0,1]^{2}$ with $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\geq C_{\mu}^{-1}$ and $(\mu_{1},\mu_{2})$ has $(s,K,1-c)$-strong thin tube for some number $K>0,$ then $(\mu_{1},\mu_{2})$ has $(s+\eta,C^{\prime}(KC_{\mu})^{C^{\prime}},1-9c^{1/2},0)$-strong thin tubes for some constant $C^{\prime}$ only depending on $u$ and $\zeta$.
+[Variant of Lemma 5.18 in [24]] Fix $u\in(0,1]$ and $u_{0}\in(0,u/3)$, $\zeta>0,C>1$ then there is an $\eta$ depending on $\zeta$ and $u$ such that if $s\in[u_{0},\varphi(u)-\zeta]$ then we have the following: If $\mu_{1},\mu_{2}$ are $(u,C)$-Frostman measures in $0,1]^{2}$ with $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\geq C_{\mu}^{-1}$ and $(\mu_{1},\mu_{2})$ has (s,K,1c)(s,K,1-c)-strong thin tube for some number $K>0,$ then $(\mu_{1},\mu_{2})$ has $(s+\eta,C^{\prime}(KC_{\mu})^{C^{\prime}},1-9c^{1/2},0)$-strong thin tubes for some constant $C^{\prime}$ only depending on $u$ and $\zeta$.
 Proof.
@@
 and furthermore, $Y=\{(x,y):x\in X,y\in Y_{x}\}$ is compact.
-For some constant $C_{0},$ only depending on the size of the support of $\mu_{1}$ and $\mu_{2}$, we have that by Proposition 6.10 and the method of constructing $D_{s},$ the measure $\mu\circ V_{x}$ is $(t\to D_{s}(t)-\eta)$-adapted at any scales in the range $[0,(K/C_{0})^{-C^{\prime}\eta^{-1}}]$ for any $\eta$ that is small enough.
+For some constant $C_{0},$ only depending on the size of the support of $\mu_{1}$ and $\mu_{2}$, we have that by Proposition 5.10 and the method of constructing $D_{s},$ the measure $\mu\circ V_{x}$ is $(t\to D_{s}(t)-\eta)$-adapted at any scales in the range $0,(K/C_{0})^{-C^{\prime}\eta^{-1}}]$ for any $\eta$ that is small enough.
 The fact that $\Sigma_{\tau}(D_{s};t,d)$ is continuous as a function in $t$ follows by Lemma 5.8, so by applying Proposition 5.11, we obtain that if $\eta$ is small enough in terms of $u$ and $\zeta,$ then
@@
 In particular, we begin by applying Proposition 5.6 to the measures $\mu_{1}\times\nu$ and $\mu_{2}\times\nu,$ to see that they have $(s,2C,1,0)$-strong thin tubes.
-Using Lemma 6.12 with $C_{\mu}=C_{d}C^{\frac{1}{s}},$ we obtain $\eta$ and $C^{\prime}$ only depending on $s$ and $\zeta$ such that if $t<(\varphi(2s)-\zeta/M)$, for some constant $M$ only depending on $s$ and $\zeta,$ which we choose later, and for some $K$ also chosen later, if $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has $(t,K,1-c,0)$-strong thin tubes then it also has $(t+\eta,C^{\prime}\min\{c,K\}^{C^{\prime}},1-9c^{1/2},0)$-strong thin tubes.
+Using Lemma 5.12 with $C_{\mu}=C_{d}C^{\frac{1}{s}},$ we obtain $\eta$ and $C^{\prime}$ only depending on $s$ and $\zeta$ such that if $t<(\varphi(2s)-\zeta/M)$, for some constant $M$ only depending on $s$ and $\zeta,$ which we choose later, and for some $K$ also chosen later, if $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has (t,K,1c,0)(t,K,1-c,0)-strong thin tubes then it also has $(t+\eta,C^{\prime}\min\{c,K\}^{C^{\prime}},1-9c^{1/2},0)$-strong thin tubes.
 Since $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has $(s,2C,1,0)$-strong thin tubes, then it also has $(s,3C,1-c_{0},0)$-strong thin tubes, where
@@
 and $(\rho_{y})_{X^{\prime}_{y}}(B(x,r))<r^{\varphi(2s)-\zeta/2}$ for every $x<\delta_{0}$, for simplicity. This implies that
-[…]
+$(\rho_{y})_{X_{y}^{\prime}}\text{ is }(\varphi(2s)-\zeta/2,\delta_{0}^{-1})\text{-Frostman}$ (57)
 Define the map $f(x,y,z,t)=\sqrt{(x-z)^{2}+(y-t)^{2}}$ and the measure $\kappa$ supported in $\mathbb{R}^{4}$ as $\mu_{1}\times\nu\times\mu_{2}\times\nu$.
@@
 ∎
-7 Proof of Theorem 1.2
-Since $H^{s}(A)>0$, by the Frostman lemma [12, Theorem 2.7], there exists a probability $(s,1)$-Frostman measure $\mu$ supported on $A$.
+6 Proof of Theorems 1.5-1.7
+Proof of Theorem 1.5.
+Since $H^{s}(A)>0$, by the Frostman lemma [13, Theorem 2.7], there exists a probability $(s,1)$-Frostman measure $\mu$ supported on $A$.
 Let $X$ and $Y$ be i.i.d.
 random variables that both have distribution measure $\mu$.
-Lemma 5.1 gives us
+Theorem 1.2 gives us
 $\max\big{\{}H_{n}(X+Y),\,H_{n}(\phi(X,Y))\big{\}}\geq(1+\epsilon)ns$
@@
 $\log\#\{D_{n}(x)\mid x\in\phi(A,A)\}\geq H_{n}(\phi(X,Y)).$
-Let $\mathcal{N}_{2^{-n}}(A+A)$ and $\mathcal{N}_{2^{-n}}(\phi(A,A))$ be the least number of intervals of length $2^{-n}$ needed to cover $A+A$ and $\phi(A,A)$, respectively. It is clear that
-$\log\mathcal{N}_{2^{-n}}(A+A)\geq H_{n}(X+Y),~{}~{}\log\mathcal{N}_{2^{-n}}(\phi(A,A))\geq H_{n}(\phi(X,Y)).$
+Recall that $\mathcal{E}_{2^{-n}}(A+A)$ and $\mathcal{E}_{2^{-n}}(\phi(A,A))$ are the least numbers of intervals of length $2^{-n}$ needed to cover $A+A$ and $\phi(A,A)$, respectively.
+It is clear that
+$\log\mathcal{E}_{2^{-n}}(A+A)\geq H_{n}(X+Y),~~\log\mathcal{E}_{2^{-n}}(\phi(A,A))\geq H_{n}(\phi(X,Y)).$
 Hence,
-$\max\left(\frac{\log\mathcal{N}_{2^{-n}}(A+A)}{n},\frac{\log\mathcal{N}_{2^{-n}}(\phi(A,A))}{n}\right)$
-$\quad\geq\max\left(\frac{H_{n}(X+Y)}{n},\frac{H_{n}(\phi(X,Y))}{n}\right)\geq(1+\epsilon)s.$
-When $n\to\infty$, one obtains
-$\max(\underline{\dim}_{B}(A+A),\underline{\dim}_{B}(\phi(A,A)))\geq(1+\epsilon)s.$
-This completes the proof.
-Appendix A Appendix - Examples
-In this appendix, we show details of some examples. Sometimes, we will write $\iota(x)=-x\log x$ for convenience.
-Example A.1.
-Given $0\leq s\leq 1$, and given $n$ sufficiently large, we will construct an $(s,O(1))$-Frostman random variable $X$ with $range(X)\subseteq[0,1)$, and compare $H_{n}(X)$, $H_{n}(X+Y)$, $H_{n}(X^{2}+Y^{2})$ and $H_{n}(XY)$, where $Y$ is an i.i.d.
-copy of $X$.
-Let $m=m(n)\in\mathbb{N}$ and $l=l(n)\in\mathbb{N}\cup\{0\}$ be parameters to be chosen later. Take
+$\max\left\{\frac{\log\mathcal{E}_{2^{-n}}(A+A)}{n},\frac{\log\mathcal{E}_{2^{-n}}(\phi(A,A))}{n}\right\}\geq\frac{1}{n}\max\left\{H_{n}(X+Y),H_{n}(\phi(X,Y))\right\}\geq(1+\epsilon)s.$
+When $n\to\infty$, we obtain
+$\max\{\underline{\dim}_{B}(A+A),\,\overline{\dim}_{B}(\phi(A,A))\}\geq(1+\epsilon)s,$
+and
+$\max\{\overline{\dim}_{B}(A+A),\,\underline{\dim}_{B}(\phi(A,A))\}\geq(1+\epsilon)s.$
+This completes the proof. ∎
+Proof of Theorem 1.6.
+We still choose a small number $\eta=\eta(\phi)$ with $0<\eta<1/10$ satisfying $A_{\eta\delta}+A_{\eta\delta}$ and $\phi(A_{\eta\delta},A_{\eta\delta})$ are contained in the $(A+A)_{\delta/10}$ and $(\phi(A,A))_{\delta/10}$, respectively, for any $\delta>0$.
+Since $A$ is $\delta$-separated, the Lebesgue measure of $A_{\eta\delta}$ is $\gg_{\phi}\delta^{1-s}$.
+Let $X$ be a random variable that has uniform distribution on $A_{\eta\delta}$.
+Combining the non-concentration condition, one deduces that
+$P(X\in B(x,r))\ll_{\phi}\begin{cases}r\delta^{s-1},\quad&\text{if }0<r\leq\delta,\\ Cr^{s},\quad&\text{if }\delta<r\leq 1,\\ 1,\quad&\text{if }r>1,\end{cases}$
+for any ball $B(x,r)$. In all the cases, the right-hand side can be bounded by $Cr^{s}$. So $X$ is $(s,O_{\phi}(C))$-Frostman.
+Let $Y$ be a random variable that is i.i.d.
+with $X$.
+Now similar relations as in (58), (61) and (62) hold.
+As a replacement of (60), we apply Theorem 1.2 to get an $\epsilon^{\prime}=\epsilon^{\prime}(s,\phi)$ such that
+$\max\{H_{n}(X+Y),\,H_{n}(\phi(x,y))\}\geq n(s+\epsilon^{\prime})+O_{\phi,s}(\log C)+O_{\phi,s}(1).$
+for all $n\in\mathbb{N}$. Putting everything together, we arrive at the conclusion that
+$\max\{\mathcal{E}_{\delta}(A+A),\,\mathcal{E}_{\delta}(\phi(A,A))\}\gg_{\phi,s}C^{-O_{\phi,s}(1)}\delta^{-\epsilon}(\#A)$
+for all $\delta>0$. ∎
+Proof of Theorem 1.7.
+Let $\epsilon=\epsilon(s)$ to be determined later.
+For any sets $E\subseteq{\mathbb{R}}$, $F\subseteq{\mathbb{R}}^{2}$, and any positive number $\lambda$, denote by $E_{\lambda}$ and $F_{\lambda}$ the rectangular $\lambda$-neighbourhood of $E$ and $F$, respectively.
+More concretely, $E_{\lambda}=\bigcup\nolimits_{x\in E}(x-\lambda,x+\lambda)$ and $F_{\lambda}=\bigcup\nolimits_{(x,y)\in F}(x-\lambda,x+\lambda)\times(y-\lambda,y+\lambda)$.
+Note that $f_{1}(x,y)=x+y$ and $f_{2}(x,y)=\phi(x,y)=\alpha x^{2}+\beta y^{2}$ are $O_{\phi}(1)$-Lipchitz on $0,1]^{2}$.
+We can choose a small number $\eta=\eta(\phi)$ with $0<\eta<1/10$ such that $A_{\eta\delta}+_{G}A_{\eta\delta}$ and $\alpha\cdot A_{\eta\delta}^{\wedge 2}+\beta\cdot A_{\eta\delta}^{\wedge 2}$ are contained in the $(A+_{G}A)_{\delta/10}$ and $(\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2})_{\delta/10}$, respectively, for any $\delta>0$.
+Then
+$\begin{split}\mathcal{E}_{\delta}(A+_{G}A)&\geq\frac{1}{2}\mathcal{E}_{\delta}(A_{\eta\delta}+_{G}A_{\eta\delta}),\\ \mathcal{E}_{\delta}(\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2})&\geq\frac{1}{2}\mathcal{E}_{\delta}(\alpha\cdot A_{\eta\delta}^{\wedge 2}+_{G}\beta\cdot A_{\eta\delta}^{\wedge 2}).\end{split}$ (58)
+Moreover, since $A$ is $\delta$-separated, we have
+$\mathcal{E}_{\delta}(A_{\eta\delta})=\mathcal{E}_{\delta}(A)=\#A=\lceil\delta^{-s}\rceil.$ (59)
+Now consider the set $G_{\eta\delta}\subseteq A_{\eta\delta}\times A_{\eta\delta}$.
+Since $A$ is $\delta$-separated, and $|G|\geq\delta^{\epsilon}|A|^{2}\geq\delta^{-2s+\epsilon}$, the $2$-dimensional Lebesgue measure of $G_{\eta\delta}$ is $\gg\eta^{2}\delta^{2-2s+\epsilon}\gg_{\phi}\delta^{2-2s+\epsilon}$.
+Let $(X,Y)$ be a random variable that has uniform distribution on $G_{\eta\delta}$. Combining the non-concentration condition, one can verify that
+[…]
+for any pair of balls $B(x,r_{1})$ and $B(y,r_{2})$.
+In all the cases, the right-hand side can be bounded by $C^{2}\delta^{-\epsilon}r_{1}^{s}r_{2}^{s}$.
+So $(X,Y)$ is jointly $(s,s,C^{2}\delta^{-\epsilon})$-Frostman.
+By Theorem 1.4, there exists an $\epsilon^{\prime}=\epsilon^{\prime}(s)$ such that
+$\begin{split}&\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\\ &\qquad\qquad\geq n(s+\epsilon^{\prime})+O_{s}(\log_{+}C)+O_{s}(\epsilon\log(1/\delta))+O_{\phi,s}(1).\end{split}$ (60)
+for all $n\in\mathbb{N}$. In particular, we take $n$ such that $2^{-n-1}<\delta\leq 2^{-n}$.
+By Lemma 2.4(ii), it follows that
+$\begin{split}&\log\#\,range_{+}(D_{n}(X))\geq H_{n}(X),\quad\log\#\,range_{+}(D_{n}(Y))\geq H_{n}(Y)\\ &\qquad\qquad\quad\log\#\,range_{+}(D_{n}(X+Y))\geq H_{n}(X+Y),\\ &\qquad\quad\log\#\,range_{+}(D_{n}(\alpha x^{2}+\beta Y^{2})\big{\}}\geq H_{n}(\alpha X^{2}+\beta Y^{2}).\end{split}$ (61)
+Moreover,
+$\begin{split}&\quad\mathcal{E}_{\delta}(A_{\eta\delta})\geq\mathcal{E}_{2^{-n}}(A_{\eta\delta})\geq\frac{1}{2}\max\left\{\#\,range_{+}(D_{n}(X)),\,\#\,range_{+}(D_{n}(Y))\right\},\\ &\qquad\;\;\mathcal{E}_{\delta}(A_{\eta\delta}+_{G}A_{\eta\delta})\geq\mathcal{E}_{2^{-n}}(A_{\eta\delta}+_{G}A_{\eta\delta})\geq\frac{1}{2}\#\,range_{+}(D_{n}(X+Y)),\\ &\mathcal{E}_{\delta}(\alpha\cdot A_{\eta\delta}^{\wedge 2}+_{G}\beta\cdot A_{\eta\delta}^{\wedge 2})\geq\mathcal{E}_{2^{-n}}(\alpha\cdot A_{\eta\delta}^{\wedge 2}+_{G}\beta\cdot A_{\eta\delta}^{\wedge 2})\geq\frac{1}{2}\#\,range_{+}(\alpha X^{2}+\beta Y^{2}).\end{split}$ (62)
+Putting (58)-(62) together, one arrives at
+$\max\{\mathcal{E}_{\delta}(A),\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\alpha\cdot A^{2}+_{G}\beta\cdot A^{2})\}$
+$\qquad\gg 2^{n(s+\epsilon^{\prime})+O_{s}(\log_{+}C)+O_{s}(\epsilon\log(1/\delta))+O_{\phi,s}(1))}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-(s+\epsilon^{\prime}-O_{s}(\epsilon))}.$ (63)
+Now we can take $\epsilon=\epsilon(s)>0$ to be sufficiently small so that $\epsilon^{\prime}-O_{s}(\epsilon)\geq\epsilon$ on the right-hand side of (63). Then
+$\max\{\mathcal{E}_{\delta}(A),\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\alpha\cdot A^{2}+_{G}\beta\cdot A^{2})\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-(s+\epsilon)}$
+for any $\delta>0$.
+Finally, by (59) one has $\mathcal{E}_{\delta}(A)\ll\delta^{-s}$. This gives
+$\max\{\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\alpha\cdot A^{2}+_{G}\beta\cdot A^{2})\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A)$
+whenever $\delta$ is sufficiently small depending on $\phi,s$ and $C$. The proof is completed. ∎
+7 Examples
+In this section, we show details of some examples. Sometimes, we will write $\iota(x)=-x\log x$ for convenience.
+Example 7.1.
+Given $0\leq s\leq 1$, and given $n$ sufficiently large, we will construct i.i.d.
+random variables $X$ and $Y$ whose ranges are contained in $[0,1]$, such that $(X,Y)$ is marginally $(s,O(1);s,O(1))$-Frostman, and compare $H_{n}(X)$, $H_{n}(X+Y)$, $H_{n}(X^{2}+Y^{2})$ and $H_{n}(XY)$.
+Indeed, let $m=m(n)\in\mathbb{N}$ and $l=l(n)\in\mathbb{N}\cup\{0\}$ be parameters to be chosen later. We suppose that $(X,Y)$ has the uniform distribution on $A\times A$, with
 $A=\bigcup\limits_{j=0}^{2^{m}-1}\left[j2^{-m},\,j2^{-m}+2^{-m-l}\right).$
-Denote by $\mathcal{L}^{d}$ the $d$-dimensional Lesbegue measure. Then $\mathcal{L}^{1}(A)=2^{-l}$. Let $X$ be a random variable having uniform distribution on $A$, i.e.,
+Denote by $\mathcal{L}^{d}$ the $d$-dimensional Lesbegue measure. Then $\mathcal{L}^{1}(A)=2^{-l}$. We see that both $X$ and $Y$ have Note that $X$ has uniform distribution on $A$, i.e.,
 $P(X\in E)=\frac{\mathcal{L}^{1}(E\cap A)}{\mathcal{L}^{1}(A)}$
@@
 In the following, we will proof the following results.
-Claim 1. The random variable $X$ is $\left(s,O(1)\right)$-Frostman whenever $0\leq s\leq\frac{m}{l+m}$.
+Claim 1. The random variable $(X,Y)$ is marginally $\left(s_{1},O(1);s_{2},O(1)\right)$-Frostman whenever $0\leq s_{1},s_{2}\leq\frac{m}{l+m}$.
 Claim 2. We have
@@
 $\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}\leq(2+o(1))m=o(n)$
-when $l/m$ tends to infinity. For $l=0$ and $s=1$, we assert by Claim 2 that $H_{n}(X)=n=ns$. Combining Corollary 2.22, Lemma 2.36 and (11), we conclude that
+when $l/m$ tends to infinity. For $l=0$ and $s=1$, we assert by Claim 2 that $H_{n}(X)=n=ns$. Combining (12), (13) and Lemma 2.34, we conclude that
 $\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}=(1+o(1))ns.$
@@
 By Lemma 2.4 (ii) and independence, one sees that
-$H(X^{\prime}+Y^{\prime})\leq\log|range_{+}(X^{\prime}+Y^{\prime})|=m+O(1),$
+$H(X^{\prime}+Y^{\prime})\leq\log\#\,range_{+}(X^{\prime}+Y^{\prime})=m+O(1),$
 and
@@
 Proof of Claim 4.
-Recall that $X^{\prime}=D_{m+l}(X)$ and $Y^{\prime}=D_{m+l}(Y)$. Since the function $(x,y)\mapsto x^{2}+y^{2}$ on $[0,1]^{2}$ is $O(1)$-Lipschitz, we obtain by Lemma 2.19 that
+Recall that $X^{\prime}=D_{m+l}(X)$ and $Y^{\prime}=D_{m+l}(Y)$. Since the function $(x,y)\mapsto x^{2}+y^{2}$ on $0,1]^{2}$ is $O(1)$-Lipschitz, we obtain by Lemma 2.18 that
 $H_{n}(X^{2}+Y^{2})=H_{n}(X^{\prime 2}+Y^{\prime 2})+O(1)$
@@
 $=\sum\limits_{k=0}^{2^{2m+1}}\iota\Big{(}P(X^{\prime 2}+Y^{\prime 2}=k2^{-2m})\Big{)}=\sum\limits_{k\in\mathcal{K}}\iota\big{(}2^{-2m}T(k)\big{)},$
-where $\mathcal{K}$ denotes the set consisting of integers $k$ such that $k=j_{1}^{2}+j_{2}^{2}$ for some $10\leq j_{1},j_{2}\leq 2^{m}-$ 1, and
+where $\mathcal{K}$ denotes the set consisting of integers $k$ such that $k=j_{1}^{2}+j_{2}^{2}$ for some $0\leq j_{1},j_{2}\leq 2^{m}-1$, and
 $T(k)=\sum\limits_{0\leq j_{1},j_{2}\leq 2^{m}-1\atop j_{1}^{2}+j_{2}^{2}=k}1.$
@@
 Proof of Claim 5.
-Since the function $(x,y)\mapsto xy$ on $[0,1]^{2}$ is $O(1)$-Lipschitz, we obtain by Lemma 2.19 that
+Since the function $(x,y)\mapsto xy$ on $0,1]^{2}$ is $O(1)$-Lipschitz, we obtain by Lemma 2.18 that
 $H_{n}(XY)=H_{n}(X^{\prime}Y^{\prime})+O(1)$
@@
 $H_{n}(X^{\prime}Y^{\prime})=\sum\limits_{k\in\mathcal{K}^{\prime}}\iota\big{(}2^{-2m}T^{\prime}(k)\big{)},$
-where $\mathcal{K}^{\prime}$ denotes the set consisting of integers $k$ such that $k=j_{1}j_{2}$ for some $10\leq j_{1},j_{2}\leq 2^{m}-$ 1, and
+where $\mathcal{K}^{\prime}$ denotes the set consisting of integers $k$ such that $k=j_{1}j_{2}$ for some $0\leq j_{1},j_{2}\leq 2^{m}-1$, and
 $T^{\prime}(k)=\sum\limits_{0\leq j_{1},j_{2}\leq 2^{m}-1\atop j_{1}j_{2}=k}1.$
@@
 Now, similar arguments as in the proof of Claim 4 leads to the conclusion that $H_{n}(XY)=(2+o(1))m$. ∎
-Example A.2.
-Let $X$ and $Y$ be the independent $(s,O(1))$-Frostman random variables defined in A.1, with $n=m+l$ and $s=\frac{m}{l+m}$.
+Example 7.2.
+Let $X$ and $Y$ be the independent $(s,O(1))$-Frostman random variables defined in 7.1, with $n=m+l$ and $s=\frac{m}{l+m}$.
 Consider $\tilde{X}=\sqrt{X}$ and $\tilde{Y}=\sqrt{Y}$.
-By Example A.1, one sees that
+By Example 7.1, one sees that
 $H_{n}(\tilde{X}^{2}+\tilde{Y}^{2})=H_{n}(X+Y)=(1+o(1))ns.$
 Therefore, the term involving maximum on the left-hand side of (2) can not be replaced by a single term $H_{n}(\phi(X,Y))$.
-Example A.3.
-Let $X$, $Y$ be the independent $(s,O(1))$-Frostman random variables defined in A.1, with $n=m+l$ and $s=\frac{m}{l+m}$.
+Example 7.3.
+Let $X$, $Y$ be the independent $(s,O(1))$-Frostman random variables defined in Example 7.1, with $n=m+l$ and $s=\frac{m}{l+m}$.
 Consider $\tilde{X}=1+X$, which is $(s,O(1))$-Frostman with range contained in $[1,2)$.
@@
 Take $\phi(x,y)=(x+y)^{2}$, which is a degenerate quadratic form.
-By Example A.1, one can deduce that
+By Example 7.1, one can deduce that
 $H_{n}(\tilde{X})=H_{n}(Y)=ns,\qquad H_{n}(\tilde{X}+Y)=(1+o(1))ns.$
@@
 So, the requirement that $\phi$ is non-degenerate in Theorem 1.2 is essential.
-Example A.4.
+Example 7.4.
 Assume that the random variables $X$ and $Y$ are independent, $X$ has uniform distribution on the set
 $A=\bigcup\limits_{j=0}^{2^{m}-1}\left[1+j2^{-m},\,1+j2^{-m}+2^{-2m}\right),$
-and $Y$ has uniform distribution on $[0,2^{-2m}]$. One can verify that $X$ is $(1/2,O(1))$-Frostman, and $Y$ is $(1/2,O(2^{m}))$-Frostman. We take $s=1/2$ and consider $n=2m$.
+and $Y$ has uniform distribution on $0,2^{-2m}]$. One can verify that $X$ is $(1/2,O(1))$-Frostman, and $Y$ is $(1/2,O(2^{m}))$-Frostman. We take $s=1/2$ and consider $n=2m$.
 Claim. We have $\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}=ns+O(1)$.
@@
 The conclusion follows by noting that $H_{n}(X^{\prime})=m=ns$. ∎
-Example A.5.
-This example is an variant of Example A.4.
+Example 7.5.
+This example is an variant of Example 7.4.
 Let $W=(X,Y)$, and suppose that the joint distribution of $(X,Y)$ is the uniform distribution on the set $A=\bigcup\nolimits_{j=0}^{2^{m}-1}\big{[}j2^{-m},\,j2^{-m}+2^{-2m}\big{)}^{2}$.
@@
 Moreover, for the random variable $W$, its Frostman constant $C$ can grow with $m$.
-If we consider $n=2m$ in Theorem 2.32, if $21<s<3/$ 2, then one has $H_{n}(|W_{1}-W_{2}|)=(1+o(1))n/2<nt$ for some suitably chosen $t$ with $s/2<t<s-1/2$.
-That is to say, the term $-\log C$ in Theorem 2.32 is essential.
+If we consider $n=2m$ in Theorem 2.30, if $1<s<3/2$, then one has $H_{n}(|W_{1}-W_{2}|)=(1+o(1))n/2<nt$ for some suitably chosen $t$ with s/2<t<s1/2s/2<t<s-1/2.
+That is to say, the term $\log C$ in Theorem 2.30 is essential.
 Proof of Claim 1.
@@
 $H_{n}\big{(}|(X_{1}^{\prime},Y_{1}^{\prime})-(X_{2}^{\prime},Y_{2}^{\prime})|\big{)}=H_{n}(\sqrt{2}|X_{1}^{\prime}-X_{2}^{\prime}|)=H_{m}(|X_{1}^{\prime}-X_{2}^{\prime}|)+O(1)=(1+o(1))m.$
-Example A.6.
+Example 7.6.
+This example is a further variant of Examples 7.4 and 7.5.
+Let $W=(X,Y)$, and suppose that the joint distribution of $(X,Y)$ is the uniform distribution on the set $A=\bigcup\nolimits_{j=0}^{2^{m}-1}\big{(}j2^{-m},\,j2^{-m}+2^{-2m}\big{)}\times\big{(}-j2^{-m},\,-j2^{-m}+2^{-2m}\big{)}$.
+Then the marginal distributions of $X$ and $Y$ are the uniform distribution on the sets $A_{0}=\bigcup\nolimits_{j=0}^{2^{m}-1}\big{(}j2^{-m},\,j2^{-m}+2^{-2m}\big{)}$ and $A_{0}^{\prime}=\bigcup\nolimits_{j=0}^{2^{m}-1}\big{(}-j2^{-m},\,-j2^{-m}+2^{-2m}\big{)}$, respectively.
+So both $X$ and $Y$ are $(1/2,O(1))$-Frostman.
+Claim 1. The random variable $(X,Y)$ is jointly $(1/4,1/4,O(1))$-Frostman, but is not conditionally $(s_{1},O(1);s_{2},O(2))$-Frostman for some $(s_{1},s_{2})\neq(0,0)$.
+Claim 2. Taking $n=2m$, one has $H_{n}(X+Y)=H_{n}(X^{2}-Y^{2})=O(1)$, and $H_{n}(X,Y)=m+O(1)$.
+In particular, if we apply Theorem 1.4 with $(X,Y)$, take $\alpha=-\beta$, $s_{1}=s_{2}=s=1/4$ and $n=2m$, then the second and third entropies in the maximum-term nearly vanish.
+So the first term $(H_{n}(X)+H_{n}(Y))/2$ is essential.
+In this example, it is $m+O(1)$, which is obviously larger than $ns=m/2$.
+Furthermore, we can not substitute $\frac{1}{2}H(X,Y)=m/2$ for $(H_{n}(X)+H_{n}(Y))/2$ as in Theorem 1.3.
+Proof of Claim 1.
+Note that $\mathcal{L}^{2}(A)=2^{-3m}$ and $\mathcal{L}^{1}(A_{0})=2^{-m}$. Let $E$ be a Borel set and $B(y,r)$ be any ball. When $0<r<2^{-2m}$, we have
+$P(X\in E,Y\in B(y,r))\ll\mathcal{L}^{1}(E\cap A_{0})\cdot r\cdot 2^{3m},$
+while
+$P(X\in E)\leq\mathcal{L}^{1}\cdot 2^{m}.$
+It follows that
+$P(X\in E,Y\in B(y,r))\ll r^{1/2}P(X\in E).$
+For any pair of balls $B(x,r_{1})$ and $B(y,r_{2})$, one has
+$P(X\in B(x,r_{1}),\,Y\in B(y,r_{2}))\ll\begin{cases}r_{1}r_{2}2^{3m},\quad&\text{if }0<r_{1}\leq r_{2}\leq 2^{-2m},\\ r_{1}2^{m},\quad&\text{if }0<r_{1}\leq 2^{-2m}<r_{2},\\ 2^{-m},\quad&\text{if }2^{-2m}<r_{1}\leq 2^{-m},\,r_{2}\geq r_{1},\\ r_{1},\quad&\text{if }2^{-m}<r_{1}\leq 1,\,r_{2}\geq r_{1},\\ 1,\quad&\text{if }r_{2}\geq r_{1}>1,\end{cases}$
+and the other cases with $r_{2}<r_{1}$ can be obtained by symmetry. One can verify that all the terms on the right-hand side of above formula can be bounded by $r_{1}^{1/4}r_{2}^{1/4}$.
+For the conditional Frostman condition, we take $E=B(x,2^{-2m})$ and $F=B(y,2^{-2m})$ such that $E\times F\subseteq A$. Then
+$P(X\in E)=2^{-m}=P(X\in E,\,Y\in B(x,2^{-2m}))=2^{-m}.$
+The only possible positive $s_{2}$ satisfying $2^{-m}\leq 2^{-m}(2^{-2m})^{s_{2}}$ is 0. Now it is not hard to see that $(X,Y)$ is $(s_{1},O(1);s_{2},O(1))$-Frostman only when $(s_{1},s_{2})=(0,0)$. ∎
+Proof of Claim 2.
+We use similar arguments as in the proof of Claim 1 in Example 7.5.
+Let $(X^{\prime},Y^{\prime})=D_{2m}(X,Y)$.
+Then $(X^{\prime},Y^{\prime})$ is the discrete uniform distribution on $\{(j2^{-m},-j2^{-m}:\,j=0,1,\ldots,2^{m}-1\}$.
+So $H_{n}(X^{\prime},Y^{\prime})=m$.
+Moreover, $range(X^{\prime}+Y^{\prime})=range(X^{\prime 2}-Y^{\prime 2})=\{0\}$, which implies that $H_{n}(X^{\prime}+Y^{\prime})=H_{n}(X^{\prime 2}-Y^{\prime 2})=0$.
+Applying similar relation as in (66), one asserts that $H_{n}(X+Y)=H_{n}(X^{2}-Y^{2})=O(1)$.
+Similarly, $H_{n}(X,Y)=m+O(1)$.
+∎
+Example 7.7.
+Let
+$B=\bigcup\limits_{i=0}^{2^{m}-1}\bigcup\limits_{j=0}^{2^{m}-1}[i2^{-m}+j2^{-2m},i2^{-m}+(j+1)2^{-2m})\times[j2^{-m}+i2^{-2m},i2^{-m}+(i+1)2^{-2m}).$
+Then, for each given $x\in[0,1)$ (or $y\in[0,1)$), the set $\{y:\,(x,y)\in B\}$ (or $\{x:\,(x,y)\in B\}$, respectively) is exactly an interval of length $2^{-2m}$. Now Let
+$A=\bigcup\limits_{i^{\prime},j^{\prime}=0}^{2^{m}-1}(i2^{-m},j2^{-m})+2^{-2m}\cdot B.$
+Here $2^{-2m}\cdot B=\{(2^{-2m}x,2^{-2m}y):\,(x,y)\in B\}$. Suppose that $(X,Y)$ is a random variable that has uniform distribution on $A$.
+Claim 1. The random variable $(X,Y)$ is marginally $(1/2,O(1);1/2,O(1))$-Frostman.
+Claim 2. The random variable $(X,Y)$ is conditionally $(1/4,O(1);1/4,O(1))$-Frostman.
+Claim 3. The random variable $(X,Y)$ is jointly $(1/2,1/2,O(1))$-Frostman.
+Proof of Claim 1.
+One can check that both $X$ and $Y$ have the uniform distribution on the set $A_{0}=\bigcup\nolimits_{j=0}^{2^{m}-1}\left[j2^{-m},\,j2^{-m}+2^{-m-l}\right)$.
+By Example 7.1, they are $(1/2,O(1))$-Frostman.
+∎
+Proof of Claim 2.
+Note that $\mathcal{L}^{2}(A)=(2^{-4m})^{2}\cdot(2^{m})^{2}\cdot(2^{m})^{2}=2^{-4m}$. And $\mathcal{L}^{1}(A_{0})=2^{-m}$. Let $E$ be any Borel set, and $B(y,r)$ be any ball. Calculation reveals that
+$P(X\in E)=\mathcal{L}^{1}(E\cap A_{0})\cdot 2^{m},$
+and
+$P(X\in E,y\in B(y,r))\ll\begin{cases}\mathcal{L}^{1}(E\cap A_{0})\cdot r\cdot 2^{4m},\quad&\text{if }0<r\leq 2^{-4m},\\ \mathcal{L}^{1}(E\cap A_{0})\cdot 2^{-4m}\cdot 2^{4m},\quad&\text{if }2^{-4m}<r\leq 2^{-m},\\ \mathcal{L}^{1}(E\cap A_{0})\cdot 2^{-4m}\frac{r}{2^{-m}}\cdot 2^{4m},\quad&\text{if }2^{-m}<r\leq 1,\\ \mathcal{L}^{1}(E\cap A_{0})\cdot 2^{-4m}2^{m}\cdot 2^{4m},\quad&\text{if }r>1.\end{cases}$
+For all the cases, it satisfies that
+$P(X\in E,y\in B(x,r))\ll r^{1/4}P(X\in E).$
+Hence $Y$ is conditionally $(1/4,O(1))$-Frostman given $X$. Similarly, one can deduce that $X$ is conditionally $(1/4,O(1))$-Frostman given $Y$. Claim 2 then follows. ∎
+Proof of Claim 3.
+For any pair of balls $B(x,r_{1})$ and $B(y,r_{2})$, calculation reveal that
+$P(X\in B(x,y_{1}),y\in B(y,r_{2}))$
+[…]
+The rest cases with $r_{1}>r_{2}$ can be obtained by symmetry. In all cases, we have
+$P(X\in B(x,y_{1}),y\in B(y,r_{2}))\ll r_{1}^{1/2}r_{2}{1/2}.$
+The proof is completed. ∎
+Example 7.8.
 By Lemma 2.21, we have the following statement: Let $k,l\in\mathbb{Q}$ be given. Then
@@
 In particular, we consider the situation that $(k,l)=(1,\sqrt{2})$.
-Let $X,Y$ be the random variables defined in Example A.1, and $X^{\prime}=D_{m+l}(X)$, $Y^{\prime}=D_{m+l}(Y)$, where the parameters are chosen as the following: for any given number $\eta>0$, let $m$ be any sufficiently large natural number, $l$ be a natural number with $(1+\eta/2)m\leq l\leq(1+\eta)m$, and $n=m+l$.
-By Example A.1, the random variables $X$ and $Y$ are $(s,O(1))$-Frostman, where $s=\frac{m}{l+m}\in[\frac{1}{2+\eta},1/2)$.
+Let $X,Y$ be the random variables defined in Example 7.1, and $X^{\prime}=D_{m+l}(X)$, $Y^{\prime}=D_{m+l}(Y)$, where the parameters are chosen as the following: for any given number $\eta>0$, let $m$ be any sufficiently large natural number, $l$ be a natural number with $(1+\eta/2)m\leq l\leq(1+\eta)m$, and $n=m+l$.
+By Example 7.1, the random variables $X$ and $Y$ are $(s,O(1))$-Frostman, where $s=\frac{m}{l+m}\in[\frac{1}{2+\eta},1/2)$.
 Also, the equalities in (65) hold.
@@
 Then $|i|,|i^{\prime}|<2^{m}$ and $|i+\sqrt{2}i^{\prime}|<2^{-(1+\eta/2)m}$.
-Combining the fact that $(i,i^{\prime})\neq(0,0)$, one can deduce that neither $i$ nor $i^{\prime}$ can be 00.
+Combining the fact that $(i,i^{\prime})\neq(0,0)$, one can deduce that neither $i$ nor $i^{\prime}$ can be 0.
 However, the irrational $\sqrt{2}$ is an algebraic number of degree $2$. By Liouville’s approximation theorem, there exists an absolute positive constant $C$ such that
@@
 The proof is completed. ∎
-Appendix B Appendix - Existence of Certain Random Variable
+Appendix A Appendix - Existence of Certain Random Variable
 In this appendix, we prove two lemmas concerning the existence of certain random variables.
-Lemma B.1.
-Let $(A,B,C)$ be a given random variable.
-Let $h$ be a function such that $h(C)$ is a well-defined discrete random variable.
-Then there exists a random variable $(A^{\prime},B^{\prime},C^{\prime},B^{\prime\prime})$ such that $(A^{\prime},B^{\prime},C^{\prime})\equiv(A,B,C)$, $(B^{\prime\prime},h(C^{\prime}))\equiv(B,h(C))$, and $(B^{\prime\prime}|A^{\prime},B^{\prime},h(C^{\prime}))\equiv(B^{\prime\prime}|h(C^{\prime}))$.
-Moreover, $B^{\prime\prime}$ and $B^{\prime}$ are conditionally i.i.d.
-given $h(C^{\prime})$.
-Furthermore, $(A^{\prime},B^{\prime})$ and $B^{\prime\prime}$ are conditionally independent given $h(C^{\prime})$.
+Lemma A.1.
+Let $(X,Y,Z)$ be a given random variable.
+Let $h$ be a function such that $h(Z)$ is a well-defined discrete random variable.
+Then there exists a random variable $(X^{\prime},Y^{\prime},Z^{\prime},Y^{\prime\prime})$ such that $(X^{\prime},Y^{\prime},Z^{\prime})\equiv(X,Y,Z)$, $(Y^{\prime\prime},h(Z^{\prime}))\equiv(Y,h(Z))$, and $(Y^{\prime\prime}|X^{\prime},Y^{\prime},h(Z^{\prime}))\equiv(Y^{\prime\prime}|h(Z^{\prime}))$.
+Moreover, $Y^{\prime\prime}$ and $Y^{\prime}$ are conditionally i.i.d.
+given $h(Z^{\prime})$.
+Furthermore, $(X^{\prime},Y^{\prime})$ and $Y^{\prime\prime}$ are conditionally independent given $h(Z^{\prime})$.
 Proof.
-Suppose that $A,B,C$ are defined as measurable functions from a probability space $(\Omega,\mathcal{B},P)$ to some other measurable space.
+Suppose that $X,Y,Z$ are defined as measurable functions from a probability space $(\Omega,\mathcal{B},P)$ to some other measurable space.
 Now we construct a new probability space $(\Omega^{\prime},\mathcal{B}^{\prime},P^{\prime})$.
 Let
-$\Omega=\bigcup_{c\in range(h(C))}\Omega_{c},\quad\Omega_{c}:=\{h(C)=c\}=\{\omega\in\Omega:\,h(C(\omega))=c\}.$
+$\Omega=\bigcup_{z\in range(h(Z))}\Omega_{z},\quad\Omega_{z}:=\{h(Z)=z\}=\{\omega\in\Omega:\,h(Z(\omega))=z\}.$
 Construct
-$\Omega^{\prime}=\bigcup_{c}(\Omega_{c}\times\Omega_{c}).$
-Then $(\omega,\nu)\in\Omega^{\prime}$ if and only if $\omega,\nu\in\Omega_{c}$ for some $c$, if and only if $h(C(\omega))=h(C(\nu))$.
-The $\sigma$-algebra $\mathcal{B}^{\prime}$ can be generated by the related Cartisian products of Borel sets in $\mathcal{B}$, and the probability measure $P^{\prime}$ can be induced by the following: if $E\times F$ is a Borel subset of $\Omega_{c}\times\Omega_{c}$ for some $c$, then
-$P^{\prime}(E\times F)=\begin{cases}\frac{P(E)P(F)}{P(\Omega_{c})},\quad&\text{if }P(\Omega_{c})>0,\\ 0,&\text{if }P(\Omega_{c})=0.\end{cases}$
-In the following, we assume without loss of generality that $P(\Omega_{c})>0$ for all $c\in range(h(C))$.
-Now for any random variable $D$ defined on $(\Omega,\mathcal{B},P)$, let $D^{\prime}$ be the following random variable on $(\Omega^{\prime},\mathcal{B}^{\prime},P^{\prime})$:
-$D^{\prime}(\omega,\nu)=D(\omega),\quad(\omega,\nu)\in\Omega^{\prime}.$
-By this definition, one sees that $\big{(}f(D))^{\prime}=f(D^{\prime})$ for any admissible function $f$. In the following, we verify that $D^{\prime}\equiv D$. For any Borel set $L$,
-$P^{\prime}(D^{\prime}\in L)$ $=P^{\prime}(\{(\omega,\nu)\in\Omega^{\prime}:\,D^{\prime}(\omega,\nu)\in L)$
-$=\sum\limits_{c}P^{\prime}(\{(\omega,\nu)\in\Omega_{c}\times\Omega_{c}:\,D(\omega)\in L\})$
-$=\sum\limits_{c}\frac{P(D\in L,h(C)=c)P(h(C)=c)}{P(h(C)=c)}=P(D\in L).$
-In particular, we can take $D=(A,B,C)$ to get $(A^{\prime},B^{\prime},C^{\prime})\equiv(A,B,C)$.
+$\Omega^{\prime}=\bigcup_{z}(\Omega_{z}\times\Omega_{z}).$
+Then $(\omega,\nu)\in\Omega^{\prime}$ if and only if $\omega,\nu\in\Omega_{z}$ for some $z$, if and only if $h(Z(\omega))=h(Z(\nu))$.
+The $\sigma$-algebra $\mathcal{B}^{\prime}$ can be generated by the related Cartisian products of Borel sets in $\mathcal{B}$, and the probability measure $P^{\prime}$ can be induced by the following: if $E\times F$ is a Borel subset of $\Omega_{z}\times\Omega_{z}$ for some $z$, then
+$P^{\prime}(E\times F)=\begin{cases}\frac{P(E)P(F)}{P(\Omega_{z})},\quad&\text{if }P(\Omega_{z})>0,\\ 0,&\text{if }P(\Omega_{z})=0.\end{cases}$
+In the following, we assume without loss of generality that $P(\Omega_{z})>0$ for all $z\in range(h(Z))$.
+Now for any random variable $W$ defined on $(\Omega,\mathcal{B},P)$, let $W^{\prime}$ be the following random variable on $(\Omega^{\prime},\mathcal{B}^{\prime},P^{\prime})$:
+$W^{\prime}(\omega,\nu)=W(\omega),\quad(\omega,\nu)\in\Omega^{\prime}.$
+By this definition, one sees that $\big{(}f(W))^{\prime}=f(W^{\prime})$ for any admissible function $f$. In the following, we verify that $W^{\prime}\equiv W$. For any Borel set $L$,
+$P^{\prime}(W^{\prime}\in L)$ $=P^{\prime}(\{(\omega,\nu)\in\Omega^{\prime}:\,W^{\prime}(\omega,\nu)\in L)$
+$=\sum\limits_{z}P^{\prime}(\{(\omega,\nu)\in\Omega_{z}\times\Omega_{z}:\,W(\omega)\in L\})$
+$=\sum\limits_{z}\frac{P(W\in L,h(Z)=z)P(h(Z)=z)}{P(h(Z)=z)}=P(W\in L).$
+In particular, we can take $W=(X,Y,Z)$ to get $(X^{\prime},Y^{\prime},Z^{\prime})\equiv(X,Y,Z)$.
 All their corresponding marginal distributions are also the same.
-One has $h(C)\equiv h(C^{\prime})$ as well.
-Next, we define $B^{\prime\prime}$ as a function on $(\Omega^{\prime},\mathcal{B}^{\prime},P^{\prime})$, given by
-$B^{\prime\prime}(\omega,\nu)=B(\nu),\quad(\omega,\nu)\in\Omega^{\prime}.$
+One has $h(Z)\equiv h(Z^{\prime})$ as well.
+Next, we define $Y^{\prime\prime}$ as a function on $(\Omega^{\prime},\mathcal{B}^{\prime},P^{\prime})$, given by
+$Y^{\prime\prime}(\omega,\nu)=Y(\nu),\quad(\omega,\nu)\in\Omega^{\prime}.$
 Then, for any Borel sets $L_{1}$, $L_{2}$,
-$P^{\prime}(B^{\prime\prime}\in L_{1},D^{\prime}\in L_{2},h(C^{\prime})=c)=P^{\prime}(\{(\omega,\nu)\in\Omega^{\prime}:\,B^{\prime\prime}(\omega,\nu)\in L_{1},D^{\prime}(\omega,\nu)\in L_{2},h(C^{\prime}(\omega,\nu))=c\})$
-$\qquad=P^{\prime}(\{(\omega,\nu)\in\Omega_{c}\times\Omega_{c}:\,B(\nu)\in L_{1},D(\omega)\in L_{2}\})=\frac{P(D\in L_{2},h(C)=c)P(B\in L_{1},h(C)=c)}{P(h(C)=c)}.$
-When we take $D$ to be trivial (a constant function), then $P^{\prime}(B^{\prime\prime}\in L_{1},h(C^{\prime})=c)=P(B\in L_{1},h(C)=c)$ and then $(B^{\prime\prime},h(C^{\prime}))\equiv(B,h(C))$.
-Moreover, the previous argument implies that $(D^{\prime},h(C^{\prime}))\equiv(D,h(C))$.
+$P^{\prime}(Y^{\prime\prime}\in L_{1},W^{\prime}\in L_{2},h(Z^{\prime})=z)=P^{\prime}(\{(\omega,\nu)\in\Omega^{\prime}:\,Y^{\prime\prime}(\omega,\nu)\in L_{1},W^{\prime}(\omega,\nu)\in L_{2},h(Z^{\prime}(\omega,\nu))=z\})$
+$\qquad=P^{\prime}(\{(\omega,\nu)\in\Omega_{z}\times\Omega_{z}:\,Y(\nu)\in L_{1},W(\omega)\in L_{2}\})=\frac{P(W\in L_{2},h(Z)=z)P(Y\in L_{1},h(Z)=z)}{P(h(Z)=z)}.$
+When we take $W$ to be trivial (a constant function), then $P^{\prime}(Y^{\prime\prime}\in L_{1},h(Z^{\prime})=z)=P(B\in L_{1},h(Z)=z)$ and then $(Y^{\prime\prime},h(Z^{\prime}))\equiv(Y,h(Z))$.
+Moreover, the previous argument implies that $(W^{\prime},h(Z^{\prime}))\equiv(W,h(Z))$.
 We can calculate further to obtain
-$P^{\prime}(B^{\prime\prime}\in L_{1},D^{\prime}\in L_{2},h(C^{\prime})=c)$ $=\frac{P(D\in L_{2},h(C)=c)P(B\in L_{1},h(C)=c)}{P(h(C)=c)}$
-$=\frac{P^{\prime}(D^{\prime}\in L_{2},h(C^{\prime})=c)P^{\prime}(B^{\prime\prime}\in L_{1},h(C^{\prime})=c)}{P^{\prime}(h(C^{\prime})=c)}$
-$=P^{\prime}(D^{\prime}\in L_{2},h(C^{\prime})=c)P^{\prime}(B^{\prime\prime}\in L_{1}|h(C^{\prime})=c).$
-So $P^{\prime}(B^{\prime\prime}\in L_{1}|D^{\prime}\in L_{2},h(C^{\prime})=c)=P^{\prime}(B^{\prime\prime}\in L_{1}|h(C^{\prime})=c)$, i.e.
-$(B^{\prime\prime}|D^{\prime},h(C^{\prime}))\equiv(B^{\prime\prime}|h(C^{\prime}))$.
-Taking $D=(A,B)$, one gets $(B^{\prime\prime}|A^{\prime},B^{\prime},h(C^{\prime}))\equiv(B^{\prime\prime}|h(C^{\prime}))$.
-Moreover, if we take $D=B$, then $(B^{\prime\prime}|B^{\prime},h(C^{\prime}))\equiv(B^{\prime\prime}|h(C^{\prime}))$.
-That is to say, $B^{\prime\prime}$ and $B^{\prime}$ are conditionally independent given $h(C^{\prime})$.
-Since $(B^{\prime\prime},h(C^{\prime}))\equiv(B,h(C))\equiv(B^{\prime},h(C^{\prime}))$, one also deduces that $B^{\prime\prime}$, $B$ are conditionally i.i.d.
-given $h(C^{\prime})$.
-Furthermore, it follows from $(B^{\prime\prime}|A^{\prime},B^{\prime},h(C^{\prime}))\equiv(B^{\prime\prime}|h(C^{\prime}))$ that $(A^{\prime},B^{\prime})$ and $B^{\prime\prime}$ are conditionally independent given $h(C^{\prime})$.
+$P^{\prime}(Y^{\prime\prime}\in L_{1},W^{\prime}\in L_{2},h(Z^{\prime})=z)$ $=\frac{P(W\in L_{2},h(Z)=z)P(Y\in L_{1},h(Z)=z)}{P(h(Z)=z)}$
+$=\frac{P^{\prime}(W^{\prime}\in L_{2},h(Z^{\prime})=z)P^{\prime}(Y^{\prime\prime}\in L_{1},h(Z^{\prime})=z)}{P^{\prime}(h(Z^{\prime})=z)}$
+$=P^{\prime}(W^{\prime}\in L_{2},h(Z^{\prime})=z)P^{\prime}(Y^{\prime\prime}\in L_{1}|h(Z^{\prime})=z).$
+So $P^{\prime}(Y^{\prime\prime}\in L_{1}|D^{\prime}\in L_{2},h(Z^{\prime})=z)=P^{\prime}(Y^{\prime\prime}\in L_{1}|h(Z^{\prime})=z)$, i.e.
+$(Y^{\prime\prime}|W^{\prime},h(Z^{\prime}))\equiv(Y^{\prime\prime}|h(Z^{\prime}))$.
+Taking $W=(X,Y)$, one gets $(Y^{\prime\prime}|X^{\prime},Y^{\prime},h(Z^{\prime}))\equiv(Y^{\prime\prime}|h(Z^{\prime}))$.
+Moreover, if we take $W=Y$, then $(Y^{\prime\prime}|Y^{\prime},h(Z^{\prime}))\equiv(Y^{\prime\prime}|h(Z^{\prime}))$.
+That is to say, $Y^{\prime\prime}$ and $Y^{\prime}$ are conditionally independent given $h(Z^{\prime})$.
+Since $(Y^{\prime\prime},h(Z^{\prime}))\equiv(Y,h(Z))\equiv(Y^{\prime},h(Z^{\prime}))$, one also deduces that $Y^{\prime\prime}$, $B$ are conditionally i.i.d.
+given $h(Z^{\prime})$.
+Furthermore, it follows from $(Y^{\prime\prime}|X^{\prime},Y^{\prime},h(Z^{\prime}))\equiv(Y^{\prime\prime}|h(Z^{\prime}))$ that $(X^{\prime},Y^{\prime})$ and $Y^{\prime\prime}$ are conditionally independent given $h(Z^{\prime})$.
 Without confusion, we will simplify this $P^{\prime}$ as $P$ in applications for convenience. ∎
-Lemma B.2.
+Lemma A.2.
 Let $X$ be a random variable, and $E$ be an event with $P(E)>0$.
@@
 $H_{n}(f(X_{E}))\leq\frac{H_{n}(f(X))}{P(E)}.$
-Furthermore, if $X$ is $(s,C)$-Frostman, then $X^{\prime}$ is $(s,C/P(E))$-Frostman.
+Furthermore, if $X$ is $(s,C)$-Frostman, then $X_{E}$ is $(s,C/P(E))$-Frostman.
 Proof.
@@
 Thang Pham and Steven Senger would like to thank the Vietnam Institute for Advanced Study in Mathematics (VIASM) and the Erdős Center for the hospitality and for the excellent working condition.
+We would like to thank Tuomas Orponen for his comments on the first version of this paper and for telling us about the references [1, 22].

### v3

@@
-On an entropy inequality for quadratic forms and applications
+Frostman random variables, entropy inequalities, and applications
 Abstract
 We introduce Frostman conditions for bivariate random variables and study discretized entropy sum–product phenomena in both independent and dependent regimes.
-In the independent setting, we show that for any non-degenerate rational quadratic form $\phi(x,y)$ and $0<s<1$, there exists some positive constant $\epsilon=\epsilon(\phi,s)$ such that any independent marginally $(s,C;s,C)$-Frostman random variable $(X,Y)$ satisfies
-$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon)$
-for sufficiently large $n$.
+We prove that for any non-degenerate rational quadratic form $\phi(x,y)$ and parameter $0<s<1$, there exists a positive constant $\epsilon=\epsilon(\phi,s)$ such that
+$\max\{H_{n}(X+Y),H_{n}(\phi(X,Y))\}\geq n(s+\epsilon)$
+for sufficiently large $n$, where the precise conditions on $(X,Y)$ depend on the Frostman level.
 The proof introduces a novel multi‑step entropy framework, combining the submodularity formula, the discretized entropy Balog-Szemerédi-Gowers theorem, and state-of-the-art results on the Falconer distance problem, to reduce general forms to a diagonal core case.
-In the dependent setting, we obtain analogous bounds for special polynomials.
-As an application, we derive that there exists $\epsilon=\epsilon(s)>0$ such that, for any $\delta$-separated set $A\subset[0,1]$ of cardinality $\delta^{-s}$ satisfying some non-concentration condition and any subset $G\subseteq A\times A$ with $|G|\geq\delta^{\epsilon}|A|^{2}$, the $\delta$-covering numbers satisfy
-$\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2})\gg_{\alpha,\beta,s}C^{-O_{s}(1)}\delta^{-\epsilon}|A|,$
-whenever $\delta$ is small enough.
+As an application, we derive a result on a discretized sum-product type problem.
+In particular, for a $\delta$-separated set $A\subset[0,1]$ of cardinality $\delta^{-s}$, satisfying some non-concentration conditions, and a dense subset $G\subset A\times A$, there exists $\epsilon=\epsilon(s,\phi)>0$ such that
+$\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\phi_{G}(A,A))\gg\delta^{-\epsilon}(\#A)$
+for all $\delta$ small enough. Here by $\mathcal{E}_{\delta}(A)$ we mean the $\delta$-covering number of $A$, $\{A+_{G}A:=\{x+y\colon(x,y)\in G\}$, and $\phi_{G}(A):=\{\phi(x,y)\colon(x,y)\in G\}$.
 1 Introduction
@@
 More precisely, we are interested in bounding the following quantity $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ from below, where $\phi(x,y)=a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ is a non-degenerate rational quadratic form, i.e.
-$a_{2}^{2}-4a_{1}a_{3}\neq 0$, $X$ and $Y$ are independent $(s,C)$-Frostman random variables whose ranges are bounded, and $s\in(0,1)$.
-Throughout this paper, the quadratic polynomial $\phi(x,y)$ is called special if there exist univariate real-analytic functions $h,a,b$ such that $\phi(x,y)=h\bigl{(}a(x)+b(y)\bigr{)}$.
-A direct computation shows that a special quadratic polynomial is one of the following two forms: $\phi(x,y)=\alpha x^{2}+\beta y^{2}$ and $\phi(x,y)=\gamma xy$, for $\alpha,\beta,\gamma\in\mathbb{R}^{*}$.
+$a_{2}^{2}-4a_{1}a_{3}\neq 0$, and $(X,Y)$ is a bivariate Frostman random variable with bounded range.
 For $\phi(x,y)=x\cdot y$, we recall the following theorem due to Máthé and O’Regan in [12].
@@
 Moreover, we say that the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, if $X$ is $(s_{1},C_{1})$-Frostman given $Y$, and $Y$ is $(s_{2},C_{2})$-Frostman given $X$.
-These definitions are not equivalent. At a common exponent level $(s_{1},s_{2})$, we have the implications
+These definitions are not equivalent. At a common exponent level $(s_{1},s_{2})$, they form a hierarchy:
 $\boxed{\text{\parbox{116.65646pt}{\centering independent\\ and marginally Frostman\@add@centering}}}\;\Rightarrow\;\boxed{\text{conditionally Frostman}}\;\Rightarrow\;\boxed{\text{jointly Frostman}}\;\Rightarrow\;\boxed{\text{marginally Frostman}}$
@@
   • a dependent pair that is conditionally $(1/4,O(1);1/4,O(1))$-Frostman, but $(1/2,1/2;O(1))$-jointly Frostman (Example 7.6).
-Our first theorem addresses the case that $X$ and $Y$ are independent real random variables, $(X,Y)$ forms marginally $(s_{1},C_{1};S_{2},C_{2})$-Frostman, and $\phi(x,y)$ is a non-degenerate rational quadratic form.
-Theorem 1.2.
+Remarkably, different levels of this hierarchy yield entropy inequalities with different strengths and applications.
+Our results demonstrate how the Frostman hierarchy affects entropy bounds. We call a quadratic form diagonal if $\phi(x,y)=\alpha x^{2}+\beta y^{2}$, and the main results read as follows.
+Theorem 1.2 (Jointly Frostman, diagonal forms).
+Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$, $C\geq 1$ and $c>0$.
+There exist a positive constant $\epsilon=\epsilon(s)$ and an integer $N=N(\alpha,\beta,s,C,c)$ such that the following statement holds.
+For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
+$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\geq n(s+\epsilon)$
+for all $n>N$.
+Theorem 1.3 (Conditionally Frostman, general forms).
 Let $\phi(x,y)$ be a non-degenerate rational quadratic form, $0<s<1$ and $C_{1},C_{2},c>0$.
-There exist a positive constant $\epsilon=\epsilon(\phi,s)$ and an integer $N=N(\phi,s,C_{1},C_{2},c)$ such that the following statement holds: Let $X$ and $Y$ be independent real random variables whose ranges are contained in [c,c][-c,c].
+There exist positive constants $\epsilon=\epsilon(s)$, $\epsilon^{\prime}=\epsilon^{\prime}(\phi,s)$ and an integer $N=N(\phi,s,C_{1},C_{2},c)$ such that the following statement holds.
+Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$.
+Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
+Then, for all $n>N$,
+$\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq\begin{cases}n(s+\epsilon),\quad&\text{if }\phi(x,y)\text{ is diagonal},\\ n(s+\epsilon^{\prime}),&\text{if }\phi(x,y)\text{ is non-diagonal}.\end{cases}$
+Here, a constant depends on $\phi$ means that it depends on the coefficients $a_{1}$, $a_{2}$ and $a_{3}$.
+Theorem 1.4 (Marginally Frostman with independence).
+Let $\phi(x,y)$ be a non-degenerate rational quadratic form, $0<s<1$ and $C_{1},C_{2},c>0$.
+There exist positive constants $\epsilon=\epsilon(s)$, $\epsilon^{\prime}=\epsilon^{\prime}(\phi,s)$ and an integer $N=N(\phi,s,C_{1},C_{2},c)$ such that the following statement holds.
+Let $X$ and $Y$ be independent real random variables whose ranges are contained in [c,c][-c,c].
 Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};S_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
-Then
-$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon)$ (2)
-for all $n>N$.
-Here, a constant depends on $\phi$ means that it depends on the coefficients $a_{1}$, $a_{2}$ and $a_{3}$.
-To show the sharpness of this theorem, we summarize the main statements of several examples in Section 7:
+Then, for all $n>N$,
+$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq\begin{cases}n(s+\epsilon),\quad&\text{if }\phi(x,y)\text{ is diagonal},\\ n(s+\epsilon^{\prime}),&\text{if }\phi(x,y)\text{ is non-diagonal}.\end{cases}$ (2)
+To show the sharpness of these results, we summarize the main statements of several examples in Section 7:
   1. 1.
     Any $(s,C)$-Frostman random variable $Z$ has the property that $H_{n}(Z)\geq ns-\log C+O(1)$ (see Example 7.1 for an example matching this bound).
-The lower bound in (2) says that the discretized entropies of $X+Y$ and $\phi(X,Y)$ cannot be small simultaneously.
-We show in Examples 7.1 and 7.2 that at least one of the terms $H_{n}(X+Y)$ or $H_{n}(\phi(X,Y))$ can be as small as $(1+o(1))ns$.
+We show in Examples 7.1 and 7.2 that there are i.i.d.
+random variables $X$ and $Y$ such that at least one of the terms $H_{n}(X+Y)$ or $H_{n}(\phi(X,Y))$ can be as small as $(1+o(1))ns$.
   2. 2.
@@
   3. 3.
-    In the statement of Theorem 1.2, the fact that NN depends on $C$ , and the condition that $X$ and $Y$ are independent are both essential (see Examples 7.4 and 7.5, respectively).
-  4. 4.
     For any $0<s<1$ and any sufficiently large $n\in\mathbb{N}$, there exists an $(s,O(1))$-Frostman random variable $X$ with range contained in $[0,1]$, such that
@@
     as $n\rightarrow\infty$, where $Y$ is i.i.d. with $X$. In other words, the term $ns$ can not be replaced by $H_{n}(X)$ in the right-hand side of (2).
+  4. 4.
+    In Theorems 1.2 and 1.3, the necessity of different terms in these maxima (for example, $\frac{H_{n}(X,Y)}{2}$ or $\frac{H_{n}(X)+H_{n}(Y)}{2}$) reflects subtle aspects of how dependency affects entropy, we provide an example showing it cannot be simplified in the Statement of Theorem 1.2 (Example 7.5).
   5. 5.
+    In the statement of Theorem 1.4, the fact that NN depends on $C$, and the condition that $X$ and $Y$ are independent are both essential (see Examples 7.4 and 7.5, respectively).
+  6. 6.
     The value of $\epsilon$ is not explicit, and an inspection of the proofs suggests it would be small.
+  7. 7.
+    Theorem 1.2 can be extended to a more general form of
+    $\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(f(X)+g(Y))\right\},$
+    where $f$ and $g$ are polynomials with $\deg(f)=\deg(g)\geq 2$.
+Establishing such a result would require extending most of the arguments in this paper, and we plan to provide further updates in a subsequent version.
 We pause here to add several further remarks.
-Our proof of Theorem 1.2 presents a connection between $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ and the entropy of the “distance” function, i.e.
+Our proof of Theorems 1.21.4 presents a connection between $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ and the entropy of the “distance” function, i.e.
 $H_{n}\big{(}(X_{1}-Y_{1})^{2}+(X_{2}-Y_{2})^{2}\big{)}$, where $(X_{1},Y_{1})$, $(X_{2},Y_{2})$ are i.i.d.
 copies of $(X,Y)$.
-When $a_{1}=a_{3}=0$ and $a_{2}=1$, i.e., $\phi(x,y)=xy$, Máthé and O’Regan [12] established a different type of connection by replacing the entropy of the “distance” function with that of the “slope” function.
+When $a_{1}=a_{3}=0$ and $a_{2}=1$, i.e.
+$\phi(x,y)=xy$, Máthé and O’Regan [12] established a different type of connection by replacing the entropy of the “distance” function with that of the “slope” function.
 By combining this with a generalization of Marstrand’s theorem due to Orponen, Shmerkin, and Wang in [21], they were able to prove Theorem 1.4 for the full range of $s$ for this specific bilinear form function.
@@
 It would be interesting to explore more connections between $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ and the entropy of other functions.
-It is worth noting that the proof of Theorem 1.2 can be extended to more general settings.
-In this version, we present two results for dependent random variables and special polynomials.
-Further results will be included in a forthcoming update.
-Theorem 1.3.
-Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$ and $C_{1},C_{2},c>0$.
-There exist a positive constant $\epsilon=\epsilon(s)$ and an integer $N=N(\alpha,\beta,s,C_{1},C_{2},c)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$.
-Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
-Then
-$\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\geq n(s+\epsilon),$
-and
-$\max\left\{\frac{H_{n}(X+Y)+H_{n}(X-Y)}{2},\,H_{n}(XY)\right\}\geq n(s+\epsilon)$
-for all $n>N$.
-Both Theorems 1.2 and 1.3 are consequences of the following theorem, which only requires a jointly Frostman condition.
-Theorem 1.4.
-Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$, $C\geq 1$ and $c>0$.
-There exist a positive constant $\epsilon=\epsilon(s)$ and an integer $N=N(\alpha,\beta,s,C,c)$ such that the following statement holds: For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
-$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\geq n(s+\epsilon)$
-for all $n>N$.
-Example 7.6 will illustrate that the first term $\frac{H_{n}(X)+H_{n}(Y)}{2}$ in above formula cannot be removed, and cannot be replaced by $\frac{1}{2}H(X,Y)$ as in Theorem 1.3.
 In the following, we discuss some applications of the entropy theorems.
-Covering number: non-degenerate quadratic polynomials.
-The next theorem presents an application of Theorem 1.2. Here, by $\underline{\dim}_{B}$ and $H^{s}$, we mean the lower box dimension and the $s$-Hausdorff measure, respectively.
-Theorem 1.5.
-Let $\phi(x,y)$ be a non-degenerate rational quadratic form.
-For $0<s<1$, there exists a positive constant $\epsilon=\epsilon(s{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},\phi})$ such that for every Borel set $A\subset{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[0,1]}$ with $H^{s}(A)>0$, we have
-$\max\left\{\underline{\dim}_{B}(A+A),\overline{\dim}_{B}(\phi(A,A))\right\}\geq s+\epsilon,$
-and
-$\max\left\{\overline{\dim}_{B}(A+A),\underline{\dim}_{B}(\phi(A,A))\right\}\geq s+\epsilon.$
-For $A\subset\mathbb{R}$ and $\delta>0$, by $\mathcal{E}_{\delta}(A)$ we mean the $\delta$-covering number of $A$. The above theorem can be rewritten in the language of covering numbers.
-Theorem 1.6.
-Let $\phi(x,y)$ be a non-degenerate rational quadratic form.
-Given $0<s<1$ and $C\geq 1$.
-Then there exist positive constants $\epsilon=\epsilon(\phi,s)$ such that the following holds for all $\delta>0$.
-Let $A\subset[0,1]$ be any $\delta$-separated set of cardinality $\lceil\delta^{-s}\rceil$.
-Suppose that for all intervals $I$ of length $l$ at least $\delta$, we have the non-concentration estimate
-$(A\cap I)\leq Cl^{s}(\#A).$
-Then we have
-$\mathcal{E}_{\delta}(A+A)+\mathcal{E}_{\delta}(\phi(A,A))\gg_{\phi,s}C^{-O_{\phi,s}(1)}\delta^{-\epsilon}(\#A).$
-These two applications are of independent interest, particularly when $\phi(x,y)$ is a non-bilinear form, and might be considered as an application of the Falconer distance problem.
-If $\phi(x,y)=xy$, then stronger results have been obtained by using a variety of techniques, see [5, 11, 12, 17] and references therein for current progress.
-Given this line of research, it is reasonable to expect different proofs of Theorem 1.5 with other methods.
-In the prime field setting, such a result has been proved by using tools from incidence geometry, see [9, Theorem 1.5] and [15, Theorem 1.3], for example.
-When $\phi(x,y)=x^{2}+y^{2}$, after posting this paper to Arxiv, Tuomas Orponen pointed out that one can use his proof of [22, Corollary 1.3] and recent results in [1] or [23] to obtain a quantitative bound for this special case.
-Fractional covering number: special quadratic polynomials.
+For $A\subset\mathbb{R}$ and $\delta>0$, by $\mathcal{E}_{\delta}(A)$ we mean the $\delta$-covering number of $A$.
 For a set $G\subseteq A\times A$, we denote
-$A+_{G}A=\{x+y:\,(x,y)\in G\},\qquad\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2}=\{\alpha x^{2}+\beta y^{2}:\,(x,y)\in G\}.$
+$A+_{G}A=\{x+y:\,(x,y)\in G\},\qquad\phi_{G}(A,A)=\{\phi(x,y):\,(x,y)\in G\},$
+$\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2}=\{\alpha x^{2}+\beta y^{2}:\,(x,y)\in G\}.$
 The next theorem presents an application of Theorem 1.2 on fractional covering numbers.
@@
 $(A\cap I)\leq Cl^{s}(\#A).$
-Then, for all $G\subset A\times A$ with $|G|\geq\delta^{\epsilon}|A|^{2}$, then
+Then, for all $G\subseteq A\times A$ with $G\geq\delta^{\epsilon}(\#A)^{2}$, we have
 $\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2})\gg_{\alpha,\beta,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A).$
-Sketch the main ideas in the proof of Theorem 1.2.
+Thanks to Theorem 1.3, extend this result to all non-degenerate rational quadratic forms is possible, but requiring minimum degree constraints.
+Theorem 1.6.
+Let $\phi(x,y)$ be a non-degenerate rational quadratic form, $0<s<1$ and $C\geq 1$.
+Then there exist positive constants $\epsilon=\epsilon(\phi,s)$ and $\delta_{0}=\delta_{0}(\alpha,\beta,s,C)$ such that the following holds for all $0<\delta\leq\delta_{0}$.
+Let $A\subset{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[0,1]}$ be any $\delta$-separated set of cardinality $\lceil\delta^{-s}\rceil$.
+Suppose that for all intervals $I$ of length $l$ at least $\delta$, we have the non-concentration estimate
+$(A\cap I)\leq Cl^{s}(\#A).$
+Then, for any subset $G\subseteq A\times A$ such that
+$\{y:\,(x,y)\in G\}\geq\delta^{\epsilon}(\#A)\quad\text{for any}\quad x\in A,$
+$\{x:\,(x,y)\in G\}\geq\delta^{\epsilon}(\#A)\quad\text{for any}\quad y\in A,$
+we have
+$\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\phi_{G}(A,A))\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A).$
+In contrast to Theorem 1.6, the next result, an application of Theorem 1.4, holds for every scale $\delta>0$ and is stable under taking subsets: it applies to any $B\subset A$ with $B\geq\delta^{\varepsilon}\#A$.
+Theorem 1.7.
+Let $\phi(x,y)$ be a non-degenerate rational quadratic form, $0<s<1$ and $C\geq 1$.
+Then there exist positive constants $\epsilon=\epsilon(\phi,s)$ such that the following holds for all $\delta>0$.
+Let $A\subset{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[0,1]}$ be any $\delta$-separated set of cardinality $\lceil\delta^{-s}\rceil$.
+Suppose that for all intervals $I$ of length $l$ at least $\delta$, we have the non-concentration estimate
+$(A\cap I)\leq Cl^{s}(\#A).$
+Then, for any subset $B\subseteq A$ satisfying $B\geq\delta^{\epsilon}(\#A)$, we have
+$\mathcal{E}_{\delta}(B+B)+\mathcal{E}_{\delta}(\phi(B,B))\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A).$
+These applications are of independent interest, particularly when $\phi(x,y)$ is a non-bilinear form, and might be viewed as consequences of the Falconer distance problem.
+If $\phi(x,y)=xy$, then stronger results have been obtained by using a variety of techniques, see [5, 11, 12, 17] and references therein for current progress.
+Given this line of research, it is reasonable to expect different proofs of Theorems 1.51.7 with other methods.
+In the prime field setting, such a result has been proved by using tools from incidence geometry, see [9, Theorem 1.5] and [15, Theorem 1.3], for example.
+When $\phi(x,y)=x^{2}+y^{2}$ ($\alpha=\beta=1$) and $G=A\times A$, after posting this paper to Arxiv, Tuomas Orponen pointed out that one can use his proof of [22, Corollary 1.3] and recent results in [1] or [23] to obtain a quantitative lower bound for the following estimate
+$\mathcal{E}_{\delta}(A+A)+\mathcal{E}_{\delta}(A^{\wedge 2}+A^{\wedge 2})\gg_{\alpha,\beta,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A).$
+We now sketch the main ideas of proofs in this paper.
 Our base case is corresponding to the diagonal forms $\phi(X,Y)=\alpha X^{2}+\beta Y^{2}$ (Lemma 4.1), with independent $(s,C)$-Frostman random variables $X,Y$.
@@
 We denote the range of $X$ by $range(X)$, which collects all the possible values that $X$ can take.
-In this paper, all the measurable spaces will be taken to be subspaces of ${\mathbb{R}}^{m}$, equipped with the Borel $\sigma$-algebra or its sub-$\sigma$-algebra, and the Euclidean norm $|\cdot|$.
+In this paper, all the measurable spaces will be taken to be subspaces of ${\mathbb{R}}^{m}$, equipped with the Borel $\sigma$-algebra or its sub-$\sigma$-algebra, the Euclidean norm $|\cdot|$, and the Lebesgue measure $\mathcal{L}^{m}(\cdot)$.
 The dimension $m$ is always regarded as an absolute constant.
@@
 $H_{n}(\alpha X^{2})=H_{n}(X)+O\big{(}\max\{1,\log c_{1},\log(1/c_{2})\}\big{)}+O(\big{|}\log|\alpha|\big{|}).$
+Example 2.17.
+Let $m\geq 1$.
+Suppose that $A$ is an $m\times m$ real matrix and $X,X_{1},X_{2}$ are multivariate real random variables of dimension $m$.
+Note that $|A(X_{1}-X_{2})|\leq\|A\|\cdot|X_{1}-X_{2}|$, where $\|A\|$ is the operator norm of $A$.
+Lemma 2.15 shows that $H_{n}(AX)\leq H_{n}(X)+O_{A}(1)$.
+In particular, if $A$ is invertible, then $H_{n}(AX)=H_{n}(X)+O_{A}(1)$.
 The following is the discretized version of submodularity.
@@
 Let $k,l\geq 0$ and $X_{1},\ldots,X_{k}$, $Y_{1},\ldots,Y_{l}$ be real random variables. Then for any $n\in\mathbb{N}$,
-$H_{n}(X_{1}+\ldots+X_{k},\,Y_{1}+\ldots,Y_{l})$
-$\quad=H(D_{n}(X_{l})+\ldots+D_{n}(X_{k}),\,D_{n}(Y_{1})+\ldots+D_{n}(Y_{l}))+O_{k,l}(1).$
+$H_{n}(X_{1}+\ldots+X_{k},\,Y_{1}+\ldots+Y_{l})$
+$\quad=H(D_{n}(X_{1})+\ldots+D_{n}(X_{k}),\,D_{n}(Y_{1})+\ldots+D_{n}(Y_{l}))+O_{k,l}(1).$
 Proof.
@@
 for all $n\in\mathbb{N}$, provided that $X$ and $Y$ are independent. In the rest of this paper, we may apply Lemma 2.20 without mention it.
-In [16], Ruzsa studied some entropy inequalities involving sums of random variables. In the following, more Plünnecke-Ruzsa type inequalities for discretized entropy are presented.
+In [16], Ruzsa studied some entropy inequalities that involving sums of random variables. In the following, more inequalities for discretized entropy are presented.
 Lemma 2.21.
@@
 By Lemma 2.18,
-$H_{n}(X+Y+Z)+H_{n}(X,Y,Z)\leq H_{n}(X+Z,Y)+H_{n}(X,Y+Z)+O(1).$
+$H_{n}(X+Y+Z)+H_{n}(X,Y,Z)\leq H_{n}(X+Z,Y)+H_{n}(X,Y+Z)+O(1).$ (14)
 Combining Lemma 2.4(iv) and the independence, the first inequality follows. Next, it follows from (13) that
@@
 Another application of Lemma 2.18 gives
-$H_{n}(X-Y)+H_{n}(X,Y,Z)\leq H_{n}(X+Z,Y+Z)+H_{n}(X,Y)+O(1).$
+$H_{n}(X-Y)+H_{n}(X,Y,Z)\leq H_{n}(X+Z,Y+Z)+H_{n}(X,Y)+O(1).$ (15)
 Combining independence, the second inequality involving XYX-Y then follows. ∎
+Without independence, the follow lemma can be used as an alternative.
 Lemma 2.22.
-Let $X$ and $Y$ be independent real random variables and $k,l\in\mathbb{Q}$. There exist a non-negative integers $d=d(k,l)$ such that
-$H_{n}(kX+lY)\leq dH_{n}(X+Y)-(d-1)(H_{n}(X)-H_{n}(Y))+O_{k,l}(1)$
+Let $X,Y$ and $Z$ be real random variables that are not necessarily independent. Then, for all $n\in\mathbb{N}$, we have
+$H_{n}(X+Y)\leq H_{n}(X+Y+Z)+H_{n}(X,Y)+H_{n}(Z)-H_{n}(X,Y,Z)+O(1),$
+and
+$H_{n}(X+Y)$ $\leq H_{n}(X+Z)+H_{n}(Y+Z)-2H_{n}(X,Y,Z)$
+$\quad+H_{n}(X,Y)+H_{n}(X)+H_{n}(Y)+H_{n}(Z)+O(1).$
+Proof.
+Lemma 2.18 shows that
+$H_{n}(X+Y)+H_{n}(X,Y,Z)\leq H_{n}(X+Y+Z,Z)+H_{n}(X,Y)+O(1).$
+Combining (14), The lemma then follows. ∎
+An important task of this subsection is to prove the following theorem on the relationship between $H_{n}(kX+lY)$ and $H_{n}(X+Y)$.
+Theorem 2.23.
+Let $k,l\in\mathbb{Q}^{\ast}$, and $X$ and $Y$ be real random variables that are not necessarily independent. Then there exist integers $d_{1}=d_{1}(k,l)>0$ and $d_{2}=d_{2}(k,l)\geq 0$ such that
+$H_{n}(kX+lY)-\frac{H_{n}(X)+H_{n}(Y)}{2}\leq d_{1}\left(H_{n}(X+Y)-\frac{H_{n}(X)+H_{n}(Y)}{2}\right)$
+$\qquad\qquad\qquad+d_{2}\big{(}H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\big{)}+O_{k,l}(1)$
 for all $n\in\mathbb{N}$. Here by $kX$ we mean the product of $k$ and the random variable $X$.
+The proof is divided into several lemmas as below.
+Lemma 2.24.
+Let $X$ and $Y$ be real random variables that are not necessarily independent. Then, for all $n\in\mathbb{N}$,
+$H_{n}(X-Y)$ $\leq 3H_{n}(X+Y)+4H_{n}(X)+4H_{n}(Y)-5H_{n}(X,Y)+O(1).$
 Proof.
-We first prove that
-$H_{n}(kX+lY)\leq(d_{1}+d_{2}+1)H_{n}(X+Y)$
-$\qquad-d_{1}H_{n}(X)-d_{2}H_{n}(Y)+(d_{1}+d_{2}+1)I_{n}(X,Y)+O_{k,l}(1)$ (14)
-for some non-negative integers $d_{1}=d_{1}(k,l)$ and $d_{2}=d_{2}(k,l)$.
-The proof is trivial when $k=l=0$. When $l=0$ and $k\neq 0$, we have
-$H_{n}(kX)=H_{n}(X)+O_{k}(1)\leq H_{n}(X+Y)+O_{k}(1)$
-by (13). Similar conclusion holds when $k=0$. In the following, we assume that $k,l\neq 0$.
-First, let us prove the case that $k,l\in\mathbb{Z}^{\ast}$. We apply double induction on $|k|$ and $|l|$. When $|k|=|l|=1$, we claim that
-$H_{n}(kX+lY)\leq 3H_{n}(X+Y)-H_{n}(X)-H_{n}(Y)+O(1).$ (15)
-Indeed, when $k=l=1$, one can deduce (15) from (13).
-When $k=1$ and l=1l=-1, we take $X,Y,Z$ to be independent random variables such that $X$ and $Z$ are identically distributed.
-Then Corollary 2.20 gives
-$H_{n}(X-Y)\leq H_{n}(X+Z)+H_{n}(Y+Z)-H_{n}(Z)+O(1).$
-By Corollary 2.20 again, we have
-$H_{n}(X+Z)\leq H_{n}(X+Y)+H_{n}(Z+Y)-H_{n}(Y)+O(1).$ (16)
-Then (15) follows.
-Since $H_{n}(W)=H_{n}(-W)$ for any random variable $W$, the cases that k=1k=-1 and $l=\pm 1$ are also satisfied.
-Since $H_{n}(W)=H_{n}(-W)$ for any random variable $W$, the cases that k=1k=-1 and $l=\pm 1$ are also satisfied.
-Now let us assume that the conclusion has been proved for some $k$ and $l$. Without loss of generality, we assume that $k>0$. We aim to prove the conclusion for $k+1$ and $l$. Lemma 2.17 gives us that
-$H_{n}((k+1)X+lY+Z)+H_{n}(X,Y,Z)\leq H_{n}(Z+lY,X)+H_{n}(X+Z,kX+lY)+O_{k,l}(1),$
-where the random variable $Z$ is taken such that $X,Y,Z$ are independent and $X$, $Z$ are identically distributed. Combining (13), we obtain that
-$H_{n}((k+1)X+lY)\leq H_{n}((k+1)X+lY+Z)+O_{k,l}(1)$
-$\leq H_{n}(Z+lY,X)+H_{n}(X+Z,kX+lY)-H_{n}(X,Y,Z)+O_{k,l}(1)$
-$\leq H_{n}(Z+lY)+H_{n}(X+Z)+H_{n}(kX+lY)-H_{n}(Y)-H_{n}(Z)+O_{k,l}(1).$
-Combining the fact
-$H_{n}(X+Z)\leq H_{n}(X+Y)+H_{n}(Z+Y)-H_{n}(Y)+O(1)$
-and the inductive hypothesis on $k$, one sees that the conclusion holds with some $d_{1}(k,l)$ and $d_{2}(k,l)$ satisfying
-$d_{1}(k+1,l)=d_{1}(1,l)+d_{1}(k,l)+1,\quad d_{2}(k,l)=d_{2}(1,l)+d_{2}(k,l)+2.$
-By symmetry, we also have the induction on $l$, and the proof of (14) is completed for the situation that $k,l\in\mathbb{Z}^{\ast}$.
-Second, when $k,l\in\mathbb{Q}^{\ast}$, we can write $k=k^{\prime}/t$ and $l=l^{\prime}/t$ for some $k^{\prime},l^{\prime},t\in\mathbb{Z}^{\ast}$. It follows that
-$H_{n}(kX+lY)$ $=H_{n}(k^{\prime}X+l^{\prime}Y)+O_{t}(1)$
-$\leq(d_{1}+d_{2}+1)H_{n}(X+Y)-d_{1}H_{n}(X)-d_{2}H_{n}(Y)+O_{k^{\prime},l^{\prime},t}(1)$
-for some $d_{1}=d_{1}(k^{\prime},l^{\prime})$ and $d_{2}=d_{2}(k^{\prime},l^{\prime})$. Then (14) follows.
-Next, in view of (15), we can add multiple scalars of $H_{n}(X+Y)-H_{n}(X)$ and $H_{n}(X+Y)-H_{n}(Y)$ on the right-hand side of (14) to get
-$H_{n}(kX+lY)\leq(2d_{1}+2d_{2}+1)H_{n}(X+Y)-(2d_{1}+2d_{2})\frac{H_{n}(X)+H_{n}(Y)}{2}+O_{k,l}(1).$
-The lemma then follows by taking $d=2d_{1}+2d_{2}+1$. ∎
-Lemma 2.21 implies the following statement: For given $k,l\in\mathbb{Q}$, one has
-$H_{n}(kX+lY)-H_{n}(X)\ll_{k,l}(H_{n}(X+Y)-H_{n}(X))+O_{k,l}(1)$
-for any $n\in\mathbb{N}$ and any i.i.d.
-random variables $X$, $Y$ with bounded range.
-However, when $k,l$ are $\mathbb{Q}$-linearly independent, such statement may fail.
+For a given $n\in\mathbb{N}$, let $X^{\prime}$ be a random variable that is conditionally i.i.d.
+with $X$ given $D_{n}(Y)$.
+Then $D_{n}(X)$ and $D_{n}(X^{\prime})$ are also conditionally i.i.d.
+given $D_{n}(Y)$, and $(D_{n}(X),D_{n}(Y))\equiv(D_{n}(X^{\prime}),D_{n}(Y))$.
+By Lemma 2.8,
+$H_{n}(X,X^{\prime},Y)=2H_{n}(X,Y)-H_{n}(Y).$ (16)
+Recall Lemma 2.20. By (15), one has
+$H_{n}(X-Y)+H_{n}(X,X^{\prime},Y)\leq H_{n}(X+X^{\prime})+H_{n}(Y+X^{\prime})+H_{n}(X,Y)+O(1).$ (17)
+In view of the second inequality in Lemma 2.22, we get
+$\begin{split}H_{n}(X+X^{\prime})&\leq H_{n}(X+Y)+H_{n}(X^{\prime}+Y)-2H_{n}(X,X^{\prime},Y)\\ &\quad+H_{n}(X,X^{\prime})+H_{n}(X)+H_{n}(X^{\prime})+H_{n}(Y)+O(1).\end{split}$ (18)
+Combining (16)-(18), the lemma then follows. ∎
+Remark 2.25.
+When $X$ and $Y$ are independent, one has
+$H_{n}(X-Y)$ $\leq 3H_{n}(X+Y)-H_{n}(X)-H_{n}(Y)+O(1).$
+for all $n\in\mathbb{N}$.
+For simplicity, we will denote the discretized mutual information by
+$\mathcal{I}_{n}(X,Y)=H_{n}(X)+H_{n}(Y)-H_{n}(X,Y).$
+Lemma 2.26.
+Let $k\in\mathbb{Z}^{\ast}$, and $X$, $Y$ be real random variables that are not necessarily independent. Then there are integers $d_{1}=d_{1}(k)>0$ and $d_{2}=d_{2}(k)\geq 0$ such that
+$H_{n}(kX+Y)-\frac{H_{n}(X)+H_{n}(Y)}{2}$ $\leq d_{1}\left(H_{n}(X+Y)-\frac{H_{n}(X)+H_{n}(Y)}{2}\right)+d_{2}\mathcal{I}_{n}(X,Y)+O_{k}(1).$
+for all $n\in\mathbb{N}$.
+Proof.
+For $k=1$, the conclusion holds for $d_{1}=1$ and $d_{2}=0$.
+For k=1k=-1, Lemma 2.24 shows that we can take $d_{1}=3$ and $d_{2}=5$.
+Now we apply induction on $|k|$.
+Assume that the conclusion has been proved for $k$ with $|k|=m$, where $m\geq 1$.
+We consider the situation $k=\pm(m+1)$.
+Let $X^{\prime}$ be as in the proof of Lemma 2.24. By the first inequality in Lemma 2.22,
+$\begin{split}&H_{n}\big{(}\pm(m+1)X+Y\big{)}\leq H_{n}\big{(}\pm(m+1)X+Y\pm mX^{\prime}\big{)}\\ &\qquad\qquad\qquad+H_{n}(X,Y)+H_{n}(X^{\prime})-H_{n}(X,X^{\prime},Y)+O_{m}(1).\end{split}$ (19)
+Note that
+$\pm(m+1)X+Y\pm mX^{\prime}=\pm(m+1)(X\pm Y)\pm m(X^{\prime}\mp Y).$
+Applying Lemma 2.18 yields
+$H_{n}\big{(}\pm(m+1)X+Y\pm mX^{\prime}\big{)}+H_{n}(X,X^{\prime},Y)$
+$\qquad\leq H_{n}(\pm mX^{\prime}+Y,X)+H_{n}(X\pm Y,X^{\prime}\mp Y)+O(1)$
+$\qquad\leq H_{n}(\pm mX+Y)+H_{n}(X+Y)+H_{n}(X-Y)+H_{n}(X)+O(1).$ (20)
+Combining Lemma 2.24, (16), (19) and (20), it follows that
+$H_{n}\big{(}\pm(m+1)X+Y\big{)}\leq H_{n}(\pm mX+Y)+4H_{n}(X+Y)+6H_{n}(X)+6H_{n}(Y)-8H_{n}(X,Y).$
+Therefore, the conclusion follows with
+$d_{1}(\pm(m+1))=d_{1}(\pm m)+4,\qquad d_{1}(\pm(m+1))=d_{1}(\pm m)+8.$
+Lemma 2.27.
+Let $k,l\in\mathbb{Z}^{\ast}$, and $X$, $Y$ be real random variables that are not necessarily independent. Then there are integers $d_{1}=d_{1}(k,l)>0$ and $d_{2}=d_{2}(k,l)\geq 0$ such that
+$H_{n}(kX+lY)-\frac{H_{n}(X)+H_{n}(Y)}{2}$ $\leq d_{1}\left(H_{n}(X+Y)-\frac{H_{n}(X)+H_{n}(Y)}{2}\right)+d_{2}\mathcal{I}_{n}(X,Y)+O(1).$
+for all $n\in\mathbb{N}$.
+Proof.
+Applying Lemma 2.26 with the random variable $(X,lY)$, we obtain that
+$H_{n}(kX+lY)-\frac{H_{n}(X)+H_{n}(Y)}{2}$
+$\qquad\leq d_{1}(k,1)\left(H_{n}(X+lY)-\frac{H_{n}(X)+H_{n}(Y)}{2}\right)+d_{2}(k,1)\mathcal{I}_{n}(X,Y)+O_{k,l}(1).$
+Applying Lemma 2.26 again, with the random variable $(Y,X)$, we find that
+$H_{n}(lY+X)-\frac{H_{n}(X)+H_{n}(Y)}{2}$
+$\qquad\leq d_{1}(l,1)\left(H_{n}(X+Y)-\frac{H_{n}(X)+H_{n}(Y)}{2}\right)+d_{2}(l,1)\mathcal{I}_{n}(X,Y)+O_{l}(1).$
+Hence, the conclusion follows by taking
+$d_{1}(k,l)=d_{1}(k,1)d_{1}(l,1),\qquad d_{2}(k,l)=d_{1}(k,1)d_{2}(l,1)+d_{2}(k,1).$
+Proof of Theorem 2.23.
+Since $k,l\in\mathbb{Q}^{\ast}$, we can write $k=k^{\prime}/t$ and $l=l^{\prime}/t$ for some $k^{\prime},l^{\prime},t\in\mathbb{Z}^{\ast}$.
+It satisfies that $H_{n}(kX+lY)=H_{n}(k^{\prime}X+l^{\prime}Y)+O_{t}(1)$.
+The conclusion follows from Lemma 2.27.
+∎
+Remark 2.28.
+Theorem 2.23 implies the following statement: For all given $k,l\in\mathbb{Q}$, if $X$ and $Y$ are independent random variables, then
+$H_{n}(kX+lY)-\frac{H_{n}(X)+H_{n}(Y)}{2}\ll_{k,l}H_{n}(X+Y)-\frac{H_{n}(X)+H_{n}(Y)}{2}+O_{k,l}(1)$
+for any $n\in\mathbb{N}$. Here, the case that $k=0$ or $l=0$ can be proved by using (13).
+However, when $k,l$ are $\mathbb{Q}$-linearly independent, the above formula may fail.
 In Example 7.7, we would show the following: For any $K>0$, there exists some $N\in\mathbb{N}$ such that: for any $n>N$, there is a random variable $X$ with range contained in $[0,1]$ satisfying
@@
 where $Y$ is i.i.d. with $X$.
+Lemma 2.29.
+Let $X,Y$ be independent random variables, and $f,g$ be weakly $\kappa$-bi-Lipschitz real-valued functions. Suppose that $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ are i.i.d. copies of $(X,Y)$. Then
+$H_{n}(f(X_{1})+f(X_{2})+g(Y_{1})+g(Y_{2}))$ $\leq 3H_{n}(f(X)+g(Y))-H_{n}(X)-H_{n}(Y)+O(\log\kappa),$
+$H_{n}(f(X_{1})-f(X_{2})+g(Y_{1})-g(Y_{2}))$ $\leq 7H_{n}(f(X)+g(Y))-3H_{n}(X)-3H_{n}(Y)+O(\log\kappa),$
+$H_{n}(f(X_{1})+f(X_{2})-g(Y_{1})-g(Y_{2}))$ $\leq 9H_{n}(f(X)+g(Y))-4H_{n}(X)-4H_{n}(Y)+O(\log\kappa).$
+Proof.
+Note that $f(X_{1}),f(X_{2}),g(Y_{1}),g(Y_{2})$ are independent, and $(X_{1},Y_{1})\equiv(X_{2},Y_{2})\equiv(X,Y)$. By using Lemma 2.21 repeatedly, combining Lemma 2.15, we get
+$H_{n}(f(X_{1})\pm f(X_{2})+g(Y_{1})\pm g(Y_{2}))$
+$\leq H_{n}(f(X_{1})\pm f(X_{2})+g(Y_{1}))+H_{n}(f(X_{1})+g(Y_{1})\pm g(Y_{2}))-H_{n}(f(X_{1})+g(Y_{1}))+O(1)$
+$\leq H_{n}(f(X_{1})+g(Y_{1}))+H_{n}(\pm f(X_{2})+g(Y_{1}))-H_{n}(g(Y_{1}))$
+$\qquad+H_{n}(f(X_{1})+g(Y_{1}))+H_{n}(f(X_{1})\pm g(Y_{2}))-H_{n}(f(X_{1}))$
+$\qquad-H_{n}(f(X_{1})+g(Y_{1}))+O(1)$
+$=H_{n}(f(X)+g(Y))+2H_{n}(f(X)\pm g(Y))-H_{n}(X)-H_{n}(Y)+O(\log\kappa).$
+By Remark 2.25, it satisfies that
+$H_{n}(f(X)-g(Y))$ $\leq 3H_{n}(f(X)+g(Y))-H_{n}(f(X))-H_{n}(q(Y))+O(1)$
+$=3H_{n}(f(X)+g(Y))-H_{n}(X)-H_{n}(Y)+O(\log\kappa).$
+The above formulae still hold if $g$ is replaced by g-g. The lemma then follows from calculation. ∎
 2.4 Frostman measures
@@
 given $D_{n}(Y^{\prime})$.
-By Lemmas 2.8 and 2.18,
+By Lemmas 2.8 and 2.19,
 $H_{n}(X^{\prime},X^{\prime\prime},Z,Y^{\prime})=2H_{n}(X^{\prime},X^{\prime\prime},Y^{\prime})-H_{n}(X^{\prime},X^{\prime\prime})$
@@
 […]
-Using Lemmas 2.15 and 2.17, we have that
+Using Lemmas 2.15 and 2.18, we have that
 $H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime}),Y^{\prime})+H_{n}(X^{\prime},Y^{\prime},X^{\prime\prime},Y^{\prime\prime})$
@@
 Thus, the estimate (28) follows. ∎
-4 Proof of Theorems 1.2-1.4 with $s>1/2$
+4 Proof of Theorems 1.21.4 with $s>1/2$
 4.1 Diagonal forms with independent random variables
@@
 Then
-$\begin{split}&\max\{H_{n}(X+Y),~H_{n}(\alpha X^{2}+\beta Y^{2})\}\geq n(s+\epsilon_{1})\\ &\qquad\qquad\qquad+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi,s}(1).\end{split}$ (35)
+$\begin{split}&\max\{H_{n}(X+Y),~H_{n}(\alpha X^{2}+\beta Y^{2})\}\geq n(s+\epsilon_{1})\\ &\qquad\qquad\qquad+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi,s}(1).\end{split}$ (39)
 for all $n\in\mathbb{N}$.
@@
 $H_{n}(X_{1}+X_{2})+H_{n}(Y_{1}+Y_{2})\leq 4H_{n}(X+Y)-H_{n}(X)-H_{n}(Y)+O(1).$ (42)
-For the third term, we use Lemma 2.20 repeatedly, and combine Lemma 2.15 and formula (36), to get
-$H_{n}(p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}))$
-$\leq H_{n}(p(X_{1})+p(X_{2})+q(Y_{1}))+H_{n}(p(X_{1})+q(Y_{1})+q(Y_{2}))-H_{n}(p(X_{1})+q(Y_{1}))+O(1)$
-$\leq H_{n}(p(X_{1})+q(Y_{1}))+H_{n}(p(X_{2})+q(Y_{1}))-H_{n}(q(Y_{1}))$
-$\qquad+H_{n}(p(X_{1})+q(Y_{2}))+H_{n}(p(X_{1})+q(Y_{1}))-H_{n}(p(X_{1}))$
-$\qquad-H_{n}(p(X_{1})+q(Y_{1}))+O(1)$
-$=3H_{n}(p(X)+q(Y))-H_{n}(X)-H_{n}(Y)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1).$ (39)
+For the third term, Lemma 2.29 shows that
+$H_{n}(p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}))\leq 3H_{n}(p(X)+q(Y))$
+$\qquad\qquad\qquad-H_{n}(X)-H_{n}(Y)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1).$ (43)
 For the fourth term, applying Lemma 2.21 shows that
@@
 $\qquad+H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})+O(\log_{+}c_{1})+O_{\phi}(1).$
-Compared with (39), the second entropy can be bounded by
-$H_{n}(p(X_{1})+p(X_{2})-q(Y_{1})-q(Y_{2}))\leq 3H_{n}(p(X_{1})-q(Y_{1}))-H_{n}(X_{1})-H_{n}(Y_{1})+O(1).$
-The term $H_{n}(p(X_{1})-q(Y_{1}))$ can be bounded by using (15), which gives
-$H_{n}(p(X_{1})-q(Y_{1}))\leq 3H_{n}(p(X_{1})+q(Y_{1}))-H_{n}(p(X_{1}))-H_{n}(q(Y_{1}))+O(1)$
-$\quad=3H_{n}(p(X)+q(Y))-H_{n}(X)-H_{n}(Y)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1).$
+Compared with (43), Lemma 2.29 shows that
+$H_{n}(p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}))\leq 9H_{n}(p(X)+q(Y))$
+$\qquad\qquad\qquad-4H_{n}(X)-4H_{n}(Y)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1).$
 It leads to
@@
 Theorem 1.2’.
-Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$, $C\geq 1$ and $c>0$.
+Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$, $C,c>0$.
 There is a positive constant $\epsilon_{2}=\epsilon_{2}(s)$ such that the following statement holds: For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
@@
 The proof is completed. ∎
-Next, we prove Theorem 1.3 in the following form.
+4.3 The general forms
+Any non-degnerate rational quadratic from $\phi(x,y)=a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ can be diagonalized by rational linear change of variables, i.e.,
+$\phi(x,y)=\alpha(\xi_{11}x+\xi_{12}y)^{2}+\beta(\xi_{21}x+\xi_{22}y)^{2}$ (56)
+for some $\alpha,\beta\in\mathbb{Q}^{\ast}$ and $\xi_{11},\xi_{12},\xi_{21},\xi_{22}\in\mathbb{Q}$ with $\xi_{11}\xi_{22}-\xi_{12}\xi_{21}\neq 0$.
+One can verify that $\phi(x,y)$ is diagonal if and only if exactly two of the coefficients $\xi_{11},\xi_{12},\xi_{21},\xi_{22}$ are zero.
+For non-diagonal case, at most one of $\xi_{11},\xi_{12},\xi_{21},\xi_{22}$ may be zero.
+In the following, we will show that, if exactly one of these numbers is zero, then we can transform $\phi(x,y)$ into another form like (56) with all corresponding coefficients non-zero.
+Now we assume without loss of generality that $\xi_{11},\xi_{21},\xi_{22}\neq 0$ and $\xi_{21}=0$.
+By dividing $\xi_{11}$ and $\xi_{12}$ in the two brackets, respectively, we may rewrite $\phi(x,y)=\alpha^{\prime}x^{2}+\beta^{\prime}(x+d^{\prime}y)^{2}$, with $\alpha^{\prime},\beta^{\prime},d^{\prime}\in\mathbb{Q}^{\ast}$.
+Now, let us take some $a,c\in\mathbb{Q}^{\ast}$ satisfying $a\neq c$, $a\neq 1/d^{\prime}$, $c\neq 1/d^{\prime}$ and
+$(d^{\prime}a-1)(d^{\prime}c-1)=-\alpha^{\prime}/\beta^{\prime}.$
+One can check that
+$\phi(x,y)=\frac{\beta^{\prime}d^{\prime}(1-cd^{\prime})}{a-c}(ax+y)^{2}+\frac{\beta^{\prime}d^{\prime}(ad^{\prime}-1)}{a-c}(cx+y)^{2}.$
+All coefficient in above formula are non-zero rational numbers.
+Furthermore, we can assume in (56) without loss of generality that $\xi_{11}\neq-\xi_{21}$ and $\xi_{12}\neq-\xi_{22}$, because non-zero constants can be multiplied inside both brackets.
+In all situations, all the coefficients depend on $a_{1},a_{2},a_{3}$.
+Now we prove Theorem 1.3 in the form of the following two theorems, which hold for more general forms.
 Theorem 1.3’.
 Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$ and $C_{1},C_{2},c>0$.
-There is a positive constants $\epsilon_{3}=\epsilon_{3}(s)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$.
+There is a positive constant $\epsilon_{3}=\epsilon_{3}(s)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$.
 Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
 Then
-(i) For all $n\in\mathbb{N}$,
 $\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\geq n(s+\epsilon_{3})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{s}(1).$
-(ii) For all $n\in\mathbb{N}$,
-$\max\left\{\frac{H_{n}(X+Y)+H_{n}(X-Y)}{2},\,H_{n}(XY)\right\}\geq n(s+\epsilon_{3})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{s}(1).$
+for all $n\in\mathbb{N}$.
 Proof.
-(i) Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, by Lemma 2.22 it is also jointly $(s_{1},s_{2},C_{1}C_{2})$-Frostman.
+Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, by Lemma 2.30 it is also jointly $(s_{1},s_{2},C_{1}C_{2})$-Frostman.
 Applying Theorem 1.2’, there is some $\epsilon_{2}=\epsilon_{2}(s)$ such that
@@
 The conclusion then follows with $\epsilon_{3}=\epsilon_{2}/2$.
-(ii) Notice that
-$xy=\frac{1}{4}(x+y)^{2}-\frac{1}{4}(x-y)^{2}.$
-Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, by Lemma 2.23 it is also jointly $(s_{1},s_{2},C^{\prime\prime})$-Frostman, with $\log C^{\prime\prime}=\log C_{1}C_{2}+O(1)$.
-Applying Theorem 1.4’ with $\alpha=1/4$, $\beta=-1/4$ and the random variable (X+Y,XY)(X+Y,X-Y), we obtain
-$\max\left\{\frac{H_{n}(X+Y)+H_{n}(X-Y)}{2},\,H_{n}(2X),\,H_{n}(XY)\right\}$ (52)
-$\qquad\qquad\geq n(s+\epsilon_{2})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1)$
-for some $\epsilon_{2}=\epsilon_{2}(s)$.
-If we switch the role of $X$ and $Y$, then similar conclusion holds, with $H_{n}(2X)$ replaced by $H_{n}(2Y)$.
-Therefore, the left-hand side of (52) can be replaced by
-$\max\left\{\frac{H_{n}(X+Y)+H_{n}(X-Y)}{2},\,\frac{H_{n}(2X)+H_{n}(2Y)}{2},\,H_{n}(XY)\right\}.$
-By Lemma 2.24, combining $s_{1}+s_{2}\geq 2s$, the second item satisfies
-$H_{n}(X,Y)\geq\frac{H_{n}(2X)+H_{n}(2Y)+O(1)}{2}+ns+O(\log_{+}C_{1}C_{2})+O_{s}(1).$
-Combining the fact
-$H_{n}(X,Y)=H_{n}(X+Y,X-Y)+O(1)\leq H_{n}(X+Y)+H_{n}(X-Y)+O(1),$
-the first conclusion follows by taking $\epsilon_{3}=\epsilon_{2}/2$. ∎
-4.3 The general forms
-Now we are ready to prove Theorem 1.2.
-It actually holds for any non-degenerate real quadratic form $\phi(x,y)=a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ that is congruent over $\mathbb{Q}$ to a non-degenerate real diagonal form, i.e.,
-$\phi(x,y)=\alpha(\xi_{11}x+\xi_{12}y)^{2}+\beta(\xi_{21}x+\xi_{22}y)^{2}$ (53)
-for some $\alpha,\beta\in{\mathbb{R}}^{\ast}$ and $\xi_{11},\xi_{12},\xi_{21},\xi_{22}\in\mathbb{Q}$ with $\xi_{11}\xi_{22}-\xi_{12}\xi_{21}\neq 0$.
-Here, all the constants $\alpha,\beta,\xi_{11},\xi_{12},\xi_{21},\xi_{22}$ depends on the coefficients $a_{1},a_{2},a_{3}$ in $\phi(x,y)$.
+$\log C^{\prime\prime}=\log C_{1}C_{2}+O_{\phi}(1)$.
+Theorem 1.3”.
+Suppose that $\phi(x,y)$ has the form (56), where $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $\xi_{11},\xi_{12},\xi_{13},\xi_{14}\in\mathbb{Q}^{\ast}$, $\xi_{11}\neq-\xi_{21}$, $\xi_{12}\neq-\xi_{22}$ and $\xi_{11}\xi_{22}-\xi_{12}\xi_{21}\neq 0$.
+Let $0<s<1$ and $C_{1},C_{2},c>0$.
+There is a positive constants $\epsilon_{4}=\epsilon_{4}(\phi,s)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$.
+Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
+Then
+$\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{4})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1).$
+for all $n\in\mathbb{N}$.
+Proof.
+Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, Lemma 2.30 tells that it is both marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman and jointly $(s_{1},s_{2},C_{1}C_{2})$-Frostman.
+It follows from Lemma 2.32(i) and the fact $s_{1}+s_{2}\geq 2s$ that
+$\min\big{\{}H_{n}(X)+H_{n}(Y),\,H_{n}(X,Y)\big{\}}\geq 2ns-\log C_{1}C_{2}+O(1).$ (57)
+In view of Lemma 2.32(ii), one has
+$\frac{H_{n}(X)+H_{n}(Y)}{2}\leq H_{n}(X,Y)-ns+O(\log_{+}C_{1}C_{2})+O(1).$
+It follows that
+$H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\leq H_{n}(X,Y)-2ns+O(\log_{+}C_{1}C_{2})+O_{\phi}(1).$ (58)
+Consider the random variable $(\xi_{11}X+\xi_{12}Y,\,\xi_{21}X+\xi_{22}Y)$.
+Its range is contained in $c^{\prime\prime},c^{\prime\prime}]^{2}$, with $c^{\prime\prime}=O_{\phi}(c)$.
+And Lemma 2.31 shows that it is jointly $(s_{1},s_{2},C^{\prime\prime})$-Frostman, with $\log C^{\prime\prime}=\log C_{1}C_{2}+O_{\phi}(1)$, since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman.
+By Lemma 1.2’, there is some $\epsilon_{2}=\epsilon_{2}(s)>0$ such that
+$\max\left\{\frac{H_{n}(\xi_{11}X+\xi_{12}Y)+H_{n}(\xi_{21}X+\xi_{22}Y)}{2},\,H_{n}\big{(}(\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big{)},\,H_{n}(\phi(X,Y))\right\}$
+$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\geq n(s+\epsilon_{2})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1)$ (59)
+for all $n\in\mathbb{N}$.
+Notice that $\xi_{11}\neq-\xi_{21}$ and $\xi_{12}\neq-\xi_{22}$. It follows from Theorem 2.23 that
+$H_{n}\big{(}(\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big{)}-\frac{H_{n}(X)+H_{n}(Y)}{2}$
+$\qquad\leq d_{1}\left(H_{n}\big{(}X+Y\big{)}-\frac{H_{n}(X)+H_{n}(Y)}{2}\right)+d_{2}\big{(}H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\big{)}+O_{\phi}(1)$ (60)
+for some $d_{1}=d_{1}(\phi)>0$ and $d_{2}=d_{2}(\phi)\geq 0$. By (57), (58) and (60), one gets
+$\begin{split}&d_{1}\cdot H_{n}(X+Y)+2d_{2}\cdot\frac{1}{2}H_{n}(X,Y)\geq H_{n}\big{(}(\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big{)}\\ &\qquad\qquad\qquad+(d_{1}+2d_{2}-1)ns+O(\log_{+}C_{1}C_{2})+O_{\phi}(1).\end{split}$ (61)
+Similar arguments show that
+$\begin{split}&d_{1}^{\prime}\cdot H_{n}(X+Y)+2d_{2}^{\prime}\cdot\frac{1}{2}H_{n}(X,Y)\geq H_{n}(\xi_{11}X+\xi_{12}Y)\\ &\qquad\qquad\qquad+(d_{1}^{\prime}+2d_{2}^{\prime}-1)ns+O(\log_{+}C_{1}C_{2})+O_{\phi}(1).\end{split}$ (62)
+for some $d_{1}^{\prime}=d_{1}^{\prime}(\phi)>0$ and $d_{2}^{\prime}=d_{2}^{\prime}(\phi)\geq 0$. And also
+$\begin{split}&d_{1}^{\prime\prime}\cdot H_{n}(X+Y)+2d_{2}^{\prime\prime}\cdot\frac{1}{2}H_{n}(X,Y)\geq H_{n}(\xi_{21}X+\xi_{22}Y)\\ &\qquad\qquad\qquad+(d_{1}^{\prime\prime}+2d_{2}^{\prime\prime}-1)ns+O(\log_{+}C_{1}C_{2})+O_{\phi}(1).\end{split}$ (63)
+for some $d_{1}^{\prime\prime}=d_{1}^{\prime\prime}(\phi)>0$ and $d_{2}^{\prime\prime}=d_{2}^{\prime\prime}(\phi)\geq 0$.
+Now the theorem follows by combining (59) and (61)-(63), with $\epsilon_{4}=\epsilon_{2}/\max\{d_{1}+2d_{2},d_{1}^{\prime}+2d_{2}^{\prime},d_{1}^{\prime\prime}+2d_{2}^{\prime\prime}\}$. ∎
+In the following, Theorem 1.4 will be proved in the following form.
 Theorem 1.4’ Let $\phi(x,y)$ be in the form (56).
-Let $0<s<1$, $C\geq 1$ and $c>0$.
+Let $0<s<1$, $C_{1},C_{2},c>0$.
 There is a positive constant $\epsilon_{4}=\epsilon_{4}(\phi,s)$ such that the following statement holds: Let $X$ and $Y$ be independent real random variables with ranges contained in [c,c][-c,c].
@@
 Then
-$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(x,y))\right\}\geq n(s+\epsilon)+O_{s}(\log_{+}C)+O(\log_{+}c)+O_{\phi,s}(1)$
+$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{4})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1)$
 for all $n\in\mathbb{N}$.
 Proof.
-By Lemma 2.23, the random variable $Z:=(\xi_{11}X+\xi_{12}Y,\xi_{21}X+\xi_{22}Y)$ is jointly $(s_{1},s_{2},C^{\prime\prime})$-Frostman, where $\log C^{\prime\prime}=\log C_{1}C_{2}+O_{\phi}(1)$.
-Moreover, the range of $Z$ is contained in $c^{\prime\prime},c^{\prime\prime}]^{2}$ with $c^{\prime\prime}=O_{\phi}(c)$.
-Applying Theorem 1.4’ with $Z$, we obtain some positive number $\epsilon_{2}=\epsilon_{2}(s)$ such that
-$\max\left\{\frac{H_{n}(\xi_{11}X+\xi_{12}Y)+H_{n}(\xi_{21}X+\xi_{22}Y)}{2},\,H_{n}\big{(}(\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big{)},\,H_{n}\big{(}\phi(X,Y)\big{)}\right\}$
-$\qquad\geq n(s+\epsilon_{2})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1)$ (54)
-for all $n\in\mathbb{N}$.
-Next, using Lemma 2.21, let us assume that $d=d(\xi_{11},\xi_{12},\xi_{21},\xi_{22})=d(\phi)$ is the positive number satisfying
-$H_{n}\big{(}(\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big{)}\leq dH_{n}(X+Y)-(d-1)\frac{H_{n}(X)+H_{n}(Y)}{2}+O(1).$ (55)
-Combining Lemma 2.24(i) and the fact that $s_{1}+s_{2}\geq 2s$, we see that the right-hand side of (55) is
-$\leq dH_{n}(X+Y)-(d-1)\frac{(ns_{1}-\log C_{1}+O(1))+(ns_{2}-\log C_{2}+O(1))}{2}+O(1)$
-$\leq dH_{n}(X+Y)-(d-1)ns+O_{\phi}(\log_{+}C_{1}C_{2})+O_{\phi}(1).$
-This leads to
-$H_{n}(X+Y)\geq\frac{H_{n}\big{(}(\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big{)}+(d-1)ns}{d}+O_{\phi}(\log_{+}C_{1}C_{2})+O_{\phi}(1).$
-Similarly, we can deduce that
-$H_{n}(X+Y)\geq\frac{H_{n}(\xi_{11}X+\xi_{12}Y)+(d^{\prime}-1)ns}{d^{\prime}}+O_{\phi}(\log_{+}C_{1}C_{2})+O_{\phi}(1),$
-and also
-$H_{n}(X+Y)\geq\frac{H_{n}(\xi_{21}X+\xi_{22}Y)+(d^{\prime\prime}-1)ns}{d^{\prime\prime}}+O_{\phi}(\log_{+}C_{1}C_{2})+O_{\phi}(1)$
-for some positive numbers $d^{\prime}=d^{\prime}(\phi)$ and $d^{\prime\prime}=d^{\prime\prime}(\phi)$. Now we deduce from (54) that
-$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{5})+O_{\phi,s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1)$
-for $\epsilon_{4}=\epsilon_{2}/(d+d^{\prime}+d^{\prime\prime})$ and all $n\in\mathbb{N}$. Theorem 1.2 then follows. ∎
+When $\phi(x,y)$ is diagonal, the conclusion can be deduced from Lemma 4.1.
+In the following, we concentrate on the non-diagonal case.
+By Lemma 2.30, the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman.
+Applying Theorem 1.3”, there is some positive constant $\epsilon_{4}=\epsilon_{4}(\phi,s)$ such that
+$\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{4})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1).$
+Combining (13), one has
+$H_{n}(X,Y)\leq H_{n}(X+Y)+H_{n}(X)+O(1)\leq 2H_{n}(X+Y)+O(1).$
+The theorem then follows. ∎
 Remark 4.3.
-In the conclusions of Lemma 4.1 and Theorems 1.2’-1.4’, the implied constants in the terms $O_{s}(\log_{+}C_{1}C_{2})$, $O_{\phi,s}(\log_{+}C_{1}C_{2})$ or $O_{s}(\log_{+}C)$ can be independent of $s$ if either of the following two situations happens: (1) we are only considering the case $1/2<s<1$; (2) the range of random variables are bounded away from 0 by some absolute constant.
+In the conclusions of Theorems 1.2’-1.4’, the implied constants in the terms $O_{s}(\log_{+}C)$ or $O_{s}(\log_{+}C_{1}C_{2})$ can be independent of $s$ if either of the following two situations happens: (1) we are only considering the case $s\gg 1$; (2) the range of random variables are bounded away from 0 by some absolute constant.
 We preserve the dependence on $s$ in these terms for the uniformity with the case $0<s\leq 1/2$.
@@
 ∎
-6 Proof of Theorems 1.5-1.7
-Proof of Theorem 1.5.
-Since $H^{s}(A)>0$, by the Frostman lemma [13, Theorem 2.7], there exists a probability $(s,1)$-Frostman measure $\mu$ supported on $A$.
-Let $X$ and $Y$ be i.i.d.
-random variables that both have distribution measure $\mu$.
-Theorem 1.2 gives us
-$\max\big{\{}H_{n}(X+Y),\,H_{n}(\phi(X,Y))\big{\}}\geq(1+\epsilon)ns$
-for sufficiently large $n$.
-By Lemma 2.4(ii), we deduce that
-$\log\#\{D_{n}(x)\mid x\in A+A\}\geq H_{n}(X+Y),$
-and
-$\log\#\{D_{n}(x)\mid x\in\phi(A,A)\}\geq H_{n}(\phi(X,Y)).$
-Recall that $\mathcal{E}_{2^{-n}}(A+A)$ and $\mathcal{E}_{2^{-n}}(\phi(A,A))$ are the least numbers of intervals of length $2^{-n}$ needed to cover $A+A$ and $\phi(A,A)$, respectively.
-It is clear that
-$\log\mathcal{E}_{2^{-n}}(A+A)\geq H_{n}(X+Y),~~\log\mathcal{E}_{2^{-n}}(\phi(A,A))\geq H_{n}(\phi(X,Y)).$
-Hence,
-$\max\left\{\frac{\log\mathcal{E}_{2^{-n}}(A+A)}{n},\frac{\log\mathcal{E}_{2^{-n}}(\phi(A,A))}{n}\right\}\geq\frac{1}{n}\max\left\{H_{n}(X+Y),H_{n}(\phi(X,Y))\right\}\geq(1+\epsilon)s.$
-When $n\to\infty$, we obtain
-$\max\{\underline{\dim}_{B}(A+A),\,\overline{\dim}_{B}(\phi(A,A))\}\geq(1+\epsilon)s,$
-and
-$\max\{\overline{\dim}_{B}(A+A),\,\underline{\dim}_{B}(\phi(A,A))\}\geq(1+\epsilon)s.$
-This completes the proof. ∎
-Proof of Theorem 1.6.
-We still choose a small number $\eta=\eta(\phi)$ with $0<\eta<1/10$ satisfying $A_{\eta\delta}+A_{\eta\delta}$ and $\phi(A_{\eta\delta},A_{\eta\delta})$ are contained in the $(A+A)_{\delta/10}$ and $(\phi(A,A))_{\delta/10}$, respectively, for any $\delta>0$.
-Since $A$ is $\delta$-separated, the Lebesgue measure of $A_{\eta\delta}$ is $\gg_{\phi}\delta^{1-s}$.
-Let $X$ be a random variable that has uniform distribution on $A_{\eta\delta}$.
-Combining the non-concentration condition, one deduces that
-$P(X\in B(x,r))\ll_{\phi}\begin{cases}r\delta^{s-1},\quad&\text{if }0<r\leq\delta,\\ Cr^{s},\quad&\text{if }\delta<r\leq 1,\\ 1,\quad&\text{if }r>1,\end{cases}$
-for any ball $B(x,r)$. In all the cases, the right-hand side can be bounded by $Cr^{s}$. So $X$ is $(s,O_{\phi}(C))$-Frostman.
-Let $Y$ be a random variable that is i.i.d.
-with $X$.
-Now similar relations as in (58), (61) and (62) hold.
-As a replacement of (60), we apply Theorem 1.2 to get an $\epsilon^{\prime}=\epsilon^{\prime}(s,\phi)$ such that
-$\max\{H_{n}(X+Y),\,H_{n}(\phi(x,y))\}\geq n(s+\epsilon^{\prime})+O_{\phi,s}(\log C)+O_{\phi,s}(1).$
-for all $n\in\mathbb{N}$. Putting everything together, we arrive at the conclusion that
-$\max\{\mathcal{E}_{\delta}(A+A),\,\mathcal{E}_{\delta}(\phi(A,A))\}\gg_{\phi,s}C^{-O_{\phi,s}(1)}\delta^{-\epsilon}(\#A)$
-for all $\delta>0$. ∎
-Proof of Theorem 1.7.
-Let $\epsilon=\epsilon(s)$ to be determined later.
+6 Proof of Theorems 1.51.7
 For any sets $E\subseteq{\mathbb{R}}$, $F\subseteq{\mathbb{R}}^{2}$, and any positive number $\lambda$, denote by $E_{\lambda}$ and $F_{\lambda}$ the rectangular $\lambda$-neighbourhood of $E$ and $F$, respectively.
 More concretely, $E_{\lambda}=\bigcup\nolimits_{x\in E}(x-\lambda,x+\lambda)$ and $F_{\lambda}=\bigcup\nolimits_{(x,y)\in F}(x-\lambda,x+\lambda)\times(y-\lambda,y+\lambda)$.
-Note that $f_{1}(x,y)=x+y$ and $f_{2}(x,y)=\phi(x,y)=\alpha x^{2}+\beta y^{2}$ are $O_{\phi}(1)$-Lipchitz on $0,1]^{2}$.
-We can choose a small number $\eta=\eta(\phi)$ with $0<\eta<1/10$ such that $A_{\eta\delta}+_{G}A_{\eta\delta}$ and $\alpha\cdot A_{\eta\delta}^{\wedge 2}+\beta\cdot A_{\eta\delta}^{\wedge 2}$ are contained in the $(A+_{G}A)_{\delta/10}$ and $(\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2})_{\delta/10}$, respectively, for any $\delta>0$.
+Proof of Theorem 1.5.
+Write $\phi(x,y)=\alpha x^{2}+\beta y^{2}$. Let $\epsilon=\epsilon(s)$ to be determined later.
+Note that $(x,y)\mapsto x+y$ and $(x,y)\mapsto\phi(x,y)$ are both $O_{\phi}(1)$-Lipchitz on $0,1]^{2}$.
+We can choose a small number $\eta=\eta(\phi)$ with $0<\eta<1/10$ such that $A_{\eta\delta}+_{G}A_{\eta\delta}$ and $\phi_{G}(A_{\eta\delta},A_{\eta\delta})$ are contained in the $(A+_{G}A)_{\delta/10}$ and $(\phi_{G}(A,A))_{\delta/10}$, respectively, for any $\delta>0$.
 Then
-$\begin{split}\mathcal{E}_{\delta}(A+_{G}A)&\geq\frac{1}{2}\mathcal{E}_{\delta}(A_{\eta\delta}+_{G}A_{\eta\delta}),\\ \mathcal{E}_{\delta}(\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2})&\geq\frac{1}{2}\mathcal{E}_{\delta}(\alpha\cdot A_{\eta\delta}^{\wedge 2}+_{G}\beta\cdot A_{\eta\delta}^{\wedge 2}).\end{split}$ (58)
+$\begin{split}\mathcal{E}_{\delta}(A+_{G}A)&\gg\mathcal{E}_{\delta}(A_{\eta\delta}+_{G}A_{\eta\delta}),\\ \mathcal{E}_{\delta}(\phi_{G}(A,A))&\gg\mathcal{E}_{\delta}(\phi_{G}(A_{\eta\delta},A_{\eta\delta}))).\end{split}$ (66)
 Moreover, since $A$ is $\delta$-separated, we have
 $\mathcal{E}_{\delta}(A_{\eta\delta})=\mathcal{E}_{\delta}(A)=\#A=\lceil\delta^{-s}\rceil.$ (67)
-Now consider the set $G_{\eta\delta}\subseteq A_{\eta\delta}\times A_{\eta\delta}$.
-Since $A$ is $\delta$-separated, and $|G|\geq\delta^{\epsilon}|A|^{2}\geq\delta^{-2s+\epsilon}$, the $2$-dimensional Lebesgue measure of $G_{\eta\delta}$ is $\gg\eta^{2}\delta^{2-2s+\epsilon}\gg_{\phi}\delta^{2-2s+\epsilon}$.
-Let $(X,Y)$ be a random variable that has uniform distribution on $G_{\eta\delta}$. Combining the non-concentration condition, one can verify that
-[…]
-for any pair of balls $B(x,r_{1})$ and $B(y,r_{2})$.
+Now consider the set $G_{\eta\delta}\subseteq A_{\eta\delta}\times A_{\eta\delta}$. Since $A$ is $\delta$-separated, and $|G|\geq\delta^{\epsilon}|A|^{2}\geq\delta^{-2s+\epsilon}$, then
+$\mathcal{L}^{2}(G_{\eta\delta})\gg\eta^{2}\delta^{2-2s+\epsilon}\gg_{\phi}\delta^{2-2s+\epsilon}.$ (68)
+Let $(X,Y)$ be a random variable that has uniform distribution on $G_{\eta\delta}$. For any pair of balls $B(x,r_{1})$ and $B(y,r_{2})$, we have
+$P(X\in B(x,r_{1}),Y\in B(y,r_{2}))\ll\frac{\mathcal{L}^{1}(B(x,r_{1})\cap A_{\eta\delta})\cdot\mathcal{L}^{1}(B(y,r_{2})\cap A_{\eta\delta})}{\mathcal{L}^{2}(G_{\eta\delta})}.$
+When $\delta<r_{2}\leq 1$, the concentration conditions gives
+$\mathcal{L}^{1}(B(y,r_{2})\cap A_{\eta\delta})\ll Cr_{2}^{s}(\#A)\delta\ll cr_{2}^{s}\delta^{1-s}.$
+Calculation reveals that
+$P(X\in B(x,r_{1}),Y\in B(y,r_{2}))\ll_{\phi}\begin{cases}r_{1}\cdot r_{2}\cdot\delta^{2s-2-\epsilon},\quad&\text{if }0<r_{1},r_{2}\leq\delta,\\ r_{1}\cdot Cr_{2}^{s}\delta^{1-s}\cdot\delta^{2s-2-\epsilon}&\text{if }0<r_{1}\leq\delta<r_{2}\leq 1,\\ r_{1}\cdot\delta^{1-s}\cdot\delta^{2s-2-\epsilon},\quad&\text{if }0<r_{1}\leq\delta,\,r_{2}>1,\\ Cr_{1}^{s}\delta^{1-s}\cdot Cr_{2}^{s}\delta^{1-s}\cdot\delta^{2s-2-\epsilon}&\text{if }\delta<r_{1},r_{2}\leq 1\\ Cr_{1}^{s}\delta^{1-s}\cdot\delta^{1-s}\cdot\delta^{2s-2-\epsilon},\quad&\text{if }\delta<r_{1}\leq 1<r_{2},\\ 1,\quad&\text{if }r_{1},r_{2}>1,\end{cases}$
+and the upper bounds in other cases can be obtained by symmetry.
 In all the cases, the right-hand side can be bounded by $C^{2}\delta^{-\epsilon}r_{1}^{s}r_{2}^{s}$.
-So $(X,Y)$ is jointly $(s,s,C^{2}\delta^{-\epsilon})$-Frostman.
-By Theorem 1.4, there exists an $\epsilon^{\prime}=\epsilon^{\prime}(s)$ such that
-$\begin{split}&\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\\ &\qquad\qquad\geq n(s+\epsilon^{\prime})+O_{s}(\log_{+}C)+O_{s}(\epsilon\log(1/\delta))+O_{\phi,s}(1).\end{split}$ (60)
+So $(X,Y)$ is jointly $(s,s,O_{\phi}(C^{2}\delta^{-\epsilon}))$-Frostman.
+By Theorem 1.2’, there exists an $\epsilon^{\prime}=\epsilon^{\prime}(s)$ such that
+$\begin{split}&\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\geq n(s+\epsilon^{\prime})+O_{s}(\log_{+}C)+O_{s}(\epsilon\log(1/\delta))+O_{\phi,s}(1).\end{split}$ (69)
 for all $n\in\mathbb{N}$. In particular, we take $n$ such that $2^{-n-1}<\delta\leq 2^{-n}$.
@@
 Moreover,
-$\begin{split}&\quad\mathcal{E}_{\delta}(A_{\eta\delta})\geq\mathcal{E}_{2^{-n}}(A_{\eta\delta})\geq\frac{1}{2}\max\left\{\#\,range_{+}(D_{n}(X)),\,\#\,range_{+}(D_{n}(Y))\right\},\\ &\qquad\;\;\mathcal{E}_{\delta}(A_{\eta\delta}+_{G}A_{\eta\delta})\geq\mathcal{E}_{2^{-n}}(A_{\eta\delta}+_{G}A_{\eta\delta})\geq\frac{1}{2}\#\,range_{+}(D_{n}(X+Y)),\\ &\mathcal{E}_{\delta}(\alpha\cdot A_{\eta\delta}^{\wedge 2}+_{G}\beta\cdot A_{\eta\delta}^{\wedge 2})\geq\mathcal{E}_{2^{-n}}(\alpha\cdot A_{\eta\delta}^{\wedge 2}+_{G}\beta\cdot A_{\eta\delta}^{\wedge 2})\geq\frac{1}{2}\#\,range_{+}(\alpha X^{2}+\beta Y^{2}).\end{split}$ (62)
-Putting (58)-(62) together, one arrives at
-$\max\{\mathcal{E}_{\delta}(A),\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\alpha\cdot A^{2}+_{G}\beta\cdot A^{2})\}$
+$\begin{split}&\quad\mathcal{E}_{\delta}(A_{\eta\delta})\geq\mathcal{E}_{2^{-n}}(A_{\eta\delta})\gg\max\left\{\#\,range_{+}(D_{n}(X)),\,\#\,range_{+}(D_{n}(Y))\right\},\\ &\qquad\;\;\mathcal{E}_{\delta}(A_{\eta\delta}+_{G}A_{\eta\delta})\geq\mathcal{E}_{2^{-n}}(A_{\eta\delta}+_{G}A_{\eta\delta})\gg\#\,range_{+}(D_{n}(X+Y)),\\ &\qquad\quad\mathcal{E}_{\delta}(\phi_{G}(A_{\eta\delta},A_{\eta\delta}))\geq\mathcal{E}_{2^{-n}}(\phi_{G}(A_{\eta\delta},A_{\eta\delta}))\gg\#\,range_{+}(\phi(X,Y)).\end{split}$ (71)
+Putting (66), (67), and (69)- (71) together, one arrives at
+$\max\{\mathcal{E}_{\delta}(A),\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\}$
 $\qquad\gg 2^{n(s+\epsilon^{\prime})+O_{s}(\log_{+}C)+O_{s}(\epsilon\log(1/\delta))+O_{\phi,s}(1))}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-(s+\epsilon^{\prime}-O_{s}(\epsilon))}.$ (72)
 Now we can take $\epsilon=\epsilon(s)>0$ to be sufficiently small so that $\epsilon^{\prime}-O_{s}(\epsilon)\geq\epsilon$ on the right-hand side of (72). Then
-$\max\{\mathcal{E}_{\delta}(A),\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\alpha\cdot A^{2}+_{G}\beta\cdot A^{2})\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-(s+\epsilon)}$
+$\max\{\mathcal{E}_{\delta}(A),\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-(s+\epsilon)}$ (73)
 for any $\delta>0$.
 Finally, by (67) one has $\mathcal{E}_{\delta}(A)\ll\delta^{-s}$. This gives
-$\max\{\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\alpha\cdot A^{2}+_{G}\beta\cdot A^{2})\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A)$
+$\max\{\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A)$
 whenever $\delta$ is sufficiently small depending on $\phi,s$ and $C$. The proof is completed. ∎
+Proof of Theorem 1.6.
+Similar arguments as above will be applied. We only mention the differences that is needed to make.
+Since $\{y:\,(x,y)\in G\}\geq\delta^{\epsilon}(\#A)\gg\delta^{-s+\epsilon}$ for any $x\in A$, one has $G\gg\delta^{-2s+\epsilon}$.
+Also, one has $G\leq(\#A)^{2}\ll\delta^{-2s}$.
+As an alternative of (68), we have
+$\delta^{2-2s+\epsilon}\ll_{\phi}\mathcal{L}^{2}(G_{\eta\delta})\ll_{\phi}\delta^{2-2s}.$
+And we have
+$\mathcal{E}_{\delta}(G_{\eta\delta})=\mathcal{E}_{\delta}(G)=\#G\leq(\#A)^{2}\ll\delta^{-2s}$
+instead of (67).
+Let $(X,Y)$ be a random variable that has uniform distribution on $G_{\eta\delta}$. Let $E$ be any Borel set and $B(y,r)$ be any ball. We have
+$P(X\in E)\gg_{\phi}\mathcal{L}^{1}(X\cap A_{\eta\delta})\cdot\delta^{1-s+\epsilon}\cdot\delta^{2s-2}.$
+And
+$P(X\in E,Y\in B(y,r))\ll_{\phi}\begin{cases}\mathcal{L}^{1}(X\cap A_{\eta\delta})\cdot r\cdot\delta^{2s-2-\epsilon},\quad&\text{if }0<r\leq\delta,\\ \mathcal{L}^{1}(X\cap A_{\eta\delta})\cdot Cr^{s}\delta^{1-s}\cdot\delta^{2s-2-\epsilon}&\text{if }\delta<r\leq 1,\\ \mathcal{L}^{1}(X\cap A_{\eta\delta})\cdot\delta^{1-s}\cdot\delta^{2s-2-\epsilon},&\text{if }r>1.\end{cases}$
+In all the cases,
+$P(X\in E,Y\in B(y,r))\ll C\delta^{-2\epsilon}r^{s}P(X\in E)$
+So $Y$ is conditionally $(s,O_{\phi}(C\delta^{-2\epsilon}))$-Frostman given $X$.
+Similarly $X$ is also conditionally $(s,O_{\phi}(C\delta^{-2\epsilon}))$-Frostman given $Y$.
+Then $(X,Y)$ is conditionally $(s,O_{\phi}(C\delta^{-2\epsilon});s,O_{\phi}(C\delta^{-2\epsilon}))$-Frostman.
+Now (69) can be replaced by
+$\begin{split}&\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\geq n(s+\epsilon^{\prime})+O_{s}(\log C)+O_{s}(\epsilon\log(1/\delta))+O_{\phi,s}(1)\end{split}$
+for some $\epsilon^{\prime}=\epsilon^{\prime}(\phi,s)>0$, which results from Theorem 1.3’. Moreover, we substitute
+$\log\#range_{+}(D_{n}(X,Y))\geq H_{n}(X,Y)$
+and
+$\mathcal{E}_{\delta}(G_{\eta\delta})\geq\mathcal{E}_{2^{-n}}(G_{\eta\delta})\gg\#range_{+}(D_{n}(X,Y))$
+for the first formula in (70) and (71), respectively. Hence (73) can be replaced by
+$\max\left\{\mathcal{E}_{\delta}(G)^{1/2},\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\right\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-(s+\epsilon)}.$
+Finally, since $\mathcal{E}_{\delta}(G)\ll\delta^{-2s}$, we conclude that
+$\max\left\{\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\right\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-(s+\epsilon)}$
+whenever $\delta$ is sufficiently small depending on $\phi$, $s$ and $C$. ∎
+Proof of Theorem 1.7.
+We use similar arguments as in the proof of Theorem 1.5, and work with $G=B\times B$.
+Since $B\geq\delta^{\epsilon}\#A\geq\delta^{-s+\epsilon}$, one has $\mathcal{L}^{1}(B_{\eta\delta})\gg\delta^{1-s+\epsilon}$.
+Let $(X,Y)$ be a random variable that has uniform distribution on $G_{\eta\delta}=B_{\eta\delta}\times B_{\eta\delta}$.
+It is not hard to see that $X$ and $Y$ are i.i.d..
+Combining the non-concentration condition, one deduces that
+$P(X\in B(x,r))\ll_{\phi}\begin{cases}r\cdot\delta^{s-1-\epsilon},\quad&\text{if }0<r\leq\delta,\\ Cr^{s}\delta^{1-s}\cdot\delta^{s-1-\epsilon},\quad&\text{if }\delta<r\leq 1,\\ 1,\quad&\text{if }r>1,\end{cases}$
+for any ball $B(x,r)$. In all the cases, the right-hand side can be bounded by $C\delta^{-\epsilon}r^{s}$. So $X$ is $(s,O_{\phi}(C\delta^{-\epsilon}))$-Frostman.
+As a replacement of (69), Theorem 1.4’ shows that
+$\max\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\}\geq n(s+\epsilon^{\prime})+O_{s}(\log C)+O_{\phi,s}(1)$
+for some $\epsilon^{\prime}=\epsilon^{\prime}(s)>0$ and for all $n\in\mathbb{N}$. And the similar conclusion as in (73) appears that
+$\max\{\mathcal{E}_{\delta}(B+B),\,\mathcal{E}_{\delta}(\phi(B,B))\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A)$
+for all $\delta>0$. ∎
+Remark 6.1.
+In the conclusions of Theorems 1.5-1.7, the term $C^{-O_{s}(1)}$ can be replaced by $C^{-O(1)}$ if either of the following two situations happens: (1) we are only considering the case $s\gg 1$; (2) the condition $A\subseteq[0,1]$ is replaced by $A\subseteq[1,2]$.
 7 Examples
-In this section, we show details of some examples. Sometimes, we will write $\iota(x)=-x\log x$ for convenience.
+In this section, we show details of some examples. Recall that $\mathcal{L}^{d}$ is the $d$-dimensional Here measure. Sometimes, we will write $\iota(x)=-x\log x$ for convenience.
 Example 7.1.
@@
 random variables $X$ and $Y$ whose ranges are contained in $[0,1]$, such that $(X,Y)$ is marginally $(s,O(1);s,O(1))$-Frostman, and compare $H_{n}(X)$, $H_{n}(X+Y)$, $H_{n}(X^{2}+Y^{2})$ and $H_{n}(XY)$.
-Indeed, let $m=m(n)\in\mathbb{N}$ and $l=l(n)\in\mathbb{N}\cup\{0\}$ be parameters to be chosen later. We suppose that $(X,Y)$ has the uniform distribution on $A\times A$, with
+Indeed, let $m=m(n)\in\mathbb{N}$ and $l=l(n)\in\mathbb{N}\cup\{0\}$ be parameters to be chosen later. We suppose that $(X,Y)$ has the uniform distribution on $G=A\times A$, with
 $A=\bigcup\limits_{j=0}^{2^{m}-1}\left[j2^{-m},\,j2^{-m}+2^{-m-l}\right).$
-Denote by $\mathcal{L}^{d}$ the $d$-dimensional Lesbegue measure. Then $\mathcal{L}^{1}(A)=2^{-l}$. We see that both $X$ and $Y$ have Note that $X$ has uniform distribution on $A$, i.e.,
+Then $\mathcal{L}^{1}(A)=2^{-l}$. We see that both $X$ and $Y$ have Note that $X$ has uniform distribution on $A$, i.e.,
 $P(X\in E)=\frac{\mathcal{L}^{1}(E\cap A)}{\mathcal{L}^{1}(A)}$
-for any Borel set $E$.
-In the following, we will proof the following results.
+for any Borel set $E$. In the following, we will proof several results. The $o(1)$-terms are in the situation that $n\rightarrow\infty$.
 Claim 1. The random variable $(X,Y)$ is marginally $\left(s_{1},O(1);s_{2},O(1)\right)$-Frostman whenever $0\leq s_{1},s_{2}\leq\frac{m}{l+m}$.
@@
 $A=\bigcup\limits_{j=0}^{2^{m}-1}\left[1+j2^{-m},\,1+j2^{-m}+2^{-2m}\right),$
-and $Y$ has uniform distribution on $0,2^{-2m}]$. One can verify that $X$ is $(1/2,O(1))$-Frostman, and $Y$ is $(1/2,O(2^{m}))$-Frostman. We take $s=1/2$ and consider $n=2m$.
-Claim. We have $\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}=ns+O(1)$.
-In this example, we do not obtain a positive constant $\epsilon$ in Theorem 1.2.
+and $Y$ has uniform distribution on $0,2^{-2m}]$. One can verify that $X$ is $(1/2,O(1))$-Frostman, and $Y$ is both $(1/2,O(2^{m}))$-Frostman and $(0,1)$-Frostman. We take $n=2m$.
+Claim. We have $\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}=n/2+O(1)$.
+In this example, we do not obtain a positive constant $\epsilon$ in Theorem 1.4 if $(X,Y)$ is regarded as a marginally $(1/2,O(1);1/2,O(2^{m}))$-Frostman random variable, i.e., $s=1/2$.
 The reason is that the Frostman constant for $Y$ grows with $m$.
@@
 Therefore, in the conclusion of Theorem 1.4, it is essential to require $n>N$, where $N$ depends on the Frostman constant $C$.
+More concretely, the term $O_{s}(\log_{+}C_{1}C_{2})$ in the conclusion in Theorem 1.4’ is essential.
+On the other hand, if we regard $(X,Y)$ as a marginally $(1/2,O(1);0,1)$-Frostman random variable, then $s=1/4$ and Theorem 1.4 works.
 Proof of Claim.
@@
 $H_{n}(X^{\prime 2}+Y^{\prime 2})=H_{n}(X^{\prime 2})=H_{n}(X^{\prime})+O(1).$
-The conclusion follows by noting that $H_{n}(X^{\prime})=m=ns$. ∎
+The conclusion follows by noting that $H_{n}(X^{\prime})=m=n/2$. ∎
 Example 7.5.
 This example is an variant of Example 7.4.
-Let $W=(X,Y)$, and suppose that the joint distribution of $(X,Y)$ is the uniform distribution on the set $A=\bigcup\nolimits_{j=0}^{2^{m}-1}\big{[}j2^{-m},\,j2^{-m}+2^{-2m}\big{)}^{2}$.
+Let $W=(X,Y)$, and suppose that the joint distribution of $(X,Y)$ is the uniform distribution on the set $G=\bigcup\nolimits_{j=0}^{2^{m}-1}\big{[}j2^{-m},\,j2^{-m}+2^{-2m}\big{)}^{2}$.
 Note that $X$ and $Y$ are dependent.
-The marginal distributions of $X$ and $Y$ are both the uniform distribution on $\bigcup\nolimits_{j=0}^{2^{m}-1}\big{[}j2^{-m},\,j2^{-m}+2^{-2m}\big{)}$.
-So both $X$ and $Y$ are $(1/2,O(1))$-Frostman.
 Let $W_{1}=(X_{1},Y_{1}),\,(X_{2},Y_{2})$ be i.i.d.
 copies of $(X,Y)$.
-We consider $n\leq 2m$.
-Claim 1. We have $\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}\leq m+O(1)$.
+We consider $m\leq n\leq 2m$.
+Claim 1.
+The random variable $(X,Y)$ is marginally marginally $(1/2,O(1);1/2,O(1))$-Frostman, and jointly $(1/4,1/4,O(1))$-Frostman, but is not conditionally $(s_{1},O(1);s_{2},O(1))$-Frostman for any $(s_{1},s_{2})\neq(0,0)$.
+Claim 2. We have $\max\{H_{2m}(X+Y),\,H_{2m}(X^{2}+Y^{2})\}\leq m+O(1)$, $H_{2m}(X,Y)=m+O(1)$ and $H_{2m}(X-Y)=H_{2m}(X^{2}-Y^{2})=O(1)$.
 Claim 3. The random variable $W$ is $\left(s,O(2^{m(2s-1)})\right)$-Frostman for any $1/2\leq s\leq 2$.
 Claim 4. For $m\leq n\leq 2m$, we have $H_{n}(|W_{1}-W_{2}|)=(1+o(1))m$.
-If we take $n=2m$ in Claim 1, then $\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}\leq n/2+O(1)$.
+Now we consider the situation $n=2m$.
+By Claim 2, one has $\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}\leq n/2+O(1)$.
 In such a situation, the two entropies can be small simultaneously.
-Therefore, the requirement of independence of $X$ and $Y$ is essential in Theorem 1.2.
-Moreover, for the random variable $W$, its Frostman constant $C$ can grow with $m$.
+Therefore, the requirement of independence of $X$ and $Y$ in Theorem 1.4 is essential.
+Moreover, if we apply Theorem 1.2 with the random variable (X,Y)(X,-Y), and take $\alpha=1$, $\beta=-1$, $s_{1}=s_{2}=s=1/4$, then
+$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X-Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\geq n(s+\epsilon)$
+for some $\epsilon>0$.
+By Claim 2, the second and third entropies in above maximum-term nearly vanish.
+So the first term $(H_{n}(X)+H_{n}(Y))/2$ is essential.
+In this example, it is $m+O(1)$, which is obviously larger than $ns=m/2$.
+Furthermore, we can not substitute $\frac{1}{2}H(X,Y)=m/2$ for $(H_{n}(X)+H_{n}(Y))/2$ as in Theorem 1.3.
+Furthermore, for the random variable $W$, its Frostman constant $C$ can grow with $m$.
 If we consider $n=2m$ in Theorem 2.38, if $1<s<3/2$, then one has $H_{n}(|W_{1}-W_{2}|)=(1+o(1))n/2<nt$ for some suitably chosen $t$ with s/2<t<s1/2s/2<t<s-1/2.
-That is to say, the term $\log C$ in Theorem 2.30 is essential.
+That is to say, the term $\log_{+}C$ in Theorem 2.38 is essential.
+Proof of Claim 1.
+The marginal distributions of $X$ and $Y$ are both the uniform distribution on $A=\bigcup\nolimits_{j=0}^{2^{m}-1}\big{[}j2^{-m},\,j2^{-m}+2^{-2m}\big{)}$.
+Example 7.1 shows that $(X,Y)$ is marginally $(1/2,O(1);1/2,O(1))$-Frostman.
+Note that $\mathcal{L}^{2}(G)=2^{-3m}$ and $\mathcal{L}^{1}(A)=2^{-m}$. For any pair of balls $B(x,r_{1})$ and $B(y,r_{2})$, one has
+$P(X\in B(x,r_{1}),\,Y\in B(y,r_{2}))\ll\begin{cases}r_{1}r_{2}2^{3m},\quad&\text{if }0<r_{1}\leq r_{2}\leq 2^{-2m},\\ r_{1}2^{m},\quad&\text{if }0<r_{1}\leq 2^{-2m}<r_{2},\\ 2^{-m},\quad&\text{if }2^{-2m}<r_{1}\leq 2^{-m},\,r_{2}\geq r_{1},\\ r_{1},\quad&\text{if }2^{-m}<r_{1}\leq 1,\,r_{2}\geq r_{1},\\ 1,\quad&\text{if }r_{2}\geq r_{1}>1,\end{cases}$
+and the other cases with $r_{2}<r_{1}$ can be obtained by symmetry.
+One can verify that all the terms on the right-hand side of above formula can be bounded by $r_{1}^{1/4}r_{2}^{1/4}$.
+So $(X,Y)$ is jointly $(1/4,1/4,O(1))$-Frostman.
+For the conditional Frostman condition, one can check that
+$P(X\in[0,2^{-2m}))=P(X\in[0,2^{-2m}),\,Y\in B(0,2^{-2m}))=2^{-m}.$
+The only possible positive $s_{2}$ satisfying $2^{-m}\ll 2^{-m}(2^{-2m})^{s_{2}}$ for all sufficiently large $m$ is 0.
+Now it is not hard to see that $(X,Y)$ is $(s_{1},o(1);s_{2},O(1))$-Frostman only when $(s_{1},s_{2})=(0,0)$.
+∎
 Proof of Claim 2.
@@
 Note that
-$range(X^{\prime}+Y^{\prime})\subseteq\{j2^{-m+1}:\,j=0,1,\ldots,2^{m}-1\},\text{ and}$
-$range(X^{\prime 2}+Y^{\prime 2})\subseteq\{j^{2}2^{-2m+1}:\,j=0,1,\ldots,2^{m}-1\}.$
+$range_{+}(X^{\prime}+Y^{\prime})\subseteq\{j2^{-m+1}:\,j=0,1,\ldots,2^{m}-1\},$
+$range_{+}(X^{\prime 2}+Y^{\prime 2})\subseteq\{j^{2}2^{-2m+1}:\,j=0,1,\ldots,2^{m}-1\},$
+and
+$range_{+}(X^{\prime}-Y^{\prime})=range_{+}(X^{\prime 2}-Y^{\prime 2})=\{0\}.$
 This gives
-$\max\{H_{n}(X^{\prime}+Y^{\prime}),\,H_{n}(X^{\prime 2}+Y^{\prime 2})\}\leq m.$
+$\max\{H_{n}(X^{\prime}+Y^{\prime}),\,H_{n}(X^{\prime 2}+Y^{\prime 2})\}\leq m,$
+and
+$H_{n}(X^{\prime}-Y^{\prime})=H_{n}(X^{\prime 2}-Y^{\prime 2})=0.$
+The conclusion then follows. ∎
 Proof of Claim 3.
@@
 Example 7.6.
-This example is a further variant of Examples 7.4 and 7.5.
-Let $W=(X,Y)$, and suppose that the joint distribution of $(X,Y)$ is the uniform distribution on the set $A=\bigcup\nolimits_{j=0}^{2^{m}-1}\big{(}j2^{-m},\,j2^{-m}+2^{-2m}\big{)}\times\big{(}-j2^{-m},\,-j2^{-m}+2^{-2m}\big{)}$.
-Then the marginal distributions of $X$ and $Y$ are the uniform distribution on the sets $A_{0}=\bigcup\nolimits_{j=0}^{2^{m}-1}\big{(}j2^{-m},\,j2^{-m}+2^{-2m}\big{)}$ and $A_{0}^{\prime}=\bigcup\nolimits_{j=0}^{2^{m}-1}\big{(}-j2^{-m},\,-j2^{-m}+2^{-2m}\big{)}$, respectively.
-So both $X$ and $Y$ are $(1/2,O(1))$-Frostman.
-Claim 1. The random variable $(X,Y)$ is jointly $(1/4,1/4,O(1))$-Frostman, but is not conditionally $(s_{1},O(1);s_{2},O(2))$-Frostman for some $(s_{1},s_{2})\neq(0,0)$.
-Claim 2. Taking $n=2m$, one has $H_{n}(X+Y)=H_{n}(X^{2}-Y^{2})=O(1)$, and $H_{n}(X,Y)=m+O(1)$.
-In particular, if we apply Theorem 1.4 with $(X,Y)$, take $\alpha=-\beta$, $s_{1}=s_{2}=s=1/4$ and $n=2m$, then the second and third entropies in the maximum-term nearly vanish.
-So the first term $(H_{n}(X)+H_{n}(Y))/2$ is essential.
-In this example, it is $m+O(1)$, which is obviously larger than $ns=m/2$.
-Furthermore, we can not substitute $\frac{1}{2}H(X,Y)=m/2$ for $(H_{n}(X)+H_{n}(Y))/2$ as in Theorem 1.3.
-Proof of Claim 1.
-Note that $\mathcal{L}^{2}(A)=2^{-3m}$ and $\mathcal{L}^{1}(A_{0})=2^{-m}$. Let $E$ be a Borel set and $B(y,r)$ be any ball. When $0<r<2^{-2m}$, we have
-$P(X\in E,Y\in B(y,r))\ll\mathcal{L}^{1}(E\cap A_{0})\cdot r\cdot 2^{3m},$
-while
-$P(X\in E)\leq\mathcal{L}^{1}\cdot 2^{m}.$
-It follows that
-$P(X\in E,Y\in B(y,r))\ll r^{1/2}P(X\in E).$
-For any pair of balls $B(x,r_{1})$ and $B(y,r_{2})$, one has
-$P(X\in B(x,r_{1}),\,Y\in B(y,r_{2}))\ll\begin{cases}r_{1}r_{2}2^{3m},\quad&\text{if }0<r_{1}\leq r_{2}\leq 2^{-2m},\\ r_{1}2^{m},\quad&\text{if }0<r_{1}\leq 2^{-2m}<r_{2},\\ 2^{-m},\quad&\text{if }2^{-2m}<r_{1}\leq 2^{-m},\,r_{2}\geq r_{1},\\ r_{1},\quad&\text{if }2^{-m}<r_{1}\leq 1,\,r_{2}\geq r_{1},\\ 1,\quad&\text{if }r_{2}\geq r_{1}>1,\end{cases}$
-and the other cases with $r_{2}<r_{1}$ can be obtained by symmetry. One can verify that all the terms on the right-hand side of above formula can be bounded by $r_{1}^{1/4}r_{2}^{1/4}$.
-For the conditional Frostman condition, we take $E=B(x,2^{-2m})$ and $F=B(y,2^{-2m})$ such that $E\times F\subseteq A$. Then
-$P(X\in E)=2^{-m}=P(X\in E,\,Y\in B(x,2^{-2m}))=2^{-m}.$
-The only possible positive $s_{2}$ satisfying $2^{-m}\leq 2^{-m}(2^{-2m})^{s_{2}}$ is 0. Now it is not hard to see that $(X,Y)$ is $(s_{1},O(1);s_{2},O(1))$-Frostman only when $(s_{1},s_{2})=(0,0)$. ∎
-Proof of Claim 2.
-We use similar arguments as in the proof of Claim 1 in Example 7.5.
-Let $(X^{\prime},Y^{\prime})=D_{2m}(X,Y)$.
-Then $(X^{\prime},Y^{\prime})$ is the discrete uniform distribution on $\{(j2^{-m},-j2^{-m}:\,j=0,1,\ldots,2^{m}-1\}$.
-So $H_{n}(X^{\prime},Y^{\prime})=m$.
-Moreover, $range(X^{\prime}+Y^{\prime})=range(X^{\prime 2}-Y^{\prime 2})=\{0\}$, which implies that $H_{n}(X^{\prime}+Y^{\prime})=H_{n}(X^{\prime 2}-Y^{\prime 2})=0$.
-Applying similar relation as in (66), one asserts that $H_{n}(X+Y)=H_{n}(X^{2}-Y^{2})=O(1)$.
-Similarly, $H_{n}(X,Y)=m+O(1)$.
-∎
-Example 7.7.
 Let
-$B=\bigcup\limits_{i=0}^{2^{m}-1}\bigcup\limits_{j=0}^{2^{m}-1}[i2^{-m}+j2^{-2m},i2^{-m}+(j+1)2^{-2m})\times[j2^{-m}+i2^{-2m},i2^{-m}+(i+1)2^{-2m}).$
-Then, for each given $x\in[0,1)$ (or $y\in[0,1)$), the set $\{y:\,(x,y)\in B\}$ (or $\{x:\,(x,y)\in B\}$, respectively) is exactly an interval of length $2^{-2m}$. Now Let
-$A=\bigcup\limits_{i^{\prime},j^{\prime}=0}^{2^{m}-1}(i2^{-m},j2^{-m})+2^{-2m}\cdot B.$
-Here $2^{-2m}\cdot B=\{(2^{-2m}x,2^{-2m}y):\,(x,y)\in B\}$. Suppose that $(X,Y)$ is a random variable that has uniform distribution on $A$.
+$G^{\prime}=\bigcup\limits_{i=0}^{2^{m}-1}\bigcup\limits_{j=0}^{2^{m}-1}[i2^{-m}+j2^{-2m},i2^{-m}+(j+1)2^{-2m})\times[j2^{-m}+i2^{-2m},i2^{-m}+(i+1)2^{-2m}).$
+Then, for each given $x\in[0,1)$ (or $y\in[0,1)$), the set $\{y:\,(x,y)\in G^{\prime}\}$ (or $\{x:\,(x,y)\in G^{\prime}\}$, respectively) is exactly an interval of length $2^{-2m}$. Now Let
+$G=\bigcup\limits_{i^{\prime},j^{\prime}=0}^{2^{m}-1}(i2^{-m},j2^{-m})+2^{-2m}\cdot G^{\prime}.$
+Here $2^{-2m}\cdot G^{\prime}=\{(2^{-2m}x,2^{-2m}y):\,(x,y)\in G^{\prime}\}$. Suppose that $(X,Y)$ is a random variable that has uniform distribution on $G$.
 Claim 1. The random variable $(X,Y)$ is marginally $(1/2,O(1);1/2,O(1))$-Frostman.
@@
 Proof of Claim 1.
-One can check that both $X$ and $Y$ have the uniform distribution on the set $A_{0}=\bigcup\nolimits_{j=0}^{2^{m}-1}\left[j2^{-m},\,j2^{-m}+2^{-m-l}\right)$.
+One can check that both $X$ and $Y$ have the uniform distribution on the set $A=\bigcup\nolimits_{j=0}^{2^{m}-1}\left[j2^{-m},\,j2^{-m}+2^{-m-l}\right)$.
 By Example 7.1, they are $(1/2,O(1))$-Frostman.
@@
 Example 7.7.
-By Lemma 2.21, we have the following statement: Let $k,l\in\mathbb{Q}$ be given. Then
+By Theorem 2.23, we have the following statement: Let $k,l\in\mathbb{Q}$ be given. Then
 $H_{n}(kX+lY)-H_{n}(X)\ll_{k,l}(H_{n}(X+Y)-H_{n}(X))+O_{k,l}(1)$ (77)

### v4

@@
 Abstract
-We introduce Frostman conditions for bivariate random variables and study discretized entropy sum–product phenomena in both independent and dependent regimes.
-We prove that for any non-degenerate rational quadratic form $\phi(x,y)$ and parameter $0<s<1$, there exists a positive constant $\epsilon=\epsilon(\phi,s)$ such that
+We introduce Frostman conditions for bivariate random variables and study discretized entropy sum-product phenomena in both independent and dependent settings.
+Fix $0<s<1$, and let $(X,Y)$ be a bivariate real random variable with bounded support, whose distribution satisfies a Frostman condition of dimension $s$.
+Let $\phi(x,y)$ be a polynomial obtained from a diagonal polynomial $\rho_{1}(x)+\rho_{2}(y)\in\mathbb{R}[x,y]$ of degree $d\geq 2$ by applying an invertible rational linear change of variables in $(x,y)$.
+We show that there exists $\epsilon=\epsilon(\phi,s)>0$ such that
 $\max\{H_{n}(X+Y),H_{n}(\phi(X,Y))\}\geq n(s+\epsilon)$
-for sufficiently large $n$, where the precise conditions on $(X,Y)$ depend on the Frostman level.
+for all sufficiently large $n$, where the precise assumptions on $(X,Y)$ depend on the Frostman level.
 The proof introduces a novel multi‑step entropy framework, combining the submodularity formula, the discretized entropy Balog-Szemerédi-Gowers theorem, and state-of-the-art results on the Falconer distance problem, to reduce general forms to a diagonal core case.
-As an application, we derive a result on a discretized sum-product type problem.
-In particular, for a $\delta$-separated set $A\subset[0,1]$ of cardinality $\delta^{-s}$, satisfying some non-concentration conditions, and a dense subset $G\subset A\times A$, there exists $\epsilon=\epsilon(s,\phi)>0$ such that
+As an application, we obtain discretized sum-product type estimates.
+In particular, for a $\delta$-separated set $A\subseteq[0,1]$ of cardinality $\delta^{-s}$, satisfying certain non-concentration conditions, and a dense subset $G\subseteq A\times A$, there exists $\epsilon=\epsilon(s,\phi)>0$ such that
 $\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\phi_{G}(A,A))\gg\delta^{-\epsilon}(\#A)$
-for all $\delta$ small enough. Here by $\mathcal{E}_{\delta}(A)$ we mean the $\delta$-covering number of $A$, $\{A+_{G}A:=\{x+y\colon(x,y)\in G\}$, and $\phi_{G}(A):=\{\phi(x,y)\colon(x,y)\in G\}$.
+for all $\delta$ small enough. Here $\mathcal{E}_{\delta}(A)$ denotes the $\delta$-covering number of $A$, $A+_{G}A:=\{x+y\colon(x,y)\in G\}$, and $\phi_{G}(A,A):=\{\phi(x,y)\colon(x,y)\in G\}$.
 1 Introduction
@@
 $P(X\in B(x,r))\leq Cr^{s}.$
-For each $n\in\mathbb{N}$, let $D_{n}\colon\mathbb{R}\to\mathbb{R}$ be the function defined by $D_{n}(x)=\frac{k}{2^{n}}$ if $k$ is the unique integer such that $\frac{k}{2^{n}}\leq x<\frac{k+1}{2^{n}}$.
+For each $n\in\mathbb{N}$, let $D_{n}\colon\mathbb{R}\to\mathbb{R}$ be the function defined by $D_{n}(x)=\frac{k}{2^{n}}$ where $k$ is the unique integer satisfying $\frac{k}{2^{n}}\leq x<\frac{k+1}{2^{n}}$.
 For any random variable $X$, $D_{n}(X)$ defines a discrete random variable that takes values in the set of dyadic rationals.
@@
 In this paper, we establish a general entropy inequality that both sharpens our theoretical understanding and points toward a range of applications.
-Our study is motivated by recent applications of entropy methods to discretized sum-product estimates [12] due to Máthé and O’Regan, and to the recent solution of the polynomial Freiman-Ruzsa conjecture [2, 3] due to Gowers, Green, Manners, and Tao.
+Our study is motivated by recent applications of entropy methods to discretized sum-product phenomena [12] due to Máthé and O’Regan, and to the recent solution of the polynomial Freiman-Ruzsa conjecture [2, 3] due to Gowers, Green, Manners and Tao.
 More precisely, we are interested in bounding the following quantity $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ from below, where $\phi(x,y)=a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ is a non-degenerate rational quadratic form, i.e.
@@
 Then there exists a positive constant $\epsilon=\epsilon(s)$ and an integer $N=N(s,C,c_{1})$ such that for all $n>N,$ we have
-$\max\left\{H_{n}(X+Y),\,H_{n}(X\cdot Y)\right\}\geq n(s+\epsilon).$ (1)
-To state our main theorems, we first introduce the following definitions for bivariate random variables $(X,Y)$.
+$\max\left\{H_{n}(X+Y),\,H_{n}(X\cdot Y)\right\}\geq n(s+\epsilon).$
+To state the main theorems, we first introduce the following definitions for bivariate random variables $(X,Y)$.
 Definition 1.1.
@@
 $P(X\in B(x,r_{1}),\,Y\in B(y,r_{2}))\leq Cr_{1}^{s_{1}}r_{2}^{s_{2}}$
-for any ball $B(x,r_{1})$ and $B(y,r_{2})$.
+for any balls $B(x,r_{1})$ and $B(y,r_{2})$.
 (ii) We say that the random variable $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman if $X$ and $Y$ are $(s_{1},C_{1})$-Frostman and $(s_{2},C_{2})$-Frostman, respectively.
-(iii) We say a random variable $X$ is $(s,C)$-Frostman given $Y$ if
+(iii) We say a random variable $X$ is conditionally $(s,C)$-Frostman given $Y$ if
 $P(X\in B(x,r),\,Y\in E)\leq Cr^{s}P(Y\in E)$
 for any ball $B(x,r)$ and Borel set $E$.
-Moreover, we say that the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, if $X$ is $(s_{1},C_{1})$-Frostman given $Y$, and $Y$ is $(s_{2},C_{2})$-Frostman given $X$.
-These definitions are not equivalent. At a common exponent level $(s_{1},s_{2})$, they form a hierarchy:
+Moreover, we say that the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, if $X$ is conditionally $(s_{1},C_{1})$-Frostman given $Y$, and $Y$ is conditionally $(s_{2},C_{2})$-Frostman given $X$.
+These definitions are not equivalent. At a common exponent level $(s_{1},s_{2})$, these conditions form a hierarchy:
 $\boxed{\text{\parbox{116.65646pt}{\centering independent\\ and marginally Frostman\@add@centering}}}\;\Rightarrow\;\boxed{\text{conditionally Frostman}}\;\Rightarrow\;\boxed{\text{jointly Frostman}}\;\Rightarrow\;\boxed{\text{marginally Frostman}}$
@@
 Remarkably, different levels of this hierarchy yield entropy inequalities with different strengths and applications.
-Our results demonstrate how the Frostman hierarchy affects entropy bounds. We call a quadratic form diagonal if $\phi(x,y)=\alpha x^{2}+\beta y^{2}$, and the main results read as follows.
-Theorem 1.2 (Jointly Frostman, diagonal forms).
-Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$, $C\geq 1$ and $c>0$.
-There exist a positive constant $\epsilon=\epsilon(s)$ and an integer $N=N(\alpha,\beta,s,C,c)$ such that the following statement holds.
-For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
-$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\geq n(s+\epsilon)$
+The results in this paper demonstrate how the Frostman hierarchy affects entropy bounds.
+Theorem 1.2 (Jointly Frostman, diagonal polynomials).
+Let $d\geq 2$ and $\rho_{1}$, $\rho_{2}$ be real polynomials of degree $d$.
+Let $0<s<1$ and $C,c>0$.
+There exist a positive constant $\epsilon=\epsilon(d,s)$ and an integer $N=N(\rho_{1},\rho_{2},s,C,c)$ such that the following statement holds.
+For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
+$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}\big(\rho_{1}(X)+\rho_{2}(Y)\big)\right\}\geq n(s+\epsilon)$
 for all $n>N$.
+In the theorem above, a constant depends on a polynomial $\rho$, we means that it depends on the degree and coefficients of $\rho$.
+The notation $1\ll_{s}s_{1},s_{2}$ means that $s_{1}$ and $s_{2}$ are bounded below by some positive constant depending only on $s$, but not on $\rho_{1},\rho_{2},C,c$ or $n$.
+Next, consider polynomials $\rho_{1}$ and $\rho_{2}$ of degree $d\geq 2$, together with an invertible matrix
+$\Xi=\begin{bmatrix}\xi_{11}&\xi_{12}\\ \xi_{21}&\xi_{22}\end{bmatrix}\in GL_{2}(\mathbb{Q}).$
+Let
+$\phi(x,y)=\rho_{1}(\xi_{11}x+\xi_{12}y)+\rho_{2}(\xi_{21}x+\xi_{22}y).$ (1)
+As a basic example, all non-degenerate rational quadratic forms can be written in the form of (1).
+Note that $\xi_{11}\xi_{22}-\xi_{12}\xi_{21}\neq 0$.
+We say that $\phi(x,y)$ is diagonal if exactly two of the entries $\xi_{11}$, $\xi_{12}$, $\xi_{21}$, $\xi_{22}$ are zero.
+In such a situation, one has $\phi(x,y)=\tilde{\rho}_{1}(x)+\tilde{\rho}_{2}(y)$ for some polynomials $\tilde{\rho}_{1}(x)$ and $\tilde{\rho}_{2}(y)$ of degree $d$.
+And we say that $\phi(x,y)$ is non-diagonal if at most one of the entries $\xi_{11}$, $\xi_{12}$, $\xi_{21}$, $\xi_{22}$ is zero.
 Theorem 1.3 (Conditionally Frostman, general forms).
-Let $\phi(x,y)$ be a non-degenerate rational quadratic form, $0<s<1$ and $C_{1},C_{2},c>0$.
-There exist positive constants $\epsilon=\epsilon(s)$, $\epsilon^{\prime}=\epsilon^{\prime}(\phi,s)$ and an integer $N=N(\phi,s,C_{1},C_{2},c)$ such that the following statement holds.
+Let $\phi$ be given as in (1).
+Let $0<s<1$ and $C_{1},C_{2},c>0$.
+There exist positive constants $\epsilon=\epsilon(d,s)$, $\epsilon^{\prime}=\epsilon^{\prime}(d,\Xi,s)$ and an integer $N=N(\phi,s,C_{1},C_{2},c)$ such that the following statement holds.
 Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$.
-Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
+Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
 Then, for all $n>N$,
 $\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq\begin{cases}n(s+\epsilon),\quad&\text{if }\phi(x,y)\text{ is diagonal},\\ n(s+\epsilon^{\prime}),&\text{if }\phi(x,y)\text{ is non-diagonal}.\end{cases}$
-Here, a constant depends on $\phi$ means that it depends on the coefficients $a_{1}$, $a_{2}$ and $a_{3}$.
+Here, a constant depends on the matrix $\Xi$ means that it depends on the entries of $\Xi$. And a constant depends on $\phi$ means it depends on $\rho_{1}$, $\rho_{2}$ and $\Xi$.
 Theorem 1.4 (Marginally Frostman with independence).
-Let $\phi(x,y)$ be a non-degenerate rational quadratic form, $0<s<1$ and $C_{1},C_{2},c>0$.
+Let $\phi$ be given as in (1).
+Let $0<s<1$ and $C_{1},C_{2},c>0$.
 There exist positive constants $\epsilon=\epsilon(s)$, $\epsilon^{\prime}=\epsilon^{\prime}(\phi,s)$ and an integer $N=N(\phi,s,C_{1},C_{2},c)$ such that the following statement holds.
-Let $X$ and $Y$ be independent real random variables whose ranges are contained in [c,c][-c,c].
-Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};S_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
+Let $X$ and $Y$ be two independent real random variables whose ranges are contained in [c,c][-c,c].
+Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
 Then, for all $n>N$,
@@
   2. 2.
-    The positive $\epsilon$ might not exist if one of the conditions of the theorem is not satisfied, including when $s=0$ or 11 (see Example 7.1), the Frostman constant $C$ grows with $n$ (see Example 7.4), or the polynomial is degenerate (see Example 7.3).
+    The positive $\epsilon$ might fail to exist if one of the conditions of the theorem is not satisfied, including when $s=0$ or 11 (see Example 7.1), the Frostman constant $C$ grows with $n$ (see Example 7.4), or the polynomial is degenerate (see Example 7.3).
   3. 3.
@@
   6. 6.
-    The value of $\epsilon$ is not explicit, and an inspection of the proofs suggests it would be small.
-  7. 7.
-    Theorem 1.2 can be extended to a more general form of
-    $\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(f(X)+g(Y))\right\},$
-    where $f$ and $g$ are polynomials with $\deg(f)=\deg(g)\geq 2$.
-Establishing such a result would require extending most of the arguments in this paper, and we plan to provide further updates in a subsequent version.
-We pause here to add several further remarks.
+    The value of $\epsilon$ is not explicit, and our arguments suggest that they are quantitatively small.
+We record several further remarks.
 Our proof of Theorems 1.21.4 presents a connection between $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ and the entropy of the “distance” function, i.e.
@@
 It would be interesting to explore more connections between $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ and the entropy of other functions.
-In the following, we discuss some applications of the entropy theorems.
-For $A\subset\mathbb{R}$ and $\delta>0$, by $\mathcal{E}_{\delta}(A)$ we mean the $\delta$-covering number of $A$.
+We now discuss applications of the entropy theorems.
+For $A\subseteq\mathbb{R}$ and $\delta>0$, by $\mathcal{E}_{\delta}(A)$ we mean the $\delta$-covering number of $A$.
 For a set $G\subseteq A\times A$, we denote
 $A+_{G}A=\{x+y:\,(x,y)\in G\},\qquad\phi_{G}(A,A)=\{\phi(x,y):\,(x,y)\in G\},$
-$\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2}=\{\alpha x^{2}+\beta y^{2}:\,(x,y)\in G\}.$
+$\rho_{1}(A)+_{G}\rho_{2}(A)=\{\rho_{1}(x)+\rho_{2}(y):\,(x,y)\in G\}.$
 The next theorem presents an application of Theorem 1.2 on fractional covering numbers.
 Theorem 1.5.
-Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$ and $C\geq 1$.
-Then there exist positive constants $\epsilon=\epsilon(s)$ and $\delta_{0}=\delta_{0}(\alpha,\beta,s,C)$ such that the following holds for all $0<\delta\leq\delta_{0}$.
-Let $A\subset{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[0,1]}$ be any $\delta$-separated set of cardinality $\lceil\delta^{-s}\rceil$.
+Let $d\geq 2$ and $\rho_{1}$, $\rho_{2}$ be real polynomials of degree $d$.
+Let $0<s<1$ and $C\geq 1$.
+There exist constants $\epsilon=\epsilon(d,s)>0$ and $\delta_{0}=\delta_{0}(\rho_{1},\rho_{2},s,C)>0$ such that the following holds for all $0<\delta\leq\delta_{0}$.
+Let $A\subseteq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[0,1]}$ be any $\delta$-separated set of cardinality $\lceil\delta^{-s}\rceil$.
 Suppose that for all intervals $I$ of length $l$ at least $\delta$, we have the non-concentration estimate
@@
 Then, for all $G\subseteq A\times A$ with $G\geq\delta^{\epsilon}(\#A)^{2}$, we have
-$\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2})\gg_{\alpha,\beta,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A).$
-Thanks to Theorem 1.3, extend this result to all non-degenerate rational quadratic forms is possible, but requiring minimum degree constraints.
+$\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\rho_{1}(A)+_{G}\rho_{2}(A))\gg_{\rho_{1},\rho_{2},s}C^{-O_{d,s}(1)}\delta^{-\epsilon}(\#A).$
+By Theorem 1.3, extending this result to more general polynomials is possible, but requiring minimum degree constraints (3) and (4).
 Theorem 1.6.
-Let $\phi(x,y)$ be a non-degenerate rational quadratic form, $0<s<1$ and $C\geq 1$.
-Then there exist positive constants $\epsilon=\epsilon(\phi,s)$ and $\delta_{0}=\delta_{0}(\alpha,\beta,s,C)$ such that the following holds for all $0<\delta\leq\delta_{0}$.
-Let $A\subset{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[0,1]}$ be any $\delta$-separated set of cardinality $\lceil\delta^{-s}\rceil$.
+Let $\phi$ be given as in (1).
+Let $0<s<1$ and $C\geq 1$.
+There exists $\epsilon>0$, which depends on $d,s$ (or $d,\Xi,s$) when $\phi(x,y)$ is diagonal (or non-diagonal, respectively), and $\delta_{0}=\delta_{0}(\phi,s,C)>0$ such that the following holds for all $0<\delta\leq\delta_{0}$.
+Let $A\subseteq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[0,1]}$ be any $\delta$-separated set of cardinality $\lceil\delta^{-s}\rceil$.
 Suppose that for all intervals $I$ of length $l$ at least $\delta$, we have the non-concentration estimate
@@
 Then, for any subset $G\subseteq A\times A$ such that
-$\{y:\,(x,y)\in G\}\geq\delta^{\epsilon}(\#A)\quad\text{for any}\quad x\in A,$
-$\{x:\,(x,y)\in G\}\geq\delta^{\epsilon}(\#A)\quad\text{for any}\quad y\in A,$
+$\{y:\,(x,y)\in G\}\geq\delta^{\epsilon}(\#A)\quad\text{for any}\quad x\in A,$ (3)
+$\{x:\,(x,y)\in G\}\geq\delta^{\epsilon}(\#A)\quad\text{for any}\quad y\in A,$ (4)
 we have
-$\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\phi_{G}(A,A))\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A).$
-In contrast to Theorem 1.6, the next result, an application of Theorem 1.4, holds for every scale $\delta>0$ and is stable under taking subsets: it applies to any $B\subset A$ with $B\geq\delta^{\varepsilon}\#A$.
+$\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\phi_{G}(A,A))\gg_{\phi,s}C^{-O_{d,s}(1)}\delta^{-\epsilon}(\#A).$
+In contrast to Theorem 1.6, the next result, an application of Theorem 1.4, holds for every scale $\delta>0$ and is stable under taking subsets: it applies to any $B\subseteq A$ with $B\geq\delta^{\varepsilon}(\#A)$.
 Theorem 1.7.
-Let $\phi(x,y)$ be a non-degenerate rational quadratic form, $0<s<1$ and $C\geq 1$.
-Then there exist positive constants $\epsilon=\epsilon(\phi,s)$ such that the following holds for all $\delta>0$.
-Let $A\subset{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[0,1]}$ be any $\delta$-separated set of cardinality $\lceil\delta^{-s}\rceil$.
+Let $\phi$ be given as in (1).
+Let $0<s<1$ and $C\geq 1$.
+Then there exists $\epsilon>0$, which depends on $d,s$ (or $d,\Xi,s$) when $\phi(x,y)$ is diagonal (or non-diagonal, respectively), such that the following holds for all $\delta>0$.
+Let $A\subseteq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[0,1]}$ be any $\delta$-separated set of cardinality $\lceil\delta^{-s}\rceil$.
 Suppose that for all intervals $I$ of length $l$ at least $\delta$, we have the non-concentration estimate
@@
 Then, for any subset $B\subseteq A$ satisfying $B\geq\delta^{\epsilon}(\#A)$, we have
-$\mathcal{E}_{\delta}(B+B)+\mathcal{E}_{\delta}(\phi(B,B))\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A).$
+$\mathcal{E}_{\delta}(B+B)+\mathcal{E}_{\delta}(\phi(B,B))\gg_{\phi,s}C^{-O_{d,s}(1)}\delta^{-\epsilon}(\#A).$
 These applications are of independent interest, particularly when $\phi(x,y)$ is a non-bilinear form, and might be viewed as consequences of the Falconer distance problem.
@@
 $\mathcal{E}_{\delta}(A+A)+\mathcal{E}_{\delta}(A^{\wedge 2}+A^{\wedge 2})\gg_{\alpha,\beta,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A).$
-We now sketch the main ideas of proofs in this paper.
-Our base case is corresponding to the diagonal forms $\phi(X,Y)=\alpha X^{2}+\beta Y^{2}$ (Lemma 4.1), with independent $(s,C)$-Frostman random variables $X,Y$.
-Roughly speaking, by applying the submodularity formula (Lemma 2.18) on certain algebraic relations between the sum function $X+Y$ and the distance function $(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}$, we are able to reduce the problem to bounding the entropy of the distance function from below.
-In particular, let $W$ be an $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $0<s<1$, the constant $C\geq 1$ is bounded and does not depend on $n$, and $R>0$.
-Let $W^{\prime}$ be a random variable which is i.i.d.
-with $W$.
-Then there exists a constant $\epsilon_{0}=\epsilon_{0}(s)>0$ such that
-$H_{n}(|W-W^{\prime}|^{2})\geq n\left(s+\epsilon_{0}\right)$ (3)
-for all $n\geq N_{0}$, where $N_{0}=N_{0}(s,C,R)$.
-Note that if the Frostman constant $C$ depends on $n$, then the inequality (3) might not be true (see Example 7.5).
-To prove (3), we make use of results on the Falconer distance problem.
-While classical estimates taken from [13, Section 4.2] are sufficient in the case $s>1/2$, we invoke the recent work of Shmerkin and Wang [24] for $s\leq 1/2$.
-[…]
-[…]
-(ii) $X^{\prime}$ and $X^{\prime\prime}$ are conditionally i.i.d. given $D_{n}(Y^{\prime})$;
-(iii) $Y^{\prime}$ and $Y^{\prime\prime}$ are conditionally i.i.d. given $D_{n}(X^{\prime})$;
-(iv) $(X^{\prime\prime},Y^{\prime})$ and $Y^{\prime\prime}$ are conditionally independent given $D_{n}(X^{\prime})$.
-Then, given that $D_{n}(X^{\prime},Y^{\prime})=(x,y)$ with $x,y$ being dyadic rationals satisfying $P\big{(}D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}>0$, in the proof we show that
-$\begin{split}\big{(}X^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}\text{ is $\left(s,O\left(C^{2}\cdot 2^{-ns}/P_{E}(D_{n}(Y^{\prime})=y)\right)\right)$-Frostman},\\ \big{(}Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}\text{ is $\left(s,O\left(C^{2}\cdot 2^{-ns}/P_{E}(D_{n}(X^{\prime})=x)\right)\right)$-Frostman}.\end{split}$ (4)
-The problem now is reduced to the case of random variables $\big{(}X^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}$ and $\big{(}Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}$, and the corresponding entropy
-$\max\left\{H_{n}\left(X^{\prime\prime}+Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\right),\,H_{n}\left(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\right)\right\}.$
-It is clear that we have many pairs $(x,y)$, and for each pair, the corresponding Frostman constants in (4) are not uniformly bounded.
-Besides technical issues including Lipschitz conditions and partitions of measure supports, the main difficulty arises because the Frostman constants may grow with $n$.
-Note that the idea of scaling the Frostman constants to $1$ does not help, since it would change the supports of variables from bounded to unbounded.
-We address this challenge by explicitly computing the entropy of the distance function. Instead of (3), we have the following for both small and large dimensions:
-Large dimension: Let $W$ be an $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $1/2<s<1$, $C\geq 1$ and $R>0$.
-Let $W^{\prime}$ be a random variable which is i.i.d with $W$.
-Then for any number $t$ satisfying $t\leq 1$ and s<t<2s1/2s<t<2s-1/2, we have that for all $n\in\mathbb{N},$
-$H_{n}(|W-W^{\prime}|)\geq nt-\log C+O(|\log R|)+O(|\log(2s-t-1/2)|)+O_{t}(1).$ (5)
-Small dimension: Let $X^{\prime\prime}$, $Y^{\prime\prime}$ be independent random variables with ranges contained in [c,c][-c,c], with $c>0$.
-Suppose that $X^{\prime\prime}$ and $Y^{\prime\prime}$ are $(s,C_{1})$-Frostman and $(s,C_{2})$-Frostman, respectively, where $0<s\leq 1/2$ and $C_{1},C_{2}\geq 1$.
-Let $(X^{\prime\prime\prime},Y^{\prime\prime\prime})$ be i.i.d.
-with $(X^{\prime\prime},Y^{\prime\prime})$.
-Then for any $t$ with $s<t<\frac{s+\varphi(2s)}{2}$, for all $n\in\mathbb{N}$, we have
-$H_{n}\big{(}|(X^{\prime\prime},Y^{\prime\prime})-(X^{\prime\prime\prime},Y^{\prime\prime\prime})|\big{)}\geq nt+O_{s}(\log n)+O_{s,t}(\log C_{1}C_{2})+O_{s,c}(1),$ (6)
-where
-$\varphi(u)=\frac{u}{2}+\frac{\sqrt{4+u^{2}}-2}{2},\quad(0<u\leq 1).$ (7)
-By (5) and (6), the error terms involving the Frostman constants in (4) are controlled up to logarithmic level.
-Then they can be further converted to certain entropy-terms that are feasible to handle.
-The proof of (5) is essentially the same as that of (3).
-To prove (6), we adapt the Shmerkin–Wang approach with appropriate modifications, while clarifying, simplifying, and quantifying key steps.
-Note that our inspection is not verbatim: for instance, in the thin‐tube step we must work with sets having a Cartesian‐product structure.
-This is why the estimate (6) holds for $W=X^{\prime\prime}\times Y^{\prime\prime}$ rather than in the general form of (5).
-Plugging these inequalities into our method, the theorem follows.
-Our final comment is that entropy methods have also been used in a range of problems in Additive Combinatorics and Geometric Measure Theory, for example, see [4, 7, 10, 16, 19, 24, 25, 26] and references therein.
 2 Preliminaries
@@
 Let $(\Omega,\mathfrak{B},P)$ be a probability space.
-A random variable $X$ is a measurable function from $(\Omega,\mathfrak{B},P)$ to another measuable space $(\Omega^{\prime},\mathfrak{B}^{\prime})$, on which a pushforward probability measure $\mu$ is induced, i.e., $\mu(B^{\prime})=P(X\in B^{\prime})$ for any $B^{\prime}\in\mathfrak{B}^{\prime}$.
+A random variable $X$ is a measurable function from $(\Omega,\mathfrak{B},P)$ to another measurable space $(\Omega^{\prime},\mathfrak{B}^{\prime})$, on which a pushforward probability measure $\mu$ is induced, i.e., $\mu(B^{\prime})=P(X\in B^{\prime})$ for any $B^{\prime}\in\mathfrak{B}^{\prime}$.
 We will say that $\mu$ is the associated distribution measure of $X$.
 We denote the range of $X$ by $range(X)$, which collects all the possible values that $X$ can take.
-In this paper, all the measurable spaces will be taken to be subspaces of ${\mathbb{R}}^{m}$, equipped with the Borel $\sigma$-algebra or its sub-$\sigma$-algebra, the Euclidean norm $|\cdot|$, and the Lebesgue measure $\mathcal{L}^{m}(\cdot)$.
-The dimension $m$ is always regarded as an absolute constant.
+In this paper, all the measurable spaces will be taken to be subspaces of ${\mathbb{R}}^{m}$, equipped with the Borel $\sigma$-algebra or its sub-$\sigma$-algebra, the Euclidean norm $|\cdot|$, and the Lebesgue measure $\mathcal{L}^{m}(\cdot)$, where the dimension $m$ is considered as an absolute constant.
 The range of any random variable is assumed to be a bounded set.
@@
 In particular, when $\mu=P$, the conditional probability measure given an event $E$ is denoted by $P_{E}$.
-For a random varaible $X$ and such an $E$, we denote by $X_{E}$ the random variable satisfying $X_{E}\equiv(X|E)$ (see Lemma A.2).
+For a random variable $X$ and such an event $E$, we denote by $X_{E}$ a random variable satisfying $X_{E}\equiv(X|E)$ (see Lemma A.2).
 That is to say,
@@
 Throughout this paper, all the functions are assumed to be measurable.
-And whenever we write $f(X)$ with $X$ a random variable and $f$ a function, we always mean that $domain(f)\supseteq range_{+}(X)$, where the values taken on a zero-measure set is neglectable.
+And whenever we write $f(X)$ with $X$ a random variable and $f$ a function, we always mean that $domain(f)\supseteq range_{+}(X)$, where the values taken on a zero-measure set is negligible.
 In this paper, by $f=O_{\kappa}(g)$, $f\ll_{\kappa}g$ or $g\gg_{\kappa}f$, we mean that there exists an absolute positive constant $C=C(\kappa)$ depending on the parameter(s) $\kappa$ such that $|f|\leq Cg$.
@@
 $H(\mu,\mathcal{A}|\mathcal{B})=\sum_{B\in\mathcal{B}}\mu(B)H(\mu_{B},\mathcal{A}).$
-In above definition, the convention $0\log 0=0$ is used for convenience.
-We also say $\mu(B)H(\mu_{B},\mathcal{A})=0$ when $\mu(B)=0$, although $\mu_{B}$ is not defined in this situation.
+In this definition, we use the convention $0\log 0=0$.
+We also interpret $\mu(B)H(\mu_{B},\mathcal{A})$ as 0 when $\mu(B)=0$, even though $\mu_{B}$ is not defined in this situation.
 Intuitively, the entropy of a random variable will tell us how hard it is to determine a value.
@@
 Definition 2.3 (Entropy of random variables).
-Let $X$ and $Y$ be discrete random variables, and $E$ be an event. The entropy of $X$ is defined as
+Let $X$ and $Y$ be discrete random variables, and $E$ be an event with positive probability. The entropy of $X$ is defined as
 $H(X)=-\sum_{x}P(X=x)\log P(X=x).$
@@
 $H(X,Z)+H(X,W)\geq H(X,Z,W)+H(X).$
-So, by applying Lemma 2.4 (iii) and (8), we obtain
+Hence, by applying Lemma 2.4 (iii) and (5), we obtain
 $H(Z)+H(W)+H(X|Z)+H(X|W)+H(Y|Z,W)$
@@
 $p_{1}\leq p_{2}\leq\ldots\leq p_{k}\leq p_{k+1}.$
-Let us consider the the maximum value of the function
+Let us analyze the the maximum value of the function
 $f(p_{1},\ldots,p_{k+1})=-\sum\limits_{i=1}^{k+1}p_{i}\log p_{i},$
@@
 $\frac{\partial\tilde{f}}{\partial p_{i}}(p_{1},\ldots,p_{j-1},p_{j+1},\ldots,p_{k+1})=\log\frac{1-\sum\nolimits_{l\neq j}p_{l}}{p_{i}}=\log\frac{p_{j}}{p_{i}}\leq 0.$
-That is to say, when the other k1k-1 variables are given, the small $p_{i}$ is, the larger value $\tilde{f}$ and $f$ take.
-Now one can verify that the maximum value of $f$ is taken at $(p_{1},\ldots,p_{k+1})$ with
+In other words, when the other k1k-1 variables are given, the small $p_{i}$ is, the larger value $\tilde{f}$ and $f$ take.
+Now we can verify that the maximum value of $f$ is taken at $(p_{1},\ldots,p_{k+1})$ with
 $p_{i}=\frac{\lambda}{k}\,\,(1\leq i\leq k),\qquad p_{k+1}=1-\lambda.$
-That is to say,
+Thus,
 $\sum\limits_{i=1}^{k+1}p_{i}\log p_{i}\leq k\cdot\left(-\frac{\lambda}{k}\log\frac{\lambda}{k}\right)-(1-\lambda)\log(1-\lambda)\leq\lambda\log k+1.$
@@
 2.3 Discretized entropy and properties
-For each $n\in\mathbb{N}$ and $m\in\mathbb{N}$, let us define a function $D_{n,m}:\,{\mathbb{R}}^{m}\rightarrow{\mathbb{R}}^{m}$ as the following.
-For any $(x_{1},\ldots,x_{m})\in{\mathbb{R}}^{m}$, if $k_{i}$ is the unique integer such that $\frac{k_{i}}{2^{n}}\leq x_{i}<\frac{k_{i}+1}{2^{n}}$ $(i=1,\ldots,m)$, then we define $D_{n,m}(x_{1},\ldots,x_{m})=(\frac{k_{1}}{2^{n}},\ldots,\frac{k_{m}}{2^{n}})$.
-That is to say, $D_{n,m}(x)$ locates the left endpoints of the dyadic intervals that the $x_{i}$’s fall in.
-For convenience, we will simplify $D_{n,m}$ as $D_{n}$, where the dimension $m$ is determined from the context.
+For each $n\in\mathbb{N}$ and $m\in\mathbb{N}$, define $D_{n,m}:\,{\mathbb{R}}^{m}\rightarrow{\mathbb{R}}^{m}$ by
+$D_{n,m}(x_{1},\ldots,x_{m})=\left(\frac{k_{1}}{2^{n}},\ldots,\frac{k_{m}}{2^{n}}\right),$
+where for each $i$ the integer $k_{i}$ is uniquely determined by $\frac{k_{i}}{2^{n}}\leq x_{i}<\frac{k_{i}+1}{2^{n}}$.
+So, $D_{n,m}(x)$ locates the left endpoints of the dyadic intervals that the $x_{i}$’s fall in.
+For convenience, we often write $D_{n}$ instead of $D_{n,m}$ when the dimension is clear from the context.
 For any random variable $X$, the expression $D_{n}(X)$ is a discrete random variable that takes values in the set of points whose components are all dyadic rationals.
-Moreover, if $X_{1},\ldots,X_{k}$ are independent (or identically distributed) random variables, then $D_{n}(X_{1}),\ldots,D_{n}(X_{k})$ are also independent (or identically distributed, respectively).
+Moreover, if $X_{1},\ldots,X_{k}$ are independent (or identically distributed) random variables, then $D_{n}(X_{1}),\ldots,D_{n}(X_{k})$ are independent (or identically distributed, respectively).
 For random variables that are not discrete, we can discretize them with the function $D_{n}$. For convenience, for any random variables $X$ and $Y$, we denote the discretized entropy by
@@
 When $X$ and $Y$ are independent, then $D_{n}(X)$ and $D_{n}(Y)$ are also independent, which gives
-$H_{n}(X,Y)=H(D_{n}(X),D_{n}(Y))=H(D_{n}(X))+H(D_{n}(Y))=H_{n}(X)+H_{n}(Y).$ (9)
+$H_{n}(X,Y)=H(D_{n}(X),D_{n}(Y))=H(D_{n}(X))+H(D_{n}(Y))=H_{n}(X)+H_{n}(Y).$
 Generally, one has
-$H_{n}(X,Y)\leq H_{n}(X)+H_{n}(Y).$ (10)
+$H_{n}(X,Y)\leq H_{n}(X)+H_{n}(Y).$
 Combining Lemma A.2, one can verify that the statement in Lemma 2.5 also holds for $H_{n}$: If $X$ is a random variable and $\mathcal{B}$ is a finite partition. Then
 $H_{n}(X)\geq\sum\limits_{B\in\mathcal{B}}P(B)H_{n}(X_{B}).$ (6)
-Moreover, the discretized entropy only depends on the discretised distribution. That is to say, if $D_{n}(X)$ and $D_{n}(Y)$ are identically distributed, then $H_{n}(X)=H_{n}(Y)$.
+Moreover, the discretized entropy only depends on the discretized distribution. If $D_{n}(X)$ and $D_{n}(Y)$ are identically distributed, then $H_{n}(X)=H_{n}(Y)$.
 Furthermore, if $range_{+}(X)\subseteq B(0,R)\subseteq{\mathbb{R}}^{m}$ for some $R\geq 1$, then $range_{+}(D_{n}(X))$ has at most $O_{m}\big(R^{m}2^{mn}\big)$ elements.
@@
 Definition 2.10.
-Let $\kappa$ be a positive real number, and $f$ be a function.
-(i) We say that $f$ is $\kappa$-Lipschitz if $|f(x)-f(y)|\leq\kappa|x-y|$ for any $x,y\in domain(f)$.
-(ii) We say that $f$ is $\kappa$-bi-Lipschitz if $f$ is invertible, and both $f$ and $f^{-1}$ are $\kappa$-Lipschitz.
-(iii) If $f$ is $\kappa$-Lipschitz, and there exists a decomposition of $domain(f)=\mathcal{R}_{1}\cup\ldots\cup\mathcal{R}_{l}$, where $l$ is a positive integer viewed as an absolute constant, such that each $f|_{\mathcal{R}_{i}}$ has an $\kappa$-Lipschitz inverse $(i=1,\ldots,l)$, then we say that $f$ is weakly $\kappa$-bi-Lipschitz.
+Let $\kappa$ be a positive real number, $f$ be a function, and $\mathcal{R}$ be a region contained in $domain(f)$.
+(i) We say that $f$ is $\kappa$-Lipschitz on $\mathcal{R}$ if $|f(x)-f(y)|\leq\kappa|x-y|$ for any $x,y\in\mathcal{R}$.
+(ii) We say that $f$ is $\kappa$-bi-Lipschitz on $\mathcal{R}$ if $f|_{\mathcal{R}}$ is invertible, with $f$ and $(f|_{\mathcal{R}})^{-1}$ being $\kappa$-Lipschitz on $\mathcal{R}$ and $f(\mathcal{R})$, respectively.
+(iii) If $f$ is $\kappa$-Lipschitz on $\mathcal{R}$, and there exists a decomposition $\mathcal{R}=\mathcal{R}_{1}\cup\ldots\cup\mathcal{R}_{d}$, where $d$ is a positive integer, such that each $f|_{\mathcal{R}_{i}}$ has an $\kappa$-Lipschitz inverse on $f(\mathcal{R}_{i})$ $(i=1,\ldots,d)$, then we say that $f$ is weakly $(\kappa,d)$-bi-Lipschitz.
+For simplicity, we will abbreviate weakly $(\kappa;O(1))$-bi-Lipschitz as weakly $\kappa$-bi-Lipschitz. And we may omit the phrase “on $\mathcal{R}$” if it is clear from the context.
 Example 2.11.
-Let $D$ be an open convex region in ${\mathbb{R}}^{m}$.
-Let $f:\,D\rightarrow{\mathbb{R}}$ be a function such that $|\frac{\partial f}{\partial x_{i}}(x_{1},\ldots,x_{m})|\leq M$ for all $(x_{1},\ldots,x_{m})\in D$.
-Then, for any $(x_{1},\ldots,x_{m})$ and $(y_{1},\ldots,y_{m})$ in $D$, there is some $(\xi_{1},\ldots,\xi_{m})\in D$ such that
+Let $\mathcal{R}$ be an open convex region in ${\mathbb{R}}^{m}$.
+Let $f:\,\mathcal{R}\rightarrow{\mathbb{R}}$ be a function such that $|\frac{\partial f}{\partial x_{i}}(x_{1},\ldots,x_{m})|\leq M$ for all $(x_{1},\ldots,x_{m})\in\mathcal{R}$.
+Then, for any $(x_{1},\ldots,x_{m})$ and $(y_{1},\ldots,y_{m})$ in $\mathcal{R}$, there is some $(\xi_{1},\ldots,\xi_{m})\in\mathcal{R}$ such that
 $|f(x_{1},\ldots,x_{m})-f(y_{1},\ldots,y_{m})|\leq\sum\limits_{i=1}^{m}\left|\frac{\partial f}{\partial x_{i}}(\xi_{1},\ldots,\xi_{m})\right||x_{i}-y_{i}|\leq Mm^{1/2}\left(\sum\limits_{i=1}^{m}|x_{i}-y_{i}|^{2}\right)^{1/2}.$
@@
 Example 2.12.
-Let $\alpha,c_{1},c_{2}$ be real numbers with $\alpha\neq 0$ and $0<c_{2}<c_{1}$. Let $f(x)=\alpha x^{2}$ with domain $c_{1},-c_{2}]\cup[c_{2},c_{1}]$. Note that, by the mean value theorem,
+Let $\alpha,c_{1},c_{2}$ be real numbers with $\alpha\neq 0$ and $0<c_{2}<c_{1}$. Let $f(x)=\alpha x^{2}$. Note that, by the mean value theorem,
 $2|\alpha|c_{2}\cdot|x-y|\leq|f(x)-f(y)|\leq 2|\alpha|c_{1}\cdot|x-y|$
-whenever $x,y\in[-c_{1},-c_{2}]$, or $x,y\in[c_{2},c_{1}]$. One asserts that $f$ is weakly $\max\left\{(2|\alpha|c_{2})^{-1},\,2|\alpha|c_{1}\right\}$-bi-Lipschitz.
+whenever $x,y\in[-c_{1},-c_{2}]$, or $x,y\in[c_{2},c_{1}]$. One asserts that $f$ is weakly $\max\left\{(2|\alpha|c_{2})^{-1},\,2|\alpha|c_{1}\right\}$-bi-Lipschitz on $c_{1},-c_{2}]\cup[c_{2},c_{1}]$.
+The following lemma is a generalization of above example, which follows immediately by mean value theorem.
+Lemma 2.13.
+Let $\mathcal{R}\subseteq{\mathbb{R}}$ and $f:\,\mathcal{R}\rightarrow{\mathbb{R}}$ be a differentiable function.
+Suppose that $\mathcal{R}$ is a union of $d$ intervals such that $f$ is monotonic on each of the intervals.
+Let $M=\sup\nolimits_{x\in\mathcal{R}}|f^{\prime}(x)|$ and $m=\inf\nolimits_{x\in\mathcal{R}}|f^{\prime}(x)|$.
+Then $f$ is weakly $(\max\{M,m^{-1}\},d)$-bi-Lipschitz on $\mathcal{R}$.
 Definition 2.14.
-Let $X$, $Y$ be random variables and $\kappa$ be a postive real number.
-We say that $Y$ is a $\kappa$-Lipschitz function (or $\kappa$-bi-Lipschitz function, or weakly $\kappa$-bi-Lipschitz function) of $X$ if $Y=f(X)$ for some $\kappa$-Lipschitz function (or $\kappa$-bi-Lipschitz function, or weakly $\kappa$-bi-Lipschitz function, respectively) $f$ with $domain(f)\supseteq range_{+}(X)$.
+Let $X$, $Y$ be random variables and $\kappa$ be a positive real number.
+We say that $Y$ is a $\kappa$-Lipschitz function (or $\kappa$-bi-Lipschitz function, or weakly $(\kappa,d)$-bi-Lipschitz function) of $X$ if $Y=f(X)$ for some function $f$ that is $\kappa$-Lipschitz (or $\kappa$-bi-Lipschitz, or weakly $(\kappa,d)$-bi-Lipschitz, respectively) on $range_{+}(X)$.
 If a function is $\kappa$-Lipschitz with $0<\kappa\leq 2$, then it is also $2$-Lipschitz.
@@
 Let $X$ and $Y$ be random variables. We have the following two statements.
-(i) If $f$ is $\kappa$-Lipschitz, then for all $n\in\mathbb{N}$,
+(i) If $f$ is $\kappa$-Lipschitz on $range_{+}(X)$, then for all $n\in\mathbb{N}$,
 $H_{n}(f(X)|X)\ll\log\kappa.$
-(ii) If $g$ be weakly $\kappa$-bi-Lipschitz, then for all $n\in\mathbb{N}$,
-$H_{n}(Y|g(Y)))\ll\log\kappa.$
-The above implied constants are independent of $n$, and the concrete expressions of the functions and random variables.
+(ii) If $g$ is weakly $(\kappa,d)$-bi-Lipschitz on $range_{+}(Y)$, then for all $n\in\mathbb{N}$,
+$H_{n}(Y|g(Y)))\ll\log\kappa+\log d.$
+The implied constants are independent of $n$ and of the particular functions and random variables involved.
 Proof.
@@
 When $D_{n}(g(Y))$ is determined, one knows that $g(Y)$ takes value in a ball with radius $O_{m^{\prime}}(2^{-n})$.
-By the definition of weakly $\kappa$-bi-Lipschitz functions, we can locate $Y$ to be lying in the pre-image of this ball, which is a union of at most $O(1)$ balls with radius $O_{m^{\prime}}(\kappa\cdot 2^{-n})$.
-So $D_{n}(Y)$ has at most $O_{m,m^{\prime}}(\kappa^{m})$ choices.
+By the definition of weakly $(\kappa,d)$-bi-Lipschitz functions, we can locate $Y$ to be lying in the pre-image of this ball, which is a union of at most $d$ balls with radius $O_{m^{\prime}}(\kappa\cdot 2^{-n})$.
+So $D_{n}(Y)$ has at most $O_{m,m^{\prime}}(d\kappa^{m})$ choices.
 It follows from Lemma 2.4(ii) that
-$H_{n}(Y|g(Y))\leq\log(O_{m,m^{\prime}}(\kappa^{m}))\ll\log\kappa.$
+$H_{n}(Y|g(Y))\leq\log(O_{m,m^{\prime}}(d\kappa^{m}))\ll\log\kappa+\log d.$
 The proof is completed. ∎
@@
 We have the following two statements.
-(i) Let $X$ and $Y$ be random variables such that $X$ is an $\kappa$-Lipschitz function of $Y$. Then, for all $n\in\mathbb{N}$,
+(i) Let $X$ and $Y$ be random variables such that $X$ is a $\kappa$-Lipschitz function of $Y$. Then, for all $n\in\mathbb{N}$,
 $H_{n}(X)\leq H_{n}(Y)+O(\log\kappa).$
-(ii) Let $X$ and $Y$ be random variables such that $X$ is a weakly $\kappa$-bi-Lipschitz function of $Y$. Then, for all $n\in\mathbb{N}$,
-$H_{n}(X)=H_{n}(Y)+O(\log\kappa).$
+(ii) Let $X$ and $Y$ be random variables such that $X$ is a weakly $(\kappa,d)$-bi-Lipschitz function of $Y$. Then, for all $n\in\mathbb{N}$,
+$H_{n}(X)=H_{n}(Y)+O(\log\kappa)+O(\log d).$
 Proof.
 Regarding (i), by (5), it is sufficient to prove that $H_{n}(X|Y)\ll\log\kappa$, which is guaranteed by Lemma 2.15(i). Regarding (ii), by (5) and Lemma 2.15(ii), we conclude that
-$H_{n}(Y)\leq H_{n}(X)+H_{n}(Y|X)=H_{n}(X)+O(\log\kappa).$
+$H_{n}(Y)\leq H_{n}(X)+H_{n}(Y|X)=H_{n}(X)+O(\log\kappa)+O(\log d).$
 The conclusion then follows. ∎
-Example 2.16.
-Let $X$ be a random variable with $range_{+}(X)\subseteq[-c_{1},-c_{2}]\cup[c_{2},c_{1}]$, where $0<c_{2}<c_{1}$.
-Let $\alpha\in{\mathbb{R}}^{\ast}$.
-One can deduce from Example 2.12 and Lemma 2.15 that
-$H_{n}(\alpha X^{2})=H_{n}(X)+O\big{(}\max\{1,\log c_{1},\log(1/c_{2})\}\big{)}+O(\big{|}\log|\alpha|\big{|}).$
 Example 2.17.
@@
 Lemma 2.16 shows that $H_{n}(AX)\leq H_{n}(X)+O_{A}(1)$.
-In particular, if $A$ is invertible, then $H_{n}(AX)=H_{n}(X)+O_{A}(1)$.
+When $A$ is invertible, we further have $H_{n}(AX)=H_{n}(X)+O_{A}(1)$.
+In particular, for any random variable $X$, $Y$, we have
+$H_{n}(X,Y)=H_{n}(X,X+Y)+O(1)\leq H_{n}(X)+H_{n}(X+Y)+O(1).$ (8)
 The following is the discretized version of submodularity.
@@
 Lemma 2.19.
-Let $X$, $Y$ be random variables, and $f$ be a $\kappa$-Lipschitz function with $range(f)\subseteq{\mathbb{R}}^{m}$, where $m$ is viewed as an absolute constant.
+Let $X$, $Y$ be random variables.
+Suppose that $f$ is $\kappa$-Lipschitz on $range_{+}(X)\cup range_{+}(Y)$, and $range(f)\subseteq{\mathbb{R}}^{m}$, where $m$ is viewed as an absolute constant.
 Suppose that $|X-Y|\leq K$ with probability $1$.
@@
 As a consequence,
-$\max\{H_{n}(X),H_{n}(Y)\}\leq H_{n}(X+Y)+O(1)$ (13)
-for all $n\in\mathbb{N}$, provided that $X$ and $Y$ are independent. In the rest of this paper, we may apply Lemma 2.20 without mention it.
-In [16], Ruzsa studied some entropy inequalities that involving sums of random variables. In the following, more inequalities for discretized entropy are presented.
+$\frac{H_{n}(X)+H_{n}(Y)}{2}\leq\max\{H_{n}(X),H_{n}(Y)\}\leq H_{n}(X+Y)+O(1)$ (9)
+for all $n\in\mathbb{N}$, provided that $X$ and $Y$ are independent. In the rest of this paper, we may apply Lemma 2.20 without mentioning it.
+In [16], Ruzsa studied some entropy inequalities that involve sums of random variables. In the following, more inequalities for discretized entropy are presented.
 Lemma 2.21.
-Let $X,Y$ and $Z$ be independent real random variables. Then, for all $n\in\mathbb{N}$,
+Let $X,Y$ and $Z$ be three independent real random variables. Then, for all $n\in\mathbb{N}$,
 $H_{n}(X+Y+Z)$ $\leq H_{n}(X+Z)+H_{n}(Y+Z)-H_{n}(Z)+O(1),$
@@
 Combining independence, the second inequality involving XYX-Y then follows. ∎
-Without independence, the follow lemma can be used as an alternative.
+Without independence, the following lemma can be used as an alternative.
 Lemma 2.22.
@@
 Theorem 2.23.
-Let $k,l\in\mathbb{Q}^{\ast}$, and $X$ and $Y$ be real random variables that are not necessarily independent. Then there exist integers $d_{1}=d_{1}(k,l)>0$ and $d_{2}=d_{2}(k,l)\geq 0$ such that
+Let $k,l\in\mathbb{Q}$, and $X$, $Y$ be real random variables that are not necessarily independent. Then there exist integers $d_{1}=d_{1}(k,l)>0$ and $d_{2}=d_{2}(k,l)\geq 0$ such that
 $H_{n}(kX+lY)-\frac{H_{n}(X)+H_{n}(Y)}{2}\leq d_{1}\left(H_{n}(X+Y)-\frac{H_{n}(X)+H_{n}(Y)}{2}\right)$
@@
 $H_{n}(X,X^{\prime},Y)=2H_{n}(X,Y)-H_{n}(Y).$ (12)
-Recall Lemma 2.20. By (15), one has
+By (11), one has
 $H_{n}(X-Y)+H_{n}(X,X^{\prime},Y)\leq H_{n}(X+X^{\prime})+H_{n}(Y+X^{\prime})+H_{n}(X,Y)+O(1).$ (13)
@@
 Lemma 2.26.
-Let $k\in\mathbb{Z}^{\ast}$, and $X$, $Y$ be real random variables that are not necessarily independent. Then there are integers $d_{1}=d_{1}(k)>0$ and $d_{2}=d_{2}(k)\geq 0$ such that
+Let $k\in\mathbb{Z}$, and $X$, $Y$ be real random variables that are not necessarily independent. Then there are integers $d_{1}=d_{1}(k)>0$ and $d_{2}=d_{2}(k)\geq 0$ such that
 $H_{n}(kX+Y)-\frac{H_{n}(X)+H_{n}(Y)}{2}$ $\leq d_{1}\left(H_{n}(X+Y)-\frac{H_{n}(X)+H_{n}(Y)}{2}\right)+d_{2}\mathcal{I}_{n}(X,Y)+O_{k}(1).$
@@
 Proof.
+For $k=0$, one can obtain by (8) that the conclusion holds for $d_{1}=d_{2}=1$.
 For $k=1$, the conclusion holds for $d_{1}=1$ and $d_{2}=0$.
@@
 Lemma 2.27.
-Let $k,l\in\mathbb{Z}^{\ast}$, and $X$, $Y$ be real random variables that are not necessarily independent. Then there are integers $d_{1}=d_{1}(k,l)>0$ and $d_{2}=d_{2}(k,l)\geq 0$ such that
-$H_{n}(kX+lY)-\frac{H_{n}(X)+H_{n}(Y)}{2}$ $\leq d_{1}\left(H_{n}(X+Y)-\frac{H_{n}(X)+H_{n}(Y)}{2}\right)+d_{2}\mathcal{I}_{n}(X,Y)+O(1).$
+Let $k,l\in\mathbb{Z}$, and $X$, $Y$ be real random variables that are not necessarily independent. Then there are integers $d_{1}=d_{1}(k,l)>0$ and $d_{2}=d_{2}(k,l)\geq 0$ such that
+$H_{n}(kX+lY)-\frac{H_{n}(X)+H_{n}(Y)}{2}$ $\leq d_{1}\left(H_{n}(X+Y)-\frac{H_{n}(X)+H_{n}(Y)}{2}\right)+d_{2}\mathcal{I}_{n}(X,Y)+O_{k,l}(1).$
 for all $n\in\mathbb{N}$.
 Proof.
-Applying Lemma 2.26 with the random variable $(X,lY)$, we obtain that
+When $k=l=0$, the choice $d_{1}=1$ and $d_{2}=0$ works. When $k\neq 0$ and $l=0$, the choice $d_{1}=d_{2}=1$ works. When $l\neq 0$, we apply Lemma 2.26 with the random variable $(X,lY)$ to get
 $H_{n}(kX+lY)-\frac{H_{n}(X)+H_{n}(Y)}{2}$
-$\qquad\leq d_{1}(k,1)\left(H_{n}(X+lY)-\frac{H_{n}(X)+H_{n}(Y)}{2}\right)+d_{2}(k,1)\mathcal{I}_{n}(X,Y)+O_{k,l}(1).$
-Applying Lemma 2.26 again, with the random variable $(Y,X)$, we find that
+$\qquad\leq d_{1}(k,1)\left(H_{n}(X+lY)-\frac{H_{n}(X)+H_{n}(Y)}{2}\right)+d_{2}(k,1)\mathcal{I}_{n}(X,Y)+O_{k,l}(1)$
+for some $d_{1}(k,1)>0$ and $d_{2}(k,1)\geq 0$. Applying Lemma 2.26 again, with the random variable $(Y,X)$, we find that
 $H_{n}(lY+X)-\frac{H_{n}(X)+H_{n}(Y)}{2}$
@@
 Proof of Theorem 2.23.
-Since $k,l\in\mathbb{Q}^{\ast}$, we can write $k=k^{\prime}/t$ and $l=l^{\prime}/t$ for some $k^{\prime},l^{\prime},t\in\mathbb{Z}^{\ast}$.
+Since $k,l\in\mathbb{Q}$, we can write $k=k^{\prime}/t$ and $l=l^{\prime}/t$ for some $k^{\prime},l^{\prime}\in\mathbb{Z}$ and $t\in\mathbb{Z}^{\ast}$.
 It satisfies that $H_{n}(kX+lY)=H_{n}(k^{\prime}X+l^{\prime}Y)+O_{t}(1)$.
@@
 $H_{n}(kX+lY)-\frac{H_{n}(X)+H_{n}(Y)}{2}\ll_{k,l}H_{n}(X+Y)-\frac{H_{n}(X)+H_{n}(Y)}{2}+O_{k,l}(1)$
-for any $n\in\mathbb{N}$. Here, the case that $k=0$ or $l=0$ can be proved by using (13).
+for any $n\in\mathbb{N}$.
 However, when $k,l$ are $\mathbb{Q}$-linearly independent, the above formula may fail.
@@
 Lemma 2.29.
-Let $X,Y$ be independent random variables, and $f,g$ be weakly $\kappa$-bi-Lipschitz real-valued functions. Suppose that $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ are i.i.d. copies of $(X,Y)$. Then
-$H_{n}(f(X_{1})+f(X_{2})+g(Y_{1})+g(Y_{2}))$ $\leq 3H_{n}(f(X)+g(Y))-H_{n}(X)-H_{n}(Y)+O(\log\kappa),$
-$H_{n}(f(X_{1})-f(X_{2})+g(Y_{1})-g(Y_{2}))$ $\leq 7H_{n}(f(X)+g(Y))-3H_{n}(X)-3H_{n}(Y)+O(\log\kappa),$
-$H_{n}(f(X_{1})+f(X_{2})-g(Y_{1})-g(Y_{2}))$ $\leq 9H_{n}(f(X)+g(Y))-4H_{n}(X)-4H_{n}(Y)+O(\log\kappa).$
+Let $X,Y$ be independent random variables.
+Suppose that the real-valued functions $f,g$ are weakly $(\kappa,d)$-bi-Lipschitz on $range_{+}(X)$ and $range_{+}(Y)$, respectively.
+Suppose that $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ are i.i.d.
+copies of $(X,Y)$.
+Then
+$H_{n}(f(X_{1})+f(X_{2})+g(Y_{1})+g(Y_{2}))$ $\leq 3H_{n}(f(X)+g(Y))-H_{n}(X)-H_{n}(Y)+O(\log\kappa)+O(\log d),$
+$H_{n}(f(X_{1})-f(X_{2})+g(Y_{1})-g(Y_{2}))$ $\leq 7H_{n}(f(X)+g(Y))-3H_{n}(X)-3H_{n}(Y)+O(\log\kappa)+O(\log d),$
+$H_{n}(f(X_{1})+f(X_{2})-g(Y_{1})-g(Y_{2}))$ $\leq 9H_{n}(f(X)+g(Y))-4H_{n}(X)-4H_{n}(Y)+O(\log\kappa)+O(\log d).$
 Proof.
@@
 $\qquad-H_{n}(f(X_{1})+g(Y_{1}))+O(1)$
-$=H_{n}(f(X)+g(Y))+2H_{n}(f(X)\pm g(Y))-H_{n}(X)-H_{n}(Y)+O(\log\kappa).$
+$=H_{n}(f(X)+g(Y))+2H_{n}(f(X)\pm g(Y))-H_{n}(X)-H_{n}(Y)+O(\log\kappa)+O(\log d).$
 By Remark 2.25, it satisfies that
-$H_{n}(f(X)-g(Y))$ $\leq 3H_{n}(f(X)+g(Y))-H_{n}(f(X))-H_{n}(q(Y))+O(1)$
-$=3H_{n}(f(X)+g(Y))-H_{n}(X)-H_{n}(Y)+O(\log\kappa).$
+$H_{n}(f(X)-g(Y))$ $\leq 3H_{n}(f(X)+g(Y))-H_{n}(f(X))-H_{n}(g(Y))+O(1)$
+$=3H_{n}(f(X)+g(Y))-H_{n}(X)-H_{n}(Y)+O(\log\kappa)+O(\log d).$
 The above formulae still hold if $g$ is replaced by g-g. The lemma then follows from calculation. ∎
@@
 (ii) We say that the random variable $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, if $X$ and $Y$ are $(s_{1},C_{1})$-Frostman and $(s_{2},C_{2})$-Frostman, respectively.
-(iii) We say a random variable $X$ is $(s,C)$-Frostman given $Y$, if
+(iii) We say a random variable $X$ is conditionally $(s,C)$-Frostman given $Y$, if
 $P(X\in B(x,r),\,Y\in E)\leq Cr^{s}P(Y\in E)$
 for any ball $B(x,r)$ and Borel set $E$.
-Moreover, We say that the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, if $X$ is $(s_{1},C_{1})$-Frostman given $Y$, and $Y$ is $(s_{2},C_{2})$-Frostman given $X$.
-The following properties follows directly from the definitions, whose proofs are left as an exercise.
+Moreover, We say that the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, if $X$ is conditionally $(s_{1},C_{1})$-Frostman given $Y$, and $Y$ is conditionally $(s_{2},C_{2})$-Frostman given $X$.
+The following properties follow directly from the definitions, whose proofs are left as an exercise.
 Lemma 2.30.
@@
 (iii) Suppose that $(X,Y)$ is jointly $(s_{1},s_{2},C)$-Frostman. Then it is also $(s_{1}+s_{2},C)$-Frostman.
-(iv) Suppose that $(X,Y)$ is a jointly $(s_{1},s_{2},C)$-Frostman random variable whose range is contained in $B(0,R)$, where $R>0$.
-Then $(X,Y)$ is marginally $(s_{1},CR^{s_{2}};s_{2},CR^{s_{1}})$-Frostman.
+(iv) Suppose that $(X,Y)$ is a jointly $(s_{1},s_{2},C)$-Frostman random variable whose range is contained in $B(0,R)$. Then $(X,Y)$ is marginally $(s_{1},CR^{s_{2}};s_{2},CR^{s_{1}})$-Frostman.
 The next lemma presents the Frostman conditions related to the arithmetic of random variables.
@@
 $P(aX+bY\in B(z,r))=\sum\limits_{i}P(aX+bY\in B(z,r),\,X\in I_{i}):=\sum\limits_{i}q_{i}.$
-Assume that $b\neq 0$. If $q_{i}>0$, then $Y$ is located in an interval $J_{i}$ of length at most $\frac{2(|a|+1)r}{|b|}$. So
+Assume that $b\neq 0$. If $q_{i}>0$, then $Y$ is located in an interval $J_{i}$ of length at most $\frac{(|a|+2)r}{|b|}$. Hence,
 $q_{i}\leq P(X\in I_{i},\,Y\in J_{i})\leq C_{2}\Big(\frac{(|a|+2)r}{|b|}\Big)^{s_{2}}P(X\in I_{i}).$
@@
 By symmetry, it is also $(s_{1},C^{\prime}_{1})$-Frostman with $\log C_{1}^{\prime}=\log C_{1}+O_{a,b}(1)$, provided that $a\neq 0$.
-(ii) Write $p:=P\big{(}aX+bY\in B(z_{1},r_{1}),\,cX+dY\in B(z_{2},r_{2})\big{)}$ for an arbitrary pair of balls $B(z_{1},r_{1})$ and $B(z_{2},r_{2})$.
-Calculation shows that, if $p>0$, then we must have
+(ii) Write $p:=P\big(aX+bY\in B(z_{1},r_{1}),\,cX+dY\in B(z_{2},r_{2})\big)$ for an arbitrary pair of balls $B(z_{1},r_{1})$ and $B(z_{2},r_{2})$. Calculation shows that, if $p>0$, then we must have
 $X\in B\Big(\frac{dz_{1}-bz_{2}}{ad-bc},\,\frac{|d|r_{1}+|b|r_{2}}{|ad-bc|}\Big),\qquad Y\in B\Big(\frac{-cz_{1}+az_{2}}{ad-bc},\,\frac{|c|r_{1}+|a|r_{2}}{|ad-bc|}\Big).$ (17)
-First, let us consider the case $ad\neq 0$. When $r_{1}\leq r_{2}$, let $\{J_{i}\}$ be a partition of $\mathbb{R}$} composed of intervals of length $r_{1}$. We estimate $p$ by
+First, consider the case $ad\neq 0$. When $r_{1}\leq r_{2}$, let $\{J_{i}\}$ be a partition of $\mathbb{R}$} composed of intervals of length $r_{1}$. We estimate $p$ by
 $p$ $=\sum\limits_{i}P\big(aX+bY\in B(z_{1},r_{1}),\,cX+dY\in B(z_{2},r_{2}),\,Y\in J_{i}\big)$
 $\leq\sum\limits_{i}P\big(aX+bY\in B(z_{1},r_{1}),\,Y\in J_{i}\big):=\sum\limits_{i}p_{i}.$
-One can check that, if $p_{i}>0$, then $X$ is located in an interval $I_{i}$ of length at most $\frac{2(|b|+1)r_{1}}{|a|}$. Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, we have
+One can check that, if $p_{i}>0$, then $X$ is located in an interval $I_{i}$ of length at most $\frac{(|b|+2)r_{1}}{|a|}$. Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, we have
 $p_{i}\leq P(X\in I_{i},Y\in J_{i})\leq C_{1}\Big(\frac{(|b|+2)r_{1}}{|a|}\Big)^{s_{1}}P(Y\in J_{i}).$
@@
 $p\leq\sum\limits_{i}P\big(cX+dY\in B(z_{2},r_{2}),\,X\in I_{i}^{\prime}\big):=\sum\limits_{i}p_{i}^{\prime}.$
-We apply similar arguments as above. For non-zero terms, the value of $Y$ locates in an interval $J_{i}^{\prime}$ of length at most $\frac{2(|c|+1)r_{2}}{|d|}$. So
+We apply similar arguments as above. For non-zero terms, the value of $Y$ locates in an interval $J_{i}^{\prime}$ of length at most $\frac{(|c|+2)r_{2}}{|d|}$. Therefore,
 $p_{i}^{\prime}\leq P(X_{i}\in I_{i}^{\prime},Y\in J_{i}^{\prime})\leq C_{1}\Big(\frac{(|c|+2)r_{2}}{|d|}\Big)^{s_{2}}P(X\in I_{i}^{\prime}),$
@@
 Proof.
-(i) is taken from [12, Lemma 4.8].
+(i) The conclusion comes from [12, Lemma 4.8].
 (ii) Since $X$ is conditionally $(s,C)$-Frostman given $Y$, then
@@
 Let $m\geq 1$.
-Suppose that $Z$ is a random variable with bounded range in ${\mathbb{R}}^{m}$, and $f$ is a function such that $domain(f)\supseteq range_{+}(Z)$.
-Assume that there are numbers $s,C>0$ such that $P(|f(Z)|<r)\leq Cr^{s}$ for any $r>0$.
+Suppose that $Z$ is a random variable with bounded range in ${\mathbb{R}}^{m}$.
+Assume that there are numbers $s,C>0$ such that $P(|Z|<r)\leq Cr^{s}$ for any $r>0$.
 Let $0<\eta\leq\min\{2^{-1}(2C)^{-1/s},\,1/2\}$, and denote by $E$ the event that $|Z|\geq\eta$.
 Assume that $P(E)\gg 1$.
-Then the random variable $f(Z)_{E}$ satisfies
-$H_{n}(f(Z)_{E})=\frac{H_{n}(f(Z))}{P(E)}+O(C(2\eta)^{s}n)+O_{m}(1).$
+Then the random variable $Z_{E}$ satisfies
+$H_{n}(Z_{E})=\frac{H_{n}(Z)}{P(E)}+O(C(2\eta)^{s}n)+O_{m}(1).$
 for any all $n\in\mathbb{N}$.
 Proof.
-Recall that $f(Z)_{E}$ is the random variable such that $f(Z)_{E}\equiv(f(Z)|E)$, and $P_{E}$ is the conditional probability measure given $E$ (see Lemma A.2).
-Assume that $k_{0}$ is a natural number with $(k_{0}-1)2^{-n}\leq\eta<k_{0}2^{-n}$. Combining the conditions in this lemma, one deduces that
-$H_{n}(f(Z)_{E})$ $=-\sum\limits_{k\in\mathbb{Z}^{m}\atop|k|\geq k_{0}+m}P_{E}(D_{n}(f(Z)_{E})=k2^{-n})\log P_{E}(D_{n}(f(Z)_{E})=k2^{-n})$
-$\qquad+O\left(-\sum\limits_{k_{0}-m\leq|k|<k_{0}+m}P_{E}(D_{n}(f(Z)_{E})=k2^{-n})\log P_{E}(D_{n}(f(Z)_{E})=k2^{-n})\right)$
-$=-\sum\limits_{|k|\geq k_{0}+m}\frac{P(D_{n}(f(Z))=k2^{-n})}{P(E)}\big{(}\log P(D_{n}(f(Z))=k2^{-n})-\log P(E)\big{)}+O_{m}(1)$
-$=-\frac{1}{P(E)}\sum\limits_{|k|\geq k_{0}+m}P(D_{n}(f(Z))=k2^{-n})\log P(D_{n}(f(Z))=k2^{-n})+O_{m}(1).$
-First, let us consider the case that $\eta>m2^{-n+2}$, which gives $(k_{0}+2m)2^{-n}\leq\eta+(2m+1)2^{-n}<2\eta$. Applying Lemma 2.9 with the fact
-$\sum\nolimits_{|k|<k_{0}+m}P(D_{n}(f(Z))=k2^{-n})\leq P(|f(Z)|<(k_{0}+2m)2^{-n})\leq P(|f(Z)|<2\eta)\leq C(2\eta)^{s}\leq 1/2,$
+Recall that $Z_{E}$ is the random variable such that $Z_{E}\equiv(Z|E)$, and $P_{E}$ is the conditional probability measure given $E$ (see Lemma A.2).
+Assume that $k_{0}$ is the natural number with $(k_{0}-1)2^{-n}\leq\eta<k_{0}2^{-n}$. Combining the conditions in this lemma, one deduces that
+$H_{n}(Z_{E})$ $=-\sum\limits_{k\in\mathbb{Z}^{m}\atop|k|\geq k_{0}+m}P_{E}(D_{n}(Z_{E})=k2^{-n})\log P_{E}(D_{n}(Z_{E})=k2^{-n})$
+$\qquad+O\left(-\sum\limits_{k_{0}-m\leq|k|<k_{0}+m}P_{E}(D_{n}(Z_{E})=k2^{-n})\log P_{E}(D_{n}(Z_{E})=k2^{-n})\right)$
+$=-\sum\limits_{|k|\geq k_{0}+m}\frac{P(D_{n}(Z)=k2^{-n})}{P(E)}\big(\log P(D_{n}(Z)=k2^{-n})-\log P(E)\big)+O_{m}(1)$
+$=-\frac{1}{P(E)}\sum\limits_{|k|\geq k_{0}+m}P(D_{n}(Z)=k2^{-n})\log P(D_{n}(Z)=k2^{-n})+O_{m}(1).$
+First, we analyze the case that $\eta>m2^{-n+2}$, which gives $(k_{0}+2m)2^{-n}\leq\eta+(2m+1)2^{-n}<2\eta$. Applying Lemma 2.9 with the fact
+$\sum\nolimits_{|k|<k_{0}+m}P(D_{n}(Z)=k2^{-n})\leq P(|Z|<(k_{0}+2m)2^{-n})\leq P(|Z|<2\eta)\leq C(2\eta)^{s}\leq 1/2,$
 we see that
-$-\sum\limits_{|k|<k_{0}+m}P(D_{n}(f(Z))=k2^{-n})\log P(D_{n}(f(Z))=k2^{-n})$
+$-\sum\limits_{|k|<k_{0}+m}P(D_{n}(Z)=k2^{-n})\log P(D_{n}(Z)=k2^{-n})$
 $\quad\quad\quad\leq C(2\eta)^{s}\cdot\log\big(2\eta 2^{n}+2m+3\big)+O(1)\leq C(2\eta)^{s}n+O_{m}(1).$
 Second, we consider the case that $\eta\leq m2^{-n+2}$. Then $k_{0}\leq 4m+1$ and
-$\sum\nolimits_{|k|<k_{0}+m}P(D_{n}(f(Z))=k2^{-n})\log P(D_{n}(f(Z))=k2^{-n})=O_{m}(1).$
+$\sum\nolimits_{|k|<k_{0}+m}P(D_{n}(Z)=k2^{-n})\log P(D_{n}(Z)=k2^{-n})=O_{m}(1).$
 Finally, we conclude that
-$H(f(Z)_{E})$ $=\frac{1}{P(E)}\left(H_{n}(f(Z))+\sum\limits_{|k|<k_{0}+m}P(D_{n}(f(Z))=k2^{-n})\log P(D_{n}(f(Z))=k2^{-n})\right)+O_{m}(1)$
-$=\frac{H_{n}(f(Z))}{P(E)}+O(C(2\eta)^{s}n)+O_{m}(1).$
+$H(Z_{E})$ $=\frac{1}{P(E)}\left(H_{n}(Z)+\sum\limits_{|k|<k_{0}+m}P(D_{n}(Z)=k2^{-n})\log P(D_{n}(Z)=k2^{-n})\right)+O_{m}(1)$
+$=\frac{H_{n}(Z)}{P(E)}+O(C(2\eta)^{s}n)+O_{m}(1).$
 In the following lemma, we show that the probability that the distance function takes small values is related to the Frostman condition.
@@
 Lemma 2.36 ([13]).
-Let $\mu$ be a probability measure with $\operatorname{spt}(\mu)\subset B(0,R)\subset\mathbb{R}^{2}$, for some $R>0$.
+Let $\mu$ be a probability measure with $\operatorname{spt}(\mu)\subseteq B(0,R)\subseteq\mathbb{R}^{2}$, for some $R>0$.
 Then, for any $0<t\leq 1$ and for real numbers $0<\eta<r<2R$, we have $\Delta(\mu)([r,r+\eta])\ll_{t}r^{1/2}\eta^{t}I_{t+1/2}(\mu)$.
 Lemma 2.37.
-Let $\mu$ be a $(2s,C)$-Frostman probability measure with $\operatorname{spt}(\mu)\subset B(0,R)\subset\mathbb{R}^{2}$, where $\frac{1}{4}<s\leq 1$ and $C,R>0$.
+Let $\mu$ be a $(2s,C)$-Frostman probability measure with $\operatorname{spt}(\mu)\subseteq B(0,R)\subseteq\mathbb{R}^{2}$, where $\frac{1}{4}<s\leq 1$ and $C,R>0$.
 Then for any $r,\eta,t$ with $0<\eta<r<2R$, $0<t\leq 1$ and $t<2s-\frac{1}{2}$, it satisfies that
@@
 $\qquad\leq(t+1/2)\left(\int_{0}^{2\tilde{R}}C\ell^{2s-t-3/2}d\ell+\int_{2\tilde{R}}^{\infty}\ell^{-t-3/2}d\ell\right)$
-$\qquad=O_{t}\left(\frac{C\tilde{R}^{2s-t-1/2}}{2s-t-1/2}\right)+O_{t}(1):=M.$
+$\qquad=O_{t}\left(\frac{C\tilde{R}^{2s-t-1/2}}{2s-t-1/2}\right)+O(\tilde{R}^{-t-1/2}):=M.$
 Thus
@@
 When $0<\eta<r<2R$ and $0<t\leq 1$, one obtains by Lemma 2.36 that
-$\Delta(\mu)([r,r+\eta])\ll_{t}r^{1/2}\eta^{t}I_{t+1/2}(\mu)=O_{t}\left(\frac{C\tilde{R}^{2s-t}\eta^{t}}{2s-t-1/2}\right)+O_{t}(\eta^{t}).$
+$\Delta(\mu)([r,r+\eta])\ll_{t}r^{1/2}\eta^{t}I_{t+1/2}(\mu)=O_{t}\left(\frac{C\tilde{R}^{2s-t}\eta^{t}}{2s-t-1/2}\right)+O(\eta^{t}).$
 The proof is completed. ∎
@@
 Theorem 2.38.
-Let $W$ be a $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $1/2<s<1$ and $C,R>0$.
+Let $W$ be a $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subseteq\mathbb{R}^{2}$, where $1/2<s<1$ and $C,R>0$.
 Let $W^{\prime}$ be a random variable which is i.i.d with $W$.
@@
 Theorem 2.41.
-Let $W$ be a $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subset\mathbb{R}^{2}$, where $1/2<s<1$ and $C,R>0$.
+Let $W$ be a $(2s,C)$-Frostman random variable with $range(W)\subseteq B(0,R)\subseteq\mathbb{R}^{2}$, where $1/2<s<1$ and $C,R>0$.
 Let $W^{\prime}$ be a random variable which is i.i.d with $W$.
@@
 $H_{n}(Z)\geq nt+O(\log_{+}C)+O(\log_{+}R)+O(|\log(2s-t-1/2)|)+O_{t}(1).$
-By Lemma 2.34, one has $P(|Z|<r)\leq Cr^{2s}$ for any $r>0$.
 Let $\eta$ be a parameter with $0<\eta<\min\{2^{-1}(2C)^{-1/2s},\,1/2\}$ to be determined later.
 Denote by $E$ the event that $|Z|\geq\eta$.
+By Lemma 2.34, one has $P(|Z|<\eta)\leq C\eta^{2s}$, which gives $P(E)\geq 1/2$.
 By Lemma 2.33, we deduce that
@@
 Applying Lemma A.2, one deduces that
-$H_{n}(Z^{2})$ $\geq P(E)\,H_{n}(Z_{E}^{2})\geq P(E)\left(H_{n}(Z_{E})+O\left(\log\frac{1}{\eta}\right)+O(1)\right)$
+$H_{n}(Z^{2})$ $\geq P(E)\,H_{n}(Z_{E}^{2})\geq P(E)\left(H_{n}(Z_{E})+O\left(\log\frac{1}{\eta}\right)\right)$
 $\geq P(E)\left(\frac{nt}{P(E)}+O(\log_{+}C)+O(\log_{+}R)\right$.
@@
 […]
-One can verify that the requirements involving $t$ and $\eta$ are satisfied, and $2s-t-1/2\gg_{s}1$. Also, $\log(1/\eta)<\log C+O_{s}(1)$. As a result, we conclude that
+Verification shows that the requirements involving $t$ and $\eta$ are satisfied, and $2s-t-1/2\gg_{s}1$. Also, $\log(1/\eta)<\log C+O_{s}(1)$. As a result, we conclude that
 $H_{n}(Z^{2})\geq n(s+\epsilon_{0})+O(\log_{+}C)+O(\log_{+}R)+O_{s}(1),$
@@
 Lemma 2.42.
-Let $X$, $X^{\prime}$, $Y$, $Y^{\prime}$ be independent real random variables.
+Let $X$, $X^{\prime}$, $Y$, $Y^{\prime}$ be four independent real random variables.
 Suppose that $X$ and $X^{\prime}$ are both $(s,C)$-Frostman with ranges contained in [c,c][-c,c], where $0<s\leq 1$ and $C,c>0$.
@@
 $H_{n}\big((X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}\big)\geq H_{n}\big((X-X^{\prime})^{2}\big)+O(1).$
-By Lemma 2.31, the random variable $Z:=X-X^{\prime}$ is $(s,O(C))$-Frostman.
+By Lemmas 2.30 and 2.31, the random variable $Z:=X-X^{\prime}$ is $(s,O(C))$-Frostman.
 By Lemma 2.32, one has $H_{n}(Z)\geq ns-\log C+O(1)$.
@@
 We do not know whether such a generalization has been studied in the literature, so we provide a complete proof in this section.
-The Balog-Szemerédi-Gowers theorem (or one of its many variants) says that if one has a finite subset of the integers whose sums line up a lot, then it must not have many sums.
-More precisely, given $A\subseteq\mathbb{Z},$ whose elements $a_{j}\in A$ have $\gamma|A|^{3}$ relations of the form $a_{1}+a_{2}=a_{3}+a_{4},$ there must be a large subset $A^{\prime}\subseteq A,$ where $|A^{\prime}|\geq C_{1}(\gamma)|A|$ that satisfies $|A^{\prime}+A^{\prime}|\leq C_{2}(\gamma)|A^{\prime}|,$ where the constants $C_{j}$ above are positive and only depend on $\gamma.$
 Theorem 3.1.
-Let $n$ be a natural number, and $(X,Y)$ be a given random variable. Suppose that
+Let $(X,Y)$ be a random variable, and $n$ be a natural number. Suppose that
 $H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\leq\log K_{1}.$ (20)
@@
 $H_{n}(Y^{\prime\prime}|X^{\prime},Y^{\prime})\geq H_{n}(Y)-\log K_{1}.$ (22)
-In addition, for any weakly $\kappa$-bi-Lipschitz functions $f$ and $g$ with
+In addition, for any real valued functions $f$ and $g$ that are weakly $(\kappa,d)$-bi-Lipschitz on $range_{+}(X)$ and $range_{+}(Y)$, respectively, with
 $H_{n}(f(X)+g(Y))-\frac{H_{n}(X)+H_{n}(Y)}{2}\leq\log K_{2},$ (23)
 we have that
-$H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|X^{\prime},Y^{\prime})-\frac{H_{n}(X)+H_{n}(Y)}{2}\leq 4\log K_{1}+3\log K_{2}+O(\log\kappa).$ (28)
+$H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|X^{\prime},Y^{\prime})-\frac{H_{n}(X)+H_{n}(Y)}{2}\leq 4\log K_{1}+3\log K_{2}+O(\log\kappa)+O(\log d).$ (24)
+Here $\log K_{1},\,\log K_{2},\,\log\kappa\,\gg 1$ and $d\in\mathbb{N}$.
 Remark 3.2.
@@
 Lemma 3.3.
-Let $n$ be a natural number, and $(X,Y)$ be a given random variable.
-Then there exists a random variable $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$ such that (i)–(iv) in Theorem 3.1 hold, with $X^{\prime\prime}$ and $Y^{\prime\prime}$ being conditionally independent given $D_{n}(X^{\prime},Y^{\prime})$.
+Let $(X,Y)$ be a random variable.
+Then for any $n\in\mathbb{N}$, there exists a random variable $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$ such that (i)–(iv) in Theorem 3.1 hold, with $X^{\prime\prime}$ and $Y^{\prime\prime}$ being conditionally independent given $D_{n}(X^{\prime},Y^{\prime})$.
 Proof.
@@
 Corollary 3.4.
-Let $n$ be a natural number and $(X,Y)$ be a given random variable.
+Let $(X,Y)$ be a random variable, and $n$ be a natural number.
 Let $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$ be the random variable shown in Lemma 3.3.
-For any $\kappa$-Lipschitz function $h$, the following properties hold.
+For any function $h$ that is $\kappa$-Lipschitz on $range_{+}((X,Y))$, the following properties hold.
 (i) $H_{n}(h(X^{\prime},Y^{\prime}))=H_{n}(h(X,Y))$.
@@
 $H_{n}(h(X^{\prime},Y^{\prime\prime}))=H_{n}(h(D_{n}(X^{\prime}),D_{n}(Y^{\prime\prime})))+O(\log\kappa).$
-Noting that $(Y^{\prime},D_{n}(X^{\prime}))\equiv(Y^{\prime\prime},D_{n}(X^{\prime}))$, one asserts that
+Noting that $(Y^{\prime},D_{n}(X^{\prime}))\equiv(Y^{\prime\prime},D_{n}(X^{\prime}))$, so
 $H_{n}(h(D_{n}(X^{\prime}),D_{n}(Y^{\prime})))=H_{n}(h(D_{n}(X^{\prime}),D_{n}(Y^{\prime\prime}))).$
-Combining this with (i), one gets (ii). Claim (iii) follows by a similar argument. ∎
+Combining these with (i), one gets (ii). Claim (iii) follows by a similar argument. ∎
 Lemma 3.5.
-With $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$ and $f,g$ as stated in Theorem 3.1, we always have
-$H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa).$
+Under the notations and conditions stated in Theorem 3.1, we have
+$H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa)+O(\log d).$
 Proof.
 Let $Z$ be a random variable such that $Z,Y^{\prime}$ are conditionally i.i.d. given $D_{n}(X^{\prime},X^{\prime\prime})$.
-Recall that $f,g$ are both weakly $\kappa$-bi-Lipschitz. Applying Lemma 2.15 and Lemma 2.18, one can verify that
+Recall that $f,g$ are both weakly $(\kappa,d)$-bi-Lipschitz. Applying Lemma 2.16 and Lemma 2.18, one can verify that
 $H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})+H_{n}(X^{\prime},X^{\prime\prime},Z,Y^{\prime})$
-$\qquad=H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})+H_{n}(X^{\prime},X^{\prime\prime},g(Z),Y^{\prime})+O(\log\kappa)$
-$\qquad\leq H_{n}(f(X^{\prime})+g(Z),f(X^{\prime\prime})+g(Z),Y^{\prime})+H_{n}(X^{\prime},X^{\prime\prime},Y^{\prime})+O(\log\kappa).$ (29)
+$\qquad=H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})+H_{n}(X^{\prime},X^{\prime\prime},g(Z),Y^{\prime})+O(\log\kappa)+O(\log d)$
+$\qquad\leq H_{n}(f(X^{\prime})+g(Z),f(X^{\prime\prime})+g(Z),Y^{\prime})+H_{n}(X^{\prime},X^{\prime\prime},Y^{\prime})+O(\log\kappa)+O(\log d).$ (25)
 Indeed, we have $(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})=f_{1}(f(X^{\prime})+g(Z),f(X^{\prime\prime})+g(Z),Y^{\prime})=f_{2}(X^{\prime},X^{\prime\prime},Y^{\prime})$ and and
@@
 given $D_{n}(Y^{\prime})$.
-By Lemmas 2.8 and 2.19,
+By Lemma 2.8 and Corollary 3.4,
 $H_{n}(X^{\prime},X^{\prime\prime},Z,Y^{\prime})=2H_{n}(X^{\prime},X^{\prime\prime},Y^{\prime})-H_{n}(X^{\prime},X^{\prime\prime})$
@@
 Putting formulae (20), (23), and (25)–(29) together, we get that
-$H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})\leq 2H_{n}(f(X)+g(Y))+2H_{n}(X)+2H_{n}(Y)-2H_{n}(X,Y)+O(\log\kappa).$
+$H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})$
+$\qquad\leq 2H_{n}(f(X)+g(Y))+2H_{n}(X)+2H_{n}(Y)-2H_{n}(X,Y)+O(\log\kappa)+O(\log d).$
 $\qquad\leq\big(2H_{n}(f(X)+g(Y))-H_{n}(X)-H_{n}(Y)\big)$
-$\qquad\qquad+2\big{(}H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\big{)}+H_{n}(X)+H_{n}(Y)+O(\log\kappa).$
-$\qquad\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa).$
-We are ready to present a proof of Theorem 3.1
+$\qquad\qquad+2\big(H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\big)+H_{n}(X)+H_{n}(Y)+O(\log\kappa)+O(\log d).$
+$\qquad\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa)+O(\log d).$
+We now present the proof of Theorem 3.1.
 Proof of Theorem 3.1.
@@
 $H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime}),Y^{\prime})+H_{n}(X^{\prime},Y^{\prime},X^{\prime\prime},Y^{\prime\prime})$
-$\qquad=H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime}),Y^{\prime})+H_{n}(f(X^{\prime}),Y^{\prime},X^{\prime\prime},Y^{\prime\prime})+O(\log\kappa)$
+$\qquad=H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime}),Y^{\prime})+H_{n}(f(X^{\prime}),Y^{\prime},X^{\prime\prime},Y^{\prime\prime})+O(\log\kappa)+O(\log d)$
 $\qquad\leq H_{n}(f(X^{\prime\prime})-f(X^{\prime}),f(X^{\prime})+g(Y^{\prime\prime}),Y^{\prime})+H_{n}(X^{\prime\prime},Y^{\prime\prime},Y^{\prime})+O(\log\kappa).$ (31)
@@
 $\qquad\leq H_{n}(f(X^{\prime\prime})-f(X^{\prime}),Y^{\prime})+H_{n}(f(X^{\prime})+g(Y^{\prime\prime}))$
-$\qquad\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa)+\frac{H_{n}(X)+H_{n}(Y)}{2}+\log K_{2}$
-$\qquad=\frac{3}{2}(H_{n}(X)+H_{n}(Y))+2\log K_{1}+3\log K_{2}+O(\log\kappa).$ (38)
+$\qquad\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa)+O(\log d)$ (34)
+$\qquad\qquad\qquad+\frac{H_{n}(X)+H_{n}(Y)}{2}+\log K_{2}$
+$\qquad=\frac{3}{2}(H_{n}(X)+H_{n}(Y))+2\log K_{1}+3\log K_{2}+O(\log\kappa)+O(\log d).$ (35)
 Combining (30)–(35), one concludes that
-$H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|X^{\prime},Y^{\prime})\leq\frac{H_{n}(X)+H_{n}(Y)}{2}+4\log K_{1}+3\log K_{2}+O(\log\kappa).$
+$H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|X^{\prime},Y^{\prime})\leq\frac{H_{n}(X)+H_{n}(Y)}{2}+4\log K_{1}+3\log K_{2}+O(\log\kappa)+O(\log d).$
 Thus, the estimate (24) follows. ∎
 4 Proof of Theorems 1.21.4 with $s>1/2$
-4.1 Diagonal forms with independent random variables
-In this section, we consider the diagonal real quadratic form $\phi(x,y)=\alpha x^{2}+\beta y^{2}$.
-When we write $O_{\phi}(\cdot)$, it means that the implied constant may depend on the coefficients $\alpha$ and $\beta$ that occurred in $\phi$.
+4.1 Diagonal quadratics with independent random variables
+In this subsection, we first consider the diagonal real quadratic form $\phi(x,y)=\alpha x^{2}+\beta y^{2}$.
 Lemma 4.1.
@@
 Then
-$\begin{split}&\max\{H_{n}(X+Y),~H_{n}(\alpha X^{2}+\beta Y^{2})\}\geq n(s+\epsilon_{1})\\ &\qquad\qquad\qquad+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi,s}(1).\end{split}$ (39)
-for all $n\in\mathbb{N}$.
+$H_{n}(X+Y)+H_{n}(\alpha X^{2}+\beta Y^{2})\geq n(2s+\epsilon_{1})+O(\log\mathfrak{K})$
+for all $n\in\mathbb{N}$, where
+$\log\mathfrak{K}=O\big(\log_{+}C_{1}C_{2}+\log_{+}c_{1}+\log_{+}(1/c_{2})+\big|\log|\alpha|\big|+\big|\log|\beta|\big|\big)+O_{s}(1)$
 Proof.
-Write $\phi(x,y)=p(x)+q(y)$, where $p(x)=\alpha x^{2}$ and $q(y)=\beta y^{2}$ are functions whose domains are both taken to be $c_{2},-c_{1}]\cup[c_{2},c_{1}]$. One asserts by Example 2.16 that
-$\begin{split}H_{n}(p(X))&=H_{n}(X)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1),\\ H_{n}(q(Y))&=H_{n}(Y)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1).\end{split}$ (40)
-Define $W=(\sqrt{|\alpha|}X,\sqrt{|\beta|}Y),$ with $|\alpha|X$ and $|\beta|Y$ being $(s,O_{\phi}(C_{1}))$-Frostman and $(s,O_{\phi}(C_{2}))$-Frostman, respectively.
+Write $\phi(x,y)=\rho_{1}(x)+\rho_{2}(y)$, where $\rho_{1}(x)=\alpha x^{2}$ and $\rho_{2}(y)=\beta y^{2}$. One asserts by Example 2.12 and Lemma 2.16 that
+$\begin{split}H_{n}(\rho_{1}(X))&=H_{n}(X)+O\big(\log_{+}c_{1}+\log_{+}(1/c_{2})+\big|\log|\alpha|\big|+1\big)=O(\log\mathfrak{K}),\\ H_{n}(\rho_{2}(Y))&=H_{n}(Y)+O\big(\log_{+}c_{1}+\log_{+}(1/c_{2})+\big|\log|\beta|\big|+1\big)=O(\log\mathfrak{K}).\end{split}$ (36)
+Define $W=(\sqrt{|\alpha|}X,\sqrt{|\beta|}Y)$.
 Let $W_{1}=(\sqrt{|\alpha|}X_{1},\sqrt{|\beta|}Y_{1})$, $W_{2}=(\sqrt{|\alpha|}X_{2},\sqrt{|\beta|}Y_{2})$ be i.i.d.
@@
 $\alpha(X_{1}-X_{2})^{2}+\beta(Y_{1}-Y_{2})^{2}=2(\alpha(X_{1}^{2}+X_{2}^{2})+\beta(Y_{1}^{2}+Y_{2}^{2}))-\alpha(X_{1}+X_{2})^{2}-\beta(Y_{1}+Y_{2})^{2}.$
-So $|W_{1}-W_{2}|^{2}=f_{1}(X_{1}+X_{2},Y_{1}+Y_{2},p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}))$, where
+We have $|W_{1}-W_{2}|^{2}=f_{1}(X_{1}+X_{2},Y_{1}+Y_{2},\rho_{1}(X_{1})+\rho_{1}(X_{2})+\rho_{2}(Y_{1})+\rho_{2}(Y_{2}))$, where
 $f_{1}(z_{1},z_{2},z_{3})=2z_{3}-\alpha z_{1}^{2}-\beta z_{2}^{2}.$
-From Example 2.11 and the fact that $range(X_{1}+X_{2}),range(Y_{1}+Y_{2})\subseteq[-2c_{1},2c_{1}]$, one sees that $f_{1}$ is $O_{\phi}(c_{1})$-Lipschitz.
-We also have that $|W_{1}-W_{2}|^{2}=f_{2}(X_{1}-X_{2},Y_{1}-Y_{2})$, with the $O_{\phi}(C_{1})$-Lipschitz function
+From Example 2.11 and the fact that $range(X_{1}+X_{2}),range(Y_{1}+Y_{2})\subseteq[-2c_{1},2c_{1}]$, one sees that $f_{1}$ is $O(|\alpha|c_{1}+|\beta|c_{1}+1)$-Lipschitz.
+We also have that $|W_{1}-W_{2}|^{2}=f_{2}(X_{1}-X_{2},Y_{1}-Y_{2})$, with the $O(|\alpha|c_{1}+|\beta|c_{1})$-Lipschitz function
 $f_{2}(z_{1},z_{2})=\alpha z_{1}^{2}+\beta z_{2}^{2}.$
 Moreover,
-$(X_{1},X_{2},Y_{1},Y_{2})=f_{3}(X_{1}+X_{2},Y_{1}+Y_{2},p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}),X_{1}-X_{2},Y_{1}-Y_{2}),$
+$(X_{1},X_{2},Y_{1},Y_{2})=f_{3}(X_{1}+X_{2},Y_{1}+Y_{2},\,\rho_{1}(X_{1})+\rho_{1}(X_{2})+\rho_{2}(Y_{1})+\rho_{2}(Y_{2}),\,X_{1}-X_{2},Y_{1}-Y_{2}),$
 with the $O(1)$-Lipschitz function
@@
 $H_{n}(|W_{1}-W_{2}|^{2})+H_{n}(X_{1},X_{2},Y_{1},Y_{2})$
-$\leq H_{n}(X_{1}+X_{2},Y_{1}+Y_{2},p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}))$
-$\qquad+H_{n}(X_{1}-X_{2},Y_{1}-Y_{2})+O(\log_{+}c_{1})+O_{\phi}(1)$
-$\leq H_{n}(X_{1}+X_{2})+H_{n}(Y_{1}+Y_{2})+H_{n}(p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}))$
-$\qquad+H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})+O(\log_{+}c_{1})+O_{\phi}(1).$ (41)
-We now estimate the terms in the last expression. For the first term, apply Lemma 2.21 to get that
-$H_{n}(X_{1}+X_{2})$ $\leq H_{n}(X_{1}+Y_{1})+H_{n}(X_{2}+Y_{1})-H_{n}(Y_{1})+O(1),$
-$=2H_{n}(X+Y)-H_{n}(Y)+O(1).$
-One can argue similarly for $H_{n}(Y_{1}+Y_{2})$, hence,
-$H_{n}(X_{1}+X_{2})+H_{n}(Y_{1}+Y_{2})\leq 4H_{n}(X+Y)-H_{n}(X)-H_{n}(Y)+O(1).$ (42)
-For the third term, Lemma 2.29 shows that
-$H_{n}(p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}))\leq 3H_{n}(p(X)+q(Y))$
-$\qquad\qquad\qquad-H_{n}(X)-H_{n}(Y)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1).$ (43)
-For the fourth term, applying Lemma 2.21 shows that
-$H_{n}(X_{1}-X_{2})\leq H_{n}(X_{1}+Y_{1})+H_{n}(X_{2}+Y_{1})-H_{n}(Y_{1})+O(1).$
-So similarly
-$H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})\leq 4H_{n}(X+Y)-H_{n}(X)-H_{n}(Y)+O(1).$ (44)
+$\leq H_{n}(X_{1}+X_{2},\,Y_{1}+Y_{2},\,\rho_{1}(X_{1})+\rho_{1}(X_{2})+\rho_{2}(Y_{1})+\rho_{2}(Y_{2}))$
+$\qquad+H_{n}(X_{1}-X_{2},Y_{1}-Y_{2})+O(\log\mathfrak{K})$
+$\leq H_{n}(X_{1}+X_{2})+H_{n}(Y_{1}+Y_{2})+H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})$
+$\qquad+H_{n}(\rho_{1}(X_{1})+\rho_{1}(X_{2})+\rho_{2}(Y_{1})+\rho_{2}(Y_{2}))+O(\log\mathfrak{K}),$ (37)
+We now estimate the terms in the last expression. Lemma 2.21 shows that
+$H_{n}(X_{1}\pm X_{2})$ $\leq H_{n}(X_{1}+Y_{1})+H_{n}(X_{2}+Y_{1})-H_{n}(Y_{1})+O(1),$
+$=2H_{n}(X+Y)-H_{n}(Y)+O(1),$ (38)
+and similarly
+$H_{n}(Y_{1}\pm Y_{2})\leq 2H_{n}(X+Y)-H_{n}(X)+O(1).$ (39)
+And Lemma 2.29 tells that
+$H_{n}(\rho_{1}(X_{1})+\rho_{1}(X_{2})+\rho_{2}(Y_{1})+\rho_{2}(Y_{2}))$
+$\qquad\leq 3H_{n}(\rho_{1}(X)+\rho_{2}(Y))-H_{n}(X)-H_{n}(Y)+O(\log\mathfrak{K}).$ (40)
 Combining (37)–(40) and Lemma 2.4(iv), we have that
-$H_{n}\big{(}|W_{1}-W_{2}|^{2}\big{)}+$ $5H_{n}(X)+5H_{n}(Y)\leq 8H_{n}(X+Y)+3H_{n}(\phi(X,Y))$
-$\quad+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1).$
-Since $X$ and $Y$ are independent, and are $(s_{1},C_{1})$-Frostman and $(s_{2},C_{2})$-Frostman, respectively, one asserts that $W$ is $(s_{1}+s_{2},C_{1}C_{2}\alpha^{-s_{1}}\beta^{-s_{2}})$-Frostman.
-Here $\log(\alpha^{-s_{1}}\beta^{-s_{2}})=O_{\phi}(1)$.
-Noting that $range(W)$ is contained in the ball in ${\mathbb{R}}^{2}$ centered at origin with radius $O_{\phi,c_{1}}(1)$.
+$H_{n}\big(|W_{1}-W_{2}|^{2}\big)+5H_{n}(X)+5H_{n}(Y)\leq 8H_{n}(X+Y)+3H_{n}(\phi(X,Y))+O(\log\mathfrak{K}).$
+Since $X$ and $Y$ are independent, and are $(s_{1},C_{1})$-Frostman and $(s_{2},C_{2})$-Frostman, respectively, one asserts that $W$ is $(s_{1}+s_{2},C_{1}C_{2}|\alpha|^{-s_{1}/2}|\beta|^{-s_{2}/2})$-Frostman.
+Here $\log(|\alpha|^{-s_{1}/2}|\beta|^{-s_{2}/2})=O(\log\mathfrak{K})$ since $s_{1},s_{2}$ are bounded.
+Noting that $range(W)$ is contained in the ball in ${\mathbb{R}}^{2}$ centered at origin with radius $O\big(\sqrt{|\alpha|}c_{1}+\sqrt{|\beta|}c_{1}\big)$.
 So we can deduce by Lemma 2.32 and Theorem 2.41 that
-$H_{n}\big{(}X\big{)}\geq ns_{1}-\log C_{1}+O(1),\quad H_{n}\big{(}Y\big{)}\geq ns_{2}-\log C_{2}+O(1),$
-$H_{n}\big{(}|W_{1}-W_{2}|^{2}\big{)}\geq n(s+\epsilon_{0})+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O_{\phi,s}(1)$
+$\begin{split}&H_{n}\big(X\big)\geq ns_{1}-\log C_{1}+O(1),\quad H_{n}\big(Y\big)\geq ns_{2}-\log C_{2}+O(1),\\ &\qquad\qquad\qquad H_{n}\big(|W_{1}-W_{2}|^{2}\big)\geq n(s+\epsilon_{0})+O(\log\mathfrak{K})\end{split}$ (41)
 for some $\epsilon_{0}=\epsilon_{0}(s)>0$ and all $n\in\mathbb{N}$. It follows that
 $n\left(s+\epsilon_{0}+5s_{1}+5s_{2}\right)\leq 8H_{n}(X+Y)+3H_{n}(\phi(X,Y))$
-$\qquad\qquad+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi,s}(1).$
+$\qquad\qquad+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O(\log\mathfrak{K})+O_{s}(1).$
 Recalling that $s_{1}+s_{2}\geq 2s$, we conclude that
-$\max\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\}\geq n\left(s+\frac{\epsilon_{0}}{11}\right)$
-$\qquad\qquad+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi,s}(1).$
-Hence we complete the proof of the lemma with constant $\epsilon_{1}=\epsilon_{0}/11$.
+$\max\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\}\geq n\left(s+\frac{\epsilon_{0}}{11}\right)+O(\log\mathfrak{K}).$ (42)
+On the other hand, we can also deduce by independence, (9), (36) and (41) that
+$\min\{H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\}$
+$\qquad\geq\min\Big\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,\frac{H_{n}(\alpha X^{2})+H_{n}(\beta Y^{2})}{2}\Big\}+O(1)$
+$\qquad\geq\frac{H_{n}(X)+H_{n}(Y)}{2}+O(\log\mathfrak{K})\geq ns+O(\log\mathfrak{K}).$ (43)
+Putting (42) and (43) together, we complete the proof of the lemma with constant $\epsilon_{1}=\epsilon_{0}/11$.
 Next, if $\alpha$ and $\beta$ have different signs, the result will follow with a slightly worse constant. Without loss of generality, we assume that $\alpha>0>\beta$. Then
@@
 $H_{n}(|W-W^{\prime}|^{2})+H_{n}(X_{1},X_{2},Y_{1},Y_{2})$
-$\leq H_{n}(X_{1}+X_{2},Y_{1}+Y_{2},p(X_{1})+p(X_{2})-q(Y_{1})-q(Y_{2}))+H_{n}(X_{1}-X_{2},Y_{1}-Y_{2})+O_{\phi,c_{1}}(1)$
-$\leq H_{n}(X_{1}+X_{2})+H_{n}(Y_{1}+Y_{2})+H_{n}(p(X_{1})+p(X_{2})-q(Y_{1})-q(Y_{2}))$
-$\qquad+H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})+O(\log_{+}c_{1})+O_{\phi}(1).$
+$\leq H_{n}(X_{1}+X_{2},Y_{1}+Y_{2},\rho_{1}(X_{1})+\rho_{1}(X_{2})-\rho_{2}(Y_{1})-\rho_{2}(Y_{2}))+H_{n}(X_{1}-X_{2},Y_{1}-Y_{2})+O_{\phi,c_{1}}(1)$
+$\leq H_{n}(X_{1}+X_{2})+H_{n}(Y_{1}+Y_{2})+H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})$
+$\qquad+H_{n}(\rho_{1}(X_{1})+\rho_{1}(X_{2})-\rho_{2}(Y_{1})-\rho_{2}(Y_{2}))+O(\log_{+}c_{1})+O_{\phi}(1).$
 Compared with (40), Lemma 2.29 shows that
-$H_{n}(p(X_{1})+p(X_{2})+q(Y_{1})+q(Y_{2}))\leq 9H_{n}(p(X)+q(Y))$
-$\qquad\qquad\qquad-4H_{n}(X)-4H_{n}(Y)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1).$
+$H_{n}(\rho_{1}(X_{1})+\rho_{1}(X_{2})-\rho_{2}(Y_{1})-\rho_{2}(Y_{2}))\leq 9H_{n}(\rho_{1}(X)+\rho_{2}(Y))$
+$\qquad\qquad\qquad-4H_{n}(X)-4H_{n}(Y)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O(\log\mathfrak{K}).$
 It leads to
 $H_{n}\big(|W_{1}-W_{2}|^{2}\big)+$ $8H_{n}(X)+8H_{n}(Y)\leq 8H_{n}(X+Y)+9H_{n}(\phi(X,Y))$
-$\quad+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi}(1),$
+$\quad+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O(\log\mathfrak{K}),$
 and then
 $n\left(s+\epsilon_{0}+8s_{1}+8s_{2}\right)\leq 8H_{n}(X+Y)+9H_{n}(\phi(X,Y))$
-$\qquad\qquad+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O_{\phi,s}(1).$
+$\qquad\qquad+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O(\log\mathfrak{K})+O_{s}(1).$
 Finally, the lemma will follow with constant $\epsilon_{1}=\epsilon_{0}/17$. ∎
 Remark 4.2.
-We emphasize that the parameters $C_{1},C_{2}$ in Lemma 4.1 can be unbounded as $n$ tends to infinity.
-And, when we apply the above lemma in the next subsection, the term $O(\log C_{1}C_{2})$ will lead to some entropy-terms.
+We emphasize that $C_{1},C_{2}$ in Lemma 4.1 can be unbounded as $n$ tends to infinity.
+And, when we apply the above lemma in the following subsections, the term $O(\log C_{1}C_{2})$ will lead to some entropy-terms.
 So it cannot be replaced by $O_{C_{1},C_{2}}(1)$.
-4.2 Diagonal forms with dependent random variables
-Now we prove Theorem 1.2 in the following form, where $X$ and $Y$ are allowed to be dependent.
-We still take $\phi(x,y)=\alpha x^{2}+\beta y^{2}$ in this subsection.
-And a constant depending on $\phi$ means it depends on $\alpha$ and $\beta$.
+Lemma 4.3.
+Let
+$\rho_{1}(x)=\alpha_{2}x^{2}+\alpha_{1}x+\alpha_{0},\quad\rho_{2}(y)=\beta_{2}x^{2}+\beta_{1}x+\beta_{0},$
+where $\alpha_{2},\beta_{2}\in{\mathbb{R}}^{\ast}$ and $\alpha_{1},\alpha_{0},\beta_{1},\beta_{0}\in{\mathbb{R}}$.
+Let $1/2<s<1$ and $C_{1},C_{2},c>0$.
+There exists a constant $\epsilon_{1}^{\prime}=\epsilon_{1}^{\prime}(s)>0$ such that the following statement holds: Let $X$, $Y$ be independent random variables with ranges contained in [c,c][-c,c].
+Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, where $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
+Then
+$H_{n}(X+Y)+H_{n}(\rho_{1}(X)+\rho_{2}(Y))\geq n(2s+\epsilon_{1}^{\prime})+O(\log\mathfrak{K})$
+for all $n\in\mathbb{N}$, where
+$\log\mathfrak{K}$ $=O_{s}\big(\log_{+}C_{1}+\log_{+}C_{2}+1\big)$
+$+O\big(|\log|\alpha_{2}|\big|+\big|\log|\beta_{2}|\big|+\log_{+}|\alpha_{1}|+\log_{+}|\beta_{1}|+\log_{+}c).$
+Proof.
+Note that by completing the squares, we have
+$\rho_{1}(x)=\alpha_{2}\left(x+a\right)^{2}+\alpha_{0}-\frac{\alpha_{1}^{2}}{4\alpha_{2}},\quad\rho_{2}(y)=\beta_{2}\left(x+b\right)^{2}+\beta_{0}-\frac{\beta_{1}^{2}}{4\beta_{2}},$
+where $a=\frac{\alpha_{1}}{2\alpha_{2}}$ and $b=\frac{\beta_{1}}{2\beta_{2}}$. By Lemma 2.16,
+$\begin{split}H_{n}(X+Y)&=H_{n}((X+a)+(Y+b))+O(1),\\ H_{n}(\rho_{1}(X)+\rho_{2}(Y))&=H_{n}(\alpha_{2}(X+a)^{2}+\beta_{2}(Y+b)^{2})+O(1).\end{split}$ (44)
+Let $\eta_{i}$ $(i=1,2)$ be parameters to be determined later, with $0<\eta_{i}\leq\min\{(4C_{i})^{-1/s_{i}},\,1/10\}$.
+Denote by $E$ the event that $|X+a|\geq\eta_{1}$, $|Y+b|\geq\eta_{2}$.
+Let $(X_{1},Y_{1})$ be the random variable such that $(X_{1},Y_{1})\equiv(X+a,Y+b\,|\,E)$ (see Lemma A.2).
+One can verify that $X_{1}$ and $Y_{1}$ are still independent.
+Since $(X+a,Y+b)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, one has
+$P(E)\geq 1-P(|X+a|<\eta_{1})-P(|Y+b|<\eta_{2})\geq 1-C_{1}\eta_{1}^{s_{1}}-C_{2}\eta_{2}^{s_{2}}\geq 1/2.$ (45)
+Lemma A.2 tells us that $(X_{1},Y_{1})$ is marginally $(s_{1},O(C_{1});s_{2},O(C_{2}))$-Frostman. Also, it follows from Lemma A.2, (44) and (45) that
+$\begin{split}H_{n}(X_{1}+Y_{1})&\leq\frac{H_{n}((X+a)+(Y+b))}{P(E)}\leq\frac{H_{n}(X+Y)}{1+O(C_{1}\eta_{1}^{s_{1}}+C_{2}\eta_{2}^{s_{2}})}+O(1),\\ H_{n}(\alpha_{2}X_{1}^{2}+\beta_{2}Y_{1}^{2})&\leq\frac{H_{n}(\alpha_{2}(X+a)^{2}+\beta_{2}(Y+b)^{2})}{P(E)}\leq\frac{H_{n}(\rho_{1}(X)+\rho_{2}(Y))}{1+O(C_{1}\eta_{1}^{s_{1}}+C_{2}\eta_{2}^{s_{2}})}+O(1).\end{split}$ (46)
+By construction, we see that
+$range_{+}(X_{1})\subseteq[-c_{1},-\eta_{1}]\cup[\eta_{1},c_{1}],\quad range_{+}(Y_{1})\subseteq[-c_{1},-\eta_{2}]\cup[\eta_{2},c_{1}]$
+with some $c_{1}$ satisfying $c_{1}\leq c+\max\{|a|,|b|\}$. In particular, we have $\log_{+}c_{1}=O(\log\mathfrak{K})$. By Lemma 4.1, there is some $\epsilon_{1}=\epsilon_{1}(s)\in(0,1)$ such that
+$\begin{split}&H_{n}(X_{1}+Y_{1})+H_{n}(\alpha_{2}X_{1}^{2}+\beta_{2}Y_{1}^{2})\geq n(2s+\epsilon_{1})\\ &\qquad\qquad\qquad+O(\log(1/\eta_{1}))+O(\log(1/\eta_{2}))+O(\log\mathfrak{K}).\end{split}$ (47)
+Combining (46) and (47), one deduces that
+$\begin{split}&H_{n}(X+Y)+H_{n}(\rho_{1}(X)+\rho_{2}(Y))\geq n(2s+\epsilon_{1})+O(\log\mathfrak{K})\\ &\qquad\qquad\qquad+O\big(C_{1}\eta_{1}^{s_{1}}n+C_{2}\eta_{2}^{s_{2}}n+\log(1/\eta_{1})+\log(1/\eta_{2})\big).\end{split}$
+Recall that we require $0<\eta_{i}<\min\{(4C_{i})^{-1/s_{i}},1/10\}$, $0<\epsilon_{1}<1$, and $\epsilon_{1}$ only depends on $s$.
+Now assume that the implied constant in the last term on the right-hand side of above inequality is $C^{\prime}$, with $C^{\prime}\geq 1$.
+We then take $\eta_{i}=\min\Big\{\big(\frac{\epsilon_{1}}{4C^{\prime}C_{i}}\big)^{1/s_{i}},\,\frac{1}{10}\Big\}$ $(i=1,2)$.
+Then $C^{\prime}C_{i}\eta_{i}^{s_{i}}n\leq n\epsilon_{1}/4$, and $O(\log_{+}(1/\eta_{i}))=O(\frac{1}{s_{i}}\log_{+}C_{i})+O_{s}(1)$.
+Recall $s_{1},s_{2}\gg s$.
+Therefore, the proof is completed by taking $\epsilon_{1}^{\prime}=\epsilon_{1}/2$.
+∎
+4.2 Diagonal polynomials of higher degrees with independent random variables
+We now extend the analysis from Section 4.1 to polynomials of arbitrary degree. Consider a polynomial
+$\rho(x)=\alpha_{d}x^{d}+\alpha_{d-1}x^{d-1}+\ldots+\alpha_{1}x+\alpha_{0},$
+where $d\geq 1$, $\alpha_{d}\in{\mathbb{R}}^{\ast}$ and $\alpha_{d-1},\ldots,\alpha_{0}\in{\mathbb{R}}$.
+To quantify the size of the polynomial coefficients, define
+$\mathfrak{c}(\rho)=\max\{|\alpha_{d}|^{-1},\,|\alpha_{d}|,\,|\alpha_{d-1}|,\,\ldots,\,|\alpha_{1}|,\,|\alpha_{0}|\}.$
+This quantity controls both the scale and inverse scale of $\rho$.
+For any $z\in{\mathbb{R}}$, define the difference polynomial $\Delta_{z}\rho$ to be
+$(\Delta_{z}\rho)(x)=\rho(x+z)-\rho(z)=\sum\limits_{j=0}^{d-1}\left(\sum\limits_{k=j+1}^{d}{k\choose j}\alpha_{k}z^{k-j}\right)x^{j},$
+which is a polynomial of degree d1d-1. Also assume that
+$\rho(x)=\alpha_{d}\prod\limits_{i=1}^{l}(x-z_{i})^{m_{i}},$
+where $1\leq l\leq d$, the numbers $z_{1},\ldots,z_{l}$ are distinct complex zeros of $\rho$, and $m_{1},\ldots,m_{l}$ are corresponding multiplicities satisfying $m_{1}+\ldots+m_{l}=d$.
+For any $\eta>0$, define the $\eta$-regular set of $\rho$ as
+$\mathcal{R}(\rho;\eta)=\big\{x\in{\mathbb{R}}:\,|x-z_{i}|\geq\eta\,(i=1,\ldots,l)\big\}.$
+This is the set of points in $\mathbb{R}$} that are at least distance $\eta$ from every root of $\rho$.
+The following three lemmas follow immediately from the definition or by calculation. We omit their proofs here.
+Lemma 4.4.
+Let $\rho$ be a real polynomial of degree $d\geq 1$. Let $\eta,c>0$.
+(i) The set ${\mathbb{R}}\setminus\mathcal{R}(\rho;\eta)$ is a union of at most $d$ open intervals. Each has radius at most $2\eta$.
+(ii) For all $x\in\mathcal{R}(\rho;\eta)$, we have $|\rho(x)|\geq|\alpha_{d}|\eta^{d}\geq\mathfrak{c}(\rho)^{-1}\eta^{d}$.
+(iii) For all $x\in[-c,c]$, we have $|\rho(x)|\ll_{d}\mathfrak{c}(\rho)\cdot\max\{c^{d},1\}$.
+Lemma 4.5.
+Let $\rho$ be a real polynomial of degree $d\geq 2$. The its derivative satisfies
+$\log\mathfrak{c}(\rho^{\prime})=\log\mathfrak{c}(\rho)+O_{d}(1)$
+Lemma 4.6.
+Let $\rho$ be a real polynomial of degree $d\geq 3$. Then for all $z\in{\mathbb{R}}^{\ast}$,
+$\log\mathfrak{c}(\Delta_{z}\rho)=\log\mathfrak{c}(\rho)+O\big(\big|\log|z|\big|\big)+O_{d}(1).$
+We will apply the following corollary, which is a direct consequence of Lemmas 4.4 and 4.5.
+Corollary 4.7.
+Let $\rho$ be a real polynomial of degree $d\geq 2$.
+(i) The set ${\mathbb{R}}\setminus\mathcal{R}(\rho^{\prime};\eta)$ is a union of at most d1d-1 open intervals. Each has length at most $2\eta$. And $\rho$ is monotonic on each of the intervals.
+(ii) Let $0<\eta<1$. Then for any $x\in\mathcal{R}(\rho;\eta)$,
+$\log_{+}\frac{1}{|\rho^{\prime}(x)|}\leq(d-1)\log(1/\eta)+\log\mathfrak{c}(\rho)+O_{d}(1).$
+(iii) Let $c>0$. Then for any $x\in[-c,c]$,
+$\log_{+}|\rho^{\prime}(x)|\leq(d-1)\log_{+}c+\log\mathfrak{c}(\rho)+O_{d}(1).$
+Lemma 4.8.
+Let $d\geq 2$ and $\rho_{1},\rho_{2}$ be real polynomials of degree $d$.
+Let $1/2<s<1$ and $C_{1},C_{2},c>0$.
+There exists a constant $\epsilon_{1}^{\prime\prime}=\epsilon_{1}^{\prime\prime}(d,s)>0$ such that the following statement holds: Let $X$, $Y$ be independent random variables with range contained in [c,c][-c,c].
+Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, where $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
+Then
+$H_{n}(X+Y)+H_{n}(\rho_{1}(X)+\rho_{2}(Y))\geq n(2s+\epsilon_{1}^{\prime\prime})+O(\log\mathfrak{K})$ (48)
+for all $n\in\mathbb{N}$, where
+$O(\log\mathfrak{K})=O_{d,s}\big(\log_{+}C_{1}+\log_{+}C_{2}+1\big)+O_{d}\big(\log\mathfrak{c}(\rho_{1})+\log\mathfrak{c}(\rho_{2})+\log_{+}c\big).$
+Proof.
+We use induction on the degree $d$.
+For $d=2$, the conclusion follows from Lemma 4.3.
+Assume that the conclusion holds for d1d-1, with $d\geq 3$.
+Let $\epsilon_{1}^{\prime}=\epsilon_{1}^{\prime\prime}(d-1,s)$ be the constant obtained in the inductive hypothesis.
+Without loss of generality, we assume that $0<\epsilon_{1}^{\prime}<1$.
+For $i=1,2$, let $\eta_{i}=\min\Big\{\big(\frac{\epsilon_{1}}{128dC_{i}}\big)^{1/s_{i}},\,\frac{1}{10}\Big\}$, which satisfies $dC_{i}\eta_{i}^{s_{i}}\leq\epsilon_{1}^{\prime}/128\leq 1/4$ and $\log(1/\eta_{i})=O_{s}(\log_{+}C_{i})+O_{d,s}(1)$.
+Let $E$ be the event that $X\in\mathcal{R}(\rho_{1}^{\prime};\eta_{1})$, $Y\in\mathcal{R}(\rho_{2}^{\prime};\eta_{2})$.
+On $E$, both $X$ and $Y$ avoid the $\eta$-neighborhoods of the critical points of $\rho_{1}$ and $\rho_{2}$, ensuring that $\rho_{1}^{\prime}(X)$ and $\rho_{2}^{\prime}(Y)$ are bounded away from zero.
+By Lemma 4.4(i), the complement ${\mathbb{R}}\setminus\mathcal{R}(\rho_{i}^{\prime};\eta_{i})$ is a union of d1d-1 intervals, each of length $2\eta_{i}$.
+Since $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, we have
+$P(E)$ $\geq 1-P(X\notin\mathcal{R}(\rho_{1}^{\prime};\eta_{1}))-P(Y\notin\mathcal{R}(\rho_{2}^{\prime};\eta_{2}))$
+$\geq 1-(d-1)C_{1}\eta_{1}^{s_{1}}-(d-1)C_{2}\eta_{2}^{s_{2}}\geq 1-\epsilon_{1}/64\geq 1/2.$
+We denote the conditional probability given $E$ by $P_{E}$.
+Let $(X_{0},Y_{0})$ be the random variable such that $(X_{0},Y_{0})\equiv(X,Y|E)$ (see Lemma A.2).
+One can check that $X_{0}$ and $Y_{0}$ are still independent.
+And $(X_{0},Y_{0})$ is marginally $(s_{1},O(C_{1});s_{2},O(C_{2}))$-Frostman.
+By Lemma 2.32(i),
+$H_{n}(X_{0})+H_{n}(Y_{0})\geq n(s_{1}+s_{2})-\log C_{1}C_{2}+O(1)\geq 2ns+O(\log\mathfrak{K}).$ (49)
+By (7) and Lemma 4.4, one has
+$H_{n}(X+Y)\leq n+\log_{+}c+O(1)=n+O(\log\mathfrak{K}),$
+$H_{n}(\rho_{1}(X)+\rho_{2}(Y))\leq n+\log\mathfrak{c}(\rho)+O_{d}(\log_{+}c+1)=n+O(\log\mathfrak{K}).$
+Together with Lemma A.2, one obtains
+$H_{n}(X_{0}+Y_{0})\leq(1+\epsilon_{1}^{\prime}/32)H_{n}(X+Y)\leq H_{n}(X+Y)+n\epsilon_{1}^{\prime}/32+O(\log\mathfrak{K}),$ (50)
+and
+$\begin{split}H_{n}(\rho_{1}(X_{0})+\rho_{2}(Y_{0}))&\leq(1+\epsilon_{1}^{\prime}/32)H_{n}(\rho_{1}(X)+\rho_{2}(Y))\\ &\leq H_{n}(\rho_{1}(X)+\rho_{2}(Y))+n\epsilon_{1}^{\prime}/32+O(\log\mathfrak{K}),\end{split}$ (51)
+where we have used the estimate $P(E)^{-1}\leq(1-\epsilon_{1}^{\prime}/64)^{-1}\leq 1+\epsilon_{1}^{\prime}/32$.
+Note that $range_{+}(X_{0})\subseteq\mathcal{R}(\rho_{1}^{\prime};\eta)\cap[-c,c]$ and $range_{+}(Y_{0})\subseteq\mathcal{R}(\rho_{2}^{\prime};\eta)\cap[-c,c]$.
+By Lemmas 2.13, and Corollary 4.7, we see that $\rho_{1}$ and $\rho_{2}$ are weakly $(\kappa,d-1)$-bi-Lipschitz on $range_{+}(X_{0})$ and $range_{+}(Y_{0})$, respectively, with $\log\kappa=O(\log\mathfrak{K})$.
+Now, let $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ be two i.i.d. copies of $(X_{0},Y_{0})$. In particular, $X_{1},X_{2},Y_{1},Y_{2}$ are independent. Combining Lemma 2.29 and (49),
+$H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2}))$
+$\qquad\leq 7H_{n}(\rho_{1}(X_{0})+\rho_{2}(Y_{0}))-3H_{n}(X_{0})-3H_{n}(Y_{0})+O(\log\mathfrak{K})$
+$\qquad\leq 7H_{n}(\rho_{1}(X_{0})+\rho_{2}(Y_{0}))-6ns+O(\log\mathfrak{K}).$ (52)
+We then examine the discretized conditional entropy
+$H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2})|X_{1}-X_{2},Y_{1}-Y_{2})+H_{n}(X_{2}+Y_{2}|X_{1}-X_{2},Y_{1}-Y_{2})$
+$\qquad\qquad\qquad\leq H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2}))+H_{n}(X_{2}+Y_{2}).$ (53)
+In what follows, we condition on the discretized differences. For brevity, write
+$H_{n}(\cdot|z_{1},z_{2}):=H\big(D_{n}(\cdot)|D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2}\big)$
+for any random variable depending on $(X_{1},X_{2},Y_{1},Y_{2})$. We will sum over all dyadic rationals $z_{1},z_{2}$ with $P_{E}\big(D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2}\big)>0$.
+Given $D_{n}(X_{1}-X_{2})=z_{1}$ and $D_{n}(Y_{1}-Y_{2})=z_{2}$, it follows by Corollary 4.7 that
+$\begin{split}&\log_{+}\Big(2^{n}\big|\big(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2})\big)-\big(\rho_{1}(X_{2}+z_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{2}+z_{2})-\rho_{2}(Y_{2})\big)\big|\Big)\\ &\qquad\qquad\qquad=O_{d}(\log_{+}c)+O(\log\mathfrak{c}(\rho_{1}))+O(\log\mathfrak{c}(\rho_{2}))+O_{d}(1)=O(\log\mathfrak{K})\end{split}$ (54)
+with probability $1$. Recall that
+$(\Delta_{z_{1}}\rho_{1})(x)=\rho_{1}(x+z_{1})-\rho_{1}(x),\quad(\Delta_{z_{2}}\rho_{2})(y)=\rho_{2}(y+z_{2})-\rho_{2}(y).$
+They are polynomials of degree d1d-1 in $x$ or $y$, respectively. By Lemma 2.19 and (54),
+$H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2})|z_{1},z_{2})=H_{n}((\Delta_{z_{1}}\rho_{1})(X_{2})+(\Delta_{z_{2}}\rho_{2})(Y_{2})|z_{1},z_{2})+O(\log\mathfrak{K}).$ (55)
+Since $(X_{1},X_{2})$ and $(Y_{1},Y_{2})$ are independent, then $X_{2}$ and $Y_{2}$ are conditionally independent given $D_{n}(X_{1}-X_{2})=z_{1}$, $D_{n}(Y_{1}-Y_{2})=z_{2}$.
+By Lemma A.2, there exist random variables, depending on $(z_{1},z_{2})$, that have the same distributions with $(X_{2},Y_{2}|D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2})$, respectively.
+Next, we will prove the following marginally Frostman results:
+$\begin{split}\big(X_{2}|D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2})\text{ is $\big(s_{1},\widetilde{C}_{1}\big)$-Frostman},\\ \big(Y_{2}|D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2}\big)\text{ is $\big(s_{2},\widetilde{C}_{2}\big)$-Frostman},\end{split}$
+where
+$\widetilde{C}_{1}=O\left(\frac{C_{1}^{2}\cdot 2^{-ns_{1}}}{P_{E}(D_{n}(X_{1}-X_{2})=z_{1})}\right),\quad\widetilde{C}_{2}=O\left(\frac{C_{2}^{2}\cdot 2^{-ns_{2}}}{P_{E}(D_{n}(Y_{1}-Y_{2})=z_{2})}\right).$ (56)
+The above implied constants do not depend on $(z_{1},z_{2})$.
+We present the proof for $X_{2}$, as the one for $Y_{2}$ can be done similarly.
+Indeed, by Lemmas 2.30 and 2.31, the random variable $(X_{2},X_{1}-X_{2})$ is jointly $(s_{1},s_{1},O(C_{1}))$-Frostman.
+Therefore, for any ball $B(x,r)$,
+$P_{E}(X_{2}\in B(x,r),D_{n}(X_{1}-X_{2})=z_{1})\ll C_{1}^{2}2^{-ns_{1}}r^{s_{1}}.$
+Combining independence,
+$P_{E}(X_{2}\in B(x,r)|D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2})$
+$\qquad=P_{E}(X_{2}\in B(x,r)|D_{n}(X_{1}-X_{2})=z_{1})\ll\frac{C_{1}^{2}2^{-ns_{1}}r^{s_{1}}}{P_{E}(D_{n}(X_{1}-X_{2})=z_{1})}.$
+So (56) follows.
+Now, for $i=1,2$, we take $\zeta_{i}=\min\Big\{\big(\frac{\epsilon_{1}^{\prime}}{400C_{i}}\big)^{1/s_{i}},\,1/16\Big\}$, which satisfies $\log(1/\zeta_{i})=O_{s}(\log_{+}C_{i})+O_{d,s}(1)=O(\log\mathfrak{K})$.
+If $n<\max\{\frac{1}{s_{1}}\log\frac{400C_{1}}{\epsilon_{1}^{\prime}},\,\frac{1}{s_{2}}\log\frac{400C_{2}}{\epsilon_{1}^{\prime}},\,4\}$, then either $\log C_{1}>ns_{1}+O_{d,s}(1)$ or $\log C_{2}>ns_{2}+O_{d,s}(1)$ or $n<4$.
+In view of $s_{1},s_{2}\gg s$, the term $O_{s}(\log_{+}C_{1})+O_{s}(\log_{+}C_{2})+O(1)$ may represent a negative number that is less than n-n, and (48) holds automatically.
+In the following, we assume that $n\geq\max\{\frac{1}{s_{1}}\log\frac{400C_{1}}{\epsilon_{1}^{\prime}},\,\frac{1}{s_{2}}\log\frac{400C_{2}}{\epsilon_{1}^{\prime}},\,4\}$.
+As a consequence, we have $\zeta_{i}\geq 2^{-ns_{i}-4}$ $(i=1,2)$.
+We first consider the case that $|z_{1}|\geq\zeta_{1}$ and $|z_{2}|\geq\zeta_{2}$. Note that $|z_{1}|,|z_{2}|\leq 2c$. In view of Lemma 4.6, the coefficients can be bounded by
+$\log\mathfrak{c}(\Delta_{z_{i}}\rho_{i})=\log\mathfrak{c}(\rho_{i})+O(\log_{+}c)+O(\log(1/\zeta_{i}))+O_{d}(1)=O(\log\mathfrak{K})$ (57)
+for $i=1,2$. Applying the inductive hypothesis (for polynomials of degree d1d-1) with the conditional Frostman constants from (56), we obtain
+$\begin{split}&H_{n}(X_{2}+Y_{2}|z_{1},z_{2})+H_{n}((\Delta_{z_{1}}\rho_{1})(X_{2})+(\Delta_{z_{2}}\rho_{2})(Y_{2})|z_{1},z_{2})\geq n(2s+\epsilon_{1}^{\prime})\\ &\qquad+O_{d}(\log\mathfrak{c}(\Delta_{z}\rho_{1})+\log\mathfrak{c}(\Delta_{z}\rho_{2}))+\log_{+}c)+O_{d,s}\big(\log_{+}\widetilde{C}_{1}+\log_{+}\widetilde{C}_{2}+1\big).\end{split}$ (58)
+Combining (55)-(58), we deduce that
+$\begin{split}&H_{n}(X_{2}+Y_{2}|z_{1},z_{2})+H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2})|z_{1},z_{2})\geq n(2s+\epsilon_{1}^{\prime})+O(\log\mathfrak{K})\\ &\quad+O_{d,s}\big(-n(s_{1}+s_{2})-\log P(D_{n}(X_{1}-X_{2})=z_{1})-\log P(D_{n}(Y_{1}-Y_{2})=z_{2})\big).\end{split}$ (59)
+For the case $|z_{1}|<\zeta_{1}$ or $|z_{1}|<\zeta_{2}$, we use the trivial bound that
+$H_{n}(X_{2}+Y_{2}|z_{1},z_{2})+H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2})|z_{1},z_{2})$
+$\geq H_{n}(X_{2}+Y_{2}|z_{1},z_{2})\geq\frac{1}{2}\big(H_{n}(X_{2}|z_{1},z_{2})+H_{n}(Y_{2}|z_{1},z_{2})+O(1)\big)$
+$\geq ns+O_{d,s}\big(\log_{+}C_{1}+\log_{+}C_{2}-n(s_{1}+s_{2})$
+$\qquad\qquad\qquad-\log P(D_{n}(X_{1}-X_{2})=z_{1})-\log P(D_{n}(Y_{1}-Y_{2})=z_{2})\big)+O(1),$ (60)
+which reslults from conditional independence, (9) and Lemma 2.32.
+We now average over all possible values of the discretized differences.
+Multiplying (59) and (60) by $P(D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2})$ and summing over all $(z_{1},z_{2})$, we obtain weighted sums of conditional entropies.
+The left-hand side gives
+$\sum\limits_{z_{1},z_{2}}P(D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2})H_{n}(X_{2}+Y_{2}|z_{1},z_{2})=H_{n}(X_{2}+Y_{2}|X_{1}-X_{2},Y_{1}-Y_{2})$
+by the definition of conditional entropy. For the negative logarithm terms, we use the fact that
+$-\sum\limits_{z_{1},z_{2}}P(D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2})\log P(D_{n}(X_{1}-X_{2})=z_{1})$
+$\qquad=-\sum\limits_{z_{1}}P(D_{n}(X_{1}-X_{2})=z_{1})\log P(D_{n}(X_{1}-X_{2})=z_{1})=H_{n}(X_{1}-X_{2}).$
+And other terms can be evaluated similarly. It leads to
+$H_{n}(X_{2}+Y_{2}|X_{1}-X_{2},Y_{1}-Y_{2})+H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2})|X_{1}-X_{2},Y_{1}-Y_{2})$
+$\qquad\geq n(2s+\epsilon_{1}^{\prime})P_{E}(|D_{n}(X_{1}-X_{2})|\geq\zeta_{1},|D_{n}(Y_{1}-Y_{2})|\geq\zeta_{2})$
+$\qquad\qquad+ns(P_{E}(|D_{n}(X_{1}-X_{2})|<\zeta_{1}\text{ or }|D_{n}(Y_{1}-Y_{2})|<\zeta_{2}))$
+$\qquad\qquad+O_{d,s}(-n(s_{1}+s_{2})+H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2}))+O(\log K).$ (61)
+Recall that $\zeta_{i}\geq 2^{-ns_{i}-4}$ $(i=1,2)$, which gives $\eta_{i}+2^{-n}\leq 20\eta_{i}$.
+Also note that $X_{1}-X_{2}$ and $Y_{1}-Y_{2}$ are $(s_{1},O(C_{1}))$-Frostman and $(s_{2},O(C_{2}))$-Frostman, respectively.
+We deduce that
+$P_{E}(|D_{n}(X_{1}-X_{2})|<\zeta_{1})\leq C_{1}(20\zeta_{1})^{s_{1}}\leq\epsilon_{1}^{\prime}/16,\quad P_{E}(|D_{n}(Y_{1}-Y_{2})|<\zeta_{2})\leq C_{2}(20\zeta_{2})^{s_{2}}\leq\epsilon_{1}^{\prime}/16,$
+and
+$H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})\geq n(s_{1}+s_{2})-\log_{+}C_{1}-\log_{+}C_{2}+O(1).$ (62)
+So $P_{E}(|D_{n}(X_{1}-X_{2})|\geq\zeta_{1},|D_{n}(Y_{1}-Y_{2})|\geq\zeta_{2})\geq 1-\epsilon_{1}^{\prime}/8$. Now (61) and (62) leads to
+$H_{n}(X_{2}+Y_{2}|X_{1}-X_{2},Y_{1}-Y_{2})+H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2})|X_{1}-X_{2},Y_{1}-Y_{2})$
+$\qquad\geq n(2s+\epsilon_{1}^{\prime})(1-\epsilon_{1}^{\prime}/8)+O(\log\mathfrak{K})\geq n(2s+3\epsilon_{1}^{\prime}/4)+O(\log\mathfrak{K}).$ (63)
+Finally, putting (50), (51), (4.2), (53) and (63) together, we arrive at
+$H_{n}(X+Y)+7H_{n}(\rho_{1}(X)+\rho_{2}(Y))\geq n(8s+\epsilon_{1}^{\prime}/2)+O(\log\mathfrak{K}).$
+By independence, we also have
+$H_{n}(X+Y)\geq\frac{H_{n}(X)+H_{n}(Y)}{2}+O(1)\geq ns-\log C_{1}C_{2}+O(1).$
+It follows that
+$7H_{n}(X+Y)+7H_{n}(\rho_{1}(X)+\rho_{2}(Y))\geq n(14s+\epsilon_{1}^{\prime}/2)+O(\log\mathfrak{K}).$
+The induction is completed by taking $\epsilon_{1}^{\prime\prime}=\epsilon_{1}^{\prime}/14$, i.e., $\epsilon_{1}^{\prime\prime}(d,s)=\epsilon_{1}^{\prime\prime}(d-1,s)/14$. ∎
+4.3 Diagonal polynomials with dependent random variables
+The results of Sections 4.1 and 4.2 assumed the independence of $X$ and $Y$.
+We now extend to the dependent case, proving Theorem 1.2 in full generality.
+The key tool is Theorem 3.1, which turns dependent random variables to a family of conditionally independent random variables at the cost of logarithmic error terms.
 Theorem 1.2’.
-Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$, $C,c>0$.
-There is a positive constant $\epsilon_{2}=\epsilon_{2}(s)$ such that the following statement holds: For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
-$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\geq n(s+\epsilon_{2})+O_{s}(\log_{+}C)+O(\log_{+}c)+O_{\phi,s}(1).$
+Let $d\geq 2$ and $\rho_{1}$, $\rho_{2}$ be real polynomials of degree $d$.
+Let $0<s<1$, $C,c>0$.
+There is a positive constant $\epsilon_{2}=\epsilon_{2}(d,s)$ such that the following statement holds: For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
+$\begin{split}&\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\rho_{1}(X)+\rho_{2}(Y))\right\}\geq n(s+\epsilon_{2})+O(\log\mathfrak{K})\end{split}$
+for all $n\in\mathbb{N}$, where
+$\log\mathfrak{K}=O_{d,s}(\log_{+}C+\log_{+}c)+O_{\rho_{1},\rho_{2},s}(1).$
 Proof.
-For simplicity, we write
-$\Psi:=\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}-ns,$
-and our aim is to prove that $\Psi\geq\epsilon n$ for some positive $\epsilon$. In particular,
-$\frac{H_{n}(X)+H_{n}(Y)}{2},\,\,H_{n}(X+Y),\,\,H_{n}(\alpha X^{2}+\beta Y^{2})\,\leq\,ns+\Psi.$ (45)
-By Lemma 2.30, the random variables $X$, $Y$, $(X,Y)$ are $(s_{1},C\cdot\max\{1,c\})$-Frostman, $(s_{2},C\cdot\max\{1,c\})$-Frostman and $(s_{1}+s_{2},C)$-Frostman, respectively.
-Let $\eta$ be a parameter to be chosen later, with $0<\eta<\min\{2^{-1}(1/2C)^{1/2s},\,1/10\}$.
-We denote by $E$ the event that $|(X,Y)|\geq\eta$, and by $P_{E}$ the corresponding conditional probability measure.
-Denote by $(X_{1},Y_{1})$ the random variable such that $(X_{1},Y_{1})\equiv(X,Y|E)$.
-Since
-$P(E)$ $=P(|(X,Y)|\geq\eta)=1-P\big{(}(X,Y)\in B((0,0),\eta)\big{)}$
-$\geq 1-C\eta^{s_{1}+s_{2}}\geq 1-C\eta^{2s}>\frac{1}{2},$
-Lemma A.2 shows that the random variables $X_{1}$, $Y_{1}$, $(X_{1},Y_{1})$ are $(s_{1},O(C\cdot\max\{1,c\}))$-Frostman, $(s_{2},O(C\cdot\max\{1,c\}))$-Frostman and $(s_{1}+s_{2},O(C)))$-Frostman, respectively.
-Combining Lemma 2.32(i), one sees that
-$\begin{split}H_{n}(X_{1})+H_{n}(Y_{1})&\geq n(s_{1}+s_{2})-2\log C+O(\log_{+}c)+O(1),\\ H_{n}(X_{1},Y_{1})&\geq n(s_{1}+s_{2})-\log C+O(1).\end{split}$ (46)
-Since $s_{1}+s_{2}\geq 2s$, the right-hand sides of both inequalities in (46) are at least $2ns+O(\log_{+}C)+O(\log_{+}c)+O(1)$.
-On the other hand, We can apply Lemma A.2, then combine (45) to get that
-$\begin{split}&\frac{H_{n}(X_{1})+H(Y_{1})}{2}\leq\frac{H_{n}(X)+H_{n}(Y)}{2P(E)}\leq(1+O(C\eta^{2s}))(ns+\Psi),\\ &H_{n}(X_{1}+Y_{1})\leq\frac{H_{n}(X+Y)}{P(E)}\leq(1+O(C\eta^{2s}))(ns+\Psi),\\ &H_{n}(\alpha X_{1}^{2}+\beta Y_{1}^{2})\leq\frac{H_{n}(\alpha X^{2}+\beta Y^{2})}{P(E)}\leq(1+O(C\eta^{2s}))(ns+\Psi).\end{split}$ (47)
+To quantify the entropy gain beyond the trivial $ns$, we write the entropy increment as
+$\Psi:=\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\rho_{1}(X)+\rho_{2}(Y))\right\}-ns.$
+Our goal is to show $\Psi\geq n\epsilon_{2}+O(\log\mathfrak{K})$ for some $\epsilon_{2}=\epsilon_{2}(d,s)>0$. In particular,
+$\frac{H_{n}(X)+H_{n}(Y)}{2},\,\,H_{n}(X+Y),\,\,H_{n}(\rho_{1}(X)+\rho_{2}(Y))\,\leq\,ns+\Psi.$ (64)
+By Lemma 2.30, the random variable $(X,Y)$ is $(s_{1}+s_{2},C)$-Frostman, and also marginally $(s_{1},C_{0};s_{2},C_{0})$-Frostman, with
+$C_{0}=C\cdot\max\{1,c\},$ (65)
+where $c$ comes from the range of $X$ and $Y$.
+Let $\epsilon_{1}^{\prime\prime}=\epsilon_{1}^{\prime\prime}(d,s)$ be the constant shown in Lemma 4.8.
+Without loss of generality, we assume that $0<\epsilon_{1}^{\prime\prime}<1$.
+We use a cutoff parameter $\mathfrak{r}\geq 1$ as previous.
+For $i=1,2$, take $\eta_{i}=\min\Big\{\big(\frac{\epsilon_{1}^{\prime\prime}}{4\mathfrak{r}dC_{0}}\big)^{1/s_{i}},\,\frac{1}{10}\Big\}$, which satisfies $dC_{0}\eta_{i}^{s_{i}}\leq\epsilon_{1}^{\prime\prime}/4\mathfrak{r}\leq 1/4$ and $\log(1/\eta_{i})=O(\log\mathfrak{K})$.
+Let $E$ be the event that $X\in\mathcal{R}(\rho_{1}^{\prime};\eta_{1})$, $Y\in\mathcal{R}(\rho_{2}^{\prime};\eta_{2})$.
+Since $(X,Y)$ is marginally $(s_{1},C_{0};s_{2},C_{0})$-Frostman, we deduce by Lemma 4.4 that
+$P(E)$ $\geq 1-P(X\notin\mathcal{R}(\rho_{1}^{\prime};\eta_{1}))-P(Y\notin\mathcal{R}(\rho_{2}^{\prime};\eta_{2}))$
+$\geq 1-(d-1)C_{0}\eta_{1}^{s_{1}}-(d-1)C_{0}\eta_{2}^{s_{2}}\geq 1-\epsilon_{1}^{\prime\prime}/2\mathfrak{r}\geq 1/2.$
+We denote the conditional probability given $E$ by $P_{E}$.
+Let $(X_{0},Y_{0})$ be the random variable such that $(X_{0},Y_{0})\equiv(X,Y|E)$ (see Lemma A.2).
+By this construction, the functions $\rho_{1}$ and $\rho_{2}$ are weakly $(\kappa,d-1)$-bi-Lipschitz on $range_{+}(X_{0})$ and $range_{+}(Y_{0})$, respectively, with $\log\kappa=O(\log\mathfrak{K})$.
+ALso, the random variable $(X_{0},Y_{0})$ is both $(s_{1}+s_{2},O(C))$-Frostman, which gives
+$H_{n}(X_{0})+H_{n}(Y_{0})\geq H_{n}(X_{0},Y_{0})\geq n(s_{1}+s_{2})-\log C+O(1)\geq 2ns+O(\log\mathfrak{K}).$ (66)
+On the other hand, We can apply Lemma A.2 with $P(E)^{-1}\leq 1+\epsilon_{1}^{\prime\prime}/\mathfrak{r}$, then combine (64) to get that
+$\frac{H_{n}(X_{0})+H(Y_{0})}{2},\,H_{n}(X_{0}+Y_{0}),\,H_{n}(\rho_{1}(X_{0})+\rho_{2}(Y_{0}))\,\leq(1+\epsilon_{1}^{\prime\prime}/\mathfrak{r})(ns+\Psi).$ (67)
 It follows from (66) and (67) that
-$\begin{split}&H_{n}(X_{1})+H_{n}(Y_{1})-H_{n}(X_{1},Y_{1})=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(1),\\ &H_{n}(X_{1}+Y_{1})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(1),\\ &H_{n}(\alpha X_{1}^{2}+\beta Y_{1}^{2})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(1).\end{split}$
-Now we write $\log K:=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(1)$.
-Applying Theorem 3.1 with the random variable $(X_{1},Y_{1})$, we obtain the random variable $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$, where $X^{\prime\prime}$ and $Y^{\prime\prime}$ are conditionally independent given $D_{n}(X^{\prime},Y^{\prime})$, and
-$H_{n}(X^{\prime\prime}|X^{\prime},Y^{\prime})\geq H_{n}(X_{1})-\log K,\quad H_{n}(Y^{\prime\prime}|X^{\prime},Y^{\prime})\geq H_{n}(Y_{1})-\log K.\\$ (48)
-First, let us take the functions $f$ and $g$ in Theorem 3.1 to be the identity functions. Then
-$H_{n}(X^{\prime\prime}+Y^{\prime\prime}|X^{\prime},Y^{\prime})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}\leq 7\log K+O(1).$ (49)
-Second, we take the functions $f,g$ in Theorem 3.1 to be $f(x)=\alpha x^{2}$ and $g(y)=\beta y^{2}$.
-By the construction $(X_{1},Y_{1})\equiv(X,Y||(X,Y)|>\eta)$, the ranges of $X^{\prime\prime}$ and $Y^{\prime\prime}$ are contained in $c,-\eta/\sqrt{2}]\cup[\eta/\sqrt{2},c]$ with probability $1$.
-Now, by Example 2.12, the random variables $f(X^{\prime\prime})$ and $g(Y^{\prime\prime})$ are weakly $O_{\phi}(\max\{\eta^{-1},c\})$-bi-Lipschitz functions of $X^{\prime\prime}$ and $Y^{\prime\prime}$, respectively.
-It follows that
-$H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|X^{\prime},Y^{\prime})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}\leq 7\log K+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi}(1).$ (50)
+$\begin{split}&H_{n}(X_{0})+H_{n}(Y_{0})-H_{n}(X_{0},Y_{0})\leq\log K\\ &H_{n}(X_{0}+Y_{0})-\frac{H_{n}(X_{0})+H_{n}(Y_{0})}{2}\leq\log K,\\ &H_{n}(\rho_{1}(X_{0})+\rho_{2}(Y_{0}))-\frac{H_{n}(X_{0})+H_{n}(Y_{0})}{2}\leq\log K,\end{split}$
+where $\log K=O(\Psi)+O(n\epsilon^{\prime\prime}/\mathfrak{r})+O(\log\mathfrak{K})$.
+We now apply Theorem 3.1 to $(X_{0},Y_{0})$, and obtain a quadruple $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$ with certain properties (see (i)-(iv) in Theorem 3.1).
+In particular, the random variables $X^{\prime\prime}$ and $Y^{\prime\prime}$ are conditionally independent given $D_{n}(X^{\prime},Y^{\prime})$.
+Let us take the functions $f$ and $g$ in Theorem 3.1 to be the identity functions.
+Then
+$H_{n}(X^{\prime\prime}+Y^{\prime\prime}|X^{\prime},Y^{\prime})-\frac{H_{n}(X_{0})+H_{n}(Y_{0})}{2}\leq 7\log K+O(1).$ (68)
+Again, we take the functions $f,g$ in Theorem 3.1 to be $f(x)=\rho_{1}(x)$ and $g(y)=\rho_{2}(y)$. With the weakly $(\kappa,d)$-bi-Lipschitz property, it satisfies that
+$H_{n}(\rho_{1}(X^{\prime\prime})+\rho_{2}(Y^{\prime\prime})|X^{\prime},Y^{\prime})-\frac{H_{n}(X_{0})+H_{n}(Y_{0})}{2}\leq 7\log K+O(\log\mathfrak{K}).$ (69)
 Next, conditionally given that $D_{n}(X^{\prime},Y^{\prime})=(x,y)$ with $x,y$ being dyadic rationals satisfying $P\big(D_{n}(X^{\prime},Y^{\prime})=(x,y)\big)>0$, we will prove the following results on induced conditional probability measures:
@@
 $\qquad=P_{E}(X^{\prime}\in B(z,r)|D_{n}(Y^{\prime})=y)=\frac{P_{E}(X^{\prime}\in B(z,r),D_{n}(Y^{\prime})=y)}{P_{E}(D_{n}(Y^{\prime})=y)}$
-$\qquad=\frac{P_{E}(X_{1}\in B(z,r),D_{n}(Y_{1})=y)}{P_{E}(D_{n}(Y^{\prime})=y)}\leq\frac{2P(X\in B(z,r),D_{n}(Y)=y)}{P_{E}(D_{n}(Y^{\prime})=y)}.$
+$\qquad=\frac{P_{E}(X_{0}\in B(z,r),D_{n}(Y_{0})=y)}{P_{E}(D_{n}(Y^{\prime})=y)}\leq\frac{2P(X\in B(z,r),D_{n}(Y)=y)}{P_{E}(D_{n}(Y^{\prime})=y)}.$
 Since $(X,Y)$ is jointly $(s_{1},s_{2},C)$-Frostman, we have
 $P(X\in B(z,r),D_{n}(Y)=y)\ll Cr^{s_{1}}2^{-ns_{2}}.$
-This completes the proof for the claims given in (51). We mention that these two Frostman constants in (51) are $\gg 1$.
+This completes the proof for the claims given in (70).
 As the expressions to follow get rather lengthy, we now introduce some shorthand.
 In what follows, $H_{n}(~\cdot~|x,y)$ is an abbreviation of $H(D_{n}(\cdot)~|D_{n}(X^{\prime},Y^{\prime})=(x,y)).$ By Lemma A.2, there exist random variables, depending on $(x,y)$, that have the same distributions with $(X^{\prime\prime},Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y))$, respectively.
-These random variables take the same value as $(X^{\prime\prime},Y^{\prime\prime})$ does.
-By construction, their ranges are contained in $\big{(}[-c,-\eta/\sqrt{2}]\cup[\eta/\sqrt{2},c]\big{)}^{2}$ with probability $1$.
-Also note that $(X^{\prime\prime}|D_{n}(X_{1},Y_{1})=(x,y))$ and $(Y^{\prime\prime}|D_{n}(X_{1},Y_{1})=(x,y))$ are independent.
-Recall that the constant $\epsilon_{1}=\epsilon_{1}(s)$ in Lemma 4.1 is independent of the concrete random variables.
-Finally, recall (51).
-Now we can apply Lemma 4.1 with these random variables, with $C_{1}=O\big{(}C\cdot 2^{-ns_{2}}/P_{E}(D_{n}(Y^{\prime})=y)\big{)}$, $C_{2}=O\big{(}C\cdot 2^{-ns_{1}}/P_{E}(D_{n}(X^{\prime})=x)\big{)}$, $c_{1}=c$ and $c_{2}=\eta$, to deduce that
-$\max\left\{H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y),H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|x,y)\right\}\geq n(s+\epsilon_{1})$
-$\qquad+O\big{(}\log_{+}C-n(s_{1}+s_{2})-\log P_{E}(D_{n}(X^{\prime})=x)-\log P_{E}(D_{n}(Y^{\prime})=y)\big{)}$
-$\qquad+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1)$ (52)
-for some $\epsilon_{1}=\epsilon_{1}(s)>0$. By conditional independence, and combining (13), we also have that
-$H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y)\geq\max\{H_{n}(X^{\prime\prime}|x,y),H_{n}(Y^{\prime\prime}|x,y)\}+O(1).$
-Using a similar argument, Lemma A.2, and Example 2.16, one obtains
-$H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|x,y)\geq\max\{H_{n}(\alpha X^{\prime\prime 2}|x,y),H_{n}(\beta Y^{\prime\prime 2}|x,y)\}+O(1)$
-$\qquad=\max\{H_{n}(X^{\prime\prime}|x,y),H_{n}(Y^{\prime\prime}|x,y)\}+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi}(1).$
-Then
-$\min\{H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y),H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|x,y)\}$
-$\qquad\geq\frac{H_{n}(X^{\prime\prime}|x,y)+H_{n}(Y^{\prime\prime}|x,y)}{2}+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi}(1).$ (53)
-Therefore given the condition that $D_{n}(X_{1},Y_{1})=(x,y)$, it follows from (52) and (53) that
-$H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y)+H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|x,y)-\frac{H_{n}(X^{\prime\prime}|x,y)+H_{n}(Y^{\prime\prime}|x,y)}{2}$
-$\quad\geq n(s+\epsilon_{1})+O\big{(}-n(s_{1}+s_{2})-\log P_{E}(D_{n}(X^{\prime})=x)-\log P_{E}(D_{n}(Y^{\prime})=y)\big{)}$
-$\qquad\qquad\qquad\quad+O(\log_{+}C)+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1).$
-Now we sum all of these expressions respective to each $(x,y)$ with the weight $P_{E}(D_{n}(X^{\prime},Y^{\prime})=(x,y))$. Note that
-$\sum\limits_{x,y}P_{E}(D_{n}(X^{\prime},Y^{\prime})=(x,y))H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y)$
-$\qquad=\sum\limits_{x,y}P_{E}(D_{n}(X^{\prime\prime},Y^{\prime\prime})=(x,y))H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y)=H_{n}(X^{\prime\prime}+Y^{\prime\prime}|X^{\prime},Y^{\prime}),\text{ and}$
-$-\sum\limits_{x,y}P_{E}(D_{n}(X^{\prime},Y^{\prime})=(x,y))\log P_{E}(D_{n}(X^{\prime})=x)$
-$\qquad=-\sum\limits_{x}P_{E}(D_{n}(X^{\prime})=x)\log P_{E}(D_{n}(X^{\prime})=x)=H_{n}(X^{\prime}).$
-The other terms behave similarly. Therefore,
-$H_{n}(X^{\prime\prime}+Y^{\prime\prime}|X^{\prime},Y^{\prime})+H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|X^{\prime},Y^{\prime})-\frac{H_{n}(X^{\prime\prime}|X^{\prime},Y^{\prime})+H_{n}(Y^{\prime\prime}|X^{\prime},Y^{\prime})}{2}$
-$\qquad\qquad\geq n(s+\epsilon_{1})+O(-n(s_{1}+s_{2})+H_{n}(X^{\prime})+H_{n}(Y^{\prime}))$
-$\qquad\qquad\qquad\quad+O(\log_{+}C)+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1).$ (54)
-Substituting from (48)–(50), and further combining with $(X^{\prime},Y^{\prime})\equiv(X_{1},Y_{1})$ we obtain that
-$\begin{split}&\left(\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}+7\log K\right)+\left(\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}+7\log K\right)\\ &\qquad-\frac{(H_{n}(X_{1})-\log K)+(H_{n}(Y_{1})-\log K)}{2}-ns+O(H_{n}(X_{1})+H_{n}(Y_{1})-n(s_{1}+s_{2}))\\ &\qquad\geq n\epsilon_{1}+O(\log_{+}C)+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1).\end{split}$ (55)
-Recalling (47), the fact $s_{1}+s_{2}\geq 2s$, and the notation $\log K=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(1)$, the left-hand side of (55) is
-$\leq O\left(\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}-ns\right)+15\log K$
-$=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1).$
-So there is some absolute constant $\tilde{C}>0$ such that
-$\tilde{C}\Psi\geq n\epsilon_{1}+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1).$
-Recall that we require $0<\eta<\min\{2^{-1}(1/2C)^{1/2s},1/10\}$, $0<\epsilon_{1}<1$, and $\epsilon_{1}$ only depends on $s$.
-Now assume that the implied constant in the term $O(C\eta^{2s}n)$ is $C^{\prime}$, with $C^{\prime}\geq 1$.
-We then take $\eta=\min\Big{\{}\frac{1}{2}\big{(}\frac{\epsilon_{1}}{2C^{\prime}C}\big{)}^{1/2s},\,\frac{1}{10}\Big{\}}$.
-Then $C^{\prime}C\eta^{2s}n\leq n\epsilon_{1}/2$.
-Therefore,
-$\Psi\gg n\epsilon_{1}+O_{s}(\log_{+}C)+O(\log_{+}c)+O_{\phi,s}(1).$
-The proof is completed. ∎
-4.3 The general forms
-Any non-degnerate rational quadratic from $\phi(x,y)=a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ can be diagonalized by rational linear change of variables, i.e.,
-$\phi(x,y)=\alpha(\xi_{11}x+\xi_{12}y)^{2}+\beta(\xi_{21}x+\xi_{22}y)^{2}$ (56)
-for some $\alpha,\beta\in\mathbb{Q}^{\ast}$ and $\xi_{11},\xi_{12},\xi_{21},\xi_{22}\in\mathbb{Q}$ with $\xi_{11}\xi_{22}-\xi_{12}\xi_{21}\neq 0$.
-One can verify that $\phi(x,y)$ is diagonal if and only if exactly two of the coefficients $\xi_{11},\xi_{12},\xi_{21},\xi_{22}$ are zero.
-For non-diagonal case, at most one of $\xi_{11},\xi_{12},\xi_{21},\xi_{22}$ may be zero.
-In the following, we will show that, if exactly one of these numbers is zero, then we can transform $\phi(x,y)$ into another form like (56) with all corresponding coefficients non-zero.
-Now we assume without loss of generality that $\xi_{11},\xi_{21},\xi_{22}\neq 0$ and $\xi_{21}=0$.
-By dividing $\xi_{11}$ and $\xi_{12}$ in the two brackets, respectively, we may rewrite $\phi(x,y)=\alpha^{\prime}x^{2}+\beta^{\prime}(x+d^{\prime}y)^{2}$, with $\alpha^{\prime},\beta^{\prime},d^{\prime}\in\mathbb{Q}^{\ast}$.
-Now, let us take some $a,c\in\mathbb{Q}^{\ast}$ satisfying $a\neq c$, $a\neq 1/d^{\prime}$, $c\neq 1/d^{\prime}$ and
-$(d^{\prime}a-1)(d^{\prime}c-1)=-\alpha^{\prime}/\beta^{\prime}.$
-One can check that
-$\phi(x,y)=\frac{\beta^{\prime}d^{\prime}(1-cd^{\prime})}{a-c}(ax+y)^{2}+\frac{\beta^{\prime}d^{\prime}(ad^{\prime}-1)}{a-c}(cx+y)^{2}.$
-All coefficient in above formula are non-zero rational numbers.
-Furthermore, we can assume in (56) without loss of generality that $\xi_{11}\neq-\xi_{21}$ and $\xi_{12}\neq-\xi_{22}$, because non-zero constants can be multiplied inside both brackets.
-In all situations, all the coefficients depend on $a_{1},a_{2},a_{3}$.
+Similar as previously, we can apply Lemma 4.8 with these random variables, with $C_{1}=O\big(C\cdot 2^{-ns_{2}}/P_{E}(D_{n}(Y^{\prime})=y)\big)$ and $C_{2}=O\big(C\cdot 2^{-ns_{1}}/P_{E}(D_{n}(X^{\prime})=x)\big)$, to deduce that
+$H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y)+H_{n}(\rho_{1}(X^{\prime\prime})+\rho_{2}(Y^{\prime\prime})|x,y)\geq n(2s+\epsilon_{1}^{\prime\prime})+O_{d}(\log_{+}c)+O_{\phi,s}(1)$
+$\qquad+O_{d,s}\big(\log_{+}C-n(s_{1}+s_{2})-\log P_{E}(D_{n}(X^{\prime})=x)-\log P_{E}(D_{n}(Y^{\prime})=y)\big).$
+Now we sum all of these expressions respective to each $(x,y)$ with the weight $P_{E}(D_{n}(X^{\prime},Y^{\prime})=(x,y))$, which yields
+$\begin{split}&H_{n}(X^{\prime\prime}+Y^{\prime\prime}|X^{\prime},Y^{\prime})+H_{n}(\rho_{1}(X^{\prime\prime})+\rho_{2}(Y^{\prime\prime 2})|X^{\prime},Y^{\prime})\geq n(2s+\epsilon_{1}^{\prime\prime})\\ &\qquad\qquad\qquad+O_{d,s}(-n(s_{1}+s_{2})+H_{n}(X^{\prime})+H_{n}(Y^{\prime}))+O(\log\mathfrak{K}).\end{split}$ (71)
+Substituting from (66), (68),(69), and further combining with $(X^{\prime},Y^{\prime})\equiv(X_{0},Y_{0})$ we obtain that
+$H_{n}(X_{0})+H_{n}(Y_{0})+14\log K+O(\log\mathfrak{K})\geq n(2s+\epsilon_{1}^{\prime\prime})$ (72)
+Recalling (67) and the notation $\log K=O(\Psi)+O(n\epsilon^{\prime\prime}/\mathfrak{r})+O(\log\mathfrak{K})$, the left-hand side of (72) is
+$\leq 2ns+O(\Psi)+O(n\epsilon^{\prime\prime}/\mathfrak{r})+O(\log\mathfrak{K}).$ (73)
+Combining (72) and (73), there is some absolute constant $\tilde{C}>0$ such that
+$\tilde{C}\Psi\geq n\epsilon_{1}^{\prime\prime}+O(\log\mathfrak{K})+O(n\epsilon^{\prime\prime}/\mathfrak{r}).$
+We now take $\mathfrak{r}\geq 1$ to be some absolute constant such that the last term on the right-hand side of above formula satisfies $O(n\epsilon_{1}^{\prime\prime}/\mathfrak{r})\leq n\epsilon_{1}^{\prime\prime}/2$, Therefore,
+$\Psi\geq n\epsilon_{2}+O(\log\mathfrak{K})$
+for $\epsilon_{2}=\epsilon_{1}^{\prime\prime}/2\tilde{C}$. The proof is completed. ∎
+4.4 More general polynomials
 Now we prove Theorem 1.3 in the form of the following two theorems, which hold for more general forms.
 Theorem 1.3’.
-Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$ and $C_{1},C_{2},c>0$.
-There is a positive constant $\epsilon_{3}=\epsilon_{3}(s)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$.
-Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
+Let $d\geq 2$ and $\rho_{1}$, $\rho_{2}$ be real polynomials of degree $d$.
+Let $0<s<1$ and $C_{1},C_{2},c>0$.
+There is a positive constant $\epsilon_{3}=\epsilon_{3}(d,s)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$.
+Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
 Then
-$\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\geq n(s+\epsilon_{3})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{s}(1).$
+$\begin{split}&\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\rho_{1}(X)+\rho_{2}(Y))\right\}\\ &\qquad\qquad\qquad\geq n(s+\epsilon_{3})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\rho_{1},\rho_{2},s}(1).\end{split}$
 for all $n\in\mathbb{N}$.
 Proof.
-Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, by Lemma 2.30 it is also jointly $(s_{1},s_{2},C_{1}C_{2})$-Frostman.
-Applying Theorem 1.2’, there is some $\epsilon_{2}=\epsilon_{2}(s)$ such that
-$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}$
-$\qquad\qquad\geq n(s+\epsilon_{2})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{s}(1)$
-for all $n\in\mathbb{N}$. In view of $s_{1}+s_{2}\geq 2s$, Lemma 2.32(ii) shows that
-$H_{n}(X,Y)\geq\frac{H_{n}(X)+H_{n}(Y)}{2}+ns+O(\log_{+}C_{1}C_{2})+O(1).$
+Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, by Lemma 2.30 it is also jointly $(s_{1},s_{2},C_{1}C_{2})$-Frostman and marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman.
+Applying Theorem 1.2’, there is some $\epsilon_{2}=\epsilon_{2}(d,s)$ such that
+$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\rho_{1}(X)+\rho_{2}(Y))\right\}$
+$\qquad\qquad\geq n(s+\epsilon_{2})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\rho_{1},\rho_{2},s}(1)$
+for all $n\in\mathbb{N}$.
+Here the implied constant in the term $O(\log_{+}c)$ is independent of $s$, since the marginally Frostman constants do not involve $\max\{1,c\}$ as in (65).
+In view of $s_{1}+s_{2}\geq 2s$, Lemma 2.32(ii) shows that
+$H_{n}(X,Y)\geq\frac{H_{n}(X)+H_{n}(Y)}{2}+ns-\frac{1}{2}\log C_{1}C_{2}+O(1).$
 The conclusion then follows with $\epsilon_{3}=\epsilon_{2}/2$.
-$\log C^{\prime\prime}=\log C_{1}C_{2}+O_{\phi}(1)$.
 Theorem 1.3”.
-Suppose that $\phi(x,y)$ has the form (56), where $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $\xi_{11},\xi_{12},\xi_{13},\xi_{14}\in\mathbb{Q}^{\ast}$, $\xi_{11}\neq-\xi_{21}$, $\xi_{12}\neq-\xi_{22}$ and $\xi_{11}\xi_{22}-\xi_{12}\xi_{21}\neq 0$.
+Suppose that $\phi(x,y)$ has the form (1).
 Let $0<s<1$ and $C_{1},C_{2},c>0$.
-There is a positive constants $\epsilon_{4}=\epsilon_{4}(\phi,s)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$.
-Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
+There is a positive constants $\epsilon_{4}=\epsilon_{4}(d,\Xi,s)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$.
+Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
 Then
-$\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{4})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1).$
+$\begin{split}&\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\qquad\geq n(s+\epsilon_{4})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\phi,s}(1).\end{split}$
 for all $n\in\mathbb{N}$.
@@
 Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, Lemma 2.30 tells that it is both marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman and jointly $(s_{1},s_{2},C_{1}C_{2})$-Frostman.
-It follows from Lemma 2.32(i) and the fact $s_{1}+s_{2}\geq 2s$ that
-$\min\big{\{}H_{n}(X)+H_{n}(Y),\,H_{n}(X,Y)\big{\}}\geq 2ns-\log C_{1}C_{2}+O(1).$ (57)
+Then
+$H_{n}(X)+H_{n}(Y)\geq H_{n}(X,Y)\geq 2ns-\log C_{1}C_{2}+O(1).$ (74)
 In view of Lemma 2.32(ii), one has
-$\frac{H_{n}(X)+H_{n}(Y)}{2}\leq H_{n}(X,Y)-ns+O(\log_{+}C_{1}C_{2})+O(1).$
+$\frac{H_{n}(X)+H_{n}(Y)}{2}\leq H_{n}(X,Y)-ns+\frac{1}{2}\log C_{1}C_{2}+O(1).$
 It follows that
-$H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\leq H_{n}(X,Y)-2ns+O(\log_{+}C_{1}C_{2})+O_{\phi}(1).$ (58)
+$H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\leq H_{n}(X,Y)-2ns+\log C_{1}C_{2}+O(1).$ (75)
 Consider the random variable $(\xi_{11}X+\xi_{12}Y,\,\xi_{21}X+\xi_{22}Y)$.
 Its range is contained in $c^{\prime\prime},c^{\prime\prime}]^{2}$, with $c^{\prime\prime}=O_{\phi}(c)$.
-And Lemma 2.31 shows that it is jointly $(s_{1},s_{2},C^{\prime\prime})$-Frostman, with $\log C^{\prime\prime}=\log C_{1}C_{2}+O_{\phi}(1)$, since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman.
-By Lemma 1.2’, there is some $\epsilon_{2}=\epsilon_{2}(s)>0$ such that
+And Lemma 2.31, we assume without loss of generality that it is jointly $(s_{1},s_{2},C^{\prime\prime})$-Frostman, with $\log C^{\prime\prime}=\log C_{1}C_{2}+O_{\phi}(1)$, since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman.
+By Lemma 1.2’, there is some $\epsilon_{2}=\epsilon_{2}(d,s)>0$ such that
 $\max\left\{\frac{H_{n}(\xi_{11}X+\xi_{12}Y)+H_{n}(\xi_{21}X+\xi_{22}Y)}{2},\,H_{n}\big((\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big),\,H_{n}(\phi(X,Y))\right\}$
-$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\geq n(s+\epsilon_{2})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1)$ (59)
+$\qquad\qquad\qquad\qquad\qquad\geq n(s+\epsilon_{2})+O_{s}(\log_{+}C_{1}+\log_{+}C_{2})+O(\log_{+}c)+O_{\phi,s}(1)$ (76)
 for all $n\in\mathbb{N}$.
-Notice that $\xi_{11}\neq-\xi_{21}$ and $\xi_{12}\neq-\xi_{22}$. It follows from Theorem 2.23 that
+By Theorem 2.23, we have
 $H_{n}\big((\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big)-\frac{H_{n}(X)+H_{n}(Y)}{2}$
 $\qquad\leq d_{1}\left(H_{n}\big(X+Y\big)-\frac{H_{n}(X)+H_{n}(Y)}{2}\right)+d_{2}\big(H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\big)+O_{\phi}(1)$ (77)
-for some $d_{1}=d_{1}(\phi)>0$ and $d_{2}=d_{2}(\phi)\geq 0$. By (57), (58) and (60), one gets
+for some $d_{1}=d_{1}(\Xi)>0$ and $d_{2}=d_{2}(\Xi)\geq 0$. By (74), (75) and (77), one gets
 $\begin{split}&d_{1}\cdot H_{n}(X+Y)+2d_{2}\cdot\frac{1}{2}H_{n}(X,Y)\geq H_{n}\big((\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big)\\ &\qquad\qquad\qquad+(d_{1}+2d_{2}-1)ns+O(\log_{+}C_{1}C_{2})+O_{\phi}(1).\end{split}$ (78)
@@
 $\begin{split}&d_{1}^{\prime}\cdot H_{n}(X+Y)+2d_{2}^{\prime}\cdot\frac{1}{2}H_{n}(X,Y)\geq H_{n}(\xi_{11}X+\xi_{12}Y)\\ &\qquad\qquad\qquad+(d_{1}^{\prime}+2d_{2}^{\prime}-1)ns+O(\log_{+}C_{1}C_{2})+O_{\phi}(1).\end{split}$ (79)
-for some $d_{1}^{\prime}=d_{1}^{\prime}(\phi)>0$ and $d_{2}^{\prime}=d_{2}^{\prime}(\phi)\geq 0$. And also
+for some $d_{1}^{\prime}=d_{1}^{\prime}(\Xi)>0$ and $d_{2}^{\prime}=d_{2}^{\prime}(\Xi)\geq 0$. And also
 $\begin{split}&d_{1}^{\prime\prime}\cdot H_{n}(X+Y)+2d_{2}^{\prime\prime}\cdot\frac{1}{2}H_{n}(X,Y)\geq H_{n}(\xi_{21}X+\xi_{22}Y)\\ &\qquad\qquad\qquad+(d_{1}^{\prime\prime}+2d_{2}^{\prime\prime}-1)ns+O(\log_{+}C_{1}C_{2})+O_{\phi}(1).\end{split}$ (80)
-for some $d_{1}^{\prime\prime}=d_{1}^{\prime\prime}(\phi)>0$ and $d_{2}^{\prime\prime}=d_{2}^{\prime\prime}(\phi)\geq 0$.
+for some $d_{1}^{\prime\prime}=d_{1}^{\prime\prime}(\Xi)>0$ and $d_{2}^{\prime\prime}=d_{2}^{\prime\prime}(\Xi)\geq 0$.
 Now the theorem follows by combining (76) and (78)-(80), with $\epsilon_{4}=\epsilon_{2}/\max\{d_{1}+2d_{2},d_{1}^{\prime}+2d_{2}^{\prime},d_{1}^{\prime\prime}+2d_{2}^{\prime\prime}\}$. ∎
-In the following, Theorem 1.4 will be proved in the following form.
-Theorem 1.4’ Let $\phi(x,y)$ be in the form (56).
+For Theorem 1.4, the situation that $\phi(x,y)$ is diagonal has been proved by Lemma 4.8. The general case will be proved in the following form.
+Theorem 1.4’ Suppose that $\phi(x,y)$ has the form (1).
 Let $0<s<1$, $C_{1},C_{2},c>0$.
-There is a positive constant $\epsilon_{4}=\epsilon_{4}(\phi,s)$ such that the following statement holds: Let $X$ and $Y$ be independent real random variables with ranges contained in [c,c][-c,c].
-Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, where $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
+There is a positive constant $\epsilon_{4}=\epsilon_{4}(d,\Xi,s)$ such that the following statement holds: Let $X$ and $Y$ be two independent real random variables with ranges contained in [c,c][-c,c].
+Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, where $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
 Then
-$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{4})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1)$
+$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{4})+O_{d,s}(\log_{+}C_{1}C_{2})+O_{d}(\log_{+}c)+O_{\phi,s}(1)$
 for all $n\in\mathbb{N}$.
 Proof.
-When $\phi(x,y)$ is diagonal, the conclusion can be deduced from Lemma 4.1.
-In the following, we concentrate on the non-diagonal case.
-By Lemma 2.30, the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman.
-Applying Theorem 1.3”, there is some positive constant $\epsilon_{4}=\epsilon_{4}(\phi,s)$ such that
-$\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{4})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1).$
-Combining (13), one has
-$H_{n}(X,Y)\leq H_{n}(X+Y)+H_{n}(X)+O(1)\leq 2H_{n}(X+Y)+O(1).$
+By Lemma 2.30, the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman. Theorem 1.3” provide us with some positive constant $\epsilon_{4}=\epsilon_{4}(d,\Xi,s)$ such that
+$\begin{split}&\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\qquad\geq n(s+\epsilon_{4})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\phi,s}(1).\end{split}$
+Combining (8) and (9), one has
+$H_{n}(X,Y)\leq H_{n}(X)+H_{n}(X+Y)+O(1)\leq 2H_{n}(X+Y)+O(1).$
 The theorem then follows. ∎
 Remark 4.9.
-In the conclusions of Theorems 1.2’-1.4’, the implied constants in the terms $O_{s}(\log_{+}C)$ or $O_{s}(\log_{+}C_{1}C_{2})$ can be independent of $s$ if either of the following two situations happens: (1) we are only considering the case $s\gg 1$; (2) the range of random variables are bounded away from 0 by some absolute constant.
+In the conclusions of Lemma 4.8 and Theorems 1.2’-1.4’, the implied constants in the terms $O_{d,s}(\log_{+}C)$, $O_{d,s}(\log_{+}C_{1}C_{2})$ or $O_{d,s}(\log_{+}c)$ can be independent of $s$ if we are only considering the case $s\gg 1$.
 We preserve the dependence on $s$ in these terms for the uniformity with the case $0<s\leq 1/2$.
@@
 We will use the upcoming theorem as a substitution.
-The core of the proof will follow arguments from [24] by Shmerkin and Wang.
+The core of the proof will follow the arguments from [24] by Shmerkin and Wang.
 However, there are enough differences that we cannot merely apply a single theorem of theirs as a black box, so to address this case, we will introduce various notions and results from their paper, and explicitly mention where our arguments diverge.
-Recall the definition of $\varphi$ from (7).
 Theorem 5.1.
@@
 for all $n\in\mathbb{N}$.
+Here
+$\varphi(u)=\frac{u}{2}+\frac{\sqrt{4+u^{2}}-2}{2},\quad(0<u\leq 1).$
 To compare Theorems 2.38 and 5.1, one may take $W=(X,Y)$ in Theorem 2.38. The random variables $X$ and $Y$ are required to be independent in Theorem 5.1, but may be dependent in Theorem 2.38.
-The rest of the proof of Theorem 1.4 for $s\leq 1/2$ will follow the same line as in the case $s>1/2$.
+The rest of the proof of Theorem 1.4 for $s\leq 1/2$ will follow the same lines as in the case $s>1/2$.
 So we omit the details.
@@
 Let $\mu_{1}$, $\mu_{2}$, and $\nu$ be $(s,C)$-Frostman probability measures on $[0,1]$, where $0<s\leq 1/2$ and $C\geq 1$.
-Assume that $\text{dist}(\operatorname{spt}(\mu_{1}),\operatorname{spt}(\mu_{2}))\geq C_{d}C^{-1/s}$ with $C_{d}>0$.
+Assume that $\text{dist}(\operatorname{spt}(\mu_{1}),\operatorname{spt}(\mu_{2}))\geq\hat{C}C^{-1/s}$ with $\hat{C}>0$.
 Then for any $\zeta>0$, we have that
-$H_{n}(\Delta(\mu_{1}\times\nu,\mu_{2}\times\nu))\geq n(\varphi(2s)-\zeta)+O_{s,\zeta,C_{d}}(\log C)$
+$H_{n}(\Delta(\mu_{1}\times\nu,\mu_{2}\times\nu))\geq n(\varphi(2s)-\zeta)+O_{s,\zeta,\hat{C}}(\log C)$
 for all sufficiently large $n$.
-We emphasize that the term $O_{s,\zeta,C_{d}}(\log C)$ can not be replaced by $O_{s,\zeta,C,C_{d}}(1)$, because in the proof of the main theorem, we are working with infinitely many random variables satisfying (51), where the Frostman constants can be large compared to $n$, and are not uniform in $(x,y)$.
-So only an error of the form $\log C$ can be turned into entropy-terms as in (54).
-The main difficulty in this section, which prevents us from applying results in [24] directly, is due to this reason.
+We emphasize that the logarithmic dependence on $C$ in the error term is essential.
+Because in the proof of the main theorem, we are working with infinitely many random variables satisfying (70), where the Frostman constants can be large compared to $n$, and are not uniform in $(x,y)$.
+Only an error of the form $\log C$ can be turned into entropy-terms as in (71).
+The main difficulty in this section, which prevents us from applying the results in [24] directly, is due to this reason.
 Rather than repeating their arguments wholesale with minor modifications, we provide a detailed description of how to prove variants of their results tailored to our purposes, and combine these with a few calculations to prove Theorem 5.2.
@@
 Let $0<s<1$ and $C\geq 1$.
-Let $\mu$ be an $(s,C)$-Frostman probability measure with $\operatorname{spt}(\mu)\subset[0,1]$.
+Let $\mu$ be an $(s,C)$-Frostman probability measure with $\operatorname{spt}(\mu)\subseteq[0,1]$.
 For any $0<\gamma<1/3$, we can find a partition $\{U_{1},U_{2},U_{3}\}$ of $[0,1]$ such that the following hold:
@@
 $H_{n}(Z)\geq\sum\limits_{i,j=1}^{3}P(E_{ij})H_{n}\big(Z_{E_{ij}}\big)=\sum\limits_{i,j=1}^{3}\mu(U_{i})\mu(U_{j})H_{n}\big(Z_{E_{ij}}\big).$
-Since $X,Y,X^{\prime},Y^{\prime}$ are independent, one can verify that $X_{U_{i}},Y,X_{U_{j}}^{\prime},Y^{\prime}$ are also independent, and $Z_{E_{ij}}\equiv|(X_{U_{i}},Y)-(X_{U_{j}}^{\prime},Y^{\prime})|$.
+Since $X,Y,X^{\prime},Y^{\prime}$ are independent, the random variables $X_{U_{i}},Y,X_{U_{j}}^{\prime},Y^{\prime}$ are also independent, and $Z_{E_{ij}}\equiv|(X_{U_{i}},Y)-(X_{U_{j}}^{\prime},Y^{\prime})|$.
 Now apply Theorem 5.2 on $Z_{E_{12}}$ with $\zeta=\zeta(s,t)>0$ to be determined later. We deduce that
@@
 Below, we often use the notation $B_{r}$ to denote an arbitrary ball of radius $r.$
-Lemma 5.4.
-[Part of Lemma 2.3 [24]] Let $(a_{m})_{m=1}^{\infty}$ be a sequence of positive numbers such that $\sum_{m=1}^{\infty}a_{m}<\infty$.
+Lemma 5.4 (Part of Lemma 2.3 [24]).
+Let $(a_{m})_{m=1}^{\infty}$ be a sequence of positive numbers such that $\sum_{m=1}^{\infty}a_{m}<\infty$.
 Suppose $\mu\in\mathcal{P}({\mathbb{R}}^{d})$ is $(2^{-m},s,a_{m})$-robust for $m\in[m_{0},m_{1}]$, where $m_{0}<m_{1}\in\mathbb{N}\cup\{+\infty\}$.
@@
 Fix $c\in(0,1)$, $C\geq 1$, $0<\delta_{1}\leq 1$.
-Let $(Y,\nu)$ be a compact measure space, $U\subset\mathbb{R}^{d}$ be a bounded domain, and let $F_{y}:U\rightarrow\mathbb{R}^{k}$ be either the radial projection to $y$ or the distance to $y$.
+Let $(Y,\nu)$ be a compact measure space, $U\subseteq\mathbb{R}^{d}$ be a bounded domain, and let $F_{y}:U\rightarrow\mathbb{R}^{k}$ be either the radial projection to $y$ or the distance to $y$.
 Let $\mu\in\mathcal{P}(U)$ satisfy
@@
 Fix $\tau,\epsilon>0$.
-There are a set $Y_{0}\subset Y$ with $\nu(Y_{0})>\max(1-2c^{1/2},0)$, and for each $y\in Y_{0}$ there is a set $X_{y}\subset X^{\prime}$ with $\mu(X_{y})>\max\{1-2c^{1/2},0\}$, such that
+There are a set $Y_{0}\subseteq Y$ with $\nu(Y_{0})>\max(1-2c^{1/2},0)$, and for each $y\in Y_{0}$ there is a set $X_{y}\subseteq X^{\prime}$ with $\mu(X_{y})>\max\{1-2c^{1/2},0\}$, such that
 $F_{y}(\mu_{X_{y}})\text{ is }(2^{-m},\Sigma_{\tau}(D;s,d)-O(\epsilon^{1/2}),2^{-\eta m/2})\text{-robust, provided that},$
@@
 By definition, for a fixed $x$, we have that $(\ker D(\pi_{x})(y))^{\perp}=\pi_{x}(y)^{\perp}$ which implies $V_{x}(y)=\pi_{x}(y)^{\perp}$.
-Using the assumption that $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\gtrsim C_{\mu}^{-1}$ we obtain that $\pi_{x}(B(z,R))\cap\operatorname{spt}\mu_{2}$ will lie inside a tube of length at most $O(C_{\mu}R)$ (see the following figure).
+Using the assumption that $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\gg C_{\mu}^{-1}$ we obtain that $\pi_{x}(B(z,R))\cap\operatorname{spt}\mu_{2}$ will lie inside a tube of length at most $O(C_{\mu}R)$ (see the following figure).
 Figure 1: We can see that the tube containing $x$ and $B(z,R)$ has width $\leq C_{\mu}{\rm diam}(\operatorname{spt}\mu_{2})$.
@@
 $\Sigma_{\tau}(D_{s}-\eta;u-\eta,d)=\Sigma_{\tau}(D_{s};u-\eta,d)-\eta\geq s+\xi/2-\eta>s+2\eta.$
-This is enough to apply Theorem 5.9 to $Y,\nu,X,$ and the associated projections, $\pi,$ with $\delta_{1}=(KC\mu)^{-C^{\prime}\eta^{-1}}$, provided that $\eta$ is small enough in terms of $u$ and $\zeta.$ This will provide us a $c^{\prime}<2c^{1/2}$, a set $Y^{\prime},$ and $X=\{(y,X_{y}):y\in Y,x\in X_{y}\subset\mathbb{R}^{2}\}$ such that $\mu_{2}(Y^{\prime})\geq 1-c^{\prime}$, and $\mu_{1}(X_{y})>1-c^{\prime}$ for each $y\in Y^{\prime}$, the set $X$ is compact and most importantly:
+This is enough to apply Theorem 5.9 to $Y,\nu,X,$ and the associated projections, $\pi,$ with $\delta_{1}=(KC\mu)^{-C^{\prime}\eta^{-1}}$, provided that $\eta$ is small enough in terms of $u$ and $\zeta.$ This will provide us a $c^{\prime}<2c^{1/2}$, a set $Y^{\prime},$ and $X=\{(y,X_{y}):y\in Y,x\in X_{y}\subseteq\mathbb{R}^{2}\}$ such that $\mu_{2}(Y^{\prime})\geq 1-c^{\prime}$, and $\mu_{1}(X_{y})>1-c^{\prime}$ for each $y\in Y^{\prime}$, the set $X$ is compact and most importantly:
 $\rho_{y}=(\mu_{1}\circ\pi_{y}^{-1})_{X_{y}}\text{ is }(\delta^{\prime},s+\eta,\delta^{\prime\epsilon})\text{-robust for all }\delta^{\prime}<\delta_{0},$
@@
 Note that although this does depend on $c$, we choose $C_{1}$ large enough that it will not matter.
-This implies that $(\rho_{y})_{X_{y}^{\prime}}$ is $(s+\eta,\delta_{1}^{-1})$-Frostman, so the measure $(\mu_{1}\circ P_{v})_{X^{\prime}_{y}}$ is $(s+\eta,O(\delta_{1}^{-1}))$-Frostman for any vector $v\in S^{1}$ because $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\gtrsim C_{\mu}^{-1}$, so the intersection of any tube of width $r$ going through $x\in\operatorname{spt}\mu_{1}$ can be covered by $C_{\mu}$ disjoint arcs of length $r$ of the unit circle centered at $x$ (see the following figure).
+This implies that $(\rho_{y})_{X_{y}^{\prime}}$ is $(s+\eta,\delta_{1}^{-1})$-Frostman, so the measure $(\mu_{1}\circ P_{v})_{X^{\prime}_{y}}$ is $(s+\eta,O(\delta_{1}^{-1}))$-Frostman for any vector $v\in S^{1}$ because $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\gg C_{\mu}^{-1}$, so the intersection of any tube of width $r$ going through $x\in\operatorname{spt}\mu_{1}$ can be covered by $C_{\mu}$ disjoint arcs of length $r$ of the unit circle centered at $x$ (see the following figure).
 By using $\phi_{y}=1_{X^{\prime}_{y}}$ for each $y$, we obtain that $(\mu_{1},\mu_{2})$ has $(s+\eta,10C_{\mu}^{2}\delta^{-1},1-3c^{1/2},0)$-thin tubes.
@@
 In particular, we begin by applying Proposition 5.6 to the measures $\mu_{1}\times\nu$ and $\mu_{2}\times\nu,$ to see that they have $(s,2C,1,0)$-strong thin tubes.
-Using Lemma 5.12 with $C_{\mu}=C_{d}C^{\frac{1}{s}},$ we obtain $\eta$ and $C^{\prime}$ only depending on $s$ and $\zeta$ such that if $t<(\varphi(2s)-\zeta/M)$, for some constant $M$ only depending on $s$ and $\zeta,$ which we choose later, and for some $K$ also chosen later, if $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has (t,K,1c,0)(t,K,1-c,0)-strong thin tubes then it also has $(t+\eta,C^{\prime}\min\{c,K\}^{C^{\prime}},1-9c^{1/2},0)$-strong thin tubes.
+Using Lemma 5.12 with $C_{\mu}=\hat{C}C^{\frac{1}{s}},$ we obtain $\eta$ and $C^{\prime}$ only depending on $s$ and $\zeta$ such that if $t<(\varphi(2s)-\zeta/M)$, for some constant $M$ only depending on $s$ and $\zeta,$ which we choose later, and for some $K$ also chosen later, if $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has (t,K,1c,0)(t,K,1-c,0)-strong thin tubes then it also has $(t+\eta,C^{\prime}\min\{c,K\}^{C^{\prime}},1-9c^{1/2},0)$-strong thin tubes.
 Since $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has $(s,2C,1,0)$-strong thin tubes, then it also has $(s,3C,1-c_{0},0)$-strong thin tubes, where
@@
 Instead of applying an analogous result directly, we will follow their proof of Theorem 4.12 up until the statement about robustness that they obtain in the middle of their proof, then pivot to conclude our proof.
-Specifically, similar to the proof of Lemma 5.12, using the assumption that $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\gtrsim 1$ we obtain that $\pi_{x}(B(z,R))\cap\operatorname{spt}\mu_{2}$ will lie inside a tube of width at most $O(R).$ So by the definition of strong thin tubes, we obtain that there is a set $X$ with $(\mu_{1}\times\nu)(X)>1-(\zeta/M)^{2}$ such that for each $x\in X$, there is a set $Y_{x}$ that has $(\mu_{2}\times\nu)(Y_{x})>1-(\zeta/M)^{2}$ and
+Specifically, similar to the proof of Lemma 5.12, using the assumption that $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\gg 1$ we obtain that $\pi_{x}(B(z,R))\cap\operatorname{spt}\mu_{2}$ will lie inside a tube of width at most $O(R).$ So by the definition of strong thin tubes, we obtain that there is a set $X$ with $(\mu_{1}\times\nu)(X)>1-(\zeta/M)^{2}$ such that for each $x\in X$, there is a set $Y_{x}$ that has $(\mu_{2}\times\nu)(Y_{x})>1-(\zeta/M)^{2}$ and
 $((\mu_{2}\times\nu)_{Y_{x}}\circ V_{x})(B(z,r))\leq KC_{0}r^{\varphi(2s)-\zeta/2},$
@@
 $\Sigma_{\tau}(D_{\varphi(2s)-\zeta/2}(t)-\eta_{1});2s-\eta_{1},d)\geq\varphi(2s)-\zeta/2+\xi-2\eta_{1}>\varphi(2s)+\zeta/2.$
-This is enough to apply Theorem 5.9 to $Y,\nu,X,$ and the associated projections, $\pi,$ with $\delta_{1}=(KC\mu)^{-C^{\prime}\eta_{1}^{-1}}$, provided that $\eta_{1}$ is small enough in terms of $u$ and $\zeta.$ This will provide us a $c^{\prime}<2c^{1/2}$, a set $Y^{\prime},$ and $X=\{(y,X_{y}):y\in Y,x\in X_{y}\subset\mathbb{R}^{2}\}$ such that $\mu_{2}(Y^{\prime})\geq 1-c^{\prime}$, and $\mu_{1}(X_{y})>1-c^{\prime}$ for each $y\in Y^{\prime}$, the set $X$ is compact and most importantly:
+This is enough to apply Theorem 5.9 to $Y,\nu,X,$ and the associated projections, $\pi,$ with $\delta_{1}=(KC\mu)^{-C^{\prime}\eta_{1}^{-1}}$, provided that $\eta_{1}$ is small enough in terms of $u$ and $\zeta.$ This will provide us a $c^{\prime}<2c^{1/2}$, a set $Y^{\prime},$ and $X=\{(y,X_{y}):y\in Y,x\in X_{y}\subseteq\mathbb{R}^{2}\}$ such that $\mu_{2}(Y^{\prime})\geq 1-c^{\prime}$, and $\mu_{1}(X_{y})>1-c^{\prime}$ for each $y\in Y^{\prime}$, the set $X$ is compact and most importantly:
 $\rho_{y}=(\mu_{1}\circ\pi_{y}^{-1})_{X_{y}}\text{ is }(\delta^{\prime},s+\eta,\delta^{\prime\epsilon})\text{-robust for all }\delta^{\prime}<\delta_{0},$
 where $\epsilon>0$ depends on $\eta$ and $\tau$ (which, in turn, depend on $u$ and $\zeta$), and $\delta_{0}=C_{1}^{-1}(K)^{-C_{1}},$ where $C_{1}$ is a constant also depending on $u$ and $\zeta.$
-The analogous statement we get is that there is a $c^{\prime}<2(\zeta/M)$, a set $Y$ and $X=\{(y,x):y\in Y,x\in X_{y}\subset\mathbb{R}^{2}\}$ such that $(\mu_{2}\times\nu)(Y)\geq 1-c^{\prime}$, $(\mu_{1}\times\nu)(X_{y})>1-c^{\prime}$ for each $y\in Y$, the set $X$ is compact and most importantly:
+The analogous statement we get is that there is a $c^{\prime}<2(\zeta/M)$, a set $Y$ and $X=\{(y,x):y\in Y,x\in X_{y}\subseteq\mathbb{R}^{2}\}$ such that $(\mu_{2}\times\nu)(Y)\geq 1-c^{\prime}$, $(\mu_{1}\times\nu)(X_{y})>1-c^{\prime}$ for each $y\in Y$, the set $X$ is compact and most importantly:
 $\rho_{y}=((\mu_{1}\times\nu)_{X_{y}}\circ d_{y}^{-1})\text{ is }(\delta^{\prime},\varphi(2s)-\zeta/2,\delta^{\prime\epsilon})\text{-robust for all }\delta^{\prime}<\delta_{0},$
@@
 Define the map $f(x,y,z,t)=\sqrt{(x-z)^{2}+(y-t)^{2}}$ and the measure $\kappa$ supported in $\mathbb{R}^{4}$ as $\mu_{1}\times\nu\times\mu_{2}\times\nu$.
-Let the set $X^{\prime}\subset\mathbb{R}^{4}$ be $X^{\prime}:=\{(y,x):y\in Y,x\in X_{y},d(x,y)\in X^{\prime}_{y}\}$ we have that
+Let the set $X^{\prime}\subseteq\mathbb{R}^{4}$ be $X^{\prime}:=\{(y,x):y\in Y,x\in X_{y},d(x,y)\in X^{\prime}_{y}\}$ we have that
 $\kappa(X^{\prime})=\int_{Y}(\mu_{1}\times\nu)(d_{y}^{-1}(X^{\prime}_{y})\cap X_{y})d(\mu_{2}\times\nu)(y)\geq\int_{Y}(\mu_{1}\times\nu)_{X_{y}}\rho_{y}(X^{\prime}_{y})d(\mu_{2}\times\nu)(y)\geq(1-c^{\prime})^{3}.$
@@
 Proof of Theorem 1.5.
-Write $\phi(x,y)=\alpha x^{2}+\beta y^{2}$. Let $\epsilon=\epsilon(s)$ to be determined later.
+Write $\phi(x,y)=\rho_{1}(x)+\rho_{2}(y)$. Let $\epsilon=\epsilon(s)$ to be determined later.
 Note that $(x,y)\mapsto x+y$ and $(x,y)\mapsto\phi(x,y)$ are both $O_{\phi}(1)$-Lipchitz on $0,1]^{2}$.
-We can choose a small number $\eta=\eta(\phi)$ with $0<\eta<1/10$ such that $A_{\eta\delta}+_{G}A_{\eta\delta}$ and $\phi_{G}(A_{\eta\delta},A_{\eta\delta})$ are contained in the $(A+_{G}A)_{\delta/10}$ and $(\phi_{G}(A,A))_{\delta/10}$, respectively, for any $\delta>0$.
+We can choose a small number $\eta=\eta(\phi)$ with $0<\eta<1/10$ such that $A_{\eta\delta}+_{G}A_{\eta\delta}$ and $\phi_{G}(A_{\eta\delta},A_{\eta\delta})$ are contained in $(A+_{G}A)_{\delta/10}$ and $(\phi_{G}(A,A))_{\delta/10}$, respectively, for any $\delta>0$.
 Then
-$\begin{split}\mathcal{E}_{\delta}(A+_{G}A)&\gg\mathcal{E}_{\delta}(A_{\eta\delta}+_{G}A_{\eta\delta}),\\ \mathcal{E}_{\delta}(\phi_{G}(A,A))&\gg\mathcal{E}_{\delta}(\phi_{G}(A_{\eta\delta},A_{\eta\delta}))).\end{split}$ (66)
+$\begin{split}\mathcal{E}_{\delta}(A+_{G}A)&\gg\mathcal{E}_{\delta}(A_{\eta\delta}+_{G}A_{\eta\delta}),\\ \mathcal{E}_{\delta}(\phi_{G}(A,A))&\gg\mathcal{E}_{\delta}(\phi_{G}(A_{\eta\delta},A_{\eta\delta})).\end{split}$ (83)
 Moreover, since $A$ is $\delta$-separated, we have
@@
 When $\delta<r_{2}\leq 1$, the concentration conditions gives
-$\mathcal{L}^{1}(B(y,r_{2})\cap A_{\eta\delta})\ll Cr_{2}^{s}(\#A)\delta\ll cr_{2}^{s}\delta^{1-s}.$
-Calculation reveals that
+$\mathcal{L}^{1}(B(y,r_{2})\cap A_{\eta\delta})\ll Cr_{2}^{s}(\#A)\cdot\eta\delta\ll_{\phi}Cr_{2}^{s}\delta^{1-s}.$
+The other cases can be calculated in a similar way. Computation reveals that
 $P(X\in B(x,r_{1}),Y\in B(y,r_{2}))\ll_{\phi}\begin{cases}r_{1}\cdot r_{2}\cdot\delta^{2s-2-\epsilon},\quad&\text{if }0<r_{1},r_{2}\leq\delta,\\ r_{1}\cdot Cr_{2}^{s}\delta^{1-s}\cdot\delta^{2s-2-\epsilon}&\text{if }0<r_{1}\leq\delta<r_{2}\leq 1,\\ r_{1}\cdot\delta^{1-s}\cdot\delta^{2s-2-\epsilon},\quad&\text{if }0<r_{1}\leq\delta,\,r_{2}>1,\\ Cr_{1}^{s}\delta^{1-s}\cdot Cr_{2}^{s}\delta^{1-s}\cdot\delta^{2s-2-\epsilon}&\text{if }\delta<r_{1},r_{2}\leq 1\\ Cr_{1}^{s}\delta^{1-s}\cdot\delta^{1-s}\cdot\delta^{2s-2-\epsilon},\quad&\text{if }\delta<r_{1}\leq 1<r_{2},\\ 1,\quad&\text{if }r_{1},r_{2}>1,\end{cases}$
@@
 So $(X,Y)$ is jointly $(s,s,O_{\phi}(C^{2}\delta^{-\epsilon}))$-Frostman.
-By Theorem 1.2’, there exists an $\epsilon^{\prime}=\epsilon^{\prime}(s)$ such that
-$\begin{split}&\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\geq n(s+\epsilon^{\prime})+O_{s}(\log_{+}C)+O_{s}(\epsilon\log(1/\delta))+O_{\phi,s}(1).\end{split}$ (69)
+By Theorem 1.2’ (see Section 4.3), there exists an $\epsilon^{\prime}=\epsilon^{\prime}(s)$ such that
+$\begin{split}&\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\geq n(s+\epsilon^{\prime})+O_{d,s}(\log C)+O_{d,s}(\epsilon\log(1/\delta))+O_{\phi,s}(1)\end{split}$ (86)
 for all $n\in\mathbb{N}$. In particular, we take $n$ such that $2^{-n-1}<\delta\leq 2^{-n}$.
@@
 $\max\{\mathcal{E}_{\delta}(A),\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\}$
-$\qquad\gg 2^{n(s+\epsilon^{\prime})+O_{s}(\log_{+}C)+O_{s}(\epsilon\log(1/\delta))+O_{\phi,s}(1))}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-(s+\epsilon^{\prime}-O_{s}(\epsilon))}.$ (72)
-Now we can take $\epsilon=\epsilon(s)>0$ to be sufficiently small so that $\epsilon^{\prime}-O_{s}(\epsilon)\geq\epsilon$ on the right-hand side of (72). Then
-$\max\{\mathcal{E}_{\delta}(A),\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-(s+\epsilon)}$ (73)
+$\qquad\gg 2^{n(s+\epsilon^{\prime})+O_{d,s}(\log C)+O_{d,s}(\epsilon\log(1/\delta))+O_{\phi,s}(1)}\gg_{\phi,s}C^{-O_{d,s}(1)}\delta^{-(s+\epsilon^{\prime}-O_{d,s}(\epsilon))}.$ (89)
+Now we can take $\epsilon=\epsilon(d,s)>0$ to be sufficiently small so that the term $\epsilon^{\prime}-O_{d,s}(\epsilon)\geq\epsilon$ on the right-hand side of (89). Then
+$\max\{\mathcal{E}_{\delta}(A),\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\}\gg_{\phi,s}C^{-O_{d,s}(1)}\delta^{-(s+\epsilon)}$ (90)
 for any $\delta>0$.
 Finally, by (84) one has $\mathcal{E}_{\delta}(A)\ll\delta^{-s}$. This gives
-$\max\{\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A)$
-whenever $\delta$ is sufficiently small depending on $\phi,s$ and $C$. The proof is completed. ∎
+$\max\{\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\}\gg_{\phi,s}C^{-O_{d,s}(1)}\delta^{-\epsilon}(\#A)$
+whenever $\delta$ is sufficiently small depending on $\phi$ and $s$ and $C$. The proof is completed. ∎
 Proof of Theorem 1.6.
@@
 Let $(X,Y)$ be a random variable that has uniform distribution on $G_{\eta\delta}$. Let $E$ be any Borel set and $B(y,r)$ be any ball. We have
-$P(X\in E)\gg_{\phi}\mathcal{L}^{1}(X\cap A_{\eta\delta})\cdot\delta^{1-s+\epsilon}\cdot\delta^{2s-2}.$
-And
-$P(X\in E,Y\in B(y,r))\ll_{\phi}\begin{cases}\mathcal{L}^{1}(X\cap A_{\eta\delta})\cdot r\cdot\delta^{2s-2-\epsilon},\quad&\text{if }0<r\leq\delta,\\ \mathcal{L}^{1}(X\cap A_{\eta\delta})\cdot Cr^{s}\delta^{1-s}\cdot\delta^{2s-2-\epsilon}&\text{if }\delta<r\leq 1,\\ \mathcal{L}^{1}(X\cap A_{\eta\delta})\cdot\delta^{1-s}\cdot\delta^{2s-2-\epsilon},&\text{if }r>1.\end{cases}$
+$P(X\in E)\gg_{\phi}\mathcal{L}^{1}(E\cap A_{\eta\delta})\cdot\delta^{1-s+\epsilon}\cdot\delta^{2s-2},$
+where the second term on the right-hand side is due to (3). Moreover,
+$P\big(X\in E,Y\in B(y,r)\big)\ll_{\phi}\begin{cases}\mathcal{L}^{1}(E\cap A_{\eta\delta})\cdot r\cdot\delta^{2s-2-\epsilon},\quad&\text{if }0<r\leq\delta,\\ \mathcal{L}^{1}(E\cap A_{\eta\delta})\cdot Cr^{s}\delta^{1-s}\cdot\delta^{2s-2-\epsilon}&\text{if }\delta<r\leq 1,\\ \mathcal{L}^{1}(E\cap A_{\eta\delta})\cdot\delta^{1-s}\cdot\delta^{2s-2-\epsilon},&\text{if }r>1.\end{cases}$
 In all the cases,
@@
 Now (86) can be replaced by
-$\begin{split}&\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\geq n(s+\epsilon^{\prime})+O_{s}(\log C)+O_{s}(\epsilon\log(1/\delta))+O_{\phi,s}(1)\end{split}$
-for some $\epsilon^{\prime}=\epsilon^{\prime}(\phi,s)>0$, which results from Theorem 1.3’. Moreover, we substitute
+$\begin{split}&\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\geq n(s+\epsilon^{\prime})+O_{d,s}(\log C)+O_{d,s}(\epsilon\log(1/\delta))+O_{\phi,s}(1)\end{split}$
+for some $\epsilon^{\prime}=\epsilon^{\prime}(\phi,s)>0$, which results from Theorem 1.3’ (see Section 4.4). Moreover, we substitute
 $\log\#range_{+}(D_{n}(X,Y))\geq H_{n}(X,Y)$
@@
 for the first formula in (87) and (88), respectively. Hence (90) can be replaced by
-$\max\left\{\mathcal{E}_{\delta}(G)^{1/2},\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\right\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-(s+\epsilon)}.$
+$\max\left\{\mathcal{E}_{\delta}(G)^{1/2},\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\right\}\gg_{\phi,s}C^{-O_{d,s}(1)}\delta^{-(s+\epsilon)}.$
 Finally, since $\mathcal{E}_{\delta}(G)\ll\delta^{-2s}$, we conclude that
-$\max\left\{\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\right\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-(s+\epsilon)}$
+$\max\left\{\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\right\}\gg_{\phi,s}C^{-O_{d,s}(1)}\delta^{-(s+\epsilon)}$
 whenever $\delta$ is sufficiently small depending on $\phi$, $s$ and $C$. ∎
@@
 Combining the non-concentration condition, one deduces that
-$P(X\in B(x,r))\ll_{\phi}\begin{cases}r\cdot\delta^{s-1-\epsilon},\quad&\text{if }0<r\leq\delta,\\ Cr^{s}\delta^{1-s}\cdot\delta^{s-1-\epsilon},\quad&\text{if }\delta<r\leq 1,\\ 1,\quad&\text{if }r>1,\end{cases}$
+$P(X\in B(x,r))\ll_{\phi}\begin{cases}r\cdot\delta^{s-1-\epsilon},\quad&\text{if }0<r\leq\delta,\\ Cr^{s}\delta^{1-s}\cdot\delta^{s-1-\epsilon},\quad&\text{if }\delta<r\leq 1,\\ 1,\quad&\text{if }r>1\end{cases}$
 for any ball $B(x,r)$. In all the cases, the right-hand side can be bounded by $C\delta^{-\epsilon}r^{s}$. So $X$ is $(s,O_{\phi}(C\delta^{-\epsilon}))$-Frostman.
-As a replacement of (69), Theorem 1.4’ shows that
+As a replacement of (86), Theorem 1.4’ (see Section 4.4) shows that
 $\max\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\}\geq n(s+\epsilon^{\prime})+O_{s}(\log C)+O_{\phi,s}(1)$
 for some $\epsilon^{\prime}=\epsilon^{\prime}(s)>0$ and for all $n\in\mathbb{N}$. And the similar conclusion as in (90) appears that
-$\max\{\mathcal{E}_{\delta}(B+B),\,\mathcal{E}_{\delta}(\phi(B,B))\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A)$
+$\max\{\mathcal{E}_{\delta}(B+B),\,\mathcal{E}_{\delta}(\phi(B,B))\}\gg_{\phi,s}C^{-O_{d,s}(1)}\delta^{-\epsilon}(\#A)$
 for all $\delta>0$. ∎
 Remark 6.1.
-In the conclusions of Theorems 1.5-1.7, the term $C^{-O_{s}(1)}$ can be replaced by $C^{-O(1)}$ if either of the following two situations happens: (1) we are only considering the case $s\gg 1$; (2) the condition $A\subseteq[0,1]$ is replaced by $A\subseteq[1,2]$.
+In the conclusions of Theorems 1.5-1.7, the term $C^{-O_{s}(1)}$ can be replaced by $C^{-O(1)}$ if we are only considering the situation $s\gg 1$.
 7 Examples
-In this section, we show details of some examples. Recall that $\mathcal{L}^{d}$ is the $d$-dimensional Here measure. Sometimes, we will write $\iota(x)=-x\log x$ for convenience.
+In this section, we show details of some examples. Recall that $\mathcal{L}^{d}$ is the $d$-dimensional Lebesgue measure. Sometimes, we will write $\iota(x)=-x\log x$ for convenience.
 Example 7.1.
@@
 $A=\bigcup\limits_{j=0}^{2^{m}-1}\left[j2^{-m},\,j2^{-m}+2^{-m-l}\right).$
-Then $\mathcal{L}^{1}(A)=2^{-l}$. We see that both $X$ and $Y$ have Note that $X$ has uniform distribution on $A$, i.e.,
-$P(X\in E)=\frac{\mathcal{L}^{1}(E\cap A)}{\mathcal{L}^{1}(A)}$
-for any Borel set $E$. In the following, we will proof several results. The $o(1)$-terms are in the situation that $n\rightarrow\infty$.
-Claim 1. The random variable $(X,Y)$ is marginally $\left(s_{1},O(1);s_{2},O(1)\right)$-Frostman whenever $0\leq s_{1},s_{2}\leq\frac{m}{l+m}$.
+Then $\mathcal{L}^{1}(A)=2^{-l}$.
+We see that both $X$ and $Y$ have uniform distribution on $A$.
+In the following, we will prove several results.
+The $o(1)$-terms are in the situation that $n\rightarrow\infty$.
+Claim 1. The random variable $(X,Y)$ is marginally $\left(s_{1},O(1);s_{2},O(1)\right)$-Frostman, whenever $0\leq s_{1},s_{2}\leq\frac{m}{l+m}$.
 Claim 2. We have
@@
 Proof of Claim 1.
-Let us consider the probability that $X$ takes a value in any ball $B(x,r)$. Calculation reveals that
+Consider the probability that $X$ takes a value in any ball $B(x,r)$. Calculation reveals that
 $P\big(X\in B(x,r)\big)\ll\begin{cases}r2^{l},\quad&\text{if }0<r\leq 2^{-l-m},\\ 2^{-m},\quad&\text{if }2^{-l-m}<r\leq 2^{-m},\\ r\quad&\text{if }2^{-m}\leq r<1,\\ 1&\text{if }r\geq 1.\end{cases}$
-One can verify that all the terms on the right-hand side of above formula can be bounded by $r^{\frac{m}{l+m}}$.
+Calculation shows that all the terms on the right-hand side of above formula can be bounded by $r^{\frac{m}{l+m}}$.
 Indeed, the random variable $X$ is $(s,O(1))$-Frostman for any $0\leq s\leq\frac{m}{l+m}$.
@@
 By Lemma 2.4 (ii) and independence, one sees that
-$H(X^{\prime}+Y^{\prime})\leq\log\#\,range_{+}(X^{\prime}+Y^{\prime})=m+O(1),$
-and
-$H(X^{\prime}+Y^{\prime})\geq H(X^{\prime})=m.$
+$m=H(X^{\prime})\leq H(X^{\prime}+Y^{\prime})\leq\log\#\,range_{+}(X^{\prime}+Y^{\prime})=m+O(1).$
 Therefore, $H_{n}(X^{\prime}+Y^{\prime})=m+O(1)$.
@@
 After excluding the zero terms, one can obtain that
-$H_{n}(X^{\prime 2}+Y^{\prime 2})=\sum\limits_{i=0}^{2^{n+1}}\iota\Big{(}P(D_{n}(X^{\prime 2}+Y^{\prime 2})\quad=i2^{-n})\Big{)}$
-$=\sum\limits_{k=0}^{2^{2m+1}}\iota\Big{(}P(X^{\prime 2}+Y^{\prime 2}=k2^{-2m})\Big{)}=\sum\limits_{k\in\mathcal{K}}\iota\big{(}2^{-2m}T(k)\big{)},$
+$H_{n}(X^{\prime 2}+Y^{\prime 2})=\sum\limits_{i=0}^{2^{n+1}}\iota\Big(P(D_{n}(X^{\prime 2}+Y^{\prime 2})=i2^{-n})\Big)$
+$\qquad=\sum\limits_{k=0}^{2^{2m+1}}\iota\Big(P(X^{\prime 2}+Y^{\prime 2}=k2^{-2m})\Big)=\sum\limits_{k\in\mathcal{K}}\iota\big(2^{-2m}T(k)\big),$
 where $\mathcal{K}$ denotes the set consisting of integers $k$ such that $k=j_{1}^{2}+j_{2}^{2}$ for some $0\leq j_{1},j_{2}\leq 2^{m}-1$, and
@@
 $\log T(k)\ll\frac{\log k}{\log\log k}\ll\frac{m}{\log m}.$
-On the other hand, one has $T(k)\geq 1$ for all $k\in\mathcal{K}$. So
+On the other hand, one has $T(k)\geq 1$ for all $k\in\mathcal{K}$. Therefore,
 $\log\left(2^{-2m}T(k)\right)=-2m+O\Big(\frac{m}{\log m}\Big).$
@@
 $\max\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\}=(1+o(1))ns.$
-So, the requirement that $\phi$ is non-degenerate in Theorem 1.4 is essential.
+Hence, the requirement that $\phi$ is non-degenerate in Theorem 1.4 is essential.
 Example 7.4.
@@
 $A=\bigcup\limits_{j=0}^{2^{m}-1}\left[1+j2^{-m},\,1+j2^{-m}+2^{-2m}\right),$
-and $Y$ has uniform distribution on $0,2^{-2m}]$. One can verify that $X$ is $(1/2,O(1))$-Frostman, and $Y$ is both $(1/2,O(2^{m}))$-Frostman and $(0,1)$-Frostman. We take $n=2m$.
+and $Y$ has uniform distribution on $0,2^{-2m}]$. Verification gives $X$ is $(1/2,O(1))$-Frostman, and $Y$ is both $(1/2,O(2^{m}))$-Frostman and $(0,1)$-Frostman. We take $n=2m$.
 Claim. We have $\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}=n/2+O(1)$.
@@
 Furthermore, we can not substitute $\frac{1}{2}H(X,Y)=m/2$ for $(H_{n}(X)+H_{n}(Y))/2$ as in Theorem 1.3.
-Furthermore, for the random variable $W$, its Frostman constant $C$ can grow with $m$.
+In addition, for the random variable $W$, its Frostman constant $C$ can grow with $m$.
 If we consider $n=2m$ in Theorem 2.38, if $1<s<3/2$, then one has $H_{n}(|W_{1}-W_{2}|)=(1+o(1))n/2<nt$ for some suitably chosen $t$ with s/2<t<s1/2s/2<t<s-1/2.
@@
 and the other cases with $r_{2}<r_{1}$ can be obtained by symmetry.
-One can verify that all the terms on the right-hand side of above formula can be bounded by $r_{1}^{1/4}r_{2}^{1/4}$.
+Calculation reveals that all the terms on the right-hand side of above formula can be bounded by $r_{1}^{1/4}r_{2}^{1/4}$.
 So $(X,Y)$ is jointly $(1/4,1/4,O(1))$-Frostman.
@@
 $P\big(X\in B(x,r)\big)\ll\begin{cases}r^{2}2^{3m},\quad&\text{if }0<r\leq 2^{-2m+1/2},\\ 2^{-m},\quad&\text{if }2^{-2m+1/2}<r\leq 2^{-m+1/2},\\ r\quad&\text{if }2^{-m+1/2}\leq r<1,\\ 1&\text{if }r\geq 1.\end{cases}$
-One can verify that all the terms on the right-hand side of above formula can be bounded by $2^{m(2s-1)}r^{s}$ for any $1/2\leq s\leq 2$.
+One can check that all the terms on the right-hand side of above formula can be bounded by $2^{m(2s-1)}r^{s}$ for any $1/2\leq s\leq 2$.
 The random variable $(X,Y)$ is $\left(s,O(2^{m(2s-1))}\right)$-Frostman.
@@
 Proof of Claim 3.
-For any pair of balls $B(x,r_{1})$ and $B(y,r_{2})$, calculation reveal that
-$P(X\in B(x,y_{1}),y\in B(y,r_{2}))$
+For any pair of balls $B(x,r_{1})$ and $B(y,r_{2})$, calculations reveal that
+$P(X\in B(x,y_{1}),y\in B(y,r_{2}))\ll M(r_{1})M(r_{2})\cdot 2^{4m}$
+where
 […]
-The rest cases with $r_{1}>r_{2}$ can be obtained by symmetry. In all cases, we have
-$P(X\in B(x,y_{1}),y\in B(y,r_{2}))\ll r_{1}^{1/2}r_{2}{1/2}.$
+In all cases, we have
+$P(X\in B(x,y_{1}),y\in B(y,r_{2}))\ll r_{1}^{1/2}r_{2}^{1/2}.$
 The proof is completed. ∎
@@
 $=P^{\prime}(W^{\prime}\in L_{2},h(Z^{\prime})=z)P^{\prime}(Y^{\prime\prime}\in L_{1}|h(Z^{\prime})=z).$
-So $P^{\prime}(Y^{\prime\prime}\in L_{1}|D^{\prime}\in L_{2},h(Z^{\prime})=z)=P^{\prime}(Y^{\prime\prime}\in L_{1}|h(Z^{\prime})=z)$, i.e.
+It follows that $P^{\prime}(Y^{\prime\prime}\in L_{1}|D^{\prime}\in L_{2},h(Z^{\prime})=z)=P^{\prime}(Y^{\prime\prime}\in L_{1}|h(Z^{\prime})=z)$, i.e.
 $(Y^{\prime\prime}|W^{\prime},h(Z^{\prime}))\equiv(Y^{\prime\prime}|h(Z^{\prime}))$.
@@
 Acknowledgements
-Thang Pham and Steven Senger would like to thank the Vietnam Institute for Advanced Study in Mathematics (VIASM) and the Erdős Center for the hospitality and for the excellent working condition.
+Thang Pham and Steven Senger would like to thank the Vietnam Institute for Advanced Study in Mathematics (VIASM) and the Erdős Center for the hospitality and for the excellent working conditions.
 We would like to thank Tuomas Orponen for his comments on the first version of this paper and for telling us about the references [1, 22].

### v5

@@
-Frostman random variables, entropy inequalities, and applications
+arXiv:2507.15196v5[math.CA] 02 Mar 2026
 Abstract
@@
 Fix $0<s<1$, and let $(X,Y)$ be a bivariate real random variable with bounded support, whose distribution satisfies a Frostman condition of dimension $s$.
-Let $\phi(x,y)$ be a polynomial obtained from a diagonal polynomial $\rho_{1}(x)+\rho_{2}(y)\in\mathbb{R}[x,y]$ of degree $d\geq 2$ by applying an invertible rational linear change of variables in $(x,y)$.
-We show that there exists $\epsilon=\epsilon(\phi,s)>0$ such that
+Let $\phi(x,y)$ be a polynomial obtained from a diagonal polynomial $\rho_{1}(x)+\rho_{2}(y)\in\mathbb{R}[x,y]$ of degree $d\geq 2$ by applying a change of variables $\Xi\in GL_{2}(\mathbb{Q})$ in $(x,y)$.
+We show that there exists $\epsilon=\epsilon(d,\Xi,s)>0$ such that
 $\max\{H_{n}(X+Y),H_{n}(\phi(X,Y))\}\geq n(s+\epsilon)$
 for all sufficiently large $n$, where the precise assumptions on $(X,Y)$ depend on the Frostman level.
-The proof introduces a novel multi‑step entropy framework, combining the submodularity formula, the discretized entropy Balog-Szemerédi-Gowers theorem, and state-of-the-art results on the Falconer distance problem, to reduce general forms to a diagonal core case.
-As an application, we obtain discretized sum-product type estimates.
+The proof introduces a novel multi‑step entropy framework, combining the state-of-the-art results on the Falconer distance problem, a discretized entropy Balog–Szemerédi–Gowers mechanism, and new entropy inequalities adapted to dependent variables, to reduce general polynomials of arbitrary degree to a diagonal quadratic case.
+As applications, we obtain innovative discretized sum-product type estimates along dense graphs.
 In particular, for a $\delta$-separated set $A\subseteq[0,1]$ of cardinality $\delta^{-s}$, satisfying certain non-concentration conditions, and a dense subset $G\subseteq A\times A$, there exists $\epsilon=\epsilon(s,\phi)>0$ such that
@@
 Our study is motivated by recent applications of entropy methods to discretized sum-product phenomena [19] due to Máthé and O’Regan, and to the recent solution of the polynomial Freiman-Ruzsa conjecture [8, 9] due to Gowers, Green, Manners and Tao.
-More precisely, we are interested in bounding the following quantity $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ from below, where $\phi(x,y)=a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ is a non-degenerate rational quadratic form, i.e.
-$a_{2}^{2}-4a_{1}a_{3}\neq 0$, and $(X,Y)$ is a bivariate Frostman random variable with bounded range.
-For $\phi(x,y)=x\cdot y$, we recall the following theorem due to Máthé and O’Regan in [12].
-Let $X$ and $Y$ be independent $(s,C)$-Frostman random variables whose ranges are contained in $c_{1},c_{1}]$, with $0<s<1$, and $C,c_{1}\geq 1$.
-Then there exists a positive constant $\epsilon=\epsilon(s)$ and an integer $N=N(s,C,c_{1})$ such that for all $n>N,$ we have
+More precisely, we seek a quantitative lower bound for
+$\max\bigl\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\bigr\},$
+where $(X,Y)$ is a bivariate random variable satisfying one of the Frostman-type nonconcentration hypotheses introduced in Definition 1.1, and $\phi:\mathbb{R}^{2}\to\mathbb{R}$ belongs to a broad class of polynomial forms of degree $d\geq 2$ (described below) obtained from a diagonal expression $\rho_{1}(x)+\rho_{2}(y)$ by composing with a rational linear change of variables.
+In particular, this class contains the non-degenerate rational quadratic forms.
+For $\phi(x,y)=xy$, we recall the following theorem due to Máthé and O’Regan in [19].
+Let $X$ and $Y$ be independent $(s,C)$-Frostman random variables whose ranges are contained in [c,c][-c,c], with $0<s<1$, and $C,c\geq 1$.
+Then there exists a positive constant $\epsilon=\epsilon(s)$ and an integer $N=N(s,C,c)$ such that for all $n>N,$ we have
 $\max\left\{H_{n}(X+Y),\,H_{n}(X\cdot Y)\right\}\geq n(s+\epsilon).$
@@
 for any ball $B(x,r)$ and Borel set $E$.
-Moreover, we say that the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, if $X$ is conditionally $(s_{1},C_{1})$-Frostman given $Y$, and $Y$ is conditionally $(s_{2},C_{2})$-Frostman given $X$.
-These definitions are not equivalent. At a common exponent level $(s_{1},s_{2})$, these conditions form a hierarchy:
+Moreover, we say that the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman if $X$ is conditionally $(s_{1},C_{1})$-Frostman given $Y$, and $Y$ is conditionally $(s_{2},C_{2})$-Frostman given $X$.
+While the usual $(s,C)$-Frostman condition for measures on $\mathbb{R}$ is standard, the three bivariate variants introduced in Definition 1.1 (joint/marginal/conditional) and the way they are used as a hierarchy in an entropy-expansion framework are, to the best of our knowledge, not part of the standard entropy sum-product toolkit.
+They are introduced here to separate non-concentration from dependence: different levels of the hierarchy genuinely lead to different entropy inequalities and different applications.
+At a common exponent level $(s_{1},s_{2})$, these conditions form the following chain:
 $\boxed{\text{\parbox{116.65646pt}{\centering independent\\ and marginally Frostman\@add@centering}}}\;\Rightarrow\;\boxed{\text{conditionally Frostman}}\;\Rightarrow\;\boxed{\text{jointly Frostman}}\;\Rightarrow\;\boxed{\text{marginally Frostman}}$
@@
   • a dependent pair that is conditionally $(1/4,O(1);1/4,O(1))$-Frostman, but $(1/2,1/2;O(1))$-jointly Frostman (Example 8.6).
-Remarkably, different levels of this hierarchy yield entropy inequalities with different strengths and applications.
+1.1 Main results on entropy estimates
 The results in this paper demonstrate how the Frostman hierarchy affects entropy bounds.
+[…]
 Theorem 1.2 (Jointly Frostman, diagonal polynomials).
@@
 Let $0<s<1$ and $C,c>0$.
-There exist a positive constant $\epsilon=\epsilon(d,s)$ and an integer $N=N(\rho_{1},\rho_{2},s,C,c)$ such that the following statement holds.
-For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
-$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}\big(\rho_{1}(X)+\rho_{2}(Y)\big)\right\}\geq n(s+\epsilon)$
+There exists a positive constant $\epsilon=\epsilon(d,s)$ and an integer $N=N(\rho_{1},\rho_{2},s,C,c)$ such that the following statement holds.
+For any bivariate real random variable $(X,Y)$ that is jointly $(s,s,C)$-Frostman and has range contained in $c,c]^{2}$, we have
+$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}\big(\rho_{1}(X)+\rho_{2}(Y)\big)\right\}\geq n(s+\epsilon)$ (1)
 for all $n>N$.
-In the theorem above, a constant depends on a polynomial $\rho$, we means that it depends on the degree and coefficients of $\rho$.
-The notation $1\ll_{s}s_{1},s_{2}$ means that $s_{1}$ and $s_{2}$ are bounded below by some positive constant depending only on $s$, but not on $\rho_{1},\rho_{2},C,c$ or $n$.
 Next, consider polynomials $\rho_{1}$ and $\rho_{2}$ of degree $d\geq 2$, together with an invertible matrix
@@
 In such a situation, one has $\phi(x,y)=\tilde{\rho}_{1}(x)+\tilde{\rho}_{2}(y)$ for some polynomials $\tilde{\rho}_{1}(x)$ and $\tilde{\rho}_{2}(y)$ of degree $d$.
-And we say that $\phi(x,y)$ is non-diagonal if at most one of the entries $\xi_{11}$, $\xi_{12}$, $\xi_{21}$, $\xi_{22}$ is zero.
+We say that $\phi(x,y)$ is non-diagonal if at most one of the entries $\xi_{11}$, $\xi_{12}$, $\xi_{21}$, $\xi_{22}$ is zero.
 Theorem 1.3 (Conditionally Frostman, general forms).
-Let $\phi$ be given as in (1).
-Let $0<s<1$ and $C_{1},C_{2},c>0$.
-There exist positive constants $\epsilon=\epsilon(d,s)$, $\epsilon^{\prime}=\epsilon^{\prime}(d,\Xi,s)$ and an integer $N=N(\phi,s,C_{1},C_{2},c)$ such that the following statement holds.
+Let $\phi$ be given as in (2), $0<s<1$ and $C,c>0$.
+There exist positive constants $\epsilon=\epsilon(d,s)$, $\epsilon^{\prime}=\epsilon^{\prime}(d,\Xi,s)$, and an integer $N=N(\phi,s,C,c)$ such that the following statement holds.
 Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$.
-Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
-Then, for all $n>N$,
-$\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq\begin{cases}n(s+\epsilon),\quad&\text{if }\phi(x,y)\text{ is diagonal},\\ n(s+\epsilon^{\prime}),&\text{if }\phi(x,y)\text{ is non-diagonal}.\end{cases}$
-Here, a constant depends on the matrix $\Xi$ means that it depends on the entries of $\Xi$. And a constant depends on $\phi$ means it depends on $\rho_{1}$, $\rho_{2}$ and $\Xi$.
-Theorem 1.4 (Marginally Frostman with independence).
-Let $\phi$ be given as in (1).
-Let $0<s<1$ and $C_{1},C_{2},c>0$.
-There exist positive constants $\epsilon=\epsilon(s)$, $\epsilon^{\prime}=\epsilon^{\prime}(\phi,s)$ and an integer $N=N(\phi,s,C_{1},C_{2},c)$ such that the following statement holds.
-Let $X$ and $Y$ be two independent real random variables whose ranges are contained in [c,c][-c,c].
-Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
-Then, for all $n>N$,
-$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq\begin{cases}n(s+\epsilon),\quad&\text{if }\phi(x,y)\text{ is diagonal},\\ n(s+\epsilon^{\prime}),&\text{if }\phi(x,y)\text{ is non-diagonal}.\end{cases}$ (2)
+Suppose that $(X,Y)$ is conditionally $(s,C;s,C)$-Frostman.
+Then
+$\max\left\{H_{n}(X+Y),\,H_{n}\big(\phi(X,Y)\big)\right\}\geq\begin{cases}n(s+\epsilon),\quad&\text{if }\phi(x,y)\text{ is diagonal},\\ n(s+\epsilon^{\prime}),&\text{if }\phi(x,y)\text{ is non-diagonal}\end{cases}$ (3)
+for all $n>N$.
 To show the sharpness of these results, we summarize the main statements of several examples in Section 8:
@@
     Any $(s,C)$-Frostman random variable $Z$ has the property that $H_{n}(Z)\geq ns-\log C+O(1)$ (see Example 8.1 for an example matching this bound).
-We show in Examples 7.1 and 7.2 that there are i.i.d.
+We show in Examples 8.1 and 8.2 that there are i.i.d.
 random variables $X$ and $Y$ such that at least one of the terms $H_{n}(X+Y)$ or $H_{n}(\phi(X,Y))$ can be as small as $(1+o(1))ns$.
   2. 2.
-    The positive $\epsilon$ might fail to exist if one of the conditions of the theorem is not satisfied, including when $s=0$ or 11 (see Example 7.1), the Frostman constant $C$ grows with $n$ (see Example 7.4), or the polynomial is degenerate (see Example 7.3).
-  3. 3.
     For any $0<s<1$ and any sufficiently large $n\in\mathbb{N}$, there exists an $(s,O(1))$-Frostman random variable $X$ with range contained in $[0,1]$, such that
     $\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}=(1+o(1))H_{n}(X)$
-    as $n\rightarrow\infty$, where $Y$ is i.i.d. with $X$. In other words, the term $ns$ cannot be replaced by $H_{n}(X)$ in the right-hand side of (2).
+    as $n\rightarrow\infty$, where $Y$ is i.i.d. with $X$. In other words, the term nsns cannot be replaced by $H_{n}(X)$ in the right-hand side of (1) and (3).
+  3. 3.
+    Examples 8.1, 8.3, 8.4 and 8.5 in Section 8 show that the hypotheses in Theorem 1.3 are essentially sharp.
+In particular, if one of the conditions is dropped, then a positive uniform gain may no longer hold.
   4. 4.
-    In Theorems 1.2 and 1.3, the necessity of different terms in these maxima (for example, $\frac{H_{n}(X,Y)}{2}$ or $\frac{H_{n}(X)+H_{n}(Y)}{2}$) reflects subtle aspects of how dependency affects entropy, we provide an example showing it cannot be simplified in the statement of Theorem 1.2 (Example 7.5).
+    In Theorems 1.2, the necessity of different term $\frac{H_{n}(X)+H_{n}(Y)}{2}$ in the maxima reflects subtle aspects of how dependency affects entropy.
+Example 8.5 shows that this term cannot be discarded.
   5. 5.
-    In the statement of Theorem 1.4, the fact that NN depends on $C$, and the condition that $X$ and $Y$ are independent are both essential (see Examples 7.4 and 7.5, respectively).
-  6. 6.
-    The value of $\epsilon$ is not explicit, and our arguments suggest that they are quantitatively small.
-We record several further remarks.
-Our proof of Theorems 1.21.4 presents a connection between $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ and the entropy of the “distance” function, i.e.
-$H_{n}\big((X_{1}-Y_{1})^{2}+(X_{2}-Y_{2})^{2}\big)$, where $(X_{1},Y_{1})$, $(X_{2},Y_{2})$ are i.i.d.
-copies of $(X,Y)$.
-When $a_{1}=a_{3}=0$ and $a_{2}=1$, i.e.
-$\phi(x,y)=xy$, Máthé and O’Regan [12] established a different type of connection by replacing the entropy of the “distance” function with that of the “slope” function.
-By combining this with a generalization of Marstrand’s theorem due to Orponen, Shmerkin, and Wang in [21], they were able to prove Theorem 1.4 for the full range of $s$ for this specific bilinear form function.
-We do not see how their method could be adapted to the form of $\phi$ with $a_{1},a_{3}\neq 0$.
-Moreover, in terms of applications, i.e.
-sum-product type problems over $\mathbb{R}$ or $\mathbb{F}_{p}$, the two functions $xy$ and $a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ appear to exhibit distinct behaviors.
-It would be interesting to explore more connections between $\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}$ and the entropy of other functions.
-We now discuss applications of the entropy theorems.
-For $A\subseteq\mathbb{R}$ and $\delta>0$, by $\mathcal{E}_{\delta}(A)$ we mean the $\delta$-covering number of $A$.
+    We do not attempt to optimize $\epsilon$ or $\epsilon^{\prime}$. The argument produces a positive gain depending on the parameters of the problem.
+1.2 New discretized sum-product type estimates along dense graphs
+We now discuss some novel applications of the entropy theorems.
+For a finite set $A\subseteq\mathbb{R}$ and $\delta>0$, by $\mathcal{E}_{\delta}(A)$ we mean the $\delta$-covering number of $A$.
+We say $A$ is $\delta$-separated if the distance between any two distinct elements in $A$ is at least $\delta$.
 For a set $G\subseteq A\times A$, we denote
@@
 $\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\rho_{1}(A)+_{G}\rho_{2}(A))\gg_{\rho_{1},\rho_{2},s}C^{-O_{d,s}(1)}\delta^{-\epsilon}(\#A).$
-By Theorem 1.3, extending this result to more general polynomials is possible, but requiring minimum degree constraints (3) and (4).
+Theorem 1.5 extends this statement to more general polynomials under the natural minimum-degree conditions (4) and (5).
 Theorem 1.5.
@@
 $\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\phi_{G}(A,A))\gg_{\phi,s}C^{-O_{d,s}(1)}\delta^{-\epsilon}(\#A).$
-In contrast to Theorem 1.6, the next result, an application of Theorem 1.4, holds for every scale $\delta>0$ and is stable under taking subsets: it applies to any $B\subseteq A$ with $B\geq\delta^{\varepsilon}(\#A)$.
+In contrast to Theorem 1.5, the next result holds for every scale $\delta>0$ and is stable under taking subsets: it applies to any $B\subseteq A$ with $B\geq\delta^{\varepsilon}(\#A)$.
 Theorem 1.6.
@@
 $\mathcal{E}_{\delta}(B+B)+\mathcal{E}_{\delta}(\phi(B,B))\gg_{\phi,s}C^{-O_{d,s}(1)}\delta^{-\epsilon}(\#A).$
-These applications are of independent interest, particularly when $\phi(x,y)$ is a non-bilinear form, and might be viewed as consequences of the Falconer distance problem.
-If $\phi(x,y)=xy$, then stronger results have been obtained by using a variety of techniques, see [5, 11, 12, 17] and references therein for current progress.
-Given this line of research, it is reasonable to expect different proofs of Theorems 1.51.7 with other methods.
-In the prime field setting, such a result has been proved by using tools from incidence geometry, see [9, Theorem 1.5] and [15, Theorem 1.3], for example.
-When $\phi(x,y)=x^{2}+y^{2}$ ($\alpha=\beta=1$) and $G=A\times A$, after posting this paper to arXiv, Tuomas Orponen pointed out that one can use his proof of [22, Corollary 1.3] and recent results in [1] or [23] to obtain a quantitative lower bound for the following estimate
-$\mathcal{E}_{\delta}(A+A)+\mathcal{E}_{\delta}(A^{\wedge 2}+A^{\wedge 2})\gg_{\alpha,\beta,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A).$
+Theorems 1.41.6 can be stated in a more analytic form, namely, by taking $A\subset[0,1]$ to be a union of closed intervals with total measure $\delta^{1-s}$.
+This observation has been made explicitly in [11].
+For $\phi(x,y)=xy$ and a complete graph $G$, the value of $\epsilon$ has been explicitly computed and improved using a variety of techniques, see [11, 18, 19, 31] and references therein for current progress.
+Note that Theorem 1.6 does not follow from the special case $B=A$, since an arbitrary subset $B$ of $A$ may not satisfy the non-concentration estimate $(B\cap I)\leq Cl^{s}(\#B)$, a phenomenon quite different from the discrete setting.
+In the discrete setting, Erdős and Szemerédi [7] proposed the following strong conjecture: For every $c>0$ and $\epsilon>0$, there exists a threshold $n_{0}$ such that for all $n\geq n_{0}$, any $n$-element subset $A\subseteq\mathbb{N}$, and any graph $G$ with $n$ vertices and at least $n^{1+c}$ edges,
+$(A+_{G}A)+\#(A\cdot_{G}A)\geq(\#A)^{1+c-\epsilon}.$
+When $G$ is the complete graph (i.e., $c=1$), this is the famous Erdős–Szemerédi sum-product conjecture, which remains wide open.
+Current records can be found in [2, 32].
+For arbitrary graphs $G$ with at least $n^{1+c}$ edges, Alon, Ruzsa, and Solymosi [1] refuted the strong conjecture above by constructing examples with simultaneously small sumsets and product sets along a graph.
+They also established that
+$(A+_{G}A)+\#(A\cdot_{G}A)\gg\frac{(\#E(G))^{3/2}}{(\#A)^{7/4}}.$
+Using the continuous-to-discrete adaptable-set framework developed in [14], as applications of Theorems 1.4, 1.5, and 1.6, we obtain analogous results for much more general families of polynomials in the setting of adaptable sets.
+Roughly speaking, if $A\subseteq[0,1]$ is a non-concentrated/adaptable $n$-point set and $G\subseteq A\times A$ is almost complete in the sense that every vertex has degree at least $n^{1-o(1)}$, then the restricted sumset $A+_{G}A$ and the restricted polynomial image $\phi_{G}(A,A)$ cannot both be small.
+It would be interesting to determine whether these results (Theorems 7.3, 7.4, and 7.5) can be proved independently using purely combinatorial methods.
+There is a substantial parallel literature on sum-product and polynomial expansion over finite fields, especially over prime fields $\mathbb{F}_{p}$, where nontrivial lower bounds typically require the size of $A$ to be small relative to $p$.
+Bourgain, Katz, and Tao [3] established the first nontrivial sum-product estimate in $\mathbb{F}_{p}$, showing that $(A+A)+\#(AA)\geq(\#A)^{1+\epsilon}$ for some $\epsilon>0$ when #A=o(p)\#A=o(p).
+The current record lower bound is $(\#A)^{\frac{5}{4}}$ due to Mohammadi and Stevens [22].
+In a more general setting, Bukh and Tsimerman [4] proved that for any polynomial $f\in\mathbb{F}_{p}[X]$ of degree $d\geq 2$ and any $A\subset\mathbb{F}_{p}$ with $A\leq\sqrt{p}$,
+[…]
+If $f(x)=x^{2}$, then Pham, Vinh, and de Zeeuw [29] improved the lower bound further to $(\#A)^{\frac{6}{5}}$.
+For a non-degenerate quadratic polynomial $\phi(x,y)$, Koh, Mojarrad, Pham, and Valculescu [16] proved that $(A+A)+\#(\phi(A,A))\gg(\#A)^{\frac{6}{5}}$.
+Bukh and Tsimerman [4] also established a one-sided sum-product estimate along graphs: for sets $A,B\subset\mathbb{F}_{p}$ and a graph $G\subset A\times B$ with $A,\#B\leq p^{\frac{1+\frac{1}{29}}{2}}$ and $G\geq\frac{1}{K}\#A\cdot\#B$, one has
+$(A+A)+\#(A\cdot_{G}B)\gg\frac{(\#B)^{\frac{1}{29}}}{K}\#A.$
+1.3 Main ideas and novelties
+In this subsection, we provide an overview of the approaches for Theorems 1.2 and 1.3.
+The essential starting point is to find some random variable that not only links with both $X+Y$ and $\phi(X,Y)$, but also produce entropy gain.
+In the case $\phi(x,y)=xy$, Máthé and O’Regan [19] use the slope (or direction) variable.
+Their key geometric input is a sharp lower bound for the set of directions spanned by segments joining two planar sets, due to Orponen–Shmerkin–Wang [28].
+In particular, if $A\subseteq[1,2]$ has Hausdorff dimension $s$, then applying the direction-set theorem to $E=A\times A$ and $F=(-A)\times(-A)$ yields that the corresponding slope set has dimension at least $\min\{1,2s\}$.
+Equivalently, one obtains a large set of projective ratios
+$\left\{\frac{a+b}{c+d}:a,b,c,d\in A\right\}.$
+In entropy language, this gives a lower bound for the discretized entropy of the slope random variable $(X+Y)/(Z+W)$ built from i.i.d.
+variables $X,Y,Z,W$ supported on $A$.
+The second step of [19] is a Shannon-type (submodularity) inequality that controls the entropy of this slope variable in terms of the entropies of $X+Y$ and $XY$ (and the base entropy $H(X)$), and hence forces entropy growth in at least one of these simpler expressions.
+For a general polynomial $\phi$, adapting the methods from [19] would require direction-set estimates for a corresponding nonlinear direction map, and existing projection theorems do not provide the kind of uniform quantitative control needed in this setting.
+Accordingly, we work with a distance function variable instead.
+Our proofs contain five main steps as shown below.
+(A) Entropy of distance variable (Theorem 2.38 and Theorem 5.1).
+We establish a distance-entropy lower bound at dyadic scale $2^{-n}$, typically for
+$Z:=|W-W^{\prime}|\qquad\text{or}\qquad Z^{2}:=|W-W^{\prime}|^{2},$
+where $W^{\prime}$ is an i.i.d. copy of a planar random variable $W$ whose distribution is $(2s,C)$-Frostman and supported in a ball $B(0,R)\subseteq\mathbb{R}^{2}$.
+Roughly speaking, when $s>\tfrac{1}{2}$, the distribution measure $\mu$ of $W$ has dimension $2s>1$ in the plane, i.e.
+it is thick enough for Fourier analytic distance-set technology to apply.
+Concretely, we use Mattila’s estimate (Lemma 2.36), which controls short-interval mass of the distance measure $\Delta(\mu)$ in terms of a Riesz energy:
+$\Delta(\mu)([r,r+\eta])\ \ll_{t}\ r^{1/2}\,\eta^{t}\,I_{t+1/2}(\mu),$
+for any $0<t\leq 1$ and $0\leq\eta\leq r$.
+This directly implies that the distance measure is $t$-Frostman (Lemma 2.37) with $t<2s-\frac{1}{2}$.
+To get an entropy gain, we would need $s<t$, which explains the required threshold of $s>\frac{1}{2}$ by Mattila’s methods.
+In the range $s\leq\tfrac{1}{2}$, we instead use a distance-entropy input based on the work of Shmerkin and Wang [35].
+This input is available only under a product structure: it treats distance measures between $\mu_{1}\times\nu$ and $\mu_{2}\times\nu$, and in random-variable language, it corresponds to $W=(X,Y)$ with $X$ and $Y$ independent and each marginal $(s,C)$-Frostman.
+The way this enters our entropy scheme is not a routine black-box reduction.
+Rather, we adapt the robust projection / thin-tube framework of [35] and carry out an additional conversion step that turns their robustness conclusions into a fixed-scale dyadic entropy lower bound at scale $2^{-n}$, with all losses made explicit.
+A central issue is quantitative bookkeeping under conditioning.
+In later steps, we condition on discretized events, producing families of conditional measures whose Frostman constants are highly non-uniform and may grow with the scale, in the sense that $\log C$ can be comparable to $n$.
+For this reason we need error terms of the form $O(\log C)$ rather than $O_{C}(1)$, so that these losses can be absorbed into entropy terms when we average over the conditioning parameters.
+(B) Diagonal quadratic $\alpha X^{2}+\beta Y^{2}$ with independent random variables (Lemma 4.1).
+Write $W=(\sqrt{|\alpha|}X,\sqrt{|\beta|}Y)$ and let $(X_{1},Y_{1})$, $(X_{2},Y_{2})$ be i.i.d. copies of $(X,Y)$, with $W_{i}=(\sqrt{|\alpha|}X_{i},\sqrt{|\beta|}Y_{i})$. Then
+$|W_{1}-W_{2}|^{2}=\alpha(X_{1}-X_{2})^{2}\pm\beta(Y_{1}-Y_{2})^{2}.$
+So $|W_{1}-W_{2}|^{2}$ is a Lipschitz function of $(X_{1}-X_{2},\;Y_{1}-Y_{2})$ on the relevant bounded ranges.
+Moreover, by a polarization identity, $|W_{1}-W_{2}|^{2}$ is also a Lipschitz function of the triple
+$\bigl(X_{1}+X_{2},\;Y_{1}+Y_{2},\;\alpha(X_{1}^{2}+X_{2}^{2})\pm\beta(Y_{1}^{2}+Y_{2}^{2})\bigr).$
+Applying discretized submodularity (with Lipschitz error terms) yields an upper bound for $H_{n}(|W_{1}-W_{2}|^{2})$ in terms of the entropies of these sum/difference expressions.
+Plünnecke–Ruzsa type entropy inequalities for independent variables then control those entropies by a linear combination of $H_{n}(X+Y)$, $H_{n}(\alpha X^{2}+\beta Y^{2})$, and the marginal entropies $H_{n}(X)$ and $H_{n}(Y)$, up to logarithmic losses.
+Combining this with the distance-entropy lower bound for $H_{n}(|W_{1}-W_{2}|^{2})$ forces a gain in $H_{n}(X+Y)+H_{n}(\alpha X^{2}+\beta Y^{2})$.
+(C) Diagonal polynomial $\rho_{1}(x)+\rho_{2}(y)$ with independent random variables (Lemma 4.8).
+A Plünnecke–Ruzsa type entropy inequality (Lemma 2.29), combined with a local weak bi-Lipschitz decomposition of $\rho_{1},\rho_{2}$ on suitable regular regions, permits a two-point differencing scheme: after conditioning on dyadic differences, one reduces degree $d$ to degree d1d-1.
+More concretely, on an event where $\rho_{1}$ and $\rho_{2}$ are weakly bi-Lipschitz, Lemma 2.29 relates $H_{n}(\rho_{1}(X)+\rho_{2}(Y))$ to the two-point expression $H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2}))$, where $(X_{1},Y_{1}),(X_{2},Y_{2})$ are i.i.d.
+copies.
+Conditioning on $D_{n}(X_{1}-X_{2})=z_{1}$ and $D_{n}(Y_{1}-Y_{2})=z_{2}$, and using Lipschitz control, one identifies (up to logarithmic losses) this conditional entropy with
+$(\Delta_{z_{1}}\rho_{1})(X_{2})+(\Delta_{z_{2}}\rho_{2})(Y_{2}),\qquad(\Delta_{z}\rho)(x):=\rho(x+z)-\rho(x),$
+where the difference polynomials have degree d1d-1.
+For gaps bounded away from 0 one applies the inductive hypothesis (with the quadratic base case given by Step (B)) to the conditional law $(X_{2},Y_{2}\,|\,D_{n}(X_{1}-X_{2})=z_{1},\ D_{n}(Y_{1}-Y_{2})=z_{2})$, which remains independent and marginally $(s_{1},\widetilde{C}_{1};\,s_{2},\widetilde{C}_{2})$-Frostman, with
+$\widetilde{C}_{1}=O\!\left(\frac{C_{1}^{2}\,2^{-ns_{1}}}{P_{E}(D_{n}(X_{1}-X_{2})=z_{1})}\right),\qquad\widetilde{C}_{2}=O\!\left(\frac{C_{2}^{2}\,2^{-ns_{2}}}{P_{E}(D_{n}(Y_{1}-Y_{2})=z_{2})}\right).$
+Finally, a weighted average over $(z_{1},z_{2})$ converts the $\log\widetilde{C}_{i}$ terms into entropies of the dyadic difference variables (e.g.
+$H_{n}(X_{1}-X_{2})$ and $H_{n}(Y_{1}-Y_{2})$), allowing one to close the induction after discarding very small differences.
+There are three technical challenges in this step.
+First, one needs a local weak bi-Lipschitz decomposition for $\rho$ away from its critical set(s), so that entropy can be transferred between $X$ and $\rho(X)$ with only logarithmic losses in degree and coefficient size.
+Second, the conditional Frostman constants $\widetilde{C}_{1},\widetilde{C}_{2}$ depend on $(z_{1},z_{2})$ and can be very large; it is therefore essential that all losses enter only through $\log\widetilde{C}_{i}$, so that after averaging they become controllable entropy terms.
+Third, the complexity of $\Delta_{z}\rho$ deteriorates as $|z|\to 0$, so one discards very small differences in a controlled way while retaining enough mass/entropy to complete the degree reduction (see Section 4.2).
+(D) Diagonal polynomial $\rho_{1}(x)+\rho_{2}(y)$ with dependent random variables (Theorem 1.2).
+A discretized entropy Balog–Szemerédi–Gowers mechanism (Theorem 3.1), allowing weakly bi-Lipschitz transformations, reduces the dependent case to a controlled form of conditional independence at dyadic scale.
+The losses are quantified in terms of the discretized mutual information $H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)$ and the logarithmic errors coming from conditioning on $D_{n}(\cdot)$.
+In particular, this dependence penalty is exactly what forces the baseline term $\frac{H_{n}(X)+H_{n}(Y)}{2}$ to appear in Theorem 1.2:: in the jointly Frostman setting one cannot treat it as $O(1)$ and must keep it inside the main maximum.
+More concretely, from $(X,Y)$ we construct auxiliary variables $(X^{\prime\prime},X^{\prime},Y^{\prime},Y^{\prime\prime})$ such that $X^{\prime\prime}$ and $Y^{\prime\prime}$ are conditionally independent given $(D_{n}(X^{\prime}),D_{n}(Y^{\prime}))$.
+Conditioning on $(D_{n}(X^{\prime}),D_{n}(Y^{\prime}))=(x,y)$ then yields a family of independent conditional laws $(X^{\prime\prime},Y^{\prime\prime}\,|\,(D_{n}(X^{\prime}),D_{n}(Y^{\prime}))=(x,y))$ whose marginals are Frostman with constants
+$C_{1}^{\prime}=O\!\left(\frac{C\,2^{-ns_{2}}}{P_{E}\!\big(D_{n}(Y^{\prime})=y\big)}\right),\qquad C_{2}^{\prime}=O\!\left(\frac{C\,2^{-ns_{1}}}{P_{E}\!\big(D_{n}(X^{\prime})=x\big)}\right).$
+as obtained from the joint Frostman hypothesis on $(X,Y)$.
+Applying the independent diagonal-polynomial estimate to each conditional law and then averaging over $(x,y)$ converts the $\log C_{i}^{\prime}$ losses into entropies of the dyadic conditioning data, allowing one to close the argument.
+There are two technical challenges in this step.
+First, the Frostman constants $C_{1}^{\prime},C_{2}^{\prime}$ vary with $(x,y)$ and may be large, so one must keep all dependence only through logarithms and then turn these into controllable entropy terms after averaging.
+Second, the conditioning is only on dyadic discretizations $D_{n}(\cdot)$ rather than on the underlying variables, so the usual Markov-chain interpretation in BSG theorem is not immediate: to prove Theorem 3.1 one must build an explicit coupling and verify that the required conditional i.i.d./independence relations persist after discretization, with the correct logarithmic losses.
+(E) General polynomial $\phi(x,y)$ (Theorems 1.3).
+The passage from diagonal forms $\rho_{1}(x)+\rho_{2}(y)$ to general forms $\phi(x,y)$ obtained by an invertible rational linear change of variables relies on controlling how discretized entropy behaves under rational linear combinations when the variables may be dependent.
+In particular, under conditional Frostman hypotheses, an invertible rational change of coordinates produces a pair of linear forms whose joint law remains non-concentrated (jointly Frostman with controlled constants), allowing us to reduce the general form to the diagonal core.
+A key input here is a collection of discretized Ruzsa-type entropy inequalities that remain valid without independence, culminating in the quantitative comparison theorem for rational coefficients (Theorem 2.23).
+Concretely, for $k,\ell\in\mathbb{Q}$, Theorem 2.23 bounds the “excess entropy”
+$H_{n}(kX+\ell Y)-\frac{H_{n}(X)+H_{n}(Y)}{2}$
+by a linear combination (with coefficients depending on $k,\ell$) of the corresponding excess for $X+Y$ and a dependence term measured by the discretized mutual information
+$I_{n}(X;Y):=H_{n}(X)+H_{n}(Y)-H_{n}(X,Y),$
+up to $O_{k,\ell}(1)$ errors.
+In particular, this dependence bookkeeping clarifies the different baseline terms in our main results: in the merely jointly Frostman regime we must work relative to $\tfrac{1}{2}(H_{n}(X)+H_{n}(Y))$, whereas in the conditionally Frostman regime it is natural to use a baseline involving $\tfrac{1}{2}H_{n}(X,Y)$, since $\tfrac{1}{2}(H_{n}(X)+H_{n}(Y))-\tfrac{1}{2}H_{n}(X,Y)=\tfrac{1}{2}I_{n}(X;Y)$.
+These comparison inequalities are stable for rational coefficient relations but can fail for $\mathbb{Q}$-linearly independent real coefficients (Example 8.7).
+For this reason, our general-form results are stated for $\Xi\in\mathrm{GL}_{2}(\mathbb{Q})$, and the reduction to the diagonal core is carried out using rational linear algebra rather than arbitrary real changes of variables.
+Finally, in the dependent setting the joint entropy $H_{n}(X,Y)$ arises naturally and must be eliminated from the final max-form lower bounds.
+A key new ingredient in our proof is a quantitative entropy-comparability statement (Proposition 4.10), showing that under conditional $(s,C;s,C)$-Frostman assumptions,
+$H_{n}(X,Y)=H_{n}\big(X+Y,\ \phi(X,Y)\big)+O_{\phi,s,C,c}(1),$
+[…]
 2 Preliminaries
@@
 And whenever we write $f(X)$ with $X$ a random variable and $f$ a function, we always mean that $domain(f)\supseteq range_{+}(X)$, where the values taken on a zero-measure set is negligible.
-In this paper, by $f=O_{\kappa}(g)$, $f\ll_{\kappa}g$ or $g\gg_{\kappa}f$, we mean that there exists an absolute positive constant $C=C(\kappa)$ depending on the parameter(s) $\kappa$ such that $|f|\leq Cg$.
-All logarithms we use here have base $2$. Moreover, we denote $\log_{+}x:=\max\{\log x,\,0\}$ for any $x>0$.
+In this paper, we denote by #S\#S the cardinality of a finite set $S$.
+By $f=O_{\kappa}(g)$, $f\ll_{\kappa}g$ or $g\gg_{\kappa}f$, we mean that there exists an absolute positive constant $C=C(\kappa)$ depending on the parameter(s) $\kappa$ such that $|f|\leq Cg$.
+All logarithms we use here have base $2$.
+Moreover, we denote $\log_{+}x:=\max\{\log x,\,0\}$ for any $x>0$.
 2.2 Entropy and basic properties
@@
 We also interpret $\mu(B)H(\mu_{B},\mathcal{A})$ as 0 when $\mu(B)=0$, even though $\mu_{B}$ is not defined in this situation.
-Intuitively, the entropy of a random variable will tell us how hard it is to determine a value.
 Before stating the first lemma, we introduce the following notation:
 $Z(\mathcal{A})=\{A\in\mathcal{A}\mid\mu(A)>0\},\quad\mathcal{A}\land\mathcal{B}=\{A\cap B\mid A\ \in\mathcal{A},B\in\mathcal{B}\}.$
-Moreover, the function $\iota(x)=-x\log x$ is non-negative valued, bounded and concave on $[0,1]$.
+Also notice that the function $\iota(x)=-x\log x$ is non-negative valued, bounded and concave on $[0,1]$.
 It is increasing on $0,e^{-1}]$, and decreasing on $e^{-1},1]$.
@@
 (v) $H(\mu,\mathcal{A}\land\mathcal{B}|\mathcal{C})\leq H(\mu,\mathcal{A}|\mathcal{C})+H(\mu,\mathcal{B}|\mathcal{C})$.
-(vi) For any Borel set $Y$, we have $\mu(Y)H(\mu_{Y},\mathcal{A})\leq H(\mu,\mathcal{A}).$
+(vi) For any Borel set $E$, we have $\mu(E)H(\mu_{E},\mathcal{A})\leq H(\mu,\mathcal{A}).$
 For discrete random variables $X$ and $Y$ defined on $(\Omega,\mathfrak{B},P)$, the families
@@
 $H(X|Y)=\sum_{y}P(Y=y)H(X|Y=y).$
+Intuitively, the entropy of a random variable will tell us how hard it is to determine a value.
 For the random variable $Z=(X,Y)$, we denote $H(Z)=H(X,Y)$, instead of $H\big((X,Y)\big)$.
@@
 $\frac{\partial\tilde{f}}{\partial p_{i}}(p_{1},\ldots,p_{j-1},p_{j+1},\ldots,p_{k+1})=\log\frac{1-\sum\nolimits_{l\neq j}p_{l}}{p_{i}}=\log\frac{p_{j}}{p_{i}}\leq 0.$
-In other words, when the other k1k-1 variables are given, the small $p_{i}$ is, the larger value $\tilde{f}$ and $f$ take.
+In other words, when the other k1k-1 variables are given, the smaller $p_{i}$ is, the larger value $\tilde{f}$ and $f$ take.
 Now we can verify that the maximum value of $f$ is taken at $(p_{1},\ldots,p_{k+1})$ with
@@
 where for each $i$ the integer $k_{i}$ is uniquely determined by $\frac{k_{i}}{2^{n}}\leq x_{i}<\frac{k_{i}+1}{2^{n}}$.
-So, $D_{n,m}(x)$ locates the left endpoints of the dyadic intervals that the $x_{i}$’s fall in.
+That is to say, $D_{n,m}(x)$ locates the left endpoints of the dyadic intervals that the $x_{i}$’s fall in.
 For convenience, we often write $D_{n}$ instead of $D_{n,m}$ when the dimension is clear from the context.
@@
 And define $H_{n}(\mu)=H_{n}(\mu,\mathcal{D}_{n})$.
-When $\mu$ is the distribution measure associated to $X$, the above two defintions coincide.
+When $\mu$ is the distribution measure associated to $X$, the above two definitions coincide.
 Some properties still hold if we replace $H$ by $H_{n}$. For example,
@@
 Lemma 2.13.
-Let $\mathcal{R}\subseteq{\mathbb{R}}$ and $f:\,\mathcal{R}\rightarrow{\mathbb{R}}$ be a differentiable function.
+Let $\mathcal{R}\subseteq{\mathbb{R}}$ and $f$ be a differentiable function with $domain(f)\supseteq\mathcal{R}$.
 Suppose that $\mathcal{R}$ is a union of $d$ intervals such that $f$ is monotonic on each of the intervals.
@@
 (ii) If $g$ is weakly $(\kappa,d)$-bi-Lipschitz on $range_{+}(Y)$, then for all $n\in\mathbb{N}$,
-$H_{n}(Y|g(Y)))\ll\log\kappa+\log d.$
+$H_{n}(Y|g(Y))\ll\log\kappa+\log d.$
 The implied constants are independent of $n$ and of the particular functions and random variables involved.
@@
 $H_{n}(f(X))\leq H_{n}(f(Y))+H_{n}(f(X)|f(Y)).$
-When $D_{n}(f(Y)))=t$ is determined, one has $|f(Y)-t|=O_{m}(2^{-n})$ and then $|f(X)-t|\leq\kappa K+O_{m}(2^{-n})$.
+When $D_{n}(f(Y))=t$ is determined, one has $|f(Y)-t|=O_{m}(2^{-n})$ and then $|f(X)-t|\leq\kappa K+O_{m}(2^{-n})$.
 So $D_{n}(f(X))$ has at most $O\big((\kappa K\cdot 2^{n}+O_{m}(1))^{m}\big)$ choices.
@@
 Therefore, the conclusion follows with
-$d_{1}(\pm(m+1))=d_{1}(\pm m)+4,\qquad d_{1}(\pm(m+1))=d_{1}(\pm m)+8.$
+$d_{1}(\pm(m+1))=d_{1}(\pm m)+4,\qquad d_{2}(\pm(m+1))=d_{2}(\pm m)+8.$
 Lemma 2.27.
@@
 for any $n\in\mathbb{N}$.
-However, when $k,l$ are $\mathbb{Q}$-linearly independent, the above formula may fail.
+However, when two real numbers $k,l$ are $\mathbb{Q}$-linearly independent, the above formula may fail.
 In Example 8.7, we would show the following: For any $K>0$, there exists some $N\in\mathbb{N}$ such that: for any $n>N$, there is a random variable $X$ with range contained in $[0,1]$ satisfying
@@
 where $Y$ is i.i.d. with $X$.
+The last lemma in this subsection is a Plünnecke–Ruzsa type inequality with functions.
 Lemma 2.29.
@@
 Then
-$H_{n}(f(X_{1})+f(X_{2})+g(Y_{1})+g(Y_{2}))$ $\leq 3H_{n}(f(X)+g(Y))-H_{n}(X)-H_{n}(Y)+O(\log\kappa)+O(\log d),$
-$H_{n}(f(X_{1})-f(X_{2})+g(Y_{1})-g(Y_{2}))$ $\leq 7H_{n}(f(X)+g(Y))-3H_{n}(X)-3H_{n}(Y)+O(\log\kappa)+O(\log d),$
-$H_{n}(f(X_{1})+f(X_{2})-g(Y_{1})-g(Y_{2}))$ $\leq 9H_{n}(f(X)+g(Y))-4H_{n}(X)-4H_{n}(Y)+O(\log\kappa)+O(\log d).$
+$H_{n}(f(X_{1})+f(X_{2})+g(Y_{1})+g(Y_{2}))$ $\leq 3H_{n}(f(X)+g(Y))-H_{n}(X)-H_{n}(Y)+O(\log\kappa+\log d),$
+$H_{n}(f(X_{1})-f(X_{2})+g(Y_{1})-g(Y_{2}))$ $\leq 7H_{n}(f(X)+g(Y))-3H_{n}(X)-3H_{n}(Y)+O(\log\kappa+\log d),$
+$H_{n}(f(X_{1})+f(X_{2})-g(Y_{1})-g(Y_{2}))$ $\leq 9H_{n}(f(X)+g(Y))-4H_{n}(X)-4H_{n}(Y)+O(\log\kappa+\log d).$
 Proof.
@@
 $\qquad-H_{n}(f(X_{1})+g(Y_{1}))+O(1)$
-$=H_{n}(f(X)+g(Y))+2H_{n}(f(X)\pm g(Y))-H_{n}(X)-H_{n}(Y)+O(\log\kappa)+O(\log d).$
+$=H_{n}(f(X)+g(Y))+2H_{n}(f(X)\pm g(Y))-H_{n}(X)-H_{n}(Y)+O(\log\kappa+\log d).$
 By Remark 2.25, it satisfies that
 $H_{n}(f(X)-g(Y))$ $\leq 3H_{n}(f(X)+g(Y))-H_{n}(f(X))-H_{n}(g(Y))+O(1)$
-$=3H_{n}(f(X)+g(Y))-H_{n}(X)-H_{n}(Y)+O(\log\kappa)+O(\log d).$
+$=3H_{n}(f(X)+g(Y))-H_{n}(X)-H_{n}(Y)+O(\log\kappa+\log d).$
 The above formulae still hold if $g$ is replaced by g-g. The lemma then follows from calculation. ∎
@@
 Definition. Let $s,s_{1},s_{2}\geq 0$ and $C,C_{1},C_{2}>0$.
-(i) We say that the random variable $(X,Y)$ is jointly $(s_{1},s_{2},C)$-Frostman, if
+(i) We say that the random variable $(X,Y)$ is jointly $(s_{1},s_{2},C)$-Frostman if
 $P(X\in B(x,r_{1}),\,Y\in B(y,r_{2}))\leq Cr_{1}^{s_{1}}r_{2}^{s_{2}}$
 for any ball $B(x,r_{1})$ and $B(y,r_{2})$.
-(ii) We say that the random variable $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, if $X$ and $Y$ are $(s_{1},C_{1})$-Frostman and $(s_{2},C_{2})$-Frostman, respectively.
-(iii) We say a random variable $X$ is conditionally $(s,C)$-Frostman given $Y$, if
+(ii) We say that the random variable $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman if $X$ and $Y$ are $(s_{1},C_{1})$-Frostman and $(s_{2},C_{2})$-Frostman, respectively.
+(iii) We say a random variable $X$ is conditionally $(s,C)$-Frostman given $Y$ if
 $P(X\in B(x,r),\,Y\in E)\leq Cr^{s}P(Y\in E)$
 for any ball $B(x,r)$ and Borel set $E$.
-Moreover, We say that the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, if $X$ is conditionally $(s_{1},C_{1})$-Frostman given $Y$, and $Y$ is conditionally $(s_{2},C_{2})$-Frostman given $X$.
+Moreover, We say that the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman if $X$ is conditionally $(s_{1},C_{1})$-Frostman given $Y$, and $Y$ is conditionally $(s_{2},C_{2})$-Frostman given $X$.
+When $X$ is conditionally $(s,C)$-Frostman given $Y$, the conditional distribution of $X$ given $Y=y$ is $(s,C)$-Frostman for $P$-a.e. $y$.
 The following properties follow directly from the definitions, whose proofs are left as an exercise.
@@
 By symmetry, it is also $(s_{1},C^{\prime}_{1})$-Frostman with $\log C_{1}^{\prime}=\log C_{1}+O_{a,b}(1)$, provided that $a\neq 0$.
-(ii) Write $p:=P\big(aX+bY\in B(z_{1},r_{1}),\,cX+dY\in B(z_{2},r_{2})\big)$ for an arbitrary pair of balls $B(z_{1},r_{1})$ and $B(z_{2},r_{2})$. Calculation shows that, if $p>0$, then we must have
+(ii) Write $p:=P\big(aX+bY\in B(z_{1},r_{1}),\,cX+dY\in B(z_{2},r_{2})\big)$ for an arbitrary pair of balls $B(z_{1},r_{1})$ and $B(z_{2},r_{2})$. Calculation shows that if $p>0$, then we must have
 $X\in B\Big(\frac{dz_{1}-bz_{2}}{ad-bc},\,\frac{|d|r_{1}+|b|r_{2}}{|ad-bc|}\Big),\qquad Y\in B\Big(\frac{-cz_{1}+az_{2}}{ad-bc},\,\frac{|c|r_{1}+|a|r_{2}}{|ad-bc|}\Big).$ (18)
@@
 $\leq\sum\limits_{i}P\big(aX+bY\in B(z_{1},r_{1}),\,Y\in J_{i}\big):=\sum\limits_{i}p_{i}.$
-One can check that, if $p_{i}>0$, then $X$ is located in an interval $I_{i}$ of length at most $\frac{(|b|+2)r_{1}}{|a|}$. Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, we have
+One can check that if $p_{i}>0$, then $X$ is located in an interval $I_{i}$ of length at most $\frac{(|b|+2)r_{1}}{|a|}$. Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, we have
 $p_{i}\leq P(X\in I_{i},Y\in J_{i})\leq C_{1}\Big(\frac{(|b|+2)r_{1}}{|a|}\Big)^{s_{1}}P(Y\in J_{i}).$
@@
 We apply similar arguments as above. For non-zero terms, the value of $Y$ locates in an interval $J_{i}^{\prime}$ of length at most $\frac{(|c|+2)r_{2}}{|d|}$. Therefore,
-$p_{i}^{\prime}\leq P(X_{i}\in I_{i}^{\prime},Y\in J_{i}^{\prime})\leq C_{1}\Big(\frac{(|c|+2)r_{2}}{|d|}\Big)^{s_{2}}P(X\in I_{i}^{\prime}),$
+$p_{i}^{\prime}\leq P(X_{i}\in I_{i}^{\prime},Y\in J_{i}^{\prime})\leq C_{2}\Big(\frac{(|c|+2)r_{2}}{|d|}\Big)^{s_{2}}P(X\in I_{i}^{\prime}),$
 and then
-$p$ $\leq C_{1}\Big(\frac{(|c|+2)r_{2}}{|d|}\Big)^{s_{2}}P\left(X\in B\Big(\frac{dz_{1}-bz_{2}}{ad-bc},\,\frac{(|d|+|b|)r_{1}}{|ad-bc|}\Big)\right)$
-$\leq C_{1}\Big(\frac{(|c|+2)r_{2}}{|d|}\Big)^{s_{2}}\cdot C_{1}\Big(\frac{dz_{1}-bz_{2}}{ad-bc},\,\frac{(|d|+|b|)r_{1}}{|ad-bc|}\Big)^{s_{1}}.$
+$p$ $\leq C_{2}\Big(\frac{(|c|+2)r_{2}}{|d|}\Big)^{s_{2}}P\left(X\in B\Big(\frac{dz_{1}-bz_{2}}{ad-bc},\,\frac{(|d|+|b|)r_{1}}{|ad-bc|}\Big)\right)$
+$\leq C_{2}\Big(\frac{(|c|+2)r_{2}}{|d|}\Big)^{s_{2}}\cdot C_{1}\Big(\frac{dz_{1}-bz_{2}}{ad-bc},\,\frac{(|d|+|b|)r_{1}}{|ad-bc|}\Big)^{s_{1}}.$
 We conclude that $(X,Y)$ is jointly $(s_{1},s_{2},C^{\prime\prime})$-Frostman with $\log C^{\prime\prime}=\log C_{1}C_{2}+O_{a,b,c,d}(1)$.
@@
 Proof.
-(i) The conclusion comes from [12, Lemma 4.8].
+(i) The conclusion comes from [19, Lemma 4.8].
 (ii) Since $X$ is conditionally $(s,C)$-Frostman given $Y$, then
@@
 $H_{n}(X,Y)-H_{n}(Y)\geq ns_{1}-\log C_{1}+O(1),\quad H_{n}(X,Y)-H_{n}(X)\geq ns_{2}-\log C_{2}+O(1).$
-The conclusion then follows.
+The conclusion then follows. ∎
 The function $x\mapsto x^{2}$ is weakly-bi-Lipschitz in a region that is bounded away from zero.
@@
 $H_{n}(Z_{E})=\frac{H_{n}(Z)}{P(E)}+O(C(2\eta)^{s}n)+O_{m}(1).$
-for any all $n\in\mathbb{N}$.
+for all $n\in\mathbb{N}$.
 Proof.
@@
 $\Delta(\mu)(B)=\int\mu(\{y\in\Omega:|x-y|\in B\})d\mu(x)$
-for any Borel set $B$. That is to say, $\Delta(\mu)$ is the image of $\mu\times\mu$ under the distance map $(x,y)\mapsto|x-y|$. We now recall a result from Mattila in [13, Section 4.2, page 61].
+for any Borel set $B$. That is to say, $\Delta(\mu)$ is the image of $\mu\times\mu$ under the distance map $(x,y)\mapsto|x-y|$. We now recall a result from Mattila in [20, Section 4.2, page 61].
 Lemma 2.36 ([20]).
@@
 $\qquad\leq(t+1/2)\left(\int_{0}^{2\tilde{R}}C\ell^{2s-t-3/2}d\ell+\int_{2\tilde{R}}^{\infty}\ell^{-t-3/2}d\ell\right)$
-$\qquad=O_{t}\left(\frac{C\tilde{R}^{2s-t-1/2}}{2s-t-1/2}\right)+O(\tilde{R}^{-t-1/2}):=M.$
+$\qquad\ll\frac{C\tilde{R}^{2s-t-1/2}}{2s-t-1/2}+\tilde{R}^{-t-1/2}:=M.$
 Thus
@@
 When $0<\eta<r<2R$ and $0<t\leq 1$, one obtains by Lemma 2.36 that
-$\Delta(\mu)([r,r+\eta])\ll_{t}r^{1/2}\eta^{t}I_{t+1/2}(\mu)=O_{t}\left(\frac{C\tilde{R}^{2s-t}\eta^{t}}{2s-t-1/2}\right)+O(\eta^{t}).$
+$\Delta(\mu)([r,r+\eta])\ll_{t}r^{1/2}\eta^{t}I_{t+1/2}(\mu)=O_{t}\left(\frac{C\tilde{R}^{2s-t}\eta^{t}}{2s-t-1/2}+\eta^{t}\right).$
 The proof is completed. ∎
 In this subsection, we show some results of the discretized entropy of the distance function of certain measures, together with the square of the distance.
-These theorems are the key ingredients of the proof of Theorem 1.4.
+These theorems are the key ingredients of the proof of Theorems 1.2 and 1.3.
 Theorem 2.38.
@@
 Then, for any $n\in\mathbb{N}$,
-$H_{n}\big((X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}\big)\geq ns+O_{s}(\log_{+}C)+O_{s}(\log n),$
+$H_{n}\big((X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}\big)\geq ns+O\left(\frac{1}{s}(\log_{+}C+\log n)+1\right),$
 and
-$H_{n}\big(\sqrt{(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}}\big)\geq ns+O_{s}(\log_{+}C)+O(\log_{+}c)+O_{s}(\log n).$
+$H_{n}\big(\sqrt{(X-X^{\prime})^{2}+(Y-Y^{\prime})^{2}}\big)\geq ns+O\left(\frac{1}{s}(\log_{+}C+\log n)+1\right)+O(\log_{+}c).$
 Proof.
@@
 3 Discretized entropy Balog-Szemerédi-Gowers theorem
-In this section, we follow the method introduced by Tao in [26, Theorem 3.1] to obtain a generalized discretized entropy version of the Balog-Szemerédi-Gowers theorem.
-We do not know whether such a generalization has been studied in the literature, so we provide a complete proof in this section.
+In this section, we follow the method introduced by Tao in [36, Theorem 3.1] to obtain a generalized discretized entropy version of the Balog-Szemerédi-Gowers theorem with weakly bi-Lipschitz functions.
 Theorem 3.1.
@@
 we have that
-$H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|X^{\prime},Y^{\prime})-\frac{H_{n}(X)+H_{n}(Y)}{2}\leq 4\log K_{1}+3\log K_{2}+O(\log\kappa)+O(\log d).$ (24)
+$H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|X^{\prime},Y^{\prime})-\frac{H_{n}(X)+H_{n}(Y)}{2}\leq 4\log K_{1}+3\log K_{2}+O(\log\kappa+\log d).$ (25)
 Here $\log K_{1},\,\log K_{2},\,\log\kappa\,\gg 1$ and $d\in\mathbb{N}$.
@@
 Under the notations and conditions stated in Theorem 3.1, we have
-$H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa)+O(\log d).$
+$H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa+\log d).$
 Proof.
@@
 $H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})+H_{n}(X^{\prime},X^{\prime\prime},Z,Y^{\prime})$
-$\qquad=H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})+H_{n}(X^{\prime},X^{\prime\prime},g(Z),Y^{\prime})+O(\log\kappa)+O(\log d)$
-$\qquad\leq H_{n}(f(X^{\prime})+g(Z),f(X^{\prime\prime})+g(Z),Y^{\prime})+H_{n}(X^{\prime},X^{\prime\prime},Y^{\prime})+O(\log\kappa)+O(\log d).$ (25)
+$\qquad=H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})+H_{n}(X^{\prime},X^{\prime\prime},g(Z),Y^{\prime})+O(\log\kappa+\log d)$
+$\qquad\leq H_{n}(f(X^{\prime})+g(Z),f(X^{\prime\prime})+g(Z),Y^{\prime})+H_{n}(X^{\prime},X^{\prime\prime},Y^{\prime})+O(\log\kappa+\log d).$ (26)
 Indeed, we have $(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})=f_{1}(f(X^{\prime})+g(Z),f(X^{\prime\prime})+g(Z),Y^{\prime})=f_{2}(X^{\prime},X^{\prime\prime},Y^{\prime})$ and and
@@
 $H_{n}(f(X^{\prime})-f(X^{\prime\prime}),Y^{\prime})$
-$\qquad\leq 2H_{n}(f(X)+g(Y))+2H_{n}(X)+2H_{n}(Y)-2H_{n}(X,Y)+O(\log\kappa)+O(\log d).$
+$\qquad\leq 2H_{n}(f(X)+g(Y))+2H_{n}(X)+2H_{n}(Y)-2H_{n}(X,Y)+O(\log\kappa+\log d).$
 $\qquad\leq\big(2H_{n}(f(X)+g(Y))-H_{n}(X)-H_{n}(Y)\big)$
-$\qquad\qquad+2\big(H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\big)+H_{n}(X)+H_{n}(Y)+O(\log\kappa)+O(\log d).$
-$\qquad\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa)+O(\log d).$
+$\qquad\qquad+2\big(H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\big)+H_{n}(X)+H_{n}(Y)+O(\log\kappa+\log d).$
+$\qquad\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa+\log d).$
 We now present the proof of Theorem 3.1.
@@
 $H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime}),Y^{\prime})+H_{n}(X^{\prime},Y^{\prime},X^{\prime\prime},Y^{\prime\prime})$
-$\qquad=H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime}),Y^{\prime})+H_{n}(f(X^{\prime}),Y^{\prime},X^{\prime\prime},Y^{\prime\prime})+O(\log\kappa)+O(\log d)$
-$\qquad\leq H_{n}(f(X^{\prime\prime})-f(X^{\prime}),f(X^{\prime})+g(Y^{\prime\prime}),Y^{\prime})+H_{n}(X^{\prime\prime},Y^{\prime\prime},Y^{\prime})+O(\log\kappa).$ (31)
+$\qquad=H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime}),Y^{\prime})+H_{n}(f(X^{\prime}),Y^{\prime},X^{\prime\prime},Y^{\prime\prime})+O(\log\kappa+\log d)$
+$\qquad\leq H_{n}(f(X^{\prime\prime})-f(X^{\prime}),f(X^{\prime})+g(Y^{\prime\prime}),Y^{\prime})+H_{n}(X^{\prime\prime},Y^{\prime\prime},Y^{\prime})+O(\log\kappa+\log d).$ (32)
 Recall that $D_{n}(X^{\prime\prime})$ and $D_{n}(Y^{\prime\prime})$ are conditionally independent given $D_{n}(X^{\prime},Y^{\prime})$. Using Lemma 2.8 and the assumption (21), one has
@@
 $\qquad\leq H_{n}(f(X^{\prime\prime})-f(X^{\prime}),Y^{\prime})+H_{n}(f(X^{\prime})+g(Y^{\prime\prime}))$
-$\qquad\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa)+O(\log d)$ (34)
+$\qquad\leq H_{n}(X)+H_{n}(Y)+2\log K_{1}+2\log K_{2}+O(\log\kappa+\log d)$ (35)
 $\qquad\qquad\qquad+\frac{H_{n}(X)+H_{n}(Y)}{2}+\log K_{2}$
-$\qquad=\frac{3}{2}(H_{n}(X)+H_{n}(Y))+2\log K_{1}+3\log K_{2}+O(\log\kappa)+O(\log d).$ (35)
+$\qquad=\frac{3}{2}(H_{n}(X)+H_{n}(Y))+2\log K_{1}+3\log K_{2}+O(\log\kappa+\log d).$ (36)
 Combining (31)–(36), one concludes that
-$H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|X^{\prime},Y^{\prime})\leq\frac{H_{n}(X)+H_{n}(Y)}{2}+4\log K_{1}+3\log K_{2}+O(\log\kappa)+O(\log d).$
+$H_{n}(f(X^{\prime\prime})+g(Y^{\prime\prime})|X^{\prime},Y^{\prime})\leq\frac{H_{n}(X)+H_{n}(Y)}{2}+4\log K_{1}+3\log K_{2}+O(\log\kappa+\log d).$
 Thus, the estimate (25) follows. ∎
-4 Proof of Theorems 1.21.4 with $s>1/2$
+4 Proof of Theorems 1.2 and 1.3 with $s>1/2$
 4.1 Diagonal quadratics with independent random variables
@@
 Write $\phi(x,y)=\rho_{1}(x)+\rho_{2}(y)$, where $\rho_{1}(x)=\alpha x^{2}$ and $\rho_{2}(y)=\beta y^{2}$. One asserts by Example 2.12 and Lemma 2.16 that
-$\begin{split}H_{n}(\rho_{1}(X))&=H_{n}(X)+O\big(\log_{+}c_{1}+\log_{+}(1/c_{2})+\big|\log|\alpha|\big|+1\big)=O(\log\mathfrak{K}),\\ H_{n}(\rho_{2}(Y))&=H_{n}(Y)+O\big(\log_{+}c_{1}+\log_{+}(1/c_{2})+\big|\log|\beta|\big|+1\big)=O(\log\mathfrak{K}).\end{split}$ (36)
+$\begin{split}H_{n}(\rho_{1}(X))&=H_{n}(X)+O\big(\log_{+}c_{1}+\log_{+}(1/c_{2})+\big|\log|\alpha|\big|+1\big)=H_{n}(X)+O(\log\mathfrak{K}),\\ H_{n}(\rho_{2}(Y))&=H_{n}(Y)+O\big(\log_{+}c_{1}+\log_{+}(1/c_{2})+\big|\log|\beta|\big|+1\big)=H_{n}(Y)+O(\log\mathfrak{K}).\end{split}$ (37)
 Define $W=(\sqrt{|\alpha|}X,\sqrt{|\beta|}Y)$.
@@
 Combining (38)–(41) and Lemma 2.4(iv), we have that
-$H_{n}\big(|W_{1}-W_{2}|^{2}\big)+5H_{n}(X)+5H_{n}(Y)\leq 8H_{n}(X+Y)+3H_{n}(\phi(X,Y))+O(\log\mathfrak{K}).$
+$H_{n}\big(|W_{1}-W_{2}|^{2}\big)+5H_{n}(X)+5H_{n}(Y)\leq 8H_{n}(X+Y)+3H_{n}(\phi(X,Y))+O(\log\mathfrak{K}).$ (42)
 Since $X$ and $Y$ are independent, and are $(s_{1},C_{1})$-Frostman and $(s_{2},C_{2})$-Frostman, respectively, one asserts that $W$ is $(s_{1}+s_{2},C_{1}C_{2}|\alpha|^{-s_{1}/2}|\beta|^{-s_{2}/2})$-Frostman.
-Here $\log(|\alpha|^{-s_{1}/2}|\beta|^{-s_{2}/2})=O(\log\mathfrak{K})$ since $s_{1},s_{2}$ are bounded.
-Noting that $range(W)$ is contained in the ball in ${\mathbb{R}}^{2}$ centered at origin with radius $O\big(\sqrt{|\alpha|}c_{1}+\sqrt{|\beta|}c_{1}\big)$.
-So we can deduce by Lemma 2.32 and Theorem 2.41 that
-$\begin{split}&H_{n}\big(X\big)\geq ns_{1}-\log C_{1}+O(1),\quad H_{n}\big(Y\big)\geq ns_{2}-\log C_{2}+O(1),\\ &\qquad\qquad\qquad H_{n}\big(|W_{1}-W_{2}|^{2}\big)\geq n(s+\epsilon_{0})+O(\log\mathfrak{K})\end{split}$ (41)
-for some $\epsilon_{0}=\epsilon_{0}(s)>0$ and all $n\in\mathbb{N}$. It follows that
-$n\left(s+\epsilon_{0}+5s_{1}+5s_{2}\right)\leq 8H_{n}(X+Y)+3H_{n}(\phi(X,Y))$
-$\qquad\qquad+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O(\log\mathfrak{K})+O_{s}(1).$
-Recalling that $s_{1}+s_{2}\geq 2s$, we conclude that
+Here $\log(|\alpha|^{-s_{1}/2}|\beta|^{-s_{2}/2})=O(\big|\log|\alpha|\big|+\big|\log|\beta|\big|)$ since $s_{1},s_{2}$ are bounded.
+Notice that $range(W)$ is contained in the ball in ${\mathbb{R}}^{2}$ centered at origin with radius $O\big((\sqrt{|\alpha|}+\sqrt{|\beta|})c_{1}\big)$.
+Also recall $s_{1}+s_{2}\geq 2s$.
+We can deduce by Lemma 2.32 and Theorem 2.41 that
+$\begin{split}H_{n}(X)+H_{n}(Y)&\geq 2ns-\log C_{1}C_{2}+O(1),\\ H_{n}\big(|W_{1}-W_{2}|^{2}\big)&\geq n(s+\epsilon_{0})+O(\log\mathfrak{K})\end{split}$ (43)
+for some $\epsilon_{0}=\epsilon_{0}(s)>0$ and all $n\in\mathbb{N}$. It follows from (42) and (43) that
+$n\left(11s+\epsilon_{0}\right)\leq 8H_{n}(X+Y)+3H_{n}(\phi(X,Y))+O(\log\mathfrak{K}).$
+We conclude that
 $\max\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\}\geq n\left(s+\frac{\epsilon_{0}}{11}\right)+O(\log\mathfrak{K}).$ (44)
@@
 $\qquad\geq\min\Big\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,\frac{H_{n}(\alpha X^{2})+H_{n}(\beta Y^{2})}{2}\Big\}+O(1)$
-$\qquad\geq\frac{H_{n}(X)+H_{n}(Y)}{2}+O(\log\mathfrak{K})\geq ns+O(\log\mathfrak{K}).$ (43)
-Putting (42) and (43) together, we complete the proof of the lemma with constant $\epsilon_{1}=\epsilon_{0}/11$.
+$\qquad=\frac{H_{n}(X)+H_{n}(Y)}{2}+O(\log\mathfrak{K})\geq ns+O(\log\mathfrak{K}).$ (45)
+Putting (44) and (45) together, we complete the proof of the lemma with $\epsilon_{1}=\epsilon_{0}/11$.
 Next, if $\alpha$ and $\beta$ have different signs, the result will follow with a slightly worse constant. Without loss of generality, we assume that $\alpha>0>\beta$. Then
@@
 Compared with (41), Lemma 2.29 shows that
-$H_{n}(\rho_{1}(X_{1})+\rho_{1}(X_{2})-\rho_{2}(Y_{1})-\rho_{2}(Y_{2}))\leq 9H_{n}(\rho_{1}(X)+\rho_{2}(Y))$
-$\qquad\qquad\qquad-4H_{n}(X)-4H_{n}(Y)+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O(\log\mathfrak{K}).$
+$H_{n}(\rho_{1}(X_{1})+\rho_{1}(X_{2})-\rho_{2}(Y_{1})-\rho_{2}(Y_{2}))$
+$\qquad\leq 9H_{n}(\rho_{1}(X)+\rho_{2}(Y))-4H_{n}(X)-4H_{n}(Y)+O(\log\mathfrak{K}).$
 It leads to
-$H_{n}\big(|W_{1}-W_{2}|^{2}\big)+$ $8H_{n}(X)+8H_{n}(Y)\leq 8H_{n}(X+Y)+9H_{n}(\phi(X,Y))$
-$\quad+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O(\log\mathfrak{K}),$
+$H_{n}\big(|W_{1}-W_{2}|^{2}\big)+8H_{n}(X)+8H_{n}(Y)\leq 8H_{n}(X+Y)+9H_{n}(\phi(X,Y))+O(\log\mathfrak{K}),$
 and then
-$n\left(s+\epsilon_{0}+8s_{1}+8s_{2}\right)\leq 8H_{n}(X+Y)+9H_{n}(\phi(X,Y))$
-$\qquad\qquad+O(\log_{+}C_{1}C_{2})+O(\log_{+}c_{1})+O(\log_{+}(1/c_{2}))+O(\log\mathfrak{K})+O_{s}(1).$
-Finally, the lemma will follow with constant $\epsilon_{1}=\epsilon_{0}/17$. ∎
+$n\left(17s+\epsilon_{0}\right)\leq 8H_{n}(X+Y)+9H_{n}(\phi(X,Y))+O(\log\mathfrak{K}).$
+Finally, the lemma will follow with $\epsilon_{1}=\epsilon_{0}/17$. ∎
 Remark 4.2.
@@
 Let
-$\rho_{1}(x)=\alpha_{2}x^{2}+\alpha_{1}x+\alpha_{0},\quad\rho_{2}(y)=\beta_{2}x^{2}+\beta_{1}x+\beta_{0},$
+$\rho_{1}(x)=\alpha_{2}x^{2}+\alpha_{1}x+\alpha_{0},\quad\rho_{2}(y)=\beta_{2}y^{2}+\beta_{1}y+\beta_{0},$
 where $\alpha_{2},\beta_{2}\in{\mathbb{R}}^{\ast}$ and $\alpha_{1},\alpha_{0},\beta_{1},\beta_{0}\in{\mathbb{R}}$.
@@
 $\log\mathfrak{K}$ $=O_{s}\big(\log_{+}C_{1}+\log_{+}C_{2}+1\big)$
-$+O\big(|\log|\alpha_{2}|\big|+\big|\log|\beta_{2}|\big|+\log_{+}|\alpha_{1}|+\log_{+}|\beta_{1}|+\log_{+}c).$
+$+O\big(|\log|\alpha_{2}|\big|+\big|\log|\beta_{2}|\big|+\log_{+}|\alpha_{1}|+\log_{+}|\beta_{1}|+\log_{+}c\big).$
+Here, and from now on, the notation $1\ll_{s}s_{1},s_{2}$ means that $s_{1}$ and $s_{2}$ are bounded below by some positive constant depending only on $s$, but not on other quantities such as $\rho_{1},\rho_{2},C_{1},C_{2},c$ or $n$.
 Proof.
-Note that by completing the squares, we have
-$\rho_{1}(x)=\alpha_{2}\left(x+a\right)^{2}+\alpha_{0}-\frac{\alpha_{1}^{2}}{4\alpha_{2}},\quad\rho_{2}(y)=\beta_{2}\left(x+b\right)^{2}+\beta_{0}-\frac{\beta_{1}^{2}}{4\beta_{2}},$
+By completing the squares, we have
+$\rho_{1}(x)=\alpha_{2}\left(x+a\right)^{2}+\alpha_{0}-\frac{\alpha_{1}^{2}}{4\alpha_{2}},\quad\rho_{2}(y)=\beta_{2}\left(y+b\right)^{2}+\beta_{0}-\frac{\beta_{1}^{2}}{4\beta_{2}},$
 where $a=\frac{\alpha_{1}}{2\alpha_{2}}$ and $b=\frac{\beta_{1}}{2\beta_{2}}$. By Lemma 2.16,
 $\begin{split}H_{n}(X+Y)&=H_{n}((X+a)+(Y+b))+O(1),\\ H_{n}(\rho_{1}(X)+\rho_{2}(Y))&=H_{n}(\alpha_{2}(X+a)^{2}+\beta_{2}(Y+b)^{2})+O(1).\end{split}$ (46)
-Let $\eta_{i}$ $(i=1,2)$ be parameters to be determined later, with $0<\eta_{i}\leq\min\{(4C_{i})^{-1/s_{i}},\,1/10\}$.
+Let $\eta_{i}$ $(i=1,2)$ be parameters to be determined later, with $0<\eta_{i}\leq\min\big\{(4C_{i})^{-1/s_{i}},\,1/10\big\}$.
 Denote by $E$ the event that $|X+a|\geq\eta_{1}$, $|Y+b|\geq\eta_{2}$.
@@
 with some $c_{1}$ satisfying $c_{1}\leq c+\max\{|a|,|b|\}$. In particular, we have $\log_{+}c_{1}=O(\log\mathfrak{K})$. By Lemma 4.1, there is some $\epsilon_{1}=\epsilon_{1}(s)\in(0,1)$ such that
-$\begin{split}&H_{n}(X_{1}+Y_{1})+H_{n}(\alpha_{2}X_{1}^{2}+\beta_{2}Y_{1}^{2})\geq n(2s+\epsilon_{1})\\ &\qquad\qquad\qquad+O(\log(1/\eta_{1}))+O(\log(1/\eta_{2}))+O(\log\mathfrak{K}).\end{split}$ (47)
+$\begin{split}&H_{n}(X_{1}+Y_{1})+H_{n}(\alpha_{2}X_{1}^{2}+\beta_{2}Y_{1}^{2})\geq n(2s+\epsilon_{1})+O\big(\log(1/\eta_{1})+\log(1/\eta_{2})+\log\mathfrak{K}\big).\end{split}$ (49)
 Combining (48) and (49), one deduces that
-$\begin{split}&H_{n}(X+Y)+H_{n}(\rho_{1}(X)+\rho_{2}(Y))\geq n(2s+\epsilon_{1})+O(\log\mathfrak{K})\\ &\qquad\qquad\qquad+O\big(C_{1}\eta_{1}^{s_{1}}n+C_{2}\eta_{2}^{s_{2}}n+\log(1/\eta_{1})+\log(1/\eta_{2})\big).\end{split}$
-Recall that we require $0<\eta_{i}<\min\{(4C_{i})^{-1/s_{i}},1/10\}$, $0<\epsilon_{1}<1$, and $\epsilon_{1}$ only depends on $s$.
-Now assume that the implied constant in the last term on the right-hand side of above inequality is $C^{\prime}$, with $C^{\prime}\geq 1$.
+$\begin{split}&H_{n}(X+Y)+H_{n}(\rho_{1}(X)+\rho_{2}(Y))\geq n(2s+\epsilon_{1})+O(\log\mathfrak{K})\\ &\qquad\qquad\qquad\qquad+O\big(C_{1}\eta_{1}^{s_{1}}n+C_{2}\eta_{2}^{s_{2}}n+\log(1/\eta_{1})+\log(1/\eta_{2})\big).\end{split}$ (50)
+Recall that we require $0<\eta_{i}<\min\big\{(4C_{i})^{-1/s_{i}},1/10\big\}$, $0<\epsilon_{1}<1$, and $\epsilon_{1}$ only depends on $s$.
+Now assume that the implied constant in the last term on the right-hand side of (50) is $C^{\prime}$, with $C^{\prime}\geq 1$.
 We then take $\eta_{i}=\min\Big\{\big(\frac{\epsilon_{1}}{4C^{\prime}C_{i}}\big)^{1/s_{i}},\,\frac{1}{10}\Big\}$ $(i=1,2)$.
-Then $C^{\prime}C_{i}\eta_{i}^{s_{i}}n\leq n\epsilon_{1}/4$, and $O(\log_{+}(1/\eta_{i}))=O(\frac{1}{s_{i}}\log_{+}C_{i})+O_{s}(1)$.
+Then $C^{\prime}C_{i}\eta_{i}^{s_{i}}n\leq n\epsilon_{1}/4$, and $O(\log_{+}(1/\eta_{i}))=O\big(\frac{1}{s_{i}}\log_{+}C_{i}\big)+O_{s}\big(\frac{1}{s_{i}}\big)$.
 Recall $s_{1},s_{2}\gg s$.
@@
 $\mathfrak{c}(\rho)=\max\{|\alpha_{d}|^{-1},\,|\alpha_{d}|,\,|\alpha_{d-1}|,\,\ldots,\,|\alpha_{1}|,\,|\alpha_{0}|\}.$
-This quantity controls both the scale and inverse scale of $\rho$.
 For any $z\in{\mathbb{R}}$, define the difference polynomial $\Delta_{z}\rho$ to be
-$(\Delta_{z}\rho)(x)=\rho(x+z)-\rho(z)=\sum\limits_{j=0}^{d-1}\left(\sum\limits_{k=j+1}^{d}{k\choose j}\alpha_{k}z^{k-j}\right)x^{j},$
-which is a polynomial of degree d1d-1. Also assume that
+$(\Delta_{z}\rho)(x)=\rho(x+z)-\rho(x)=\sum\limits_{j=0}^{d-1}\left(\sum\limits_{k=j+1}^{d}{k\choose j}\alpha_{k}z^{k-j}\right)x^{j},$
+which is a polynomial in $x$ of degree d1d-1. Also assume that
 $\rho(x)=\alpha_{d}\prod\limits_{i=1}^{l}(x-z_{i})^{m_{i}},$
@@
 Let $\rho$ be a real polynomial of degree $d\geq 1$. Let $\eta,c>0$.
-(i) The set ${\mathbb{R}}\setminus\mathcal{R}(\rho;\eta)$ is a union of at most $d$ open intervals. Each has radius at most $2\eta$.
+(i) The set ${\mathbb{R}}\setminus\mathcal{R}(\rho;\eta)$ is a union of at most $d$ open intervals. Each has length at most $2\eta$.
 (ii) For all $x\in\mathcal{R}(\rho;\eta)$, we have $|\rho(x)|\geq|\alpha_{d}|\eta^{d}\geq\mathfrak{c}(\rho)^{-1}\eta^{d}$.
@@
 Lemma 4.5.
-Let $\rho$ be a real polynomial of degree $d\geq 2$. The its derivative satisfies
-$\log\mathfrak{c}(\rho^{\prime})=\log\mathfrak{c}(\rho)+O_{d}(1)$
+Let $\rho$ be a real polynomial of degree $d\geq 2$. Then its derivative satisfies
+$\log\mathfrak{c}(\rho^{\prime})\leq\log\mathfrak{c}(\rho)+O_{d}(1).$
 Lemma 4.6.
 Let $\rho$ be a real polynomial of degree $d\geq 3$. Then for all $z\in{\mathbb{R}}^{\ast}$,
-$\log\mathfrak{c}(\Delta_{z}\rho)=\log\mathfrak{c}(\rho)+O\big(\big|\log|z|\big|\big)+O_{d}(1).$
-We will apply the following corollary, which is a direct consequence of Lemmas 4.4 and 4.5.
-Corollary 4.7.
-Let $\rho$ be a real polynomial of degree $d\geq 2$.
-(i) The set ${\mathbb{R}}\setminus\mathcal{R}(\rho^{\prime};\eta)$ is a union of at most d1d-1 open intervals. Each has length at most $2\eta$. And $\rho$ is monotonic on each of the intervals.
-(ii) Let $0<\eta<1$. Then for any $x\in\mathcal{R}(\rho;\eta)$,
+$\log\mathfrak{c}(\Delta_{z}\rho)\leq\log\mathfrak{c}(\rho)+O_{d}\big(\big|\log|z|\big|+1\big).$
+The following lemma is important for our application.
+Lemma 4.7.
+Let $\rho$ be a real polynomial of degree $d\geq 2$. Suppose that $0<\eta<1$ and $c>0$.
+(i) The set ${\mathbb{R}}\setminus\mathcal{R}(\rho^{\prime};\eta)$ is a union of at most d1d-1 open intervals. Each has length at most $2\eta$.
+(ii) The set $\mathcal{R}(\rho^{\prime};\eta)\cap[-c,c]$, if non-empty, is a union of at most $d$ closed intervals (including singleton set). And $\rho$ is monotonic on each of the intervals.
+(iii) For any $x\in\mathcal{R}(\rho^{\prime};\eta)$,
 $\log_{+}\frac{1}{|\rho^{\prime}(x)|}\leq(d-1)\log(1/\eta)+\log\mathfrak{c}(\rho)+O_{d}(1).$
-(iii) Let $c>0$. Then for any $x\in[-c,c]$,
+(iv) For any $x\in[-c,c]$,
 $\log_{+}|\rho^{\prime}(x)|\leq(d-1)\log_{+}c+\log\mathfrak{c}(\rho)+O_{d}(1).$
+(v) The function $\rho$ is weakly $(\kappa,d)$-bi-Lipschitz on the region $\mathcal{R}(\rho^{\prime};\eta)\cap[-c,c]$, with
+$\log\kappa\leq\log\mathfrak{c}(\rho)+O_{d}\big(\log_{+}c+\log(1/\eta)+1\big),$
+if this region is non-empty.
+Proof.
+The statements (i)-(iv) are direct consequences of Lemmas 4.4 and 4.5. The statement (v) follows by combining (ii)-(iv) and Lemma 2.13. ∎
 Lemma 4.8.
@@
 for all $n\in\mathbb{N}$, where
-$O(\log\mathfrak{K})=O_{d,s}\big(\log_{+}C_{1}+\log_{+}C_{2}+1\big)+O_{d}\big(\log\mathfrak{c}(\rho_{1})+\log\mathfrak{c}(\rho_{2})+\log_{+}c\big).$
+$\log\mathfrak{K}=O_{d,s}\big(\log_{+}C_{1}+\log_{+}C_{2}+1\big)+O_{d}\big(\log\mathfrak{c}(\rho_{1})+\log\mathfrak{c}(\rho_{2})+\log_{+}c\big).$
 Proof.
@@
 Without loss of generality, we assume that $0<\epsilon_{1}^{\prime}<1$.
-For $i=1,2$, let $\eta_{i}=\min\Big\{\big(\frac{\epsilon_{1}}{128dC_{i}}\big)^{1/s_{i}},\,\frac{1}{10}\Big\}$, which satisfies $dC_{i}\eta_{i}^{s_{i}}\leq\epsilon_{1}^{\prime}/128\leq 1/4$ and $\log(1/\eta_{i})=O_{s}(\log_{+}C_{i})+O_{d,s}(1)$.
+To quantify the entropy gained beyond the trivial amount $2ns$, we write the entropy increment as
+$\Psi:=H_{n}(X+Y)+H_{n}\big(\rho_{1}(X)+\rho_{2}(Y)\big)-2ns.$ (52)
+Our goal is to show $\Psi\geq n\epsilon_{1}^{\prime\prime}+O(\log\mathfrak{K})$ for some $\epsilon_{1}^{\prime\prime}=\epsilon_{1}^{\prime\prime}(d,s)>0$.
+Let $\mathfrak{r}=\mathfrak{r}(d,s)\geq 1$ be a parameter to be determined later.
+For $i=1,2$, let $\eta_{i}=\min\Big\{\big(\frac{\epsilon_{1}^{\prime}}{4\mathfrak{r}dC_{i}}\big)^{1/s_{i}},\,\frac{1}{10}\Big\}$, which satisfies $dC_{i}\eta_{i}^{s_{i}}\leq\epsilon_{1}^{\prime}/4\mathfrak{r}\leq 1/4$ and $\log(1/\eta_{i})=O_{s}(\log_{+}C_{i})+O_{d,s}(1)=O(\log\mathfrak{K})$.
 Let $E$ be the event that $X\in\mathcal{R}(\rho_{1}^{\prime};\eta_{1})$, $Y\in\mathcal{R}(\rho_{2}^{\prime};\eta_{2})$.
-On $E$, both $X$ and $Y$ avoid the $\eta$-neighborhoods of the critical points of $\rho_{1}$ and $\rho_{2}$, ensuring that $\rho_{1}^{\prime}(X)$ and $\rho_{2}^{\prime}(Y)$ are bounded away from zero.
-By Lemma 4.4(i), the complement ${\mathbb{R}}\setminus\mathcal{R}(\rho_{i}^{\prime};\eta_{i})$ is a union of d1d-1 intervals, each of length $2\eta_{i}$.
+By Lemma 4.7(i), the complement ${\mathbb{R}}\setminus\mathcal{R}(\rho_{i}^{\prime};\eta_{i})$ is a union of at most d1d-1 intervals, each of length at most $2\eta_{i}$.
 Since $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, we have
 $P(E)$ $\geq 1-P(X\notin\mathcal{R}(\rho_{1}^{\prime};\eta_{1}))-P(Y\notin\mathcal{R}(\rho_{2}^{\prime};\eta_{2}))$
-$\geq 1-(d-1)C_{1}\eta_{1}^{s_{1}}-(d-1)C_{2}\eta_{2}^{s_{2}}\geq 1-\epsilon_{1}/64\geq 1/2.$
-We denote the conditional probability given $E$ by $P_{E}$.
-Let $(X_{0},Y_{0})$ be the random variable such that $(X_{0},Y_{0})\equiv(X,Y|E)$ (see Lemma A.2).
-One can check that $X_{0}$ and $Y_{0}$ are still independent.
-And $(X_{0},Y_{0})$ is marginally $(s_{1},O(C_{1});s_{2},O(C_{2}))$-Frostman.
-By Lemma 2.32(i),
+$\geq 1-(d-1)C_{1}\eta_{1}^{s_{1}}-(d-1)C_{2}\eta_{2}^{s_{2}}\geq 1-\epsilon_{1}^{\prime}/2\mathfrak{r}\geq 1/2.$
+We denote the conditional probability given $E$ by $P_{E}$. Let $(X_{0},Y_{0})$ be the random variable such that $(X_{0},Y_{0})\equiv(X,Y|E)$ (see Lemma A.2). Combining (52), one obtains
+$H_{n}(X_{0}+Y_{0})+H_{n}(\rho_{1}(X_{0})+\rho_{2}(Y_{0}))$
+$\qquad\leq\big(1+O(\epsilon_{1}^{\prime}/\mathfrak{r})\big)\big(H_{n}(X+Y)+H_{n}(\rho_{1}(X)+\rho_{2}(Y))\big)$
+$\qquad\leq 2ns+O(n\epsilon_{1}^{\prime}/\mathfrak{r}+\Psi),$ (53)
+where we have used the estimate $P(E)^{-1}\leq(1-\epsilon_{1}^{\prime}/2\mathfrak{r})^{-1}\leq 1+O(\epsilon_{1}^{\prime}/\mathfrak{r})$.
+One can also check that $X_{0}$ and $Y_{0}$ are still independent. And $(X_{0},Y_{0})$ is marginally $(s_{1},O(C_{1});s_{2},O(C_{2}))$-Frostman. In view of Lemma 2.32(i),
 $H_{n}(X_{0})+H_{n}(Y_{0})\geq n(s_{1}+s_{2})-\log C_{1}C_{2}+O(1)\geq 2ns+O(\log\mathfrak{K}).$ (54)
-By (7) and Lemma 4.4, one has
-$H_{n}(X+Y)\leq n+\log_{+}c+O(1)=n+O(\log\mathfrak{K}),$
-$H_{n}(\rho_{1}(X)+\rho_{2}(Y))\leq n+\log\mathfrak{c}(\rho)+O_{d}(\log_{+}c+1)=n+O(\log\mathfrak{K}).$
-Together with Lemma A.2, one obtains
-$H_{n}(X_{0}+Y_{0})\leq(1+\epsilon_{1}^{\prime}/32)H_{n}(X+Y)\leq H_{n}(X+Y)+n\epsilon_{1}^{\prime}/32+O(\log\mathfrak{K}),$ (50)
-and
-$\begin{split}H_{n}(\rho_{1}(X_{0})+\rho_{2}(Y_{0}))&\leq(1+\epsilon_{1}^{\prime}/32)H_{n}(\rho_{1}(X)+\rho_{2}(Y))\\ &\leq H_{n}(\rho_{1}(X)+\rho_{2}(Y))+n\epsilon_{1}^{\prime}/32+O(\log\mathfrak{K}),\end{split}$ (51)
-where we have used the estimate $P(E)^{-1}\leq(1-\epsilon_{1}^{\prime}/64)^{-1}\leq 1+\epsilon_{1}^{\prime}/32$.
+It follows from (10) that
+$H_{n}(X_{0}+Y_{0})\geq\frac{H_{n}(X_{0})+H_{n}(Y_{0})}{2}+O(1)\geq ns+O(\log\mathfrak{K}).$ (55)
 Note that $range_{+}(X_{0})\subseteq\mathcal{R}(\rho_{1}^{\prime};\eta)\cap[-c,c]$ and $range_{+}(Y_{0})\subseteq\mathcal{R}(\rho_{2}^{\prime};\eta)\cap[-c,c]$.
-By Lemmas 2.13, and Corollary 4.7, we see that $\rho_{1}$ and $\rho_{2}$ are weakly $(\kappa,d-1)$-bi-Lipschitz on $range_{+}(X_{0})$ and $range_{+}(Y_{0})$, respectively, with $\log\kappa=O(\log\mathfrak{K})$.
-Now, let $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ be two i.i.d. copies of $(X_{0},Y_{0})$. In particular, $X_{1},X_{2},Y_{1},Y_{2}$ are independent. Combining Lemma 2.29 and (49),
+By Lemma 4.7(v), we see that $\rho_{1}$ and $\rho_{2}$ are weakly $(\kappa,d)$-bi-Lipschitz on $range_{+}(X_{0})$ and $range_{+}(Y_{0})$, respectively, with $\log\kappa=O(\log\mathfrak{K})$.
+Then we also have
+$H_{n}(\rho_{1}(X_{0})+\rho_{2}(Y_{0}))\geq\frac{H_{n}(\rho_{1}(X_{0}))+H_{n}(\rho_{2}(Y_{0}))}{2}+O(1)$
+$\qquad=\frac{H_{n}(X_{0})+H_{n}(Y_{0})}{2}+O(\log\mathfrak{K})\geq ns+O(\log\mathfrak{K}).$ (56)
+Combining (53), (55) and (56), one obtains
+$H_{n}(X_{0}+Y_{0}),\,H_{n}(\rho_{1}(X_{0})+\rho_{2}(Y_{0}))\,\leq ns+O(n\epsilon_{1}^{\prime}/\mathfrak{r}+\Psi+\log\mathfrak{K}).$ (57)
+Now, let $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ be two i.i.d. copies of $(X_{0},Y_{0})$. In particular, $X_{1},X_{2},Y_{1},Y_{2}$ are independent. By Lemma 2.21,
+$H_{n}(X_{1}-X_{2})$ $\leq H_{n}(X_{1}+Y_{1})+H_{n}(X_{2}+Y_{1})-H_{n}(Y_{1})+O(1)$
+$=2H_{n}(X_{0}+Y_{0})-H_{n}(Y_{0})+O(1).$
+Similarly, one has $H_{n}(Y_{1}-Y_{2})\leq 2H_{n}(X_{0}+Y_{0})-H(X_{0})+O(1)$. Combining these two inequalities together with (54) and (57), we get
+$H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})\leq 2ns+O(n\epsilon_{1}^{\prime}/\mathfrak{r}+\Psi+\log\mathfrak{K}).$ (58)
+Combining Lemma 2.29, (54) and (55), one sees that
 $H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2}))$
-$\qquad\leq 7H_{n}(\rho_{1}(X_{0})+\rho_{2}(Y_{0}))-3H_{n}(X_{0})-3H_{n}(Y_{0})+O(\log\mathfrak{K})$
-$\qquad\leq 7H_{n}(\rho_{1}(X_{0})+\rho_{2}(Y_{0}))-6ns+O(\log\mathfrak{K}).$ (52)
+$\qquad\leq 7H_{n}(\rho_{1}(X_{0})+\rho_{2}(Y_{0}))-3H_{n}(X_{0})-3H_{n}(Y_{0})+O(\log\kappa+\log d)$
+$\qquad\leq ns+O(n\epsilon_{1}^{\prime}/\mathfrak{r}+\Psi+\log\mathfrak{K}).$ (59)
 We then examine the discretized conditional entropy
-$H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2})|X_{1}-X_{2},Y_{1}-Y_{2})+H_{n}(X_{2}+Y_{2}|X_{1}-X_{2},Y_{1}-Y_{2})$
-$\qquad\qquad\qquad\leq H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2}))+H_{n}(X_{2}+Y_{2}).$ (53)
+$\begin{split}&H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2})|X_{1}-X_{2},Y_{1}-Y_{2})+H_{n}(X_{2}+Y_{2}|X_{1}-X_{2},Y_{1}-Y_{2})\\ &\,\,\leq H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2}))+H_{n}(X_{2}+Y_{2})\leq 2ns+O(n\epsilon_{1}^{\prime}/\mathfrak{r}+\Psi+\log\mathfrak{K}),\end{split}$ (60)
+which follows from Lemma 2.4(iv), (57) and (4.2).
 In what follows, we condition on the discretized differences. For brevity, write
 $H_{n}(\cdot|z_{1},z_{2}):=H\big(D_{n}(\cdot)|D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2}\big)$
-for any random variable depending on $(X_{1},X_{2},Y_{1},Y_{2})$. We will sum over all dyadic rationals $z_{1},z_{2}$ with $P_{E}\big(D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2}\big)>0$.
-Given $D_{n}(X_{1}-X_{2})=z_{1}$ and $D_{n}(Y_{1}-Y_{2})=z_{2}$, it follows by Corollary 4.7 that
-$\begin{split}&\log_{+}\Big(2^{n}\big|\big(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2})\big)-\big(\rho_{1}(X_{2}+z_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{2}+z_{2})-\rho_{2}(Y_{2})\big)\big|\Big)\\ &\qquad\qquad\qquad=O_{d}(\log_{+}c)+O(\log\mathfrak{c}(\rho_{1}))+O(\log\mathfrak{c}(\rho_{2}))+O_{d}(1)=O(\log\mathfrak{K})\end{split}$ (54)
+for any random variable that is a function of $(X_{1},X_{2},Y_{1},Y_{2})$.
+We will sum over all dyadic rationals $z_{1},z_{2}$ with $P_{E}\big(D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2}\big)>0$.
+In particular, one has $z_{1},z_{2}\in[-2c-1,2c+1]$.
+Given $D_{n}(X_{1}-X_{2})=z_{1}$ and $D_{n}(Y_{1}-Y_{2})=z_{2}$, it follows by Lemma 4.7(iv) that
+$\log_{+}\Big(2^{n}\cdot\big|\big(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2})\big)-\big(\rho_{1}(X_{2}+z_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{2}+z_{2})-\rho_{2}(Y_{2})\big)\big|\Big)$
+$\qquad\qquad\qquad=O_{d}(\log_{+}c)+O(\log\mathfrak{c}(\rho_{1}))+O(\log\mathfrak{c}(\rho_{2}))+O_{d}(1)=O(\log\mathfrak{K})$ (61)
 with probability $1$. Recall that
@@
 We present the proof for $X_{2}$, as the one for $Y_{2}$ can be done similarly.
-Indeed, by Lemmas 2.30 and 2.31, the random variable $(X_{2},X_{1}-X_{2})$ is jointly $(s_{1},s_{1},O(C_{1}))$-Frostman.
+Indeed, by Lemmas 2.30 and 2.31, the random variable $(X_{2},X_{1}-X_{2})$ is jointly $(s_{1},s_{1},O(C_{1}^{2}))$-Frostman.
 Therefore, for any ball $B(x,r)$,
@@
 $\qquad=P_{E}(X_{2}\in B(x,r)|D_{n}(X_{1}-X_{2})=z_{1})\ll\frac{C_{1}^{2}2^{-ns_{1}}r^{s_{1}}}{P_{E}(D_{n}(X_{1}-X_{2})=z_{1})}.$
-So (56) follows.
-Now, for $i=1,2$, we take $\zeta_{i}=\min\Big\{\big(\frac{\epsilon_{1}^{\prime}}{400C_{i}}\big)^{1/s_{i}},\,1/16\Big\}$, which satisfies $\log(1/\zeta_{i})=O_{s}(\log_{+}C_{i})+O_{d,s}(1)=O(\log\mathfrak{K})$.
-If $n<\max\{\frac{1}{s_{1}}\log\frac{400C_{1}}{\epsilon_{1}^{\prime}},\,\frac{1}{s_{2}}\log\frac{400C_{2}}{\epsilon_{1}^{\prime}},\,4\}$, then either $\log C_{1}>ns_{1}+O_{d,s}(1)$ or $\log C_{2}>ns_{2}+O_{d,s}(1)$ or $n<4$.
-In view of $s_{1},s_{2}\gg s$, the term $O_{s}(\log_{+}C_{1})+O_{s}(\log_{+}C_{2})+O(1)$ may represent a negative number that is less than n-n, and (48) holds automatically.
-In the following, we assume that $n\geq\max\{\frac{1}{s_{1}}\log\frac{400C_{1}}{\epsilon_{1}^{\prime}},\,\frac{1}{s_{2}}\log\frac{400C_{2}}{\epsilon_{1}^{\prime}},\,4\}$.
-As a consequence, we have $\zeta_{i}\geq 2^{-ns_{i}-4}$ $(i=1,2)$.
-We first consider the case that $|z_{1}|\geq\zeta_{1}$ and $|z_{2}|\geq\zeta_{2}$. Note that $|z_{1}|,|z_{2}|\leq 2c$. In view of Lemma 4.6, the coefficients can be bounded by
-$\log\mathfrak{c}(\Delta_{z_{i}}\rho_{i})=\log\mathfrak{c}(\rho_{i})+O(\log_{+}c)+O(\log(1/\zeta_{i}))+O_{d}(1)=O(\log\mathfrak{K})$ (57)
+So (63) follows. We also mention that $\widetilde{C}_{i}\gg C_{i}$ $(i=1,2)$.
+Now, for $i=1,2$, we take $\zeta_{i}=\min\Big\{\big(\frac{\epsilon_{1}^{\prime}}{\mathfrak{r}C_{i}}\big)^{1/s_{i}},\,1/16\Big\}$, which satisfies $\log(1/\zeta_{i})=O_{s}(\log_{+}C_{i})+O_{d,s}(1)=O(\log\mathfrak{K})$.
+If $n<\max\{\frac{1}{s_{1}}\log\frac{\mathfrak{r}C_{1}}{\epsilon_{1}^{\prime}},\,\frac{1}{s_{2}}\log\frac{\mathfrak{r}C_{2}}{\epsilon_{1}^{\prime}},\,4\}$, then either $\log C_{1}>ns_{1}+O_{d,s}(1)$ or $\log C_{2}>ns_{2}+O_{d,s}(1)$ or $n<4$.
+In view of $s_{1},s_{2}\gg s$, the term $O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2}+1)$ may represent a negative number that is less than 10n-10\,n, and (51) holds automatically.
+In the following, we assume that $n\geq\max\{\frac{1}{s_{1}}\log\frac{\mathfrak{r}C_{1}}{\epsilon_{1}^{\prime}},\,\frac{1}{s_{2}}\log\frac{\mathfrak{r}C_{2}}{\epsilon_{1}^{\prime}},\,4\}$.
+As a consequence, we have $\zeta_{i}\geq 2^{-n-4}$ $(i=1,2)$.
+We first consider the case that $|z_{1}|\geq\zeta_{1}$ and $|z_{2}|\geq\zeta_{2}$. Recall that $|z_{1}|,|z_{2}|\leq 2c+1$. In view of Lemma 4.6, the coefficients can be bounded by
+$\log\mathfrak{c}(\Delta_{z_{i}}\rho_{i})\leq\log\mathfrak{c}(\rho_{i})+O_{d}\big(\log_{+}c+\log(1/\zeta_{i})+1\big)=O(\log\mathfrak{K})$ (64)
 for $i=1,2$. Applying the inductive hypothesis (for polynomials of degree d1d-1) with the conditional Frostman constants from (63), we obtain
-$\begin{split}&H_{n}(X_{2}+Y_{2}|z_{1},z_{2})+H_{n}((\Delta_{z_{1}}\rho_{1})(X_{2})+(\Delta_{z_{2}}\rho_{2})(Y_{2})|z_{1},z_{2})\geq n(2s+\epsilon_{1}^{\prime})\\ &\qquad+O_{d}(\log\mathfrak{c}(\Delta_{z}\rho_{1})+\log\mathfrak{c}(\Delta_{z}\rho_{2}))+\log_{+}c)+O_{d,s}\big(\log_{+}\widetilde{C}_{1}+\log_{+}\widetilde{C}_{2}+1\big).\end{split}$ (58)
+$\begin{split}&H_{n}(X_{2}+Y_{2}|z_{1},z_{2})+H_{n}\big((\Delta_{z_{1}}\rho_{1})(X_{2})+(\Delta_{z_{2}}\rho_{2})(Y_{2})|z_{1},z_{2}\big)\geq n(2s+\epsilon_{1}^{\prime})\\ &\qquad+O_{d,s}\big(\log_{+}\widetilde{C}_{1}+\log_{+}\widetilde{C}_{2}+1\big)+O_{d}\big(\log\mathfrak{c}(\Delta_{z}\rho_{1})+\log\mathfrak{c}(\Delta_{z}\rho_{2})+\log_{+}c\big).\end{split}$ (65)
 Combining (62)-(65), we deduce that
-$\begin{split}&H_{n}(X_{2}+Y_{2}|z_{1},z_{2})+H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2})|z_{1},z_{2})\geq n(2s+\epsilon_{1}^{\prime})+O(\log\mathfrak{K})\\ &\quad+O_{d,s}\big(-n(s_{1}+s_{2})-\log P(D_{n}(X_{1}-X_{2})=z_{1})-\log P(D_{n}(Y_{1}-Y_{2})=z_{2})\big).\end{split}$ (59)
-For the case $|z_{1}|<\zeta_{1}$ or $|z_{1}|<\zeta_{2}$, we use the trivial bound that
+$\begin{split}&H_{n}(X_{2}+Y_{2}|z_{1},z_{2})+H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2})|z_{1},z_{2})\geq n(2s+\epsilon_{1}^{\prime})\\ &\qquad+O(\log\mathfrak{K})+O_{d,s}\big(-2ns-\log P_{E}(D_{n}(X_{1}-X_{2})=z_{1})-\log P_{E}(D_{n}(Y_{1}-Y_{2})=z_{2})\big).\end{split}$ (66)
+For the case $|z_{1}|<\zeta_{1}$ or $|z_{2}|<\zeta_{2}$, we use the trivial bound that
 $H_{n}(X_{2}+Y_{2}|z_{1},z_{2})+H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2})|z_{1},z_{2})$
-$\geq H_{n}(X_{2}+Y_{2}|z_{1},z_{2})\geq\frac{1}{2}\big(H_{n}(X_{2}|z_{1},z_{2})+H_{n}(Y_{2}|z_{1},z_{2})+O(1)\big)$
-$\geq ns+O_{d,s}\big(\log_{+}C_{1}+\log_{+}C_{2}-n(s_{1}+s_{2})$
-$\qquad\qquad\qquad-\log P(D_{n}(X_{1}-X_{2})=z_{1})-\log P(D_{n}(Y_{1}-Y_{2})=z_{2})\big)+O(1),$ (60)
-which reslults from conditional independence, (9) and Lemma 2.32.
+$\geq H_{n}(X_{2}+Y_{2}|z_{1},z_{2})\geq\frac{H_{n}(X_{2}|z_{1},z_{2})+H_{n}(Y_{2}|z_{1},z_{2})}{2}+O(1)$
+$\geq ns+O\big(\log_{+}C_{1}+\log_{+}C_{2}-2ns$
+$\qquad\qquad-\log P_{E}(D_{n}(X_{1}-X_{2})=z_{1})-\log P_{E}(D_{n}(Y_{1}-Y_{2})=z_{2})\big)+O(1),$ (67)
+which results from conditional independence, (10), (63) and Lemma 2.32.
 We now average over all possible values of the discretized differences.
-Multiplying (59) and (60) by $P(D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2})$ and summing over all $(z_{1},z_{2})$, we obtain weighted sums of conditional entropies.
-The left-hand side gives
-$\sum\limits_{z_{1},z_{2}}P(D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2})H_{n}(X_{2}+Y_{2}|z_{1},z_{2})=H_{n}(X_{2}+Y_{2}|X_{1}-X_{2},Y_{1}-Y_{2})$
+Multiplying (66) and (67) by $P_{E}(D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2})$ and summing over all $(z_{1},z_{2})$, we obtain weighted sums of conditional entropies.
+The first term on the left-hand side gives
+$\sum\limits_{z_{1},z_{2}}P_{E}(D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2})H_{n}(X_{2}+Y_{2}|z_{1},z_{2})=H_{n}(X_{2}+Y_{2}|X_{1}-X_{2},Y_{1}-Y_{2})$
 by the definition of conditional entropy. For the negative logarithm terms, we use the fact that
-$-\sum\limits_{z_{1},z_{2}}P(D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2})\log P(D_{n}(X_{1}-X_{2})=z_{1})$
-$\qquad=-\sum\limits_{z_{1}}P(D_{n}(X_{1}-X_{2})=z_{1})\log P(D_{n}(X_{1}-X_{2})=z_{1})=H_{n}(X_{1}-X_{2}).$
+$-\sum\limits_{z_{1},z_{2}}P_{E}(D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2})\log P_{E}(D_{n}(X_{1}-X_{2})=z_{1})$
+$\qquad=-\sum\limits_{z_{1}}P_{E}(D_{n}(X_{1}-X_{2})=z_{1})\log P_{E}(D_{n}(X_{1}-X_{2})=z_{1})=H_{n}(X_{1}-X_{2}).$
 And other terms can be evaluated similarly. It leads to
 $H_{n}(X_{2}+Y_{2}|X_{1}-X_{2},Y_{1}-Y_{2})+H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2})|X_{1}-X_{2},Y_{1}-Y_{2})$
-$\qquad\geq n(2s+\epsilon_{1}^{\prime})P_{E}(|D_{n}(X_{1}-X_{2})|\geq\zeta_{1},|D_{n}(Y_{1}-Y_{2})|\geq\zeta_{2})$
-$\qquad\qquad+ns(P_{E}(|D_{n}(X_{1}-X_{2})|<\zeta_{1}\text{ or }|D_{n}(Y_{1}-Y_{2})|<\zeta_{2}))$
-$\qquad\qquad+O_{d,s}(-n(s_{1}+s_{2})+H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2}))+O(\log K).$ (61)
-Recall that $\zeta_{i}\geq 2^{-ns_{i}-4}$ $(i=1,2)$, which gives $\eta_{i}+2^{-n}\leq 20\eta_{i}$.
+$\qquad\geq n(2s+\epsilon_{1}^{\prime})\cdot P_{E}(|D_{n}(X_{1}-X_{2})|\geq\zeta_{1},|D_{n}(Y_{1}-Y_{2})|\geq\zeta_{2})$
+$\qquad\qquad+ns\cdot P_{E}(|D_{n}(X_{1}-X_{2})|<\zeta_{1}\text{ or }|D_{n}(Y_{1}-Y_{2})|<\zeta_{2})$
+$\qquad\qquad+O_{d,s}\big(-2ns+H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})\big)+O(\log\mathfrak{K}).$ (68)
+Recall that $\zeta_{i}\geq 2^{-n-4}$ $(i=1,2)$, which gives $\zeta_{i}+2^{-n}\leq 20\zeta_{i}$.
 Also note that $X_{1}-X_{2}$ and $Y_{1}-Y_{2}$ are $(s_{1},O(C_{1}))$-Frostman and $(s_{2},O(C_{2}))$-Frostman, respectively.
 We deduce that
-$P_{E}(|D_{n}(X_{1}-X_{2})|<\zeta_{1})\leq C_{1}(20\zeta_{1})^{s_{1}}\leq\epsilon_{1}^{\prime}/16,\quad P_{E}(|D_{n}(Y_{1}-Y_{2})|<\zeta_{2})\leq C_{2}(20\zeta_{2})^{s_{2}}\leq\epsilon_{1}^{\prime}/16,$
-and
-$H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})\geq n(s_{1}+s_{2})-\log_{+}C_{1}-\log_{+}C_{2}+O(1).$ (62)
-So $P_{E}(|D_{n}(X_{1}-X_{2})|\geq\zeta_{1},|D_{n}(Y_{1}-Y_{2})|\geq\zeta_{2})\geq 1-\epsilon_{1}^{\prime}/8$. Now (61) and (62) leads to
+$P_{E}(|D_{n}(X_{1}-X_{2})|<\zeta_{1})\leq C_{1}(20\zeta_{1})^{s_{1}}\ll\epsilon_{1}^{\prime}/\mathfrak{r},\quad P_{E}(|D_{n}(Y_{1}-Y_{2})|<\zeta_{2})\leq C_{2}(20\zeta_{2})^{s_{2}}\ll\epsilon_{1}^{\prime}/\mathfrak{r}.$
+So $P_{E}(|D_{n}(X_{1}-X_{2})|\geq\zeta_{1},|D_{n}(Y_{1}-Y_{2})|\geq\zeta_{2})=1+O(\epsilon_{1}^{\prime}/\mathfrak{r})$. Now (58) and (68) leads to
 $H_{n}(X_{2}+Y_{2}|X_{1}-X_{2},Y_{1}-Y_{2})+H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2})|X_{1}-X_{2},Y_{1}-Y_{2})$
-$\qquad\geq n(2s+\epsilon_{1}^{\prime})(1-\epsilon_{1}^{\prime}/8)+O(\log\mathfrak{K})\geq n(2s+3\epsilon_{1}^{\prime}/4)+O(\log\mathfrak{K}).$ (63)
-Finally, putting (50), (51), (4.2), (53) and (63) together, we arrive at
-$H_{n}(X+Y)+7H_{n}(\rho_{1}(X)+\rho_{2}(Y))\geq n(8s+\epsilon_{1}^{\prime}/2)+O(\log\mathfrak{K}).$
-By independence, we also have
-$H_{n}(X+Y)\geq\frac{H_{n}(X)+H_{n}(Y)}{2}+O(1)\geq ns-\log C_{1}C_{2}+O(1).$
-It follows that
-$7H_{n}(X+Y)+7H_{n}(\rho_{1}(X)+\rho_{2}(Y))\geq n(14s+\epsilon_{1}^{\prime}/2)+O(\log\mathfrak{K}).$
-The induction is completed by taking $\epsilon_{1}^{\prime\prime}=\epsilon_{1}^{\prime}/14$, i.e., $\epsilon_{1}^{\prime\prime}(d,s)=\epsilon_{1}^{\prime\prime}(d-1,s)/14$. ∎
+$\qquad\geq n(2s+\epsilon_{1}^{\prime})+O_{d,s}(n\epsilon_{1}^{\prime}/\mathfrak{r}+\Psi)+O(\log\mathfrak{K}).$ (69)
+Finally, putting (60) and (69) together, we arrive at
+$2ns+O(n\epsilon_{1}^{\prime}/\mathfrak{r}+\Psi+\log\mathfrak{K})\geq n(2s+\epsilon_{1}^{\prime})+O_{d,s}(n\epsilon_{1}^{\prime}/\mathfrak{r}+\Psi)+O(\log\mathfrak{K}).$
+That is to say, there is some constant $\mathfrak{C}=\mathfrak{C}(d,s)>0$ such that
+$\mathfrak{C}\Psi\geq n\epsilon_{1}^{\prime}+O_{d,s}(n\epsilon_{1}^{\prime}/\mathfrak{r})+O(\log\mathfrak{K}).$
+Now we take $\mathfrak{r}=\mathfrak{r}(d,s)\geq 1$ so that the second term on the right-hand side of above formula satisfies $O_{d,s}(n\epsilon_{1}^{\prime}/\mathfrak{r})\leq n\epsilon_{1}^{\prime}/2$.
+Then the induction is completed by taking $\epsilon_{1}^{\prime\prime}=\epsilon_{1}^{\prime}/2\mathfrak{C}$, i.e., $\epsilon_{1}^{\prime\prime}(d,s)=\epsilon_{1}^{\prime\prime}(d-1,s)/2\mathfrak{C}(d,s)$.
+∎
 4.3 Diagonal polynomials with dependent random variables
@@
 We now extend to the dependent case, proving Theorem 1.2 in full generality.
-The key tool is Theorem 3.1, which turns dependent random variables to a family of conditionally independent random variables at the cost of logarithmic error terms.
+The key tool is Theorem 3.1, which turns dependent random variables to a family of conditionally independent random variables with controllable error terms.
 Theorem 1.2’.
@@
 Let $0<s<1$, $C,c>0$.
-There is a positive constant $\epsilon_{2}=\epsilon_{2}(d,s)$ such that the following statement holds: For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
-$\begin{split}&\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\rho_{1}(X)+\rho_{2}(Y))\right\}\geq n(s+\epsilon_{2})+O(\log\mathfrak{K})\end{split}$
+There is a constant $\epsilon_{2}=\epsilon_{2}(d,s)>0$ such that the following statement holds: For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
+$\begin{split}&\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}\big(\rho_{1}(X)+\rho_{2}(Y)\big)\right\}\geq n(s+\epsilon_{2})+O(\log\mathfrak{K})\end{split}$
 for all $n\in\mathbb{N}$, where
@@
 Proof.
-To quantify the entropy gain beyond the trivial $ns$, we write the entropy increment as
-$\Psi:=\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\rho_{1}(X)+\rho_{2}(Y))\right\}-ns.$
+Similar as previous, we write
+$\Psi:=\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}\big(\rho_{1}(X)+\rho_{2}(Y)\big)\right\}-ns.$
 Our goal is to show $\Psi\geq n\epsilon_{2}+O(\log\mathfrak{K})$ for some $\epsilon_{2}=\epsilon_{2}(d,s)>0$. In particular,
@@
 Without loss of generality, we assume that $0<\epsilon_{1}^{\prime\prime}<1$.
-We use a cutoff parameter $\mathfrak{r}\geq 1$ as previous.
+Let $\mathfrak{r}=\mathfrak{r}(d,s)\geq 1$ be a parameter to be determined later.
 For $i=1,2$, take $\eta_{i}=\min\Big\{\big(\frac{\epsilon_{1}^{\prime\prime}}{4\mathfrak{r}dC_{0}}\big)^{1/s_{i}},\,\frac{1}{10}\Big\}$, which satisfies $dC_{0}\eta_{i}^{s_{i}}\leq\epsilon_{1}^{\prime\prime}/4\mathfrak{r}\leq 1/4$ and $\log(1/\eta_{i})=O(\log\mathfrak{K})$.
 Let $E$ be the event that $X\in\mathcal{R}(\rho_{1}^{\prime};\eta_{1})$, $Y\in\mathcal{R}(\rho_{2}^{\prime};\eta_{2})$.
-Since $(X,Y)$ is marginally $(s_{1},C_{0};s_{2},C_{0})$-Frostman, we deduce by Lemma 4.4 that
+Since $(X,Y)$ is marginally $(s_{1},C_{0};s_{2},C_{0})$-Frostman, we deduce by Lemma 4.7(i) that
 $P(E)$ $\geq 1-P(X\notin\mathcal{R}(\rho_{1}^{\prime};\eta_{1}))-P(Y\notin\mathcal{R}(\rho_{2}^{\prime};\eta_{2}))$
@@
 Let $(X_{0},Y_{0})$ be the random variable such that $(X_{0},Y_{0})\equiv(X,Y|E)$ (see Lemma A.2).
-By this construction, the functions $\rho_{1}$ and $\rho_{2}$ are weakly $(\kappa,d-1)$-bi-Lipschitz on $range_{+}(X_{0})$ and $range_{+}(Y_{0})$, respectively, with $\log\kappa=O(\log\mathfrak{K})$.
-ALso, the random variable $(X_{0},Y_{0})$ is both $(s_{1}+s_{2},O(C))$-Frostman, which gives
+By Lemma 4.7(v), the functions $\rho_{1}$ and $\rho_{2}$ are weakly $(\kappa,d)$-bi-Lipschitz on $range_{+}(X_{0})$ and $range_{+}(Y_{0})$, respectively, with $\log\kappa=O(\log\mathfrak{K})$.
+Also, the random variable $(X_{0},Y_{0})$ is $(s_{1}+s_{2},O(C))$-Frostman, which gives
 $H_{n}(X_{0})+H_{n}(Y_{0})\geq H_{n}(X_{0},Y_{0})\geq n(s_{1}+s_{2})-\log C+O(1)\geq 2ns+O(\log\mathfrak{K}).$ (72)
@@
 $\begin{split}&H_{n}(X_{0})+H_{n}(Y_{0})-H_{n}(X_{0},Y_{0})\leq\log K\\ &H_{n}(X_{0}+Y_{0})-\frac{H_{n}(X_{0})+H_{n}(Y_{0})}{2}\leq\log K,\\ &H_{n}(\rho_{1}(X_{0})+\rho_{2}(Y_{0}))-\frac{H_{n}(X_{0})+H_{n}(Y_{0})}{2}\leq\log K,\end{split}$
-where $\log K=O(\Psi)+O(n\epsilon^{\prime\prime}/\mathfrak{r})+O(\log\mathfrak{K})$.
-We now apply Theorem 3.1 to $(X_{0},Y_{0})$, and obtain a quadruple $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$ with certain properties (see (i)-(iv) in Theorem 3.1).
+where $\log K=O(\Psi+n\epsilon_{1}^{\prime\prime}/\mathfrak{r}+\log\mathfrak{K})$.
+We now apply Theorem 3.1 to $(X_{0},Y_{0})$, and obtain a quadruple $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$ such that (i) $(X^{\prime},Y^{\prime})\equiv(X_{0},Y_{0})$; (ii) $X^{\prime}$ and $X^{\prime\prime}$ are conditionally i.i.d.
+given $D_{n}(Y^{\prime})$; (iii) $Y^{\prime}$ and $Y^{\prime\prime}$ are conditionally i.i.d.
+given $D_{n}(X^{\prime})$; (iv) $(X^{\prime\prime},Y^{\prime})$ and $Y^{\prime\prime}$ are conditionally independent given $D_{n}(X^{\prime})$.
 In particular, the random variables $X^{\prime\prime}$ and $Y^{\prime\prime}$ are conditionally independent given $D_{n}(X^{\prime},Y^{\prime})$.
@@
 where the implied constants do not depend on $(x,y)$. We present the proof for $X^{\prime\prime}$, as the one for $Y^{\prime\prime}$ can be done similarly.
-For any ball $B(z,r)$, using conditions (i)–(iv) in Theorem 3.1, we have
+For any ball $B(z,r)$, using conditions (i)–(iv) on $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$, we have
 $P_{E}(X^{\prime\prime}\in B(z,r)|D_{n}(X^{\prime},Y^{\prime})=(x,y))=P_{E}(X^{\prime\prime}\in B(z,r)|D_{n}(Y^{\prime})=y)$
@@
 In what follows, $H_{n}(~\cdot~|x,y)$ is an abbreviation of $H(D_{n}(\cdot)~|D_{n}(X^{\prime},Y^{\prime})=(x,y)).$ By Lemma A.2, there exist random variables, depending on $(x,y)$, that have the same distributions with $(X^{\prime\prime},Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y))$, respectively.
-Similar as previously, we can apply Lemma 4.8 with these random variables, with $C_{1}=O\big(C\cdot 2^{-ns_{2}}/P_{E}(D_{n}(Y^{\prime})=y)\big)$ and $C_{2}=O\big(C\cdot 2^{-ns_{1}}/P_{E}(D_{n}(X^{\prime})=x)\big)$, to deduce that
+Similar as previously, we can apply Lemma 4.8 with these random variables with $C_{1}=O\big(C\cdot 2^{-ns_{2}}/P_{E}(D_{n}(Y^{\prime})=y)\big)$ and $C_{2}=O\big(C\cdot 2^{-ns_{1}}/P_{E}(D_{n}(X^{\prime})=x)\big)$, together with the fact $s_{1}+s_{2}\geq 2s$, to deduce that
 $H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y)+H_{n}(\rho_{1}(X^{\prime\prime})+\rho_{2}(Y^{\prime\prime})|x,y)\geq n(2s+\epsilon_{1}^{\prime\prime})+O_{d}(\log_{+}c)+O_{\phi,s}(1)$
-$\qquad+O_{d,s}\big(\log_{+}C-n(s_{1}+s_{2})-\log P_{E}(D_{n}(X^{\prime})=x)-\log P_{E}(D_{n}(Y^{\prime})=y)\big).$
+$\qquad+O_{d,s}\big(\log_{+}C-2ns-\log P_{E}(D_{n}(X^{\prime})=x)-\log P_{E}(D_{n}(Y^{\prime})=y)\big).$
 Now we sum all of these expressions respective to each $(x,y)$ with the weight $P_{E}(D_{n}(X^{\prime},Y^{\prime})=(x,y))$, which yields
-$\begin{split}&H_{n}(X^{\prime\prime}+Y^{\prime\prime}|X^{\prime},Y^{\prime})+H_{n}(\rho_{1}(X^{\prime\prime})+\rho_{2}(Y^{\prime\prime 2})|X^{\prime},Y^{\prime})\geq n(2s+\epsilon_{1}^{\prime\prime})\\ &\qquad\qquad\qquad+O_{d,s}(-n(s_{1}+s_{2})+H_{n}(X^{\prime})+H_{n}(Y^{\prime}))+O(\log\mathfrak{K}).\end{split}$ (71)
-Substituting from (66), (68),(69), and further combining with $(X^{\prime},Y^{\prime})\equiv(X_{0},Y_{0})$ we obtain that
-$H_{n}(X_{0})+H_{n}(Y_{0})+14\log K+O(\log\mathfrak{K})\geq n(2s+\epsilon_{1}^{\prime\prime})$ (72)
-Recalling (67) and the notation $\log K=O(\Psi)+O(n\epsilon^{\prime\prime}/\mathfrak{r})+O(\log\mathfrak{K})$, the left-hand side of (72) is
-$\leq 2ns+O(\Psi)+O(n\epsilon^{\prime\prime}/\mathfrak{r})+O(\log\mathfrak{K}).$ (73)
-Combining (72) and (73), there is some absolute constant $\tilde{C}>0$ such that
-$\tilde{C}\Psi\geq n\epsilon_{1}^{\prime\prime}+O(\log\mathfrak{K})+O(n\epsilon^{\prime\prime}/\mathfrak{r}).$
-We now take $\mathfrak{r}\geq 1$ to be some absolute constant such that the last term on the right-hand side of above formula satisfies $O(n\epsilon_{1}^{\prime\prime}/\mathfrak{r})\leq n\epsilon_{1}^{\prime\prime}/2$, Therefore,
+$\begin{split}&H_{n}(X^{\prime\prime}+Y^{\prime\prime}|X^{\prime},Y^{\prime})+H_{n}(\rho_{1}(X^{\prime\prime})+\rho_{2}(Y^{\prime\prime})|X^{\prime},Y^{\prime})\geq n(2s+\epsilon_{1}^{\prime\prime})\\ &\qquad\qquad\qquad+O_{d,s}(-2ns+H_{n}(X^{\prime})+H_{n}(Y^{\prime}))+O(\log\mathfrak{K}).\end{split}$ (77)
+Substituting from (74), (75), and further combining with $(X^{\prime},Y^{\prime})\equiv(X_{0},Y_{0})$, the left-hand side of (77) is bounded by
+$H_{n}(X_{0})+H_{n}(Y_{0})+14\log K+O(\log\mathfrak{K}).$
+Recalling (73) and the notation $\log K=O(\Psi+n\epsilon_{1}^{\prime\prime}/\mathfrak{r}+\log\mathfrak{K})$, we arrive at
+$2ns+O(\Psi+n\epsilon_{1}^{\prime\prime}/\mathfrak{r}+\log\mathfrak{K})\geq n(2s+\epsilon_{1}^{\prime\prime})+O_{d,s}(\Psi+n\epsilon_{1}^{\prime\prime}/\mathfrak{r})+O(\log\mathfrak{K}).$
+That is to say, there is some $\mathfrak{C}=\mathfrak{C}(d,s)>0$ such that
+$\mathfrak{C}\Psi\geq n\epsilon_{1}^{\prime\prime}+O_{d,s}(n\epsilon_{1}^{\prime\prime}/\mathfrak{r})+O(\log\mathfrak{K}).$
+We now take $\mathfrak{r}=\mathfrak{r}(d,s)\geq 1$ to be some constant such that the second term on the right-hand side of above formula satisfies $O_{d,s}(n\epsilon_{1}^{\prime\prime}/\mathfrak{r})\leq n\epsilon_{1}^{\prime\prime}/2$, Therefore,
 $\Psi\geq n\epsilon_{2}+O(\log\mathfrak{K})$
-for $\epsilon_{2}=\epsilon_{1}^{\prime\prime}/2\tilde{C}$. The proof is completed. ∎
+for $\epsilon_{2}=\epsilon_{1}^{\prime\prime}/2\mathfrak{C}$. The proof is completed. ∎
 4.4 More general polynomials
-Now we prove Theorem 1.3 in the form of the following two theorems, which hold for more general forms.
+We now study more general polynomials. We first prove several theorems that are weaker than Theorem 1.3, but contain explicit error terms.
 Theorem 1.3’.
-Let $d\geq 2$ and $\rho_{1}$, $\rho_{2}$ be real polynomials of degree $d$.
+Suppose that $\phi(x,y)$ has the form (2).
 Let $0<s<1$ and $C_{1},C_{2},c>0$.
-There is a positive constant $\epsilon_{3}=\epsilon_{3}(d,s)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$.
+There is a constant $\epsilon_{3}>0$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$.
 Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
 Then
-$\begin{split}&\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\rho_{1}(X)+\rho_{2}(Y))\right\}\\ &\qquad\qquad\qquad\geq n(s+\epsilon_{3})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\rho_{1},\rho_{2},s}(1).\end{split}$
-for all $n\in\mathbb{N}$.
+$\begin{split}&\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\qquad\geq n(s+\epsilon_{3})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\phi,s}(1)\end{split}$
+for all $n\in\mathbb{N}$. Here $\epsilon_{3}=\epsilon_{3}(d,s)$ when $\phi(x,y)$ is diagonal, and $\epsilon_{3}=\epsilon_{3}(d,\Xi,s)$ when $\phi(x,y)$ is non-diagonal.
 Proof.
 Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, by Lemma 2.30 it is also jointly $(s_{1},s_{2},C_{1}C_{2})$-Frostman and marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman.
-Applying Theorem 1.2’, there is some $\epsilon_{2}=\epsilon_{2}(d,s)$ such that
-$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\rho_{1}(X)+\rho_{2}(Y))\right\}$
-$\qquad\qquad\geq n(s+\epsilon_{2})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\rho_{1},\rho_{2},s}(1)$
-for all $n\in\mathbb{N}$.
-Here the implied constant in the term $O(\log_{+}c)$ is independent of $s$, since the marginally Frostman constants do not involve $\max\{1,c\}$ as in (65).
-In view of $s_{1}+s_{2}\geq 2s$, Lemma 2.32(ii) shows that
-$H_{n}(X,Y)\geq\frac{H_{n}(X)+H_{n}(Y)}{2}+ns-\frac{1}{2}\log C_{1}C_{2}+O(1).$
-The conclusion then follows with $\epsilon_{3}=\epsilon_{2}/2$.
-Theorem 1.3”.
-Suppose that $\phi(x,y)$ has the form (1).
-Let $0<s<1$ and $C_{1},C_{2},c>0$.
-There is a positive constants $\epsilon_{4}=\epsilon_{4}(d,\Xi,s)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$.
-Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
-Then
-$\begin{split}&\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\qquad\geq n(s+\epsilon_{4})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\phi,s}(1).\end{split}$
-for all $n\in\mathbb{N}$.
-Proof.
-Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, Lemma 2.30 tells that it is both marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman and jointly $(s_{1},s_{2},C_{1}C_{2})$-Frostman.
-Then
-$H_{n}(X)+H_{n}(Y)\geq H_{n}(X,Y)\geq 2ns-\log C_{1}C_{2}+O(1).$ (74)
-In view of Lemma 2.32(ii), one has
-$\frac{H_{n}(X)+H_{n}(Y)}{2}\leq H_{n}(X,Y)-ns+\frac{1}{2}\log C_{1}C_{2}+O(1).$
+In view of $s_{1}+s_{2}\geq 2s$, applying Lemma 2.4(iv) and Lemma 2.32 yields
+$H_{n}(X)+H_{n}(Y)\geq H_{n}(X,Y)\geq 2ns-\log C_{1}C_{2}+O(1),$ (78)
+and
+$H_{n}(X,Y)\geq\frac{H_{n}(X)+H_{n}(Y)}{2}+ns-\frac{1}{2}\log C_{1}C_{2}+O(1).$ (79)
 It follows that
 $H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\leq H_{n}(X,Y)-2ns+\log C_{1}C_{2}+O(1).$ (80)
-Consider the random variable $(\xi_{11}X+\xi_{12}Y,\,\xi_{21}X+\xi_{22}Y)$.
+When $\phi(x,y)$ is diagonal, we can write $\phi(x,y)=\tilde{\rho}_{1}(x)+\tilde{\rho}_{2}(y)$ with $\tilde{\rho}_{1}$, $\tilde{\rho}_{2}$ being polynomials of degree $d$.
+Applying Theorem 1.2’ yields $\epsilon_{2}=\epsilon_{2}(d,s)>0$ such that
+$\begin{split}&\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\tilde{\rho_{1}}(X)+\tilde{\rho_{2}}(Y))\right\}\\ &\qquad\qquad\geq n(s+\epsilon_{2})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\phi,s}(1)\end{split}$ (81)
+for all $n\in\mathbb{N}$.
+Here the implied constant in the term $O_{d}(\log_{+}c)$ is independent of $s$, since the marginally Frostman constants do not involve $\max\{1,c\}$ as in (71).
+The conclusion then follows by combining (79) and (81), with $\epsilon_{3}(d,s)=\epsilon_{2}(d,s)/2$.
+When $\phi(x,y)$ is non-diagonal in the form (2), we consider the random variable $(\xi_{11}X+\xi_{12}Y,\,\xi_{21}X+\xi_{22}Y)$.
 Its range is contained in $c^{\prime\prime},c^{\prime\prime}]^{2}$, with $c^{\prime\prime}=O_{\phi}(c)$.
-And Lemma 2.31, we assume without loss of generality that it is jointly $(s_{1},s_{2},C^{\prime\prime})$-Frostman, with $\log C^{\prime\prime}=\log C_{1}C_{2}+O_{\phi}(1)$, since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman.
+By Lemma 2.31, we assume without loss of generality that it is jointly $(s_{1},s_{2},C^{\prime\prime})$-Frostman, with $\log C^{\prime\prime}=\log C_{1}C_{2}+O_{\phi}(1)$, since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman.
 By Lemma 1.2’, there is some $\epsilon_{2}=\epsilon_{2}(d,s)>0$ such that
@@
 for some $d_{1}=d_{1}(\Xi)>0$ and $d_{2}=d_{2}(\Xi)\geq 0$. By (78), (80) and (83), one gets
-$\begin{split}&d_{1}\cdot H_{n}(X+Y)+2d_{2}\cdot\frac{1}{2}H_{n}(X,Y)\geq H_{n}\big((\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big)\\ &\qquad\qquad\qquad+(d_{1}+2d_{2}-1)ns+O(\log_{+}C_{1}C_{2})+O_{\phi}(1).\end{split}$ (78)
+$\begin{split}&d_{1}\cdot H_{n}(X+Y)+2d_{2}\cdot\frac{1}{2}H_{n}(X,Y)\geq H_{n}\big((\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big)\\ &\qquad\qquad\qquad+(d_{1}+2d_{2}-1)\big(ns-\frac{1}{2}\log_{+}C_{1}C_{2}\big)+O_{\phi}(1).\end{split}$ (84)
 Similar arguments show that
-$\begin{split}&d_{1}^{\prime}\cdot H_{n}(X+Y)+2d_{2}^{\prime}\cdot\frac{1}{2}H_{n}(X,Y)\geq H_{n}(\xi_{11}X+\xi_{12}Y)\\ &\qquad\qquad\qquad+(d_{1}^{\prime}+2d_{2}^{\prime}-1)ns+O(\log_{+}C_{1}C_{2})+O_{\phi}(1).\end{split}$ (79)
+$\begin{split}&d_{1}^{\prime}\cdot H_{n}(X+Y)+2d_{2}^{\prime}\cdot\frac{1}{2}H_{n}(X,Y)\geq H_{n}(\xi_{11}X+\xi_{12}Y)\\ &\qquad\qquad\qquad+(d_{1}^{\prime}+2d_{2}^{\prime}-1)\big(ns-\frac{1}{2}\log_{+}C_{1}C_{2}\big)+O_{\phi}(1)\end{split}$ (85)
 for some $d_{1}^{\prime}=d_{1}^{\prime}(\Xi)>0$ and $d_{2}^{\prime}=d_{2}^{\prime}(\Xi)\geq 0$. And also
-$\begin{split}&d_{1}^{\prime\prime}\cdot H_{n}(X+Y)+2d_{2}^{\prime\prime}\cdot\frac{1}{2}H_{n}(X,Y)\geq H_{n}(\xi_{21}X+\xi_{22}Y)\\ &\qquad\qquad\qquad+(d_{1}^{\prime\prime}+2d_{2}^{\prime\prime}-1)ns+O(\log_{+}C_{1}C_{2})+O_{\phi}(1).\end{split}$ (80)
-for some $d_{1}^{\prime\prime}=d_{1}^{\prime\prime}(\Xi)>0$ and $d_{2}^{\prime\prime}=d_{2}^{\prime\prime}(\Xi)\geq 0$.
-Now the theorem follows by combining (76) and (78)-(80), with $\epsilon_{4}=\epsilon_{2}/\max\{d_{1}+2d_{2},d_{1}^{\prime}+2d_{2}^{\prime},d_{1}^{\prime\prime}+2d_{2}^{\prime\prime}\}$. ∎
-For Theorem 1.4, the situation that $\phi(x,y)$ is diagonal has been proved by Lemma 4.8. The general case will be proved in the following form.
-Theorem 1.4’ Suppose that $\phi(x,y)$ has the form (1).
+$\begin{split}&d_{1}^{\prime\prime}\cdot H_{n}(X+Y)+2d_{2}^{\prime\prime}\cdot\frac{1}{2}H_{n}(X,Y)\geq H_{n}(\xi_{21}X+\xi_{22}Y)\\ &\qquad\qquad\qquad+(d_{1}^{\prime\prime}+2d_{2}^{\prime\prime}-1)\big(ns-\frac{1}{2}\log_{+}C_{1}C_{2}\big)+O_{\phi}(1)\end{split}$ (86)
+for some $d_{1}^{\prime\prime}=d_{1}^{\prime\prime}(\Xi)>0$ and $d_{2}^{\prime\prime}=d_{2}^{\prime\prime}(\Xi)\geq 0$. Now the theorem follows by combining (82) and (84)-(86), with
+$\epsilon_{3}(d,\Xi,s)=\frac{\epsilon_{2}(d,s)}{\max\{d_{1}(\Xi)+2d_{2}(\Xi),\,d_{1}^{\prime}(\Xi)+2d_{2}^{\prime}(\Xi),\,d_{1}^{\prime\prime}(\Xi)+2d_{2}^{\prime\prime}(\Xi)\}}.$
+Theorem 1.3”.
+Suppose that $\phi(x,y)$ has the form (2).
 Let $0<s<1$, $C_{1},C_{2},c>0$.
-There is a positive constant $\epsilon_{4}=\epsilon_{4}(d,\Xi,s)$ such that the following statement holds: Let $X$ and $Y$ be two independent real random variables with ranges contained in [c,c][-c,c].
+There is a positive constant $\epsilon_{3}>0$ such that the following statement holds: Let $X$ and $Y$ be independent real random variables with ranges both contained in [c,c][-c,c].
 Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, where $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$.
 Then
-$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{4})+O_{d,s}(\log_{+}C_{1}C_{2})+O_{d}(\log_{+}c)+O_{\phi,s}(1)$
-for all $n\in\mathbb{N}$.
+$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{4})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\phi,s}(1)$
+for all $n\in\mathbb{N}$. Here $\epsilon_{3}=\epsilon_{3}(d,s)$ when $\phi(x,y)$ is diagonal, and $\epsilon_{3}=\epsilon_{3}(d,\Xi,s)$ when $\phi(x,y)$ is non-diagonal.
 Proof.
-By Lemma 2.30, the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman. Theorem 1.3” provide us with some positive constant $\epsilon_{4}=\epsilon_{4}(d,\Xi,s)$ such that
-$\begin{split}&\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\qquad\geq n(s+\epsilon_{4})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\phi,s}(1).\end{split}$
+By Lemma 2.30, the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman. So the conclusion of Theorem 1.3’ holds. , i.e.,
+$\begin{split}&\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\qquad\geq n(s+\epsilon_{3})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\phi,s}(1).\end{split}$
 Combining (9) and (10), one has
@@
 Remark 4.9.
-In the conclusions of Lemma 4.8 and Theorems 1.2’-1.4’, the implied constants in the terms $O_{d,s}(\log_{+}C)$, $O_{d,s}(\log_{+}C_{1}C_{2})$ or $O_{d,s}(\log_{+}c)$ can be independent of $s$ if we are only considering the case $s\gg 1$.
+In the conclusions of Lemma 4.8 and Theorems 1.2’, 1.3’, 1.3”, the implied constants in the terms $O_{d,s}(\log_{+}C)$, $O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})$ or $O_{d,s}(\log_{+}c)$ can be independent of $s$ if we are only considering the case $s\gg 1$.
 We preserve the dependence on $s$ in these terms for the uniformity with the case $0<s\leq 1/2$.
-5 Proof of Theorem 1.4 with $s\leq 1/2$
-The situation is more subtle for this case.
-It is hard to expect Theorem 2.38 holds true when the dimension gets smaller.
-We will use the upcoming theorem as a substitution.
-The core of the proof will follow the arguments from [24] by Shmerkin and Wang.
-However, there are enough differences that we cannot merely apply a single theorem of theirs as a black box, so to address this case, we will introduce various notions and results from their paper, and explicitly mention where our arguments diverge.
+To prove Theorem 1.3, we need a further property regarding the discretized joint entropy $H_{n}(X,Y)$.
+Proposition 4.10.
+Let $\phi(x,y)$ be given as in (2). Let $(X,Y)$ be a conditionally $(s,C;s,C)$-Frostman bivariate real random variable with range contained in $c,c]^{2}$, where $s,C,c>0$. Then
+$H_{n}(X,Y)=H_{n}\big(X+Y,\,\phi(X,Y)\big)+O_{\phi,s,C,c}(1)$
+for any $n\in\mathbb{N}$.
+Proposition 4.10 will be proved within several lemmas, in which same notations and conditions are used.
+Let $F:\,{\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$ be defined by $F(x,y)=(x+y,\phi(x,y))$.
+Write $\mathfrak{J}(x,y)$ for its Jacobian matrix, and $J(x,y)$ for its Jacobian determinant.
+A direct computation yields
+$J(x,y)$ $=\frac{\partial\phi}{\partial y}(x,y)-\frac{\partial\phi}{\partial x}(x,y)$
+$=(\xi_{12}-\xi_{11})\rho_{1}^{\prime}(\xi_{11}x+\xi_{12}y)+(\xi_{22}-\xi_{21})\rho_{2}^{\prime}(\xi_{21}x+\xi_{22}y),$
+which is a polynomial of degree at most d1d-1.
+Lemma 4.11.
+The Jacobian determinant $J(x,y)$ is not the zero polynomial.
+Proof.
+Suppose on the contrary that $J\equiv 0$. Consider the linear change of variables
+$(u,v):=(\xi_{11}x+\xi_{12}y,\ \xi_{21}x+\xi_{22}y).$
+Since $\Xi$ is invertible, $(x,y)\mapsto(u,v)$ is an invertible linear map, hence polynomial identities in $(x,y)$ are equivalent to polynomial identities in $(u,v)$.
+In $(u,v)$-coordinates, the equality $J\equiv 0$ appears as
+$(\xi_{12}-\xi_{11})\,\rho_{1}^{\prime}(u)+(\xi_{22}-\xi_{21})\,\rho_{2}^{\prime}(v)\equiv 0.$
+Such a sum can vanish identically only if both summands are constant and opposite.
+In view of $d\geq 2$, one sees that $\rho_{1}^{\prime}$ and $\rho_{2}^{\prime}$ are non-constant.
+It follows that $\xi_{12}=\xi_{11}$ and $\xi_{22}=\xi_{21}$.
+This contradicts the fact that $\xi_{11}\xi_{22}-\xi_{12}\xi_{21}\neq 0$.
+∎
+By Lemma 4.11, the critical set
+$\mathcal{Z}:=\{(x,y)\in{\mathbb{R}}^{2}:\ J(x,y)=0\}.$
+is a (possibly reducible) algebraic curve. For any $\delta>0$, we denote by $\mathcal{N}_{\delta}(\mathcal{Z})$ the $\delta$-neighbourhood of the set $\mathcal{Z}$ in ${\mathbb{R}}^{2}$.
+Lemma 4.12.
+It satisfies that
+$P\big((X,Y)\in\mathcal{N}_{\delta}(\mathcal{Z})\big)\ll_{\phi,c}C\delta^{s}.$
+for all $\delta\in(0,1]$.
+Proof.
+Write $\mathcal{R}:=[-c,c]^{2}$ and $\Gamma:=\mathcal{Z}\cap\mathcal{R}$.
+Since $\Gamma$ is a compact semialgebraic set of dimension at most $1$, it admits a finite $C^{1}$-stratification.
+In particular, there exist integers $l=l(\phi),l^{\prime}=l^{\prime}(\phi)\geq 0$, $C^{1}$ arcs $\gamma_{1},\dots,\gamma_{l}\subseteq\mathcal{R}$, and points $p_{1},\dots,p_{l^{\prime}}\in\mathcal{R}$ such that
+$\Gamma=\Big(\bigcup_{i=1}^{l}\gamma_{i}\Big)\cup\{p_{1},\dots,p_{l^{\prime}}\},$
+and for each $i$ the arc $\gamma_{i}$ is the graph of a $C^{1}$ function either of the form $\{(x,f_{i}(x)):x\in I_{i}\}$ or $\{(g_{i}(y),y):y\in J_{i}\}$ on some compact intervals $I_{i},J_{i}$.
+Let
+$L:=1+\max_{1\leq i\leq l}\{\|f_{i}^{\prime}\|_{\infty}+\|g_{i}^{\prime}\|_{\infty}\}\ll_{\phi,c}1.$
+Then for each graph arc we have the inclusions
+$\mathcal{N}_{\delta}\big(\{(x,f_{i}(x))\}\big)\cap\mathcal{R}\subseteq\{(x,y)\in\mathcal{R}:\ |y-f_{i}(x)|\leq L\delta\},$
+$\mathcal{N}_{\delta}\big(\{(g_{i}(y),y)\}\big)\cap\mathcal{R}\subseteq\{(x,y)\in\mathcal{R}:\ |x-g_{i}(y)|\leq L\delta\}.$
+For an arc of the form $\{(x,f_{i}(x))\}$, it follows by the law of total expectation and the conditionally Frostman condition that
+$P\big((X,Y)\in\mathcal{N}_{\delta}(\gamma_{i})\big)\leq P\big(|Y-f_{i}(X)|\leq L\delta\big)=\mathbb{E}\Big(P\big(|Y-f_{i}(X)|\leq L\delta\mid X\big)\Big)\leq C(L\delta)^{s}.$
+Similarly, for an arc of the form $\{(g_{i}(y),y)\}$ the same bound holds. Finally, for each point $p_{j}=(p_{j,1},p_{j,2})$, we have
+$P\big((X,Y)\in B(p_{j},\delta)\big)\leq P\big(|Y-p_{j,2}|\leq\delta\big)\leq C\delta^{s},$
+since $(X,Y)$ is also marginally $(s,C;s,C)$-Frostman by Lemma 2.30.
+Summing over the finitely many arcs and points yields
+$P\big((X,Y)\in\mathcal{N}_{\delta}(\mathcal{Z})\big)\leq l\cdot C(L\delta)^{s}+l^{\prime}\cdot C\delta^{s}\ll_{\phi,c}C\delta^{s}.$
+When $\mathcal{Z}$ is empty, it is not hard to prove Proposition 4.10.
+In the following, we assume $\mathcal{Z}$ to be non-empty, and apply the Łojasiewicz inequality (see [21, Theorem 2.12] for example) on our Jacobian determinant $J(x,y)$ and the distance function to $\mathcal{Z}$.
+Lemma 4.13 (Łojasiewicz).
+There exist constants $\gamma=\gamma(\phi,c)>0$ and $L=L(\phi,c)>0$ such that
+$\text{dist}\big((x,y),\mathcal{Z}\big)\leq L|J(x,y)|^{1/\gamma}$
+for all $(x,y)\in[-c,c]^{2}$.
+Corollary 4.14.
+It satisfies that
+$P\big(|J(X,Y)|\leq 2^{-k}\big)\ll_{\phi,c}C2^{-ks/\gamma}$
+for all $k\in\mathbb{N}$. In particular, if we define the integer-valued random variable $K:=\max\Big\{0,\ \big\lfloor-\log|J(X,Y)|\big\rfloor\Big\}$, then
+$P(K\geq k)\ll_{\phi,c}C2^{-ks/\gamma}$
+for all $k\geq 0$.
+Proof.
+By Lemma 4.13 and then Lemma 4.12, we have
+$P(|J(X,Y)|\leq 2^{-k})\leq P\big((X,Y)\in\mathcal{N}_{L2^{-k/\gamma}}(\mathcal{Z})\big)\ll_{\phi,c}C(L2^{-k/\gamma})^{s}.$
+The bound for $K$ follows since $\{K\geq k\}\subseteq\{|J(X,Y)|\leq 2^{-k}\}$. ∎
+Next, let us consider the Jacobian matrix $\mathfrak{J}(x,y)$.
+In the region $(x,y)\in[-c-1,c+1]$, the operator norm satisfies $\|\mathfrak{J}(x,y)\|\leq\mathfrak{L}$ for some constant $\mathfrak{L}=\mathfrak{L}(\phi,c)>0$.
+Similarly, the Jacobian determinant $J$ is also $\mathfrak{L}^{\prime}$-Lipschitz function on $c-1,c+1]^{2}$ for some constant $\mathfrak{L}^{\prime}=\mathfrak{L}^{\prime}(\phi,c)$, i.e.,
+$|J(x,y)-J(x^{\prime},y^{\prime})|\leq\mathfrak{L}^{\prime}\,|(x,y)-(x^{\prime},y^{\prime})|$ (87)
+for all $(x,y),(x^{\prime},y^{\prime})\in[-c-1,c+1]^{2}$.
+For fixed $(u,v)\in{\mathbb{R}}^{2}$, solving $F(x,y)=(u,v)$ means
+$x+y=u,\qquad\phi(x,u-x)=v.$
+The second equation is a univariate polynomial equation in $x$ of degree at most $d$.
+Hence for each fixed $u$ and each $v$, there are at most $d$ solutions for $x$ unless the polynomial in $x$ degenerates to an identically constant polynomial, which can occur only for finitely many $u$ and thus on a target set of Lebesgue measure 0.
+Therefore the fiber multiplicity function $N(u,v):=\#\{(x,y)\in[-c,c]^{2}:F(x,y)=(u,v)\}$ satisfies
+$N(u,v)\leq d\quad\text{for Lebesgue-a.e.\ }(u,v)\in{\mathbb{R}}^{2}.$ (88)
+Let $\mathcal{Q}_{n}$ denote the family of dyadic squares in ${\mathbb{R}}^{2}$ of side length $2^{-n}$.
+For $(u,v)\in{\mathbb{R}}^{2}$, let $Q_{n}(u,v)\in\mathcal{Q}_{n}$ be the unique dyadic square containing $(u,v)$.
+Lemma 4.15.
+For given $n\in\mathbb{N}$, let $B\in\mathcal{Q}_{n}$ be a dyadic square in the target $(u,v)$-plane. For $k\geq 0$, define
+$E_{n,k}(B):=\Big\{(x,y)\in[-c,c]^{2}:\ F(x,y)\in B\ \text{and}\ 2^{-k-1}<|J(x,y)|\leq 2^{-k}\Big\}.$
+Assume $k\leq n-\big\lceil\log\mathfrak{L}^{\prime}\big\rceil-5$. Then the number of dyadic input squares $Q\in\mathcal{Q}_{n}$ in the $(x,y)$-plane that intersect $E_{n,k}(B)$ has the upper bound
+$\{Q\in\mathcal{Q}_{n}:Q\cap E_{n,k}(B)\neq\emptyset\}\ll_{\phi,c}2^{k}.$
+Here the implied constant is independent of $n,k,B$.
+Proof.
+Let $\mathcal{S}$ be the set of dyadic input squares $Q\in\mathcal{Q}_{n}$ intersecting $E_{n,k}(B)$, and set $U:=\bigcup_{Q\in\mathcal{S}}Q$.
+Then the squares in $\mathcal{S}$ are disjoint and the area satisfies
+$\mathcal{L}^{2}(U)=2^{-2n}(\#\mathcal{S}).$ (89)
+Fix any $Q\in\mathcal{S}$.
+By definition, there exists $(x_{Q},y_{Q})\in Q\cap[-c,c]^{2}$ with $F(x_{Q},y_{Q})\in B$ and $|J(x_{Q},y_{Q})|>2^{-k-1}$.
+By (87) and the fact $k\leq n-\big\lceil\log\mathfrak{L}^{\prime}\big\rceil-5$, we have
+$|J(x,y)-J(x_{Q},y_{Q})|\leq\mathfrak{L}^{\prime}\,|(x,y)-(x_{Q},y_{Q})|\leq\mathfrak{L}^{\prime}\,\sqrt{2}\,2^{-n}\leq 2^{-k-3}$
+for all $(x,y)\in Q$, and then
+$|J(x,y)|\geq|J(x_{Q},y_{Q})|-2^{-k-3}\geq 2^{-k-1}-2^{-k-3}\geq 2^{-k-2}.$
+Hence $|J(x,y)|\geq 2^{-k-2}$ for all $(x,y)\in U$. Combining (89), one then deduce that
+$\iint_{U}|J(x,y)|\,dxdy\geq 2^{-k-2}\mathcal{L}^{2}(U)\gg 2^{-2n-k}(\#\mathcal{S}).$ (90)
+Next, since $\|\mathfrak{J}\|\leq\mathfrak{L}$ on $c-1,c+1]^{2}$, the image of each $Q$ has diameter at most $\mathfrak{L}\,\sqrt{2}2^{-n}$.
+Because $F(x_{Q},y_{Q})\in B$, it follows that $F(Q)$ is contained in a fixed enlargement $B^{+}$ of $B$ by an $O(\mathfrak{L}2^{-n})$ margin, hence $F(U)\subseteq B^{+}$, with the area satisfying
+$\mathcal{L}^{2}(B^{+})\ll_{\phi,c}2^{-2n}.$ (91)
+Now applying the area formula to the Lipschitz map $F$ as a change of variable on the measurable set $U$, combining (88) and (91), we get
+$\iint_{U}|J(x,y)|\,dxdy\leq\iint_{F(U)}N(u,v)\,dudv\leq d\,\mathcal{L}^{2}(F(U))\leq d\,\mathcal{L}^{2}(B^{+})\ll_{\phi,c}2^{-2n}.$ (92)
+Putting (90) and (92) together, one concludes $S\ll_{\phi,c}2^{k}$. The lemma then follows. ∎
+Proof of Proposition 4.10.
+For simplicity, we write
+$Z_{n}:=D_{n}(X,Y),\qquad W_{n}:=D_{n}\big(X+Y,\phi(X,Y)\big).$
+By mean value inequality, one can see that $W_{n}$ is an $\mathfrak{L}$-Lipschitz function of $Z_{n}$, where, as previous, $\mathfrak{L}=\mathfrak{L}(\phi,c)$ is an upper bound for the operator norm of the Jacobian matrix $\mathfrak{J}(x,y)$ in the region $c-1,c+1]^{2}$.
+Lemma 2.16 shows that
+$H_{n}(W_{n})\leq H_{n}(Z_{n})+O_{\phi,c}(1).$
+For the opposite direction, in view of (6), it is sufficient to prove that
+$H(Z_{n}|W_{n})=O_{\phi,s,C,c}(1).$ (93)
+Recall the random variable $K$ from Corollary 4.14, we will use the inequality
+$H(Z_{n}\mid W_{n})\leq H(K)+H(Z_{n}\mid W_{n},K)$ (94)
+First, by Corollary 4.14 one has
+$p_{k}:=P(K=k)\leq P(K\geq k)\leq M2^{-ks/\gamma},\quad(k\geq 0)$
+for some constant $M=M(\phi,C,c)\geq 1$. Here $\gamma$ is the constant in Lemma 4.13 that only depends on $\phi$ and $c$. Now calculation gives
+$\mathbb{E}(K)=\sum_{k\geq 1}P(K\geq k)\leq M\sum_{k\geq 1}2^{-ks/\gamma}\ll_{\phi,s,C,c}1.$ (95)
+Moreover, take $k_{1}=\lceil(\gamma/s)\log(8M)\rceil$. For any $k\geq k_{1}$, one has $p_{k}<1/e$, so by monotonicity one has
+$p_{k}\log p_{k}\leq M2^{-ks/\gamma}(-\log M+ks/\gamma)\leq(Mks/\gamma)\,2^{-ks/\gamma}.$
+It follows that
+$\sum_{k\geq k_{1}}p_{k}\log p_{k}\leq(Ms/\gamma)\sum\limits_{k\geq k_{1}}k2^{-ks/\gamma}\ll_{\phi,s,C,c}1.$
+Moreover,
+$\sum_{0\leq k<k_{1}}p_{k}\log p_{k}\leq k_{1}\ll_{\phi,s,C,c}1.$
+since $p_{k}\log p_{k}\leq 1$. It follows that
+$H(K)=-\sum\limits_{k\geq 0}p_{k}\log p_{k}\ll_{\phi,s,C,c}1.$ (96)
+Second, we fix $n$ and condition on the event $\{W_{n}=(u,v),\,K=k\}$. The value $W_{n}=(u,v)$ fixes the target dyadic square $B:=Q_{n}(u,v)$ intersecting the range of $(X+Y,\phi(X,Y))$.
+When $k\leq n-\big\lceil\log\mathfrak{L}^{\prime}\big\rceil-5$, Lemma 4.15 shows that the number of dyadic input squares $Q\in\mathcal{Q}_{n}$, which can intersect the range of $(X,Y)$ while still being consistent with $F(X,Y)\in B$ and the Jacobian bound, is at most $O_{\phi,c}(2^{k})$.
+Therefore,
+$H(Z_{n}\mid W_{n}=w,K=k)\leq\log\big(O_{\phi,c}(2^{k})\big)=k+O_{\phi,c}(1).$
+When $k>n-\big\lceil\log\mathfrak{L}^{\prime}\big\rceil-5$, we use the trivial bound for the discretized variables $Z_{n}$ in $c,c]^{2}$:
+$H(Z_{n}\mid W_{n}=w,K=k)\leq H(Z_{n})\leq 2n+O_{c}(1)\leq 2k+O_{\phi,c}(1),$
+where we have used Lemma 2.4(iv), (8) and the fact that $n<k+\big\lceil\log\mathfrak{L}^{\prime}\big\rceil+5$. Combining the two regimes, we see that uniformly for all $k\geq 0$,
+$H(Z_{n}\mid W_{n}=w,K=k)\leq 2k+O_{\phi,C,c}(1).$
+Taking the expectation over $K$, we obtain by (95) that
+$H(Z_{n}\mid W_{n},K)\leq 2\mathbb{E}(K)+O_{\phi,C,c}(1)=O_{\phi,s,C,c}(1).$ (97)
+Combing (94), (96) and (97), we obtain (93). The proposition then follows. ∎
+Finally, we complete the proof of Theorem 1.3.
+Proof of Theorem 1.3.
+Since $(X,Y)$ is conditionally $(s,C;s,C)$-Frostman, an application of Proposition 4.10 leads to
+$\max\{H_{n}(X+Y),\,\phi(X,Y)\}\geq\frac{1}{2}H(X,Y)+O_{\phi,s,C,c}(1).$
+Inserting the above inequality to the first-term in the maxima in the conclusion of Theorem 1.3’ with the conditionally $(s,C;s,C)$-Frostman random variable $(X,Y)$, we obtain
+$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{3})+O_{\phi,s,C,c}(1).$
+Theorem 1.3 then follows. ∎
+5 Proof of Theorems 1.2 and 1.3 with $s\leq 1/2$
+When $s\leq 1/2$, as noted in the introduction, the distance–entropy input used for $s>1/2$ cannot yield a gain beyond the Frostman exponent.
+In this range, we instead use a product-structure distance mechanism, following Shmerkin and Wang [35], and this is the reason for the independence hypothesis in Theorem 5.1.
+Define the increasing function
+$\varphi(u)=\frac{u}{2}+\frac{\sqrt{4+u^{2}}-2}{2},\quad(0<u\leq 1).$
 Theorem 5.1.
-Let $X$, $Y$ be independent random variables with ranges contained in [c,c][-c,c], with $c>0$.
-Suppose that $X$ and $Y$ are $(s,C_{1})$-Frostman and $(s,C_{2})$-Frostman, respectively, where $0<s\leq 1/2$ and $C_{1},C_{2}\geq 1$.
+Let $X$, $Y$ be independent random variables with ranges contained in [c,c][-c,c], where $c>0$.
+Suppose that $(X,Y)$ is marginally $(s,C;s,C)$-Frostman, where $0<s\leq 1/2$ and $C>0$.
 Let $(X^{\prime},Y^{\prime})$ be i.i.d.
@@
 Then for any $t$ with $s<t<\frac{s+\varphi(2s)}{2}$, we have
-$H_{n}\big(|(X,Y)-(X^{\prime},Y^{\prime})|\big)\geq nt+O_{s}(\log n)+O_{s,t}(\log C_{1}C_{2})+O_{s,c}(1)$
+$H_{n}\big(|(X,Y)-(X^{\prime},Y^{\prime})|\big)\geq nt+O_{s}(\log n+\log_{+}c)+O_{s,t}(\log_{+}C+1)$
 for all $n\in\mathbb{N}$.
-Here
-$\varphi(u)=\frac{u}{2}+\frac{\sqrt{4+u^{2}}-2}{2},\quad(0<u\leq 1).$
+The proof has three parts.
+We first reduce Theorem 5.1 to an entropy lower bound for suitable distance measures, which is stated as Theorem 5.2.
+Then, we prove Theorem 5.2 by adapting the robust projection and thin-tube arguments of Shmerkin and Wang to our entropy setting, while keeping explicit logarithmic dependence on the Frostman constant.
+Finally, we return to the random-variable formulation and deduce Theorem 5.1.
 To compare Theorems 2.38 and 5.1, one may take $W=(X,Y)$ in Theorem 2.38. The random variables $X$ and $Y$ are required to be independent in Theorem 5.1, but may be dependent in Theorem 2.38.
-The rest of the proof of Theorem 1.4 for $s\leq 1/2$ will follow the same lines as in the case $s>1/2$.
+The rest of the proof of Theorems 1.2 and 1.3 for $s\leq 1/2$ will follow the same lines as in the case $s>1/2$.
 So we omit the details.
@@
 $H_{n}(\Delta(\mu_{1}\times\nu,\mu_{2}\times\nu))\geq n(\varphi(2s)-\zeta)+O_{s,\zeta,\hat{C}}(\log C)$
-for all sufficiently large $n$.
+for all $n\in\mathbb{N}$.
 We emphasize that the logarithmic dependence on $C$ in the error term is essential.
-Because in the proof of the main theorem, we are working with infinitely many random variables satisfying (70), where the Frostman constants can be large compared to $n$, and are not uniform in $(x,y)$.
-Only an error of the form $\log C$ can be turned into entropy-terms as in (71).
+Because in the proof of the main theorem, we are working with infinitely many random variables satisfying (63) or (76), where the Frostman constants can be large compared to $n$, and are not uniform in $(z_{1},z_{2})$ or $(x,y)$.
+Only an error of the form $\log C$ can be turned into entropy-terms as in (68) or (77), respectively.
 The main difficulty in this section, which prevents us from applying the results in [35] directly, is due to this reason.
-Rather than repeating their arguments wholesale with minor modifications, we provide a detailed description of how to prove variants of their results tailored to our purposes, and combine these with a few calculations to prove Theorem 5.2.
 5.1 Proof of Theorem 5.1 from Theorem 5.2
@@
 Proof of Theorem 5.1.
-For simplicity, we denote $Z:=|(X,Y)-(X^{\prime},Y^{\prime})|$, and take $C=C_{1}C_{2}$.
-The random variables $\frac{X}{2c}+\frac{1}{2}$ and $\frac{Y}{2c}+\frac{1}{2}$ are both $\big(s,O((2c)^{s}C)\big)$-Frostman, with ranges contained in $[0,1]$.
-Moreover,
-$H_{n}(|(X,Y)-(X^{\prime},Y^{\prime})|)=H_{n}\left(\left|\left(\frac{X}{2c}+\frac{1}{2},\frac{Y}{2c}+\frac{1}{2}\right)-\left(\frac{X^{\prime}}{2c}+\frac{1}{2},\frac{Y^{\prime}}{2c}+\frac{1}{2}\right)\right|\right)+O_{c}(1).$
-Without loss of generality, we assume at the beginning of the proof that both $X$ and $Y$ are $(s,C)$-Frostman, with range contained in $[0,1]$. Our aim is to prove that
-$H_{n}\big(Z\big)\geq nt+O_{s}(\log n)+O_{s,t}(\log C).$
+If $c>1$, then we consider the independent random variables $\frac{X}{2c}+\frac{1}{2}$ and $\frac{Y}{2c}+\frac{1}{2}$.
+They are both $\big(s,O((2c)^{s}C)\big)$-Frostman, and have range contained in $[0,1]$.
+Also,
+$H_{n}(|(X,Y)-(X^{\prime},Y^{\prime})|)\geq H_{n}\left(\left|\left(\frac{X}{2c}+\frac{1}{2},\,\frac{Y}{2c}+\frac{1}{2}\right)-\left(\frac{X^{\prime}}{2c}+\frac{1}{2},\,\frac{Y^{\prime}}{2c}+\frac{1}{2}\right)\right|\right)+O(1).$
+Then we can work with $\big(\frac{X}{2c}+\frac{1}{2},\,\frac{Y}{2c}+\frac{1}{2}\big)$ instead of $(X,Y)$.
+Writing $Z:=|(X,Y)-(X^{\prime},Y^{\prime})|$, we assume without loss of generality at the beginning of the proof that the ranges of $X$ and $Y$ are both contained in $[0,1]$.
+Our aim is to prove that
+$H_{n}\big(Z\big)\geq nt+O_{s}(\log n)+O_{s,t}(\log C+1).$
 Let $\gamma=\gamma(s,t)\in(0,1/3)$ be a parameter to be determined later.
 We apply Lemma 5.3 on the distribution measure of $X$, denoted by $\mu$.
-Then there is a partition $\{U_{1},U_{2},U_{3}\}$ of $[0,1]$ such that $\text{dist}(U_{1},U_{2})\gg_{s,t}C^{-1/s}$, both $\mu_{U_{1}}$ and $\mu_{U_{2}}$ are $(s,O(C))$-Frostman, and $1/2-\gamma\leq\mu(U_{1}),\mu(U_{2})\leq 1/2$.
+Then there is a partition $\{U_{1},U_{2},U_{3}\}$ of $[0,1]$ such that $\text{dist}(U_{1},U_{2})\gg(\gamma/C)^{1/s}$, both $\mu_{U_{1}}$ and $\mu_{U_{2}}$ are $(s,O(C))$-Frostman, and $1/2-\gamma\leq\mu(U_{1}),\mu(U_{2})\leq 1/2$.
 Note that $range\big((X,Y,X^{\prime},Y^{\prime})\big)\subseteq[0,1]^{4}$.
@@
 By (7), we have
-$H_{n}(Z)\geq\sum\limits_{i,j=1}^{3}P(E_{ij})H_{n}\big(Z_{E_{ij}}\big)=\sum\limits_{i,j=1}^{3}\mu(U_{i})\mu(U_{j})H_{n}\big(Z_{E_{ij}}\big).$
+$H_{n}(Z)\geq\sum\limits_{i,j=1}^{3}P(E_{ij})H_{n}\big(Z_{E_{ij}}\big)=\sum\limits_{i,j=1}^{3}\mu(U_{i})\mu(U_{j})H_{n}\big(Z_{E_{ij}}\big),$
+where $Z_{E_{ij}}$ is the random variable such that $Z_{E_{ij}}\equiv(Z|E_{ij})$.
 Since $X,Y,X^{\prime},Y^{\prime}$ are independent, the random variables $X_{U_{i}},Y,X_{U_{j}}^{\prime},Y^{\prime}$ are also independent, and $Z_{E_{ij}}\equiv|(X_{U_{i}},Y)-(X_{U_{j}}^{\prime},Y^{\prime})|$.
+Here $X_{U_{i}}$ and $X_{U_{j}}$ denote the random variables such that $X_{U_{i}}\equiv(X|X\in U_{i})$ and $X^{\prime}_{U_{j}}\equiv(X^{\prime}|X^{\prime}\in U_{j})$.
 Now apply Theorem 5.2 on $Z_{E_{12}}$ with $\zeta=\zeta(s,t)>0$ to be determined later. We deduce that
-$H_{n}(Z_{E_{21}})=H_{n}(Z_{E_{12}})\geq n(\varphi(2s)-\zeta)+O_{s,t}(\log C)$
+$H_{n}(Z_{E_{21}})=H_{n}(Z_{E_{12}})\geq n(\varphi(2s)-\zeta)+O_{s,t}(\log C+1)$
 for all $n\in\mathbb{N}$. Moreover, by Lemma 2.42, one obtains that
-$H_{n}(Z_{E_{11}}),\,H_{n}(Z_{E_{22}})\,\geq ns+O_{s}(\log n)+O_{s}(\log C)$
+$H_{n}(Z_{E_{11}}),\,H_{n}(Z_{E_{22}})\,\geq ns+O_{s}\big(\log n+\log C\big)+O(1)$
 for all $n\in\mathbb{N}$.
@@
 $H_{n}(Z)$ $\geq\sum\limits_{i,j=1}^{2}\mu(U_{i})\mu(U_{j})H_{n}\big(Z_{E_{ij}}\big)$
-$\geq 2abn(\varphi(2s)-\zeta)+(a^{2}+b^{2})ns+O_{s}(\log n)+O_{s,t}(\log C)$
-$\geq 2(1/2-\gamma)^{2}(s+\varphi(2s)-\zeta)n+O_{s}(\log n)+O_{s,t}(\log C).$
+$\geq 2abn(\varphi(2s)-\zeta)+(a^{2}+b^{2})ns+O_{s}(\log n)+O_{s,t}(\log C+1)$
+$\geq 2(1/2-\gamma)^{2}(s+\varphi(2s)-\zeta)n+O_{s}(\log n)+O_{s,t}(\log C+1).$
 To complete the proof, we consider the inequality
@@
 5.2 Proof of Theorem 5.2
-We begin by recording the results, or slight variants thereof, that we use from Shmerkin and Wang, [24], as well as the requisite material to state them.
-We then use these to explain how to prove Theorem 5.2.
-Our proof basically runs parallel to the proof of their Theorem 5.2, but with slightly different hypotheses and a different conclusion.
+We prove Theorem 5.2 by adapting parts of the framework of Shmerkin and Wang [35] to an entropy setting.
+To keep the exposition self-contained, we first recall the definitions and auxiliary results from [35] that we use, together with a few minor variants stated in the form convenient for us.
+After this, we follow the main line of their argument, but we work under slightly different hypotheses and our conclusion is an entropy lower bound rather than a robustness statement.
+To guide the reader, we indicate what is imported and what is new.
+We import from [35] the notions of robust measures, thin tubes, and adaptedness, and we use the quantitative estimates in [35] built on these concepts.
+What we modify is the bookkeeping of hypotheses and constants.
+In particular, we track the dependence on the Frostman constant $C$ throughout, in a form uniform in the conditioning parameters that appear later.
+What is new here is the final conversion step, where we turn the robustness information delivered by the thin-tube argument into a dyadic entropy lower bound at a fixed scale, with an explicit loss of size $O(\log C)$.
 5.2.1 Material from Shmerkin and Wang
@@
 Moreover if either $\mu$ or $\nu$ are trivial, then $(\mu,\nu)$ trivially have $(t,K,c,\delta)$-strong thin tubes.
-While it is clear that if $(\mu,\nu)$ have $(t,K,c,\delta)$-strong thing tubes, they have $(t,K,c,\delta)$-thin tubes, the following result from [24] shows that a partial converse also holds.
-Lemma 5.5.
-[Lemma 5.9 in [24]] Suppose $(\mu,\nu)$ have $(t,K,1-c,\delta)$-thin tubes.
+While it is clear that if $(\mu,\nu)$ have $(t,K,c,\delta)$-strong thin tubes, they have $(t,K,c,\delta)$-thin tubes, the following result from [35] shows that a partial converse also holds.
+Lemma 5.5 (Lemma 5.9 in [35]).
+Suppose $(\mu,\nu)$ have $(t,K,1-c,\delta)$-thin tubes.
 Then they have $(t,K^{\prime},c^{\prime},\delta)$-strong thin tubes, where $c^{\prime}>1-3c$, $K^{\prime}=2K$ if $c<1/4$, and $c^{\prime}>(1-\sqrt{c})/2$, $K^{\prime}=(1-\sqrt{c})^{-1}K$ if $c\geq 1/4$.
@@
 We need something slightly different, so we state and prove explicitly what we will use here.
-For $v\in\mathbb{R}^{2},$ we will define $P_{v}$ as $P_{v}(x)=x\cdot v,$ so $P_{v}\mu$ is the pullback of $\mu$ through the map $P_{v},$ which means
+For $v\in\mathbb{R}^{2},$ we will define $P_{v}$ as $P_{v}(x)=x\cdot v,$ so $P_{v}\mu$ is the pullforward of $\mu$ through the map $P_{v},$ which means
 $P_{v}(\mu)(B(x,r))=\mu\{y:y\cdot v\in B(x,r)\}.$
-Proposition 5.6.
-[Variant of Proposition 5.11 in [24]] Let $\mu_{1},\mu_{2},\nu$ be $(s,C)$-Frostman probability measures in [0,1],[0,1],then $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has $(s,2C,1,0)$-strong thin tubes.
+Proposition 5.6 (Variant of Proposition 5.11 in [35]).
+Let $\mu_{1},\mu_{2},\nu$ be $(s,C)$-Frostman probability measures in $[0,1],$ then $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has $(s,2C,1,0)$-strong thin tubes.
 Proof.
@@
 We use these new definitions to state the following, which is Lemma 5.16 in Shmerkin and Wang.
-Lemma 5.7.
-[Lemma 5.16 in [24]] Let $\rho\in\mathcal{P}(S^{1})$ be a $(\delta,s,K)$-measure and fix $\epsilon>0$.
+Lemma 5.7 (Lemma 5.16 in [35]).
+Let $\rho\in\mathcal{P}(S^{1})$ be a $(\delta,s,K)$-measure and fix $\epsilon>0$.
 If $\delta<K^{-C^{\prime}\epsilon^{-1}}$ for some absolute constant $C^{\prime}$, then $\eta(s,t)$ can be chosen so that $\rho$ is $(D_{s}-\epsilon;1)$-adapted at scale $\delta$.
@@
 The next result is Lemma 5.17 from Shmerkin and Wang, which tells us about the continuity of $\Sigma_{\tau}(D;t,d).$
-Lemma 5.8.
-[Lemma 4.7 from [24]] Fix $\tau$ and $D$ as above. If $D$ is (Lipschitz) continuous, then $s\mapsto\Sigma_{\tau}(D;s,d)$ is (locally Lipschitz) continuous on $(0,d)$.
+Lemma 5.8 (Lemma 4.7 from [35]).
+Fix $\tau$ and $D$ as above. If $D$ is (Lipschitz) continuous, then $s\mapsto\Sigma_{\tau}(D;s,d)$ is (locally Lipschitz) continuous on $(0,d)$.
 This result is almost identical to Theorem 4.8 in Shmerkin and Wang, but we modify the hypothesis in (98) below to have a constant, $C_{\mu}$ instead of 1.
@@
 $V_{F}(x):=\left(\text{ker}(DF(x))\right)^{\perp}.$
-Theorem 5.9.
-[Variant of Theorem 4.8 in [24]] Let $1\leq k\leq d,$ be natural numbers, $D:[0,d]\to[0,k]$ be a continuous function.
+Theorem 5.9 (Variant of Theorem 4.8 in [35]).
+Let $1\leq k\leq d$ be natural numbers, $D:[0,d]\to[0,k]$ be a continuous function.
 Fix $c\in(0,1)$, $C\geq 1$, $0<\delta_{1}\leq 1$.
@@
 Moreover, $\{(x,y):y\in Y_{0},x\in X_{y}\}$ is compact.
-This is a technical result taken directly from Shmerkin and Wang.
-Lemma 5.10.
-Let $\rho\in\mathcal{P}(S^{1})$ be a $(\delta,s,K)$-measure and fix $\epsilon>0$.
-If $\delta<K^{-C^{\prime}\epsilon^{-1}}$ for some absolute constant $C^{\prime}$, then $\eta(s,t)$ can be chosen so that $\rho$ is $(D_{s}-\epsilon,1)$-adapted at scale $\delta.$
 The next result is the combinatorial engine for this part of their argument, which we do not modify.
-Proposition 5.11.
-[Proposition 5.17 in [24]] Given $\zeta>0$ and $u\in(0,1],$ there are $\xi>0$ and $\tau>0,$ both depending on $u$ and $\zeta,$ such that if $s\in(0,\varphi(u)-\zeta]$ then $\Sigma_{\tau}(D_{s};u,2)\geq s+\xi.$
+Proposition 5.10 (Proposition 5.17 in [35]).
+Given $\zeta>0$ and $u\in(0,1],$ there are $\xi>0$ and $\tau>0,$ both depending on $u$ and $\zeta,$ such that if $s\in(0,\varphi(u)-\zeta]$ then $\Sigma_{\tau}(D_{s};u,2)\geq s+\xi.$
 This lemma will be applied repeatedly below. It is a variant of Lemma 5.18 from Shmerkin and Wang, but there are enough differences that we include a detailed proof.
-Lemma 5.12.
-[Variant of Lemma 5.18 in [24]] Fix $u\in(0,1]$ and $u_{0}\in(0,u/3)$, $\zeta>0,C>1$ then there is an $\eta$ depending on $\zeta$ and $u$ such that if $s\in[u_{0},\varphi(u)-\zeta]$ then we have the following: If $\mu_{1},\mu_{2}$ are $(u,C)$-Frostman measures in $0,1]^{2}$ with $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\geq C_{\mu}^{-1}$ and $(\mu_{1},\mu_{2})$ has (s,K,1c)(s,K,1-c)-strong thin tube for some number $K>0,$ then $(\mu_{1},\mu_{2})$ has $(s+\eta,C^{\prime}(KC_{\mu})^{C^{\prime}},1-9c^{1/2},0)$-strong thin tubes for some constant $C^{\prime}$ only depending on $u$ and $\zeta$.
+Lemma 5.11 (Variant of Lemma 5.18 in [35]).
+Fix $u\in(0,1]$ and $u_{0}\in(0,u/3)$, $\zeta>0,C>1$ then there is an $\eta$ depending on $\zeta$ and $u$ such that if $s\in[u_{0},\varphi(u)-\zeta]$ then we have the following: If $\mu_{1},\mu_{2}$ are $(u,C)$-Frostman measures in $0,1]^{2}$ with $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\geq C_{\mu}^{-1}$ and $(\mu_{1},\mu_{2})$ has (s,K,1c)(s,K,1-c)-strong thin tube for some number $K>0$ and $c>0$, then $(\mu_{1},\mu_{2})$ has $(s+\eta,C^{\prime}(KC_{\mu})^{C^{\prime}},1-9c^{1/2},0)$-strong thin tubes for some constant $C^{\prime}$ only depending on $u$ and $\zeta$.
 Proof.
@@
 where $\epsilon>0$ depends on $\eta$ and $\tau$ (which, in turn, depend on $u$ and $\zeta$), and $\delta_{0}=C_{1}^{-1}K^{-C_{1}},$ where $C_{1}$ is a constant also depending on $u$ and $\zeta.$
-Next, we apply Lemma 5.4 to $\rho_{y}$ to obtain that for each $y$ there is a set $X^{\prime}_{y}\subseteq S_{1}$ for which $(\rho_{y})_{X^{\prime}_{y}}(B(x,r))<r^{s+\eta}$ for every $x<\delta_{0}$ and
+Next, we apply Lemma 5.4 to $\rho_{y}$ to obtain that for each $y$ there is a set $X^{\prime}_{y}\subseteq S^{1}$ for which $(\rho_{y})_{X^{\prime}_{y}}(B(x,r))<r^{s+\eta}$ for every $x<\delta_{0}$ and
 $\rho(X^{\prime}_{y})>1-\sum_{m=-\log\delta_{0}}^{\infty}2^{-m\epsilon}\geq 1-\frac{2\delta_{0}}{1-2^{-\epsilon}}\geq 1-c^{1/2}.$
@@
 ∎
-5.2.2 Modifications to their argument
+5.2.2 Modifications to their arguments
 Theorem 5.2 is proved by following the same argument that Shmerkin and Wang prove their Theorem 5.2 in [35], except that their conclusion is about Hausdorff dimension, whereas ours is about entropy.
@@
 From there, Shmerkin and Wang prove statements about Hausdorff dimension, but here we will diverge and write out the details on how to obtain our desired conclusion about entropy.
-Proof.
+Proof of Theorem 5.2.
 We verify that the hypotheses in the statement allow us to argue as Shmerkin and Wang did in the proof of their Proposition 5.15.
@@
 $\kappa(X^{\prime})=\int_{Y}(\mu_{1}\times\nu)(d_{y}^{-1}(X^{\prime}_{y})\cap X_{y})d(\mu_{2}\times\nu)(y)\geq\int_{Y}(\mu_{1}\times\nu)_{X_{y}}\rho_{y}(X^{\prime}_{y})d(\mu_{2}\times\nu)(y)\geq(1-c^{\prime})^{3}.$
-Finally, using (82), and Lemma 2.32, we can calculate the entropy as
+Finally, since (99) hold for all $n$, and by Lemma 2.32, we can bounds the entropy as
 $H_{n}(\Delta(\mu_{1}\times\nu,\mu_{2}\times\nu))$ $=H(\kappa,f^{-1}(\mathcal{D}_{n}))$
 $\geq\kappa(X^{\prime})H(\kappa_{X^{\prime}},f^{-1}(\mathcal{D}_{n}))$
-$\geq(1-c^{\prime})^{3}(n(\varphi(2s)-\zeta/2)-\log(\delta_{0})^{-1}).$
+$\geq(1-c^{\prime})^{3}(n(\varphi(2s)-\zeta/2)-\log\delta_{0}^{-1}).$
 Since $\delta_{0}^{-1}=C_{1}\min\{K,\zeta^{-1}\}^{C_{1}}$ and $K=(C^{\prime}\min\{c_{0}^{-1},C\})^{C^{\prime O(1/\eta)}},$ we have that $\log\delta_{0}^{-1}=C_{2}\log C+C_{2}$ for some large constant $C_{2}$.
 Furthermore we can choose $M$ large enough in terms of $\zeta$ and $s$ so that $(1-c^{\prime})^{3}(\varphi(2s)-\zeta/2)>(1-2\zeta/M)^{3}(\varphi(2s)-\zeta/2)>(\varphi(2s)-\zeta),$ which concludes the proof.
-∎
 6 Proof of Theorems 1.41.6
@@
 Proof of Theorem 1.4.
-Write $\phi(x,y)=\rho_{1}(x)+\rho_{2}(y)$. Let $\epsilon=\epsilon(s)$ to be determined later.
-Note that $(x,y)\mapsto x+y$ and $(x,y)\mapsto\phi(x,y)$ are both $O_{\phi}(1)$-Lipchitz on $0,1]^{2}$.
+Write $\phi(x,y)=\rho_{1}(x)+\rho_{2}(y)$. Let $\epsilon=\epsilon(d,s)$ to be determined later.
+Note that $(x,y)\mapsto x+y$ and $(x,y)\mapsto\phi(x,y)$ are both $O_{\phi}(1)$-Lipschitz on $0,1]^{2}$.
 We can choose a small number $\eta=\eta(\phi)$ with $0<\eta<1/10$ such that $A_{\eta\delta}+_{G}A_{\eta\delta}$ and $\phi_{G}(A_{\eta\delta},A_{\eta\delta})$ are contained in $(A+_{G}A)_{\delta/10}$ and $(\phi_{G}(A,A))_{\delta/10}$, respectively, for any $\delta>0$.
@@
 $P(X\in B(x,r_{1}),Y\in B(y,r_{2}))\ll\frac{\mathcal{L}^{1}(B(x,r_{1})\cap A_{\eta\delta})\cdot\mathcal{L}^{1}(B(y,r_{2})\cap A_{\eta\delta})}{\mathcal{L}^{2}(G_{\eta\delta})}.$
-When $\delta<r_{2}\leq 1$, the concentration conditions gives
+When $\delta<r_{2}\leq 1$, the concentration condition gives
 $\mathcal{L}^{1}(B(y,r_{2})\cap A_{\eta\delta})\ll Cr_{2}^{s}(\#A)\cdot\eta\delta\ll_{\phi}Cr_{2}^{s}\delta^{1-s}.$
@@
 So $(X,Y)$ is jointly $(s,s,O_{\phi}(C^{2}\delta^{-\epsilon}))$-Frostman.
-By Theorem 1.2’ (see Section 4.3), there exists an $\epsilon^{\prime}=\epsilon^{\prime}(s)$ such that
+By Theorem 1.2’ (see Section 4.3), there exists an $\epsilon^{\prime}=\epsilon^{\prime}(d,s)$ such that
 $\begin{split}&\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\geq n(s+\epsilon^{\prime})+O_{d,s}(\log C)+O_{d,s}(\epsilon\log(1/\delta))+O_{\phi,s}(1)\end{split}$ (103)
@@
 $\begin{split}&\quad\mathcal{E}_{\delta}(A_{\eta\delta})\geq\mathcal{E}_{2^{-n}}(A_{\eta\delta})\gg\max\left\{\#\,range_{+}(D_{n}(X)),\,\#\,range_{+}(D_{n}(Y))\right\},\\ &\qquad\;\;\mathcal{E}_{\delta}(A_{\eta\delta}+_{G}A_{\eta\delta})\geq\mathcal{E}_{2^{-n}}(A_{\eta\delta}+_{G}A_{\eta\delta})\gg\#\,range_{+}(D_{n}(X+Y)),\\ &\qquad\quad\mathcal{E}_{\delta}(\phi_{G}(A_{\eta\delta},A_{\eta\delta}))\geq\mathcal{E}_{2^{-n}}(\phi_{G}(A_{\eta\delta},A_{\eta\delta}))\gg\#\,range_{+}(\phi(X,Y)).\end{split}$ (105)
-Putting (83), (84), and (86)- (88) together, one arrives at
+Putting (100), (101), and (103)–(105) together, one arrives at
 $\max\{\mathcal{E}_{\delta}(A),\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\}$
@@
 $\max\{\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\}\gg_{\phi,s}C^{-O_{d,s}(1)}\delta^{-\epsilon}(\#A)$
-whenever $\delta$ is sufficiently small depending on $\phi$ and $s$ and $C$. The proof is completed. ∎
+whenever $\delta$ is sufficiently small depending on $\phi$, $s$ and $C$. The proof is completed. ∎
 Proof of Theorem 1.5.
-Similar arguments as above will be applied. We only mention the differences that is needed to make.
+Arguments similar to those above will be applied. We only highlight the key differences.
 Since $\{y:\,(x,y)\in G\}\geq\delta^{\epsilon}(\#A)\gg\delta^{-s+\epsilon}$ for any $x\in A$, one has $G\gg\delta^{-2s+\epsilon}$.
@@
 $\begin{split}&\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\geq n(s+\epsilon^{\prime})+O_{d,s}(\log C)+O_{d,s}(\epsilon\log(1/\delta))+O_{\phi,s}(1)\end{split}$
-for some $\epsilon^{\prime}=\epsilon^{\prime}(\phi,s)>0$, which results from Theorem 1.3’ (see Section 4.4). Moreover, we substitute
+for some $\epsilon^{\prime}>0$, which results from Theorems 1.3’ (see Section 4.4). Moreover, we substitute
 $\log\#range_{+}(D_{n}(X,Y))\geq H_{n}(X,Y)$
@@
 Proof of Theorem 1.6.
-We use similar arguments as in the proof of Theorem 1.5, and work with $G=B\times B$.
+This argument is also similar to that of the proof of Theorem 1.4, but here we work with $G=B\times B$.
 Since $B\geq\delta^{\epsilon}\#A\geq\delta^{-s+\epsilon}$, one has $\mathcal{L}^{1}(B_{\eta\delta})\gg\delta^{1-s+\epsilon}$.
 Let $(X,Y)$ be a random variable that has uniform distribution on $G_{\eta\delta}=B_{\eta\delta}\times B_{\eta\delta}$.
-It is not hard to see that $X$ and $Y$ are i.i.d..
+One checks that $X$ and $Y$ are i.i.d..
 Combining the non-concentration condition, one deduces that
@@
 for any ball $B(x,r)$. In all the cases, the right-hand side can be bounded by $C\delta^{-\epsilon}r^{s}$. So $X$ is $(s,O_{\phi}(C\delta^{-\epsilon}))$-Frostman.
-As a replacement of (86), Theorem 1.4’ (see Section 4.4) shows that
-$\max\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\}\geq n(s+\epsilon^{\prime})+O_{s}(\log C)+O_{\phi,s}(1)$
-for some $\epsilon^{\prime}=\epsilon^{\prime}(s)>0$ and for all $n\in\mathbb{N}$. And the similar conclusion as in (90) appears that
+As a replacement of (103), Lemma 4.8 and Theorem 1.3” (see Section 4.4) shows that
+$\max\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\}\geq n(s+\epsilon^{\prime})+O_{d,s}(\log C)+O_{d,s}(\epsilon\log(1/\delta))+O_{\phi,s}(1)$
+for some $\epsilon^{\prime}>0$ and for all $n\in\mathbb{N}$. Finally, a similar conclusion as in (107) appears, namely that
 $\max\{\mathcal{E}_{\delta}(B+B),\,\mathcal{E}_{\delta}(\phi(B,B))\}\gg_{\phi,s}C^{-O_{d,s}(1)}\delta^{-\epsilon}(\#A)$
@@
 Remark 6.1.
-In the conclusions of Theorems 1.5-1.7, the term $C^{-O_{s}(1)}$ can be replaced by $C^{-O(1)}$ if we are only considering the situation $s\gg 1$.
+In the conclusions of Theorems 1.41.6, the term $C^{-O_{d,s}(1)}$ can be replaced by $C^{-O_{d}(1)}$ if we are only considering the situation $s\gg 1$.
+7 Erdős–Szemerédi sum-product conjecture for adaptable sets
+In this section, we translate Theorems 1.41.6 into purely discrete cardinality statements for finite sets.
+The key inputs are that uniformly $s$-adaptable sets automatically satisfy the required non-concentration hypothesis, whereas under the weaker assumption of $s$-adaptability one may pass to a large non-concentrated subset via a pruning argument.
+7.1 Definitions and results
+Definition 7.1.
+Let $0<s<1$. Let $A\subseteq[0,1]$ be finite with $|A|=N$, and let $\delta:=N^{-1/s}$. Define the probability measure
+$d\mu_{A}^{s}(x):=\frac{1}{N}\sum_{a\in A}\frac{1}{2\delta}\,\mathbf{1}_{[a-\delta,a+\delta]}(x)\,dx.$ (108)
+We say that $A$ is $s$-adaptable with constant $C_{0}$ if
+$I_{s}(\mu_{A}^{s}):=\iint_{\mathbb{R}^{2}}|x-y|^{-s}\,d\mu_{A}^{s}(x)\,d\mu_{A}^{s}(y)\leq C_{0},$ (109)
+where $C_{0}$ is independent of $N$.
+Definition 7.2.
+We say that $A$ is uniformly $s$-adaptable with constant $C_{0}$ if for every $a\in A$,
+$\sum_{\begin{subarray}{c}a^{\prime}\in A\\ a^{\prime}\neq a\end{subarray}}|a-a^{\prime}|^{-s}\leq C_{0}\,N.$ (110)
+We now list some examples of uniformly $s$-adaptable sets.
+Let
+$A_{N}=\left\{\frac{i}{N}:i=1,2,\dots,N\right\}\subseteq[0,1].$
+Then, $A_{N}$ is $\delta$-separated because $1/N\geq\delta=N^{-1/s}$ for $s<1$. For any $a=i/N$,
+$\sum_{a^{\prime}\neq a}|a-a^{\prime}|^{-s}\ll\sum_{k=1}^{N-1}\left(\frac{k}{N}\right)^{-s}=N^{s}\sum_{k=1}^{N-1}k^{-s}\leq N^{s}\left(1+\int_{1}^{N}x^{-s}\,dx\right)\ll_{s}N.$
+So $A_{N}$ is uniformly $s$-adaptable with a constant $C_{0}=C_{0}(s)$. For non-lattice-structure sets, one can take
+$A_{N}=\left\{\frac{i}{N}+\xi_{i}:i=1,\dots,N\right\}\cap[0,1],$
+with $|\xi_{i}|\leq\frac{1}{100N}$ and $\xi_{i}$ monotone. Many other examples can be found in [14].
+Theorem 7.3.
+Let $d\geq 2$ and let $\rho_{1},\rho_{2}$ be real polynomials of degree $d$.
+Let $0<s<1$ and $C_{0}\geq 1$.
+Let $\epsilon=\epsilon(d,s)>0$ and $\delta_{0}=\delta_{0}(\rho_{1},\rho_{2},s,C_{0})>0$ be the constants from Theorem 1.4.
+Let $A\subseteq[0,1]$ be any finite set satisfying $\delta:=(\#A)^{-1/s}\leq\delta_{0}$.
+Suppose that $A$ is $\delta$-separated and uniformly $s$-adaptable with constant $C_{0}$.
+If $G\subseteq A\times A$ satisfies the global density condition
+$G\geq\delta^{\epsilon}(\#A)^{2},$ (111)
+then
+$(A+_{G}A)\;+\;\#(\rho_{1}(A)+_{G}\rho_{2}(A))\ \gg_{\rho_{1},\rho_{2},s}\ C_{0}^{-O_{d,s}(1)}\,\delta^{-\epsilon}(\#A).$ (112)
+Theorem 7.4.
+Let $\phi$ be given as in (2), $0<s<1$ and $C_{0}\geq 1$. Let $\epsilon=\epsilon(d,s)>0$ or $\epsilon=\epsilon(d,\Xi,s)>0$, and $\delta_{0}=\delta_{0}(\phi,s,C_{0})>0$ be constants from Theorem 1.5.
+Let $A\subseteq[0,1]$ be any finite set satisfying $\delta:=(\#A)^{-1/s}\leq\delta_{0}$.
+Suppose that $A$ is $\delta$-separated and uniformly $s$-adaptable with constant $C_{0}$.
+If $G\subseteq A\times A$ satisfies the two-sided minimum-degree conditions
+$\{y\in A:(x,y)\in G\}\geq\delta^{\epsilon}(\#A)\,\,(\forall x\in A),\quad\#\{x\in A:(x,y)\in G\}\geq\delta^{\epsilon}(\#A)\,\,(\forall y\in A),$ (113)
+then
+$(A+_{G}A)+\#(\phi_{G}(A,A))\ \gg_{\phi,s}\ C_{0}^{-O_{d,s}(1)}\,\delta^{-\epsilon}\,(\#A).$
+Theorem 7.5.
+Let $\phi$ be given as in (2), $0<s<1$, $C_{0}\geq 1$, and $c_{\mathrm{sep}}>1$.
+Let $\epsilon=\epsilon(d,s)>0$ or $\epsilon=\epsilon(d,\Xi,s)>0$ be the constant from Theorem 1.6.
+Let $A\subseteq[0,1]$ be any finite set and let $\delta:=(\#A)^{-1/s}$.
+Suppose that $A$ is $c_{\mathrm{sep}}\delta$-separated, and also $s$-adaptable with constant $C_{0}$.
+Then, there exists a subset $A^{\sharp}\subseteq A$ with
+$A^{\sharp}\gg_{s,c_{\mathrm{sep}}}\,\#A$
+such that for every subset $B\subseteq A^{\sharp}$ with
+$B\geq\delta^{\epsilon}\,(\#A),$
+one has
+$(B+B)+\#(\phi(B,B))\ \gg_{\phi,s,\mathrm{sep}}\ C_{0}^{-O_{d,s}(1)}\,\delta^{-\epsilon}\,(\#A).$
+The role of $A^{\sharp}$ is precisely to restore a non-concentration property not guaranteed by mere $s$-adaptability.
+Bounded discrete $s$-energy permits configurations of the form $A=A_{1}\cup A_{2}$, where $A_{1}$ is a large, highly concentrated block (e.g.
+a long arithmetic progression at spacing $\delta$) and $A_{2}$ is a dispersed remainder.
+Such sets can be $s$-adaptable yet fail the non-concentration estimate on intervals containing $A_{1}$.
+Taking $B=A_{1}$ then produces a large subset for which the hypotheses needed to force sum/polynomial expansion may fail, so one cannot expect a Theorem 7.5–type conclusion to hold for all $B\subseteq A$.
+Lemma 7.6 prunes away the points with large local energy and outputs a large subset $A^{\sharp}\subseteq A$ which is $\delta^{\sharp}$-separated and satisfies the required non-concentration condition at scale $\delta^{\sharp}$.
+This is why Theorem 7.5 is formulated for $B\subseteq A^{\sharp}$: once $A^{\sharp}$ satisfies non-concentration, one can apply the subset-stable expansion result (Theorem 1.6, with $A=A^{\sharp}$ and $\delta=\delta^{\sharp}$) to obtain the desired conclusion for every large $B\subseteq A^{\sharp}$.
+7.2 Proof of Theorems 7.3 and 7.4
+Proof of Theorem 7.3.
+We verify the non-concentration hypothesis required for Theorem 1.4.
+Write N:=#AN:=\#A for simplicity.
+Let $I\subseteq\mathbb{R}$ be an interval with length $l\geq\delta$, and set $m:=\#(A\cap I)$.
+If $m\leq 1$ we are done.
+Otherwise, fix $a\in A\cap I$.
+Then for every $a^{\prime}\in A\cap I$ with $a^{\prime}\neq a$ we have $|a-a^{\prime}|\leq l$, hence
+$C_{0}N\ \geq\ \sum_{\begin{subarray}{c}a^{\prime}\in A\\ a^{\prime}\neq a\end{subarray}}|a-a^{\prime}|^{-s}\ \geq\ \sum_{\begin{subarray}{c}a^{\prime}\in A\cap I\\ a^{\prime}\neq a\end{subarray}}|a-a^{\prime}|^{-s}\ \geq\ (m-1)\,l^{-s}.$
+Therefore, $m\leq 1+C_{0}Nl^{s}$.
+Since $l\geq\delta$ implies $Nl^{s}\geq N\delta^{s}=1$, we obtain $m\leq 2C_{0}Nl^{s}$.
+So, the required non-concentration constant is $O(C_{0})$.
+Therefore, Theorem 1.4 gives
+$\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\rho_{1}(A)+_{G}\rho_{2}(A))\ \gg_{\rho_{1},\rho_{2},s}\ C_{0}^{-O_{d,s}(1)}\,\delta^{-\epsilon}\,\#A.$
+Using $|S|\geq\mathcal{E}_{\delta}(S)$ for finite sets $S$, the theorem follows. ∎
+Proof of Theorem 7.4.
+As in the proof of Theorem 7.3, the non-concentration constant is $O(C_{0})$. Applying Theorem 1.5 gives
+$\mathcal{E}_{\delta}(A+_{G}A)+\mathcal{E}_{\delta}(\phi_{G}(A,A))\ \geq\ C_{0}^{-O_{d,s}(1)}\,\delta^{-\epsilon}\,\#A.$
+Using $|S|\geq\mathcal{E}_{\delta}(S)$ for finite sets $S$, the theorem follows. ∎
+7.3 Proof of Theorem 7.5
+We first prove a pruning lemma that produces a large subset satisfying a non-concentration condition from $s$-adaptability.
+Lemma 7.6.
+Fix $0<s<1$, $C_{0}\geq 1$, and $c_{\mathrm{sep}}>1$.
+Let $A\subseteq[0,1]$ be any finite set with N:=#AN:=\#A, and let $\delta:=N^{-1/s}$.
+Suppose that $A$ is $c_{\mathrm{sep}}\,\delta$-separated, and is $s$-adaptable with constant $C_{0}$.
+Let $K\geq\max\left\{\frac{1}{1-c_{sep}^{-s}},2\right\}$, and define
+$N_{K}:=\Big\lceil\Big(1-\frac{1}{K}\Big)N\Big\rceil,\qquad\delta_{K}:=N_{K}^{-1/s}.$
+Then, there exists a subset $A_{K}\subseteq A$ with $A_{K}=N_{K}$ such that
+(i) $A_{K}$ is $\delta_{K}$-separated with $\delta_{K}\leq c_{\mathrm{sep}}\delta$,
+(ii) for every interval $I\subseteq\mathbb{R}$ of length $l\geq\delta_{K}$,
+$(A_{K}\cap I)\leq C_{\mathrm{nc}}\,l^{s}\,\#A_{K},$ (114)
+where $C_{\mathrm{nc}}\leq C(s)\,K\,C_{0}$ for some constant $C(s)>0$ depending only on $s$.
+Proof.
+For $a\in A$, define the discrete energy
+$E(a):=\sum_{\begin{subarray}{c}a^{\prime}\in A\\ a^{\prime}\neq a\end{subarray}}|a-a^{\prime}|^{-s}.$
+We first compare the continuous energy $I_{s}(\mu_{A}^{s})$ with the discrete sum.
+Since $A$ is $c_{\mathrm{sep}}\delta$-separated, for any distinct $a,a^{\prime}\in A$ and any $x\in[a-\delta/4,a+\delta/4]$, $y\in[a^{\prime}-\delta/4,a^{\prime}+\delta/4]$ we have
+$|x-y|\leq|a-a^{\prime}|+\frac{\delta}{2}\leq\frac{3}{2}\,|a-a^{\prime}|.$
+Using the definition (108), we have
+$\iint|x-y|^{-s}\,d\mu_{A}^{s}(x)\,d\mu_{A}^{s}(y)\;\geq\;\Big(\frac{1}{2\delta N}\Big)^{2}\Big(\frac{\delta}{2}\Big)^{2}\Big(\frac{3}{2}\Big)^{-s}\,|a-a^{\prime}|^{-s}=c(s)\,\frac{|a-a^{\prime}|^{-s}}{N^{2}},$
+where $c(s):=2^{-4}(3/2)^{-s}>0$. Summing over all ordered pairs $a\neq a^{\prime}$ gives
+$I_{s}(\mu_{A}^{s})\geq c(s)\,\frac{1}{N^{2}}\sum_{\begin{subarray}{c}a,a^{\prime}\in A\\ a\neq a^{\prime}\end{subarray}}|a-a^{\prime}|^{-s}=c(s)\,\frac{1}{N^{2}}\sum_{a\in A}E(a).$
+Since $I_{s}(\mu_{A}^{s})\leq C_{0}$, we conclude that
+$\frac{1}{N}\sum_{a\in A}E(a)\leq C^{\prime}(s)\,C_{0}\,N,$
+for $C^{\prime}(s):=c(s)^{-1}$.
+A direct computation shows that at most $N/K$ points satisfy $E(a)>K\,C^{\prime}(s)\,C_{0}\,N$. Hence there is a subset $A^{\prime}\subseteq A$ with $A^{\prime}\geq(1-1/K)N$ such that
+$E(a)\leq K\,C^{\prime}(s)\,C_{0}\,N\qquad\text{for all }a\in A^{\prime}.$
+Choose any subset $A_{K}\subseteq A^{\prime}$ with $A_{K}=N_{K}$.
+Then, $A_{K}$ is $\delta_{K}$-separated, since $A_{K}\subseteq A$ and $\min_{a\neq a^{\prime}}|a-a^{\prime}|\geq c_{\mathrm{sep}}\delta\geq\delta_{K}$.
+To prove (114), we use similar arguments as in the proof of Theorem 7.3 to get, for any interval $I$ with length $l\geq\delta_{K}$,
+$(A_{K}\cap I)\leq 1+K\,C^{\prime}(s)\,C_{0}\,N\,l^{s}\leq C(s)\,K\,C_{0}\,l^{s}\,N_{K}$
+for some $C(s)>0$. This completes the proof. ∎
+Proof of Theorem 7.5.
+Let $\epsilon>0$ be the exponent from Theorem 1.6. Since Theorem 1.6 remains valid if $\epsilon$ is replaced by a smaller positive number, we may assume that
+$0<\epsilon<s.$ (115)
+Write N:=#AN:=\#A. Choose
+$K_{*}\ :=\ \left\lceil\max\left\{2,\ \frac{1}{1-c_{\mathrm{sep}}^{-s}}\right\}\right\rceil\qquad\text{and set}\qquad N_{*}\ :=\ \left\lceil\left(1-\frac{1}{K_{*}}\right)N\right\rceil.$
+Apply Lemma 7.6 with $K=K_{*}$, and let
+$A^{\sharp}\ :=\ A_{K_{*}}\subseteq A,\qquad\#A^{\sharp}=N_{*},\qquad\delta^{\sharp}\ :=\ (N_{*})^{-1/s}.$
+Then, Lemma 7.6 gives:
+(i) $A^{\sharp}$ is $\delta^{\sharp}$-separated and $\delta^{\sharp}\leq c_{\mathrm{sep}}\delta$;
+(ii) for every interval $I\subseteq{\mathbb{R}}$ with length $l\geq\delta^{\sharp}$,
+$(A^{\sharp}\cap I)\ \leq\ C_{\mathrm{nc}}\,l^{s}\,\#A^{\sharp},\qquad\text{with}\qquad C_{\mathrm{nc}}\ \leq\ C(s)\,K_{*}\,C_{0}.$ (116)
+Moreover,
+$A^{\sharp}=N_{*}\ \geq\ \left(1-\frac{1}{K_{*}}\right)N.$ (117)
+Let $B\subseteq A^{\sharp}$ with $B\geq\delta^{\epsilon}\#A$. We claim that $B$ is large enough to apply Theorem 1.6 at scale $\delta^{\sharp}$. Indeed, using (115) and $A^{\sharp}\leq\#A$,
+$(\delta^{\sharp})^{\epsilon}\#A^{\sharp}=(N_{*})^{1-\epsilon/s}\leq N^{1-\epsilon/s}=\delta^{\epsilon}(\#A)\leq\#B.$
+Hence
+[…]
+We may therefore apply Theorem 1.6 to the $\delta^{\sharp}$-separated set $A^{\sharp}$, using the non-concentration estimate (116), and the subset $B$ satisfying (118). This implies
+$\mathcal{E}_{\delta^{\sharp}}(B+B)\ +\ \mathcal{E}_{\delta^{\sharp}}(\phi(B,B))\ \gg_{\phi,s}\ C_{\mathrm{nc}}^{-O_{s}(1)}\,(\delta^{\sharp})^{-\epsilon}\,\#A^{\sharp}.$
+Since $S\geq\mathcal{E}_{\delta^{\sharp}}(S)$ for any finite set $S$, we obtain
+$(B+B)\ +\ \#(\phi(B,B))\ \gg_{\phi,s}\ C_{\mathrm{nc}}^{-O_{s}(1)}\,(\delta^{\sharp})^{-\epsilon}\,\#A^{\sharp}.$ (119)
+It remains to compare $(\delta^{\sharp})^{-\epsilon}\#A^{\sharp}$ with $\delta^{-\epsilon}\#A$. By (117),
+$(\delta^{\sharp})^{-\epsilon}(\#A^{\sharp})=(N_{*})^{1+\epsilon/s}\geq\left(1-\frac{1}{K_{*}}\right)^{1+\epsilon/s}N^{1+\epsilon/s}=\left(1-\frac{1}{K_{*}}\right)^{1+\epsilon/s}\,\delta^{-\epsilon}\,(\#A).$
+Also, from (116) we have $C_{\mathrm{nc}}\leq C(s)K_{*}C_{0}$, so
+$C_{\mathrm{nc}}^{-O_{s}(1)}\ \geq\ C_{0}^{-O_{s}(1)}\,c_{0}(s,c_{\mathrm{sep}})$
+for some $c_{0}(s,c_{\mathrm{sep}})>0$ depending only on $s$ and $c_{\mathrm{sep}}$. Combining with (119) gives
+$(B+B)\ +\ \#(\phi(B,B))\ \geq_{\phi,s}\ C_{0}^{-O_{s}(1)}\,\Big(c_{0}(s,c_{\mathrm{sep}})\left(1-\tfrac{1}{K_{*}}\right)^{1+\epsilon/s}\Big)\,\delta^{-\epsilon}\,(\#A).$
+Finally, write
+$c(s,c_{\mathrm{sep}})\ :=\ \min\left\{1-\frac{1}{K_{*}},\ \ c_{0}(s,c_{\mathrm{sep}})\left(1-\frac{1}{K_{*}}\right)^{1+\epsilon/s}\right\}.$
+Then, (117) implies $A^{\sharp}\geq c(s,c_{\mathrm{sep}})\,(\#A)$, and the last display implies the desired lower bound. This completes the proof. ∎
 8 Examples
-In this section, we show details of some examples. Recall that $\mathcal{L}^{d}$ is the $d$-dimensional Lebesgue measure. Sometimes, we will write $\iota(x)=-x\log x$ for convenience.
+In this section, we will show details of some examples. A short summary is listed in advance in Table 1 below.
+Example Construction What it demonstrates
+8.1 $X,Y$ i.i.d., uniform on $A=\bigcup\nolimits_{j=0}^{2^{m}-1}[j2^{-m},j2^{-m}+2^{-m-l})$ $\bullet$ lower bounds in Lem 2.32(i) are tight $\bullet$ Thm 1.3 fails when $s=0$ or 11 $\bullet$ nsns cannot be replaced by $H_{n}(X)$ in Thms 1.2 and 1.3 $\bullet$ single term $H_{n}(X+Y)$ can be small
+8.2 $\tilde{X}=\sqrt{X}$, $\tilde{Y}=\sqrt{Y}$ from above $\bullet$ single term $)H_{n}(\phi(\tilde{X},\tilde{Y}$)) can be small
+8.3 $\tilde{X}=1+X$, $\tilde{Y}=Y$ from above $\phi(x,y)=(x+y)^{2}$ $\bullet$ Thm 1.3 fails if $\phi$ can be reduced to a univariate function
+8.4 $X$ uniform on intervals $Y$ uniform on $0,2^{-2m}] \bullet$ Thm 1.3 fails when $C$ grows with $n$
+8.5 $(X,Y)$ uniform on $\bigcup_{j=0}^{2^{m}-1}[j2^{-m},j2^{-m}+2^{-2m}]^{2}$ $\bullet$ Thm 1.3” fails without independence $\bullet$ $\frac{1}{2}(H_{n}(X)+H_{n}(Y))$ in Thm 1.2 cannot be discarded $\bullet$ $\log^{+}C$ is essential in Thm 2.38 $\bullet$ jointly $(1/4;1/4)$ but marginally $(1/2;1/2)$
+8.6 $(X,Y)$ uniform on carefully constructed $G$ $\bullet$ conditionally (1/4; 1/4) but jointly (1/2; 1/2)
+8.7 $X$, $Y$ as in 8.1 $\bullet$ $H_{n}(X+\sqrt{2}Y)-H_{n}(X)$ are much larger than $H_{n}(X+Y)-H_{n}(X)$ $\bullet$ Thm 2.23 fails when $k$ and $l$ are $\mathbb{Q}$-independent.
+Table 1: Summary of examples demonstrating sharpness and necessity of conditions in main theorems.
+Examples 7.1–7.5 show various conditions are essential; Examples 7.5 and 7.6 demonstrate the Frostman hierarchy structure; Example 7.7 shows rational arithmetic is crucial.
+Recall that $\mathcal{L}^{d}$ is the $d$-dimensional Lebesgue measure. Sometimes, we will write $\iota(x)=-x\log x$ for convenience.
 Example 8.1.
@@
 Then $\mathcal{L}^{1}(A)=2^{-l}$.
-We see that both $X$ and $Y$ have uniform distribution on $A$.
+We see that $X$ and $Y$ are independent, and both have uniform distribution on $A$.
 In the following, we will prove several results.
@@
 Moreover, when $n\leq m$, it follows from Claims 3 and 4 that
-$\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}=(1+o(1))H_{n}(X).$
-In such situation, the number $l$ can be any non-negative integer, and $X$ is $(s,O(1))$-Frostman for any $s$ with $0\leq s\leq\frac{m}{m+l}$.
+$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\right\}$
+$\qquad=\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}=(1+o(1))H_{n}(X).$
+In such situations, the number $l$ can be any non-negative integer, and $X$ is $(s,O(1))$-Frostman for any $s$ with $0\leq s\leq\frac{m}{m+l}$.
 So all the $s$ from $(0,1)$ are covered.
-Therefore, in the right-hand side of (2), the term $ns$ can not be replaced by $H_{n}(X)$.
+Therefore, in the right-hand side of (1) and (3), the term $ns$ can not be replaced by $H_{n}(X)$.
 Furthermore, if we take $n=m+l$ and $s=0$, then one deduces by Claims 2, 3, and 4 that
@@
 $P\big(X\in B(x,r)\big)\ll\begin{cases}r2^{l},\quad&\text{if }0<r\leq 2^{-l-m},\\ 2^{-m},\quad&\text{if }2^{-l-m}<r\leq 2^{-m},\\ r\quad&\text{if }2^{-m}\leq r<1,\\ 1&\text{if }r\geq 1.\end{cases}$
-Calculation shows that all the terms on the right-hand side of above formula can be bounded by $r^{\frac{m}{l+m}}$.
+Computation shows that all the terms on the right-hand side of above formula can be bounded by $r^{\frac{m}{l+m}}$.
 Indeed, the random variable $X$ is $(s,O(1))$-Frostman for any $0\leq s\leq\frac{m}{l+m}$.
@@
 $H_{n}(X^{2}+Y^{2})=H_{n}(X^{\prime 2}+Y^{\prime 2})+O(1)\geq H_{n}(X^{\prime 2})+O(1).$
-It is sufficent to show that $H_{n}(X^{\prime 2})=(1+o(1))n$.
+It is sufficient to show that $H_{n}(X^{\prime 2})=(1+o(1))n$.
 For $n\leq i\leq 2^{n}-1$, one has
@@
 Example 8.2.
-Let $X$ and $Y$ be the independent $(s,O(1))$-Frostman random variables defined in 7.1, with $n=m+l$ and $s=\frac{m}{l+m}$.
+Let $X$ and $Y$ be the independent $(s,O(1))$-Frostman random variables defined in 8.1, with $n=m+l$ and $s=\frac{m}{l+m}$.
 Consider $\tilde{X}=\sqrt{X}$ and $\tilde{Y}=\sqrt{Y}$.
@@
 Consider $\tilde{X}=1+X$, which is $(s,O(1))$-Frostman with range contained in $[1,2)$.
-Let $\tilde{Y}$ be an i.i.d.
-copy of $X$.
+And $\tilde{X}$, $Y$ are independent.
 Take $\phi(x,y)=(x+y)^{2}$, which is a degenerate quadratic form.
@@
 This gives
-$\max\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\}=(1+o(1))ns.$
-Hence, the requirement that $\phi$ is non-degenerate in Theorem 1.4 is essential.
+$\max\{H_{n}(\tilde{X}+Y),\,H_{n}(\phi(\tilde{X},Y))\}=(1+o(1))ns.$
+Therefore, in Theorem 1.3, it is essential to require that $\phi$ cannot be reduced to a univariate function.
 Example 8.4.
@@
 Claim. We have $\max\{H_{n}(X+Y),\,H_{n}(X^{2}+Y^{2})\}=n/2+O(1)$.
-In this example, we do not obtain a positive constant $\epsilon$ in Theorem 1.4 if $(X,Y)$ is regarded as a marginally $(1/2,O(1);1/2,O(2^{m}))$-Frostman random variable, i.e., $s=1/2$.
+In this example, we do not obtain a positive constant $\epsilon$ in Theorem 1.3 if $(X,Y)$ is regarded as independent marginally $(1/2,O(1);1/2,O(2^{m}))$-Frostman random variable, i.e., $s=1/2$.
 The reason is that the Frostman constant for $Y$ grows with $m$.
@@
 Therefore, in the conclusion of Theorem 1.3, it is essential to require $n>N$, where $N$ depends on the Frostman constant $C$.
-More concretely, the term $O_{s}(\log_{+}C_{1}C_{2})$ in the conclusion in Theorem 1.4’ is essential.
+In other words, the term $O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})$ in the conclusion in Theorems 1.3’ and 1.3” is essential.
 On the other hand, if we regard $(X,Y)$ as a marginally $(1/2,O(1);0,1)$-Frostman random variable, then $s=1/4$ and Theorem 1.3 works.
@@
 Example 8.5.
-This example is an variant of Example 7.4.
+This example is a variant of Example 8.4.
 Let $W=(X,Y)$, and suppose that the joint distribution of $(X,Y)$ is the uniform distribution on the set $G=\bigcup\nolimits_{j=0}^{2^{m}-1}\big[j2^{-m},\,j2^{-m}+2^{-2m}\big)^{2}$.
@@
 Claim 1.
-The random variable $(X,Y)$ is marginally marginally $(1/2,O(1);1/2,O(1))$-Frostman, and jointly $(1/4,1/4,O(1))$-Frostman, but is not conditionally $(s_{1},O(1);s_{2},O(1))$-Frostman for any $(s_{1},s_{2})\neq(0,0)$.
+The random variable $(X,Y)$ is marginally $(1/2,O(1);1/2,O(1))$-Frostman, and jointly $(1/4,1/4,O(1))$-Frostman, but is not conditionally $(s_{1},O(1);s_{2},O(1))$-Frostman for any $(s_{1},s_{2})\neq(0,0)$.
 Claim 2. We have $\max\{H_{2m}(X+Y),\,H_{2m}(X^{2}+Y^{2})\}\leq m+O(1)$, $H_{2m}(X,Y)=m+O(1)$ and $H_{2m}(X-Y)=H_{2m}(X^{2}-Y^{2})=O(1)$.
@@
 In such a situation, the two entropies can be small simultaneously.
-Therefore, the requirement of independence of $X$ and $Y$ in Theorem 1.4 is essential.
+Therefore, the requirement of independence of $X$ and $Y$ in Theorem 1.3” is essential.
 Moreover, if we apply Theorem 1.2 with the random variable (X,Y)(X,-Y), and take $\alpha=1$, $\beta=-1$, $s_{1}=s_{2}=s=1/4$, then
@@
 for some $\epsilon>0$.
-By Claim 2, the second and third entropies in above maximum-term nearly vanish.
+By Claim 2, the second and third entropies in above maximum term nearly vanish.
 So the first term $(H_{n}(X)+H_{n}(Y))/2$ is essential.
-In this example, it is $m+O(1)$, which is obviously larger than $ns=m/2$.
-Furthermore, we can not substitute $\frac{1}{2}H(X,Y)=m/2$ for $(H_{n}(X)+H_{n}(Y))/2$ as in Theorem 1.3.
+In this example, it is $m+O(1)$, which is larger than $ns=m/2$.
 In addition, for the random variable $W$, its Frostman constant $C$ can grow with $m$.
-If we consider $n=2m$ in Theorem 2.38, if $1<s<3/2$, then one has $H_{n}(|W_{1}-W_{2}|)=(1+o(1))n/2<nt$ for some suitably chosen $t$ with s/2<t<s1/2s/2<t<s-1/2.
+Consider $n=2m$ in Theorem 2.38.
+If $1<s<3/2$, then one has $H_{n}(|W_{1}-W_{2}|)=(1+o(1))n/2<nt$ for some suitably chosen $t$ with s/2<t<s1/2s/2<t<s-1/2.
 That is to say, the term $\log_{+}C$ in Theorem 2.38 is essential.
@@
 For the conditional Frostman condition, one can check that
-$P(X\in[0,2^{-2m}))=P(X\in[0,2^{-2m}),\,Y\in B(0,2^{-2m}))=2^{-m}.$
+$P\big(X\in[0,2^{-2m})\big)=P\big(X\in[0,2^{-2m}),\,Y\in B(0,2^{-2m})\big)=2^{-m}.$
 The only possible positive $s_{2}$ satisfying $2^{-m}\ll 2^{-m}(2^{-2m})^{s_{2}}$ for all sufficiently large $m$ is 0.
@@
 Let
-$G^{\prime}=\bigcup\limits_{i=0}^{2^{m}-1}\bigcup\limits_{j=0}^{2^{m}-1}[i2^{-m}+j2^{-2m},i2^{-m}+(j+1)2^{-2m})\times[j2^{-m}+i2^{-2m},i2^{-m}+(i+1)2^{-2m}).$
+$G^{\prime}=\bigcup\limits_{i=0}^{2^{m}-1}\bigcup\limits_{j=0}^{2^{m}-1}[i2^{-m}+j2^{-2m},i2^{-m}+(j+1)2^{-2m})\times[j2^{-m}+i2^{-2m},j2^{-m}+(i+1)2^{-2m}).$
 Then, for each given $x\in[0,1)$ (or $y\in[0,1)$), the set $\{y:\,(x,y)\in G^{\prime}\}$ (or $\{x:\,(x,y)\in G^{\prime}\}$, respectively) is exactly an interval of length $2^{-2m}$. Now Let
 $G=\bigcup\limits_{i^{\prime},j^{\prime}=0}^{2^{m}-1}(i2^{-m},j2^{-m})+2^{-2m}\cdot G^{\prime}.$
-Here $2^{-2m}\cdot G^{\prime}=\{(2^{-2m}x,2^{-2m}y):\,(x,y)\in G^{\prime}\}$. Suppose that $(X,Y)$ is a random variable that has uniform distribution on $G$.
+Here $2^{-2m}\cdot G^{\prime}=\{(2^{-2m}x,2^{-2m}y):\,(x,y)\in G^{\prime}\}$.
+Suppose that $(X,Y)$ is a random variable that has uniform distribution on $G$.
+It is not hard to see that $X$ and $Y$ are dependent.
 Claim 1. The random variable $(X,Y)$ is marginally $(1/2,O(1);1/2,O(1))$-Frostman.
@@
 Proof of Claim 1.
-One can check that both $X$ and $Y$ have the uniform distribution on the set $A=\bigcup\nolimits_{j=0}^{2^{m}-1}\left[j2^{-m},\,j2^{-m}+2^{-m-l}\right)$.
+One can check that both $X$ and $Y$ have the uniform distribution on the set $A_{0}=\bigcup\nolimits_{j=0}^{2^{m}-1}\left[j2^{-m},\,j2^{-m}+2^{-2m}\right)$.
 By Example 8.1, they are $(1/2,O(1))$-Frostman.
@@
 Proof of Claim 2.
-Note that $\mathcal{L}^{2}(A)=(2^{-4m})^{2}\cdot(2^{m})^{2}\cdot(2^{m})^{2}=2^{-4m}$. And $\mathcal{L}^{1}(A_{0})=2^{-m}$. Let $E$ be any Borel set, and $B(y,r)$ be any ball. Calculation reveals that
+Note that $\mathcal{L}^{2}(G)=(2^{-4m})^{2}\cdot(2^{m})^{2}\cdot(2^{m})^{2}=2^{-4m}$. And $\mathcal{L}^{1}(A_{0})=2^{-m}$. Let $E$ be any Borel set, and $B(y,r)$ be any ball. Calculation reveals that
 $P(X\in E)=\mathcal{L}^{1}(E\cap A_{0})\cdot 2^{m},$
@@
 $P(X\in E,y\in B(y,r))\ll\begin{cases}\mathcal{L}^{1}(E\cap A_{0})\cdot r\cdot 2^{4m},\quad&\text{if }0<r\leq 2^{-4m},\\ \mathcal{L}^{1}(E\cap A_{0})\cdot 2^{-4m}\cdot 2^{4m},\quad&\text{if }2^{-4m}<r\leq 2^{-m},\\ \mathcal{L}^{1}(E\cap A_{0})\cdot 2^{-4m}\frac{r}{2^{-m}}\cdot 2^{4m},\quad&\text{if }2^{-m}<r\leq 1,\\ \mathcal{L}^{1}(E\cap A_{0})\cdot 2^{-4m}2^{m}\cdot 2^{4m},\quad&\text{if }r>1.\end{cases}$
-For all the cases, it satisfies that
+For all the cases, we have that
 $P(X\in E,y\in B(x,r))\ll r^{1/4}P(X\in E).$
@@
 Proof of Claim 3.
-For any pair of balls $B(x,r_{1})$ and $B(y,r_{2})$, calculations reveal that
-$P(X\in B(x,y_{1}),y\in B(y,r_{2}))\ll M(r_{1})M(r_{2})\cdot 2^{4m}$
-where
+For any pair of balls $B(x,r_{1})$ and $B(y,r_{2})$, calculation reveal that
+$P(X\in B(x,y_{1}),y\in B(y,r_{2}))\ll$
+$\qquad\begin{cases}r_{1}\cdot r_{2}\cdot 2^{4m},\quad&\text{if }0<r_{1}\leq r_{2}\leq 2^{-4m},\\ r_{1}\cdot 2^{-4m}\cdot 2^{4m},\quad&\text{if }0<r_{1}\leq 2^{-4m}<r_{2}\leq 2^{-m},\\ r_{1}\cdot 2^{-4m}\frac{r_{2}}{2^{-m}}\cdot 2^{4m},\quad&\text{if }0<r_{1}\leq 2^{-4m},\,2^{-m}<r_{2}\leq 1,\\ r_{1}\cdot 2^{-4m}2^{m}\cdot 2^{4m},\quad&\text{if }0<r_{1}\leq 2^{-4m},\,r_{2}>1,\end{cases}$
+$\qquad\begin{cases}2^{-4m}\cdot 2^{-4m}\cdot 2^{4m},\quad&\text{if }2^{-4m}<r_{1}\leq r_{2}\leq 2^{-3m},\\ 2^{-4m}\cdot 2^{-4m}\frac{r_{2}}{2^{-3m}}\cdot 2^{4m},\quad&\text{if }2^{-4m}<r_{1}\leq 2^{-3m}<r_{2}\leq 2^{-2m},\\ 2^{-4m}\cdot 2^{-4m}2^{m}\cdot 2^{4m},\quad&\text{if }2^{-4m}<r_{1}\leq 2^{-3m},\,2^{-2m}<r_{2}\leq 2^{-m},\\ 2^{-4m}\cdot 2^{-4m}2^{m}\frac{r_{2}}{2^{-m}}\cdot 2^{4m},\quad&\text{if }2^{-4m}<r_{1}\leq 2^{-3m},\,2^{-m}<r_{2}\leq 1,\\ 2^{-4m}\cdot 2^{-4m}2^{m}2^{m}\cdot 2^{4m},\quad&\text{if }2^{-4m}<r_{1}\leq 2^{-3m},\,r_{2}>1,\end{cases}$
+$\qquad\begin{cases}2^{-4m}\frac{r_{1}}{2^{-3m}}\cdot 2^{-4m}\frac{r_{2}}{2^{-3m}}\cdot 2^{4m},\quad&\text{if }2^{-3m}<r_{1}\leq r_{2}\leq 2^{-2m},\\ 2^{-4m}\frac{r_{1}}{2^{-3m}}\cdot 2^{-4m}2^{m}\cdot 2^{4m},\quad&\text{if }2^{-3m}<r_{1}\leq 2^{-2m}<r_{2}\leq 2^{-m},\\ 2^{-4m}\frac{r_{1}}{2^{-3m}}\cdot 2^{-4m}2^{m}\frac{r_{2}}{2^{-m}}\cdot 2^{4m},\quad&\text{if }2^{-3m}<r_{1}\leq 2^{-2m},\,2^{-m}<r_{2}\leq 1,\\ 2^{-4m}\frac{r_{1}}{2^{-3m}}\cdot 2^{-4m}2^{m}2^{m}\cdot 2^{4m},\quad&\text{if }2^{-3m}<r_{1}\leq 2^{-2m},\,r_{2}>1,\end{cases}$
+$\qquad\begin{cases}2^{-4m}2^{m}\cdot 2^{-4m}2^{m}\cdot 2^{4m},\quad&\text{if }2^{-2m}<r_{1}\leq r_{2}\leq 2^{-m},\\ 2^{-4m}2^{m}\cdot 2^{-4m}2^{m}\frac{r_{2}}{2^{-m}}\cdot 2^{4m},\quad&\text{if }2^{-2m}<r_{1}\leq 2^{-m}<r_{2}\leq 1,\\ 2^{-4m}2^{m}\cdot 2^{-4m}2^{m}2^{m}\cdot 2^{4m},\quad&\text{if }2^{-2m}<r_{1}\leq 2^{-m},\,r_{2}>1,\end{cases}$
 […]
-In all cases, we have
+The rest of the cases with $r_{1}>r_{2}$ can be obtained by symmetry. In all cases, we have
 $P(X\in B(x,y_{1}),y\in B(y,r_{2}))\ll r_{1}^{1/2}r_{2}^{1/2}.$
@@
 with $X$.
-In this example, we show that the above statement may fail when $k,l$ are $\mathbb{Q}$-independent.
+In this example, we show that the above statement may fail when $k,l$ are $\mathbb{Q}$-independent real numbers.
 In particular, we consider the situation that $(k,l)=(1,\sqrt{2})$.
@@
 The proof is completed. ∎
-Appendix A Appendix - Existence of Certain Random Variable
+Appendix A Appendix - Existence of Certain Random Variables
 In this appendix, we prove two lemmas concerning the existence of certain random variables.
@@
 Construct
-$\Omega^{\prime}=\bigcup_{z}(\Omega_{z}\times\Omega_{z}).$
-Then $(\omega,\nu)\in\Omega^{\prime}$ if and only if $\omega,\nu\in\Omega_{z}$ for some $z$, if and only if $h(Z(\omega))=h(Z(\nu))$.
-The $\sigma$-algebra $\mathcal{B}^{\prime}$ can be generated by the related Cartisian products of Borel sets in $\mathcal{B}$, and the probability measure $P^{\prime}$ can be induced by the following: if $E\times F$ is a Borel subset of $\Omega_{z}\times\Omega_{z}$ for some $z$, then
+$\Omega^{\prime}=\bigcup_{z}(\Omega_{z}\times\Omega_{z}),$
+which is a disjoint union.
+Then $(\omega,\nu)\in\Omega^{\prime}$ if and only if $\omega,\nu\in\Omega_{z}$ for some $z$, which is also equivalent to $h(Z(\omega))=h(Z(\nu))$.
+The $\sigma$-algebra $\mathcal{B}^{\prime}$ can be generated by the related Cartesian products of Borel sets in $\mathcal{B}$, and the probability measure $P^{\prime}$ can be induced by the following: if $E\times F$ is a Borel subset of $\Omega_{z}\times\Omega_{z}$ for some $z$, then
 $P^{\prime}(E\times F)=\begin{cases}\frac{P(E)P(F)}{P(\Omega_{z})},\quad&\text{if }P(\Omega_{z})>0,\\ 0,&\text{if }P(\Omega_{z})=0.\end{cases}$
@@
 Thang Pham and Steven Senger would like to thank the Vietnam Institute for Advanced Study in Mathematics (VIASM) and the Erdős Center for the hospitality and for the excellent working conditions.
-We would like to thank Tuomas Orponen for his comments on the first version of this paper and for telling us about the references [1, 22].
+We are grateful to Tuomas Orponen for a series of insightful questions and valuable comments on the first version of this paper, which led us to substantially strengthen the results and greatly improve the exposition.
+In particular, he pointed out that the discretized sum-product problem for $\phi(x,y)=x^{2}+y^{2}$ and a complete graph $G$ could be approached using the techniques from [6], [30], [26], [27].